A Complete Bibliography of Publications in Network Security

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

26 December 2021 Version 1.04

Title word cross-reference

$1.6 [1468]. $107 [1192]. 128 [790]. $19m [2417]. $1bn [3899]. 2 [1981]. $24.95 [149]. $24.99 [881]. $3 [274]. $30m [2325]. $33 [1609]. $39.95 [4065]. 40 [674]. 56 [1029]. $66 [374]. $69.95 [3334]. $8 [1232].

-bit [674, 790, 1029]. -oR [323]. -peer [1981].

.HTR [1229]. .info [2025]. .NET [2247, 4319].

/bin/mail [152].

0 [149, 3305, 3344, 3345, 3794, 3814, 3901, 3956].

0-471-05318-X [149]. 0-471-16378-3 [1178].

1 [146, 184, 268, 315, 476, 604, 648, 650, 838, 1010, 1024, 1132, 1141, 1212, 1225, 1328, 1381, 1725, 1892, 1931, 2538, 3197, 3232, 3277, 3304, 3435, 3668, 3742, 3813, 3844, 3933]. 1-57870-185-6 [1482]. 1-78022-784-1 [3977]. 1-84969-412-5 [3435]. 1.9bn [3879]. 10 [527, 836]. 100 [103]. 1000 [991]. 10Gbps [2974]. 11 [549, 860]. 12a [566]. 13 [602, 879]. 14 [902]. 15 [921]. 17799 [2542]. 19 [4348, 4361, 4465]. 1997 [799, 819]. 1998 [1178].

2 [168, 208, 294, 346, 386, 520, 628, 669, 670, 859, 1025, 1042, 1149, 1241, 1242, 1346, 1765, 1942, 2554, 3210, 3295, 3326, 3327, 3679, 3700, 3784, 3845, 3955, 4029]. 2.0 [2660, 2870]. 2.2 [1287]. 2.x [759]. 2000 [1473, 1931, 1942, 2205]. 2001 [1723]. 2003 [2095]. 2006 [2605, 2652]. 2018 [4087, 4099, 4132]. 2019 [4146, 4197, 4210, 4222, 4281, 4293, 4338]. 2020 [4304, 4378]. 20m [4016]. 26.5m [2575]. 2600 [473, 530]. 29.95 [1617]. 2FA [2984, 3725, 4260]. 2nd [3295].

3 [234, 369, 689, 792, 880, 893, 3243, 3413, 3414, 3509, 3833, 3911]. 3-2-1 [4453]. 3.0 [1668]. 300PL [1272]. 31.99 [1763]. 32-bit [1483]. 3Com [1000].

4 [390, 705, 866, 1045, 3252, 3363, 3374, 3383, 3384, 3402, 3466, 3476, 3477, 3519, 3555, 3751, 3823, 3873, 3891, 3998]. 4.0 [1230, 1282, 1725, 1765, 4216, 4273]. 4.01 [1079]. 4.1x [349].

5 [727, 3045, 3253, 3266, 3403, 3435, 3699, 3710, 3863, 4009]. 5.0 [956]. 5.2.1 [13]. 50 [2230]. 5a [407]. 5b [430]. 5G [4122, 4344, 4428].

6 [228, 452, 754, 2965, 3221, 3222, 3392, 3454, 3720, 3772, 3834, 3891, 3922, 3988].

7 [471, 774, 3185, 3209, 3657, 3658, 3689, 3783, 3881, 3892, 4008]. 7xx [1181].

8 [492, 798, 3276, 3373, 3393, 3741, 3966]. 802.11 [2940]. 8e6 [2877].

9 [506, 817, 1549, 3287, 3678, 3752, 3761, 3771, 3824, 3864, 3976]. 9/11 [1803]. 9000 [630]. 95 [337, 401, 876]. 97 [683]. 978 [3197, 3209, 3210, 3221, 3222, 3232, 3243, 3252, 3253, 3266, 3276, 3277, 3287, 3295, 3304, 3305, 3326, 3327, 3344, 3345, 3363, 3373, 3374, 3383, 3384, 3392, 3393, 3402, 3403, 3413, 3414, 3454, 3466, 3476, 3477, 3509, 3519, 3555, 3657, 3658, 3668, 3678, 3679, 3689, 3699, 3700, 3710, 3720, 3741, 3742, 3751, 3752, 3761, 3771, 3772, 3783, 3784, 3794, 3813, 3814, 3823, 3824, 3833, 3834, 3844, 3845, 3863, 3864, 3873, 3881, 3891, 3892, 3901, 3911, 3922, 3933, 3955, 3956, 3966, 3976, 3988, 3998, 4008, 4009]. 978-0-12-385059-1 [3197]. 978-0-12-397181-4 [3751]. 978-0-12-411630-6 [3988]. 978-0-12-415933-4 [3402]. 978-0-12-416600-4 [3476, 3477]. 978-0-12-418676-7 [3614]. 978-0-12-418683-5 [3614]. 978-0-12-419967-5 [3699]. 978-0-12-420114-9 [3678]. 978-0-12-800006-9 [3976]. 978-0-12-800216-2 [3679]. 978-0-12-800353-4 [3998]. 978-0-12-801275-8 [3741]. 978-0-12-801867-5 [3730]. 978-0-12-802437-9 [3824]. 978-0-12-802714-1 [3844]. 978-0-12-802754-7 [3783]. 978-0-12-802864-3 [3833]. 978-0-12-802916-9 [3803]. 978-0-12-803306-7 [3881]. 978-0-12-803340-1 [3933]. 978-0-12-803451-4 [3873]. 978-0-12-803843-7 [4008]. 978-0-12-804171-0 [3814]. 978-0-12-804172-7 [3804]. 978-0-12-804180-2 [3784]. 978-0-12-804755-2 [3845]. 978-0-12-805105-4 [3823]. 978-0-12-805197-9 [3864]. 978-0-13-407695-9 [3771]. 978-0-13-417145-6 [3772]. 978-0-19-877957-5 [4009]. 978-0-691-14175-6 [3922]. 978-0-691-17654-3 [3911]. 978-0-7356-6337-4 [3425]. 978-0-7356-6339-8 [3425]. 978-1-107-02443-4 [3444]. 978-1-107-61377-5 [3444]. 978-1-118-01474-5 [3266]. 978-1-118-02647-2 [3295].

978-1-118-95850-6 [3720]. 978-1-4493-0500-0 [3344, 3345]. 978-1-4493-0538-3 [3352, 3353]. 978-1-4493-0539-0 [3352, 3353]. 978-1-4493-1943-4 [3383]. 978-1-4493-3389-8 [3393]. 978-1-4919-4946-7 [3658]. 978-1-4939-6455-0 [3901]. 978-1-5107-2332-0 [3956]. 978-1-58705-916-2 [3210]. 978-1-58714-109-6 [3221, 3222]. 978-1-58714-227-7 [3209]. 978-1-59327-266-1 [3232]. 978-1-59327-288-3 [3243]. 978-1-59327-290-6 [3392]. 978-1-59327-294-4 [3252]. 978-1-59327-385-9 [3287]. 978-1-59327-424-5 [3403]. 978-1-59327-554-9 [3731]. 978-1-59327-581-5 [3710]. 978-1-59327-590-7 [3689]. 978-1-59327-601-0 [3794]. 978-1-59327-602-7 [3892]. 978-1-59327-640-9 [3752]. 978-1-59327-649-2 [3700]. 978-1-59327-651-5 [3863]. 978-1-59327-666-9 [3761]. 978-1-59327-703-1 [3813]. 978-1-59327-757-4 [3891]. 978-1-59327-759-8 [3966]. 978-1-59327-793-2 [3955]. 978-1-59749-472-4 [3519]. 978-1-59749-572-1 [3304]. 978-1-59749-594-3 [3184]. 978-1-59749-595-0 [3184]. 978-1-59749-615-5 [3253]. 978-1-59749-627-8 [3276]. 978-1-59749-635-3 [3413]. 978-1-59749-655-1 [3277]. 978-1-59749-661-2 [3326, 3327]. 978-1-59749-665-0 [3305]. 978-1-59749-727-5 [3334]. 978-1-59749-948-4 [3374]. 978-1-59749-951-4 [3384, 3466].

978-1-59749-957-6 [3454]. 978-1-59749-961-3 [3414]. 978-1-59749-993-4 [3555]. 978-1-68050-288-6 [4065]. 978-1-78398-214-1 [3668]. 978-1-78548-004-1 [3742]. 978-1-78548-012-6 [3834]. 978-1-84928-390-8 [3373]. 978-1-84928-398-4 [3363]. 978-1-84951-742-3 [3509]. 978-1-84969-510-7 [3657].

= [2734].

A-V [2412]. Aaron [3771]. abandons [1774]. Abhinav [3509]. Abilities [1593]. About.com [1515]. Abraham [3823]. Abstracts [1792]. abuse [86, 792, 1177, 2781, 4101, 4416, 4461, 4473]. Abusing [3395, 4126]. Academia [1947]. Academic [3751, 3988]. accelerate [3746]. Accelerated [2321]. accelerator [1049]. accelerators [875, 931]. Accellion [4466]. Accenture [4185]. Accenture/Ponemon [4185]. Access [21, 46, 47, 162, 210, 431, 446, 524, 618, 630, 664, 700, 747, 812, 873, 890, 895, 913, 1030, 1048, 1215, 1253, 1472, 1488, 1597, 2083, 2253, 2337, 2355, 2747, 3202, 3267, 3337, 3470, 3967, 3969]. accessing [3638]. accidentally [3064]. account [2755]. Accountability [4446]. accountable [174]. Accounting [3963]. accounts [300, 3447, 4107]. accused [2052]. accuses [3871]. ACD [4245]. ACE [178, 561, 700]. ACE/Client [178]. ACE/Server [561, 700]. a’changing [2558]. Achieving [4248]. Achilles [2397]. acquire [1953]. acquires [662, 1610]. acquisition [118, 965, 1061]. Acrobat [1161]. across [424, 2654, 2986, 3713]. Act [473, 530, 604, 628, 706, 828, 1529, 2009, 2280, 2382, 2652, 4372]. Action [910, 933, 2085, 2554, 3963, 4159]. Active [1366, 1463, 2144, 3318, 4179, 4245]. Actively [3829]. ActiveX [699, 896, 903, 1266, 1612]. activists [1936]. activity [2673, 3134]. actors [4371, 4416]. actually [4394]. Adam [3771]. adapter [179]. Adapters [1339]. Adapting [3982]. adaptive [1160, 2073, 3267]. ADCs [3489]. add [2124, 2411]. add-on [2124]. add-ons [2411]. Added [1018, 1410, 1891]. Addison [1763]. Address [506, 846, 1138, 1419]. addressed [221]. addresses [225]. Addressing [2430, 2784, 2797, 3166, 3753, 3893]. adds [2330]. Adhering [4380]. ADKs [1491]. admin [3786]. administration [1251]. Administrator [77, 2218, 3185]. administrators [1618, 2023, 2328, 3184]. admins [3035]. AdminSuite [779]. Admission [1728]. admits [1866]. Adobe [1161, 3003]. adopt [538, 1565]. adopts [3611]. ads [26, 2303, 3538]. ADSL [1253]. Advanced [912, 1001, 1210, 1293, 1508, 1523, 1788, 1956, 2795, 3018, 3070, 3127, 3156, 3236, 3567, 3692, 3989, 4329]. advances [548]. advantage [31, 1526]. Adventures [1971]. adversaries [3989]. Advertising [3000, 3762]. advice [4137]. advocate [233]. advocates [595]. Adware [1984, 2155, 2791]. AdWords [2810]. AES [1501, 3018]. affair [3744]. affect [3579]. affects [4006]. affordable [1106]. Africa [245]. After [1670, 1967, 2141, 2386, 2957, 3460, 3582, 4061, 4159]. aftermath [4061]. Again [20, 90, 994, 1581, 1821, 2070, 2704, 2800, 3441, 3654]. Against [114, 218, 503, 655, 901, 914, 933, 975, 988, 1157, 1218, 1321, 1453, 1515, 1625, 1641, 1757, 1768, 2233, 2275, 2715, 2982, 2995, 3077, 3451, 3570, 3607, 3611, 3616, 3675, 3841, 3989, 4013, 4030, 4068, 4239, 4367, 4391, 4406, 4421, 4426, 4442]. age [268, 294, 1277, 2327, 3929, 3988, 4200, 4462]. agencies [377, 2268, 3723, 3860, 4444]. agency [278, 1208, 3841]. agenda [1022]. agent [925, 4162]. agile [3721, 3774, 4056]. agility [4330, 4424]. Agreement [247, 459, 575, 724]. AgTech [4249]. aha [1279]. Aharoni [3243]. Ahead [1599, 1676, 2078, 2380, 3587, 3671]. AI [1738, 1800, 4127, 4250, 4329]. aid [468]. aim [623, 2316]. aims [3323]. aircraft [4267]. airforce [1039]. airport [1881]. airwaves [2654]. Airways [3687]. AIX [216, 943, 1247, 1446]. AJAX [2660]. AJAX/web [2660]. Ajin [3823]. Al [3228]. Al-Qaeda [3228]. Alarm [1256, 3811]. alarms [1848, 1886]. Alcatel [2868]. alert [119]. Alerts [1698, 2084, 2336]. Alexander [1972]. Alexandre [3221, 3222]. algorithm [396, 674, 762]. Algorithms [1088, 1507]. alive [552]. Allan [3881]. alleged [3003]. allegedly [3380]. alleviate [1996]. Alliance [250, 360, 485, 599, 752, 781, 783, 809, 830, 968, 1139, 1318, 1408, 1450, 1635, 2335]. alliances [623]. allies [725]. allow [535, 964, 985, 1215, 1423]. allowed [807, 1489]. Allows [583, 1633]. ally [251]. almost [1731]. Along [1530]. Alphabetic [1103]. already [4397]. also [3417, 3676]. Alternate [2588]. Alternative [123, 921, 1226, 2602, 2698, 2699, 2795]. alternatives [146, 168, 2508, 2873, 3116]. Always [1596]. amasses [3601]. Amazon [1692, 3585]. ambushed [1260]. AMD [4323]. amendment [1243]. amendments [1262]. America [122, 177, 326, 707, 824, 1038, 1200, 3956]. American [601, 795, 2333]. amidst [343]. Amit [3752]. among [3119, 3687, 3986, 4254, 4270]. Amongst [1723]. Amsterdam [1881]. AMTSO [3288]. analysing [3224]. Analysis [706, 727, 799, 1579, 1910, 1979, 2126, 2142, 2168, 2205, 2218, 2555, 2598, 2750, 2798, 2832, 2973, 3055, 3127, 3226, 3232, 3334, 3378, 3392, 3576, 3636, 3648, 3823, 3944, 4075]. analysis-led [3576]. Analyst [642, 2092]. analytics [3522, 4174, 4200]. analytics-driven [4174]. Anatomy [1133, 2892, 2929, 3017, 3692, 3885]. Andre [3210]. Andrea [3373]. Andrew [3392, 3853]. Android [3216, 3241, 3254, 3324, 3341, 3364, 3375, 3381, 3387, 3464, 3479, 3552, 3638, 3710, 3727, 3772, 3823]. Animated [2667]. Anna [1758]. announce [75]. announced [811]. announces [118, 752, 965, 2377, 4017]. annoyance [2495, 3154]. Annual [489, 4132, 4281]. Anomaly [3834]. Anonymity [3456, 3629, 3807]. Anonymous [3154]. ANR [225]. ANSI [126]. answer [546, 1781, 2140, 4091, 4296]. answers [366, 625]. Anti [285, 366, 526, 650, 670, 683, 721, 742, 803, 969, 1103, 1201, 1434, 1449, 1454, 1538, 1616, 1687, 1887, 2056, 2374, 2477, 2490, 2870, 2886, 2910, 2933, 3131, 3293, 3390]. anti-bug [366]. anti-forensics [2886]. anti-hacker [1687]. Anti-hackers [1538]. anti-malware [2933]. Anti-phishing [2477, 2910]. anti-piracy [3293]. Anti-social [2870]. Anti-spam [803]. anti-spyware [2374]. Anti-virus [285, 526, 650, 670, 683, 721, 742, 969, 1103, 1201, 1434, 1449, 1454, 1616, 1887, 2056, 2374, 2490, 3131, 3390]. antivirus [2277]. Anton [3374, 3403, 3413, 3761]. Antonio [3834]. any [2364, 4115]. anything [1731, 2085]. anyway [4112]. AOL [786, 1365, 1636]. AOL4FREE [771]. AOLGOLD [405]. Apache [1952, 2029, 4029]. Apex [3526]. API [4080, 4112, 4226]. APIs [3478]. app [3329, 4011, 4359]. Appeal [1675, 2361]. appearance [3130]. appeased [595]. appetite [1527]. Apple [2278, 2516, 2573, 3026, 3085, 3422, 3847]. Applet [877]. applets [731, 849, 914, 1487]. appliance [1389, 2333, 2413, 3355, 3405]. Applicable [3444]. Application [1390, 1595, 1986, 2036, 2202, 2536, 2599, 2670, 2729, 2737, 3020, 3269, 3295, 3471, 3489, 3578, 3720, 3794, 3957, 4056, 4318]. application-based [2202]. Applications [359, 682, 848, 897, 947, 1004, 1018, 1472, 1890, 2110, 2251, 2262, 2270, 2452, 2570, 2660, 2663, 2821, 2963, 3005, 3039, 3126, 3553, 3639, 3823, 4019, 4046]. Applied [3498]. Appointed [2138]. Apprenticeships [4296]. approach [2405, 2642, 2678, 2772, 2773, 3094, 3197, 3267, 3363, 3406, 3429, 3457, 3597, 3630, 3673, 3899, 4047, 4152, 4161, 4176]. Approaches [1930, 2195, 2706, 2808, 2852, 3007]. approval [1611]. approvals [909]. Apps [2184, 2502, 3045, 3238, 3384, 3466, 3769, 3817]. APT [3240, 4400]. APTs [3268, 3420, 3592, 3641]. Aquilina [3519]. Aracil [3834]. arbitrary [885, 964]. Architecture [1678, 2015, 2021, 2770, 2987, 3363, 3375]. archive [1464]. area [147]. arena [1008]. arguments [4418]. arm [2167, 2853]. armies [2072, 2524]. arms [310]. army [295, 1274, 1854, 1917, 2640]. arrest [89, 185, 1626, 2308]. arrested [154]. arrests [1078, 2458, 3302, 3931]. arrival [2327]. arrives [1388, 2136]. art [1378, 2540, 4057]. Article [1581]. Artificial [3937, 4070, 4239]. Arvind [3853]. Ascend [992]. Ashley [3744]. ask [2133]. asked [826]. asks [794, 1716, 2824]. ASN.1 [2212]. ASP [1594]. aspects [599, 2409]. assault [2436]. assesses [642]. Assessing [166, 1146, 2218, 2339, 3234, 3990, 4061]. Assessment [254, 425, 798, 868, 1212, 1241, 1547, 1911, 1917, 1987, 2059, 2132, 2202, 2366, 2405, 2431, 4120]. asset [2769, 3895]. assets [564, 3866, 4366]. assigner [325]. Associates [970, 1144, 1337]. assuming [3948]. assurance [1261, 3321]. assurances [4207]. AssureNet [662]. Assuring [3604]. Asus [4196]. AT&T [137, 182, 305, 335, 716, 824, 3085, 4352]. AT&T/Mondex [716]. Atlantic [50]. ATM [207, 302, 403, 432, 1145, 1234, 3082]. atsar [1377]. attached [2248]. Attachment [2713]. Attachment-based [2713]. Attack [407, 430, 437, 452, 508, 614, 654, 905, 1133, 1224, 1334, 1352, 1380, 1441, 1443, 1654, 1720, 1741, 2014, 2079, 2184, 2245, 2481, 2510, 2578, 2648, 2676, 2720, 2829, 2832, 2890, 2892, 2926, 2929, 2934, 2940, 2982, 3035, 3053, 3063, 3146, 3260, 3264, 3284, 3432, 3445, 3448, 3461, 3469, 3563, 3588, 3591, 3654, 3705, 3783, 3820, 3846, 3868, 3871, 3885, 3899, 3926, 3952, 4074, 4155, 4204, 4218, 4278, 4371, 4397, 4405]. attacked [656, 2025, 2077, 3238, 3241, 3400, 3473, 3964]. Attacker [885]. Attackers [959, 1267, 2790, 2800, 3482, 3643, 3994, 4096, 4334, 4358, 4421]. Attacking [344, 1131, 3375, 4348]. Attacks [128, 346, 565, 589, 655, 711, 833, 933, 1353, 1373, 1399, 1401, 1402, 1444, 1633, 1731, 1753, 1834, 1853, 1895, 1922, 1927, 2107, 2142, 2207, 2271, 2274, 2527, 2582, 2634, 2685, 2715, 2716, 2796, 2957, 2968, 2995, 3051, 3117, 3155, 3218, 3342, 3361, 3459, 3516, 3542, 3560, 3584, 3611, 3612, 3634, 3671, 3686, 3697, 3738, 3799, 3801, 3805, 3865, 3883, 3888, 3931, 3935, 3949, 3953, 3989, 3996, 3999, 4000, 4005, 4013, 4025, 4032, 4035, 4049, 4052, 4058, 4073, 4093, 4138, 4143, 4232, 4275, 4309, 4331, 4359, 4406, 4410, 4440, 4471]. attacks [625, 686, 2383, 2536, 2729, 2737, 3257, 3395, 4175, 4178, 4214, 4415, 4442, 4452]. Attempt [1453, 4456]. attempts [1376, 4214]. attitude [2022, 2589, 3124, 3579]. attracts [1798]. attribution [3949]. Attrition [1685, 1710]. auctioned [2854]. Audit [1023, 1725, 1765, 2347, 2661, 3167, 3587]. audited [2342]. Auditing [649, 1225, 1242, 1279, 1369, 1457, 1725, 1765, 3039]. auditors [3427]. audits [2917, 3219]. August [2165]. Aurora [3022]. Aussies [1358]. Australia [1303, 3336]. Australian [68, 810, 878, 4371]. Authenticated [1660]. Authenticating [818, 2395]. Authentication [202, 385, 392, 463, 525, 536, 577, 722, 732, 768, 813, 832, 989, 1312, 1406, 1410, 1431, 1502, 1525, 2113, 2498, 2543, 2654, 2699, 2701, 2731, 3019, 3188, 3330, 3609, 3714, 3777, 4152, 4447]. authenticity [1125]. Author [1171, 1178, 1758, 2309]. Authorities [3801, 3841, 4167]. authority [1436]. authorization [520]. authorizer [306, 597]. authors [2162, 2223]. Auto [367].

Automated [452, 920, 3823, 4259].

Automatic [974, 1306, 2285]. Automating [2882, 3803, 3959]. automation [2719, 3934, 4174, 4318]. Automotive [3745, 4304]. automountd [1215]. AV [2487]. availability [1128, 2262, 3378, 4067]. available [30, 665, 714, 875, 888, 2396]. Avast [4157]. avenues [2184, 2619]. Avoid [1421, 2446, 2590, 3125, 3370, 4175]. Avoiding [3133, 3447, 3786, 3826, 4379, 4463]. avoids [2980]. awake [1886]. Award [140, 1542]. aware [3182]. Awareness [1209, 2306, 2520, 3155, 3458, 3699, 4020, 4071]. away [95]. AWS [4047]. AX [2144]. Axent [118, 662, 899]. axes [2876]. Azlan [360].

b [586, 1869, 2390]. BA [4109]. Baby [73]. Bace [1482]. Back [1403, 1459, 1757, 1772, 1832, 1919, 2031, 2063, 2160, 2171, 2278, 2428, 2629, 2713, 2940, 3025, 3056, 3389, 3694, 3734, 3780, 3805, 4453]. back-up [3694]. back-ups [4453]. Backdoor [1877]. backdoors [3528]. backed [256, 3111, 3985]. Background [2119, 2130]. backing [1651]. BackOffice [1153]. backs [1584, 1819]. Backup [258, 501, 1432, 1877]. Bad [494, 585, 1790, 1982, 2044, 2164, 2864, 3289, 3786, 4005, 4201]. badder [2149]. Badman [1442]. badware [2944]. BAE [2292]. bail [1732]. Bair [3933]. Baker [453]. balance [1610, 3147, 4123, 4173, 4330]. balanced [4067]. Balancing [1114, 1622, 3663, 4382]. Balkanizes [1188]. ball [1543]. Baltimore [1105, 1161, 1384, 1851]. ban [1326]. Bank [122, 300, 399, 808, 1078, 1572, 1796, 1809, 2392, 2438, 2471, 2564, 3974]. Banking [159, 264, 417, 445, 573, 891, 947, 1085, 1104, 1494, 2755, 3171]. Banks [855, 881, 2712, 3400]. banned [2979, 3857]. bans [3433]. bar [626, 1254, 2390]. Barbarians [2430]. barbecues [844]. Barclays [541, 1809]. bare [4177]. Baron [239]. Barr [1243]. barrage [4194]. barred [2606]. barrier [1301]. Based [769, 833, 897, 898, 947, 1018, 1068, 1073, 1127, 1201, 1390, 1418, 1461, 1547, 1678, 1891, 2117, 2175, 2202, 2713, 2871, 2954, 2995, 3226, 3246, 3486, 3502, 3560, 3866, 3901, 3957, 3981, 4012]. baseline [2689]. Bash [3657]. Basic [2328]. Basics [1212, 1494, 2171, 3277, 3326, 3327, 3476, 3477, 3615, 3734]. bates [1845]. Battle [299, 406, 498, 837, 3033, 3041, 3066, 3837, 3841, 3847]. battles [3758, 3775, 3781, 4184]. bay [284, 2611]. BayRS [463]. Baysecure [561]. BBC [2253, 2909]. Be [399, 444, 477, 648, 1392, 1437, 1538, 1559, 1702, 1726, 1811, 1838, 1858, 1935, 2199, 2342, 2348, 2357, 2384, 2549, 3026, 3111, 3204, 3279, 3337, 3417, 3505, 3682, 3815, 3855, 3857, 4010, 4122, 4254, 4331, 4364, 4450]. beaconing [4329]. beams [123, 1921]. Beat [537, 3427]. beating [4247]. beats [2256]. BEC [4105]. become [1272, 1937, 2274, 3810, 3919]. becomes [62, 398, 504, 593, 3899, 3909]. becoming [3633, 4114]. bed [72]. beef [1341]. Been [1530, 1982]. before [382, 868, 1413, 3712]. beginners [2761]. Begins [1791]. behaving [4288]. Behaviour [1283, 1894, 3226, 3640]. Behavioural [2973]. Behind [1459, 2365, 2543, 2627, 2862, 3641, 3933]. Being [1447, 1677, 2313, 2680, 3451, 3738, 4452, 4465]. beleagured [2281]. Bell [50, 200, 951]. bells [73, 2421]. Benchmark [1956]. benchmarks [1962, 2035]. benefit [1185]. benefits [2326, 2475, 2985, 3269, 3609]. BERR [2802]. Best [1162, 1451, 1559, 1671, 1767, 2115, 2117, 2289, 2416, 2494, 2573, 2600, 2863, 3279, 3306, 3501, 4137, 4251, 4261]. beta [3179]. better [78, 130, 660, 1145, 1641, 2520, 2537, 2539, 2685, 2813, 3430, 4250, 4398]. betting [2256]. between [575, 903, 1925, 2199, 2409, 4031, 4173, 4230, 4330, 4352]. Beware [2918]. Beyond [1054, 2028, 2454, 2496, 2622, 3224, 3267, 3774, 4308]. Bible [3304]. Bid [1844]. Bidding [2448]. Big [1187, 1530, 1666, 1982, 2144, 2158, 2169, 2179, 2188, 2192, 2196, 2210, 2220, 2238, 2265, 2290, 2304, 2317, 2351, 2367, 2643, 2814, 3282, 3342, 3344–3346, 3522, 3537, 3589, 3661, 3771, 3772, 3913, 3923, 4009]. bigger [2149, 2751]. biggest [1704, 2185, 2228, 3601, 3821, 3883, 4174, 4328]. biggest-ever [3601]. Bill [797, 1188, 1193, 1260, 1262, 1505, 1528, 1832, 1966, 3699, 3772, 3791]. Billion [274, 887, 1232, 1235]. bin [152, 510]. bind [11, 435, 983, 1549, 1627, 1976]. Binding [1437]. Bio [1480]. Bio-logically [1480]. Biometric [1431, 1573, 1881]. Biometrics [1825, 1849, 1928, 2407, 2483, 2633, 2641]. birth [212]. Bit [674, 790, 1029, 1483, 4055]. Bit9 [3410]. Bitcoin [3853, 4025]. Bitdefender [4271, 4293]. bite [2383]. bites [2160]. BitLocker [2777]. Bits [387, 695, 837, 1235, 2456]. bitten [3026, 3410]. Biz [1662]. Black [239, 427, 2016, 2152, 2297, 2845, 3101, 3230, 3350, 3448, 3689]. Blackberry [2433]. blackhat [2727, 2738]. Blackhats [2106, 2316]. blackmail [840, 2256]. Blame [2160, 2178]. blamed [1098, 2389, 4025, 4130]. blames [3441, 4143]. blanket [1608]. Blaster [2148, 2161]. blasts [1601, 2447]. Blg [2196]. blindness [3104]. Bliss [1769, 2381, 4439]. Blitz [1662, 2953]. Blizzard [1007]. block [2292, 2517, 3267, 4014]. blockchain [4030, 4430]. Blockers [1894]. Bloomberg [4117]. blows [2930, 3263]. Blue [1094, 1757]. Bluejacking [2204]. blues [258, 1094]. Bluetooth [1913, 2098, 2231, 2383, 2531, 2817, 3223]. blunder [3399]. Blurring [4031]. BNP [2879]. Board [60, 330, 904, 1117, 1316, 1614, 2708]. Boardroom [4446]. boardrooms [3506]. boards [494]. boat [953]. Body [2180]. Bolero.net [1497]. bolstered [1719]. bomb [858]. bombard [1554]. bomber [2292]. bombs [126]. Bonn [2959]. Bonneau [3853]. Book [3184, 3197, 3209, 3210, 3221, 3222, 3232, 3243, 3252, 3253, 3266, 3276, 3277, 3287, 3295, 3304, 3305, 3326, 3327, 3344, 3345, 3373, 3374, 3383, 3384, 3413, 3414, 3425, 3435, 3444, 3454, 3466, 3476, 3477, 3519, 3555, 3614, 3657, 3658, 3668, 3678, 3679, 3689, 3699, 3700, 3710, 3720, 3741, 3742, 3751, 3752, 3761, 3771, 3772, 3783, 3784, 3794, 3813, 3814, 3823, 3824, 3833, 3834, 3844, 3845, 3853, 3863, 3864, 3873, 3881, 3891, 3892, 3901, 3911, 3922, 3955, 3956, 3966, 3976, 3977, 3988, 3998, 4008, 4009, 4065, 4089]. Book [149, 479, 799, 819, 881, 1178, 1482, 1617, 1673, 1763, 1767, 3334, 3352, 3353, 3363, 3392, 3393, 3402, 3403, 3509, 3700, 3730, 3731, 3803, 3804, 3933, 3944, 3945]. Books [881]. Bookshelf [4065]. Boom [131, 176, 1849]. booming [2140]. Boost [192, 1559, 2615, 3282]. boosting [3829]. boot [426, 2591]. Bootable [1748]. Border [3905]. Borders [1639, 2259, 2664]. BorderWare [415]. borne [610, 2993]. boss [2351, 2367]. Bot [2422, 2801, 3056, 4427]. both [2805, 3020]. Botherders [3041]. botmaster [2562]. Botnet [2524, 2640, 2672, 2909, 2946, 3112, 3129, 3194, 3341, 3411, 3888, 4167, 4190, 4325]. Botnets [2545, 2571, 2627, 2889, 3033, 3066, 3073, 3201, 3227, 3356, 3806, 4162]. Bots [1581, 2594, 2935]. Bottom [2195]. Bottom-up [2195]. bounce [907]. boundaries [2428, 2496, 3163, 4031]. bounty [2255]. Bouzefrane [3742]. Bovine [900]. Box [427, 1500, 2016, 2401, 4010]. boycott [1536]. BR [155]. brain [3030]. branch [1536]. Branden [3374]. Branding [1544]. Brandon [3892, 3966]. brawn [3030]. Brazil [244]. Brazilians [1252]. Breach [2065, 2673, 2743, 3219, 3373, 3499, 3643, 3659, 3691, 3808, 3873, 3874, 3974, 4029, 4096, 4182, 4378, 4444]. breached [300, 3842, 3855, 4109, 4435]. Breaches [388, 974, 1420, 1918, 2228, 2553, 2779, 2825, 2900, 3573, 3687, 3738, 3768, 3848, 4043, 4063, 4135, 4181, 4197, 4242, 4254, 4466]. break [160, 440, 498, 991, 1037, 3861]. break-in [440]. Breaking [232, 748, 1621, 2018, 2428, 2817, 3786]. breaks [615, 4230]. Breakthrough [765, 951, 1855]. Bredolab [3129]. breeding [3533]. Brett [3933]. Brexit [3913]. Brian [3772, 3911]. bridged [872]. Bridges [1682]. Bridging [3785]. bridles [2456]. Brief [1541, 1588, 1605, 1628, 1646, 1658, 1669, 1690, 1713, 1735, 1756, 1776, 1799, 1823, 1847, 1868, 1968, 2026, 2038, 2055, 2064, 2080, 2094, 2109, 2123, 2137, 2150, 2163, 2173, 2186, 2200, 2215, 2227, 2244, 2258, 2269, 2282, 2295, 2310, 2323, 2334, 2349, 2362, 2378, 2391, 2403, 2415, 2427, 2439, 2449, 2461, 2470, 2482, 2493, 2503, 2515, 2525, 2535, 2552, 2565, 2576, 2586, 2596, 2607, 2618, 2628, 2638, 2650, 2659, 2669, 2677, 2686, 2696, 2705, 2714, 2726, 2735, 2745, 2757, 2768, 2780, 2792, 2803, 2816, 2827, 2837, 3186, 3196, 3198, 3220, 3242, 3565]. brief [2847, 2859, 2869, 2880, 2891, 2901, 2911, 2923, 2937, 2947, 2960, 2971, 2983, 2994, 3006, 3016, 3024, 3034, 3044, 3054, 3065, 3079, 3090, 3102, 3113, 3123, 3132, 3143, 3153, 3162, 3173, 3183, 3208, 3231, 3251, 3265, 3275, 3286, 3294, 3303, 3314, 3325, 3333, 3343, 3351, 3362, 3372, 3382, 3391, 3401, 3412, 3424, 3434, 3443, 3453, 3465, 3475, 3485, 3496, 3507, 3518, 3529, 3543, 3554, 3574, 3583, 3593, 3603, 3613, 3625, 3635, 3645, 3656, 3667, 3677, 3688, 3698, 3709, 3719, 3729, 3740, 3750, 3760, 3770, 3782, 3793, 3802, 3812, 3822, 3832, 3843, 3852]. brief [368, 1196, 3862, 3872, 3880, 3890, 3900, 3910, 3921, 3932, 3943, 3954, 3965, 3975, 3987, 3997, 4007, 4018, 4026, 4036, 4045, 4053, 4064, 4076, 4088, 4100, 4111, 4121, 4133, 4147, 4160, 4172, 4186, 4198, 4211, 4223, 4234, 4246, 4258, 4272, 4282, 4294, 4305, 4317, 4327, 4339, 4353, 4363, 4376, 4388, 4402, 4414, 4425, 4437, 4448, 4460, 4470]. Briefing [423, 894]. brigade [3556]. brightening [2996]. Bring [79, 2315, 3558, 3663]. Bringing [302, 3436, 3471, 3521]. brings [968, 2335, 2603]. British [482, 2370, 3687]. Broad [250]. broadband [1343].

Broadcom [4323]. broaden [2361]. broader [3281]. Brokat [1104]. brokering [1386]. brother [1187, 1530, 2814]. Browser [866, 1731, 2948, 2953, 2954, 3058, 3200]. browser-based [2954]. Browsers [1034, 2611, 2698, 2995, 3053, 3071]. Browsezilla [2611]. Bruce [149, 2083, 3197, 3955]. Bruised [2278]. Brussel [979]. BT [85, 180, 191, 193, 1449, 1629, 2615]. Buchanan [1608]. buck [95, 1915]. Budapest [1816]. budget [2252, 3011]. budgeting [2540]. Budgets [1138]. budgie [2408]. Buffer [631, 633, 673, 759, 841, 908, 964, 1013, 1081, 1097, 1151, 1154, 1180, 1378]. Bug [366, 692, 805, 1366, 1397, 1775, 1995, 2270, 2371, 2448, 3032, 3287, 3572, 3623, 4218, 4315]. Bug-fixed [2270]. bugged [1435]. BugNet [1366]. Bugs [804, 854, 1423, 2267]. Bugtraq [2077]. Build [1827, 2954, 4080]. Building [529, 1531, 1682, 2394, 2644, 3211, 3290, 3291, 3317, 3598, 3680, 3699, 3797, 4056]. buildings [194, 1760, 3877]. builds [4056]. built [3732]. Bull [247]. bullet [3358]. bulletin [60, 330, 494]. bulletproof [3586]. Bullish [2370]. Bulter [242]. Bunni [1845]. Bunny [242]. buoyant [2320]. Bureau [2685]. burn [4020]. bury [3981]. bus [3290]. Bush [1774, 1783, 2389]. Business [134, 219, 460, 485, 623, 638, 645, 848, 971, 1016, 1017, 1302, 1328, 1346, 1408, 1451, 1471, 1531, 1641, 1657, 1660, 1880, 1905, 1929, 1944, 2028, 2045, 2416, 2451, 2452, 2528, 2533, 2534, 2557, 2685, 2747, 2766, 2896, 2898, 2988, 3114, 3452, 3471, 3579, 3607, 3721, 3765, 3838, 3991, 3998, 4066, 4137, 4352]. Businesses [180, 482, 537, 840, 1455, 1468, 1603, 2155, 3588, 3605, 3835, 3849, 3875, 4078, 4236, 4247, 4283, 4328, 4427]. busted [2148, 2344]. Butterworth [3998]. Butterworth-Heinemann [3998]. buy [1729]. Buyer [1988]. buys [617, 1384, 1406, 1800, 2361, 2615, 3100].

BYOD [3296, 3394, 3437, 3478, 3617]. BYOE [4014]. bytecode [475].

C [209, 1867, 3373, 3891]. C# [3966]. C/S [209]. C2Net [930]. CA [1202, 1233, 1545, 2388, 3399]. Cable [418, 421]. cables [220, 2681]. Cache [25]. cached [681]. cafes [1818]. cage [849, 1019, 2408]. Cahoot [2332]. Caibua [273]. Caja [2759]. Calendar [1600, 1623, 1642, 1664, 1730, 1750, 1770, 1793, 1815, 1840, 1864, 1885, 1899, 1916, 1933, 1948, 1961, 1977, 1991, 2007, 2020, 2032, 2047, 2062, 2075, 2120, 2134, 2145, 2159, 2170, 2182, 2211, 2742, 2823, 2843, 2908, 2920, 2932, 2955, 2978, 2990, 3012, 3031, 3040, 3050, 3099, 3110, 3128, 3150, 3170, 3180, 3193, 3217, 3229, 3239, 3249, 3261, 3272, 3283, 3292, 3311, 3322, 3331, 3340, 3349, 3359, 3369, 3379, 3398, 3409, 3421, 3431, 3440, 3450, 3462, 3483, 3493, 3504, 3515, 3527, 3540, 3551, 3571, 3610, 3622, 3652, 3664, 3706]. Calendar [1333, 1562, 2090, 2105, 2197, 2224, 2240, 2254, 2266, 2279, 2291, 2305, 2318, 2331, 2345, 2358, 2372, 2387, 2399, 2423, 2434, 2445, 2457, 2467, 2478, 2489, 2501, 2511, 2522, 2532, 2546, 2560, 2572, 2585, 2593, 2604, 2614, 2625, 2635, 2645, 2656, 2665, 2675, 2683, 2693, 2702, 2711, 2722, 2732, 2753, 2764, 2775, 2787, 2799, 2812, 2833, 2856, 2865, 2875, 2887, 2897, 2943, 2967, 3002, 3021, 3062, 3075, 3087, 3120, 3139, 3159, 3205, 3301, 3388, 3472, 3562, 3580, 3590, 3886, 3951, 3962, 4140, 4166, 4241, 4265, 4277, 4299, 4370]. calendar [3600, 3632, 3642, 3674, 3685, 3695, 3716, 3726, 3737, 3747, 3757, 3767, 3779, 3790, 3800, 3809, 3819, 3830, 3840, 3859, 3870, 3878, 3898, 3908, 3918, 3930, 3940, 3972, 3984, 3995, 4004, 4015, 4024, 4034, 4042, 4051, 4060, 4072, 4084, 4095, 4106, 4116, 4128, 4153, 4180, 4193, 4206, 4217, 4229, 4253, 4289, 4311, 4322, 4333, 4347, 4357, 4383, 4396, 4408, 4420, 4432, 4443, 4455, 4475]. Calender [1707]. California [1741]. Californian [503, 1794, 2562]. Call [895, 986, 1247, 2280, 2537, 2539, 3556, 4372]. Callas [3492]. Called [1157, 1834]. Calling [59, 3491, 3611]. calling-card [59]. calls [246, 400, 748, 1251, 2040, 2146, 2826, 3114, 3763]. Cambridge [3444]. Cameron [3988]. Campaign [274]. Can [105, 371, 495, 1571, 1572, 1653, 1731, 1787, 1811, 1863, 1889, 1939, 2073, 2085, 2212, 2275, 2285, 2287, 3336, 3559, 3588, 3723, 3866, 3959, 3978, 4047, 4082, 4090, 4203, 4247, 4283, 4364, 4417, 4426, 4449, 4463]. Canada [189]. Canadian [1586]. Canal [1912]. cancelled [3581, 4465]. Can’t [1439, 1624, 2287, 2472, 3026, 3691, 3724, 4332]. capabilities [829, 912, 915]. capability [1490]. capable [557]. capital [483]. Cappuccino [2612]. Captcha [2734]. Car [3811, 3813]. Card [59, 179, 540, 916, 1050, 1248, 1320, 1406, 1418, 1431, 1467, 1502, 1504, 1522, 1525, 1830, 1946, 2437, 2743, 2840, 3084]. Card-Based [1418]. cardholder [2689]. cards [1304, 1374, 1567, 1583, 1660]. CardSystems [2435]. Cardwell [3845]. Care [1280, 1575, 2070, 2425, 2898, 3439]. careers [2805]. careless [2533]. Carnivore [1527, 1553, 1593]. Carnivores [1511]. Carpe [2989]. Carraig [1327]. carry [1564, 1753]. carrying [2592]. Cars [3745, 3764, 4389]. Carvey [3334]. Case [355, 453, 922, 1094, 1912, 2466, 2613, 3838]. cash [12, 464, 1069, 1165, 2936]. castle [3457]. casts [2743]. casual [1191]. cat [3038]. catapult [626]. catastrophic [763]. catch [753, 1738]. catchers [3912]. Catching [2848, 3010, 3058, 3339]. Cats [1723]. CATT [1806]. Caught [889, 1171, 1706, 2264, 2321, 3328]. cause [959, 1324, 3015, 3108]. caused [4029]. causes [1237, 1334, 2469, 2755, 2909, 3078]. Caution [1616]. CCTV [1826, 2523, 4038]. cdmount [1446]. CDs [456, 1748]. Cell [412, 985]. cell-phone [412]. Cellular [53, 63, 89, 117, 241, 353, 737, 898, 3244, 3357, 3481, 4248]. Censoring [1781]. censorship [1667, 2077, 3199]. central [3478]. centralised [4418]. centralizes [1142]. Centrally [2234]. Centre [115, 263, 1554, 1746, 2054, 4132]. centres [1476, 2100, 3828, 4283]. Centric [1958]. Cenzic [3005]. CEO [1907]. CERT [1698, 1720, 1828, 2052, 2081]. certain [347]. certainties [3420]. Certificate [1436, 1445, 3263, 3399]. Certificates [938, 1248, 1385, 1437, 1643, 3213]. certification [143, 607, 796, 810]. certifications [2647]. Certified [1031]. certifies [584]. CGI [510, 694]. chain [3978, 4191, 4298, 4392, 4444]. Chaining [2274]. Chairman [1972]. challenge [210, 490, 935, 2169, 2794, 2977, 3268, 3661]. challenged [3023]. Challenger [13]. Challenges [231, 1330, 2261, 3141, 3166, 3258, 3394, 3585, 3587, 3893, 4012, 4091, 4164, 4273, 4345]. Challenging [1393]. Champagne [1602]. champion [1542]. Chandler [3771]. Change [576, 836, 987, 1420, 2069, 2643, 3608, 3627, 3721, 3939, 4400]. Changes [27, 1342, 1421, 1860, 2040, 4295]. Changing [1950, 2455, 2487, 2623, 2664, 2763]. channel [2634]. Channels [2088]. chaos [604, 3296]. Characterising [2850, 2861]. charge [88]. charged [189, 2309]. charges [114]. charitable [2359]. Charles [3393]. chart [2187, 2201]. charts [2791]. Chauhan [3730]. cheap [3763]. cheaper [1797]. Check [379, 752, 1207, 1343, 2203, 2321, 2414, 4105]. Checkfree [314]. checking [3049]. checklist [3968]. Checks [97, 1162, 2119]. Chell [3720]. chemical [3264]. cheques [332]. Cherry [3741]. Chet [3614]. chicks [1814]. Child [190, 327, 1047, 2165, 4279]. child-safe [1047]. Children [1508, 4332]. China [353, 442, 1078, 1382, 1539, 1665, 1818, 1951, 3072, 3423, 3441, 3582, 4117, 4129, 4232]. Chinese [1233, 1383, 2926, 2944, 3070, 3370, 3433, 4348]. Chip [1107, 1958]. Chipping [1510]. chips [1255, 1747, 4117, 4323]. chkey [782].

Chooses [892, 1854]. Choosing [381]. Chris [3232, 3772, 3944]. Christine [3751]. Christmas [1842]. Christopher [3413]. Church [261]. Chuvakin [3374, 3413]. CIA [39, 1038]. CIO [1697, 4087]. Circle [3492]. Cisco [236, 505, 621, 725, 809, 936, 1046, 1080, 1119, 1181, 1199, 1555, 1657, 1924, 1937, 2085, 2108, 2214, 2229, 2271, 2376, 2413, 2447, 2476, 2815, 2968, 3209, 3210, 3221, 3222, 4043]. CISO [4369]. CISOs [4259]. CISSP [1701, 3414, 3824]. Citicorp [300]. Citizen [1958, 2680]. Citizens [1726]. Citrix [4301, 4313]. City [2835]. Civil [1511]. civilian [377]. CKS [546]. claim [763, 2777]. claims [331, 399, 3601, 3889, 4117, 4371]. clamp [1504]. clarity [3547]. class [460, 2402, 3049]. Classic [3457]. classification [3702, 3895]. clean [550, 1715, 2230]. Cleaning [3376]. clear [2437]. clearances [2226]. cleared [1692]. clears [1262]. Click [535, 1632, 3136]. clickjacking [3086]. Client [33, 178, 223, 424, 2117, 2699, 2716, 3166, 3215]. client-side [2716]. Client/server [223, 424]. clients [583, 3060]. Climate [3721]. Clinton [1251, 1505]. Clip [1378]. Clipper [3, 109, 292]. cloaking [1739]. clock [3524]. Close [2392]. closed [1873, 2263]. closer [2335]. closes [736, 1921]. closing [2569, 4296]. closure [2065]. Cloud [2984, 2996, 3070, 3146, 3174, 3190, 3215, 3258, 3363, 3376, 3415, 3426, 3488, 3514, 3549, 3628, 3722, 3723, 3746, 3789, 3818, 3957, 4012, 4139, 4187, 4252, 4263, 4330, 4343, 4368, 4415]. cloud-based [3957, 4012]. clouds [2883, 3386]. Club [473, 530]. CMA [2280]. CMS [3533]. CNN [826]. Co [105, 1972, 2611, 3524]. Co-Chairman [1972]. co-exist [105]. co-operation [3524]. Co. [2235]. Cocoa [3771]. Code [297, 352, 793, 1254, 1288, 1451, 1488, 1733, 1746, 1757, 1787, 1915, 2002, 2039, 2213, 2344, 2682, 3042, 3093, 3827, 4044, 4384]. codes [1396, 2207]. coding [619, 2140]. Cohen [970, 2089]. Cold [4138]. Colin [3458, 3514]. collaborate [29]. collaboration [1285, 2271]. Collaborative [1371]. collect [2844, 4090]. collection [40, 2189, 3959]. Collective [2808, 3611]. college [240]. Colley [3720]. collide [3212, 3913]. Collingbourne [3252]. colours [4059]. Columbia [1543]. Column [1704]. Combatants [612]. combated [1066]. Combating [2222]. Combatting [63]. Combination [2322, 2999, 3609]. combine [461, 871]. combined [976]. combines [2243]. COMDEX [1101, 1109]. come [20, 27, 102, 1032, 1704, 2648, 3591, 3871, 4041, 4321]. comeback [3169]. comes [2153, 2327, 2676, 2899, 3791, 3846, 3929, 4078, 4439]. Coming [1102, 1581, 2500, 3927]. command [862, 2672]. commandeered [69]. commands [885, 964, 3073]. Commerce [37, 249, 538, 541, 613, 640, 677, 735, 939, 972, 973, 977, 986, 990, 995, 1024, 1042, 1050, 1102, 1107, 1159, 1165, 1193, 1219, 1258, 1299, 1301, 1307, 1338, 1348, 1358, 1360, 1361, 1370, 1391, 1392, 1409, 1427, 1429, 1486, 1567, 1607, 1658, 1678, 1873, 1915]. Commercial [112, 1108]. commercialisation [2666]. Commissioners [1577]. commit [73, 2514]. committee [647, 1773, 4155]. common [924, 1336, 2825, 2910, 4379, 4454]. comms [6]. communicating [2496]. communication [3926, 4039]. Communications [30, 156, 287, 507, 604, 628, 645, 706, 766, 845, 1983, 3675, 3912]. community [2]. Compact [1203]. Companies [36, 374, 378, 538, 679, 696, 829, 928, 969, 991, 1017, 1062, 1084, 1085, 1112, 1185, 1205, 1219, 1233, 1271, 1299, 1302, 1325, 1356, 1386, 1387, 1402, 1408, 1420, 1428, 1953, 2261, 2972, 3586]. Company [490, 504, 540, 617, 622, 965, 1033, 1061, 1343, 1369, 1404, 1406, 2065, 2348, 3095]. Compaq [1214]. comparison [903]. comparisons [3107]. compete [2647]. competition [417, 4173]. competitive [1526, 3662]. competitor [398]. complacent [4078]. complains [155]. complement [3019]. completed [869]. Completely [847]. Complex [3844, 4069, 4187]. complexity [3147, 3755, 4235]. Compliance [927, 1271, 2401, 2797, 2818, 2988, 3137, 3374, 3491, 3958, 4081, 4176, 4251, 4270]. compliant [1205, 1716]. compliants [1798]. complies [636]. Components [3393]. comprehensive [1558]. Compress [441]. compromise [658, 821, 1379, 1489, 2577, 4242]. compromised [1583, 2274, 2931, 4021]. compromises [2144]. compromising [680, 1423, 3660, 3935]. Compsec [2000]. CompTIA [3266]. CompuServe [458, 844]. Computer [86, 167, 634, 655, 731, 754, 777, 799, 1039, 1178, 1420, 1475, 1614, 1785, 1950, 2003, 2217, 2280, 2327, 2479, 2830, 2958, 3197, 3508, 3545, 4008, 4372]. Computerized [2172]. Computers [115, 240, 1090, 1498, 3977]. Computing [815, 948, 1509, 1580, 1635, 1902, 2087, 2301, 2538, 2554, 2569, 2857, 3166, 3174, 3258, 4114, 4404]. concentrate [669]. concept [2741, 2890]. concern [1348, 2014, 3022, 3821]. concerned [1846]. concerns [125, 634, 992, 2078, 2286, 2303, 2694, 4184]. Concert [133]. concerted [3591]. Conference [263, 1723, 2152]. conferences [4465]. conferencing [1137]. Conficker [2935, 2959]. confidence [3882, 4270]. confident [440, 3919]. confidential [1695]. Confidentiality [945, 1659, 2228, 3049, 4067]. configuration [1142, 3061]. confirms [1366]. conflicts [1401]. confront [2435]. confronted [4194]. Confusion [2017, 3514]. Congress [797, 837, 3370]. Congressional [647, 1243]. Conjunction [1987]. Connect [824]. Connected [3596, 3764, 4149, 4225, 4389]. Connection [282, 847, 868, 1347, 3931]. connections [919, 1414]. connectivity [123, 139, 618, 1362]. conquer [2341, 3045]. Conrad [3414, 3824]. Cons [1278, 3020]. Consensus [1768]. consequences [3104]. Conservative [2436]. Considerations [389, 1507, 1744, 2099]. considered [1314]. considers [2257]. consigned [4010]. consolidation [3297]. Consortium [292, 341, 760, 788, 1336]. constitution [917]. constraints [946]. consultancy [521]. Consultant [2084, 2085]. Consulting [2086]. Consumers [1223, 1546, 1554, 3768]. contact [1567, 4283]. contain [456, 3780]. containers [3893]. contains [329]. contemporary [2405]. Content [611, 775, 949, 1191, 1305, 1456, 1596, 4365]. contest [366]. context [3537, 3670]. continents [872]. continue [1217, 2685, 2926]. continues [327, 886, 1166, 2199, 2922, 3033]. Continuity [1929, 2045, 2451, 2452, 2557, 3721, 4137]. Continuous [1271, 3672, 4081]. contract [1125, 1144, 1449]. Control [81, 268, 294, 315, 689, 1048, 1166, 1267, 1485, 1744, 2122, 2672, 2997, 3267, 3271, 3307, 3394, 3471, 3798, 3916, 3967, 3969, 4423, 4441]. controlled [2234]. Controller [2624, 3905]. Controlling [2759]. controls [939, 1266, 1761]. Controversy [109, 343, 1111, 2447, 2909]. conundrum [3320, 3536]. Convenience [1430]. Convention [1676]. Conventional [1393, 2069, 4441]. converge [4273]. converged [2443, 2505, 3211, 3212]. Convergence [2214, 2335, 2678, 3958]. Converging [2731]. conversion [673]. converted [4465]. convicted [239]. Cook [3402]. Cookbook [3383, 3509, 3668]. cookie [1425]. Cookies [723, 958, 1984]. Cooking [4071]. Cool [1898, 3892]. cooperate [1704]. cooperates [74]. cope [3516]. copiers [1191]. coping [3608]. copyright [637, 1305, 3052]. copyrights [100]. Coreflood [3194]. Corera [3977]. Corey [3535]. Coriolis [881]. Cornerstone [1369, 3923]. coronavirus [4426]. coronavirus-related [4426]. Corp. [781, 2084]. Corporate [116, 157, 293, 748, 813, 1177, 1640, 2121, 2797, 2975, 3117, 3247, 3712]. Corporates [1594, 2097, 2110, 3686]. Corporation [1557, 1596]. Corporations [592, 1706]. Correction [3233]. Corrupt [455]. Cost [1204, 1468, 1552, 1896, 2110, 2228, 2385, 2779, 2900, 3095, 3560, 3749, 3763, 4185, 4316]. Cost-cutting [1552]. costly [2495, 4463]. costs [188, 3076, 3218, 4073]. COTS [678]. Could [535, 840, 1397, 1441, 1920, 2273, 2956, 3279, 3579, 3883]. Council [1221, 1656]. councillors [47]. count [36]. countdown [2236]. Counter [1837]. counterfeit [155]. Countering [1394, 3106]. counterintelligence [2639]. Countermeasures [3833]. Counterpane [2615]. countersurveillance [3419]. Counting [3763]. countries [30]. courier [1559]. couriers [1559]. course [2112, 2277]. Court [608, 1772, 1785, 1794, 2962]. courting [1477]. covered [4312]. Covering [1809, 2088, 3907]. Covert [2088]. Covid [4348, 4361, 4465]. Covid-19 [4348, 4361, 4465]. CPRM [1604]. Crack [378, 1834, 1965, 2135, 2658, 2801, 2828]. cracked [469, 674, 695, 2053, 3152, 3484]. Cracker [1785]. Crackers [1395, 2751]. Cracking [18, 1029, 1915, 3013, 3043, 3103, 3189, 3244, 3270, 3329]. cracks [297, 1607]. Craig [3344, 3345, 3813]. crash [1046, 1693, 1760, 3108]. crashes [2359]. crater [2472]. crawlers [2206]. crawling [2121]. Crayon [1462]. craze [870]. create [367, 3978, 4016]. created [1390, 4343]. Creates [572, 596, 1343, 1783, 3332, 3624]. Creating [1768, 2916, 3828]. Creation [1433]. Credential [4461, 4464, 4473]. credentials [3601, 3831]. credibility [2394, 3797]. Credit [540, 916, 1374, 1504, 1522, 1583, 1946, 2437, 3084]. Credit-card [540]. Crime [18, 38, 98, 537, 677, 1039, 1178, 1239, 1420, 1478, 1481, 1521, 1557, 1571, 2208, 2474, 2551, 2836, 2939, 3841, 4120]. Crimes [453, 1331, 3531]. Criminals [2432, 3030, 3213, 3542, 3607, 4390]. cripple [3213, 3411]. crippling [1641]. criteria [3789]. Critical [1, 1261, 1471, 2386, 2452, 2544, 2758, 2977, 3307, 3342, 3523, 4013, 4155, 4263, 4313, 4366, 4386, 4405, 4406, 4449]. criticised [2723, 3860]. Criticized [292, 331, 697, 843, 929, 1614]. criticizes [668]. cron [173]. Cross [483, 679, 1873, 2002, 2578, 2602, 2940, 3328, 3651, 3679]. cross-layer [2940]. cross-platform [679]. Cross-site [1873, 2002, 2578, 2602, 3328]. Crossfire [1706]. crossroads [2807, 2819]. crumbling [3825]. Crume [1617]. cry [1824, 3971]. Crying [1766]. Cryptic [2082]. Crypto [794, 954, 1020, 1022, 1036, 1166, 1188, 1367, 1478, 1723, 1781, 1782, 1805, 1817, 1855, 2153, 2406, 2424, 2440, 2559, 3464, 4035]. Crypto-currencies [4035]. Cryptocurrency [3853]. Cryptographic [368, 704, 875, 931, 967, 990, 1049, 1115, 1196, 1723, 1956, 2632]. Cryptography [231, 507, 917, 1206, 1292, 1364, 1509, 1658, 1891, 1974, 1985, 4164, 4404]. CryptoLocker [3517, 3539]. cryptosystem [563, 1051, 1074]. crystal [1543]. CSAPP [715]. CSI [1960, 2867]. CSO [4378]. CSOs [2225]. CSS [3658]. CTO [2153, 2447, 2524]. Cuadra [3834]. Cuadra-Sánchez [3834]. CUDA [3402]. CueCat [1537]. Cultural [3107, 3385]. Culture [268, 294, 2034, 2045, 2736, 4079]. Cumming [3700]. Cup [2574]. Cupboard [1860]. curbs [1269]. cure [2382, 3865]. Curiosity [1537]. Curmudgeon [493]. currencies [4035]. Current [711, 2082]. cursor [2667]. curve [1074]. Custom [1890, 2251]. Customer [1467, 1692, 1863, 2471, 2854, 2979]. customer-targeted [2471]. customers [345, 1671, 2477, 2534, 3085]. cutting [1552, 2972]. cutting-edge [2972]. Cyber [95, 135, 622, 624, 1264, 1399, 1478, 1481, 1542, 1571, 1625, 1641, 1681, 1773, 1818, 1837, 1895, 1927, 2096, 2436, 2474, 2921, 2926, 2936, 2968, 2982, 3149, 3155, 3161, 3262, 3282, 3316, 3444, 3459, 3497, 3498, 3530, 3576, 3582, 3584, 3591, 3607, 3611, 3659, 3671, 3686, 3701, 3732, 3788, 3807, 3836, 3879, 3919, 3926, 3959, 3996, 4001, 4043, 4049, 4052, 4058, 4066, 4077, 4085, 4114, 4125, 4132, 4143, 4159, 4179, 4197, 4204, 4210, 4216, 4240, 4245, 4257, 4275, 4286, 4304, 4309, 4320, 4345, 4375, 4390, 4406, 4407, 4415, 4431, 4433, 4446, 4471, 4472]. cyber [887, 2339, 2939, 3317, 3864, 3896, 3958, 4063, 4296, 4439]. Cyber-Accountability [4446]. cyber-attack [2926, 3591, 3926]. Cyber-attacks [1927, 3459, 3584, 3611, 3671, 3686, 3996, 4049, 4052, 4058]. Cyber-breaches [4043]. cyber-buck [95]. cyber-cafes [1818]. cyber-crime [1478, 2939]. cyber-criminals [3607]. cyber-distributor [135]. Cyber-Elephant [4446]. Cyber-espionage [2921]. cyber-resilience [3317]. cyber-risk [4125]. Cyber-Sabotage [1641]. cyber-safe [622]. Cyber-security [1773, 3149, 3161, 3262, 3282, 3498, 3530, 3576, 3659, 3732, 3836, 3879, 3896, 3919, 3958, 3959, 4063, 4066, 4085, 4114]. cyber-squatters [1264]. Cyber-swine [2936]. Cyber-terrorism [1895]. Cyber-terrorist [2436]. Cyber-terrorists [1625, 1837]. cyber-threats [2339]. Cyber-Warfare [3497]. CyberAngels [316]. cyberattack [2899]. CyberCash [314, 354]. cybercops [324]. Cybercrime [988, 1587, 1629, 1676, 1768, 1816, 2617, 2839, 3038, 3091, 3247, 3313, 3451, 3513, 3599, 3749, 3841, 3854, 3998, 4185, 4189, 4256, 4283, 4336, 4426]. Cyberflex [1110]. cyberfraud [503]. CyberGuard [204, 443, 466, 539, 558, 810, 830]. CyberSAFE [13]. Cybersecurity [1994, 2022, 2040, 2174, 2858, 2898, 3566, 4054, 4352]. CyberSource [135]. Cyberspace [129, 272, 548, 775, 881, 1295, 1557, 1932, 1993, 2919, 4230]. cyberspying [3072]. Cybersquatting [1218, 1415]. cyberterrorism [1276]. cyberthreat [3915]. cyberthreats [3160, 4266]. Cyberwar [2924, 3423, 3566]. cycles [731]. Cylink [1234]. Cypress [4323]. czar [794]. Czech [1516].

D [3017]. daemon [351, 606, 1118]. Dafydd [3295]. Dalziel [3783, 3784, 3804, 3814, 3823, 3845]. damage [1129, 3234, 3567, 4073, 4188]. damages [1192]. damaging [3799]. Danger [1879, 2883, 2886, 3897, 4178]. Dangerous [126, 1672, 1910, 2393, 3856]. Dangers [1395, 3704, 4303]. Dario [2088, 2100]. Dark [1915, 2582, 2986, 3569, 3682, 3831, 3973]. Darkness [1534]. darknet [2954, 3707]. darknets [3289]. Data [5, 6, 66, 103, 134, 199, 248, 255, 256, 318, 386, 441, 619, 747, 766, 767, 850, 883, 948, 969, 1217, 1235, 1249, 1255, 1476, 1577, 1630, 1695, 1820, 1829, 1877, 1921, 1980, 2142, 2533, 2583, 2588, 2673, 2743, 2754, 2766, 2979, 3015, 3064, 3085, 3138, 3158, 3178, 3204, 3219, 3344, 3345, 3446, 3460, 3470, 3487, 3512, 3537, 3579, 3589, 3605, 3643, 3665, 3694, 3702, 3768, 3833, 3848, 3861, 3895, 3902, 3913, 3959, 3992, 3993, 4009, 4090, 4168, 4170, 4176, 4181, 4242, 4254, 4263, 4307, 4334, 4378, 4418, 4471]. data [19, 432, 1195, 1929, 2199, 2302, 2396, 2472, 2529, 2566, 2568, 2592, 2601, 2610, 2820, 2975, 3175, 3202, 3246, 3346, 3394, 3428, 3490, 3499, 3522, 3536, 3547, 3585, 3618, 3631, 3638, 3649, 3659, 3661, 3712, 3715, 3734, 3746, 3808, 3873, 3923, 3980, 4079, 4083, 4103, 4135, 4364, 4365, 4451]. data-security [1255]. Database [431, 1284, 1655, 1700, 2116, 2128, 2472, 2636, 2700, 2814, 2832, 3007]. datacentre [3106, 3134, 3248, 4199, 4290]. datamedia [118]. DataVault [3202]. date [3694]. dating [3739]. Dave [3892]. David [3243, 3794, 3976]. Davie [3197]. Davies [3863]. Davison [3976]. Dawn [4009]. Day [1616, 1720, 1898, 1969, 2514, 3612]. days [1709, 2473, 3920]. DCAs [1394]. DCE [867]. DCIM [3416]. DCMS [4197]. DCOM [2158]. DCS [27]. DDOS [1925, 1982, 2010, 2025, 2101, 2256, 2400, 2829, 3273, 3278, 3358, 3432, 3445, 3542, 3634, 3697, 3738, 3756, 3883, 4078, 4093, 4178, 4306, 4340, 4410]. De-perimeterisation [2404, 2903]. dead [2373, 2463, 3355]. deadlock [1037]. deadly [2307]. Deal [716, 1970]. Dealing [2730, 2886, 3511, 4395]. Death [1657, 2177, 2907, 3420, 3725]. Debate [327, 704, 795, 967, 994, 1575, 2332, 2388, 3078, 3332]. debates [478]. Debian [1154]. debit [1467]. debt [4395]. debunks [2174]. debut [764]. DEC [532]. decade [4199, 4337]. decade-old [4199]. deceit [2359]. December [2201]. Decency [604, 628, 706]. Deception [1813, 1990, 3988]. Deciding [1436]. Decisions [269, 1634]. declared [1751]. Decrypting [4251]. decryption [2824]. DeCSS [1624, 1675, 1794]. dedicated [812, 3981]. Deep [1313, 2453]. deepens [2160]. deeper [3281]. defaced [1709, 1795]. Defacing [2760]. Defamatory [1648]. DefCon [2297, 3101, 3230]. defeat [902, 2325, 2400]. defeated [3759]. defeating [4214]. defectors [4158]. Defence [452, 980, 1224, 1371, 1621, 1813, 2118, 2190, 2454, 2599, 2730, 2982, 3255, 3306, 3468, 3611, 3916, 4245, 4340, 4467]. defence-in-depth [2454, 2599]. defences [960, 2832, 3395, 3586, 3989]. Defend [1917, 3783]. Defender [202, 1106, 1440]. Defending [2115, 2700, 2860, 3256, 3607, 4068]. Defense [2206, 4413]. define [1336]. defines [313]. defining [3866, 4438]. definitive [1618]. defraud [1822]. delay [2077, 4182]. Delayed [50, 1722]. Delays [1699]. deliberate [3780]. Delivering [3684]. delivers [1201, 1902]. delivery [1062, 3020, 3489]. Dell [1063]. Deluxe [3266]. demand [42, 420, 2343]. demands [1404, 3753, 3963]. demo [456]. democracy [4126]. demonstrate [1920, 1949]. demonstrates [2815]. Demystifying [2017, 4250]. Denial [437, 654, 867, 905, 943, 959, 1119, 1198, 1230, 1334, 1350, 1443, 1477, 1720, 1834, 1925, 2101, 2957, 3068, 3257, 3805, 3992]. Denial-of-service [437, 867, 905, 943, 959, 1119, 1230, 1334, 1350, 1443, 1477, 1720, 1834, 1925]. denies [1373]. Denning [2426]. Denny [3741]. Department [953, 1261]. departments [2444, 4085]. Deploy [1004]. Deploying [2130, 2353]. Deployment [447, 1203, 1833, 2421, 3500, 4150]. deploys [2333]. depth [2118, 2454, 2599, 2942]. DERA [1696]. derived [963]. Derrick [3184]. DeSanto [3976]. descriptions [1964]. deserves [262]. Design [82, 2099, 2599, 2940, 2998, 3003, 3200, 4380, 4394]. Designing [2028]. designs [515]. Desktop [512, 666, 924, 2881]. Desperados [1989]. despite [440]. Dessart [3210]. destined [2181]. destroy [3366, 3643]. destroying [3806]. destruction [3428]. destructive [4224]. Detail [2131]. details [1156, 1467, 2437, 2564, 2575, 2778]. detect [774]. detected [488]. Detecting [3226, 3713, 3806]. Detection [745, 746, 902, 920, 934, 974, 1032, 1127, 1256, 1270, 1345, 1421, 1482, 1523, 1789, 1833, 1834, 1938, 1955, 1987, 2017, 2073, 2074, 2093, 2246, 2730, 2850, 2861, 2894, 2940, 2951, 2980, 3269, 3486, 3534, 3701, 3834, 3865, 4329, 4355, 4403, 4451]. detections [2952]. detects [1524]. deter [3236]. determining [3903]. deterred [607]. detox [3649]. develop [470, 542, 1085, 1233]. developer [188, 3418]. developers [3077, 3581, 3978]. Developing [293, 2804]. development [3479, 3774, 4011, 4029, 4377]. Developments [213, 499, 2153]. develops [8, 2982]. Device [259, 285, 615, 643, 789, 1202, 2139, 2216, 2521, 3548, 4268, 4300]. Devices [199, 295, 1906, 1973, 2529, 2579, 2584, 2589, 3298, 3308, 3509, 3544, 3596, 3602, 3638, 3646, 3672, 3727, 3778, 3969, 4006, 4021, 4057, 4225]. devil [233, 2425]. devil-may-care [2425]. devious [2198]. Devon [3243]. DevOps [4094]. DevSecOps [4436]. DHCP [1923, 2115]. DHS [2723, 2992]. diagnostic [521]. dial [524, 1903]. dial-up [1903]. Diary [3287]. Did [3108, 3874]. DIDS [2073]. dies [2424]. difference [1925]. Different [1274, 2195]. Difficult [1702]. Diffie [888]. Diffusion [3844]. Dig [2453, 2472]. Digital [355, 476, 576, 578, 1049, 1086, 1155, 1158, 1214, 1231, 1248, 1332, 1355, 1385, 1405, 1437, 1440, 1505, 1557, 1575, 1989, 2125, 2189, 2250, 2820, 3127, 3213, 3326, 3327, 3514, 3525, 3556, 3575, 3587, 3733, 3742, 3911, 3979, 3988, 4134, 4136, 4170, 4413, 4418, 4430]. dilemma [148, 3669]. Dimension [1894]. dimensions [4150]. dip [533]. dire [2280]. direct [252, 1195]. Director [2241]. directories [510]. directors [1571]. Directory [1366, 3318]. dirty [3376]. disappear [2292]. disappearing [3982]. disappointed [1325]. disarray [1964]. Disaster [34, 94, 948, 1477, 2000, 2363, 3488, 3561, 4137]. discipline [599, 3178]. disclose [4142]. Disclosure [794, 1967, 2024, 2185, 2388, 2555, 2653, 3078, 3096]. discontinues [1685]. discovered [751, 1135, 1516, 1883, 2921]. Discovering [1005]. discovers [2959]. Discovery [2139, 2300, 2330, 2555, 2655, 2663, 3788, 4083, 4103]. discredits [2890]. disgruntled [2956]. dish [1644]. disinfect [3869]. disks [3376]. dismantles [2547]. dismissed [114]. disorganized [2008]. dispute [1217]. Disputes [1392]. disrupt [1041, 3273]. disruption [3624]. Dissecting [3200, 3947, 4319]. disseminating [1192]. disservices [1164]. distance [400]. distraction [3240]. distribute [1140]. Distributed [1925, 2073, 2101, 2709, 3805, 3992]. Distributed-Denial-of-Service [1925]. distributes [415]. Distributing [835]. Distribution [174, 2699, 4039, 4233]. Distributor [135, 1304]. Divide [1337, 2341, 3045]. divided [739]. DIY [3073, 4429]. DMCA [1762, 1790]. DNC [3871, 3874]. DNS [1439, 1549, 1633, 2107, 2115, 2640, 2834, 3014, 3557, 3588, 3595, 3754, 3829, 3846, 3881, 3970, 4073, 4297, 4471]. DNS-related [4073]. DNSSEC [1576, 3080, 3577]. Do [106, 371, 648, 1439, 1459, 1460, 1814, 2069, 2070, 2140, 2212, 2219, 2326, 2802, 2828, 2844, 2853, 3618, 3914, 3939, 4203, 4247, 4262]. Doctrine [2919]. Document [329, 423, 1559, 2697, 2851, 3947]. Documenting [2089]. Documents [1087, 1553, 1593, 2144]. DoD [678, 1726, 2723]. dodge [1936, 3358]. dodging [2506]. dodgy [4374]. Does [230, 1023, 1459, 1679, 1884, 2338, 3874, 4394]. Doesn’t [1824, 1838, 3348]. doing [2385, 3037, 3752]. dollar [887]. Dollars [2893]. domain [325, 422, 1218, 2025, 2624, 3418, 3651, 4471]. domains [3442, 4285]. domestic [1726]. dominant [416]. Dominic [3720]. Domino [721]. done [1311]. Donn [1178]. Don’t [1462, 1617, 1729, 1890, 2273, 3337, 3579, 4175]. Doom [1462]. door [3001, 3091, 3686, 3780, 4038, 4461, 4473]. doors [4202]. Dorothy [2426]. Dot [1832]. Dot-NET [1832]. dots [2941, 4176]. double [1778, 1821, 3097]. doubled [1631]. doubt [243, 2569]. Doubts [594, 3653]. down [18, 378, 734, 1397, 1403, 1504, 1709, 1711, 1818, 1965, 2113, 2195, 2568, 3109, 3129, 3201, 3219, 3338, 3432, 3442, 3494, 3973, 4086, 4119, 4279, 4325, 4349]. Downadup [2888]. download [382, 1220, 1283, 1463, 1874, 2778]. Downloader [2013]. downloading [2893]. downplays [2777]. downtime [4463]. dozens [750, 4312]. Dr [2153]. Draft [160, 1325, 1367]. Dramatic [393, 4254]. dream [4201]. drive [197, 2343, 2375, 2778]. drive-by [2778]. driven [832, 3735, 4174, 4341]. driver [285]. drivers [2365]. drives [1867, 2355, 3643, 3676]. drops [2636, 3517]. drug [3494]. DTI [763, 1420]. dtmail [1122]. dual [467]. Due [4465]. Dulaney [3266]. dumb [4259]. duped [1643]. during [3926]. Dutch [1796]. DVD [1310]. dxconsole [476]. dying [3796]. dynamic [2640]. DynaSoft [74].

E-Biz [1662]. e-book [3352, 3353, 3614]. E-Business [1408, 1451, 1471, 1641, 1657, 1660, 1880, 2028]. E-Commerce [735, 972, 977, 986, 990, 995, 1050, 1102, 1107, 1165, 1193, 1219, 1258, 1299, 1301, 1307, 1338, 1348, 1358, 1360, 1361, 1370, 1392, 1409, 1427, 1429, 1486, 1658, 1678, 1873, 1915]. e-crime [2551]. e-discovery [2655]. E-government [1286]. e-ISBN [3184]. e-learning [3176, 3291]. E-mail [79, 120, 126, 148, 149, 232, 288, 310, 394, 402, 528, 567, 581, 720, 736, 741, 834, 846, 851, 865, 874, 932, 945, 949, 1099, 1157, 1164, 1177, 1388, 1416, 1499, 1524, 1545, 1636]. E-merchants [1300]. E-money [1584]. E-payment [1319]. E-voting [2739]. E-warehouse [1852]. E-word [1329]. E3 [810]. E security [1674]. Eagle [580, 659]. Early [2084]. Earth [3679]. eased [477]. easier [3934]. East [2924, 3285, 3361, 4278]. Easy [427, 722, 2441, 3164, 3482, 3727]. eating [723]. eats [1652]. eBay [2308]. ebook [3425]. EBPP [1428]. Ebsworth [3480]. EC [1716, 1994]. ECB [4254]. ECHELON [1243, 1693, 1726]. ECHR [1716]. ECMAScript [3891]. economy [1641]. ecosystem [3263, 3807]. Ed [3425]. edge [2312, 2972, 3662]. Edgescan [4326]. EDI [91, 450, 898]. EDI-based [898]. Edinburgh [1723]. Edited [3444, 3751, 3803, 3901, 4008]. Edition [3197, 3232, 3266, 3276, 3295, 3334, 3374, 3383, 3555, 3658, 3751, 3771, 3772, 3824, 3892, 4008].

Editor [2138]. Editorial [904, 1117, 1316]. EDS [558]. educate [1376]. Education [1, 1277, 1617, 2541, 2852, 4020]. Edward [3853]. Effect [1680]. Effective [1, 1945, 2421, 2462, 3247, 4103, 4227]. Effectiveness [56]. effects [1641, 2655]. effort [307, 900, 1344, 2933]. efforts [2797]. eID [2840]. eight [1754]. election [2389, 3871, 4397]. election-voting [2389]. elections [3273]. electoral [3952]. electric [283]. electrical [2931]. electricity [4096]. Electronic [12, 138, 148, 149, 249, 260, 279, 287, 318, 332, 464, 483, 514, 541, 639, 643, 645, 677, 939, 1020, 1024, 1042, 1086, 1087, 1391, 1457, 1559, 1570, 1575, 1650, 2397, 3393, 3419, 4365]. electronically [1088]. Elenkov [3710]. Elephant [4446]. Eliminate [162, 506]. eliminated [488]. Eliminating [1419]. elliptic [1074]. else [1743]. Elsevier [3742]. elusive [2276]. Email [1760, 1797, 1812, 1844, 2494, 3559, 3563, 4012, 4222]. emails [2198, 3961, 4063, 4374]. embarrassment [1905]. Embassy [3992]. Embedded [1882, 2365]. embraces [364]. emerge [594, 3759, 3851]. emergence [592]. emergency [2866]. Emerges [1576, 3032, 3168]. Emmett [3266]. Emotet [4224]. employee [1475, 2337]. Employees [982, 1641, 1704, 2806, 2956, 3826]. Employer [1529]. EMV [1611]. enable [1161, 1339]. Enabled [501, 1637]. enablement [4352]. enables [716, 1110, 1253, 1503]. Enabling [1370, 3296, 3310, 3936]. encode [2826]. encouraged [1060]. Encouraging [4228]. Encrypt [441, 747, 1235]. Encrypted [835, 845, 911, 2584, 4212]. encrypting [4136]. Encryption [11, 151, 199, 292, 301, 303, 362, 408, 413, 414, 420, 459, 477, 498, 505, 514, 540, 579, 583, 636, 647, 657, 658, 666, 674, 679, 685, 695– 697, 714, 717, 765, 768, 787, 790, 791, 795, 797, 843, 850, 856, 876, 878, 888, 947, 968, 972, 994, 1006, 1021, 1052, 1058, 1082, 1173, 1210, 1249, 1269, 1293, 1309, 1317, 1325, 1354, 1382, 1403, 1411, 1508, 1659, 1719, 1732, 1788, 1956, 2153, 2241, 2561, 2881, 3018, 3084, 3107, 3320, 3484, 3487, 3511, 3541, 3628, 3758, 3781, 3857, 3923, 4039, 4319, 4429]. encryption [1037, 1223, 1344, 3663]. encryptor [361, 1088, 1234, 1501]. Encryptors [1186]. encrypts [248]. Encyclopedia [3393]. End [229, 551, 696, 1861, 2441, 2473, 3093, 3147, 3670, 4318, 4393]. end-to-end [4318, 4393]. end-user [3093, 3147]. end-users [2441]. Endorsements [164]. Endpoint [2567, 2740, 4145, 4201]. endpoints [2995, 4392]. ends [99, 3564]. enemy [2914, 2964, 4165]. Energy [1741, 3298, 3591, 3799, 3985]. enforces [1162]. Enforcing [2770]. Engebretson [3277]. engine [362, 1070, 1245, 1284, 2781, 2980]. engineer [191, 495]. Engineering [267, 1723, 1724, 2033, 2198, 3135, 3176, 3637, 3901, 3941]. Engineering-Based [3901]. engines [2548, 2774]. Enhanced [500, 561, 664, 718, 726, 784, 1089, 1450, 1593, 2987]. Enhancement [1412]. Enhancements [466, 793, 1009, 1414]. enhances [1318, 1452]. ENISA [4159]. enlists [1520]. enough [2766, 3541, 3743]. ensnare [3941]. Ensuring [4406]. Entegrity [1069]. enter [1872]. Enterasys [2447]. Entercept [1609]. Enterprise [522, 559, 682, 897, 948, 1032, 1228, 1886, 2078, 2083, 2361, 2483, 2821, 3144, 3209, 3457, 3512, 3648, 3784, 4038, 4091, 4412, 4424]. Enterprise-grade [4091]. enterprise-wide [522]. Enterprises [783, 1680, 2995, 4444]. Entertainment [1007]. Entitlement [2773]. entitlements [2762]. Entrust [1125, 1183, 1220]. envelope [851, 2658]. Environment [33, 924, 1279, 1390, 1880, 2339, 2443, 2644, 2841, 2870, 3059, 3211, 3308, 3597, 4047]. environments [646, 2581, 2917, 3561, 4069, 4187]. epidemic [2788, 3091, 4201]. epilepsy [2790]. equals [4381]. Equant [1519]. Equifax [4029, 4062]. equipment [3004, 3008, 3370, 4406]. era [71, 418]. erase [3605]. Erasmus [3720]. Eric [3414, 3678, 3824]. Ericsson [74]. Ernst [781]. Erratum [2725]. error [1014, 3399]. ESB [3290]. escalating [4440]. Escape [1703, 2342, 2395]. Escrow [478, 593, 797, 993]. eSignature [3929]. especially [3552]. Espionage [111, 2480, 2921, 3599, 3616]. Espionage-linked [2480]. Essential [939, 2903, 3845]. Essentials [4240, 4342]. Essinger [1763]. Establish [365, 1697]. Establishes [1829]. Establishing [147]. Estonia [2924, 3348]. Estonian [3992]. Ethernet [179]. Ethical [1839, 1889, 2465, 3225, 3354]. ethically [4417]. eToken [1551]. eTrust [703]. EU [99, 1157, 1217, 1631, 1656, 1717, 1983, 2199, 2230, 2789, 3203, 3605, 3712]. EU-US [1217]. EUNet [282]. EUR [2230]. Europe [248, 303, 808, 968, 1060, 1319, 1656, 2512, 3001, 3107, 3448, 4159]. European [160, 979, 1354, 1355, 1407, 2051, 2574, 2840, 3161, 3665, 3985]. Europol [3718, 4120]. Eurovision [2259]. evaluates [2779]. evaluating [2277, 4441]. evaluation [3081]. evaluations [650, 670]. Evan [3253]. evasion [3156]. even [601, 1937]. Event [1960, 1961, 4451]. Eventful [2232]. Events [35, 58, 84, 107, 113, 127, 150, 171, 186, 211, 235, 270, 296, 319, 348, 372, 391, 410, 434, 454, 474, 496, 509, 531, 553, 568, 588, 605, 629, 651, 671, 691, 708, 728, 756, 776, 820, 839, 861, 923, 941, 998, 1026, 1043, 1056, 1076, 1095, 1116, 1134, 1150, 1169, 1179, 1197, 1213, 1227, 1244, 1265, 1281, 1296, 1315, 1333, 1562, 1600, 1623, 1642, 1664, 1684, 1707, 1730, 1750, 1770, 1793, 1815, 1840, 1864, 1885, 1899, 1916, 1933, 1948, 1977, 1991, 2007, 2020, 2032, 2047, 2062, 2075, 2090, 2105]. Events [2120, 2134, 2145, 2159, 2170, 2182, 2197, 2211, 2224, 2240, 2254, 2266, 2279, 2291, 2305, 2318, 2331, 2345, 2358, 2372, 2387, 2399, 2423, 2434, 2445, 2457, 2467, 2478, 2489, 2501, 2511, 2522, 2532, 2546, 2560, 2572, 2585, 2593, 2604, 2614, 2625, 2635, 2645, 2656, 2665, 2675, 2683, 2693, 2702, 2711, 2722, 2732, 2742, 2753, 2764, 2775, 2787, 2799, 2812, 2823, 2833, 2843, 2856, 2865, 2875, 2887, 2897, 2908, 2920, 2932, 2943, 2955, 2967, 2978, 2990, 3002, 3012, 3021, 3031, 3040, 3050, 3062, 3075, 3087, 3099, 3110, 3120, 3128]. Events [3139, 3150, 3159, 3170, 3180, 3205, 3217, 3229, 3239, 3249, 3261, 3272, 3283, 3292, 3301, 3311, 3322, 3331, 3340, 3349, 3359, 3369, 3379, 3388, 3398, 3409, 3421, 3431, 3440, 3450, 3462, 3472, 3483, 3493, 3504, 3515, 3527, 3540, 3551, 3562, 3571, 3580, 3590, 3600, 3610, 3622, 3632, 3642, 3652, 3664, 3674, 3685, 3695, 3706, 3716, 3726, 3737, 3747, 3757, 3767, 3779, 3790, 3800, 3809, 3819, 3830, 3840, 3850, 3859, 3870, 3878, 3886, 3898, 3908, 3918, 3930, 3940, 3951, 3962, 3972, 3984, 3995, 4004, 4015, 4024, 4034, 4042, 4051]. Events [2882, 4060, 4072, 4084, 4095, 4106, 4116, 4128, 4140, 4153, 4166, 4180, 4193, 4206, 4217, 4229, 4241, 4253, 4265, 4277, 4289, 4299, 4311, 4322, 4333, 4347, 4357, 4370, 4383, 4396, 4408, 4420, 4432, 4443, 4455, 4465, 4475]. Ever [1767, 2395, 2710, 3564, 3601, 3909]. ever-wider [3909]. Every [846, 3238]. everybody [2086]. everyman [2874]. everyone [2783]. everything [1009, 1560, 1908, 2004]. eviction [1264]. Evidence [2189, 4037]. evolution [1293, 2485, 2737, 2829, 3481, 3578, 3681, 3756, 3787]. Evolutionary [2471]. evolve [2926]. evolving [2710, 3405]. ex [1261, 1641]. ex-staffers [1261]. ex-technical [1641]. Exactly [838, 859, 880, 1479]. exaggerated [3495]. example [1679]. exceeding [986]. Excel [557]. exchange [579, 1156]. executes [885, 1488]. execution [2871]. Executive [2708, 3882]. Exim [4358]. exist [105, 1884]. existence [4433]. expand [1200]. expanded [1175]. expands [363, 1234, 1404, 3801]. expensive [330]. experience [616, 1744]. Experienced [1420]. experimental [3081]. Experiments [1879]. Expert [929, 1732, 2174, 2658]. Experts [763, 1425, 2858, 3623]. explained [52, 2206]. explicit [309]. explodes [916, 2788]. Exploit [680, 1136, 2136, 3422, 4361]. Exploitable [1377, 4415]. exploitation [1627, 4279]. exploited [942, 2213, 4337, 4409, 4452]. Exploiting [2870, 2963, 3906]. exploits [2267, 2480, 2793, 3146, 3371, 3395, 3623]. Explore [1577]. Explorer [693, 709, 866, 1045, 1079, 1124, 2281, 2364, 3032]. Exploring [2782]. explosion [1102]. Export [368, 615, 704, 790, 807, 825, 878, 909, 1166, 1188, 1196, 1269, 1309, 1317, 1367, 1478]. Exportable [301, 408]. exports [292, 946]. exposed [1084, 1648, 2036, 3460, 4224, 4283]. exposing [4427]. exposure [2388, 2597, 2905]. expreserve [569]. Express [601, 2333, 3092]. Expressions [3383]. Extended [443, 4295]. Extending [637, 2622]. extensions [3200]. extent [489]. exterminate [2293]. External [1045, 2825]. Extinguish [1914]. Extortion [1645, 3052, 4216]. extortionists [2066]. extra [2443]. extradited [2549]. extranet [1357, 1519]. extranets [2353]. extremely [2386]. eye [2407]. Eying [2633].

FAA [1920]. Face [260, 840, 1254, 1259, 2623, 2763, 3237, 3799, 3855]. Facebook [3063, 3086, 3377, 3422, 3941, 4194]. faced [2261]. faces [4179]. facility [4456]. facing [2155, 3993]. fact [2153, 3722, 4192]. factor [1410, 2543, 2701, 3188, 3609]. Facts [3855, 3939]. Fadyushin [3435]. fail [688, 966, 1138, 1695, 3769, 4262]. Failed [2979]. failing [1560, 3516, 3573, 4142]. fails [1228, 2300, 2935, 3004, 3274, 3432, 4219]. failure [1352]. failures [1198]. Faircloth [3276]. fake [3052, 4107]. Fall [1791]. fallacy [3839, 4205, 4393]. fallout [4457]. falls [4441]. False [1848, 2884, 3905]. familiar [3479]. Family [1006, 2149, 2413, 2603]. fantasy [1168]. far [266, 740, 2412]. farewell [1844]. Fashion [2294]. fast [4410]. faster [2050, 2162]. fattens [2229]. Fault [139, 201, 376, 1645]. fault-tolerance [376]. Fault-tolerant [139]. Favour [1418, 1805]. FBI [343, 375, 845, 889, 1420, 1493, 1553, 1672, 1746, 2241, 2458, 2684, 3360, 3811, 3847]. fcagent [1118]. FDA [4268]. fear [592, 677, 937, 2319, 3377]. Fears [215, 571, 613, 1038, 1400, 1495, 2303, 3072, 3722]. features [340, 718, 784, 2587, 2840]. February [2232]. federal [278, 1208, 1512, 1697, 3723]. federated [2335, 2355, 2466, 2794, 2982]. Feds [787, 2945]. feed [3579, 4175]. feeling [1718]. Feldman [3414, 3824]. Felten [3853]. femtocells [3244]. fence [2344]. fest [2469]. Festival [2152]. feud [1006]. Few [3182]. Fibre [8, 194, 220, 2681]. fibre-optic [8, 220]. fiction [2153, 4192]. Field [3519]. fierce [3078]. Fifth [1704, 3197, 3768, 3771]. fight [274, 887, 1665, 1800, 2278, 2428, 3077, 3451, 4122, 4239, 4391, 4406]. Fighting [130, 1178, 1812, 3356, 3957, 4261]. fights [214, 218, 326, 1506, 1757]. File [717, 942, 971, 1001, 1062, 1252, 1452, 1618, 2333, 2963, 3052, 3490, 3546, 4416]. Fileless [3935, 4050, 4261]. Files [802, 828, 911, 1351, 1432, 1636, 2347, 2769, 2811, 2966, 3643]. filling [2670]. film [1687]. filter [290]. Filtering [912, 1456, 1744]. Final [837, 2236, 3130]. finally [305, 1058, 3796]. Finance [1780]. Financial [102, 256, 1015, 1096, 1211, 1544, 1704, 1723, 2239, 3313, 3573, 3919]. Find [1425, 2106]. Finding [1211, 2058, 2371, 3597, 4067, 4094]. Finds [961, 1420, 1467, 4315, 4423]. Fine [2513, 4188]. Fine-tuning [2513]. fined [60]. Finest [1598]. Finger [2957, 3423]. fingerprint [639, 789, 1906, 3525]. fingerprints [3127]. finite [767]. Finjan [699]. Finland [1231]. fire [217, 438, 1148, 2502, 3328, 3556, 3791, 4082, 4300]. FireEye [3641]. firestorm [979]. Firewall [49, 140, 339, 341, 357, 359, 365, 403, 415, 443, 451, 467, 472, 490, 499, 523, 529, 597, 621, 626, 688, 726, 769, 784, 815, 834, 886, 912, 1008, 1010, 1025, 1054, 1067, 1128, 1142, 1203, 1356, 1381, 1387, 1441, 1466, 1476, 1556, 2067, 2243, 2508, 2985, 3083, 3105, 3165, 3167, 3195, 3929]. FireWall-1 [1381]. firewalled [2581]. Firewalls [54, 217, 284, 370, 526, 686, 688, 748, 785, 799, 819, 1063, 1064, 1558, 1834, 1987, 2005, 2366, 2464, 2599, 2670, 3221, 3222, 3397, 3500, 3780, 4424, 4450]. firm [765, 1406, 1851, 2979, 3601, 3675, 3717]. Firms [255, 1714, 1801, 1803, 2332, 2712, 2846, 3219, 3264, 3361, 3370, 3482, 3516, 3573, 3591, 3592, 3758, 3792, 3799, 3820, 3861, 3927, 3942, 4090, 4107, 4142, 4182, 4232, 4301, 4384, 4435, 4467]. First [399, 451, 557, 600, 705, 808, 869, 1078, 1113, 1167, 1650, 1833, 2106, 2136, 3312, 3996]. Fish [1747]. Five [704, 1419, 2348, 2517, 2562, 3447, 3592, 3769, 4446]. Fix [950, 952, 1779, 2267, 2544, 2992]. Fixed [1838, 2270]. Fixes [216, 805, 854, 2091, 3061, 3381]. Fixing [2272, 3727]. Flame [3332]. Flaw [3, 376, 456, 687, 693, 770, 865, 1182, 1441, 2136, 2185, 2544, 2834, 2866, 3003, 3042, 3078, 3464, 3563, 3666, 3708, 4006, 4119, 4301, 4313, 4409]. Flawed [413, 993, 1699]. Flaws [763, 952, 1700, 1952, 2097, 2158, 2169, 2242, 2348, 2723, 2724, 2855, 2969, 3077, 3097, 3553]. Flexible [1305, 3430]. flies [3010]. Flood [141, 602, 2940, 3805, 4068]. flooded [741]. flooding [589]. Florida [4456]. flow [1380, 2750, 3902]. flown [2408]. flows [1069]. flu [2936, 2989]. fluid [4462]. flux [2839]. fly [749]. fm fls [570]. Focus [1017, 2566, 3117, 3151, 3915]. focused [2588]. foe [2698]. foil [1761]. fold [3634]. Following [3479]. font [1180]. Fool [2464]. Fool-proofing [2464]. Fooling [2056]. foothold [1618]. footprint [4021]. footprinting [2925]. footsteps [3479].

Force [4469]. forced [1424]. ForceField [701]. forceful [2183]. forces [314, 930, 1497, 2065]. fore [2605]. forecast [416]. Forecasting [3976]. Forensic [2886, 3115, 3174, 3334, 3636, 3639, 3746, 3955]. Forensics [1332, 2003, 2250, 2631, 2820, 2886, 3326, 3327, 3519, 3545, 3575, 3587, 3614, 3679]. forged [126]. Forgery [506, 1419]. Forget [1890]. forking [3572]. forks [2230]. form [679, 1033, 1408, 3899, 3909]. formats [2704]. formed [156, 357, 760]. Former [2241]. forms [1086, 3759]. fort [3838]. Forte [2088, 2100]. Fortezza [364, 377]. forthcoming [3665]. fortress [20, 55, 295, 345]. fortresses [3732]. Fortune [991]. Forum [1657, 3228]. forums [2547]. fought [265]. found [3, 25, 242, 408, 484, 684, 692, 693, 865, 1001, 1512, 1752, 3464, 4323]. foundation [2417, 3702]. Founder [1851, 1972, 2153]. Four [872, 3634, 3769, 4328]. Four-fold [3634]. Fowler [3873]. fpkg2swpkg [632]. fragile [3807]. fragment [1426]. Fragmentation [346, 1273, 1834]. Frame [198, 361, 502, 1186, 1236]. framebusting [3255]. Frametrapping [3255]. Framework [547, 561, 3690, 3823, 4023, 4067, 4227, 4446]. franca [4364]. Francisco [1960, 2835]. Francois [3210]. Fraud [53, 63, 89, 130, 218, 412, 428, 737, 745, 898, 916, 1143, 1221, 1300, 1504, 1522, 1554, 1631, 1798, 1800, 1946, 2385, 3522, 3848, 4035]. fraudster [753]. fraudsters [3145]. fraudulent [1643, 3058, 3339, 4063]. Fred [970, 2089]. Free [92, 505, 712, 772, 1164, 1220, 1226, 1248, 1278, 1288, 1547, 1691, 1696, 1712, 1852, 1903, 2110, 2303, 2456, 2878, 3219]. Free-Net [92]. FreeBSD [962, 963]. Freedom [261, 473, 530, 2070, 2641]. freely [888]. French [297, 696]. friend [2698]. Friendly [4082]. fries [1817]. Front [87, 1975]. Frontier [1226, 2001, 3319, 3937]. fronts [1365, 3063]. frustrate [845]. FTC [1592]. FTP [395, 907, 1027, 1057, 1136, 1151, 3067]. ftpd [863, 1268]. ftpd/rlogind [863]. FUD [3722]. fuels [2241]. Full [311, 968, 1267, 2185]. Fuller [819]. Fun [2671]. functional [4205]. functions [1035]. Fund [45]. fundamentally [2069]. Fundamentals [3804]. funding [3572]. funds [4398]. Further [218, 417, 581, 590, 594, 786, 825, 1188, 1357, 1359, 2374, 2890, 3263]. Future [104, 228, 375, 627, 918, 1210, 1308, 1985, 2003, 2100, 2234, 2707, 2996, 3025, 3165, 3502, 3505, 3621, 3745, 4093, 4249, 4287, 4428, 4438, 4472]. Fuzzing [2838, 2938].

G7 [210]. gab [2033]. gags [2296]. gain [1268, 4126]. Gaining [1618, 3950]. gains [810]. Galitsky [1972]. game [879, 1836, 1857, 3038, 3427, 3884]. gamer [2066]. games [265, 754]. Gang [1681, 2736, 3601]. gap [2653, 3856, 4296]. gaps [3785, 4029, 4427]. Gardner [3699]. Gartner [3240, 3589]. gatekeepers [3946]. gates [1866, 2377, 2430, 2514]. Gateway [564, 681, 896, 3105, 3290, 4080]. gatherers [2100]. Gathering [4390]. gathers [2375]. Gauntlet [915, 1752]. gazing [1543]. GCHQ [3675]. GDPR [3835, 3928, 3958, 4041, 4098, 4237]. GE [481]. Gear [2127]. gears [2616]. Gemalto [4171]. genealogy [2674]. General [871]. generate [1848]. Generation [136, 308, 1010, 1025, 1297, 2151, 2728, 2902, 3397, 3500, 3814, 4319]. Gentle [2019]. Geoffrey [3881]. Georgia [3380]. German [1570, 2255, 2308]. Germany [458, 1649, 1819, 1922, 2574]. Get [537, 1158, 1396, 1611, 1780, 1897, 2149, 2198, 2219, 2472, 2571, 2637, 3000, 3274, 3506, 4284]. Gets [92, 469, 519, 856, 1125, 1144, 1145, 1206, 1481, 1651, 1754, 3312, 3523]. Getting [397, 1241, 1540, 1620, 1955, 2063, 2264, 2284, 2380, 2518, 2521, 2530, 2590, 2673, 2813, 3456, 3694, 4078, 4151, 4451]. Gevers [3864]. Ghosts [2314]. giant [1921]. giants [290]. gift [2033]. Gigabit [1389]. Gigger [1867]. Giuliani [2357]. give [95, 804, 928]. Given [394, 1946]. gives [426, 1267, 1449, 1485]. Glance [2048]. GlimpseHTTP [884]. glitch [777]. glitches [1098]. Global [70, 138, 156, 507, 640, 742, 814, 1068, 1183, 1187, 1232, 1288, 1303, 1317, 1493, 1674, 1768, 1917, 2363, 3080, 3091, 3536, 4001, 4304]. GlobalSign [1689]. Glype [3395]. Gmail [2303, 2846]. GNS3 [3731]. GNU [3700]. go [230, 1670, 1814, 2551, 2633, 2844, 2853, 4124]. goals [2506]. Godfathers [2450]. Goes [244, 304, 1187, 1570, 1796, 1893, 2044, 2203, 2559, 2752, 3696, 3703, 3883]. Going [117, 1419, 1771, 2164, 2896, 4013]. gold [2206, 3655]. Golden [1862]. Goldfeder [3853]. Gone [2302, 2468]. Good [187, 286, 585, 648, 755, 1811, 2680, 2822, 2927, 3390, 3816, 3983]. Goodbye [3714, 4276]. Google [2296, 2303, 2778, 2846, 2912, 2934, 3014, 3023, 3064, 3207, 3495, 3679, 3941, 4119]. Googling [2206]. Gordon [3977]. govern [2896]. Governance [2885, 3363, 3373, 4176]. Government [64, 260, 289, 478, 583, 656, 810, 878, 892, 1082, 1286, 1391, 1806, 1827, 1936, 1962, 2037, 2268, 2306, 2343, 2636, 2754, 3194, 3641, 3758, 3820, 3860, 3879, 3992, 4017, 4044, 4085, 4207, 4371, 4398, 4444]. Governments [210, 459, 1403, 2039, 3207, 3262, 3743, 4421]. Goyvaerts [3383]. GPRS [2001]. grabbing [3883]. grade [3004, 4091]. Gradient [517]. Graffiti [1569]. Gragido [3976]. Graham [3700]. Graham-Cumming [3700]. grail [3777]. Grant [3751]. granted [1317]. Graphics [373]. graphs [4426]. Gray [3966]. great [95, 442]. greater [2306, 3506, 4427]. greatest [1348, 2014]. greynet [2626]. Grid [2087, 2538, 2554, 2931, 2946, 3417, 3498]. GridOne [3019]. grips [2521, 4151]. Ground [313, 658, 1577]. Group [156, 679, 794, 796, 867, 881, 984, 1270, 1496, 1667, 2053, 2096]. Groups [260, 1829, 2933, 3293, 4256, 4400]. grow [1232, 1878, 2480, 2922, 3697]. Growing [161, 353, 1569, 3313, 3915, 4087]. grown [2223]. grown-up [2223]. grows [40, 3552, 3851, 4221]. Growth [6, 70, 159, 412, 421, 450, 886, 3341, 3756, 4145, 4190]. GSM [1110, 1115, 2044, 2260, 3189]. guaranteed [336]. Guaranteeing [1054]. guarantees [1416, 3797]. Guard [1641, 1682]. Guardicore [4424]. Gudaitis [3305, 3988]. Guerilla [3864]. GUI [2411]. guidance [2767]. Guide [1103, 1691, 1988, 3243, 3266, 3414, 3519, 3606, 3629, 3824, 4390]. Guidelines [1130, 1335, 1355, 2029, 3161, 4085]. guilty [191, 242]. Gumblar [2934, 3035]. guru [1998, 2563].

H [362, 1619, 3844]. H-P [362]. Haaster [3864]. habits [3786]. Hack [734, 1310, 1626, 1814, 1831, 1856, 2249, 2274, 3022, 3053, 3181, 3653, 3871, 3956, 4260, 4312, 4457, 4464]. Hacked [1383, 1535, 1601, 1951, 1980, 2051, 2765, 3717, 4271, 4450, 4456]. Hacker [57, 88, 153, 170, 189, 190, 212, 262, 297, 619, 803, 1090, 1253, 1270, 1373, 1582, 1687, 1992, 2053, 2066, 2285, 2465, 2517, 2549, 3280, 3295, 3380, 3720, 3813, 3927, 4390]. Hacker-proof [619, 1090]. hacker-tracker [212]. Hackers [22, 106, 214, 237, 240, 299, 311, 324, 326, 433, 607, 656, 750, 786, 804, 857, 906, 914, 975, 1002, 1059, 1321, 1395, 1468, 1485, 1520, 1536, 1538, 1617, 1625, 1670, 1719, 1742, 1753, 1820, 1865, 1889, 1920, 1921, 1949, 2122, 2152, 2292, 2402, 3111, 3204, 3225, 3284, 3285, 3354, 3410, 3579, 3623, 3764, 3845, 3867, 3884, 3924, 3985, 3991, 4030, 4107, 4247, 4348]. Hackers/viruses [1468]. Hacking [393, 409, 582, 777, 870, 1041, 1129, 1667, 1839, 2012, 2206, 2588, 2862, 2912, 3039, 3082, 3119, 3164, 3277, 3305, 3384, 3441, 3466, 3476, 3477, 3619, 3646, 3675, 3717, 3730, 3778, 3912, 4035, 4126, 4129, 4238, 4389, 4447]. hacks [1470, 3360, 3380]. Hacktivism [3234, 3703]. Hacktivists [2079, 3293]. Hades [1898]. Hagen [3352, 3353]. half [3172]. Halifax [2468]. Hallem [3503]. halt [1424]. halves [1836, 1857]. hampered [613]. hand [4344]. hand-in-hand [4344]. Handbook [3295, 3720, 3813, 4008]. handheld [1498, 2584]. handhelds [1249]. handle [2, 4090]. handler [823]. Hands [11, 3092, 3817]. Hands-on [3092]. hang [2338]. happen [1326, 1920, 2436]. happening [1240]. happenings [1762]. Happens [1460]. Happy [1858]. hard [1850, 2382, 3643, 3676, 3896, 4043, 4410]. hardback [3444]. hardball [2490]. hardcore [3703]. harder [2050]. hardest [1077]. Hardware [636, 1501, 1891, 3896, 3981]. Hardware-based [1891, 3981]. Hardy [3772]. Harlan [3334]. Harmful [2110]. Harrington [3679]. Harris [643]. Harvey [4087]. Hathaway [2945]. hats [2845]. haul [3601, 4168]. Having [723]. Havoc [1766, 2755]. Hawkins [2084]. Haxor [1814]. Hazards [80]. HDD [3428]. HDLC [338]. HDLC/SDLC [338]. head [1158]. head$ [2193]. Headline [1420, 3883]. headline-grabbing [3883]. headlines [2543]. healing [65]. Health [10, 1280, 1575]. Healthcare [2958, 3439, 3748, 3817, 3960, 3979, 4074, 4290, 4345, 4398]. heart [2559, 4011]. Heartbleed [3563, 3572]. heat [3582]. heated [343, 3332]. heating [1036]. heavily [697, 929]. Heinemann [3998]. held [2712]. Helix [3503]. hell [3133]. Hellman [396, 888]. Hello [1863, 4276]. help [537, 1339, 1520, 2218, 2824, 3077, 3354, 4047, 4070]. Helping [2218]. Helps [78, 255, 855, 1637, 3925]. Henry [3783, 3784, 3804, 3814, 3823, 3845]. Herdict [2709]. Here [1530, 1618, 1928, 2510, 2913, 4450, 4465]. heterogeneous [3755]. Hewlett [1257, 1967]. hi [2308]. hi-jacker [2308]. Hidden [1893, 2985, 4452]. hide [1900, 3846, 4190]. Hiding [2862, 3933]. hierarchical [3257]. hierarchy [1743, 2420, 3049]. High [44, 59, 62, 72, 96, 667, 696, 733, 852, 1128, 1579, 2142, 2262, 2929, 3218, 3573]. high-end [696]. high-level [2929]. High-profile [3573]. High-Publicity [1579]. High-rise [44]. High-speed [72]. High-tech [59]. Higher [3142, 3218]. Highest [2226, 3854]. highjackers [1761]. highlights [1109, 3101, 3230, 3350, 4157]. highly [4224]. highway [119, 4418]. hijack [4167]. hijacker [889]. hijacking [183]. Hillegass [3771]. hindered [677, 1495]. hindrance [4070]. hire [433]. hires [480]. Hiring [3867]. Hiscox [4210]. History [34, 388, 3977, 4010]. Hit [750, 786, 1059, 1100, 1571, 1613, 1653, 1982, 2245, 2694, 2845, 3293, 3361, 3422, 3592, 3666, 3974, 4035, 4043, 4254, 4410]. Hit-List [1613]. hits [797, 1111, 1563, 1759, 3739, 3909, 4290]. hitting [2050]. HMRC [4086]. hoax [187, 276, 591]. Hoaxes [661, 1875]. hold [1566, 4365]. Holden [3922]. Holding [2566, 3805, 3838]. holds [228, 2835]. Hole [513, 751, 1633, 1666, 1752, 1872, 2447, 3061, 3082, 3422]. Holes [216, 315, 346, 369, 390, 407, 430, 452, 471, 492, 506, 527, 549, 566, 586, 602, 676, 1306, 1900, 2036, 2169, 2179, 2196, 2210, 2220, 2238, 2265, 2290, 2304, 2317, 2368]. holiday [1865]. holistic [2678, 3429]. Hollywood [1965]. Holmes [4009]. Holographic [789]. Holt [3988]. holy [3777]. Home [42, 1085, 1518, 1842, 2620, 2747, 2844, 4428]. Homeland [1783, 1966, 2191]. homepage [608]. homework [2140]. Honey [1957, 3010]. Honeynet [1742]. Honeynets [2142, 2316, 2639]. honeypot [1963, 2051, 2124, 2142, 3046, 4405]. Honeypots [2006, 2130, 2131]. Hong [1219]. Honig [3392]. Hooda [3209]. Hoodie [4401]. hook [2477]. hooked [2590]. hooks [386]. hope [2364, 2739]. hopefuls [4209]. horizon [2232]. horror [1733, 1870]. horse [738, 771, 1100, 1123]. Hosmer [3614].

Hospitals [3801, 3810]. Host [28, 1127, 1543, 3486]. Host-based [1127, 3486]. hostage [3875]. hostile [914, 3308]. Hosting [1835, 3059, 3586]. hosts [4415]. hours [3035]. House [2257]. How-to [3435]. Howling [1766]. HP [395, 436, 457, 517, 555, 574, 630, 936, 1027, 1120, 1172, 1198, 1340, 1615, 1907, 2267]. HP-UX [395, 555, 1027, 1120, 1172, 1198]. HSBC [2534]. HTML [832, 1844, 2002, 3045]. HTML-driven [832]. HTML5 [3658]. HTTP [1181, 2117]. HTTP-Based [2117]. httpd [694]. HTTPS [3743]. Huawei [4184]. Human [471, 2337, 4346]. Humans [448, 2734]. hunt [2121, 2255]. Hunter [2448, 3287]. Hunting [2441, 3948, 3983, 4227, 4390]. hurt [2273, 2287, 3724]. Hutchens [3668]. Huw [3252]. Hybrid [1941, 3045, 3168, 4355]. Hybris [1563]. hype [1733, 2739, 3240]. Hyperbole [1598, 2017]. hypervisor [2871]. hypervisor-based [2871]. hyping [3927].

IAAC [1802]. Iain [3976]. IAM [3621, 4226]. IBM [9, 40, 140, 196, 386, 470, 702, 724, 943, 1051, 1247, 1272, 1344, 1565, 1825, 1829, 1904, 2214, 2335, 2603, 2733, 4316, 4375, 4469]. IBM/ObserveIT [4316]. ICAF [1802]. ICE T [592]. ICL [251, 1126]. ICL/TimeStep [1126]. ICS [4416]. ICS-specific [4416]. ICSPA [3524]. ID [355, 1863, 2636, 2743, 2761, 2981]. IDC [489]. idea [585, 2822]. ideal [4264]. Identification [1830, 1906]. identifies [639, 1324]. identify [3071]. Identifying [3510, 3795]. identities [4430]. Identity [1644, 1958, 2335, 2337, 2465, 2466, 2506, 2580, 2629, 2630, 2794, 3144, 3321, 3438, 3532, 3662, 3733, 3742, 4380]. identity-as-a-service [4380]. IDN [2393]. IDS [1924, 2067, 2093, 2124, 2289, 2347, 2411, 2421, 4048]. IDSs [1848]. IE [1488, 2114, 2386, 2544, 3042]. IE5 [1283]. IE7 [2587]. IEEE [2940]. If [2140, 3691]. ifconfig [822]. Ignorance [1769, 2381, 4439]. II [55, 1832, 1857, 2009, 2128, 2131]. III [295]. IIoT [4236]. IIS [513, 1725, 1765]. IKE [1202]. Illegal [775, 2481]. illicit [3973]. ILOVEYOU [1453]. IM [2668]. Image [639, 1254, 1715, 1908]. images [25]. Imaging [3955]. IMAP [710]. Imminent [4110]. Impact [432, 1509, 1548, 2558, 2958, 3298, 3467, 3659, 3967, 4058, 4237]. Impending [3791]. implement [4421]. Implementation [877, 1182, 1956, 2818, 3105]. Implementing [2465, 2832]. implications [1744, 2130, 2444, 2706, 3437, 3455, 3847, 4404]. Importance [1677, 1678, 1891, 2250, 2966, 3512, 3670, 4285]. Important [1438, 2168, 3011, 3354, 3896]. impressive [886]. Improve [793, 1697, 1866, 1911, 2832, 3959]. Improved [46, 257, 766]. improves [1297]. Improving [491, 1092, 3378]. IMSI [3912]. In-session [2914]. in-the-wild [3046]. Inc. [373]. incarnations [3858]. Incident [566, 586, 1294, 2157, 2328, 2462, 2542, 2608, 2642, 2748, 2772, 2873, 2966, 3029, 3545, 4099]. Incidents [1579, 1778, 2626, 4022]. incorporated [1086]. incorporates [539]. Increase [393, 610, 638, 1518, 2935, 2981, 3121, 3342, 3371, 3552, 3634, 4334, 4359]. Increased [16, 260, 972, 1420, 1430, 1726, 3459]. increases [196, 2637]. increasing [3076, 4309]. increasingly [2968]. independent [600]. Index [3534, 4469]. India [91, 2437, 3072]. Indian [260, 1513, 1590]. Indianapolis [1482]. indicators [4295]. indicted [2692]. indictments [4129]. indicts [3582]. Industrial [111, 874, 3561, 3619, 3678, 3916, 4257, 4264]. Industry [367, 725, 1520, 1536, 1575, 1651, 1806, 2008, 2010, 2022, 2156, 3501, 3502, 3745, 3953, 4216, 4273, 4340]. inevitable [2903]. infected [1869, 2811]. infecting [557]. infection [771, 2980]. infections [2959, 3121, 3302]. infested [3089]. infiltrating [4162]. infiltration [2716]. infinity [2406]. InfinityBlack [4349]. Info [205]. Info-Terrorism [205]. Infohighway [245, 278]. Information [119, 176, 210, 368, 455, 465, 473, 502, 530, 564, 813, 838, 859, 880, 1020, 1039, 1083, 1192, 1196, 1259, 1277, 1328, 1346, 1369, 1655, 1858, 1945, 1947, 2084, 2133, 2152, 2339, 2708, 2748, 2770, 2842, 2972, 3069, 3084, 3163, 3190, 3347, 3406, 3460, 3499, 3520, 3537, 3615, 3699, 3702, 3766, 3949, 4008, 4203]. infosec [1175, 1718, 1790, 1813, 2320, 2426, 2563, 2616, 3804]. infosecurity [1892, 2092, 3505]. Infowar [1175]. infrastracture [8]. Infrastructure [726, 773, 1200, 1261, 1576, 1774, 2061, 2499, 2977, 3080, 3342, 3523, 3924, 4012, 4013, 4038, 4155, 4204, 4386, 4405, 4449]. Infrastructures [1417]. infringements [2, 1194]. initiative [121, 725, 1429, 3992]. initiatives [4207]. Injection [2002, 2527, 2536, 2760]. injections [2738]. inky [2756]. INN [806]. Innovation [3936, 4200, 4344]. inoculation [1517]. inquiry [458]. Ins [1789]. insecure [511, 2144, 2270, 3364, 3596]. insecurities [973, 3474]. insecurity [215, 347, 489, 3254]. Insider [2784, 3633, 4098, 4221, 4316, 4341, 4367, 4391]. Insight [2076, 2086]. insights [1710]. Instant [1871, 1934, 2233, 2416, 2577, 2644, 3435, 3486]. Instilling [4079]. institute [3368, 4185]. institutions [1704, 3919]. insurance [1496]. insurers [735]. Integralis [339, 462, 987]. integrate [386]. Integrated [485, 743, 1031, 1127, 1184, 2276, 2413, 2772]. integration [1003, 1030, 1160, 1302, 1385, 2042, 2111, 2446]. Integrity [125, 336, 705, 2975, 3546, 4067, 4406]. Intel [1850, 3100, 4323, 4384]. intellectual [268, 294, 3914]. Intelligence [325, 1640, 1882, 2100, 2808, 3681, 3730, 3735, 3803, 3925, 3937, 3980, 4070, 4146, 4239, 4390, 4469]. intelligence-driven [3735]. Intelligent [1129, 1456, 1937, 3609, 3673, 4069]. intent [3158]. interaction [3046]. Intercept [3977]. interception [1545, 2679]. Interchange [318]. interest [219]. interface [2795]. interim [4017]. Internal [24, 424, 686, 1896, 2430, 2491, 2613]. Internals [3710]. International [246, 799, 1514, 1563, 1691, 3444, 3524]. Internationally [1639]. Internationally-Minded [1639]. Internet [36, 37, 39, 62, 87, 90, 140, 157, 195, 196, 200, 213, 219, 221, 222, 224, 226, 243, 244, 246, 252, 257, 282, 305, 331, 334, 339, 352, 378, 379, 384, 385, 400, 423, 444, 447, 457, 460, 465, 481, 486, 501, 515–517, 537, 538, 542, 573, 574, 610, 612, 613, 618, 620, 621, 635, 637, 638, 640, 641, 646, 659, 681, 682, 693, 712, 719, 738– 740, 761, 783, 793, 825, 829, 833, 847, 848, 853, 910, 916, 1004, 1104, 1617, 3323, 4120]. Internet [19, 54, 78, 100–102, 122, 142, 263, 264, 291, 293, 313, 315, 346, 347, 369, 390, 407, 409, 428, 430, 452, 471, 492, 506, 527, 549, 550, 564, 566, 586, 601, 602, 623, 627, 649, 690, 918, 973, 1015, 1016, 1045, 1050, 1071, 1079, 1101, 1143, 1221, 1250, 1321, 1328, 1343, 1346, 1362, 1363, 1376, 1401, 1450, 1469, 1483, 1493, 1539, 1641, 1691, 1717, 1744, 1798, 1861, 1983, 2070, 2230, 2281, 2306, 2441, 2460, 2676, 3032, 3080, 3203, 3432, 3456, 3484, 3531, 3599, 3602, 3745, 3888, 3894, 4027, 4141, 4154, 4338]. Internet [2, 26, 31, 55, 112, 145, 163, 261, 449, 453, 493, 508, 748, 775, 799, 824, 837, 939, 1024, 1042, 1093, 1132, 1149, 1239, 1314, 1400, 1532, 1559, 1576, 1656, 1763, 2054, 2336, 2364, 3684, 3869, 3946, 3969, 3970, 4002, 4040, 4126, 4173]. Internet-based [833]. Internet-borne [610]. Internet-enabled [501]. Internet/intranet [517, 618, 621, 681, 918]. InterNetNews [698]. Interpreters [510]. INTERSHOP [1031]. Interview [2253, 3480, 3492, 3503, 3535]. intranet [517, 618, 621, 681, 686, 819, 918, 1003]. Intranets [680, 740]. introduced [1501]. introduces [699, 970]. Introduction [125, 2119, 2247, 3497, 3508, 3520, 4009]. intruder [1215]. intruders [1321, 1872]. Intrusion [902, 920, 1032, 1127, 1256, 1345, 1482, 1523, 1789, 1833, 1834, 1955, 1987, 2017, 2073, 2181, 2243, 2246, 2300, 2717, 2850, 2861, 2951, 2952, 3486, 4355]. intrusions [2234]. Invasion [608, 2377, 2747]. invent [2460]. inventory [1075]. investigates [1592, 2479]. Investigating [3039, 3127, 3531, 4471]. Investigation [1470, 2909, 3115, 3174, 3639]. Investigations [3522, 3746, 3788, 4378]. investigators [777]. Investigatory [3791]. Investing [195]. Investment [428, 519, 1634, 2484, 2999]. invests [7, 71, 1825]. involvement [3070]. ioctls [822]. iOS [236, 1046, 1080, 1119, 1199, 3794, 3823]. IoT [3705, 3978, 3986, 4021, 4044, 4057, 4091, 4123, 4142, 4145, 4150, 4151, 4157, 4161– 4164, 4171, 4192, 4201, 4248, 4344, 4387, 4412]. IP [128, 139, 183, 228, 373, 406, 506, 563, 589, 769, 852, 1005, 1182, 1273, 1419, 1426, 1501, 1578, 1647, 2209, 2321, 2379, 2486, 2600, 3009, 3028, 3502, 3647, 3993, 4273]. IP-based [3502]. IPCS [1443]. IpFrags [1834]. iPod [2529, 2728]. IPS [2181, 2289, 3004]. IPSEC [1202, 1290, 2749]. IPSec/IKE [1202]. IPV6 [1009, 3177, 3209, 3235, 3279, 3352, 3353, 3455]. IPX [769]. Iran [3459, 4230]. Iranians [4278]. IRC [105, 1100]. Ireland [808, 1542, 2085]. Irina [591]. IRIX [512, 653, 758, 823, 842, 882, 999, 1012, 1118]. IronBridge [545]. IronWall [545]. irrational [3887]. ISA [1473]. ISBN [3184, 3197, 3209, 3210, 3221, 3222, 3232, 3243, 3252, 3253, 3266, 3276, 3277, 3287, 3295, 3304, 3305, 3326, 3327, 3344, 3345, 3373, 3374, 3383, 3384, 3393, 3413, 3414, 3425, 3435, 3444, 3454, 3466, 3476, 3477, 3509, 3519, 3555, 3614, 3657, 3658, 3668, 3678, 3679, 3689, 3699, 3700, 3710, 3720, 3741, 3742, 3751, 3752, 3761, 3771, 3772, 3783, 3784, 3794, 3813, 3814, 3823, 3824, 3833, 3834, 3844, 3845, 3863, 3864, 3873, 3881, 3891, 3892, 3901, 3911, 3922, 3955, 3956, 3966, 3976, 3977, 3988, 3998, 4008, 4009, 4065]. ISBN [149, 1178, 1482, 3334, 3352, 3353, 3363, 3392, 3402, 3403, 3730, 3731, 3803, 3804, 3933]. ISDN [42, 256]. Islamic [1753]. ISN [1297]. isn’t [2472, 3727, 4439]. ISO [2542]. iSolve [1504]. ISP [600, 760, 1342, 1897]. ISP-independent [600]. ISPs [750, 1335, 1649, 2183, 3888]. ISS [1473, 2024, 2447, 2524, 2603]. ISSE [2605]. issue [689, 1091, 3733]. issued [497, 687, 894, 4386]. Issues [21, 221, 225, 981, 995, 1438, 1696, 2140, 2192, 2209, 2740, 2767, 2866, 3042, 3323, 3385, 3407, 3817, 4129, 4194, 4196, 4268]. issuing [1643]. ISTE [3742, 3834]. IT/OT [4102, 4113]. Italian [1744, 3974]. Italy [193]. ither [1619]. ITIL [2642]. ITSEC [479]. itself [3599]. ITT [468]. IV [345]. IVES [137]. Ivy [2390].

J [3454]. Jabadoo [866]. Jabbusch [3305]. Jack [3305]. jacker [2308]. Jail [2371]. Jake [453]. Jakobsson [3901]. James [1763, 3519]. jamming [3357]. Jan [3383]. Japan [657]. Japanese [299, 617, 1270, 1384, 2392]. Jason [3731]. Java [475, 692, 731, 768, 770, 804, 809, 849, 877, 903, 997, 1018, 1068, 1390, 1487, 3367, 3422, 3568]. Java-based [1068, 1390]. JavaScript [398, 1326, 2759, 3658, 4065]. Javier [3834]. JCP [853]. Jeff [1617]. Jeffreys [3945]. Jeffreys-Jones [3945]. Jelle [3864]. Jennifer [3305]. jeopardized [945]. Jeremy [3276]. Jessop [2082]. Jet [1284]. jihad [3396]. Jim [3243]. Joanap [4167]. job [188, 1964, 2343, 2631, 3067, 4112]. Joe [4065]. Joel [3678]. John [149, 1178, 2082, 2523, 3326, 3327, 3524, 3700, 3933, 3976, 4008]. join [314, 930, 1303, 1497, 1802]. joined [3877]. joined-up [3877]. Joining [2941, 4176]. joins [715, 788, 809, 830]. Joint [481, 558, 1011]. Jon [3492]. Jones [3945]. Joseph [3853]. Josh [3476, 3477]. Joshua [3414, 3824, 3922]. Journalists [582, 3460]. JPEGS [1940]. jScript [1045]. Judge [262, 2481]. judges [67]. July [4099]. jump [4010]. Juniper [3780]. Junk [1636]. Junkie [4]. Jurisdiction [1093, 1532]. Just [1239, 1599, 1766, 2006, 2152, 2385, 2584, 3364, 4020, 4201]. Justice [1261, 2652]. Justin [3668, 3689].

Kali [3604]. Kaminsky [2848]. Kane [642]. Karamanian [3210]. karma [604]. Karyotis [3844]. Kaspersky [2824, 3633, 3696, 4257]. Kaufman [3197, 3844]. Kaufmann [3402, 4008]. KCMS [554]. Kearns [3243]. Keep [78, 149, 284, 1716, 3588, 4332, 4368]. Keeping [120, 2396, 2542, 2611, 3670, 3711, 3895, 4365, 4366]. keeps [712, 1560, 1978]. Keith [3657]. Kennedy [3243]. kept [414, 3501, 3618]. Kerberos [142, 469, 1444]. Kernighan [3911]. Kevin [185, 819, 2084, 3413, 3845]. Kevvie [3873]. Key [104, 413, 414, 478, 579, 593, 595, 665, 695, 797, 814, 909, 929, 931, 938, 993, 1014, 1022, 1035, 1364, 1398, 1417, 1507, 1544, 2632, 2699, 3057, 3469, 3693, 3766, 3769, 3789, 3836, 3960, 4164, 4319, 4429]. Keyboard [3933]. keychain [2830]. Keyloggers [2687, 3126, 3226, 4320]. keys [387, 1817, 3126, 4430]. keystroke [2147]. Khouzani [3844]. kid [4014]. Kido [2913]. Kids [1637, 1688]. Kilger [3988]. Kill [932, 2162, 3348]. killed [1537]. King [2085]. kingdom [1663]. kiss [3714]. kit [1205, 2795]. kits [532]. Klein [3287]. Klez [1919]. KMPG [3482]. Knapp [3678]. knife [2862]. knocked [3228]. knocking [4461]. Know [1617, 1861, 1950, 2043, 2127, 2273, 3618, 3827, 3945, 4165]. Knowing [2914, 4103]. known [388, 1374, 1679, 2863]. knows [2086]. Kong [1219]. Korea [214, 2957, 3653, 4025, 4167, 4467]. Korean [4158, 4351]. Korn [652]. Kournikova [1758]. KPMG [480, 4087]. Kristin [3772]. Kumar [3730]. Kyas [799]. Kyrgyzstan [2899].

L0phtCrack [1668, 2965]. Lab

[25, 1239, 1506, 2982, 3435]. Labs [951]. Lack [1208, 1358, 2042, 2111, 3602]. lacking [1512]. LaMacchia [114]. LAN [17, 28, 46, 108, 110, 123, 132, 139, 371, 449, 521, 577, 1275, 1339, 1893, 2043, 2526]. LAN/WAN [108]. land [1570]. Landscape [1979, 2418, 3546, 3568, 3736, 3915, 4293, 4472]. Langill [3678]. language [398, 628, 730, 1233]. LANs [14, 227, 1888, 2069, 2609]. Lantern [1817, 1843]. lapses [366]. Laptop [2193, 2868, 3360]. Large [783, 858, 1476, 1668, 2066, 2940, 4301]. large-scale [2940]. largest [2492]. Larry [3197]. Laser [123]. last [1514, 1963, 2384, 3319, 3920]. Latches [1327]. late [2023]. latency [3125]. Latest [128, 307, 685, 805, 870, 1901, 2014, 2408, 3381, 3687, 3708, 4254]. Latin [1200]. launch [93, 2933, 3612]. Launched [49, 700, 742, 1083, 1250, 1389, 1411, 1693, 1712, 1721, 2142, 3013]. launches [48, 544, 600, 702, 853, 1427, 1451, 2112, 2202, 2703, 2733, 2789, 2910, 3014, 3260, 3879, 4044]. Laundry [1579]. Laurent [3742]. Law [232, 269, 1586, 2146, 2556, 3444]. lawful [2679]. laws [795, 1173, 1309, 1587, 1647, 3460, 3861]. Lawsuit [1636, 1722]. Lawyers [1395]. lax [17, 266, 1015, 2684]. Layer [371, 2488, 2771, 2940, 3471]. Layered [1574, 3306, 4152]. layoffs [1641]. Layton [3803]. ld.so [801]. lead [1441, 3848, 4466]. leader [1104]. Leaders [725, 1944, 2435]. leading [3986]. leads [2255, 3539, 3572, 4119, 4145]. League [2390]. leak [85, 1952, 2879, 3643, 3952, 4334]. leaked [1367, 1907]. Leaks [3085, 3158, 3302, 3675, 3960, 4384]. leap [2332]. Learn [1077, 1787, 2991]. Learned [688, 2049]. Learning [3176, 3191, 3291, 3658, 4135, 4203, 4355, 4474]. learnt [1803]. least [3336, 3691]. leave [1015, 1084, 4283]. leaves [153, 2340, 3001, 3181, 3563, 3856, 4218]. led [3576]. Lee [3363]. Left [511, 1559, 1634, 4189]. legacy [2181]. Legal [167, 1132, 1149, 1585, 1762, 1794]. Legally [1084, 1437]. legislates [1625]. legislation [1021, 1271, 1514, 3712]. legislators [795]. legislature [10]. length [1014]. Lenin [755]. Less [1679, 1959, 3707]. lesser [1679]. lesser-known [1679]. lesson [330]. Lessons [688, 1077, 2049, 4135, 4405]. let [2285, 4020]. Lethal [2329]. Letter [2141]. letters [126]. level [138, 767, 850, 1279, 2226, 2366, 2929, 3446]. levels [2498, 4181]. Levithan [3383]. Lewis [2523]. Liability [1194]. liable [1402]. liberation [87]. Liberties [1511]. Liberty [2335]. Librarians [542]. libraries [730]. libXt [729]. licence [396, 790]. license [1317]. Licensed [1786]. lid [2930]. Life [1863, 2001, 2451, 3035, 3420, 3817]. lifecycle [4200]. lifts [1269]. light [1744, 2582, 4452]. lightning [549]. Lights [1534, 3885]. Like [438, 1863, 2271, 3403]. likely [3641]. limelight [1900]. Limited [807, 1617, 2252]. limits [1209, 1292, 2598]. line [2229, 2451, 3570]. lingering [2569]. lingua [4364]. Link [9, 180, 336, 826, 1088, 1624, 3330, 4392]. link-state [3330]. Linked [1194, 1876, 2480]. Linking [1313, 1675]. links [44, 194, 338, 483, 1360]. Linux [684, 1154, 1182, 1490, 1653, 1680, 1755, 1777, 1869, 1922, 1935, 1995, 1998, 2299, 2300, 3604, 3666, 3668, 4337]. Lion [1652]. Liska [3881]. List [1613, 1712, 2319, 2583, 2879]. listed [4465]. LISTSERV [1464]. literature [1792]. little [427, 1094]. Living [552, 3513, 4199]. Lloyds [2477]. LLP [781]. loadable [1088]. loadmodule [349]. lobby [1566]. Local [147, 895, 1379, 1489, 3801]. location [2175]. Lock [414, 1090, 1841, 3469]. lockdown [2624]. locked [448]. Locking [3338, 3816, 4473]. LOCKOUT [842]. Log [2176, 2462, 2769, 2966, 3127, 3413, 4418]. logger [2147]. Logging [2015, 3413]. Logica [1105]. logical [1833]. logically [1480]. login [842, 1080]. login/scheme [842]. LOGINOUT [1028]. Logjam [3708]. London [1888, 2481, 3977]. Long [400, 942, 3666, 3708]. long-term [3666, 3708]. longer [387, 1736, 3067, 3095, 3406]. longstanding [3858]. Look [368, 938, 1196, 1508, 2100, 2453, 2543, 2721, 3281, 3389]. Looking [1928, 2913, 3479, 4224, 4472]. looks [647, 4267]. Looming [1656, 3561]. loose [1372]. Lords [2712]. lose [340, 374]. loses [396]. Losing [1715, 3841, 4346]. Loss [2766, 3015, 3138, 3178, 3712, 3833, 4270]. Losses [1420, 2754, 3112, 3227, 4110]. lost [1430, 3456]. lots [2110]. lotteries [1314]. Lotus [31, 356, 429, 721, 952]. louder [2530]. Love [1435, 3191, 4124]. Low [634, 766, 1204, 2366, 2432, 3046, 3305, 3646]. Low-cost [1204]. low-interaction [3046]. Low-level [2366]. Lowdown [1701]. Lowther [3305]. Ltd [1327, 1368]. Lucent [785, 2868]. Ludloff [3344, 3345]. Lumeta [2330]. lunch [2303]. lurch [2340]. Lurid [3260]. LYDownload.c [802]. Lynx [802]. Lyons [3524].

M [1567, 1568, 1607, 3844]. M-commerce [1567, 1607]. M-spam [1568]. Mac [1715, 2419, 2516, 2550, 3026, 3312]. machine [692, 1298, 2314, 2854, 4203, 4355]. machines [2928, 4259]. Macmillan [1482]. Macro [439, 488, 497, 596, 926, 1135, 3947]. Macs [3131]. made [144, 427, 1797]. Madison [3744]. maelstrom [369]. mafia [2836].

Mafiaboy [1754]. magic [384, 1817, 1843]. Magistrate [587]. magnet [1688]. Magnetics [1368]. Mail [28, 79, 120, 126, 148, 149, 152, 232, 288, 310, 394, 402, 514, 528, 567, 581, 720, 736, 741, 834, 846, 851, 865, 874, 932, 945, 949, 1099, 1157, 1164, 1177, 1388, 1416, 1457, 1499, 1524, 1545, 1636, 1734, 4358]. mailers [1226]. MailSafe [1368]. mainframe [139]. mainframes [9]. mainstream [3883]. maintain [3748]. Maintaining [2570, 4330]. Major [194, 616, 1951, 2970, 3118, 3367, 3624, 3643, 3810, 3888, 3974, 4063, 4232, 4315, 4384, 4466]. Makan [3657]. Make [356, 623, 755, 1205, 1687, 2073, 3004, 3393, 3559, 3573, 3700, 3883, 3904, 4219, 4236]. Make/O’Reilly [3393]. Makes [193, 771, 825, 1090, 1855, 1945, 2292, 2551, 2878, 3130, 3169, 3188, 3324, 3348, 3971]. Making [334, 2084, 2938, 3416, 3482, 3628, 3854, 4066, 4286]. Malcolm [3956]. Malformed [1229]. Maliciou [2088]. Malicious [328, 1396, 1787, 1853, 2298, 2765, 3053, 3158, 3200, 3947, 4196, 4247, 4329]. Malin [3988]. Malvertising [3739]. Malware [1779, 2217, 2325, 2327, 2471, 2487, 2619, 2666, 2674, 2707, 2713, 2763, 2785, 2798, 2808, 2892, 2933, 2981, 2993, 3000, 3015, 3052, 3057, 3089, 3108, 3213, 3216, 3312, 3341, 3387, 3392, 3479, 3495, 3519, 3538, 3539, 3570, 3696, 3713, 3787, 3844, 3909, 3935, 3947, 4068, 4212, 4224, 4328]. Malware-infested [3089]. Man [2147, 3058, 3339]. Man-in-the-Browser [3058]. Man-in-the-Middle [3058, 3339]. manage [3627, 3942, 4463]. Managed [516, 1409, 1503, 1911, 1988, 2776, 3365, 3490]. Management [121, 727, 754, 877, 931, 937, 1065, 1068, 1075, 1930, 1958, 2129, 2176, 2268, 2276, 2337, 2409, 2462, 2465, 2518, 2542, 2580, 2616, 2629, 2630, 2632, 2642, 2678, 2772, 2773, 2786, 2794, 2807, 2819, 2882, 2916, 3124, 3144, 3253, 3318, 3413, 3500, 3501, 3545, 3742, 3798, 3804, 3815, 3902, 3917, 3925, 3969, 4092, 4285, 4365, 4379, 4418, 4436, 4451]. Manager [15, 699, 700, 899, 2411, 2972, 4228]. managers [966, 1561, 1850, 2626]. Managing [648, 669, 689, 705, 727, 754, 774, 798, 817, 836, 860, 879, 902, 921, 937, 955, 980, 996, 1011, 1023, 1040, 1053, 1073, 1091, 1114, 1133, 1148, 1168, 1176, 1189, 1209, 1224, 1240, 1259, 1277, 1292, 1311, 1380, 1678, 2117, 2762, 2797, 2805, 3037, 3069, 3134, 3408, 3662, 3734, 4163, 4392]. Manual [3444]. manufacturer [11]. Manufacturing [4233]. many [2, 229, 1848, 3707, 4150, 4179, 4260, 4465]. mapping [4057]. March [2918]. Marconi [1607]. Marcus [3295]. mark [548]. market [6, 69, 117, 133, 333, 499, 617, 886, 1211, 1232, 1289, 1384, 1428, 1630, 1719, 2092, 2630, 2703, 3494, 4349]. Marketing [558, 1195, 1598]. marketplace [1158]. marketplaces [3973]. Markets [3831, 4362]. Markus [3901]. marriage [4124]. Marshal [2877]. Marsicano [3772]. Martijn [3864]. Mary [3344, 3345]. Maryline [3742]. mask [3738]. Masking [3631]. mass [117]. masses [2311, 2632, 2838]. Massive [421, 1644, 2980, 3361, 3653, 4168]. MasterCard [291, 461]. matching [2498]. material [309]. Math [3752]. mathematical [2406]. Mathematics [3922]. Mati [3243]. Matter [1727, 2949, 3178, 3480, 3651, 4213]. matters [3577, 4238]. maturity [4352]. maverick [3978]. Max [3988]. Maximize [1064]. Maximizing [2661]. may [444, 477, 1173, 1326, 1564, 1704, 2425, 3111, 3204, 3505, 3861]. Maybe [1621, 3360]. mayor [2357]. MBR [2782]. McAfee [426, 550, 831, 871, 1594, 2548, 2857, 2918, 3100, 3452]. McCain [1188]. McDonalds [1627]. McFarland [3209]. MCI [200]. McNeally [2514]. McVeigh [981]. MD5 [2890, 2907]. mean [1679, 3874, 4394]. meaning [3148]. means [1331, 1618, 3141, 3835]. Meanwhile [1618]. measures [318, 1400, 2867, 4361].

Measuring [2998, 4125]. mechanics [2679, 2718]. mechanism [2940]. mechanisms [508, 3298]. Media [706, 2412, 3784, 3826]. mediation [1190]. Medical [1673, 2690, 3548, 4268]. medium [706, 3849]. medium-size [3849]. meets [1285]. mega [572, 3017]. Mega-D [3017]. megatrends [3467]. Melissa [1170, 1171]. meltdown [4001]. meltdowns [3083]. melting [1372]. members [2879, 3582]. memory [4452]. menace [2970, 3335, 3851, 4063, 4440]. mentality [20, 55, 295, 345]. merchants [1300]. mere [3154]. Merge [479, 572, 928, 2096, 2261, 2877]. Mergent [247]. Merkle [396]. Merlin [384]. mesh [2940]. Message [287, 315, 706, 3274]. Messages [149, 153, 741, 835, 932, 2826, 3052]. messaging [69, 169, 598, 1285, 1871, 1934, 2409, 2416, 2577, 2644, 3648]. messenger [1055]. Meta [1010, 1025]. Meta-firewall [1010, 1025]. metadata [2697]. metal [4177]. Metamail [757]. Metasploit [3092, 3243, 3509]. Meteors [1561]. meters [2950]. method [3019]. methodology [472, 726, 3176]. Methods [975, 4077]. Metric [1745, 1892]. metrics [2818, 3627]. Mexicans [1540]. Michael [881, 3392, 3444, 3679, 3833]. Mick [3480]. Micro [772, 1090, 2271, 4110, 4187, 4362, 4428]. Micro-miniature [1090]. Micro-segmentation [4187, 4428]. microchip [1314]. MicroLan [715]. Micron [1722]. micropayment [716]. microscope [3548]. Microsoft [75, 93, 158, 249, 288, 331, 356, 465, 515, 547, 550, 575, 626, 736, 784, 788, 790, 855, 942, 950, 960, 1044, 1153, 1159, 1174, 1245, 1283, 1285, 1297, 1298, 1320, 1340, 1341, 1344, 1351, 1473, 1488, 1535, 1596, 1666, 1671, 1687, 1775, 1841, 1901, 1904, 2034, 2039, 2125, 2160, 2169, 2213, 2242, 2255, 2283, 2340, 2374, 2384, 2412, 2469, 2479, 2490, 2563, 2587, 2605, 2606, 2667, 2756, 2777, 2866, 2876, 2895, 3184, 3185, 3304, 3411, 3422, 3425, 3564, 4325, 4413]. Microsoft-Mail [288]. Mid [4293]. Mid-Year [4293]. middle [313, 2924, 3058, 3285, 3339, 3361, 4278]. Midian [2086]. Midyear [4110]. Migrating [1016, 2247]. migration [3177, 4343]. migrations [3723]. Mike [3384, 3466]. milestone [3958]. Military [429, 850, 1041, 1223, 1401, 1761, 3207, 3423, 4433]. Military-level [850]. Millennium [1581]. Miller [3853]. million [374, 745, 1192, 1252, 1583, 1609, 2122, 2230, 3495, 4374]. Millions [3563, 3831, 4218, 4242]. millisecond [1168]. MIME [673, 834, 1055]. Mimecast [4222]. MIMEsweeper [720]. mimic [2836]. mind [3093, 3124, 4368]. Minded [1639]. minds [3939]. mine [39]. miniature [1090]. minimise [3967, 4058]. minimising [3567]. Ministry [843]. minor [2097]. Minute [2031]. minutes [1737, 3238]. mirror [1710]. mirroring [1685]. MIS [1602]. Misconceptions [52, 2017]. Misenar [3414, 3824]. Misimpressions [1914]. miss [3378]. misses [953]. Mission [1471, 2452, 4454]. Mission-critical [1471]. Mistakes [2091]. misunderstood [2410]. Misuse [1475, 1785, 2280, 3826, 4372]. mitigate [3659, 4309, 4426]. Mitigating [3257, 3407, 4032]. Mitigation [2101, 3278, 3595, 4078, 4427]. mitigations [3387]. Mitnick [154, 185]. mixed [158, 769, 2917]. mmap [962]. Mnemonics [2180]. Mob [1481]. Mobile [6, 43, 218, 333, 1110, 1319, 1359, 1360, 1484, 1498, 1650, 1699, 1721, 1787, 1796, 1973, 2216, 2288, 2355, 2398, 2568, 2589, 2619, 2707, 2710, 3037, 3114, 3188, 3202, 3237, 3241, 3298, 3308, 3335, 3364, 3371, 3408, 3448, 3474, 3495, 3503, 3509, 3544, 3552, 3553, 3638, 3640, 3686, 3696, 3720, 3785, 3823, 3833, 3912, 4037, 4449]. mobility [1396, 3439]. mode [80, 606]. model [793, 2069, 2563, 2709, 2952, 2982, 3291].

Modelling [2895, 4049]. Models [3844]. modem [707, 811]. modems [418, 1238, 1550]. Modern [2796, 3298, 3546, 3690, 3844, 4089, 4365, 4424]. module [732]. modules [2821]. Moment [1808]. momentum [2375]. Mondex [716]. monetisation [4214]. Money [1584, 2551, 3095, 3696]. Monitor [1550, 1649, 1726, 1983, 3236]. monitored [237]. Monitoring [136, 148, 893, 1113, 1167, 1251, 1349, 1529, 1613, 1936, 2054, 2360, 3289, 3386, 3546, 3626, 3672, 3934, 4342]. monitors [2523]. monopolies [99, 160]. Monroe [2919]. Monster [4254]. monster.com [2720]. Month [1831, 1856, 3808]. months [1754, 2842]. Moon [1766]. Moraes [3221, 3222]. Morgan [3197, 3402, 3844, 4008, 4065]. Moroccan [2458]. Moskowitz [3998]. Most [312, 767, 1420, 1955, 2592, 2765, 3241, 3308, 3364, 3394, 3895, 3899, 3909, 4379, 4403, 4451]. most-attacked [3241]. motion [2601, 3490]. motivates [57]. motivations [3068]. Motorola [251, 253, 255, 519]. mountd [1081]. mounting [3237]. mouse [535, 3038]. Move [116, 1981, 2239, 2537, 3735, 4453]. movement [1647, 3323]. Moves [193, 356, 450, 952, 1037, 2171, 2523, 2531, 2557, 3801]. movies [1604]. Moving [2045, 2986, 3418, 3576, 4368]. Mozilla [2811]. MP [2436]. MPack [2695]. MPs [2280]. MS [1824, 2071, 2136, 2344]. MS-SQL [2071]. MSBlaster [2164]. MSPs [4463]. Much [1893]. muffin [2612]. Mullen [3304]. Multi [599, 713, 744, 745, 831, 1220, 3059, 3306]. Multi-discipline [599]. multi-layered [3306]. multi-million [745]. multi-platform [713, 1220]. multi-tenanted [3059]. Multi-tier [744, 831]. Multifactor [4447]. multifunction [2521]. multimedia [73, 75, 250, 2581]. multimodal [2633]. multinational [2797]. Multiple [983, 1064]. multiplies [3705]. multiply [3474]. Muninder [3209]. munition [231]. murder [777]. Music [1520, 1604]. must [3417, 3915]. muzzles [1539]. MVA [1911]. MVS [125]. mwcollectd [3046]. My [1663, 1863, 2302]. MyDoom [2219, 2235, 2296]. MyNet [546]. MySQL [3658]. mystery [1147]. Myth [2191, 2204, 2299].

NAC [2476]. Nachreiner [3535]. NAFTA [1022]. NAI [1449, 1752]. Naked [1632]. Name [325, 576, 987, 1055, 2648, 4115]. names [942]. Naming [104]. Nance [3956]. Napster [1618, 1772]. Narayanan [3853]. NAS [2248]. NASA [1980]. Nash [4087]. nasty [2695, 3762]. Natas [175]. Nate [3771]. Nation [2030, 3989, 3990, 3999, 4000, 4138, 4331, 4358, 4371, 4421, 4440]. Nation-state [3989, 3999, 4000, 4138, 4331, 4358, 4371, 4421, 4440]. National [1357, 1469, 1506, 1804, 1830, 1993, 2417, 3841, 3924, 3963, 4013, 4038, 4132]. Nations [4312]. Nationwide [91]. Native [1067, 4415]. NATO [1285, 3461, 3611]. Natural [730]. navigation [3683]. Navigator [609, 778, 805]. nCipher [930, 2153]. NCSA [584, 701]. NCSC [3996, 4245, 4266, 4281]. NDR [3048]. NDS [1174, 1912]. near [3378]. near-miss [3378]. Nearly [3172, 3553, 4006]. necessarily [219]. necessary [4092]. Necurs [4325]. Need [1009, 1587, 1866, 1914, 2280, 3877, 4048, 4081, 4124, 4240, 4356]. needed [1606, 2919]. needing [638]. Needs [192, 852, 1137, 1138, 1973, 2225, 2420, 2797, 2992, 3240]. negates [4429]. negative [2655]. neglected [3245]. negligence [4439]. Negligent [1801]. negotiate [4417]. negotiates [303, 365]. Negotiating [3036]. neighbour [883]. neighbourhood [3773]. Nerd [3771, 3772]. Net [51, 67, 92, 116, 124, 215, 281, 290, 316, 328, 342, 356, 380, 431, 482, 840, 889, 982, 1105, 1241, 1252, 1590, 1688, 1738, 1827, 1832, 2264, 2368, 2450, 2496, 2649, 2801, 2888, 3682, 3789].

NetConnect [415]. NetCop [309]. Netect [934]. Netherlands [1782]. NetRegulation.inconfusion [1072]. NetScan [970]. Netscape [224, 297, 304, 307, 320, 360, 366, 480, 481, 559, 609, 713, 718, 778, 805, 854, 1228, 1267]. NetScreen [1008, 2243]. NetSec [1960]. Netsolve [600]. NetSP [16]. NetWare [9, 32, 155, 178, 523]. NetWitness [2878]. Network [43, 50, 94, 98, 121, 136, 141, 162, 184, 203, 208, 225, 234, 249, 251, 254, 266, 269, 283, 308, 336, 359, 383, 390, 404, 422, 425, 427, 455, 500, 502, 518, 526, 642, 689, 743, 773, 831, 833, 893, 895, 928, 1040, 1061, 1068, 1113, 1114, 1133, 1137, 1147, 1167, 1189, 1292, 1438, 1458, 1655, 1741, 1745, 1828, 1833, 1859, 1943, 2000, 2001, 2003, 2021, 2028, 2060, 2072, 2085, 2138, 2225, 2251, 2411, 2497, 2521, 2631, 2663, 2716, 2750, 2872, 3061, 3125, 3267, 3508, 3626, 3660, 3668, 3678, 3775, 4028, 4359]. network [1, 24, 82, 143, 166, 344, 367, 371, 388, 389, 432, 491, 495, 565, 648, 669, 705, 727, 754, 774, 798, 817, 836, 860, 879, 902, 920, 921, 933, 937, 955, 976, 980, 996, 1011, 1023, 1041, 1053, 1054, 1073, 1075, 1091, 1092, 1127, 1131, 1139, 1146, 1148, 1168, 1176, 1209, 1224, 1240, 1251, 1256, 1259, 1277, 1279, 1280, 1288, 1306, 1311, 1324, 1330, 1494, 1917, 1924, 2248, 2276, 2324, 2330, 2366, 2376, 2455, 2496, 2613, 2664, 2906, 2952, 3091, 3105, 3232, 3256, 3281, 3299, 3310, 3321, 3469, 3534, 3560, 3753, 3829, 4096, 4202, 4250, 4403, 4406]. network [158, 182, 871, 1144, 1337, 2405, 2517, 2721, 2752, 3104, 3297, 3936, 4342, 4428, 4438, 4451]. network-based [3560]. Network-related [269]. Networking [96, 539, 719, 1205, 1518, 1807, 2617, 2870, 3259, 3639, 4031]. Networks [23, 65, 76, 77, 81, 97, 99, 147, 258, 268, 294, 353, 424, 543, 563, 597, 616, 667, 741, 769, 816, 971, 1005, 1041, 1202, 1236, 1255, 1578, 1654, 2108, 2422, 2505, 2690, 2730, 2747, 2836, 2930, 2940, 2974, 2997, 3009, 3028, 3081, 3115, 3197, 3209, 3212, 3244, 3257, 3270, 3430, 3439, 3557, 3620, 3755, 3807, 3844, 3942, 4057, 4091, 4149, 4248, 4387, 4442, 4449, 4463]. Neumann [3731]. Newcombe [3363]. newgrp [633]. NewLove [1421]. News [108, 390, 494, 560, 711, 806, 1795, 1845, 2012, 3186, 3198, 4414]. Next [442, 559, 580, 1944, 2151, 2474, 2902, 3355, 3397, 3445, 3490, 3500, 3764, 3814, 3924, 3937]. Next-generation [3397, 3500]. NFC [3725]. NFS [3407]. NHS [2958, 4016]. NIAC [2040]. Nicholas [3891]. Nicko [2153]. Nicolson [3977]. NIDS [1833]. Nightmare [1807, 1832, 2060, 3883]. Nikhil [3209]. Nikkel [3955]. Nikolay [3710]. Nimda [1759, 1779]. NIPC [1589, 1686]. Nirvana [1807, 4248]. NIS [511, 1013]. Nitro [3264]. Nixon [3658]. No [909, 1736, 2342, 2595, 3067, 3095, 3232, 3243, 3252, 3287, 3392, 3403, 3406, 3689, 3700, 3710, 3731, 3752, 3761, 3794, 3813, 3846, 3863, 3887, 3891, 3892, 3955, 3966, 4286]. Nokia [1615, 4146]. Non [1567, 2444, 3096, 4205]. Non-contact [1567]. non-disclosure [3096]. non-functional [4205]. non-US [2444]. nonsense [2938]. norm [4366]. normal [3991, 4000]. Nortel [252]. North [2957, 3653, 4025, 4158, 4167, 4351, 4467]. northern [2512]. Norton [77]. Not-so-smart [2950]. Note [1094]. notebook [1573]. notebooks [227]. notes [31, 182, 429]. nothing [2385, 3596]. notice [1264]. notices [305, 1208]. notify [2553]. novel [2987]. Novell [29, 227, 282, 416, 1174]. Novelty [2074]. Nozomi [4387]. nph [694]. nph-test-cgi [694]. NSA [762, 1166, 1175, 1638, 1726, 1777, 3473, 3484, 3528, 3675, 3676, 3952, 4315]. NSF [2325]. NSS [3195]. NT [425, 451, 487, 562, 577, 594, 626, 646, 734, 749, 751, 769, 800, 815, 956, 989, 1019, 1067, 1089, 1124, 1225, 1230, 1242, 1246, 1282]. NT-based [769]. NTFS [2588]. NtOpenProcessToken [800]. nuclear [25, 3993]. nude [25]. nudge [4431]. number [41, 59, 740, 1453]. numbers [1903, 2225]. Numeric [2975]. nut [1834]. Nutan [3730]. NY [1644]. Nynex [8].

Obama [2898]. Obfuscation [2828]. object [2998, 3049]. object-oriented [2998, 3049]. OBS [2862]. obscurity [1679, 2442]. ObserveIT [4316]. Obsession [1876]. obstructs [878]. O’Connor [3454]. October [2469]. Ode [1440]. Odyssey [1863]. Off [5, 311, 2930, 4259, 4271]. Offensive [2733, 3468]. Offer [200, 282, 784, 847, 1062, 1261, 1745]. offered [500, 536, 744, 891, 1065]. offering [913, 1456]. offerings [574, 1234]. offers [485, 490, 505, 772, 912, 914, 1243, 1304, 1362, 1363, 1671, 4207]. office [683, 1204, 1637, 1783, 1859, 2622, 2844, 3963, 4284]. officer [755]. officers [3506]. official [210]. officially [3871]. offline [3228]. Offshore [2140]. often [4429]. O’Gorman [3243]. oil [3361]. Old [1331, 1419, 1851, 1997, 2373, 2785, 4199]. older [4218]. olive [1536]. olive-branch [1536]. Ollie [3720]. Olympic [2513]. omniguard [48]. On-LAN [577]. Once [3373]. One [295, 329, 858, 1304, 1558, 1599, 1930, 1939, 2519, 3495, 3592, 3612, 4061, 4141, 4237]. One-day [3612]. one-stop [2519]. one-time [295]. OneCare [2876]. Onion [1910, 2671]. Online [10, 18, 61, 93, 111, 159, 176, 177, 280, 284, 326, 327, 343, 417, 676, 707, 746, 891, 981, 1078, 1218, 1386, 1407, 1495, 1504, 1531, 1540, 1637, 1665, 1680, 1689, 1872, 2066, 2256, 2498, 2524, 2528, 2533, 2534, 2639, 2709, 2736, 2752, 2836, 2939, 2948, 3030, 3041, 3094, 3494, 4390]. Only [1597, 2140, 4259]. ons [2411]. Ontario [2767]. onto [102]. Open [520, 713, 867, 984, 1206, 1288, 1565, 1647, 1673, 1819, 1959, 2239, 2263, 2508, 2813, 3001, 3276, 3621, 3803, 3827, 3903, 4038, 4040, 4056, 4075, 4155, 4189, 4215, 4381, 4436]. open-source [1288, 4436]. OpeN/2 [520]. opened [482, 1447]. opening [3686, 3936, 4202]. openings [804]. opens [1488, 2039]. OpenSSH [1465]. OpenVMS [1028]. Operating [416, 565, 571, 1679, 1939, 3713]. Operation [2100, 2921, 3524, 4016, 4319]. Operational [1041, 4274, 4307]. Operations [169, 2104, 3828, 3882, 4433]. operators [60]. Opinion [2239]. opportunities [3258]. Opportunity [3894]. oppose [1166]. optic [8, 44, 220, 2681]. optical [65, 72]. optimal [3439]. optimisation [2781]. Optimising [2985, 3917, 4354]. option [2424]. Optional [2098]. Options [36, 167, 1312, 2113, 2706]. opts [1917]. Oracle [209, 363, 364, 431, 1312, 1900, 2412, 3526]. Oracle7 [363]. Orange [479]. order [1774]. Orders [1726]. O’Reilly [3344, 3345, 3352, 3353, 3383, 3393, 3658]. Organisation [2910, 3213, 3855, 3883, 4079, 4375]. organisational [4391, 4441]. organisations [3172, 3284, 3665, 3748, 3856, 3960, 4022, 4052, 4286, 4417, 4466]. Organised [4120]. organization [1493, 1996, 2538, 2554]. Organizations [129, 792, 1015, 2409]. oriented [2998, 3049]. origin [1416]. OS/2 [386]. OSF [476, 867]. OSF/1 [476]. OSF/DCE [867]. OSSEC [3486]. OT [4102, 4113, 4273, 4387]. OT/IoT [4387]. Other [981, 1395, 1618, 3146, 3171, 3563, 3738, 4115, 4384]. Othmar [799]. Our [1866, 2664]. ourselves [2218, 2395]. oust [2322]. out-number [740]. outages [1101, 3775]. outbreak [2492]. Outfoxing [241]. outlines [375, 1207, 2100]. outlook [1055]. outmoded [2952]. Outs [1789]. Outsiders [991]. Outsmarting [3145]. outsource [860]. Outsourced [2360, 2475]. Outsourcing [389, 600, 1503, 2140, 2557]. outweigh [2326]. over-hyping [3927]. over-run [631, 633, 908]. over-runs [964]. overcome [4343]. Overcoming [4101]. overflow [673, 759, 841, 1013, 1081, 1097, 1154, 1180, 1378, 1464]. overflows [1151]. overhaul [1587]. overload [2084, 3491]. oversight [1243]. overstayed [2507]. overtake [28]. overview [1263, 2419, 2863, 2929]. Own [367, 761, 975, 1827, 1911, 2506, 3427, 3558, 3663, 3884]. own-goals [2506]. owner [639]. owners [2968]. ownership [4427]. Oxford [4009]. Oxley [2444]. Oxymoron [528, 1640].

P [362]. PAC [3053]. Pacific [200]. pack [1020]. package [223, 1154]. Packard [1257, 1967]. Packet [346, 3232, 3944, 4039]. packs [2695]. Packt [3435, 3509, 3657, 3668]. Pagan [819]. pagers [165]. pages [101, 812]. paid [2066]. pain [2469, 3318]. Pains [4087]. pall [2743]. palm [1308, 1454]. PAM [4262]. Panacea [1417]. Panda [3730]. PandaLabs [3451]. Pandemic [4341, 4356, 4465]. pandemic-driven [4341]. Panel [660, 1726]. panic [3367, 3623]. Panix [625]. PANs [1909]. paper [868]. paperback [3334, 3444, 4065]. Paradigm [1743]. paradox [3715]. Paranoid [3364]. parasite [2327]. Parasites [2287]. Parasitic [2785]. Parker [1178]. Parliament [3964]. parliamentary [4155]. Parsing [1014]. Part [55, 146, 168, 184, 208, 234, 268, 294, 295, 315, 345, 346, 369, 390, 407, 430, 452, 471, 492, 506, 527, 549, 566, 586, 602, 604, 628, 648, 650, 669, 670, 689, 705, 727, 754, 774, 798, 817, 836, 838, 859, 860, 879, 880, 902, 921, 1010, 1024, 1025, 1042, 1132, 1149, 1212, 1225, 1241, 1242, 1328, 1346, 1597, 1725, 1765, 1832, 1836, 1857, 1931, 1942, 2128, 2130, 2131, 2538, 2554, 2819, 2861, 3009, 4102, 4113, 4141, 4154]. partial [2959]. parties [3069]. Partly [1708]. partners [1105]. Partnership [357, 988, 1016, 1126, 1140, 1201, 1448, 3149]. party [919, 1436]. Pass [2844]. passed [2822]. Passenger [2199]. Passing [1915]. Passive [870, 3829]. passport [2658]. passwd [1122]. Password [146, 168, 295, 443, 511, 631, 836, 1342, 2135, 2395, 2507, 2621, 2751, 2756, 2902, 3043, 3116, 3725, 3798, 4228].

Passwords [56, 165, 229, 1000, 1876, 2041, 2147, 2463, 2504, 3103, 3157, 3796, 3816, 4332]. Past [1599, 4043, 4093, 4453, 4474]. Paste [1079]. PASV [1136]. Patch [2023, 2129, 2212, 2268, 2469, 2866, 2905, 2953, 3815, 4092]. patch-fest [2469]. patched [2386, 2491]. patches [322, 1874, 2242, 2724]. patching [2285, 2969, 4235]. patchy [709]. path [1180, 2301, 3330]. Pathways [576, 662, 3514]. patient [1280, 2958, 3439]. Patrick [3277]. Patriot [2009]. patterns [1548, 2963, 3795]. Paul [2085, 2086, 3803]. Pauli [3476, 3477]. pave [2624]. paves [1231]. pay [1618, 3792]. paying [3125]. payment [1031, 1307, 1319, 4033]. payments [354, 486, 1780, 4334, 4417]. payouts [3899]. pays [1521]. PBX [206]. PC [29, 94, 519, 759, 1272, 1413, 1826]. PCCIP [1261]. PCI [3374]. PCMCIA [227]. PCs [814, 1318, 4244]. PDA [1606, 1826, 2180, 2286]. PDAs [1784]. PDF [1736, 3051]. peace [3093]. Peachy [1736]. Peacomm [2649]. Peapod [1033]. Pearson [1617]. Peekabooty [1667]. Peer [1618, 1954, 1981, 2027, 2963, 3115]. peer- [1981]. Peer-To-Peer [1618, 1954, 2027, 2963, 3115]. Peering [3818]. pegged [2944]. Peggy [3751]. Pen [2647, 4016, 4139]. Pen-testing [4016, 4139]. pendulum [2070]. penetrate [240]. Penetrating [3415]. Penetration [472, 817, 857, 1970, 2016, 2058, 2086, 3243, 3276, 3277, 3435, 3509, 3555, 3594, 3604, 3606, 3637, 3646, 3657, 4082]. penguin [1652]. Pentagon [310, 1726, 1830, 2172]. People [1724, 1773, 2070, 3836, 4288]. performance [852, 1455, 3397, 3660]. performing [97, 518, 1686]. Perils [2129, 3027]. perimeter [2567, 3355, 3438, 3550, 3982, 4308]. perimeterisation [2404, 2903]. perimeters [2021]. periodic [3226]. permission [394]. permissions [1398]. perplexed [3629]. Perry [3892, 3966]. Persistent [3070, 3236, 3567, 3692]. PersonaCard [103]. Personal [846, 1558, 2979, 3246, 3638]. perspective [370, 1055, 3069, 4307, 4378]. Perspectives [1970, 2016, 2058, 2086]. Pervasive [2369]. Peterson [3197]. petition [1157]. petrol [1559]. Pfizer [2293, 4467]. PGP [52, 101, 504, 627, 697, 807, 1317, 1491]. pgs [3334, 4065]. phasing [2885]. phasing-in [2885]. phenomenon [3907]. Phillips [3413, 3772]. phish [2500, 4003, 4175]. phished [2313]. Phishing [2264, 2368, 2468, 2477, 2590, 2637, 2685, 2910, 2914, 2929, 2942, 3109, 3207, 4025, 4175, 4292]. Phone [43, 45, 53, 63, 89, 99, 130, 218, 246, 400, 412, 445, 737, 745, 748, 895, 985, 1108, 1484, 1721, 2288, 2398, 3114]. phones [898, 1498]. photonics [7]. PHP [3658]. phreaking [3214]. Physical [371, 2096, 2426]. physically [3799]. PI [1141]. pick [3775]. PICS [641]. Picture [2144, 2158, 2169, 2179, 2188, 2196, 2210, 2220, 2238, 2265, 2290, 2304, 2317, 2751]. pictures [2329, 3653]. piercing [2975]. Pigeon [1723]. Pillars [4446]. pilot [310, 541]. pimping [3204]. pinch [1718]. ping [654]. Pinning [2568]. pinpoints [2307, 2668]. Pinto [3295]. Piracy [60, 62, 83, 114, 191, 1603, 1965, 2222, 3293]. Pirates [90, 275, 1191, 1738, 2121]. Pirc [3976]. Pitfall [1986]. pitfalls [2379, 2505, 2526, 3447, 4379]. PIX [1466]. pizza [1136]. PKI [965, 1017, 1140, 1159, 1161, 1183, 1184, 1211, 1220, 1263, 1302, 1320, 1370, 1393, 1404, 1410, 1412, 1418, 1427, 1564, 1637, 1764, 2438, 3210, 3825, 4429]. PKZIP [271]. PLA [3582]. place [334, 3846]. places [1301]. Placing [2178, 3046]. plan [9, 910, 2022, 2040]. planned [332, 2226, 2245, 4465]. Planning [956, 3352, 3353, 3721]. plans [160, 177, 1337, 2363, 2814]. plasters [1666]. Platform [679, 713, 1159, 1220, 1287, 1454, 1635, 2247, 2414, 2821, 3241, 3364]. platforms [3533, 3552, 3801, 4275]. Platt [3393]. play [875, 1500, 1551, 2337, 3866]. Playboy [828]. playboy.com [1820]. Player [2019]. Players [2412, 3344, 3345]. Playing [2490, 3884]. plays [1008, 1667]. plc [2668]. please [1863]. Plot [3956]. Plug [875, 1500, 1551, 2068]. Plug-and-play [1500, 1551]. pluggable [732]. Plugging [4102, 4113]. plugs [1549]. Plus [203, 1912]. PMTU [1350]. point [663, 752, 1207, 1343, 2203, 2321, 2322, 2414, 3463]. pointed [2957]. Points [2083, 2667, 3375, 3423]. poised [2370]. poison [4456]. poisoning [2788]. PoizonB0x [1740]. Poking [1639]. Police [237, 343, 394, 1590, 1626, 2308, 2435, 2652, 3293, 3842]. Policies [79, 147, 420, 775, 966, 1058, 1177, 1478, 2166, 2454]. Policy [292, 293, 408, 647, 843, 1006, 1021, 1188, 1369, 1606, 1945, 2568, 2927, 3001, 3611]. polite [22]. Political [1688, 1708, 1999, 4107, 4126]. Politicians [1276]. politics [1036]. Poll [1805]. Ponemon [2779, 4185]. Poor [1773, 1929, 2624, 4423]. poorly [2491, 3268]. POP [710]. popular [3119, 3364, 3617, 3831, 3899]. popularity [3324, 3552]. Porn [190, 272, 612, 1524, 1749, 1812, 2734]. pornography [327, 1812, 2564]. Port [437, 491, 1092, 1585, 1739, 2184, 4054]. portable [2579]. portals [1409]. portfolio [253]. ports [3138]. PoS [4025]. Pose [1828, 2956, 4278]. posed [3045, 3903]. poses [148, 210, 1469, 1887, 3335, 3970]. positive [2884, 3211]. possible [498, 1938, 2274, 2999, 3341, 3404]. possibly [3473]. post [1637, 1803]. postage [1250]. poster [1648]. Postfix [1226]. postmortem [2205]. postponed [4465]. Pot [1957]. Potential [421, 1129, 1425, 1930, 2480]. potentially [1641]. Potter [2083]. Power [681, 1557, 1727, 3646, 4310, 4454]. Power-cached [681]. Powerful [1448]. Powergen [1467]. Powers [2009, 3791]. PowerShell [3425]. PPTP [1044].

Practical [2715, 2773, 3232, 3392, 3944, 3955, 3989, 4161]. practicality [3663]. Practice [1451, 1702, 2117, 2494, 3626, 4137]. Practices [1162, 2115, 2416, 2600, 2624, 4251, 4261, 4436]. practitioners [3773]. Pragmatic [3363, 4065]. pranksters [881]. Pray [2068]. Preacher [1992]. Preble [3771]. precedent [1672]. precious [3895]. predator [2529]. Predators [1511]. Predicted [1101, 2136]. Predictions [986, 2744, 3140]. predicts [2092]. Preemptive [933]. prefer [198]. preferences [2411]. preferred [1574]. preliminaries [1834]. premature [1952]. Preparation [3873]. prepare [2647]. prepared [2357]. prepares [2776]. Preparing [1678, 3167, 3445, 3736]. present [3118, 4063, 4093]. Preserving [3059]. President [1505]. presidential [4209, 4397]. Press [799, 819, 1907, 3209, 3210, 3221, 3222, 3232, 3243, 3252, 3287, 3392, 3403, 3425, 3444, 3689, 3700, 3710, 3731, 3742, 3751, 3752, 3761, 3794, 3813, 3834, 3863, 3891, 3892, 3911, 3922, 3955, 3966, 3988, 4009]. Press/Elsevier [3742]. pressed [459]. Pressure [88, 3605, 3825]. pressures [3106]. prevalent [3909]. Prevent [774, 4238]. Preventing [83, 2871, 3039, 3083, 3138, 4202]. Prevention [737, 898, 1353, 2093, 2181, 2474, 2717, 2966, 3178, 3854, 3865]. prevents [1273, 1475]. prey [2441]. Price [1728, 2193, 2610, 3334, 3517, 4065]. prime [3324]. PrimeLife [2789]. Princeton [770, 1805, 3911, 3922]. Principal [2084]. Principles [2189]. print [1308, 3352, 3353, 3425, 3614]. Printers [1828, 3245]. printing [2521]. priorities [2455]. prioritising [4235]. priority [2172, 3142, 3854, 4066]. Pritchard [3751]. Privacy [10, 61, 595, 643, 981, 1163, 1195, 1208, 1217, 1280, 1405, 1515, 1580, 1586, 1592, 1613, 1691, 1692, 1714, 1829, 1863, 1936, 2257, 2303, 2398, 2840, 2948, 3023, 3344, 3345, 3385, 3629, 3669, 3837, 4194, 4380]. Private [120, 149, 197, 244, 502, 543, 719, 816, 1035, 1288, 1458, 1714, 3085, 3149, 3386]. Privatize [3509, 3544]. Privilege [899, 1268, 3336, 3501]. Privileged [3134, 3144, 3447, 4101]. Proactive [2408]. probably [232, 2766]. probe [1129]. probed [1820]. Problem [274, 551, 570, 778, 841, 907, 1620, 1708, 1724, 1930, 2199, 2235, 2609, 3456, 3495, 3949, 4062, 4403, 4451, 4461]. problematic [3617]. Problems [21, 368, 401, 435, 616, 786, 1027, 1196, 1237, 1618, 2015, 2804, 2974, 3232, 3381, 3589]. Procedures [147, 2505, 3575, 4389]. proceed [4465]. process [1228, 2997, 3916]. Processes [1645, 4354]. Processing [620, 3575]. processor [3105]. procfs [821]. produces [1126]. Product [308, 382, 386, 600, 650, 659, 670, 744, 833, 1162, 1416, 1452, 2229, 2418, 4226]. production [2741]. productive [4263]. productivity [4251]. products [141, 342, 578, 743, 1232, 1382, 1729, 2111, 2239, 2322, 2354]. Profession [2375]. Professional [3555]. Professionals [1790, 2425, 3151, 3305, 4270]. Professor [668]. Profile [733, 1681, 3573]. profitable [444]. profiting [2052]. profits [1629, 3218]. Proginet [9]. Program [119, 533, 632, 758, 780, 944, 1310, 1887, 1957, 3699]. programme [143, 584, 742, 796, 830, 887, 2417, 2520]. Programmer [3403]. Programmers [2046]. Programming [80, 3402, 3771, 3772]. programs [554, 823]. progress [4219, 4286]. Project [313, 587, 1881, 2130, 2789, 3581]. projects [4262]. promiscuous [80]. Promise [228, 445, 1985, 2961, 3365, 4429]. promises [2946]. promote [1139, 1299]. promoted [43]. promotes [2867]. prompt [1080, 1400]. prone [2107]. pronounces [2373]. Proof [619, 1090, 1721, 2890]. proofing [2464]. properly [1228, 2518]. property [268, 294, 3914]. proposal [362, 739]. proposals [1261, 2257]. proposed [1529]. Pros [1278, 2226, 2616, 3020]. Prosecute [1661]. prosecuted [275]. Prosecutes [1809]. Protect [255, 881, 1191, 1255, 1321, 1453, 1455, 1576, 1604, 1890, 2125, 3394, 3557, 3743, 3914, 4069, 4263, 4449]. protected [767, 917, 1487, 3172]. Protecting [24, 1280, 1688, 1833, 2108, 2118, 2486, 2601, 2995, 3084, 3146, 3246, 3248, 3307, 3310, 3469, 3532, 3585, 3616, 4149, 4204, 4225, 4248, 4273]. Protection [66, 76, 286, 402, 422, 443, 446, 524, 581, 597, 637, 642, 645, 660, 669, 683, 772, 874, 914, 948, 1011, 1061, 1087, 1195, 1305, 1369, 1448, 1498, 1577, 1696, 1834, 1894, 1947, 1990, 2243, 2519, 2539, 2851, 2868, 2902, 2944, 2952, 3175, 3449, 3460, 3512, 3536, 3547, 3665, 4306]. protects [361, 563, 792, 813, 833, 989, 1020, 1218]. Protek [1285]. protest [828]. protests [2079]. Protocol [315, 390, 542, 834, 853, 1699, 1938, 3969]. protocols [2581, 3330, 3484]. proved [1559]. provide [829, 1061, 1125, 1186, 1220, 1302, 1356, 1386, 1608, 1689]. Providers [174, 440, 739, 1132, 1149, 1595, 3946]. provides [659, 783, 890, 1035, 1106, 1107, 1340, 1450, 1519, 3780]. Providing [567, 3905]. proving [2484]. provision [243]. provokes [2388]. prowl [2889]. Proxies [2822, 3395]. proxy [3053]. ps [321]. psyche [3280]. psychology [3135]. psychos [881]. PTK [2795]. ptrace [1247]. Public [23, 210, 304, 413, 502, 665, 814, 938, 1364, 1417, 1507, 1579, 2306, 3149, 3728, 4164, 4429]. Public-Key [413, 1507]. public-private [3149]. publication [4465]. Publicity [1579]. Published [149, 3184, 3363, 4065]. Publisher [1178]. publishes [827]. Publishing [1482, 3363, 3373, 3435, 3509, 3657, 3668, 3956]. Puerto [892]. punch [2695, 4247]. purchase [457]. purchasers [280]. purchasing [2485]. purpose [3126]. push [2124, 2306, 4033]. Pushes [10, 249, 253, 1830, 2374, 2658]. put [1059, 4117, 4295]. puts [243, 261, 3605, 4170].

Putting [229, 1055, 2691, 3225, 4011].

Pwn2Own [3055]. pwned [2915]. Python [3454, 3614, 3689, 3752].

Q&A [3458]. Q1 [2791]. Q4 [4338]. Qaeda [3228]. qualified [2140]. quality [3715]. Qualys [2401]. QualysGuard [2476]. Quantitative [1083]. Quantum [1509, 1855, 1974, 1985, 2153, 2440, 2561, 4114, 4404]. quarter [703]. Quarterly [4099]. Que [1557]. queen [2118]. quest [3547]. question [860, 2553]. questioned [911]. Questions [1174, 2133, 3203]. Quick [2048]. quicksand [3732]. Quiet [1975]. quite [1736].

R [3374, 3844, 3863, 4008]. Rabbit [4005]. race [2406]. Rack [3449]. radar [3776, 4178]. Radicchio [1338]. Radio [281]. rage [1967]. rages [967, 994]. Raggo [3833]. railways [2809]. Rainbow [713]. raised [1256]. raises [626]. raising [45, 3458]. rally [625]. Rambus [1722]. ramifications [2652]. rampant [3517]. Ramsey [777]. Ranch [3771, 3772]. range [3778, 3909]. ranked [634]. ransom [2066, 2566, 2835, 3517]. ransoms [3792, 3960]. Ransomware [3759, 3801, 3810, 3851, 3865, 3875, 3876, 3889, 3899, 3909, 3928, 3967, 3986, 4005, 4290, 4319, 4334, 4354, 4417]. rants [493]. Rapid [6, 1385]. Rapid7 [4401]. Raptor [578, 580, 785, 1063]. Rash [2114, 3086]. rate [766, 1775]. Rather [1809]. Raxco [48]. Razor [2718]. RC2 [827]. RC5 [900]. rdist [556, 841, 1097]. RDP [4218]. re [2580]. re-thinking [2580]. reach [1513, 4181]. reached [575]. reaches [247, 377]. react [4022]. reactions [3528]. Reactor [1098]. read [1487]. reader [1050]. Readiness [4210, 4391]. ready [1737, 1886, 3665, 4122, 4341]. Real [650, 670, 1091, 1681, 1879, 2142, 2252, 2440, 2563, 2793, 2886, 2928, 2936, 2940, 3020, 3232, 3240, 3396, 3510, 3903]. Real-time [2940]. real-world [2936, 3232]. reality [2204, 2299, 4105]. Really [997, 1393, 2070, 2235, 2783, 3011]. realm [1226]. reassembly [1426]. Rebecca [1482]. rebirth [3214]. rebounds [3015]. receive [966]. receives [158, 1183]. recently [1641]. recession [2956]. recognition [2808]. recommended [611, 641, 1175]. recommends [1726]. Reconnaissance [238, 2872]. record [4181]. records [794, 1673, 2556, 2854, 4242]. Recover [3509, 3544]. recoverable [567]. recovery [94, 593, 595, 909, 929, 948, 969, 1022, 2363, 3428, 3488, 4137]. Recycling [3008]. Red [955, 1733, 1746, 1757, 2845, 3814, 3882, 4059]. redefines [720]. redirects [3053]. Reduce [82, 1612, 3076, 3925, 4082, 4354]. reduced [1143]. reduces [1014, 4098]. Reducing [166, 1146, 3701, 3755, 4058, 4318]. Reference [3185]. Reflection [1808]. Reflex [596, 1368]. reform [2280]. reformat [1867]. Regin [3644]. registers [1921]. registrar [66]. registration [277, 1218]. registry [1398, 1570, 2636]. Regular [3383]. Regulation [272, 611, 1619, 2174, 2424, 4123]. Regulations [615, 636, 657, 787, 825, 1082, 1195, 1367, 1403, 1805, 2343, 4192]. regulator [4188]. regulators [3344, 3345]. rejected [484]. related [269, 995, 4073, 4426]. relationship [4352]. Relative [1929]. Relativistic [798]. relaxes [1058, 1382]. relaxing [825]. relay [198, 361, 502, 1186, 1236]. Release [1699, 1777, 1962, 2024]. released [1130, 1732]. releases [1553]. Relevance [3693]. relevant [4114]. relief [1166]. reload [1424]. Relying [4450]. remain [416, 2276, 3653, 3919]. remains [4313]. remedies [3795]. remedy [2504]. Remote [21, 46, 345, 446, 664, 747, 818, 832, 873, 890, 895, 913, 1046, 1124, 1204, 1464, 1489, 1597, 1761, 1869, 1903, 1923, 2253, 3310, 3337, 3714, 4419]. remotely [3532, 4021]. remove [826]. removed [946]. repair [746, 1306]. replace [1401, 2181]. replaced [1052]. Replacement [2011]. Replacing [2504, 3030, 3206]. replay [4032]. Report [136, 414, 979, 1053, 1104, 1723, 1784, 2152, 2280, 2626, 3423, 4021, 4099, 4146, 4157, 4159, 4210, 4222, 4233, 4293, 4304, 4326, 4338, 4375, 4378, 4387, 4413, 4423]. reported [1375, 1399, 1522]. Reporter [2152]. reporting [1554, 1603, 2910]. reports [2275, 3996, 4182]. representative [503]. reputation [3095]. reputational [4188]. request [1229]. require [1649, 3721]. required [909, 4419]. requirement [4205]. Requirements [169, 2538, 2820, 3968]. requires [1342]. Research [742, 1139, 2006, 2052, 2221, 2237, 2307, 2417, 3368, 4077]. Researcher [2371, 2815, 2930, 2946, 3042]. Researchers [2460, 2801, 2826, 2954, 2991, 3707, 4348]. resell [785]. resells [578]. reset [1466, 4419]. Resetting [3524]. Resilience [3161, 3317, 3948, 4298, 4382]. Resilient [4375]. resistance [4382]. Resolved [1392]. Resources [409, 565, 1487, 2337]. respond [774]. respondents [1399]. responder [3029]. Response [158, 566, 572, 586, 1294, 1910, 1944, 2157, 2328, 2462, 2608, 2748, 2873, 2884, 3545, 3690, 3701, 3873, 3876, 4022, 4099, 4252, 4403, 4451]. responsibilities [1175, 2770, 3029]. responsibility [1132, 1149, 2783]. responsible [2712, 3653]. Rest [2104, 2601, 2820]. Restoring [3882]. restricted [696]. restrictions [11, 477, 765]. result [1288, 3158]. resulting [1352]. results [3399]. resurgence [2785, 3171]. retail [1921, 2930]. Retailers [3915]. Rethink [4236, 4306, 4424]. Rethinking [3290, 3714]. Return [2060, 2484, 2999]. Returning [1148]. returns [2492]. Reuters [2360]. Revamping [51]. Reveal [1593, 3675]. Revealed [2057, 2879]. reveals [489, 610, 984, 2548, 4433]. revelations [3528]. revenge [1753]. revenues [462]. reverse [2963]. Review [2004, 3184, 3197, 3209, 3210, 3221, 3222, 3232, 3243, 3252, 3253, 3266, 3276, 3277, 3287, 3295, 3304, 3305, 3326, 3327, 3344, 3345, 3373, 3374, 3383, 3384, 3413, 3414, 3425, 3435, 3444, 3454, 3466, 3476, 3477, 3509, 3519, 3555, 3614, 3657, 3658, 3678, 3679, 3689, 3699, 3700, 3710, 3720, 3741, 3742, 3751, 3752, 3761, 3771, 3772, 3783, 3784, 3794, 3813, 3814, 3823, 3824, 3833, 3834, 3844, 3845, 3853, 3863, 3864, 3873, 3881, 3891, 3892, 3901, 3911, 3922, 3955, 3956, 3966, 3976, 3977, 3988, 3998, 4008, 4009, 4089]. Review [149, 184, 208, 234, 799, 819, 881, 1178, 1278, 1482, 1617, 1763, 1836, 1857, 2207, 2221, 2237, 2350, 2945, 2965, 3046, 3334, 3352, 3353, 3363, 3392, 3393, 3402, 3403, 3730, 3731, 3803, 3804, 3933, 3944, 3945, 4065, 4132, 4281]. reviewKali [3668]. reviews [650, 670]. Revised [1367]. Revisions [548, 4372]. Revisited [1345, 1974]. Revisiting [1533]. revitalising [2640]. Revolution [1705, 3481]. revolutionary [878]. rewards [2448]. Reworking [2005]. RFID [2410, 2658, 2715, 2767]. Rhodri [3945]. Rican [892]. Richard [1557]. Rickey [3864]. Right [518, 529, 535, 1439, 2058, 2472, 2556, 3503, 4022, 4123]. Rights [2125, 2355]. Ring [38, 4300]. Ringing [1860, 1886]. RIP [1528]. Rise [41, 44, 462, 1812, 2422, 2687, 2746, 2993, 3048, 3051, 3089, 3541, 3542, 3738, 4073, 4212, 4254]. rises [2900]. rising [3647, 4340]. Risk [112, 166, 727, 798, 799, 868, 996, 1083, 1114, 1146, 1444, 1465, 1533, 1679, 1784, 1828, 2294, 2498, 2518, 2548, 2550, 2557, 2617, 2678, 2730, 2916, 3076, 3114, 3253, 3705, 3712, 3734, 3858, 3866, 3894, 3903, 3917, 3925, 3978, 4075, 4082, 4125, 4170, 4295, 4309, 4318, 4344, 4354, 4412]. Risk-based [3866]. Risks [82, 1418, 1612, 1835, 2103, 2155, 2285, 2326, 2475, 2660, 2668, 2710, 2805, 3765, 3774, 4101, 4163, 4343]. Risky [4215]. rivalry [514]. rlogin [672, 864]. rlogin/Term [672]. rlogind [863]. RMS [2125]. road [312, 954, 2301, 3494, 4103]. roadblocks [4284]. roaming [829]. robbers [1078]. Robert [3803]. Robin [3658]. robot [4039]. robotic [4275]. Rocher [4254]. rocks [2834]. rocky [658]. rods [549]. Rogers [3305]. Rogue [1238, 1938, 2083, 2143, 2611]. Rogueware [2922, 2970]. ROI [2661]. Role [22, 289, 507, 1726, 1947, 2100, 2337, 2972, 3354, 3405, 3426, 3489, 3866, 4250]. roles [2708]. roll [703, 2657, 2744]. Roman [295]. room [4094]. Root [1268, 1324, 1379, 1689, 2648]. Rootkit [2509, 2815]. rootkits [2782, 2871]. Round [704, 771, 1901, 2551, 3122]. round-up [3122]. Roundup [4110]. Rountree [3184]. router [181, 992, 1046]. Routers [344, 463, 1131, 1910, 2181]. routes [2517]. Routing [2671, 3199, 3330, 4028, 4287]. Row [243]. RPC [2158, 2273, 2866]. rpc.cmsd [957]. rpc.statd [1215]. rpc.ypupdated [411]. RPKI [4287]. RSA [365, 396, 575, 674, 717, 827, 935, 1066, 3181, 3206]. RTF [1351]. Ruby [3252]. Rule [1785, 1862]. rules [10, 1382, 1794, 2024, 2277, 2655, 3665]. rumours [3121]. run [631, 633, 849, 908, 1447]. Running [2643]. runpriv [758]. runs [964, 3517]. Russ [3305]. Russia [3260, 3871, 4005, 4130, 4143]. Russian [1347, 2402, 3601, 3641, 3931, 3952, 4096, 4358]. Russians [130, 3927, 4278].

S [209, 834]. S/MIME [834]. Sabotage [1641]. Safe [23, 622, 623, 910, 986, 1047, 1260, 1262, 1266, 1555, 2140, 2499, 4134, 4332, 4366]. Safeguarding [268, 294, 4367]. safely [502]. safer [334, 1354, 1363]. Safety [911, 1054, 2496, 2537, 2809, 4019, 4046]. Sagishi [4162]. Saha [3752]. sake [930]. sale [90, 2437]. Salesforce [4101]. Sam [954]. Samba [2103]. Sambi [3209]. Same [1420, 2046, 4104]. Samia [3742]. Sammons [3326, 3327]. San [1960, 2835]. Sánchez [3834]. sanctions [4421]. sand [3570]. sandbox [1663]. Sandboxing [3570].

Sanders [3232, 3944]. Sanford [3998]. Sanjay [3209]. SANS.org [1709]. Sarbanes [2444]. SASE [4438]. Sasser [2271, 2309]. Sasser-like [2271]. SATAN [188, 230]. satellite [45, 560, 1517, 1951, 3683]. save [1432]. saves [1739]. saving [3095]. Savvy [1481]. say [1211, 1714, 1850, 2225, 3707]. Says [262, 1104, 1479, 1624, 1631, 1781, 1784, 1841, 2241, 2280, 2357, 2436, 2481, 2563, 2968, 3005, 3051, 3240, 3370, 3451, 3452, 3482, 3495, 3633, 3641, 3696, 3820, 3841, 4043]. SCADA [2758, 2977, 3620]. Scalable [2798, 2951]. scale [1591, 2940, 3619]. Scales [142, 1622]. scam [824, 1645]. Scammers [3000, 4361]. scams [3901, 3941, 4033]. scan [3267]. scanned [896, 1413]. scanner [337, 562, 598, 1108, 1388, 1524, 1902]. Scanners [2059, 2497, 2980]. Scanning [1335, 1585, 3668]. scans [309, 382]. scare [320, 609, 4288]. Scared [1657]. scares [3052]. scene [1563]. Sceptical [1723]. scheme [607, 763, 842, 1936, 2743]. Schmidt [3413]. Schmitt [3444]. Schneier [149, 1781]. schools [2390]. Science [1773, 2417, 2440]. SCM [1611]. score [837]. Scotland [38]. Scott [819]. scraping [3638]. screen [1372, 2292]. Screen-melting [1372]. screening [1300]. screensaver [1573]. script [694]. Scripted [1079]. Scripting [1266, 1873, 2002, 2578, 2602, 3328]. Scripts [3892]. scupper [2347]. SD [3904, 3968, 4124]. SD-WAN [3904, 3968, 4124]. SDI [14]. SDL [2895]. SDLC [338, 2885]. SDN [3426]. SDR [3335]. SeaChange [1033]. seal [1611]. Sean [3305]. search [2504, 2548, 2621, 2774, 2781, 2980, 3023, 3550]. searches [2250]. searching [3623]. seas [62]. season [661, 1822]. seasoned [3924]. seat [2713]. Seattle [2692]. SEC [428]. Second [703, 771, 1235, 3232, 3266, 3383, 3555, 3751, 3772, 3892, 4245]. Secondary [1929]. Secret [1980, 2057, 2504, 2547, 2826, 3501, 3977].

Secrets [226, 2072, 2801, 3922]. sector [2239, 3728, 3799]. sectors [3985]. Secure [19, 21, 47, 121, 163, 206, 209, 278, 288, 342, 354, 399, 419, 422, 431, 445, 447, 483, 486, 539, 543, 573, 618, 719, 735, 812, 825, 829, 847, 853, 873, 877, 890, 891, 913, 988, 997, 1050, 1062, 1085, 1107, 1219, 1226, 1236, 1286, 1289, 1299, 1302, 1307, 1319, 1338, 1362, 1386, 1390, 1391, 1429, 1452, 1472, 1473, 1499, 1607, 1615, 1680, 1695, 1717, 1780, 1859, 1934, 1935, 1943, 1959, 1993, 2028, 2095, 2099, 2409, 2499, 2534, 2644, 3081, 3202, 3271, 3439, 3509, 3544, 3559, 3707, 4191, 4365, 4394]. secure [470, 548, 567, 815, 936, 1428, 1559, 1777, 2082, 2333, 2682, 2830, 2841, 2857, 3175, 3290, 3291, 3489, 3588, 3630, 3670, 3680, 3723, 3754, 3789, 3869, 3904, 4021, 4047, 4080, 4177, 4418]. secured [78, 518, 635, 872, 1357, 1360, 1407, 1630]. Securely [835, 1004, 1016, 1409, 3470, 3526, 4288]. secures [14, 122, 1069, 1202, 1496, 1609, 2259]. SecureWay [702]. SecureZone [1109]. Securicor [134]. SecurID [463, 3181, 3206]. Securing [919, 1330, 1679, 1880, 2029, 2216, 2233, 2262, 2455, 2505, 2528, 2538, 2554, 2579, 2589, 2600, 2620, 2690, 2728, 2747, 2974, 2977, 2997, 3007, 3009, 3028, 3080, 3138, 3144, 3309, 3363, 3416, 3741, 3755, 3784, 3849, 3866, 3916, 3961, 4027, 4030, 4039, 4136, 4141, 4148, 4154, 4161, 4187, 4249, 4308, 4442]. Security [15, 37, 85, 203, 277, 401, 464, 513, 536, 570, 642, 667, 682, 802, 821, 1030, 1044, 1121, 1139, 1160, 1335, 1414, 1526, 1591, 1634, 1694, 1778, 1795, 1860, 1878, 1891, 1905, 1918, 1926, 1964, 1966, 1970, 1973, 1979, 1981, 1986, 1988, 2000, 2001, 2004, 2008, 2019, 2034, 2035, 2065, 2082, 2089, 2111, 2128, 2129, 2138, 2152, 2332, 2606, 2646, 2858, 2933, 3061, 3151, 3266, 3274, 3304, 3305, 3506, 3508, 3509, 3520, 3544, 3604, 3615, 3626, 3699, 3710, 3823, 4008, 4075, 4077, 4110, 4132, 4171, 4197, 4222]. Security [33, 98, 144, 149, 169, 263, 312, 389, 468, 547, 689, 699, 951, 995, 1040, 1112, 1114, 1130, 1133, 1189, 1292, 1346, 1369, 1420, 1438, 1458, 1461, 1578–1580, 1594, 1595, 1597, 1614, 1617, 1621, 1640, 1740, 1745, 1767, 1828, 1835, 1882, 1911, 1931, 1942, 1944, 1945, 2018, 2030, 2043, 2046, 2098, 2102, 2104, 2116, 2140, 2142, 2151, 2154, 2176, 2180, 2209, 2248, 2260, 2312, 2429, 2433, 2442, 2520, 2553, 2587, 2642, 2651, 2719, 2748, 2809, 2917, 2972, 2976, 3176, 3319, 3446, 3627, 3693, 3797, 4012, 4020, 4286, 4304, 4307, 4338, 4387, 4424, 4435, 4446, 4459, 4472]. security [16, 29, 36, 74, 115, 134, 138, 140, 141, 192, 195, 213, 221, 223–227, 251, 253, 254, 257, 259, 287, 302, 307, 308, 320, 322, 331, 352, 359, 379, 383, 385, 386, 404, 423– 425, 435, 440, 456, 457, 461, 484, 487, 500, 515, 517, 522, 532, 561, 572, 574, 578, 594, 613, 617, 634, 644, 646, 659, 663, 664, 676, 680, 709, 717, 718, 720, 733, 743, 744, 766, 778, 783, 809, 811, 814, 831, 848, 852, 897, 928, 1018, 1105]. security [918, 930, 949, 966, 971, 972, 976, 990, 1003, 1015, 1019, 1028, 1032, 1034, 1050, 1068, 1083, 1089, 1099, 1106, 1137, 1138, 1162, 1184, 1186, 1232, 1255, 1301, 1303, 1306, 1340, 1348, 1358, 1359, 1361, 1389, 1412, 1430, 1450, 1465, 1469, 1471, 1474, 1486, 1493– 1495, 1500, 1503, 1512, 1546–1548, 1608, 1729, 1773, 1846, 1852, 1866, 1874, 1886, 1953, 1962, 1998, 2037, 2069, 2073, 2078, 2083–2085, 2096, 2100, 2133, 2183, 2225, 2229, 2303, 2414, 2425, 2524, 2573, 2583, 2615, 2684, 2733, 2744]. security [103, 266, 291, 340, 366, 371, 648, 669, 705, 727, 752, 754, 774, 788, 796, 798, 1328, 1678, 2300, 2418, 2521, 2657, 2668, 2694, 2723, 2724, 2740, 2769, 2776, 2802, 2846, 2855, 2867, 2956, 2980, 2992, 3059, 3076, 3077, 3082, 3098, 3141, 3142, 3149, 3152, 3161, 3187, 3190, 3212, 3240, 3250, 3262, 3282, 3284, 3298, 3404, 3407, 3429, 3455, 3498, 3502, 3530, 3573, 3579, 3589, 3601, 3602, 3623, 3673, 3681, 3728, 3748, 3774, 3860, 3879, 3919, 3959, 4085, 4091, 4145, 4184, 4257, 4263, 4267, 4275, 4300, 4318, 4320, 4352, 4398, 4404, 4423]. security [82, 105, 110, 142, 143, 406, 432, 491, 773, 817, 836, 860, 879, 902, 920, 921, 937, 955, 980, 996, 1009, 1011, 1023, 1053, 1071, 1073, 1091, 1092, 1113, 1147, 1148, 1167, 1168, 1174, 1176, 1209, 1224, 1240, 1257–1259, 1277, 1278, 1280, 1311, 1679, 1704, 2194, 2218, 2352, 2379, 2411, 2417, 2472, 2483, 2484, 2526, 2567, 2568, 2652, 2739, 2772, 2786, 2797, 2818, 2842, 2853, 2885, 2896, 2966, 2988, 2989, 3011, 3026, 3037, 3069, 3094, 3147, 3163, 3177, 3211, 3306, 3418, 3430, 3537, 3659, 3763, 3798, 3836, 3838, 3839, 3905, 3949, 3981, 4330, 4368, 4377, 4406, 4407, 4441]. security [17, 147, 289, 311, 318, 427, 668, 755, 903, 1339, 1400, 1558, 2014, 2021, 2042, 2214, 2226, 2239, 2251, 2257, 2276, 2286, 2299, 2301, 2354, 2365, 2370, 2373, 2376, 2377, 2384, 2394, 2408, 2413, 2419, 2426, 2430, 2454, 2485, 2513, 2536, 2563, 2622, 2689, 2691, 2771, 2882, 2928, 2948, 3047, 3215, 3281, 3394, 3397, 3406, 3427, 3449, 3458, 3503, 3535, 3547, 3576, 3598, 3640, 3660, 3661, 3712, 3735, 3753, 3766, 3773, 3817, 3828, 3829, 3946, 4022, 4031, 4056, 4066, 4094, 4112, 4114, 4164, 4203, 4226, 4236, 4250, 4274, 4285, 4287, 4343, 4345, 4381, 4454]. security [1, 2, 79, 181, 182, 528, 599, 713, 725, 749, 751, 760, 789, 799, 865, 894, 952, 956, 1096, 1145, 1275, 1341, 1698, 1877, 2166, 2188, 2191, 2195, 2205, 2252, 2272, 2284, 2324, 2326, 2364, 2380, 2404, 2420, 2465, 2475, 2488, 2494, 2498, 2531, 2558, 2612, 2623, 2639, 2660, 2661, 2663, 2680, 2708, 2710, 2717, 2770, 2783, 2832, 2916, 2987, 2996, 3036, 3060, 3106, 3299, 3308, 3321, 3346, 3347, 3363, 3365, 3385, 3437, 3467, 3478, 3533, 3702, 3722, 3732, 3785, 3866, 3877, 3893, 3896, 3923, 3927, 3937, 4011, 4071, 4079, 4124, 4173, 4251, 4364, 4392, 4428, 4431, 4439]. security [519, 992, 1763, 1783, 2100, 2540, 2662, 2721, 2752, 3184, 3247, 3253, 3549, 3663, 3678, 3794, 3881, 3958, 3968, 4063, 4438]. security-based [1678]. security-positive [3211]. SecurlD [813]. SecurVPN [873]. see [1903, 2287, 2751, 2991, 3279, 3724]. seedy [1040]. Seek [3366, 4190]. Seeking [133, 3354, 4037]. seeks [540]. seems [4168]. seen [1641]. sees [6, 498]. segmentation [4187, 4428]. Seitz [3689]. Selecting [2701, 3789, 4342]. Self [65, 611, 2108, 4430]. Self-healing [65]. self-protecting [2108]. self-regulation [611]. self-sovereignty [4430]. Selfie [4292]. sell [177, 1063, 1698, 2438]. Selling [583, 1655, 2786, 2979]. Semantic [1895]. Sendmail [172, 298, 323, 369, 590, 606, 1121, 1198, 1226, 2103, 2169]. sends [1156]. Senior [2084, 2138, 2616]. sense [2084, 2938, 3188, 3905]. sensitive [1630, 2592, 3394, 3470, 3618, 3631]. sensor [3257]. sensors [2850, 2861]. sent [502, 1711]. sentenced [190]. SeNTry [544]. SentryLink [199]. SEO [2788]. separating [2299]. Serbian [1442]. Series [2119]. Serious [1469, 1934, 1946, 1986, 3154, 3312]. Seriously [1749, 4052]. Server [33, 49, 201, 223, 424, 465, 561, 698, 700, 784, 806, 885, 1128, 1180, 1201, 1228, 1322, 1341, 1424, 1473, 1552, 1725, 1765, 2051, 2095, 2144, 2274, 3014, 3297, 3741, 3874]. server-based [1201]. Servers [455, 711, 718, 1151, 1246, 1340, 1397, 1486, 1725, 1734, 1765, 1923, 2029, 2122, 2143, 2648, 2860, 3563, 4010, 4117, 4177, 4312, 4337, 4358]. Serverware [544]. serves [1264, 2811]. Service [93, 134, 222, 238, 311, 437, 440, 460, 516, 521, 654, 730, 736, 867, 905, 943, 959, 970, 1119, 1132, 1149, 1198, 1230, 1273, 1334, 1350, 1352, 1443, 1477, 1503, 1595, 1618, 1685, 1720, 1834, 1925, 2101, 2547, 2771, 2957, 3068, 3257, 3805, 3946, 4380]. Services [5, 131, 196, 200, 333, 447, 1110, 1164, 1544, 1904, 1988, 1998, 2102, 2154, 2429, 2524, 2615, 2776, 3246, 3313, 3363, 3365, 3563, 3585]. serving [345]. servitude [2301]. Session [183, 2914, 3905]. Sessions [1423, 2117]. SessionWall [792, 893]. SessionWall-3 [792, 893]. Set [115, 220, 333, 548, 635, 808, 869, 927, 1031, 1694, 1697, 1717, 1778, 1821, 2149, 3097]. Seth [3414, 3503, 3824]. sets [1672, 3718]. Setting [2556, 3435]. settle [291]. Setup [1463]. Seven [2277, 2841]. Several [1930, 3256]. severe [3005]. severity [1775]. Sex [453]. SGI [653, 823, 842, 882, 999, 1118, 1180]. shade [1900]. shadow [4163, 4340]. Shadows [1557, 3070, 3436, 3521]. shakes [2363]. Shane [3402]. Shannon [3209]. Share [1858]. sharers [3052]. shares [1710]. Sharing [1618, 1954, 2199, 2514, 2610, 2963, 3027, 3202, 3470, 3773]. Shark [1019]. Sharma [3209]. Sharp [4073]. Sharpening [2608]. sharpens [2294]. Shaun [3720]. shave [2392]. Shavers [3933]. Shedding [2582]. Shell [652, 1869, 3657, 3892]. Shellshock [3623]. Shema [3384, 3466]. Shetland [940]. Shields [2995]. shift [1618, 2321, 2616, 2704]. Shifts [2630, 4362]. shines [1998]. Shining [4452]. ships [2868]. Shiva [14, 252]. Shock [1870, 2034]. Shop [311, 1304, 1637]. shopping [163, 548, 573, 635, 1495]. Shore [960]. short [2277, 4009, 4441]. shortage [1559]. shortcomings [3195]. Should [669, 1053, 1813, 2043, 2351, 3682, 3815, 3857, 4010, 4022, 4236, 4259, 4288, 4331]. Shouldn’t [1702]. show [954, 1949]. Showdown [1656]. Shows [1348, 1805, 2617, 3952]. Shredder [978]. SHTTP [164]. shut [734, 4279]. shuts [1818]. shutting [4119]. Side [471, 1040, 1093, 1532, 1915, 2556, 2634, 2716, 2986]. Side-channel [2634]. sideways [2453]. Sidewinder [78]. SIEM [3191, 3578, 3597, 4276]. sights [3718]. sign [317, 358, 383, 522, 546]. sign-on [317, 358, 383, 522, 546]. Signature [722, 1505]. Signatures [1049, 1086, 1355, 1405, 1575, 1650, 1689, 4048]. signed [1816]. significant [977, 3633, 3738]. Signing [793, 1329, 1411, 3093]. Signs [1505, 4321]. Sikich [4233]. Sikorski [3392]. Silent [3492, 3858, 4320]. silicon [373, 1090]. Silk [3494]. Silvia [3352, 3353]. SIM [2874]. similar [1618]. Simmons [3373]. Simple [144, 267, 1077, 2277, 2925, 3358, 3487]. simplified [1160, 1203, 1433]. simplifies [1270]. simplify [1030, 2580]. Simplifying [4065]. simulates [1129]. Simulating [1176]. simulators [585]. simultaneously [441]. Singapore [440, 830, 1335, 1584]. Singh [3509]. Singing [1094]. Single [317, 358, 383, 522, 546, 663, 2636]. single-database [2636]. single-point [663]. Sinister [3160]. sinkholes [3356]. sinks [2156]. sins [2307]. SIP [3839]. SirCam [1734]. SITA [222]. site [379, 656, 1156, 1191, 1192, 1383, 1795, 1845, 1873, 2002, 2256, 2578, 2602, 2790, 2992, 3328, 4279]. Sites [584, 733, 945, 1002, 1047, 1084, 1194, 1208, 1375, 1740, 1753, 2765, 3053, 3831, 4086]. situation [2758]. Six [1252, 2842, 3271, 4001]. six-step [3271]. sixth [1010, 1025]. size [2609, 3849]. Sizing [3057]. Skeleton [3126]. Skewed [2347]. Skills [1075, 3151, 3845, 3856, 4296]. SKIP [764]. Sklyarov [1842]. Skyhorse [3956]. Skype [2510]. slacken [1173]. Slammer [2049, 2060, 2071]. slap [1995]. slashes [2037]. Sleuth [2795]. slim [3060]. slip [1996]. slip-ups [1996]. slow [1397]. Small [180, 2529, 2545, 3849]. smaller [1387]. Smart [1110, 1248, 1304, 1320, 1406, 1418, 1431, 1502, 1525, 1567, 1660, 2946, 2950, 3187, 3417, 3498, 3596, 3877, 4428]. smartcard [665, 717, 1003, 1035, 1854]. smartcards [614, 1601, 2634]. Smarten [3712]. Smarter [2571, 3000, 3522]. smartphone [2968, 3088, 3145, 3182]. smartphones [2298, 2537, 3118, 3765]. Smarts [3305]. SMB [3105]. SME [3597]. SMEs [2370, 4134]. smile [2862]. Smith [3813]. SMS [1883]. smurf [905]. snafu [678]. snaps [2857]. Sniffers [603]. Sniffing [34, 1238]. SNMP [340, 1291, 1886, 1900]. snoop [1323]. snooped [3064]. snooping [953, 1843, 2009, 3528]. Snort [2103, 2411]. Snowden [3675]. SOA [3290]. SOAPBOX [1683]. soar [1694, 4356]. soars [451]. SOARX [4276]. Sober [2201, 2492]. Sobig [2141, 2149, 2167]. Sobig.F [2187]. SOC [4016, 4264]. Social [267, 495, 1724, 2033, 2198, 2617, 2870, 3135, 3259, 3637, 3639, 3784, 3826, 3901, 3941]. society [22, 518, 1231]. SOCKS [144]. SoftID [525]. Software [83, 202, 236, 290, 320, 376, 513, 520, 525, 532, 644, 687, 752, 791, 978, 1048, 1098, 1184, 1199, 1214, 1248, 1317, 1325, 1411, 1594, 1633, 1917, 2082, 2277, 2324, 2662, 2813, 2831, 2878, 2886, 3097, 3404, 3548, 3558, 3717, 3761, 3774, 3857, 3903, 3925, 3978, 4162, 4199, 4202, 4377, 4459]. Solaris [321, 554, 759, 779, 822, 957, 1122, 1152, 1323, 1872]. SolarWinds [4444, 4457]. sold [3831]. solstice [779]. Solution [485, 501, 526, 721, 747, 852, 891, 949, 990, 1065, 1068, 1127, 1128, 1141, 1159, 1184, 1201, 1203, 1204, 1236, 1289, 1307, 1320, 1356, 1417, 1427, 1474, 1597, 1660, 2756, 2959]. Solutions [538, 539, 559, 620, 921, 1085, 1115, 1140, 1410, 1655, 1924, 2015, 2154, 2863, 3585, 3651, 3981, 4032, 4273]. solve [3232, 3323]. solved [2734]. solves [435]. Solving [81, 973, 4164]. some [1228, 1744, 1834, 1863, 3107]. Someren [2153]. something [1743]. somewhat [1846]. son [3644]. song [1771]. Sons [149, 1178]. Sony [3643, 3653]. Soon [601]. sophisticated [1133, 2867, 4371]. Sophos [337, 1376]. Sorting [2017]. sorts [2153]. Source [1206, 1288, 1565, 1647, 1819, 1959, 2063, 2213, 2239, 2263, 2344, 2508, 2813, 3276, 3803, 3827, 3903, 4040, 4056, 4075, 4215, 4381, 4384, 4436]. sourcing [3621]. South [2957]. southwestern [175]. sovereignty [4430]. space [1981, 4452]. Spain [1286]. Spam [492, 803, 844, 901, 1568, 2156, 2167, 2177, 2500, 2627, 2704, 2713, 2894, 2918, 2993, 3089, 3624, 4239]. Spam-borne [2993]. Spamhaus [3432]. spammer [2692]. Spammers [2293, 2479, 2774, 2810, 3056]. spams [2187]. span [2376]. Spanair [3108]. spanner [1314]. spans [599]. sparekey [32]. spares [2363]. Sparks [749]. spate [1641]. spawns [3022]. speak [1276]. spear [3207]. spear-phishing [3207]. Spec [419]. Special [2974]. specific [706, 4416]. specification [827, 1290]. specifications [461, 1904, 2840]. speech [2456, 2641]. speed [72, 96, 667, 3043]. Speedy [913, 1049, 1364]. spend [2325]. spending [1694, 1878, 3282]. Spider [1683]. Spies [284, 1837, 2121, 3542, 3977]. spikes [2512]. spinning [1978]. spirit [2691]. splash [2370]. splits [335]. spoil [2303]. sponsored [3991, 3993]. spoofing [128, 183, 589, 2987]. spot [2169, 3691]. spotlight [3448]. spotlights [2348]. spots [4148]. Spraul [3403, 3761]. spread [3121, 3538]. Spreading [1170, 3377]. spreads [175, 328, 438, 1422]. spreadsheet [557, 3491]. Sprengers [3864]. Springer [3901]. spy [4117]. SpyEye [3033, 3168]. Spying [1716, 3200]. Spyware [1984, 2146, 2155, 2369, 2374, 2384, 2397, 2495, 2512]. SQL [1341, 2071, 2536, 2738, 2760, 2800, 3741]. squabbles [940]. squad [1059]. squatters [1264]. squeaky [1715]. squeaky-clean [1715]. squeezes [2242]. squishing [1683]. Srinivas [3210]. SS7 [3906]. SSD [3043, 3428]. SSH [925, 2169, 2326]. ssh-agent [925]. SSL [224, 1423, 1445, 2528, 2749, 2999, 3300, 3323, 4251]. SSO [1302]. stability [3203]. stack [1182, 2113]. stacks [963]. staff [3866, 4341]. staffers [1261]. Stage [263, 1746, 3490]. Stagefright [3727]. staging [996]. stake [177, 1384]. stakeholders [3344, 3345]. stalkers [881]. stampedes [1734]. stamps [624]. Standard [279, 291, 381, 514, 1052, 1071, 1183, 1210, 1272, 1293, 1303, 1508, 1651, 1699, 1788, 1909, 1956, 2335, 2542, 2910, 3018]. standardisation [2352]. standardization [2642]. Standards [406, 479, 1285, 1336, 1566, 2818]. Standing [1682]. stands [1758]. Star [2440]. Starch [3232, 3243, 3252, 3287, 3392, 3403, 3689, 3700, 3710, 3731, 3752, 3761, 3794, 3813, 3863, 3891, 3892, 3955, 3966]. start [1158, 1165, 2383, 2490, 2849, 4138]. start-up [1165]. starts [2438, 4377]. statd [908]. State [953, 2030, 2249, 2435, 2570, 2839, 3107, 3111, 3299, 3330, 3347, 3985, 3989– 3991, 3993, 3999, 4000, 4138, 4171, 4222, 4257, 4274, 4331, 4358, 4371, 4421, 4440, 4459, 4462]. State-backed [3111, 3985]. State-sponsored [3991, 3993]. States [175, 1250]. station [1601]. Stats [4326]. status [231, 3181]. Stay [1420, 1618, 2510, 3041, 4104, 4263]. Staying [2556, 3671]. steal [731, 2564]. Stealing [1425, 2529]. steals [246, 2147]. Stealth [2292]. stealthy [4178]. Steams [1676]. Steganography [2057, 2706]. STEM [3751]. Step [1113, 1167, 1558, 1833, 2214, 2335, 3271, 3425, 3924]. Stepping [3798]. Steps [2360, 2396, 2841, 3487, 4001, 4129]. Steve [3655]. Steven [3383, 3853]. stewardship [3499]. Stewart [3772]. sticking [2006]. sticks [2342, 2425]. sticky [2051]. stifles [2346]. still [585, 709, 977, 1009, 1208, 1354, 2270, 2356, 2583, 2609, 2684, 3192, 3306, 3463, 3665, 4062, 4465]. stimulate [450]. sting [1672, 2316]. Stock [1630]. Stock-market [1630]. stolen [1000, 1374, 3601]. Stop [232, 1304, 1380, 1636, 2203, 2313, 2519, 3691]. storage [2248, 2579, 2853, 4354]. stores [1407]. stories [1190]. storm [2407, 2888, 2934]. story [603, 1764]. Stowe [3881]. straight [2556, 3696]. stranded [1559]. Strategic [1882, 3094, 4066]. Strategies [799, 1224, 2421, 2474, 2916, 3135, 3419, 3595, 3659, 3751]. strategy [1207, 1257, 1350, 1993, 2037, 2451, 2558, 3175, 3271, 3279, 3478, 3879, 4017]. Streamlining [4083]. Streams [2588]. Street [3305]. streetwise [3417, 3523]. strength [874, 968, 1223, 1949, 3950, 4381]. strengthen [182, 3948]. strengthened [1361]. Strengthening [2821]. stretching [2598]. strike [724, 2800, 3056]. strikes [2010, 3001]. striking [3147, 4123, 4173]. stripping [1216]. Strong [540, 583, 714, 989, 1035, 1419, 1782, 4332, 4377]. Stronger [301, 3348]. structure [2540]. structured [3011]. struggle [2268, 3748, 3942]. Struts [4029]. students [2390, 3119]. Study [1083, 1258, 1420, 2548, 2617, 3266, 3368, 3414, 3824, 4049]. stuffing [4464]. stupid [2831]. Stuttard [3295]. Stuxnet [3111, 3121, 3644]. style [2373]. Subcommittee [1262]. subject [1923]. SubSeven [1957]. Subverting [3481]. succeeds [803]. Success [1190, 2561, 3539, 3751, 4430]. successful [1133, 3591, 4150]. such [3335]. Sudhanshu [3730]. sued [762, 1007]. sues [824, 1851]. suffer [258, 1402, 4384]. suffers [458, 2720, 3086, 3263, 3643]. suid exec [652]. Suidperl [534]. Suit [231, 1515, 3597]. suitable [1476]. Suite [914, 972]. summary [1193]. Summer [2165]. Summit [210]. sun [34, 322, 323, 349, 787, 961, 1122, 1152, 1998, 3407]. SunOS [782, 822]. SunSoft [456]. Super [1154]. superhighway [71, 197, 289]. Superior [1987, 2717]. Superman [552]. Supermarket [1852]. Supernet [360]. Supplier [2058]. supplies [134]. supply [3978, 4191, 4298, 4444]. Support [115, 157, 363, 679, 931, 966, 1047, 1391, 1404, 1801, 2335, 3564]. supports [701, 713]. Supreme [608]. surely [2086]. surface [2510, 4058]. surfing [615]. surge [2627]. surges [132]. surprise [2649]. Surveillance [260, 985, 1656, 2172, 3717, 3781]. Survey [489, 610, 634, 1099, 1301, 1348, 1399, 1420, 4049, 4063, 4087, 4197]. surveys [86]. Survival [2000]. survives [1760]. sus [2989]. susceptible [686]. suspect [1626]. Swede [846]. Sweden [193]. Swedish [275]. swine [2936, 2989]. Swiss [1626]. switch [1273, 2355, 2520, 4259]. switches [1096, 1937, 2093]. Switching [96, 491, 1092, 2061]. Sybex [3266]. Symantec [1363, 1953, 2084, 2124, 2361, 2373, 2446, 2666, 2776, 2810, 3051, 3293, 3411]. symbiosis [2567]. SYN [589, 602]. Syngress [3184, 3253, 3276, 3277, 3304, 3305, 3326, 3327, 3334, 3374, 3384, 3413, 3414, 3454, 3466, 3476, 3477, 3519, 3555, 3614, 3678, 3679, 3699, 3730, 3741, 3783, 3784, 3803, 3804, 3814, 3823, 3824, 3833, 3845, 3864, 3873, 3881, 3933, 3976]. Synopsis [1074, 4075]. Synopsys [4436]. Syrian [3442]. sysdef [862]. Sysid.exe [1483]. syslog [350, 436, 1119]. System [45, 76, 136, 142, 464, 565, 598, 619, 684, 745, 877, 902, 971, 985, 1001, 1031, 1096, 1216, 1247, 1250, 1319, 1364, 1411, 1574, 1618, 1679, 1939, 2023, 2328, 2491, 2701, 2868, 2915, 3081, 3378, 3648, 3952, 4355, 4423]. systematic [472]. Systems [312, 407, 416, 430, 522, 524, 571, 715, 734, 920, 1039, 1618, 1834, 1955, 1987, 2056, 2073, 2085, 2181, 2267, 2273, 2389, 2550, 2600, 2717, 2952, 3174, 3176, 3184, 3197, 3291, 3307, 3519, 3578, 3683, 3713, 3871, 3916, 4136, 4259]. Systemware [1984].

T [3454]. tables [511, 1647]. tabs [3711]. tackle [3317, 4107]. tackles [428, 2384]. Tackling [3211, 3650, 4033, 4473]. Tactics [1940, 4297, 4340, 4389, 4400]. tail [181]. Taiwan [71]. Taiwanese [635]. Take [5, 582, 601, 858, 1749, 2557, 3432]. take-off [5]. takedown [3017]. takedowns [4336]. taken [3129, 3494, 3682, 3973, 4349]. takeover [1923]. takes [263, 745, 787, 2649, 2713, 2888, 2898, 2934, 3194, 3442, 4086, 4325]. Taking [1314, 3201, 3318, 3774, 3875]. Tales [1557]. talk [4048]. talkd [675]. Tallinn [3444]. Tangled [1557, 3315]. Tankard [3458, 3514]. Tap [394, 1721]. Tap-proof [1721]. Tapping [1593, 2681]. target [1387, 1922, 1939, 2790, 2968, 3324, 3764, 3810, 3985, 4174, 4358]. Targeted [523, 1484, 2471, 2746, 3218, 3516, 3584, 3888, 4158, 4309]. targeting [3607]. Targets [1594, 3207, 3264, 3676, 3935, 4278, 4290, 4467]. Tasini [922]. task [3157]. Taskforce [1994]. tax [3125]. taxes [1252, 3420]. Taylor [3892]. TCP [139, 563, 589, 769, 1123, 1181, 1182, 1466, 2486]. TCP/IP [139, 563, 769, 1182, 2486]. teach [2085, 3336]. Team [572, 770, 1219, 1710, 1806, 3545, 3717, 3882]. teaming [955, 3814, 4059]. Tech [59, 1658, 1800, 2712, 3305, 3758, 4107, 4232]. Technical [1011, 1279, 1482, 1641, 2261, 4029, 4395]. Techniques [1041, 1703, 1834, 2057, 2672, 3156, 4389, 4390]. Technological [2873]. Technologies [468, 785, 2894, 3853]. Technology [3, 358, 505, 636, 1163, 1272, 1434, 1454, 1499, 1854, 1906, 2346, 2589, 2616, 2718, 2749, 2871, 3235, 3433, 3481, 3502, 3630, 3711, 4274, 4307]. Teenage [1373]. teenager [2223]. teens [1590]. teeth [2156]. Telco [2044]. Telecom [828, 953, 1918, 2385]. Telecommunications [1195, 1200, 1262]. telecommuting [345]. Telecomputing [345]. telecoms [70, 160, 753]. Telekom [75]. telemetry [3416]. telephone [393, 818]. telephony [1115, 2600, 4136]. Teleworkers [41, 2622]. teleworking [4366]. Telnet [151, 351, 734, 943]. temporary [802]. Ten [1418, 3104]. tenacious [2346]. tenanted [3059]. Tenneti [3210]. tensions [3582]. Terence [3344, 3345]. term [672, 3666, 3708]. Terminal [1246]. terminology [4284]. terrain [1917]. territory [2374]. terror [1849]. Terrorism [205, 1751, 1801, 1895, 1932, 3223]. terrorist [2436]. terrorists [1625, 1837, 2241]. Terry [3305, 3988]. Test [72, 694, 2018, 2647, 3288, 3435]. tested [927]. Tester [3243, 3276]. Testing [472, 688, 726, 817, 857, 1621, 1970, 2016, 2058, 2086, 2277, 2281, 2354, 2662, 3277, 3418, 3435, 3509, 3555, 3586, 3594, 3604, 3606, 3637, 3646, 3657, 3838, 4016, 4082, 4139]. tests [2277, 2756, 3769]. text [2250]. TFN2K [1334]. Theft [59, 1644, 2506, 2575, 2761, 2981, 3768, 4035]. their [1641, 2193, 3427, 3573, 3723, 3884, 3941, 4259, 4463]. Them [723, 975, 1703, 1889, 2517, 3236, 4020, 4343]. Theory [1702]. there [22, 1240, 1926, 2235, 2284, 2364, 2630, 3136, 3329, 3596, 4247]. thief [753, 1136]. Thiel [3794]. thieves [241, 2193]. Thin [3060, 3166, 3215]. Things [1420, 2043, 3602, 3684, 3745, 3869, 3888, 3894, 3969, 3970, 4002, 4027, 4040, 4141, 4154, 4173, 4321]. think [3403, 3707, 3934]. Thinking [1802, 2580, 3166]. thinks [1274, 2010]. Third [919, 1237, 1436, 3069, 3276, 3334, 3374, 3658, 3824, 4008]. third-party [919]. thirds [3820, 3861]. Thomas [3555, 3699, 3988]. Thompson [2373]. Thomson [799]. Thor [3304]. Thorny [1093, 1532]. though [1679]. thousands [3842]. Threat [128, 145, 977, 1029, 1065, 1469, 1641, 1681, 1804, 2014, 2027, 2083, 2203, 2233, 2234, 2480, 2516, 2545, 2598, 2784, 2785, 2806, 2895, 2939, 2956, 3057, 3118, 3122, 3148, 3154, 3192, 3224, 3245, 3250, 3313, 3335, 3396, 3534, 3546, 3599, 3620, 3633, 3648, 3690, 3692, 3736, 3773, 3876, 3883, 3927, 3938, 3948, 3970, 3976, 3980, 4050, 4058, 4099, 4120, 4146, 4221, 4227, 4261, 4293, 4341, 4367, 4416, 4469]. threatened [906, 1405, 3460]. threatening [153, 456, 4216]. threatens [2398, 3727, 4301, 4444]. Threats [887, 1704, 1893, 1896, 2154, 2298, 2315, 2339, 2430, 2441, 2443, 2487, 2497, 2509, 2530, 2709, 2862, 3045, 3070, 3211, 3235– 3237, 3278, 3317, 3368, 3510, 3552, 3557, 3567, 3595, 3647, 3650, 3683, 3701, 3754, 3821, 3833, 3851, 3858, 3957, 3960, 3986, 3990, 3993, 4023, 4098, 4110, 4316, 4328, 4351, 4391, 4427, 4441]. Threatwatch [4097, 4108, 4118, 4131, 4144, 4156, 4169, 4183, 4195, 4208, 4220, 4231, 4243, 4255, 4269, 4280, 4291, 4302, 4314, 4324, 4335, 4350, 4360, 4373, 4385, 4399, 4411, 4422, 4434, 4445, 4458, 4468]. Three [335, 1583, 1709, 1720, 1953, 2906, 3001, 3659]. three-strikes [3001]. throughout [4079]. ticks [2401]. tier [744, 831]. tiered [2613].

Ties [325]. tighten [2271, 3098, 3573, 4267]. Tightening [657]. Tighter [1398]. TikTok [4303]. Tilman [3863]. Tim [981]. Time [20, 262, 295, 413, 1073, 1461, 1745, 1749, 1773, 1892, 2633, 2822, 2849, 2940, 2989, 3125, 3288, 3491, 3510, 3649, 3701, 3714, 3981, 4041, 4306, 4453, 4465]. Time-Based [1073, 1461]. Timeline [3868]. Timely [1677]. times [187, 2281, 2558, 3256, 3300, 3842]. TimeStep [1126]. Timing [2004]. Timothy [3304]. Tin [1479]. Tips [2313, 2400, 2906]. TIS [701, 724, 946, 1822]. Tmobile [2597]. Toad [1639]. Tobias [3287]. today [3135, 3137, 4136]. together [1032, 1112, 2338, 4124, 4256]. token [103, 644, 2477]. tokenisation [3084]. tokenless [3609]. Tokens [295, 3010, 3206, 3714]. told [1773, 3098]. tolerance [376]. tolerant [139, 201]. toll [2421]. tolls [707]. tomorrow [3137, 3736]. too [229, 266, 723, 1848, 1850, 2023]. Tool [254, 425, 512, 934, 1142, 1324, 1475, 2330, 2639, 2727, 2878, 2895, 2944, 3013, 4342]. toolbox [2738]. Toolkit [909, 3276, 3334, 3636, 4226]. toolkits [1220]. Tools [1278, 1834, 2101, 2132, 2151, 2218, 2322, 2432, 2529, 2926, 2938, 3528, 4390]. Top [181, 1301, 1330, 2036, 2081, 2169, 2172, 2187, 2195, 2319, 2517, 2583, 2786, 3968, 4368]. Top-down [2195]. Tops [1613, 1700, 2201, 2791, 3495]. TOR [459, 3473, 3654]. Toronto [92]. Toshiba [1855]. Total [539, 1163]. touch [4346]. ‘TouchLock’ [259]. Tough [856, 1834, 3300]. toughen [1082]. Towers [1791]. town [2500]. trace [2940]. trace-back [2940]. tracing [2697]. track [3895]. trackable [1559]. tracker [212]. tracking [2171, 2175, 2398, 2697, 3711]. tracks [2666]. trade [110]. trading [138, 1413]. traditional [1559, 2496, 3406, 4453]. Traffic [252, 1380, 1559, 1910, 2088, 2582, 3226, 3289, 3834, 4251]. trail [1059]. trailing [2858]. trails [3587]. training [4071]. Transaction [620]. Transactions [102, 419, 853, 1359, 1615, 2939]. transfer [390, 1559, 2333, 3490]. transfers [1452]. transformation [3979, 4134, 4170, 4418]. transient [3672]. Transition [468, 1956]. transmissions [850]. transparency [4381]. transportation [4149]. trap [324, 2564]. trapper [596]. travails [707]. travel [794]. travelled [312]. Treading [4286]. Treating [2806]. treatment [1206]. treatments [2315]. Treats [1582]. treaty [1816]. trek [2440]. Trend [772, 1569, 2271, 4110, 4362]. Trends [1978, 2188, 2207, 2471, 2558, 3057, 3068, 3160, 3224, 3766]. trial [261, 1758]. trials [1650]. trick [2479]. Tricks [110, 1582, 1958, 3135]. trillion [1468]. triple [220, 333]. Tripwire [1287]. Tristrata [1111]. Tritheim [1106]. Triton [4130]. TROJ XPLOREZIP [1216]. Trojan [271, 738, 771, 1100, 1123, 1442, 1485, 1755, 1869, 2190, 2564, 2755, 3179]. trojans [2013, 2397, 2566, 2594, 2746, 3171, 4050]. Trouble [1313, 1365, 1540, 2044, 2667]. Troubled [1638]. Tru64 [1214, 1967]. Tru64/DIGITAL [1214]. True [1834, 2327, 3724, 4059]. TrueCrypt [3581]. TruSecure [2202]. Trust [1295, 1318, 1531, 1572, 1596, 1743, 1763, 1797, 1884, 1889, 1902, 2306, 2454, 2949, 3426, 3480, 4047, 4213, 4462]. Trusted [1635, 2301, 2384, 2569, 2821, 2871]. Trusting [2870, 4369]. TrustWorks [1972]. trustworthy [1902, 3323, 4127]. Truth [958, 3896, 4037]. TSB [2477]. Tsunami [2363]. tuning [2513]. tunnelling [4297]. tunnels [1433]. Turk [2458]. turn [2774, 2854, 3219]. turned [975, 1726]. Turning [2769]. Turnkey [1141]. turns [449, 1647, 2070, 3762]. TV [615, 1601]. TVD [892]. Twice [1535, 1795, 2010, 3085, 4043]. Twinkle [1222]. Twitter [3073, 3098]. Two [105, 1365, 1410, 1453, 1836, 1857, 2543, 2701, 3188, 3238, 3609, 3820, 3861, 3973, 4154]. Two-factor [1410, 2543, 2701, 3188, 3609]. Two-thirds [3820, 3861]. type [497, 1744]. types [4416]. typical [1023]. Tyrone [3720].

U.S [1420]. UDP [437, 527]. UEFI [4452]. UK [7, 66, 69, 412, 478, 765, 928, 945, 1072, 1082, 1158, 1185, 1193, 1271, 1420, 1429, 1451, 1559, 1625, 1650, 1673, 1905, 2085, 2280, 2306, 2424, 2533, 2636, 2668, 2676, 2743, 2754, 2779, 2814, 3728, 3758, 3768, 3820, 3841, 3860, 3879, 3977, 4017, 4044, 4052, 4155, 4207, 4219, 4266, 4328, 4398, 4433]. Ukraine [3885, 4005, 4143]. Ultimo [1496]. unable [1874]. unacceptable [408]. unauthorized [162]. uncertain [3181]. uncertainty [2646, 2651]. Uncle [954]. Uncovered [3210]. Uncovering [4345]. uncovers [770, 3195]. uncrackable [2621]. under-performing [97, 1686]. Undercover [324, 4162]. Underground [238, 1582, 4362]. undermines [2042]. underpins [4356]. understaffed [1686]. Understanding [1480, 1911, 2963, 3280, 3407, 3525, 3620, 3693, 3754, 3891, 3901, 3911, 3999, 4427]. understood [3268]. undetected [230]. Unicode [2393]. uniform [10]. Unify [2580]. unique [4407]. uniquely [3071]. Unisys [1610]. unit [422, 3423, 4433]. unite [290, 316, 969, 1428, 1473]. United [175, 1250, 4312]. universities [2245, 4266]. University [1096, 1543, 2112, 3444, 3911, 3922, 4009]. UNIX [404, 484, 687, 744, 899, 1048, 1155, 1214, 1287, 1322, 1853]. unknown [3650]. unlawful [1192]. unleash [2283]. unleashed [1538, 2319]. unmasking [753]. unopened [1422]. Unpatched [2192]. unpredictability [980]. unprepared [4052]. unprotected [1888]. Unravelling [2153]. Unsafe [1375]. unsecured [3704]. Unseen [4110]. unsniffable [3416]. unsolvable [81]. unsolved [1147]. Unto [3373]. Untrusted [1079, 3246]. untrustworthy [2410]. unveil [2945]. unveiled [1499]. Unveiling [3569]. unveils [9, 222, 796, 815, 1478, 2413, 2778]. unwanted [932, 1157]. Update [68, 1291, 1675, 1762, 1843, 1881, 2285, 2590]. updated [64, 926]. updates [2283, 2297, 4196]. upgrade [340, 1863, 2280]. Upgrades [899, 2414, 2499]. UPnP [1841]. ups [1996, 4453]. upset [26]. Upstream [4304]. Upturn [2594]. Urge [1616]. urged [1455]. urges [660, 868]. URL [1228, 3133]. URLs [2393]. USA [269, 1170, 1196, 1482, 2009]. Usability [2526, 3147]. USB [2068, 2342, 2425, 3138]. Use [82, 112, 378, 910, 1436, 1462, 1744, 1813, 2218, 2767, 3213, 3967, 4228]. used [738, 1738, 3043, 3273, 3612, 3738, 4415]. UseLogin [1465]. Usenet [560]. User [1, 330, 385, 1030, 1156, 1312, 1502, 1660, 1876, 2541, 2629, 2641, 2852, 3093, 3134, 3147, 3693, 3831, 4101, 4115]. Users [17, 26, 119, 157, 198, 288, 405, 420, 616, 638, 712, 786, 818, 1268, 1376, 1632, 1874, 2313, 2340, 2428, 2441, 2637, 3071, 3141, 3237, 3670, 3887, 3941, 4295]. uses [717, 808, 834, 2564]. Using [387, 696, 840, 883, 1041, 2326, 2548, 2641, 2810, 2821, 2832, 3053, 3105, 3232, 3460, 3526, 3557, 3746, 3765, 3893, 4039, 4174, 4239, 4355, 4471]. usual [3452]. usually [705]. Utilities [3192]. utility [666, 1001]. UTM [2519, 3405]. UX [395, 555, 1027, 1120, 1172, 1198].

V [1668, 2412, 3403, 3761]. V.5 [298]. v11 [4459]. V6 [2209, 2321]. V7.1 [1028]. vacation [961]. Vacca [4008]. Valentine [1616, 2514]. Valerie [3699]. valid [2541]. validated [2962]. validation [1183, 1445, 4295, 4441]. Value [2142, 2336]. valued [2518]. Vampire [1703]. Vandalizes [1740]. Vanishing [1556]. variant [2148]. Variants [2141, 2346]. Variations [2131]. Vasileios [3844]. Vaulting [207]. VDI [3047]. vector [2578, 3588]. vectors [3146, 3448]. Vegas [2969]. vehicle [1559]. veil [2975]. Vendor [623, 2418, 2487]. Vendors [365, 625, 1275, 1387, 1616, 1853, 1981, 2408, 2606, 2953, 3797]. Ventana [819]. venture [75, 481, 1011]. ventures [3077].

Veracode [4459]. verdicts [2657]. verge [117]. verification [279, 722, 818, 2641]. verifier [475]. Verifying [3404]. VeriSign [1643, 1904]. Veritas [2361]. Verizon [2825]. Version [228, 429, 435, 470, 544, 1123]. versions [1354, 4218]. Versus [1575, 1912, 2195, 3147]. very [3479, 4009]. veteran [2575]. via [400, 560, 654, 821, 895, 1110, 1502, 2760, 3244, 3768]. Viagra [2293]. vicious [397]. Victim [1720, 3035, 3337]. victims [3768, 3889, 3909]. video [1137, 2222, 4300]. View [883, 1393, 2465, 2487, 2587, 2602]. Vigilantes [1670, 1989]. Vim [4089]. violations [1692]. Violent [3454]. Vipul [2718]. Viral [1875]. Virtual [285, 399, 543, 587, 692, 719, 816, 851, 1298, 1458, 1989, 2538, 2554, 2793, 2830, 2841, 2863, 2928, 3020, 3396, 4465]. Virtualisation [2727, 2987, 3074, 3319]. virtualised [2976]. Virus [40, 68, 76, 81, 119, 174, 273, 276, 285, 329, 337, 405, 426, 438, 497, 526, 551, 557, 562, 571, 585, 591, 592, 596, 598, 609, 642, 650, 661, 670, 683, 684, 712, 721, 742, 746, 772, 874, 969, 976, 1103, 1135, 1156, 1201, 1388, 1421, 1422, 1432, 1434, 1448, 1449, 1454, 1484, 1498, 1516, 1517, 1616, 1620, 1653, 1696, 1708, 1759, 1870, 1887, 1939, 1940, 1975, 2056, 2112, 2162, 2165, 2171, 2201, 2207, 2223, 2288, 2374, 2490, 2595, 3131, 3390]. Viruses [167, 299, 382, 397, 402, 439, 448, 488, 527, 535, 581, 610, 914, 926, 1447, 1468, 1480, 1506, 1564, 2346, 2384, 2432, 2578, 2958]. Virusland [1971]. Virusnet [690]. VirusPatrol [712]. Visa [291, 461, 1800]. visibility [3535, 4284]. vision [3549]. Visitor [1413]. Vista [2587, 2606, 2643, 2657]. VistaNet [894]. Visual [4238]. visuals [2330]. Vital [1947, 2337, 3926, 4391]. VLSI [1107]. VMware [2724]. vogue [2488]. Voice [169, 818, 1115, 1237, 1525, 2379, 3506, 3647, 3704, 4310]. voicemail [1907]. void [2670]. VOIP [2014, 2443, 2459, 2530, 2539, 2581, 2592, 2668, 2826, 3309, 4136]. Volatile [2820]. volrmmount [944]. Volume [2432, 3393]. voracious [1527]. voting [2389, 2676, 2739]. VPN [872, 890, 915, 1060, 1126, 1141, 1202, 1204, 1207, 1234, 1356, 1362, 1414, 1433, 1500, 2099, 2311, 2749, 2999, 3338]. VPN-1 [1141]. VPNs [365, 1185, 1597, 3408]. vs [1989, 2016, 2263, 2577, 2749, 2820, 2988, 3084, 3428, 3894, 4226]. vulnerabilites [2081]. Vulnerabilities [152, 172, 173, 351, 555, 590, 653, 732, 779, 806, 823, 842, 863, 882, 884, 924, 983, 984, 999, 1012, 1044, 1057, 1097, 1120, 1122, 1152, 1155, 1172, 1181, 1212, 1215, 1230, 1241, 1901, 1911, 1969, 1997, 2035, 2050, 2106, 2114, 2144, 2192, 2231, 2412, 2491, 2886, 3005, 3088, 3533, 3686, 3795, 3903, 4102, 4104, 4113, 4122, 4142, 4157, 4174, 4199, 4323, 4466]. Vulnerability [151, 236, 298, 321, 323, 349, 350, 373, 392, 395, 411, 436, 476, 512, 532– 534, 554, 556, 569, 606, 630– 633, 652, 672, 675, 694, 698, 710, 729, 730, 757– 759, 780, 782, 800– 802, 822, 862, 864, 867, 908, 925, 944, 957, 961– 963, 1001, 1028, 1045, 1080, 1081, 1118, 1121, 1153, 1199, 1214, 1229, 1245–1247, 1267, 1282– 1284, 1298, 1351, 1381, 1700, 1911, 1930, 1987, 2059, 2126, 2132, 2158, 2168, 2555, 2807, 2819, 2998, 3124, 3815, 4326]. vulnerability [1055, 1070, 1212, 1322–1324, 1377, 1378, 1426, 1445, 1446, 1463, 1466, 1490– 1492, 1712, 1871, 1883, 1902, 1917, 1996, 2024, 2202, 2212, 2275, 2405, 2431, 2497, 2703, 2905, 3182, 3568, 3727, 3925, 4062, 4379]. vulnerability-management [4379]. Vulnerable [614, 667, 1002, 1038, 1654, 1853, 2097, 2356, 3563, 3564, 3778, 3856, 4209, 4218, 4244]. Vyacheslav [3435].

W [1619]. W2K [1666]. W2K.Stream [1516]. W97M.Footprint [1135]. WabiSabiLabi [2703]. WAFs [3957]. wages [2320]. wait [2212, 3279]. wake [1849]. Waledac [3169]. walk [4344]. Wall [442, 467, 3082]. wallet [1165]. walls [2991, 4284]. WAN [108, 1651, 3904, 3968, 4124]. wanna [2390, 3971]. wanna-B [2390]. WannaCry [4025, 4061]. Want [1546, 1617, 1824, 1838, 2712]. wanting [484]. wants [1493, 1867]. Wapping [552]. War [170, 265, 901, 1751, 1791, 1812, 2079, 2448, 2575, 3112, 3227, 4138, 4230, 4247]. Warchalking [2204]. ward [311]. warehouse [1852]. Warfare [838, 859, 880, 1259, 3316, 3444, 3497]. warming [4001]. warn [3262]. Warning [497, 687, 2386, 4268, 4386]. warns [503, 960, 1545, 3459, 3589, 4155, 4266, 4348]. warrant [733]. warring [3285]. wars [3056]. was [1311, 3653]. washer [1644]. Washing [1579]. watch [1641, 1782, 3773]. WatchGuard [3535, 4338]. Watching [1189, 1742, 1810, 2351, 2367]. water [4286, 4456]. watering [3422]. Waters [1638]. Watters [3803]. Wave [883, 1895, 2359, 2474, 4290]. way [495, 1231, 1311, 2187, 2624, 3164, 3503, 4069, 4284]. Wayfarer [449]. Ways [2, 407, 430, 902, 2135, 3358, 4260]. Weak [1486, 3375, 3728, 4148]. weakened [985]. weakest [3463, 4392]. weakness [1374, 3780, 4300]. Weaknesses [475, 1005, 1066, 3906, 3919]. wealthy [2968]. Weaponised [3213]. Weaponising [3980, 4002]. weapons [25, 4320]. wearable [3630]. weather [248]. Web [161, 192, 201, 381, 407, 430, 559, 584, 615, 656, 668, 718, 720, 733, 772, 788, 796, 812, 897, 947, 949, 1002, 1018, 1034, 1047, 1084, 1184, 1191, 1192, 1218, 1245, 1305, 1340, 1373, 1383, 1464, 1474, 1547, 1552, 1633, 1649, 1743, 1744, 1835, 1846, 1904, 1970, 1998, 2029, 2036, 2121, 2144, 2154, 2206, 2251, 2547, 2698, 2720, 2737, 2765, 2771, 2796, 2834, 2855, 2870, 2934, 2992, 2995, 3005, 3039, 3117, 3176, 3238, 3246, 3250, 3295, 3315, 3384, 3463, 3466, 3476, 3477, 3553, 3569, 3730, 3769, 3973, 4019, 4046, 4359]. Web [716, 826, 881, 1310, 1557, 1725, 1765, 2102, 2262, 2429, 2527, 2536, 2570, 2660, 2729, 3585, 3795, 4214]. web-based [897, 947, 1018, 1547, 2995, 3246]. Web-surfing [615]. Web/E [720]. Web/E-mail [720]. Webcam [3907]. WebCom [655]. webdist.cgi [823]. WebGlimpse [884]. Webjacking [2591]. website [2781, 2925, 2980, 3783]. websites [1697, 2079, 2760, 3563, 3607, 3612, 3739]. WEF [1626]. Weidenfeld [3977]. Welcome [2001, 2507]. Well [552, 1786, 3148]. well-meaning [3148]. went [2158, 2161, 2385]. WEP [1737, 2011]. were [4465]. Wesley [1763]. Western [3591]. Wheeler [3253]. WheelGroup [781]. Where [669, 816, 1392, 1659, 1744, 2223, 2302, 3618, 3815, 4441]. whether [1436, 1704, 3653]. while [2797, 3871]. White [101, 868, 1711, 2016, 3225]. Whitehouse [286, 1478, 1993, 3720]. whitelisting [2961]. Who [1211, 1479, 1810, 1982, 2630, 3136, 3271, 3874, 3994, 4461]. whole [603]. Whom [707, 1053, 1786, 2421, 3400]. whose [2822, 4112]. Wicked [3892]. Wide [192, 522, 3778, 4155, 4189]. widely [1170]. widening [2653]. widens [2376]. wider [2184, 3909]. widespread [3022]. Wife [1632]. Wifi [2612, 3036, 3064, 3081, 3152, 3164, 4006, 4189]. wild [438, 1755, 3046, 4409]. Wiles [3305]. Wiley [149, 1178, 1658, 3295, 3720]. Wilhelm [3555]. Will [601, 1392, 1393, 1401, 2303, 2348, 2384, 2395, 2436, 3721, 3855, 3976, 3983]. Williams [3374]. willing [3792]. Wilson [3425]. win [2389]. Win2000 [2340]. Winamp [2356]. window [1045, 2904, 2905]. Windows [226, 285, 337, 401, 425, 487, 577, 646, 791, 800, 876, 942, 956, 989, 1124, 1230, 1246, 1282, 1320, 1327, 1398, 1653, 1931, 1942, 1995, 2095, 2205, 2273, 2550, 2587, 2621, 2860, 3078, 3184, 3185, 3334, 3425, 3479, 3519, 3564, 3636, 4218, 4315, 4442]. Winning [4297]. Wins [1542, 2574, 3112, 3227]. Winword [439]. wipe [1731, 2529]. WIPO [1264]. wire [2408]. wired [2731]. Wireless [5, 1249, 1289, 1299, 1338, 1361, 1414, 1470, 1472, 1580, 1592, 1615, 1651, 1654, 1705, 1807, 1859, 1880, 1888, 1906, 1909, 1921, 1926, 1943, 1963, 2043, 2069, 2113, 2127, 2139, 2151, 2166, 2175, 2188, 2246, 2272, 2292, 2294, 2297, 2431, 2526, 2581, 2597, 2609, 2688, 2699, 2731, 2940, 2991, 3081, 3257, 3270, 3329, 3789]. Wireless-based [2175]. Wireshark [3232]. wiretapping [64, 375]. wise [2637]. WISekey [1497]. wising [2597]. within [1704, 2337, 2956, 2964, 3489, 3557, 3585, 3897]. without [579, 680, 1447, 2021, 2072, 3076, 3163, 3660, 3935]. Wizard [277]. WLAN [2043]. WLANs [2292]. wmcdplay [1379]. Wnn [1492]. Woe [380]. Wolf [1766, 1824]. Women [1814, 3505, 3751]. won [3451]. wonder [3612]. won’t [1205]. Word [328, 488, 497, 1329, 3947]. WordPress [3061]. Work [1023, 1112, 1275, 1393, 1850, 2275, 2476, 2802, 3111, 3628, 3789, 3828]. workaround [350]. Worker [1349, 2620, 2835, 3408]. workers [42, 4308]. workflows [4284]. workgroups [2248]. Working [429, 737, 1548, 2355, 2640, 3310, 3406, 3532, 4256, 4419, 4462]. Workplace [1613, 1812, 3259]. Works [3761]. World [192, 650, 670, 869, 1189, 1838, 1978, 2072, 2402, 2429, 2563, 2736, 2804, 2904, 2936, 2976, 3232, 3336, 3408, 3911, 4225]. worlds [2793, 3212]. WorldView [1492]. Worldwide [663, 714, 764, 940, 1927, 2194, 3528, 4181, 4444]. Worm [1372, 1483, 1671, 1693, 1879, 1999, 2049, 2071, 2085, 2136, 2141, 2164, 2382, 2492, 2649, 2888]. wormable [4218]. Worms [1747, 1811, 1941, 2165, 2348, 2578, 2688]. worried [1589, 2766, 4331]. worries [277, 280, 405, 1586, 2583, 3074]. Worrying [2849, 4323]. Worse [1620, 2164, 2673, 4188]. Worst [1460, 3786, 3948]. worth [2484]. Would [755]. WPA [3013]. wrangles [940]. wrap [823, 3055]. Wrappers [1123]. wrecking [2217]. writer [2165]. writers [1484, 2112, 2148]. Writing [2682]. Written [1767]. wrong [2158, 2161]. WSLS [2771]. WSUS [4442]. wu [1268]. wu-ftpd [1268]. WWW [885].

X [149, 392, 1180, 1678, 2419, 2516, 2849, 3179, 3771, 4469]. X-commerce [1678]. X-Force [4469]. X.400 [598]. X.509 [1202]. X25 [597]. Xinetica [1550]. XML [1764]. XP [1804, 3561, 3564]. XPSP1 [2340].

Y2K [921, 1091, 1190, 1205]. Yahoo

[865, 906, 1434, 3538]. Year [274, 276, 1599, 1836, 1857, 2078, 2207, 2350, 2412, 3140, 3899, 3996, 4043, 4061, 4237, 4245, 4293, 4407]. Years [1419, 2562, 3920]. yes [2140]. yesterday [3151]. York [2357]. you’ll [4450]. Young [781, 2458]. You’re [1976]. yourself [881, 2472]. youths [2255]. YouTube [2694]. Yves [4254].

Zakas [3891]. zap [2035]. Zaxus [1427]. Zealand [1303]. Zergo [965]. Zero [1969, 2236, 2473, 3224, 3920, 4047, 4462]. Zerodays [3920]. Zerologon [4409]. Zeus [3109, 3130, 3168, 3171]. ZigBee [4032]. zombie [2640]. zombies [1739, 2571]. Zoombombing [4393]. Zotob [2458, 2459].

References

Steffora:1994:UEC

[1] Ann Steffora. User education critical to effective network security. Network Security, 1994(6):2, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                              //www.sciencedirect.com/science/
                                              article/pii/1353485894900280.

Steffora:1994:ICM

[2] Ann Steffora. Internet community has many ways to handle security infringements. Network Security, 1994(6):3, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900299.

Anonymous:1994:FFC

[3] Anonymous. Flaw found in Clipper technology. Network Security, 1994(6):3–4, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900302.

Anonymous:1994:J

[4] Anonymous. Junkie. Network Security, 1994(6):4, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900310.

Anonymous:1994:WDS

[5] Anonymous. Wireless data services take-off. Network Security, 1994(6):4, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900329.

Anonymous:1994:MDC

[6] Anonymous. Mobile data comms market sees rapid growth. Network Security, 1994(6):4–5, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900337.

Anonymous:1994:UIP

[7] Anonymous. UK invests in photonics. Network Security, 1994(6):5, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485894900345.

Anonymous:1994:NDF

[8] Anonymous. Nynex develops fibreoptic infrastracture. Network Security, 1994(6):5, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900353.

Anonymous:1994:PUP

[9] Anonymous. Proginet unveils plan to link IBM mainframes to Netware. Network Security, 1994(6):5, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485894900361.

Anonymous:1994:ULP

[10] Anonymous. US legislature pushes for uniform online health privacy rules. Network Security, 1994 (6):5–6, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/135348589490037X.

Anonymous:1994:ERB

[11] Anonymous. Encryption restrictions bind manufacturer’s hands. Network Security, 1994(6):6, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894900388.

Anonymous:1994:EC

[12] Anonymous. Electronic cash. Network Security, 1994(6):6, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894900396.

Anonymous:1994:CC

[13] Anonymous. CyberSAFE Challenger 5.2.1. Network Security, 1994(6):7, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/135348589490040X.

Anonymous:1994:SSS

[14] Anonymous. SDI secures Shiva LANs. Network Security, 1994(6):7, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894900418.

Anonymous:1994:SM

[15] Anonymous. Security manager. Network Security, 1994(6):7, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894900426.

Anonymous:1994:INS

[16] Anonymous. Increased NetSP security. Network Security, 1994(6):7, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894900434.

Michaels:1994:LUL

[17] Steven Michaels. LAN users are lax on security. Network Security, 1994(6):8, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900442.

Riggs:1994:COC

[18] Brian Riggs. Cracking down on online crime. Network Security, 1994(6):8, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900450.

Highland:1994:HSD

[19] Harold Joseph Highland. How secure is data over the Internet? Network Security, 1994(6):9–11, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900469.

Peterson:1994:FMT

[20] Padgett Peterson. The fortress mentality: Its time has come again. Network Security, 1994 (6):12–13, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900477.

Hancock:1994:IPS

[21] William M. Hancock. Issues and problems in secure remote access. Network Security, 1994(6):14–18, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894900485.

Lindup:1994:TRH

[22] Ken Lindup. Is there a role for hackers in polite society? Network Security, 1994(6):19, June 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900493.

Armstrong:1994:PNS

[23] Lisa Armstrong. Are public networks safe? Network Security, 1994(7):2, July 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900019.

Snell:1994:PIN

[24] Monica Snell. Protecting the internal network. Network Security, 1994(7):2, July 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900027.

English:1994:CNI

[25] Erin English. Cache of nude images found at nuclear weapons lab. Network Security, 1994(7):2–3, July 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894900035.

Snell:1994:UUI

[26] Monica Snell. Users upset over Internet ads. Network Security, 1994(7):3, July 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900043.

Anonymous:1994:CCD

[27] Anonymous. Changes to come for DCS. Network Security, 1994(7):3–4, July 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485894900051.

Anonymous:1994:LMO

[28] Anonymous. LAN mail to overtake host mail. Network Security, 1994(7):4, July 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/135348589490006X.

Anonymous:1994:NPS

[29] Anonymous. Novell and PC security collaborate. Network Security, 1994(7):4, July 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900078.

Anonymous:1994:CAM

[30] Anonymous. Communications available to more countries. Network Security, 1994(7):4, July 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900086.

Snell:1994:LNI

[31] Monica Snell. Lotus notes the Internet advantage. Network Security, 1994(7):4–5, July 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900094.

Armstrong:1994:SN

[32] Lisa Armstrong. A sparekey to Netware. Network Security, 1994(7):5, July 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894900108.

Lawrence:1994:SCS

[33] L. G. Lawrence. Security in a client server environment. Network Security, 1994(7):5–15, July 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900116.

Gordon:1994:SSH

[34] S. Gordon and I. Nedelchev. Sniffing in the sun: History of a disaster. Network Security, 1994 (7):16–19, July 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900124.

Anonymous:1994:Ea

[35] Anonymous. Events. Network Security, 1994(7):20, July 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900132.

Anonymous:1994:UCC

[36] Anonymous. US companies count Internet security options. Network Security, 1994(8):2, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900736.

Anonymous:1994:SCI

[37] Anonymous. Security for commerce on the Internet. Network Security, 1994(8):2, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900744.

Anonymous:1994:RCS

[38] Anonymous. Ring of crime in Scotland. Network Security, 1994 (8):2–3, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900752.

Anonymous:1994:CMI

[39] Anonymous. CIA to mine Internet. Network Security, 1994 (8):3, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900760.

Anonymous:1994:IVC

[40] Anonymous. IBM’s virus collection grows. Network Security, 1994(8):3, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900779.

Anonymous:1994:RNT

[41] Anonymous. Rise in number of teleworkers. Network Security, 1994 (8):4, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900787.

Anonymous:1994:HWD

[42] Anonymous. Home workers demand ISDN. Network Security, 1994(8):4, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900795.

Anonymous:1994:MPN

[43] Anonymous. Mobile phone network promoted. Network Security, 1994(8):4, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900809.

Anonymous:1994:HRO

[44] Anonymous. High-rise optic links. Network Security, 1994(8):4, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894900817.

Anonymous:1994:FRS

[45] Anonymous. Fund raising for satellite phone system. Network Security, 1994(8):4–5, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900825.

Anonymous:1994:IRL

[46] Anonymous. Improved remote LAN access. Network Security, 1994 (8):5, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900833.

Anonymous:1994:SAC

[47] Anonymous. Secure access for councillors. Network Security, 1994 (8):5, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900841.

Anonymous:1994:RLO

[48] Anonymous. Raxco launches omniguard. Network Security, 1994(8):5, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/135348589490085X.

Anonymous:1994:NFS

[49] Anonymous. New firewall server launched. Network Security, 1994 (8):5, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900868.

Anonymous:1994:BAN

[50] Anonymous. Bell Atlantic network delayed. Network Security, 1994 (8):5–6, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900876.

Anonymous:1994:RN

[51] Anonymous. Revamping the net! Network Security, 1994(8):6, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894900884.

Anonymous:1994:PME

[52] Anonymous. PGP misconceptions explained. Network Security, 1994 (8):6, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900892.

Anonymous:1994:CPFa

[53] Anonymous. Cellular phone fraud. Network Security, 1994(8):7–10, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894900906.

Hancock:1994:IF

[54] Bill Hancock. Internet firewalls. Network Security, 1994(8):10–11, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894900914.

Peterson:1994:FMP

[55] Padgett Peterson. The fortress mentality — Part II: The Internet. Network Security, 1994(8):

11–14, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900922.

Ongetta:1994:EP

[56] Silvano Ongetta. Effectiveness of passwords. Network Security, 1994 (8):14–17, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900930.

Schifreen:1994:WMH

[57] Robert Schifreen. What motivates a hacker? Network Security, 1994 (8):17–19, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900949.

Anonymous:1994:Eb

[58] Anonymous. Events. Network Security, 1994(8):19–20, August 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900957.

Anonymous:1994:HTC

[59] Anonymous. High-tech calling-card number theft. Network Security, 1994(9):2, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894901554.

Anonymous:1994:BBO

[60] Anonymous. Bulletin board operators fined for piracy. Network Security, 1994(9):2, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894901562.

Anonymous:1994:OP

[61] Anonymous. Online privacy. Network Security, 1994(9):2, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894901570.

Anonymous:1994:IBH

[62] Anonymous. The Internet becomes the high seas of piracy. Network Security, 1994(9):2–3, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894901589.

Anonymous:1994:CCP

[63] Anonymous. Combatting cellular phone fraud. Network Security, 1994 (9):3, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894901597.

Anonymous:1994:UGW

[64] Anonymous. US government wiretapping updated. Network Security, 1994(9):3, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894901600.

Anonymous:1994:SHO

[65] Anonymous. Self-healing optical networks. Network Security, 1994 (9):3–4, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894901619.

Anonymous:1994:NUD

[66] Anonymous. New UK data protection registrar. Network Security, 1994(9):4, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894901627.

Anonymous:1994:NJ

[67] Anonymous.  Net judges.  Net-

work Security, 1994(9):4, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485894901635.

Anonymous:1994:AVU

[68] Anonymous. Australian virus update. Network Security, 1994(9): 4–5, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894901643.

Anonymous:1994:UMM

[69] Anonymous. UK messaging market commandeered. Network Security, 1994(9):5, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894901651.

Anonymous:1994:GTG

[70] Anonymous. Global telecoms growth. Network Security, 1994(9):5, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/135348589490166X.

Anonymous:1994:TIS

[71] Anonymous. Taiwan invests in the superhighway era. Network Security, 1994(9):5, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894901678.

Anonymous:1994:HSO

[72] Anonymous. High-speed optical test bed. Network Security, 1994 (9):5, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894901686.

Anonymous:1994:BBC

[73] Anonymous. Baby Bells commit to multimedia. Network Security, 1994 (9):5–6, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894901694.

Anonymous:1994:ECD

[74] Anonymous. Ericsson cooperates with DynaSoft on security. Network Security, 1994(9):6, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485894901708.

Anonymous:1994:MTA

[75] Anonymous. Microsoft and Telekom announce multimedia venture. Network Security, 1994(9):6, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485894901716.

Anonymous:1994:VPS

[76] Anonymous. Virus protection system for networks. Network Security, 1994(9):6, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894901724.

Anonymous:1994:NAN

[77] Anonymous. Norton administrator for networks 1.5. Network Security, 1994(9):6, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894901732.

Cheek:1994:SHK

[78] Martin Cheek. Sidewinder helps keep Internet better secured. Network Security, 1994(9):6–7, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485894901740.

Snell:1994:PBM

[79] Monica Snell. Policies bring more security to e-mail. Network Security, 1994(9):7, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894901759.

Hancock:1994:HPM

[80] Bill Hancock. Hazards of promiscuous mode programming. Network Security, 1994(9):8–9, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894901767.

Hruska:1994:VCN

[81] Jan Hruska. Virus control on networks: Solving the unsolvable. Network Security, 1994(9):10–13, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894901775.

Hansen:1994:UND

[82] Lesley Hansen. Use of network design to reduce security risks. Network Security, 1994(9):13–17, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894901783.

Schifreen:1994:PSP

[83] Robert Schifreen. Preventing software piracy. Network Security, 1994 (9):17–19, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894901791.

Anonymous:1994:Ed

[84] Anonymous. Events. Network Security, 1994(9):20, September 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894901805.

Anonymous:1994:SLB

[85] Anonymous. Security leak at BT. Network Security, 1994(10):2, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485894900507.

Anonymous:1994:CAS

[86] Anonymous. Computer abuse surveys. Network Security, 1994 (10):2, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900515.

Anonymous:1994:ILF

[87] Anonymous. Internet liberation front. Network Security, 1994 (10):2–3, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900523.

Anonymous:1994:PCH

[88] Anonymous. Pressure to charge hacker. Network Security, 1994 (10):3, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900531.

Anonymous:1994:CPFb

[89] Anonymous. Cellular phone fraud arrest. Network Security, 1994 (10):3, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/135348589490054X.

Anonymous:1994:IPS

[90] Anonymous. Internet pirates sale again. Network Security, 1994 (10):3–4, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900558.

Anonymous:1994:NEI

[91] Anonymous. Nationwide EDI for India. Network Security, 1994 (10):4, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900566.

Anonymous:1994:TGF

[92] Anonymous. Toronto gets Free-Net. Network Security, 1994(10):4, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894900574.

Anonymous:1994:MLN

[93] Anonymous. Microsoft launch new online service. Network Security, 1994(10):4, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900582.

Anonymous:1994:PND

[94] Anonymous. PC network disaster recovery. Network Security, 1994 (10):5, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900590.

Anonymous:1994:GCB

[95] Anonymous. The great ‘cyber-bucks’ give away. Network Security, 1994 (10):5, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900604.

Anonymous:1994:HSS

[96] Anonymous. High speed switching and networking. Network Security, 1994(10):5, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900612.

Anonymous:1994:CUP

[97] Anonymous. Checks for underperforming networks. Network Security, 1994(10):5, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485894900620.

Anonymous:1994:CSN

[98] Anonymous. Crime and security network. Network Security, 1994 (10):6, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900639.

Anonymous:1994:EEM

[99] Anonymous. EU ends monopolies on phone networks. Network Security, 1994(10):6, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900647.

Armstrong:1994:IC

[100] Lisa Armstrong. Internet copyrights. Network Security, 1994 (10):6, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900655.

English:1994:IWP

[101] Erin English. Internet white pages with PGP. Network Security, 1994 (10):7, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900663.

Cheek:1994:FTC

[102] Martin Cheek. Financial transactions to come onto the Internet. Network Security, 1994(10):7, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894900671.

Armstrong:1994:PDS

[103] Lisa Armstrong. PersonaCard 100 data security token. Network Security, 1994(10):7–8, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/135348589490068X.

Hancock:1994:NKF

[104] Bill Hancock. Naming is the key to the future. Network Security, 1994 (10):8–9, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900698.

Gordon:1994:ISC

[105] Sara Gordon. IRC and security — can the two co-exist? Network Security, 1994(10):10–17, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485894900701.

Schifreen:1994:HHD

[106] Robert Schifreen. How hackers do it. Network Security, 1994(10): 17–19, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/135348589490071X.

Anonymous:1994:Ee

[107] Anonymous. Events. Network Security, 1994(10):20, October 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900728.

Anonymous:1994:LWN

[108] Anonymous. LAN/WAN news. Network Security, 1994(11):2–5, November 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894900965.

Madsen:1994:CC

[109] Wayne Madsen. The Clipper controversy. Network Security, 1994 (11):6–11, November 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900973.

Doyle:1994:TLS

[110] Frank Doyle. Tricks of the LAN security trade. Network Security, 1994 (11):12–13, November 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900981.

Madsen:1994:OIE

[111] Wayne Madsen. Online industrial espionage. Network Security, 1994 (11):14–18, November 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/135348589490099X.

Lindup:1994:CIU

[112] Ken Lindup. Commercial Internet use: a new risk? Network Security, 1994(11):19, November 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894901007.

Anonymous:1994:Ef

[113] Anonymous. Events. Network Security, 1994(11):20, November 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894901015.

Anonymous:1994:PCD

[114] Anonymous. Piracy charges dismissed against LaMacchia. Network Security, 1994(12):2, December 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900140.

Anonymous:1994:CSS

[115] Anonymous. Computers security support centre set up. Network Security, 1994(12):2–3, December 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900159.

Anonymous:1994:CMN

[116] Anonymous. Corporate move to the net. Network Security, 1994 (12):3, December 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900167.

Anonymous:1994:CVG

[117] Anonymous. Cellular on the verge of going mass market. Network Security, 1994(12):3, December 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485894900175.

Anonymous:1994:AAA

[118] Anonymous. Axent announces acquisition of datamedia. Network Security, 1994(12):4, December 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900183.

Anonymous:1994:VAP

[119] Anonymous. Virus alert program for information highway users. Network Security, 1994(12):4, December 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894900191.

Anonymous:1994:KMP

[120] Anonymous. Keeping e-mail private. Network Security, 1994 (12):4, December 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900205.

Anonymous:1994:SNM

[121] Anonymous. Secure network management initiative. Network Security, 1994(12):4, December 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900213.

Armstrong:1994:BAS

[122] Lisa Armstrong. Bank of America secures the Internet. Network Security, 1994(12):4–5, December 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900221.

Hall:1994:LBL

[123] Phil Hall. Laser beams as a LAN connectivity alternative. Network Security, 1994(12):5–6, December 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/135348589490023X.

Schwartau:1994:NN

[124] Winn Schwartau. To net or not to net? Network Security, 1994 (12):7–11, December 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900248.

Crocker:1994:IMI

[125] Norman Crocker. An introduction to MVS integrity concerns. Network Security, 1994(12):12–16, December 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894900256.

Peterson:1994:DLA

[126] Padgett Peterson. Dangerous letters: ANSI bombs and forged email. Network Security, 1994(12): 17–19, December 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485894900264.

Anonymous:1994:Eg

[127] Anonymous.  Events.  Network

Security, 1994(12):19–20, December 1994. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485894900272.

Anonymous:1995:ISA

[128] Anonymous. IP spoofing attacks latest threat. Network Security, 1995(1):2, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900599.

Anonymous:1995:OC

[129] Anonymous. Organizations in cyberspace. Network Security, 1995 (1):2, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900605.

Anonymous:1995:RBF

[130] Anonymous. Russians better at fighting phone fraud. Network Security, 1995(1):2–3, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900617.

Anonymous:1995:BVS

[131] Anonymous. A boom in VAN services. Network Security, 1995 (1):3, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900629.

Anonymous:1995:LS

[132] Anonymous. The LAN surges on. Network Security, 1995(1):3, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485898900630.

Anonymous:1995:CSM

[133] Anonymous. Concert seeking 10% of market. Network Security, 1995 (1):3, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898900642.

Anonymous:1995:SSD

[134] Anonymous. Securicor supplies data security business service. Network Security, 1995(1):3–4, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485898900654.

Anonymous:1995:CCD

[135] Anonymous. CyberSource, the cyberdistributor. Network Security, 1995 (1):4, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898900666.

Anonymous:1995:NMR

[136] Anonymous. Network monitoring report generation system. Network Security, 1995(1):4, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898900678.

Armstrong:1995:AI

[137] Lisa Armstrong. AT&T IVES. Network Security, 1995(1):4–5, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348589890068X.

Anonymous:1995:NLSa

[138] Anonymous. New level of security for global electronic trading. Network Security, 1995(1):5, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898900691.

Anonymous:1995:FTT

[139] Anonymous. Fault-tolerant TCP/ IP for LAN to mainframe connectivity. Network Security, 1995 (1):5, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900708.

Anonymous:1995:AII

[140] Anonymous. Award for IBM’s Internet security firewall. Network Security, 1995(1):5, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589890071X.

Anonymous:1995:FNS

[141] Anonymous. Flood of network security products. Network Security, 1995(1):5–6, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900721.

English:1995:SKS

[142] Erin English. MIT scales up Kerberos security system for Internet. Network Security, 1995 (1):6, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900733.

English:1995:NCP

[143] Erin English. A new certification programme for network security. Network Security, 1995 (1):6–7, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900745.

English:1995:SMS

[144] Erin English. Security made simple with SOCKS. Network Security, 1995(1):7, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900757.

Wolfe:1995:IT

[145] Henry Wolfe. The Internet threat. Network Security, 1995(1):7–8, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800800380.

Schwartau:1995:PAPa

[146] Winn Schwartau. Password alternatives — Part 1. Network Security, 1995(1):9–13, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800392.

Kovacich:1995:LAN

[147] Gerald L. Kovacich. Local area networks security: Establishing policies and procedures. Network Security, 1995(1):13–16, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800409.

Saxby:1995:EMP

[148] Stephen Saxby. Electronic monitoring poses e-mail dilemma. Network Security, 1995(1):17–18, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898900769.

Zajac:1995:BRM

[149] Bernard P. Zajac. Book review: E-Mail Security — How To Keep Your Electronic Messages Private: Bruce Schneier, published by John Wiley & Sons at $24.95. ISBN 0-471-05318-X. Network Security, 1995 (1):18–19, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900770.

Anonymous:1995:Ea

[150] Anonymous. Events. Network Security, 1995(1):20, January 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898900782.

Anonymous:1995:TEV

[151] Anonymous. Telnet encryption vulnerability. Network Security, 1995 (2):2, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901132.

Anonymous:1995:VM

[152] Anonymous. Vulnerabilities in /bin/mail. Network Security, 1995 (2):2–3, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901140.

Anonymous:1995:HLT

[153] Anonymous. Hacker leaves threatening messages. Network Security, 1995(2):3, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901159.

Lindup:1995:MA

[154] Ken Lindup. Mitnick arrested. Network Security, 1995(2):3–4, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485895901167.

Anonymous:1995:BCA

[155] Anonymous. BR complains about counterfeit NetWare. Network Security, 1995(2):4, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901175.

Anonymous:1995:GCG

[156] Anonymous. Global communications group formed. Network Security, 1995(2):4, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901183.

Anonymous:1995:SCI

[157] Anonymous. Support for corporate Internet users. Network Security, 1995(2):4, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901191.

Anonymous:1995:MNR

[158] Anonymous. Microsoft Network receives mixed response. Network Security, 1995(2):4–5, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485895901205.

Anonymous:1995:OBG

[159] Anonymous. Online banking growth. Network Security, 1995(2):5, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485895901213.

Anonymous:1995:DPB

[160] Anonymous. Draft plans to break European telecoms monopolies. Network Security, 1995(2):5, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485895901221.

Anonymous:1995:GW

[161] Anonymous. The growing Web. Network Security, 1995(2):5, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/135348589590123X.

Anonymous:1995:EUN

[162] Anonymous. Eliminate unauthorized network access. Network Security, 1995(2):5–6, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901248.

Snell:1995:SSI

[163] Monica Snell. Secure shopping on the Internet. Network Security, 1995(2):6, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901256.

Armstrong:1995:ES

[164] Lisa Armstrong. Endorsements for SHTTP. Network Security, 1995 (2):6–7, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901264.

English:1995:PTP

[165] Erin English. Passwords through pagers. Network Security, 1995 (2):7, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901272.

Hancock:1995:ARN

[166] Bill Hancock. Assessing and reducing network risk. Network Security, 1995(2):7–9, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800410.

Zajac:1995:CVL

[167] Bernard P. Zajac. Computer viruses — legal options. Network Security, 1995(2):9–10, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800422.

Schwartau:1995:PAPb

[168] Winn Schwartau. Password alternatives — Part 2. Network Security, 1995 (2):11–15, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800434.

Kovacich:1995:SRV

[169] Gerald Kovacich. Security requirements for voice messaging operations. Network Security, 1995(2): 15–18, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800800446.

Lindup:1995:HW

[170] Ken Lindup. The hacker war. Network Security, 1995(2):19, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485895901280.

Anonymous:1995:Eb

[171] Anonymous. Events. Network Security, 1995(2):20, February 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901299.

Anonymous:1995:SV

[172] Anonymous. Sendmail vulnerabilities. Network Security, 1995(3):2, March 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485895901302.

Anonymous:1995:CVa

[173] Anonymous. ‘at’ and ‘cron’ vulnerabilities. Network Security, 1995 (3):2–3, March 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901310.

Anonymous:1995:PAV

[174] Anonymous. Providers not accountable for virus distribution. Network Security, 1995(3):3, March 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901329.

Anonymous:1995:NSSa

[175] Anonymous. Natas spreads in southwestern United States. Network Security, 1995(3):3, March 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901337.

Anonymous:1995:BOI

[176] Anonymous. Boom in online information. Network Security, 1995 (3):3–4, March 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901345.

Anonymous:1995:AOP

[177] Anonymous. America Online plans to sell 5% stake. Network Security, 1995(3):4, March 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901353.

Anonymous:1995:ACN

[178] Anonymous. ACE/Client for NetWare. Network Security, 1995(3):4, March 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/135348589590137X.

Anonymous:1995:NEA

[179] Anonymous. New Ethernet adapter card. Network Security, 1995(3): 4, March 1995. CODEN NTSCF5.

ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/1353485895901361.

Anonymous:1995:SBL

[180] Anonymous. Small businesses link through BT. Network Security, 1995(3):4–5, March 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901388.

Riggs:1995:TTR

[181] Brian Riggs. Top to tail router security. Network Security, 1995(3):5, March 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485895901396.

Snell:1995:ASS

[182] Monica Snell. AT&T strengthen security of Network Notes. Network Security, 1995(3):5, March 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/135348589590140X.

Thomsen:1995:ISS

[183] Dan Thomsen. IP spoofing and session hijacking. Network Security, 1995(3):6–11, March 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800800458.

Sheffield:1995:NRPa

[184] Chris Sheffield. Network review — Part 1. Network Security, 1995 (3):11–16, March 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580080046X.

Lindup:1995:AKM

[185] Ken Lindup. The arrest of Kevin Mitnick. Network Security, 1995 (3):16–19, March 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800471.

Anonymous:1995:Ec

[186] Anonymous. Events. Network Security, 1995(3):20, March 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901418.

Anonymous:1995:GTH

[187] Anonymous. ‘Good Times’ is a hoax. Network Security, 1995(4):2, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485896897005.

Anonymous:1995:SCD

[188] Anonymous. Satan costs developer his job. Network Security, 1995(4):2, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902090.

Anonymous:1995:HCC

[189] Anonymous. Hacker charged in Canada. Network Security, 1995(4):

2, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/1353485895902104.

Anonymous:1995:CPH

[190] Anonymous. Child porn hacker sentenced. Network Security, 1995 (4):2–3, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902112.

Anonymous:1995:BEG

[191] Anonymous. BT engineer guilty of piracy. Network Security, 1995(4): 3, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/1353485896897017.

Anonymous:1995:WWW

[192] Anonymous. World Wide Web needs a security boost. Network Security, 1995(4):3, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902120.

Anonymous:1995:BMM

[193] Anonymous. BT makes moves in Sweden and Italy. Network Security, 1995(4):3–4, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902139.

Anonymous:1995:FLM

[194] Anonymous. Fibre links for major buildings. Network Security, 1995(4):4, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902147.

Anonymous:1995:IIS

[195] Anonymous. Investing in Internet security. Network Security, 1995(4):4, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902155.

Anonymous:1995:III

[196] Anonymous. IBM increases Internet services. Network Security, 1995(4):4, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902163.

Anonymous:1995:PDS

[197] Anonymous. Private drive on superhighway. Network Security, 1995 (4):4–5, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902171.

Anonymous:1995:UPF

[198] Anonymous. Users prefer frame relay. Network Security, 1995(4):5, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/135348589590218X.

Anonymous:1995:SDE

[199] Anonymous. SentryLink data encryption devices. Network Security, 1995(4):5, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895902201.

Anonymous:1995:MPB

[200] Anonymous. MCI and Pacific Bell offer Internet services. Network Security, 1995(4):5, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895902198.

Anonymous:1995:FTW

[201] Anonymous. Fault tolerant Web server. Network Security, 1995(4):6, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/135348589590221X.

Anonymous:1995:DAS

[202] Anonymous. Defender authentication software. Network Security, 1995(4):6, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895902228.

Anonymous:1995:NSPa

[203] Anonymous. Network security plus. Network Security, 1995 (4):6–7, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895902236.

Armstrong:1995:C

[204] Lisa Armstrong. CyberGuard. Network Security, 1995(4):7, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485896897029.

Armstrong:1995:IT

[205] Lisa Armstrong. Info-terrorism. Network Security, 1995(4):7–8, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485896897030.

Armstrong:1995:SP

[206] Lisa Armstrong. Secure PBX. Network Security, 1995(4):8, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485896897042.

Riggs:1995:VA

[207] Brian Riggs. Vaulting with ATM. Network Security, 1995(4):8–9, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485896897054.

Sheffield:1995:NRPb

[208] Chris Sheffield. Network review — Part 2. Network Security, 1995 (4):9–13, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897078.

Armstrong:1995:OSC

[209] Lisa Armstrong. Oracle Secure C/S. Network Security, 1995(4):9, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485896897066.

Saxby:1995:GSP

[210] Stephen Saxby. G7 Summit poses challenge to governments on public access to official information. Network Security, 1995(4):14–15, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/135348589689708X.

Anonymous:1995:Ed

[211] Anonymous. Events. Network Security, 1995(4):15–16, April 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902244.

Anonymous:1995:BHT

[212] Anonymous. The birth of the ‘hacker-tracker’. Network Security, 1995(5):2, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901922.

Anonymous:1995:DIS

[213] Anonymous. Developments in Internet security. Network Security, 1995(5):2, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901930.

Anonymous:1995:KFH

[214] Anonymous. Korea fights the hackers. Network Security, 1995(5):2, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485895901949.

Anonymous:1995:FNI

[215] Anonymous. Fears of Net insecurity. Network Security, 1995(5):2–3, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485895901957.

Anonymous:1995:FAH

[216] Anonymous. Fixes for AIX holes. Network Security, 1995(5):3, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485895901965.

Anonymous:1995:FUF

[217] Anonymous. Firewalls under fire. Network Security, 1995(5):3–4, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485895901973.

Anonymous:1995:FFA

[218] Anonymous. Further fights against mobile phone fraud. Network Security, 1995(5):4–5, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901981.

Anonymous:1995:IIN

[219] Anonymous. Internet interest not necessarily business. Network Security, 1995(5):5, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/135348589590199X.

Anonymous:1995:UFO

[220] Anonymous. US fibre-optic cables set to triple. Network Security, 1995(5):5, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902007.

Anonymous:1995:ISIa

[221] Anonymous. Internet security issues addressed. Network Security, 1995(5):5–6, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902015.

Anonymous:1995:SUI

[222] Anonymous. SITA unveils Internet service. Network Security, 1995(5):6, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902023.

Anonymous:1995:CSS

[223] Anonymous. Client/server security package. Network Security, 1995(5):6, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/135348589590204X.

Anonymous:1995:NSI

[224] Anonymous. Netscape’s SSL Internet security. Network Security, 1995(5):6, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902031.

Anonymous:1995:AANa

[225] Anonymous. ANR addresses network security issues. Network Security, 1995(5):6–7, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902058.

Anonymous:1995:SWS

[226] Anonymous. Secrets for Windows, security for Internet. Network Security, 1995(5):7, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902066.

Anonymous:1995:PSN

[227] Anonymous. PCMCIA security for notebooks and Novell LANs. Network Security, 1995(5):7, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485895902074.

English:1995:IVH

[228] Erin English. IP Version 6 holds promise for the future. Network Security, 1995(5):7–8, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897091.

Bournellis:1995:PET

[229] Cynthia Bournellis. Putting an end to too many passwords. Network Security, 1995(5):8, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897108.

English:1995:SDG

[230] Erin English. SATAN does not go undetected. Network Security, 1995(5):8–9, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/135348589689711X.

Bucholtz:1995:SCS

[231] Chris Bucholtz. Suit challenges status of cryptography as munition. Network Security, 1995(5):9, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485896897121.

Hancock:1995:SMY

[232] Bill Hancock. Stop that e-mail! You are probably breaking the law! Network Security, 1995 (5):10–12, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897133.

Gordon:1995:DA

[233] Sarah Gordon. Devil’s advocate. Network Security, 1995 (5):12–15, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897145.

Sheffield:1995:NRPc

[234] Chris Sheffield. Network review — Part 3. Network Security, 1995 (5):15–19, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897157.

Anonymous:1995:Ee

[235] Anonymous. Events. Network Security, 1995(5):20, May 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902082.

Anonymous:1995:CISa

[236] Anonymous. Cisco IOS software vulnerability. Network Security, 1995(6):2, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901701.

Anonymous:1995:HMP

[237] Anonymous. Hackers monitored police. Network Security, 1995(6):2, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/135348589590171X.

Anonymous:1995:URS

[238] Anonymous. Underground reconnaissance service. Network Security, 1995(6):2, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901728.

Anonymous:1995:BBC

[239] Anonymous. Black Baron convicted. Network Security, 1995 (6):2–3, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901736.

Anonymous:1995:HPC

[240] Anonymous. Hackers penetrate college computers. Network Security, 1995(6):3, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901744.

Anonymous:1995:OCT

[241] Anonymous. Outfoxing the cellular thieves. Network Security, 1995(6):3, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901752.

Anonymous:1995:BBF

[242] Anonymous. Bulter Bunny found guilty. Network Security, 1995(6): 3, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/1353485895901760.

Anonymous:1995:RPI

[243] Anonymous. Row puts Internet provision in doubt. Network Security, 1995(6):4, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901779.

Anonymous:1995:IGP

[244] Anonymous. Internet goes private in Brazil. Network Security, 1995(6):4, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901787.

Anonymous:1995:IA

[245] Anonymous. Infohighway in Africa? Network Security, 1995(6):4, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485895901795.

Anonymous:1995:ISIb

[246] Anonymous. Internet steals international phone calls. Network Security, 1995(6):4–5, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901809.

Anonymous:1995:MRA

[247] Anonymous. Mergent reaches agreement with Bull. Network Security, 1995(6):5, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901817.

Anonymous:1995:EEW

[248] Anonymous. Europe encrypts weather data. Network Security, 1995(6): 5, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/1353485896897169.

Anonymous:1995:MNP

[249] Anonymous. Microsoft network pushes into electronic commerce. Network Security, 1995(6):5–6, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485895901825.

Anonymous:1995:BMA

[250] Anonymous. Broad multimedia alliance. Network Security, 1995(6): 6, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/1353485895901833.

Anonymous:1995:IMA

[251] Anonymous. ICL and Motorola ally for network security. Network Security, 1995(6):6, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901841.

Anonymous:1995:NSD

[252] Anonymous. Nortel and Shiva direct Internet traffic. Network Security, 1995(6):6, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/135348589590185X.

Anonymous:1995:MPS

[253] Anonymous. Motorola pushes its security portfolio. Network Security, 1995(6):6–7, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897170.

Anonymous:1995:NSA

[254] Anonymous. Network security assessment tool. Network Security, 1995(6):7, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901868.

Anonymous:1995:MHF

[255] Anonymous. Motorola helps firms protect data. Network Security, 1995(6):7, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901876.

Anonymous:1995:FDB

[256] Anonymous. Financial data backed up by ISDN. Network Security, 1995(6):7, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901884.

Anonymous:1995:ISIc

[257] Anonymous. Internet security improved. Network Security, 1995 (6):7–8, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901892.

Bucholtz:1995:NSB

[258] Chris Bucholtz. Networks suffer backup blues. Network Security, 1995(6):8, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901906.

Anonymous:1995:TSD

[259] Anonymous. The ‘TouchLock’ security device. Network Security, 1995(6):8, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897182.

Anonymous:1995:IGF

[260] Anonymous. Indian groups face increased Government electronic surveillance. Network Security, 1995(6): 9, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/1353485896897194.

Riggs:1995:CPF

[261] Brian Riggs. Church puts freedom of Internet on trial. Network Security, 1995(6):10, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897212.

Bucholtz:1995:JSH

[262] Chris Bucholtz. Judge says hacker deserves more time. Network Security, 1995(6):10, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897200.

Bucholtz:1995:STC

[263] Chris Bucholtz. Security takes centre stage at Internet conference. Network Security, 1995 (6):10–11, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897224.

Kaneshige:1995:BI

[264] Tom Kaneshige. Banking on the Internet. Network Security, 1995(6): 11, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/1353485896897236.

Bucholtz:1995:UWG

[265] Chris Bucholtz. US war games fought. Network Security, 1995 (6):11–12, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897248.

Bucholtz:1995:NSF

[266] Chris Bucholtz. Network security far too lax. Network Security, 1995(6):12, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/135348589689725X.

Hancock:1995:SSE

[267] Bill Hancock. Simple social engineering. Network Security, 1995 (6):13–14, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897261.

Worlock:1995:CCSa

[268] David R. Worlock. The culture of control: Safeguarding intellectual property in the age of networks — Part 1. Network Security, 1995(6):14–17, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897273.

Saxby:1995:NRL

[269] Stephen Saxby. Network-related law — recent decisions from USA. Network Security, 1995(6):18–19, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485896897285.

Anonymous:1995:Ef

[270] Anonymous. Events. Network Security, 1995(6):20, June 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901914.

Anonymous:1995:PT

[271] Anonymous. PKZIP Trojan. Network Security, 1995(7):2, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485895900942.

Anonymous:1995:RPC

[272] Anonymous. Regulation of porn in cyberspace. Network Security, 1995(7):2, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900950.

Anonymous:1995:CVb

[273] Anonymous. Caibua virus. Network Security, 1995(7):2–3, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485895900969.

Anonymous:1995:CFB

[274] Anonymous. Campaign to fight $3 billion a year problem. Network Security, 1995(7):3, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900977.

Anonymous:1995:SPP

[275] Anonymous. Swedish pirates prosecuted. Network Security, 1995(7): 3, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/1353485895900993.

Anonymous:1995:YVH

[276] Anonymous. The year of the virus hoax. Network Security, 1995(7):3, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895900985.

Anonymous:1995:SWR

[277] Anonymous. Security worries with Registration Wizard. Network Security, 1995(7):4, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901000.

Anonymous:1995:FAS

[278] Anonymous. Federal agency to secure infohighway. Network Security, 1995(7):4, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901019.

Anonymous:1995:NEV

[279] Anonymous. New electronic verification standard. Network Security, 1995(7):4, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901027.

Anonymous:1995:WOP

[280] Anonymous. The worries of online purchasers. Network Security, 1995(7):4–5, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901035.

Anonymous:1995:RN

[281] Anonymous. Radio over the Net. Network Security, 1995(7):5, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485895901043.

Anonymous:1995:NEO

[282] Anonymous. Novell and EUNet offer Internet connection. Network Security, 1995(7):5, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/135348589590106X.

Anonymous:1995:EN

[283] Anonymous. The electric network. Network Security, 1995(7):5, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485895901051.

Anonymous:1995:FKO

[284] Anonymous. Firewalls to keep online spies at bay. Network Security, 1995(7):6, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901078.

Anonymous:1995:WAV

[285] Anonymous. Windows anti-virus virtual device driver. Network Security, 1995(7):6, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901094.

Anonymous:1995:PGW

[286] Anonymous. Protection as good as at the Whitehouse. Network Security, 1995(7):6, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901086.

Anonymous:1995:MSE

[287] Anonymous. Message security for electronic communications. Network Security, 1995(7):6–7, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901108.

Anonymous:1995:SMM

[288] Anonymous. Secure e-mail for Microsoft-Mail users. Network Security, 1995(7):7, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896897297.

Kaneshige:1995:GRS

[289] Thomas Kaneshige. Government’s role in superhighway security. Network Security, 1995(7):7, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901116.

Kaneshige:1995:SGU

[290] Tom Kaneshige. Software giants unite to filter the Net. Network Security, 1995(7):8, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897303.

Bucholtz:1995:VMS

[291] Chris Bucholtz. Visa and MasterCard settle on Internet security standard. Network Security, 1995(7):8–9, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897315.

Bucholtz:1995:EEC

[292] Chris Bucholtz. Encryption exports, Clipper policy criticized by consortium. Network Security, 1995(7):9, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897327.

Beseke:1995:DCI

[293] Kermit Beseke. Developing a corporate Internet policy. Network Security, 1995(7):10–13, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897339.

Worlock:1995:CCSb

[294] David R. Worlock. The culture of control: Safeguarding intellectual property in the age of networks — Part 2. Network Security, 1995(7):13–17, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896897340.

Peterson:1995:FMPa

[295] Padgett Peterson. The fortress mentality — Part III: Tokens, onetime password devices and the Roman Army. Network Security, 1995 (7):17–19, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896897352.

Anonymous:1995:Eg

[296] Anonymous. Events. Network Security, 1995(7):20, July 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901124.

Anonymous:1995:FHC

[297] Anonymous. French hacker cracks Netscape code. Network Security, 1995(8):2, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895900802.

Anonymous:1995:SVV

[298] Anonymous. Sendmail V.5 vulnerability. Network Security, 1995 (8):2, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900810.

Anonymous:1995:JBH

[299] Anonymous. Japanese battle hackers, viruses. Network Security, 1995 (8):2–3, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900829.

Anonymous:1995:CBA

[300] Anonymous. Citicorp bank accounts breached. Network Security, 1995(8):3, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900837.

Anonymous:1995:SEE

[301] Anonymous. Stronger encryption exportable. Network Security, 1995 (8):3, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900845.

Anonymous:1995:BSA

[302] Anonymous. Bringing security to ATM. Network Security, 1995 (8):3–4, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900853.

Anonymous:1995:ENE

[303] Anonymous. Europe negotiates over encryption. Network Security, 1995(8):4, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895900861.

Anonymous:1995:NGP

[304] Anonymous. Netscape goes public. Network Security, 1995(8):4, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/135348589590087X.

Anonymous:1995:AFN

[305] Anonymous. AT&T finally notices the Internet. Network Security, 1995(8):4, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895900888.

Anonymous:1995:A

[306] Anonymous. The authorizer. Network Security, 1995(8):4–5, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485895900896.

Anonymous:1995:NLSb

[307] Anonymous. Netscape’s latest security effort. Network Security, 1995(8):5, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/135348589590090X.

Anonymous:1995:NGN

[308] Anonymous. New generation network security product. Network Security, 1995(8):5, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900926.

Anonymous:1995:NSE

[309] Anonymous. NetCop scans for explicit material. Network Security, 1995(8):5, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900918.

Bucholtz:1995:PMP

[310] Chris Bucholtz. Pilot’s e-mail has Pentagon up in arms. Network Security, 1995(8):5–6, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897364.

Kaneshige:1995:FSS

[311] Thomas Kaneshige. Full service security shop to ward off hackers. Network Security, 1995 (8):6, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897376.

Kaneshige:1995:SSR

[312] Thomas Kaneshige. Security systems for the road most travelled. Network Security, 1995(8):6–7, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485896897388.

Bournellis:1995:IPD

[313] Cynthia Bournellis. Internet project defines middle ground. Network Security, 1995(8):7, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/135348589689739X.

English:1995:CCJ

[314] Erin English. CyberCash and Checkfree join forces. Network Security, 1995(8):7–8, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896897406.

Cohen:1995:IHPa

[315] Fred Cohen. Internet holes — Part 1: Internet control message protocol. Network Security, 1995 (8):8–10, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/135348589689742X.

English:1995:CUN

[316] Erin English. CyberAngels unite on the net. Network Security, 1995(8):8, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896897418.

Shaer:1995:SS

[317] Caroline Shaer. Single sign-on. Network Security, 1995(8):11–15, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485896897431.

Ongetta:1995:EDI

[318] Silvano Ongetta. Electronic data interchange — what are the security measures? Network Security, 1995 (8):16–19, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897443.

Anonymous:1995:Eh

[319] Anonymous. Events. Network Security, 1995(8):19–20, August 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900934.

Anonymous:1995:NSSb

[320] Anonymous. Netscape’s software security scare. Network Security, 1995 (9):2, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900594.

Anonymous:1995:SPV

[321] Anonymous. Solaris ps vulnerability. Network Security, 1995(9): 2–3, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900608.

Anonymous:1995:NSPb

[322] Anonymous. New security patches from Sun. Network Security, 1995 (9):3, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895900616.

Anonymous:1995:SSV

[323] Anonymous. Sun Sendmail -oR vulnerability. Network Security, 1995 (9):3–4, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895900624.

Anonymous:1995:UCT

[324] Anonymous. Undercover cybercops trap hackers. Network Security, 1995 (9):4, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895900632.

Anonymous:1995:TUI

[325] Anonymous. Ties to US intelligence for domain name assigner. Network Security, 1995 (9):4, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895900640.

Anonymous:1995:AOF

[326] Anonymous. America Online fights hackers. Network Security, 1995 (9):4, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900659.

Anonymous:1995:OCP

[327] Anonymous. Online child pornography debate continues. Network Security, 1995(9):5, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900667.

Anonymous:1995:MWS

[328] Anonymous. Malicious Word spreads to the Net. Network Security, 1995 (9):5, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900675.

Anonymous:1995:DAV

[329] Anonymous. Document about virus contains one. Network Security, 1995 (9):5, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900683.

Rees:1995:ELB

[330] Frank Rees. An expensive lesson for bulletin board user. Network Security, 1995(9):6, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900691.

Anonymous:1995:MISa

[331] Anonymous. Microsoft’s Internet security claims criticized. Network Security, 1995(9):6, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897455.

Anonymous:1995:ECP

[332] Anonymous. Electronic cheques planned. Network Security, 1995 (9):6–7, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900705.

Anonymous:1995:MSM

[333] Anonymous. Mobile services market set to triple. Network Security, 1995(9):7, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900713.

Anonymous:1995:MISb

[334] Anonymous. Making the Internet a safer place. Network Security, 1995 (9):7, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900721.

Anonymous:1995:AST

[335] Anonymous. AT&T splits into three. Network Security, 1995 (9):8, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589590073X.

Anonymous:1995:INL

[336] Anonymous. Integrity of network link guaranteed. Network Security, 1995(9):8, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485895900748.

Anonymous:1995:SWV

[337] Anonymous. Sophos’ Windows 95 virus scanner. Network Security, 1995(9):8, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485895900756.

Anonymous:1995:HSL

[338] Anonymous. DES for HDLC/ SDLC links. Network Security, 1995 (9):8, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485895900764.

Anonymous:1995:IFI

[339] Anonymous. Internet firewall from Integralis. Network Security, 1995 (9):8–9, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900772.

Barry:1995:SUL

[340] David Barry. SNMP upgrade to lose security features. Network Security, 1995 (9):9–10, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897479.

Davis:1995:FC

[341] Cindy Davis. Firewall consortium. Network Security, 1995 (9):9, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900780.

Anonymous:1995:NSN

[342] Anonymous. New secure Net products. Network Security, 1995 (9):9, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897467.

Kaneshige:1995:FPO

[343] Thomas Kaneshige. FBI to police online amidst heated controversy. Network Security, 1995(9): 10, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897480.

Hancock:1995:ANR

[344] Bill Hancock. Attacking network routers. Network Security, 1995(9): 11–12, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897492.

Peterson:1995:FMPb

[345] Padgett Peterson. The fortress mentality — Part IV: Telecomputing, telecommuting and serving remote customers. Network Security, 1995(9): 12–14, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897509.

Cohen:1995:IHPb

[346] Fred Cohen. Internet holes — Part 2: Packet fragmentation attacks. Network Security, 1995(9): 14–16, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897510.

Herschberg:1995:ICI

[347] Bob Herschberg and Edo Roos Lindgreen. Internet: a certain insecurity. Network Security, 1995(9): 16–19, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897522.

Anonymous:1995:Ei

[348] Anonymous. Events. Network Security, 1995(9):20, September 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895900799.

Anonymous:1995:SLV

[349] Anonymous. Sun 4.1x loadmodule vulnerability. Network Security, 1995(10):2, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895902570.

Anonymous:1995:SVW

[350] Anonymous. syslog vulnerability workaround. Network Security, 1995(10):2, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895902589.

Anonymous:1995:TDV

[351] Anonymous. Telnet daemon vulnerabilities. Network Security, 1995 (10):2–3, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895902597.

Anonymous:1995:ISC

[352] Anonymous. Internet security code. Network Security, 1995(10):3, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485895902600.

Anonymous:1995:GCN

[353] Anonymous. Growing cellular networks in China. Network Security, 1995(10):3–4, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895902619.

Anonymous:1995:SPC

[354] Anonymous. Secure payments from Cybercash. Network Security, 1995 (10):4, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902627.

Anonymous:1995:DIC

[355] Anonymous. The digital ID case. Network Security, 1995(10):4, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485895902635.

Anonymous:1995:MLM

[356] Anonymous. Microsoft and Lotus make net moves. Network Security, 1995(10):4–5, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902643.

Anonymous:1995:FPF

[357] Anonymous. Firewall partnership formed. Network Security, 1995 (10):5, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902651.

Anonymous:1995:SST

[358] Anonymous. Single sign-on technology. Network Security, 1995 (10):5, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902678.

Anonymous:1995:FSN

[359] Anonymous. Firewall security for network applications. Network Security, 1995(10):5, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895902686.

Anonymous:1995:AANb

[360] Anonymous. Azlan’s alliance with Netscape and supernet. Network Security, 1995(10):5, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/135348589590266X.

Anonymous:1995:FRE

[361] Anonymous. Frame relay encryptor protects. Network Security, 1995 (10):5–6, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895902694.

Anonymous:1995:HPE

[362] Anonymous. H-P’s encryption engine proposal. Network Security, 1995(10):6, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896897534.

English:1995:OES

[363] Erin English. Oracle expands support for Oracle7. Network Security, 1995(10):6, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902708.

English:1995:OEF

[364] Erin English. Oracle embraces Fortezza. Network Security, 1995 (10):7, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897558.

Bournellis:1995:RNF

[365] Cynthia Bournellis. RSA negotiates with firewall vendors to establish VPNs. Network Security, 1995(10):7, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897546.

Bucholtz:1995:NAS

[366] Chris Bucholtz. Netscape answers security lapses with anti-bug contest. Network Security, 1995 (10):7–8, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/135348589689756X.

Kaneshige:1995:AIC

[367] Thomas Kaneshige. Auto industry to create its own network. Network Security, 1995 (10):8, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897571.

Hancock:1995:ECI

[368] Bill Hancock. Export of cryptographic information from the US: a brief look at the problems. Network Security, 1995(10):9–11, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896897583.

Cohen:1995:IHPc

[369] Fred Cohen. Internet holes — Part 3: The sendmail maelstrom. Network Security, 1995(10):11–12, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485896897595.

Schultz:1995:NPF

[370] E. Eugene Schultz. A new perspective on firewalls. Network Security, 1995 (10):13–17, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896897601.

Banes:1995:PLN

[371] D. W. Banes. Physical layer network security: What your LAN can do for you. Network Security, 1995 (10):17–19, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896897613.

Anonymous:1995:Ej

[372] Anonymous. Events. Network Security, 1995(10):20, October 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895902716.

Anonymous:1995:SGI

[373] Anonymous. Silicon Graphics Inc.’s Ip vulnerability. Network Security, 1995(11):2, November 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901426.

Anonymous:1995:UCL

[374] Anonymous. 30 US companies lose $66 million. Network Security, 1995 (11):2, November 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901434.

Anonymous:1995:FOW

[375] Anonymous. FBI outlines a wiretapping future. Network Security, 1995 (11):2–3, November 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901442.

Anonymous:1995:FFT

[376] Anonymous. Flaw in fault-tolerance software. Network Security, 1995 (11):3, November 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897625.

Madsen:1995:FRC

[377] Wayne Madsen. Fortezza reaches into civilian agencies. Network Security, 1995(11):3, November 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901450.

Anonymous:1995:CCI

[378] Anonymous. Companies crack down on Internet use. Network Security, 1995(11):3, November 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901469.

Anonymous:1995:CISb

[379] Anonymous. Check in at Internet site to check security. Network Security, 1995(11):3–4, November 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901477.

Anonymous:1995:WNI

[380] Anonymous. Woe is the Net, or is it? Network Security, 1995 (11):4, November 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901485.

Anonymous:1995:CWS

[381] Anonymous. Choosing a Web standard. Network Security, 1995 (11):4, November 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901493.

Anonymous:1995:PSV

[382] Anonymous. Product scans for viruses before a download. Network Security, 1995(11):4, November 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901507.

Anonymous:1995:SSN

[383] Anonymous. Single sign-on network security. Network Security, 1995 (11):4–5, November 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901515.

Anonymous:1995:MIM

[384] Anonymous. Merlin’s Internet magic. Network Security, 1995 (11):5, November 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901523.

Anonymous:1995:IUA

[385] Anonymous. Internet user authentication security. Network Security, 1995(11):5, November 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901531.

Anonymous:1995:DSP

[386] Anonymous. OS/2 data security product to integrate IBM’s security hooks. Network Security, 1995 (11):5, November 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/135348589590154X.

Browne:1995:UKL

[387] Graham Browne. Using DES with keys longer than 56 bits. Network Security, 1995(11):5–6, November 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897637.

Hancock:1995:RHK

[388] Bill Hancock. Recent history of known network breaches. Network Security, 1995(11):6–9, November 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897649.

Leach:1995:SCN

[389] John Leach and Colin Brown Zergo. Security considerations of network outsourcing. Network Security, 1995 (11):10–14, November 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897650.

Cohen:1995:IHPd

[390] Fred Cohen. Internet holes — Part 4: Network news transfer protocol. Network Security, 1995(11): 15–19, November 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897662.

Anonymous:1995:Ek

[391] Anonymous. Events. Network Security, 1995(11):20, November 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901558.

Anonymous:1995:XAV

[392] Anonymous. X authentication vulnerability. Network Security, 1995 (12):2, December 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901566.

Anonymous:1995:DIT

[393] Anonymous. Dramatic increase in telephone hacking. Network Security, 1995(12):2, December 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901574.

Anonymous:1995:PGP

[394] Anonymous. Police given permission to tap e-mail. Network Security, 1995(12):2, December 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901582.

Anonymous:1995:HUF

[395] Anonymous. HP-UX ftp vulnerability. Network Security, 1995 (12):2–3, December 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901590.

Anonymous:1995:RLL

[396] Anonymous. RSA loses licence for Hellman–Merkle algorithm. Network Security, 1995(12):3, December 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485895901604.

Anonymous:1995:VGM

[397] Anonymous. The viruses are getting more vicious. Network Security, 1995(12):3, December 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901612.

Anonymous:1995:JLB

[398] Anonymous. JavaScript language becomes a competitor. Network Security, 1995(12):3, December 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901620.

Anonymous:1995:FVB

[399] Anonymous. First virtual bank claims to be secure. Network Security, 1995 (12):4, December 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485895901639.

Anonymous:1995:LDP

[400] Anonymous. Long distance phone calls via the Internet. Network Security, 1995(12):4, December 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901655.

Anonymous:1995:SPW

[401] Anonymous. Security problems with Windows 95. Network Security, 1995 (12):4, December 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901647.

Anonymous:1995:PVM

[402] Anonymous. Protection from viruses for e-mail. Network Security, 1995 (12):4–5, December 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901663.

Anonymous:1995:FA

[403] Anonymous. A firewall for ATM. Network Security, 1995(12):5, December 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485895901671.

Anonymous:1995:UNS

[404] Anonymous. Unix network security. Network Security, 1995 (12):5, December 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/135348589590168X.

English:1995:AVW

[405] Erin English. AOLGOLD virus worries users. Network Security, 1995 (12):5–6, December 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897674.

English:1995:ISS

[406] Erin English. IP security standards battle it out. Network Security, 1995 (12):6, December 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897686.

Cohen:1995:IHPe

[407] Fred Cohen. Internet holes — Part 5a: 50 ways to attack your web systems. Network Security, 1995 (12):7–11, December 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897704.

English:1995:EEP

[408] Erin English. Exportable encryption policy found ‘unacceptable’. Network Security, 1995 (12):7, December 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897698.

Hancock:1995:IHR

[409] Bill Hancock. Internet hacking resources. Network Security, 1995(12): 11–15, December 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896897716.

Anonymous:1995:El

[410] Anonymous. Events. Network Security, 1995(12):16, December 1995. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485895901698.

Anonymous:1996:RYV

[411] Anonymous. rpc.ypupdated vulnerability. Network Security, 1996 (1):2, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901518.

Anonymous:1996:GUC

[412] Anonymous. Growth in UK cellphone fraud. Network Security, 1996 (1):2–3, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589690152X.

Anonymous:1996:PKE

[413] Anonymous. Public-key encryption flawed in time. Network Security, 1996(1):3, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901531.

Anonymous:1996:ERK

[414] Anonymous. Encryption report kept under lock and key. Network Security, 1996(1):3, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901543.

Anonymous:1996:NDB

[415] Anonymous. NetConnect distributes BorderWare firewall. Network Security, 1996(1):3–4, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896901555.

Anonymous:1996:NOS

[416] Anonymous. Novell’s operating systems forecast to remain dominant. Network Security, 1996 (1):4, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896901567.

Anonymous:1996:FOB

[417] Anonymous. Further online banking competition. Network Security, 1996(1):4, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896901579.

Anonymous:1996:NEC

[418] Anonymous. The new era of cable modems. Network Security, 1996 (1):4–5, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896901580.

Anonymous:1996:SST

[419] Anonymous. Spec for secure transactions. Network Security, 1996 (1):5, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901592.

Anonymous:1996:UDE

[420] Anonymous. Users demand encryption policies. Network Security, 1996(1):5, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901609.

Anonymous:1996:CGP

[421] Anonymous. Cable growth potential massive. Network Security, 1996(1):5, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901610.

Anonymous:1996:SDU

[422] Anonymous. Secure domain unit for network protection. Network Security, 1996(1):5–6, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901622.

Anonymous:1996:BDI

[423] Anonymous. Briefing document for Internet security. Network Security, 1996(1):6, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901634.

Anonymous:1996:ISA

[424] Anonymous. Internal security across client/server networks. Network Security, 1996(1):6, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901646.

Anonymous:1996:NSA

[425] Anonymous. Network security assessment tool for Windows NT. Network Security, 1996(1):7, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485896901658.

Kaneshige:1996:MGB

[426] Thomas Kaneshige. McAfee gives the boot to virus. Network Security, 1996(1):7, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589690166X.

Kaneshige:1996:NSM

[427] Thomas Kaneshige. Network security made easy with little black box. Network Security, 1996(1): 7–8, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901671.

English:1996:STI

[428] Erin English. SEC tackles Internet investment fraud. Network Security, 1996(1):8, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901683.

English:1996:LWM

[429] Erin English. Lotus working on military version of Notes. Network Security, 1996(1):8, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901695.

Cohen:1996:IHPa

[430] Fred Cohen. Internet holes — Part 5b: 50 ways to attack your Web systems. Network Security, 1996 (1):9–13, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800900011.

Kaneshige:1996:OSD

[431] Thomas Kaneshige. Oracle to secure database access over the Net. Network Security, 1996(1):9, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485896901701.

Hanson:1996:IAS

[432] Lesley Hanson. The impact of ATM on security in the data network. Network Security, 1996(1): 13–17, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800900023.

Lindup:1996:HH

[433] Ken Lindup. Hackers for hire. Network Security, 1996(1):18, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485896901713.

Anonymous:1996:Ea

[434] Anonymous. Events. Network Security, 1996(1):20, January 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896901725.

Anonymous:1996:NVS

[435] Anonymous. New version solves BIND security problems. Network Security, 1996(2):2, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900410.

Anonymous:1996:HSV

[436] Anonymous. HP syslog vulnerability. Network Security, 1996 (2):2, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900422.

Anonymous:1996:UPD

[437] Anonymous. UDP port denial-of-service attack. Network Security, 1996 (2):2–3, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900434.

Anonymous:1996:VSL

[438] Anonymous. Virus spreads like wild fire. Network Security, 1996 (2):3, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900446.

Anonymous:1996:WMVa

[439] Anonymous. Winword macro viruses. Network Security, 1996(2):3–4, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485896900458.

Anonymous:1996:SSP

[440] Anonymous. Singapore’s service providers confident of security despite break-in. Network Security, 1996(2):4, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589690046X.

Anonymous:1996:CED

[441] Anonymous. Compress and encrypt data simultaneously. Network Security, 1996(2):4–5, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900471.

Anonymous:1996:NGW

[442] Anonymous. The next Great Wall of China? Network Security, 1996(2):5, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900483.

Anonymous:1996:EPP

[443] Anonymous. Extended password protection for CyberGuard Firewall. Network Security, 1996 (2):5, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900495.

Anonymous:1996:IMP

[444] Anonymous. Internet may not be profitable. Network Security, 1996 (2):5–6, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900501.

Anonymous:1996:PSP

[445] Anonymous. The promise of secure phone banking. Network Security, 1996(2):6, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900513.

Anonymous:1996:RAP

[446] Anonymous. Remote access protection. Network Security, 1996 (2):6, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900525.

Anonymous:1996:SDI

[447] Anonymous. Secure deployment of Internet services. Network Security, 1996(2):6, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900537.

Anonymous:1996:HVL

[448] Anonymous. Humans and viruses locked out. Network Security, 1996 (2):6–7, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900549.

Sterlicchi:1996:WTI

[449] John Sterlicchi. Wayfarer turns Internet into LAN. Network Security, 1996(2):7, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900550.

Sterlicchi:1996:NMS

[450] John Sterlicchi. New moves to stimulate EDI growth. Network Security, 1996(2):7–8, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900562.

Sterlicchi:1996:FNF

[451] John Sterlicchi. First NT firewall soars in. Network Security, 1996 (2):8–9, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900574.

Cohen:1996:IHPb

[452] Fred Cohen. Internet holes — Part 6: Automated attack and defence. Network Security, 1996(2): 9–14, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800900035.

Osen:1996:SCI

[453] Janet Osen. Sex, crimes and the Internet: the Jake Baker case. Network Security, 1996(2):15–22, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800900047.

Anonymous:1996:Eb

[454] Anonymous. Events. Network Security, 1996(2):24, February 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900586.

Anonymous:1996:CIN

[455] Anonymous. Corrupt information from network servers. Network Security, 1996(3):2, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901154.

Anonymous:1996:SDC

[456] Anonymous. SunSoft demo CDs contain a security threatening flaw. Network Security, 1996(3):2, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485896901166.

Anonymous:1996:ISP

[457] Anonymous. Internet security purchase for HP. Network Security, 1996(3):2–3, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901178.

Anonymous:1996:CGS

[458] Anonymous. CompuServe in Germany suffers another inquiry. Network Security, 1996(3):3, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589690118X.

Anonymous:1996:GPT

[459] Anonymous. Governments pressed Tor agreement on encryption. Network Security, 1996(3):3, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901191.

Anonymous:1996:BCI

[460] Anonymous. Business class Internet service. Network Security, 1996(3):3–4, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901208.

Anonymous:1996:VMC

[461] Anonymous. Visa and MasterCard combine security specifications. Network Security, 1996(3):4, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348589690121X.

Anonymous:1996:IRR

[462] Anonymous. Integralis revenues rise 60%. Network Security, 1996(3): 4, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1353485896901221.

Anonymous:1996:SAB

[463] Anonymous. SecurID authentication for BayRS routers. Network Security, 1996(3):4, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896901233.

Anonymous:1996:SSE

[464] Anonymous. Security system for electronic cash. Network Security, 1996(3):4, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896901245.

Anonymous:1996:MII

[465] Anonymous. Microsoft’s Internet information server. Network Security, 1996(3):4–5, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896901257.

Anonymous:1996:CNE

[466] Anonymous. CyberGuard’s new enhancements. Network Security, 1996(3):5, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901269.

Anonymous:1996:FND

[467] Anonymous. The firewall with the new dual wall. Network Security, 1996(3):5, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901270.

Sterlicchi:1996:STA

[468] John Sterlicchi. Security technologies aid ITT transition. Network Security, 1996(3):5–6, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901282.

Savarnejad:1996:KGC

[469] Atoosa Savarnejad. Kerberos gets cracked. Network Security, 1996(3): 6, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485896901294.

Savarnejad:1996:IDS

[470] Atoosa Savarnejad. IBM to develop secure version. Network Security, 1996(3):6, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901300.

Cohen:1996:IHPc

[471] Fred Cohen. Internet holes — Part 7: The human side. Network Security, 1996(3):7–10, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800900059.

Moyer:1996:SMF

[472] Philip R. Moyer and E. Eugene Schultz. A systematic methodology for firewall penetration testing. Network Security, 1996(3):11–18, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800900060.

Osen:1996:FIAa

[473] Janet Osen. The Freedom of Information Act and the 2600 Club. Network Security, 1996(3):19, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485896901312.

Anonymous:1996:Ec

[474] Anonymous. Events 20. Network Security, 1996(3):20, March 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896901324.

Anonymous:1996:WJB

[475] Anonymous. Weaknesses in Java bytecode verifier. Network Security, 1996(4):2, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901737.

Anonymous:1996:DOD

[476] Anonymous. Digital OSF/1 dxconsole vulnerability. Network Security, 1996(4):2, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901749.

Anonymous:1996:ERM

[477] Anonymous. Encryption restrictions may be eased. Network Security, 1996(4):2–3, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901750.

Anonymous:1996:UGD

[478] Anonymous. UK Government debates key escrow. Network Security, 1996(4):3, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901762.

Anonymous:1996:IOB

[479] Anonymous. ITSEC and Orange Book standards merge. Network Security, 1996(4):3, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901786.

Savarnejad:1996:NHK

[480] Atoosa Savarnejad. Netscape hires KPMG. Network Security, 1996(4):

3, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485896901774.

Anonymous:1996:NGJ

[481] Anonymous. Netscape and GE’s joint Internet venture. Network Security, 1996(4):3–4, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901798.

Anonymous:1996:NOB

[482] Anonymous. Net opened to British businesses. Network Security, 1996 (4):4, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485896901804.

Anonymous:1996:SEL

[483] Anonymous. Secure electronic links to cross capital. Network Security, 1996(4):4, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901816.

Anonymous:1996:URS

[484] Anonymous. Unix rejected, security found wanting. Network Security, 1996(4):4, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901828.

Anonymous:1996:AOI

[485] Anonymous. Alliance offers integrated business solution. Network Security, 1996(4):4–5, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589690183X.

Anonymous:1996:SIP

[486] Anonymous. Secure Internet payments. Network Security, 1996(4):5, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485896901841.

Anonymous:1996:MSW

[487] Anonymous. More security for Windows NT. Network Security, 1996(4):5, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901853.

Anonymous:1996:WMVb

[488] Anonymous. Word macro viruses detected and eliminated. Network Security, 1996(4):5–6, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901865.

Savarnejad:1996:IAS

[489] Atoosa Savarnejad. IDC annual survey reveals extent of insecurity. Network Security, 1996(4):6, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485896836890.

Savarnejad:1996:UCO

[490] Atoosa Savarnejad. US company offers firewall challenge. Network Security, 1996(4):6–7, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896836907.

Hancock:1996:INS

[491] Bill Hancock. Improving network security through port switching. Network Security, 1996(4):7–8, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485896836919.

Cohen:1996:IHPd

[492] Fred Cohen. Internet holes — Part 8: Spam. Network Security, 1996(4):8–10, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896836920.

Wood:1996:ICR

[493] Charles Cresson Wood. An Internet Curmudgeon’s rants. Network Security, 1996(4):10–11, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896836932.

Osen:1996:BNB

[494] Janet Osen. Bad news for bulletin boards. Network Security, 1996 (4):11–13, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896836944.

Hancock:1996:CYS

[495] Bill Hancock. Can you social engineer your way into your network? Network Security, 1996 (4):14–15, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896836956.

Anonymous:1996:Ed

[496] Anonymous. Events. Network Security, 1996(4):16, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901877.

Anonymous:1996:WIN

[497] Anonymous. Warning issued on new type of Word macro virus. Network Security, 1996(5):2, May 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901051.

Anonymous:1996:EBS

[498] Anonymous. Encryption battle sees possible break through. Network Security, 1996(5):2, May 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901063.

Anonymous:1996:FMD

[499] Anonymous. Firewall market developments. Network Security, 1996(5):2, May 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901075.

Anonymous:1996:ENS

[500] Anonymous. Enhanced network security offered. Network Security, 1996(5):2–3, May 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901087.

Anonymous:1996:IEB

[501] Anonymous. The Internet-enabled backup solution. Network Security, 1996(5):3, May 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901099.

Anonymous:1996:PIS

[502] Anonymous. Private information sent safely over public frame relay network. Network Security, 1996(5):3, May 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901105.

Savarnejad:1996:CRW

[503] Atoosa Savarnejad. Californian representative warns against cyberfraud. Network Security, 1996(5):4, May 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485896901117.

Savarnejad:1996:PBC

[504] Atoosa Savarnejad. PGP becomes a company. Network Security, 1996(5):4, May 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901129.

Savarnejad:1996:COF

[505] Atoosa Savarnejad. Cisco offers free encryption technology. Network Security, 1996(5):4–5, May 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901130.

Cohen:1996:IHPe

[506] Fred Cohen. Internet holes — Part 9: IP address forgery and how to eliminate it. Network Security, 1996(5):5–12, May 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/135348589681909X.

Parker:1996:RCG

[507] Tom Parker. The role of cryptography in global communications. Network Security, 1996(5):13–17, May 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485896819106.

Peterson:1996:IAM

[508] A. Padgett Peterson. Internet attack mechanisms. Network Security, 1996(5):17–19, May 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896819118.

Anonymous:1996:Ee

[509] Anonymous. Events. Network Security, 1996(5):20, May 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896901142.

Anonymous:1996:ICD

[510] Anonymous. Interpreters in CGI bin directories. Network Security, 1996(6):2, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896901336.

Anonymous:1996:NPT

[511] Anonymous. NIS+ password tables left insecure. Network Security, 1996(6):2, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896901348.

Anonymous:1996:IDT

[512] Anonymous. IRIX desktop tool vulnerability. Network Security, 1996(6):2, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348589690135X.

Anonymous:1996:SHI

[513] Anonymous. Security hole in IIS software. Network Security, 1996 (6):2–3, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901361.

Anonymous:1996:EME

[514] Anonymous. Electronic mail encryption standards’ rivalry. Network Security, 1996(6):3, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901373.

Anonymous:1996:MDI

[515] Anonymous. Microsoft’s designs for Internet security. Network Security, 1996(6):3, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901385.

Anonymous:1996:MIS

[516] Anonymous. Managed Internet service. Network Security, 1996(6):4, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485896901397.

Anonymous:1996:IIS

[517] Anonymous. Internet/intranet security from Gradient and HP. Network Security, 1996(6):4, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901415.

Anonymous:1996:PRS

[518] Anonymous. Performing Right Society’s network secured. Network Security, 1996(6):4, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901403.

Anonymous:1996:PSG

[519] Anonymous. PC Security gets investment from Motorola. Network Security, 1996(6):4–5, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901427.

Anonymous:1996:OAS

[520] Anonymous. OpeN/2 authorization software. Network Security, 1996(6):5, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901439.

Anonymous:1996:LDC

[521] Anonymous. LAN diagnostic consultancy service. Network Security, 1996(6):5, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901440.

Anonymous:1996:SSS

[522] Anonymous. Single sign-on security for enterprise-wide systems. Network Security, 1996(6):5–6, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901452.

Anonymous:1996:FTN

[523] Anonymous. Firewall targeted for NetWare. Network Security, 1996(6):6, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901464.

Anonymous:1996:PDA

[524] Anonymous. Protection for dial access systems. Network Security, 1996(6):6, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901488.

Anonymous:1996:SAS

[525] Anonymous. SoftID authentication software. Network Security, 1996(6):6, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901476.

Anonymous:1996:AVS

[526] Anonymous. Anti-virus solution for network firewalls. Network Security, 1996(6):6–7, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589690149X.

Cohen:1996:IHPf

[527] Fred Cohen. Internet holes — Part 10: UDP viruses. Network Security, 1996(6):7–10, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896884510.

Peterson:1996:MSO

[528] A. Padgett Peterson. E-mail security — an oxymoron? Network Security, 1996(6):10–13, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896884522.

Schultz:1996:BRF

[529] E. Eugene Schultz. Building the right firewall. Network Security, 1996(6):13–17, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896884534.

Osen:1996:FIAb

[530] Janet Osen. The Freedom of Information Act and the 2600 Club. Network Security, 1996(6):17–19, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485896884546.

Anonymous:1996:Ef

[531] Anonymous. Events. Network Security, 1996(6):20, June 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901506.

Anonymous:1996:VDS

[532] Anonymous. Vulnerability in DEC software security kits. Network Security, 1996(7):2, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900896.

Anonymous:1996:VDP

[533] Anonymous. Vulnerability in the dip program. Network Security, 1996(7):2, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900902.

Anonymous:1996:SV

[534] Anonymous. Suidperl vulnerability. Network Security, 1996(7):2–3, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485896900914.

Anonymous:1996:RMC

[535] Anonymous. Right mouse click could allow in viruses. Network Security, 1996(7):3, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900926.

Anonymous:1996:SAO

[536] Anonymous. Security and authentication offered. Network Security, 1996(7):3–4, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900938.

Anonymous:1996:BGH

[537] Anonymous. Businesses get help to beat Internet crime. Network Security, 1996(7):4, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589690094X.

Anonymous:1996:CAI

[538] Anonymous. Companies adopt Internet commerce solutions. Network Security, 1996(7):4, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900951.

Anonymous:1996:TSN

[539] Anonymous. Total secure networking solutions incorporates CyberGuard. Network Security, 1996(7):4–5, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485896900963.

Anonymous:1996:CCC

[540] Anonymous. Credit-card company seeks strong encryption. Network Security, 1996(7):5, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900975.

Anonymous:1996:BPE

[541] Anonymous. Barclays’ pilot for electronic commerce. Network Security, 1996(7):5, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900987.

Anonymous:1996:LDI

[542] Anonymous. Librarians develop Internet protocol. Network Security, 1996(7):5–6, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900999.

Anonymous:1996:SVP

[543] Anonymous. Secure virtual private networks. Network Security, 1996(7):6, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901002.

Anonymous:1996:SLN

[544] Anonymous. Serverware launches new version of SeNTry. Network Security, 1996(7):6, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901014.

Anonymous:1996:II

[545] Anonymous. IronBridge and IronWall. Network Security, 1996(7):6–7, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485896901026.

Anonymous:1996:CMA

[546] Anonymous. CKS MyNet, the answer to single sign-on? Network Security, 1996(7):7, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896901038.

Sterlicchi:1996:SFM

[547] John Sterlicchi. Security framework from Microsoft. Network Security, 1996(7):7–8, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896837354.

Savarnejad:1996:RSM

[548] Atoosa Savarnejad. Revisions in SET mark advances to secure cyberspace shopping. Network Security, 1996(7):8, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896837366.

Cohen:1996:IHPg

[549] Fred Cohen. Internet holes — Part 11: Internet lightning rods. Network Security, 1996(7):9–13, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/135348589683738X.

Kaneshige:1996:MMC

[550] Thomas Kaneshige. McAfee and Microsoft clean up Internet. Network Security, 1996(7):9, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896837378.

Solomon:1996:EVP

[551] Alan Solomon. The end of the virus problem. Network Security, 1996(7):13–17, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896837391.

Lindup:1996:SAW

[552] Ken Lindup. Superman is alive and well and living in Wapping. Network Security, 1996(7):18, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485896837408.

Anonymous:1996:Eg

[553] Anonymous. Events. Network Security, 1996(7):20, July 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589690104X.

Anonymous:1996:SKP

[554] Anonymous. Solaris 2.5 KCMS programs vulnerability. Network Security, 1996(8):2, August 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900306.

Anonymous:1996:HUV

[555] Anonymous. HP-UX vulnerabilities. Network Security, 1996 (8):2, August 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900318.

Anonymous:1996:RV

[556] Anonymous. The rdist vulnerability. Network Security, 1996 (8):2–3, August 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348589690032X.

Anonymous:1996:FVC

[557] Anonymous. First virus capable of infecting Excel spreadsheet. Network Security, 1996(8):3, August 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485896900331.

Anonymous:1996:JMC

[558] Anonymous. Joint marketing for CyberGuard and EDS. Network Security, 1996(8):3, August 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900343.

Anonymous:1996:EWS

[559] Anonymous. Enterprise Web solutions from Netscape and NeXT. Network Security, 1996(8):3, August 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485896900355.

Anonymous:1996:UNS

[560] Anonymous. Usenet news via satellite. Network Security, 1996 (8):3–4, August 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900367.

Anonymous:1996:BFS

[561] Anonymous. Baysecure Framework’s security enhanced by ACE/ Server. Network Security, 1996 (8):4, August 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900379.

Anonymous:1996:NNVa

[562] Anonymous. New NT virus scanner. Network Security, 1996(8):4, August 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485896900380.

Anonymous:1996:CPT

[563] Anonymous. CryptoSystem protects TCP/IP networks. Network Security, 1996(8):4–5, August 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900392.

David:1996:IGY

[564] Jon David. The Internet — gateway to your information assets. Network Security, 1996 (8):5–7, August 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896811524.

Hancock:1996:OSA

[565] Bill Hancock. Operating system attacks from network resources. Network Security, 1996(8):8–11, August 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485896811536.

Cohen:1996:IHPh

[566] Fred Cohen. Internet holes — Part 12a: Internet incident response. Network Security, 1996 (8):12–14, August 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896811548.

Peterson:1996:PSR

[567] Padgett Peterson. Providing secure, recoverable e-mail. Network Security, 1996(8):15–19, August 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/135348589681155X.

Anonymous:1996:Eh

[568] Anonymous. Events. Network Security, 1996(8):20, August 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900409.

Anonymous:1996:VE

[569] Anonymous. Vulnerability in expreserve. Network Security, 1996(9):2, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1353485896900161.

Anonymous:1996:SPF

[570] Anonymous. Security problem in fm fls. Network Security, 1996 (9):2, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900173.

Anonymous:1996:VFN

[571] Anonymous. Virus fears for new operating systems. Network Security, 1996(9):2, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900185.

Anonymous:1996:MCM

[572] Anonymous. Merge creates mega security response team. Network Security, 1996(9):2–3, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900197.

Anonymous:1996:SBS

[573] Anonymous. Secure banking and shopping on Internet. Network Security, 1996(9):3, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900203.

Anonymous:1996:ISO

[574] Anonymous. Internet security offerings from HP. Network Security, 1996(9):3, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900215.

Anonymous:1996:ARB

[575] Anonymous. Agreement reached between Microsoft and RSA. Network Security, 1996(9):3–4, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485896900227.

Anonymous:1996:NCD

[576] Anonymous. Name change for Digital Pathways. Network Security, 1996(9):4, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900239.

Anonymous:1996:LAW

[577] Anonymous. On-LAN authentication for Windows NT. Network Security, 1996(9):4, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900252.

Anonymous:1996:DRR

[578] Anonymous. Digital resells Raptor’s security products. Network Security, 1996(9):4, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900240.

Anonymous:1996:EKE

[579] Anonymous. Encryption without key exchange. Network Security, 1996 (9):4–5, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900264.

Anonymous:1996:NER

[580] Anonymous. The next Eagle from Raptor. Network Security, 1996 (9):5, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896844011.

Anonymous:1996:FPM

[581] Anonymous. Further protection from e-mail viruses. Network Security, 1996(9):5, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896900276.

Persyn:1996:JTH

[582] Tim Persyn. Journalists take to hacking. Network Security, 1996 (9):5, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896900288.

Bee:1996:UGA

[583] Andrianne Bee. US government allows selling of strong encryption to US clients. Network Security, 1996 (9):6, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896844035.

Bee:1996:NPC

[584] Andrianne Bee. NCSA programme certifies Web sites. Network Security, 1996(9):6, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896844023.

Gordon:1996:GVS

[585] Sarah Gordon. Are good virus simulators still a bad idea? Network Security, 1996(9):7–13, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485896844047.

Cohen:1996:IHPi

[586] Fred Cohen. Internet holes — Part 12(b): Internet incident response. Network Security, 1996(9): 13–17, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896844059.

Osen:1996:VMP

[587] Janet Osen. The Virtual Magistrate [Project]. Network Security, 1996 (9):18, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896844060.

Anonymous:1996:Ei

[588] Anonymous. Events. Network Security, 1996(9):20, September 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589690029X.

Anonymous:1996:TSF

[589] Anonymous. TCP SYN flooding and IP spoofing attacks. Network Security, 1996(10):2, October 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900598.

Anonymous:1996:FSV

[590] Anonymous. Further sendmail vulnerabilities. Network Security, 1996 (10):2–3, October 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900604.

Anonymous:1996:IHV

[591] Anonymous. Irina, a hoax not a virus. Network Security, 1996 (10):3, October 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900616.

Anonymous:1996:CFIa

[592] Anonymous. Corporations fear ICE T virus emergence. Network Security, 1996(10):3, October 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900628.

Anonymous:1996:KEB

[593] Anonymous. Key escrow becomes key recovery. Network Security, 1996 (10):3–4, October 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589690063X.

Anonymous:1996:FDE

[594] Anonymous. Further doubts emerge over NT security. Network Security, 1996(10):4, October 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900641.

Anonymous:1996:PAA

[595] Anonymous. Privacy advocates not appeased by key recovery. Network Security, 1996(10):4, October 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485896900653.

Anonymous:1996:RCM

[596] Anonymous. Reflex creates macro virus trapper. Network Security, 1996(10):4, October 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900665.

Anonymous:1996:AFP

[597] Anonymous. The Authorizer firewall, protection for X25 networks. Network Security, 1996(10):5, October 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485896826493.

Anonymous:1996:VSX

[598] Anonymous. Virus scanner for X.400 messaging system. Network Security, 1996(10):5, October 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896900677.

Savarnejad:1996:MDA

[599] Atoosa Savarnejad. Multi-discipline alliance spans all security aspects. Network Security, 1996(10):5, October 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/1353485896900689.

Savarnejad:1996:NLF

[600] Atoosa Savarnejad. Netsolve launches first ISP-independent outsourcing product. Network Security, 1996 (10):6, October 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/135348589682650X.

Bee:1996:SEI

[601] Adrianne Bee. Soon even the Internet will take American Express. Network Security, 1996 (10):6–7, October 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485896826511.

Cohen:1996:IHPj

[602] Fred Cohen. Internet holes — Part 13: The SYN flood. Network Security, 1996(10):7–9, October 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896826523.

David:1996:SWS

[603] Jon David. Sniffers — the whole story. Network Security, 1996(10): 9–13, October 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896826535.

Osen:1996:CDAa

[604] Janet Osen. The Communications Decency Act and the karma of chaos — Part 1. Network Security, 1996 (10):13–19, October 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485896826547.

Anonymous:1996:Ej

[605] Anonymous. Events. Network Security, 1996(10):20, October 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900690.

Anonymous:1996:SDM

[606] Anonymous. Sendmail daemon mode vulnerability. Network Security, 1996 (11):2, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900008.

Anonymous:1996:HDC

[607] Anonymous. Hackers not deterred by certification scheme. Network Security, 1996(11):2, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589690001X.

Anonymous:1996:ISC

[608] Anonymous. Invasion of Supreme Court’s homepage. Network Security, 1996(11):2, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900021.

Anonymous:1996:NNVb

[609] Anonymous. Netscape Navigator virus scare. Network Security, 1996 (11):2, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900033.

Anonymous:1996:SRI

[610] Anonymous. Survey reveals Internetborne viruses on increase. Network Security, 1996(11):3, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897812226.

McCormack:1996:CSR

[611] Michael McCormack. Content selfregulation recommended. Network Security, 1996(11):3, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485896900045.

Anonymous:1996:CIP

[612] Anonymous. Combatants of Internet porn. Network Security, 1996 (11):3–4, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900057.

Anonymous:1996:ICH

[613] Anonymous. Internet commerce hampered by security fears. Network Security, 1996(11):4, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485896900069.

Anonymous:1996:SVN

[614] Anonymous. Smartcards vulnerable to new attack. Network Security, 1996(11):4, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900070.

Anonymous:1996:TWS

[615] Anonymous. TV Web-surfing device breaks export regulations. Network Security, 1996(11):4, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485896900082.

Anonymous:1996:UEM

[616] Anonymous. Users experience major problems with networks. Network Security, 1996(11):4–5, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485896900094.

Anonymous:1996:JCB

[617] Anonymous. Japanese company buys into security market. Network Security, 1996(11):5, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900100.

Anonymous:1996:SAI

[618] Anonymous. Secure access for Internet/intranet connectivity. Network Security, 1996(11):5, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485896900124.

Anonymous:1996:HPS

[619] Anonymous. Hacker-proof system for data coding. Network Security, 1996 (11):5, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900112.

Anonymous:1996:ITP

[620] Anonymous. Internet transaction processing solutions. Network Security, 1996(11):5–6, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900136.

Anonymous:1996:CFIb

[621] Anonymous. Cisco’s firewall for Internet/intranet. Network Security, 1996(11):6, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900148.

Bee:1996:HCS

[622] Adrianne Bee. How cyber-safe is your company? Network Security, 1996 (11):6–7, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897812238.

Bee:1996:VAA

[623] Adrianne Bee. Vendor alliances aim to make Internet safe for business. Network Security, 1996 (11):7, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348589781224X.

Bee:1996:CS

[624] Adrianne Bee. Cyber stamps. Network Security, 1996(11):7–8, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485897812251.

Kaneshige:1996:VRA

[625] Thomas Kaneshige. Vendors rally for answers to Panix Attacks. Network Security, 1996(11):8, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897812263.

Kaneshige:1996:MCR

[626] Thomas Kaneshige. Microsoft’s catapult raises NT firewall bar. Network Security, 1996(11):8–9, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897812275.

David:1996:FPI

[627] Jon David. The future of PGP on the Internet. Network Security, 1996 (11):9–12, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897812287.

Osen:1996:CDAb

[628] Janet Osen. The Communications Decency Act — Part 2: The language of the Communications Decency Act. Network Security, 1996 (11):12–18, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897812299.

Anonymous:1996:Ek

[629] Anonymous. Events. Network Security, 1996(11):20, November 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589690015X.

Anonymous:1996:HAV

[630] Anonymous. HP 9000 access vulnerability. Network Security, 1996 (12):2, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900707.

Anonymous:1996:PBR

[631] Anonymous. Password buffer overrun vulnerability. Network Security, 1996(12):2, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900719.

Anonymous:1996:FPV

[632] Anonymous. fpkg2swpkg program vulnerability. Network Security, 1996 (12):2, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900720.

Anonymous:1996:NBR

[633] Anonymous. newgrp buffer over-run vulnerability. Network Security, 1996 (12):2–3, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900732.

Anonymous:1996:CSC

[634] Anonymous. Computer security concerns ranked low in survey. Network Security, 1996(12):3, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485896900744.

Anonymous:1996:ISS

[635] Anonymous. Internet shopping secured by SET for Taiwanese. Network Security, 1996(12):3, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485896900756.

Anonymous:1996:HET

[636] Anonymous. Hardware encryption technology complies with encryption regulations. Network Security, 1996 (12):4, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900768.

Anonymous:1996:ECP

[637] Anonymous. Extending copyright protection for the Internet. Network Security, 1996(12):4, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348589690077X.

Anonymous:1996:IBU

[638] Anonymous. Increase in business users needing the Internet. Network Security, 1996(12):4–5, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900781.

Anonymous:1996:EFI

[639] Anonymous. Electronic fingerprint identifies owner of image. Network Security, 1996(12):5, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485896900793.

Anonymous:1996:GIC

[640] Anonymous. Global Internet commerce. Network Security, 1996 (12):5, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900811.

Anonymous:1996:PRI

[641] Anonymous. PICS recommended for Internet. Network Security, 1996 (12):5, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348589690080X.

Anonymous:1996:KSA

[642] Anonymous. Kane Security Analyst assesses network virus protection. Network Security, 1996 (12):5, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900823.

Anonymous:1996:EPD

[643] Anonymous. Electronic privacy device from Harris. Network Security, 1996(12):6, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900835.

Anonymous:1996:NSS

[644] Anonymous. New software security token. Network Security, 1996 (12):6, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900847.

Anonymous:1996:PEB

[645] Anonymous. Protection for electronic business communications. Network Security, 1996(12):6–7, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485896900859.

Anonymous:1996:ISW

[646] Anonymous. Internet security for Windows NT environments. Network Security, 1996(12):7, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485896900860.

Madsen:1996:CCL

[647] Wayne Madsen. Congressional Committee looks at US encryption policy. Network Security, 1996(12): 7–8, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485896900872.

Cohen:1996:MNS

[648] Fred Cohen. Managing network security — Part 1: How good do you have to be? Network Security, 1996 (12):9–11, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800900072.

David:1996:AI

[649] Jon David. Auditing the Internet. Network Security, 1996(12): 11–14, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800900084.

Gordon:1996:RWA

[650] Sarah Gordon and Richard Ford. Real world anti-virus product reviews and evaluations — Part 1. Network Security, 1996(12):14–18, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485800900096.

Anonymous:1996:El

[651] Anonymous. Events. Network Security, 1996(12):20, December 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485896900884.

Anonymous:1997:KSS

[652] Anonymous. Korn shell suid exec vulnerability. Network Security, 1997(1):2, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897834897.

Anonymous:1997:VSI

[653] Anonymous. Vulnerabilities in SGI IRIX. Network Security, 1997 (1):2, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897834885.

Anonymous:1997:DSA

[654] Anonymous. Denial of service attack via ping. Network Security, 1997(1):3, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897834903.

Anonymous:1997:CAA

[655] Anonymous. Computer attacks against WebCom. Network Security, 1997 (1):3–4, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897834915.

Anonymous:1997:GWS

[656] Anonymous. Government’s Web site attacked by hackers. Network Security, 1997(1):4, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897834939.

Anonymous:1997:TER

[657] Anonymous. Tightening of encryption regulations in Japan. Network Security, 1997(1):4, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897834927.

Anonymous:1997:ECR

[658] Anonymous. Encryption compromise on rocky ground. Network Security, 1997(1):4–5, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897834940.

Anonymous:1997:EPS

[659] Anonymous. Eagle provides security for Internet product. Network Security, 1997(1):5, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897834976.

Anonymous:1997:PUB

[660] Anonymous. Panel urges better protection. Network Security, 1997(1):5, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897834964.

Anonymous:1997:SVHa

[661] Anonymous. The season for virus hoaxes. Network Security, 1997 (1):5, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897834952.

Anonymous:1997:AAA

[662] Anonymous. Axent acquires AssureNet pathways. Network Security, 1997(1):5–6, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897834988.

Anonymous:1997:WSP

[663] Anonymous. Worldwide singlepoint security. Network Security, 1997(1):6, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897835015.

Anonymous:1997:ERA

[664] Anonymous. Enhanced remote access security. Network Security, 1997(1):6, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897835003.

Anonymous:1997:PKS

[665] Anonymous. Public key smartcard available. Network Security, 1997(1):6, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589783499X.

Anonymous:1997:DEU

[666] Anonymous. Desktop encryption utility. Network Security, 1997 (1):7, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897835040.

Anonymous:1997:SVHb

[667] Anonymous. Security for vulnerable high speed networks. Network Security, 1997(1):7, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897835039.

Huynh:1997:PCW

[668] Danthanh Huynh. Professor criticizes Web security. Network Security, 1997(1):7, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897835027.

Cohen:1997:MNSa

[669] Fred Cohen. Managing network security — Part 2: Where should we concentrate protection? Network Security, 1997(1):8–11, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897835052.

Gordon:1997:RWA

[670] Sarah Gordon and Richard Ford. Real world anti-virus product reviews and evaluations — Part 2. Network Security, 1997(1):11–18, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897835064.

Anonymous:1997:Ea

[671] Anonymous. Events. Network Security, 1997(1):20, January 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900350.

Anonymous:1997:VRT

[672] Anonymous. Vulnerability in rlogin/ Term. Network Security, 1997 (2):2, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897866342.

Anonymous:1997:MCB

[673] Anonymous. MIME conversion buffer overflow. Network Security, 1997 (2):2–3, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897866354.

Anonymous:1997:RBE

[674] Anonymous. RSA’s 40-bit encryption algorithm cracked. Network Security, 1997(2):3, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589786638X.

Anonymous:1997:TV

[675] Anonymous. talkd vulnerability. Network Security, 1997(2):3, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897866378.

Anonymous:1997:HOS

[676] Anonymous. The holes in online security. Network Security, 1997 (2):3, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897866366.

Anonymous:1997:ECH

[677] Anonymous. Electronic commerce hindered by fear of crime. Network Security, 1997(2):3–4, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485897866391.

Anonymous:1997:UDC

[678] Anonymous. US DoD COTS snafu. Network Security, 1997 (2):4, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897866408.

Anonymous:1997:CFG

[679] Anonymous. Companies form group to support cross-platform encryption. Network Security, 1997 (2):4, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/1353485897900593.

Anonymous:1997:EIC

[680] Anonymous. Exploit intranets without compromising security. Network Security, 1997(2):4–5, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348589786641X.

Anonymous:1997:PCI

[681] Anonymous. Power-cached Internet/ intranet gateway. Network Security, 1997(2):5, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897866445.

Anonymous:1997:SEI

[682] Anonymous. Security for enterprise Internet applications. Network Security, 1997(2):5, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897866433.

Anonymous:1997:AVP

[683] Anonymous. Anti-virus protection for Office 97. Network Security, 1997(2):5, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897866421.

Anonymous:1997:VFL

[684] Anonymous. Virus found on Linux system. Network Security, 1997 (2):5–6, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897866457.

Bee:1997:LE

[685] Adrianne Bee. The latest on encryption. Network Security, 1997 (2):6–7, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897866469.

Savarnejad:1997:IFS

[686] Atoosa Savarnejad. Intranet firewalls susceptible to internal attacks. Network Security, 1997 (2):7, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897866482.

Savarnejad:1997:WIF

[687] Atoosa Savarnejad. Warning issued, flaw in Unix software. Network Security, 1997(2):7, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897866470.

Schultz:1997:WFF

[688] E. Eugene Schultz. When firewalls fail: Lessons learned from firewall testing. Network Security, 1997 (2):8–11, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897866494.

Cohen:1997:MNSb

[689] Fred Cohen. Managing network security — Part 3: Network security as a control issue. Network Security, 1997 (2):12–15, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897866500.

Aubrey-Jones:1997:IV

[690] David Aubrey-Jones. Internet — virusnet? Network Security, 1997 (2):15–19, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897866512.

Anonymous:1997:Eb

[691] Anonymous. Events. Network Security, 1997(2):20, February 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589790060X.

Anonymous:1997:BFJ

[692] Anonymous. Bug found in Java Virtual Machine. Network Security, 1997(3):2, March 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897830334.

Anonymous:1997:FFI

[693] Anonymous. Flaw found in Internet Explorer. Network Security, 1997(3):2, March 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897830322.

Anonymous:1997:HNT

[694] Anonymous. httpd  nph-test-cgi

script vulnerability. Network Security, 1997(3):2, March 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897830309.

Anonymous:1997:EKB

[695] Anonymous. Encryption key of 48 bits cracked. Network Security, 1997(3):2, March 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897830310.

Anonymous:1997:FCR

[696] Anonymous. French companies restricted from using high-end encryption. Network Security, 1997(3): 3, March 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S135348589783036X.

Anonymous:1997:PEH

[697] Anonymous. PGP encryption heavily criticized. Network Security, 1997(3):3, March 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897830358.

Anonymous:1997:VIS

[698] Anonymous. Vulnerability in Inter-NetNews server. Network Security, 1997(3):3, March 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897830346.

Szweda:1997:FIA

[699] Roy Szweda. Finjan introduces ActiveX Security Manager. Network Security, 1997(3):3–4, March 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897830371.

Anonymous:1997:ASA

[700] Anonymous. ACE/Server access manager launched. Network Security, 1997(3):4, March 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900337.

Szweda:1997:NSN

[701] Roy Szweda. NCSA supports new TIS ForceField. Network Security, 1997(3):4–5, March 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897830383.

Savarnejad:1997:ILS

[702] Atoosa Savarnejad. IBM launches SecureWay. Network Security, 1997(3): 5, March 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485897830395.

Savarnejad:1997:ERS

[703] Atoosa Savarnejad. eTrust to roll out in second quarter. Network Security, 1997(3):5–6, March 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897830401.

Hancock:1997:UCE

[704] Bill Hancock. The US cryptographic export debate — round five? Network Security, 1997 (3):6–7, March 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897830413.

Cohen:1997:MNSc

[705] Fred Cohen. Managing network security — Part 4: Integrity first, usually. Network Security, 1997 (3):8–10, March 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897830425.

Osen:1997:MMM

[706] Janet Osen. The medium is the message: a media specific analysis of the Communications Decency Act. Network Security, 1997 (3):10–16, March 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897830437.

Osen:1997:WMT

[707] Janet Osen. For whom the modem tolls: The travails of America Online. Network Security, 1997 (3):17–19, March 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897830449.

Anonymous:1997:Ec

[708] Anonymous. Events. Network Security, 1997(3):20, March 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900349.

Anonymous:1997:ESS

[709] Anonymous. Explorer security still patchy. Network Security, 1997(4):2, April 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897899928.

Anonymous:1997:VIP

[710] Anonymous. Vulnerability in IMAP and POP. Network Security, 1997 (4):2, April 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485897899916.

Anonymous:1997:CAN

[711] Anonymous. Current attacks on news servers. Network Security, 1997(4):2–3, April 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589789993X.

Anonymous:1997:VKI

[712] Anonymous. VirusPatrol keeps Internet users virus free. Network Security, 1997(4):3, April 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897899953.

Szweda:1997:RSN

[713] Roy Szweda. Rainbow supports Netscape’s open multi-platform security. Network Security, 1997(4):3, April 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485897899941.

Anonymous:1997:SEA

[714] Anonymous. Strong encryption available worldwide. Network Security, 1997(4):3–4, April 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897899965.

Anonymous:1997:MSJ

[715] Anonymous. MicroLan systems joins CSAPP. Network Security, 1997(4):4, April 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897899977.

Szweda:1997:AMD

[716] Roy Szweda. AT&T/Mondex deal enables Web ‘micropayments’. Network Security, 1997(4):4, April 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/1353485897900313.

Anonymous:1997:FSU

[717] Anonymous. File security uses smartcard and RSA encryption. Network Security, 1997(4):5, April 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897899990.

Anonymous:1997:ESF

[718] Anonymous. Enhanced security features for Netscape web servers. Network Security, 1997(4):5, April 1997.

CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897899989.

Anonymous:1997:SVP

[719] Anonymous. Secure virtual private networking over the Internet. Network Security, 1997(4):5–6, April 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897800001.

Anonymous:1997:MRW

[720] Anonymous. MIMEsweeper redefines web/e-mail security. Network Security, 1997(4):6, April 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897800037.

Anonymous:1997:LDA

[721] Anonymous. Lotus Domino antivirus solution. Network Security, 1997(4):6, April 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897800025.

Anonymous:1997:EAS

[722] Anonymous. Easy authentication and signature verification. Network Security, 1997(4):6, April 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897800013.

Savarnejad:1997:HYC

[723] Atoosa Savarnejad. Having your cookies and eating them too. Network Security, 1997(4):6–7, April 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897800049.

Savarnejad:1997:ITS

[724] Atoosa Savarnejad. IBM and TIS strike an agreement. Network Security, 1997(4):7–8, April 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897800050.

Savarnejad:1997:CAI

[725] Atoosa Savarnejad. Cisco allies with industry leaders for security initiative. Network Security, 1997(4):8, April 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897800062.

Moyer:1997:EFI

[726] Philip R. Moyer. Enhanced firewall infrastructure testing methodology. Network Security, 1997 (4):9–15, April 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897800074.

Cohen:1997:MNSd

[727] Fred Cohen. Managing network security — Part 5: Risk management or risk analysis. Network Security, 1997(4):15–19, April 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897800086.

Anonymous:1997:Ed

[728] Anonymous. Events. Network Security, 1997(4):20, April 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900325.

Anonymous:1997:VLa

[729] Anonymous. Vulnerability in libXt. Network Security, 1997(5):2, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897856875.

Anonymous:1997:NLS

[730] Anonymous. Natural language service libraries vulnerability. Network Security, 1997(5):2–3, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897856887.

Anonymous:1997:JAS

[731] Anonymous. Java applets that steal computer cycles. Network Security, 1997(5):3, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897856905.

Anonymous:1997:VPA

[732] Anonymous. Vulnerabilities in pluggable authentication module. Network Security, 1997(5):3, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897856899.

Anonymous:1997:HPW

[733] Anonymous. High profile Web sites warrant high security. Network Security, 1997(5):3–4, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897856917.

Anonymous:1997:NSS

[734] Anonymous. NT systems shut down by Telnet hack. Network Security, 1997(5):4, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897856930.

Anonymous:1997:SCI

[735] Anonymous. Secure e-commerce for insurers. Network Security, 1997(5):4, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897856929.

Anonymous:1997:MCM

[736] Anonymous. Microsoft closes email service. Network Security, 1997(5):4, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485897900362.

Anonymous:1997:CPF

[737] Anonymous. Cellular phone fraud prevention working? Network Security, 1997(5):4–5, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897856942.

Anonymous:1997:IUT

[738] Anonymous. Internet used as Trojan horse. Network Security, 1997(5):5, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900374.

Anonymous:1997:PDI

[739] Anonymous. Providers divided over Internet proposal. Network Security, 1997(5):5, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900386.

Anonymous:1997:IFN

[740] Anonymous. Intranets to far outnumber Internet. Network Security, 1997(5):5–6, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900398.

Anonymous:1997:NFM

[741] Anonymous. Networks flooded with e-mail messages. Network Security, 1997(5):6, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900404.

Anonymous:1997:GAV

[742] Anonymous. Global anti-virus research programme launched. Network Security, 1997(5):6, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900416.

Anonymous:1997:INSa

[743] Anonymous. Integrated network security products. Network Security, 1997(5):6, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900428.

Anonymous:1997:MTU

[744] Anonymous. Multi-tier Unix security product offered. Network Security, 1997(5):7, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589790043X.

Anonymous:1997:NDS

[745] Anonymous. New detection system takes on multi-million phone fraud. Network Security, 1997(5):7–8, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897900441.

Anonymous:1997:DRN

[746] Anonymous. Detection and repair for new online virus. Network Security, 1997(5):8, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900453.

Anonymous:1997:RAS

[747] Anonymous. Remote access solution to encrypt data. Network Security, 1997(5):8, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900465.

Sechrist:1997:IPC

[748] Douglas Sechrist. Internet phone calls breaking through corporate firewalls. Network Security, 1997(5):8–9, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897900477.

Savarnejad:1997:SFN

[749] Atoosa Savarnejad. Sparks fly over NT security. Network Security, 1997(5):9–10, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900490.

Breitrose:1997:HHI

[750] Charlie Breitrose. Hackers hit ISPs by the dozens. Network Security, 1997(5):9, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900489.

Savarnejad:1997:ANS

[751] Atoosa Savarnejad. Another NT security hole discovered. Network Security, 1997(5):10–11, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900507.

Capen:1997:CPA

[752] Ian Capen. Check Point announces security software alliance. Network Security, 1997(5):11, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900519.

Anonymous:1997:CTL

[753] Anonymous. To catch a thief … unmasking the telecoms fraudster. Network Security, 1997 (5):12–13, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900520.

Cohen:1997:MNSe

[754] Fred Cohen. Managing network security — Part 6: Computer games and network security management. Network Security, 1997 (5):14–18, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900532.

Lindup:1997:WLM

[755] Ken Lindup. Would Lenin make a good security officer? Network Security, 1997(5):19, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900544.

Anonymous:1997:Ee

[756] Anonymous. Events. Network Security, 1997(5):20, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900556.

Anonymous:1997:MV

[757] Anonymous. Metamail vulnerability. Network Security, 1997(6):2, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897895141.

Anonymous:1997:IRP

[758] Anonymous. IRIX runpriv program vulnerability. Network Security, 1997(6):2–3, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897895153.

Anonymous:1997:VSX

[759] Anonymous. Vulnerability with Solaris 2.x PC buffer overflow. Network Security, 1997(6):3, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897895165.

Savarnejad:1997:ISC

[760] Atoosa Savarnejad. ISP security consortium formed. Network Security, 1997(6):4, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897895189.

Anonymous:1997:IO

[761] Anonymous. Internet is on its own. Network Security, 1997(6):4, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485897895177.

Anonymous:1997:NSA

[762] Anonymous. NSA sued for algorithm. Network Security, 1997(6):4, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485897900076.

Anonymous:1997:ECD

[763] Anonymous. Experts claim DTI’s scheme has catastrophic flaws. Network Security, 1997(6):4–5, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485897895190.

Anonymous:1997:WDS

[764] Anonymous. Worldwide debut for SKIP. Network Security, 1997(6): 5, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1353485897895220.

Anonymous:1997:BUF

[765] Anonymous. Breakthrough for UK firm with US encryption restrictions. Network Security, 1997(6):5, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897895219.

Anonymous:1997:LDR

[766] Anonymous. Low data rate communications security improved. Network Security, 1997(6):5, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897895207.

Anonymous:1997:DPM

[767] Anonymous. Data protected at most finite level. Network Security, 1997(6):5–6, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897895232.

Anonymous:1997:EAJ

[768] Anonymous. Encryption and authentication for Java. Network Security, 1997(6):6, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897895256.

Anonymous:1997:NBF

[769] Anonymous. NT-based firewall for mixed TCP/IP and IPX networks. Network Security, 1997(6):6, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897895244.

Savarnejad:1997:PTU

[770] Atoosa Savarnejad. Princeton team uncovers flaw in Java. Network Security, 1997(6):6–7, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897895268.

Savarnejad:1997:AMS

[771] Atoosa Savarnejad. AOL4FREE makes second round as Trojan horse infection. Network Security, 1997(6):7, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897895281.

Breitrose:1997:TMO

[772] Charlie Breitrose. Trend micro offers free virus protection from the Web. Network Security, 1997(6):7, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348589789527X.

Hansen:1997:NIS

[773] Lesley Hansen. Network infrastructure security. Network Security, 1997(6):8–12, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897895293.

Cohen:1997:MNSf

[774] Fred Cohen. Managing network security — Part 7: Prevent, detect and respond. Network Security, 1997(6):13–16, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589789530X.

Saxby:1997:PCI

[775] Stephen Saxby. Policies for cyberspace — illegal content on the Internet. Network Security, 1997 (6):16–19, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897895311.

Anonymous:1997:Ef

[776] Anonymous. Events. Network Security, 1997(6):20, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900088.

Breitrose:1997:CHG

[777] Charlie Breitrose. Computer hacking glitch for Ramsey murder investigators. Network Security, 1997(7):2, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897898625.

Anonymous:1997:NNS

[778] Anonymous. Netscape Navigator security problem. Network Security, 1997(7):2, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897898613.

Anonymous:1997:SSA

[779] Anonymous. Solaris solstice Admin-Suite vulnerabilities. Network Security, 1997(7):3, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897898637.

Anonymous:1997:VP

[780] Anonymous. Vulnerability in the at(1) program. Network Security, 1997(7):3–4, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897898649.

Szweda:1997:EYL

[781] Roy Szweda. Ernst & Young LLP — WheelGroup Corp. alliance. Network Security, 1997(7):4, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485897898662.

Anonymous:1997:SCVa

[782] Anonymous. SunOS chkey vulnerability. Network Security, 1997(7):4, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485897898650.

Anonymous:1997:API

[783] Anonymous. Alliance provides Internet security to large enterprises. Network Security, 1997(7):4–5, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485897898674.

Savarnejad:1997:MOE

[784] Atoosa Savarnejad. Microsoft to offer enhanced firewall features in new server. Network Security, 1997(7):5, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897898698.

Anonymous:1997:LTR

[785] Anonymous. Lucent Technologies to resell Raptor’s firewalls. Network Security, 1997(7):5, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897898686.

Anonymous:1997:AUH

[786] Anonymous. AOL users hit further problems with hackers. Network Security, 1997(7):6, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897898716.

Szweda:1997:STF

[787] Roy Szweda. Sun takes on Feds over US encryption regulations. Network Security, 1997(7):6, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897898704.

Breitrose:1997:MJW

[788] Charlie Breitrose. Microsoft joins Web security consortium. Network Security, 1997(7):6–7, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897898728.

Szweda:1997:HFS

[789] Roy Szweda. Holographic fingerprint security device. Network Security, 1997(7):7, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897898741.

Anonymous:1997:UEL

[790] Anonymous. US export licence for 128-bit encryption for Microsoft. Network Security, 1997(7):7, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348589789873X.

Szweda:1997:ESW

[791] Roy Szweda. Encryption software for Windows. Network Security, 1997(7):8, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897898777.

Anonymous:1997:SPO

[792] Anonymous. SessionWall-3 protects organizations from abuse. Network Security, 1997(7):8, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897898765.

Anonymous:1997:EIM

[793] Anonymous. Enhancements to improve model for signing Internet code.

Network Security, 1997(7):8, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897898753.

Savarnejad:1997:GAD

[794] Atoosa Savarnejad. Group asks for disclosure of travel records of crypto czar. Network Security, 1997(7):8–9, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897898789.

Breitrose:1997:ALD

[795] Charlie Breitrose. American legislators debate encryption laws. Network Security, 1997(7):9–10, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897898790.

Breitrose:1997:WSG

[796] Charlie Breitrose. Web security group unveils certification programme. Network Security, 1997(7):10, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897898807.

Madsen:1997:KEE

[797] Wayne Madsen. Key Escrow Encryption Bill hits Congress. Network Security, 1997(7):11–12, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897898819.

Cohen:1997:MNSg

[798] Fred Cohen. Managing network security — Part 8: Relativistic risk assessment. Network Security, 1997 (7):13–14, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897898820.

Schultz:1997:BRIa

[799] Eugene Schultz. Book review: Internet security: Risk analysis, strategies, and firewalls: Othmar Kyas, International Thomson Computer Press, 1997. Network Security, 1997(7):15, July 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897898832.

Anonymous:1997:WNN

[800] Anonymous. Windows NT NtOpen-ProcessToken vulnerability. Network Security, 1997(8):2, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485897832242.

Anonymous:1997:VLb

[801] Anonymous. Vulnerability in ld.so. Network Security, 1997(8):2, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485897832230.

Anonymous:1997:SVL

[802] Anonymous. Security vulnerability in Lynx temporary files and LYDownload.c. Network Security, 1997(8):3, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897832254.

Sterlicchi:1997:ASH

[803] John Sterlicchi. Anti-spam hacker succeeds. Network Security, 1997 (8):3–4, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897832266.

Anonymous:1997:JBG

[804] Anonymous. Java bugs give openings to hackers. Network Security, 1997(8):4, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589783228X.

Anonymous:1997:NFL

[805] Anonymous. Netscape fixes latest Navigator bug. Network Security, 1997(8):4, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897832278.

Anonymous:1997:INSb

[806] Anonymous. INN news server vulnerabilities. Network Security, 1997 (8):4–5, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897832291.

Anonymous:1997:PAL

[807] Anonymous. PGP allowed limited export. Network Security, 1997(8):5, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897832321.

Anonymous:1997:BIU

[808] Anonymous. Bank of Ireland uses SET first in Europe. Network Security, 1997(8):5, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348589783231X.

Anonymous:1997:CJJ

[809] Anonymous. Cisco joins Java security alliance. Network Security, 1997(8):5, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897832308.

Szweda:1997:CGE

[810] Roy Szweda. CyberGuard gains E3 certification from Australian Government. Network Security, 1997 (8):5–6, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897832333.

Anonymous:1997:NSM

[811] Anonymous. New security modem announced. Network Security, 1997(8):6, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897832357.

Anonymous:1997:SAD

[812] Anonymous. Secure access to dedicated Web pages. Network Security, 1997(8):6, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897832345.

Anonymous:1997:SAP

[813] Anonymous. SecurlD authentication protects corporate information. Network Security, 1997 (8):6–7, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897832369.

Anonymous:1997:GPK

[814] Anonymous. Global public key security for PCs. Network Security, 1997(8):7, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897832382.

Szweda:1997:SCU

[815] Roy Szweda. Secure Computing unveils NT firewall. Network Security, 1997(8):7, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897832370.

Hancock:1997:VPN

[816] Bill Hancock. Virtual private networks: What, why, when, where and how. Network Security, 1997 (8):8–11, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897832394.

Cohen:1997:MNSh

[817] Fred Cohen. Managing network security — Part 9: Penetration testing? Network Security, 1997 (8):12–15, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897832400.

Shaw:1997:VVA

[818] Andrew Shaw. Voice verification — authenticating remote users over the telephone. Network Security, 1997 (8):16–18, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897832412.

Schultz:1997:BRIb

[819] E. Eugene Schultz. Book review: Intranet firewalls: Scott Fuller and Kevin Pagan, Ventana Press, 1997. Network Security, 1997(8):19, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897832424.

Anonymous:1997:Eg

[820] Anonymous. Events. Network Security, 1997(8):20, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900064.

Anonymous:1997:SCP

[821] Anonymous. Security compromise via procfs. Network Security, 1997 (9):2, September 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589790009X.

Anonymous:1997:SSI

[822] Anonymous. SunOS, Solaris ifconfig ioctls vulnerability. Network Security, 1997(9):2, September 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900106.

Anonymous:1997:SIW

[823] Anonymous. SGI, IRIX webdist.cgi, handler and wrap programs vulnerabilities. Network Security, 1997 (9):2–3, September 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900118.

Szweda:1997:ASC

[824] Roy Szweda. AT&T sues Connect America and others for Internet scam. Network Security, 1997 (9):3, September 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348589790012X.

Anonymous:1997:FRE

[825] Anonymous. Further relaxing of export regulations makes Internet more secure. Network Security, 1997 (9):3, September 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900131.

Szweda:1997:CAR

[826] Roy Szweda. CNN asked to remove Web link. Network Security, 1997(9):4, September 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900143.

Anonymous:1997:RPR

[827] Anonymous. RSA publishes RC2 specification. Network Security, 1997 (9):4, September 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900155.

Szweda:1997:PFT

[828] Roy Szweda. Playboy files Telecom Act protest. Network Security, 1997 (9):4–5, September 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900167.

Anonymous:1997:CPS

[829] Anonymous. Companies provide secure Internet roaming capabilities. Network Security, 1997(9): 5, September 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900179.

Anonymous:1997:CJA

[830] Anonymous. CyberGuard joins alliance programme in Singapore. Network Security, 1997(9):5, September 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897900180.

Anonymous:1997:MTN

[831] Anonymous. Multi-tier network security from McAfee. Network Security, 1997(9):5, September 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900192.

Anonymous:1997:HDR

[832] Anonymous. HTML-driven remote authentication. Network Security, 1997(9):6, September 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900209.

Anonymous:1997:PPN

[833] Anonymous. Product protects network from Internet-based attacks. Network Security, 1997 (9):6, September 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900210.

Savarnejad:1997:MFU

[834] Atoosa Savarnejad. E-mail firewall uses S/MIME protocol. Network Security, 1997(9):6–7, September 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897802528.

Breitrose:1997:DEM

[835] Charlie Breitrose. Distributing encrypted messages more securely. Network Security, 1997(9):7, September 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348589780253X.

Cohen:1997:MNSi

[836] Fred Cohen. Managing network security — Part 10: Change your password. Network Security, 1997 (9):8–11, September 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897802541.

Osen:1997:BBF

[837] Janet Osen. In the battle of the bits — final score: Internet 1, Congress 0. Network Security, 1997 (9):12–16, September 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897802553.

Schwartau:1997:WEIa

[838] Winn Schwartau. What exactly is information warfare? — Part 1. Network Security, 1997(9):17– 19, September 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897802565.

Anonymous:1997:Eh

[839] Anonymous. Events. Network Security, 1997(9):20, September 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900222.

Anonymous:1997:BUN

[840] Anonymous. Businesses using Net could face blackmail. Network Security, 1997(10):2, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900945.

Anonymous:1997:BOP

[841] Anonymous. Buffer overflow problem in rdist. Network Security, 1997(10):2, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900957.

Anonymous:1997:SIL

[842] Anonymous. SGI IRIX LOCK-OUT and login/scheme vulnerabilities. Network Security, 1997 (10):2–3, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900969.

Anonymous:1997:MEP

[843] Anonymous. Ministry’s encryption policy criticized. Network Security, 1997(10):3–4, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900970.

Szweda:1997:CBS

[844] Roy Szweda. CompuServe barbecues spam. Network Security, 1997 (10):4, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900982.

Szweda:1997:ECF

[845] Roy Szweda. Encrypted communications frustrate FBI. Network Security, 1997(10):4, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900994.

Anonymous:1997:ESP

[846] Anonymous. Every Swede to have personal e-mail address. Network Security, 1997(10):4–5, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897901008.

Anonymous:1997:CSI

[847] Anonymous. Completely secure Internet connection on offer. Network Security, 1997(10):5, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348589790101X.

Anonymous:1997:ISB

[848] Anonymous. Internet security for business to business applications. Network Security, 1997(10):5, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897901021.

Anonymous:1997:JAR

[849] Anonymous. Java applets run from a ‘cage’. Network Security, 1997 (10):5–6, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897901033.

Anonymous:1997:MLE

[850] Anonymous. Military-level encryption for all data transmissions. Network Security, 1997 (10):6, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897901045.

Breitrose:1997:VEM

[851] Charlie Breitrose. Virtual envelope for e-mail. Network Security, 1997(10):6, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897901057.

Anonymous:1997:HPS

[852] Anonymous. High performance solution to IP security needs. Network Security, 1997(10):6–7, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485897901069.

Anonymous:1997:JLS

[853] Anonymous. JCP launches secure Internet transactions protocol. Network Security, 1997(10): 7–8, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897857327.

Wehde:1997:NFB

[854] E. D. Wehde. Netscape fixes bugs. Network Security, 1997 (10):8, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897857339.

Wehde:1997:MHB

[855] E. D. Wehde. Microsoft helps banks. Network Security, 1997 (10):8–9, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897857340.

Wehde:1997:UGT

[856] E. D. Wehde. US gets tough on encryption. Network Security, 1997 (10):9–10, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897857352.

Schultz:1997:HPT

[857] Eugene Schultz. Hackers and penetration testing. Network Security, 1997(10):10, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897857364.

Lindup:1997:TOL

[858] Ken Lindup. Take one large bomb. Network Security, 1997(10): 10–11, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897857376.

Schwartau:1997:WEIb

[859] Winn Schwartau. What exactly is information warfare? — Part 2. Network Security, 1997(10): 12–16, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897857388.

Cohen:1997:MNSj

[860] Fred Cohen. Managing network security — Part 11: To outsource or not to outsource, that is the question! Network Security, 1997 (10):17–19, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589785739X.

Anonymous:1997:Ei

[861] Anonymous. Events. Network Security, 1997(10):20, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897901070.

Anonymous:1997:SCVb

[862] Anonymous. sysdef command vulnerability. Network Security, 1997 (11):2, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900611.

Anonymous:1997:FRV

[863] Anonymous. ftpd/rlogind vulnerabilities. Network Security, 1997 (11):2, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900623.

Anonymous:1997:VR

[864] Anonymous. Vulnerability in rlogin. Network Security, 1997(11):2–3, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1353485897900635.

Szweda:1997:FFY

[865] Roy Szweda. Flaw found in Yahoo email security …. Network Security, 1997(11):3, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900647.

Szweda:1997:LJE

[866] Roy Szweda. … and Jabadoo in Explorer 4 browser. Network Security, 1997(11):3, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900659.

Anonymous:1997:OGO

[867] Anonymous. Open Group OSF/ DCE denial-of-service vulnerability. Network Security, 1997(11): 3–4, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900660.

Anonymous:1997:WPU

[868] Anonymous. White paper urges risk assessment before connection. Network Security, 1997(11):4, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897900672.

Anonymous:1997:WFS

[869] Anonymous. World’s first SET completed. Network Security, 1997 (11):5, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900684.

Anonymous:1997:PHL

[870] Anonymous. Passive hacking, the latest craze. Network Security, 1997 (11):5, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900696.

Anonymous:1997:MNG

[871] Anonymous. McAfee and Network General combine. Network Security, 1997(11):5, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900702.

Anonymous:1997:FCB

[872] Anonymous. Four continents bridged by secured VPN. Network Security, 1997(11):5–6, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900714.

Anonymous:1997:SSR

[873] Anonymous. SecurVPN for secure remote access. Network Security, 1997 (11):6, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900726.

Anonymous:1997:ISM

[874] Anonymous. Industrial strength e-mail virus protection. Network Security, 1997(11):6, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900738.

Anonymous:1997:NPP

[875] Anonymous. New ‘plug and play’ cryptographic accelerators available. Network Security, 1997(11):6, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348589790074X.

Anonymous:1997:EW

[876] Anonymous. Encryption for Windows 95. Network Security, 1997 (11):7, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900751.

Anonymous:1997:AMS

[877] Anonymous. Applet Management System for secure Java implementation. Network Security, 1997 (11):7, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900763.

Rees:1997:AGO

[878] Frank Rees. Australian Government obstructs the export of revolutionary encryption. Network Security, 1997 (11):7–8, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900775.

Cohen:1997:MNSk

[879] Fred Cohen. Managing network security — Part 13: The network security game. Network Security, 1997 (11):9–11, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348589787328X.

Schwartau:1997:WEIc

[880] Winn Schwartau. What exactly is information warfare? — Part 3. Network Security, 1997(11): 12–18, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897873291.

Dwan:1997:BRW

[881] Berni Dwan. Book review: Web psychos, stalkers and pranksters how to protect yourself in cyberspace: Banks, Michael A., Coriolis Group Books, $24.99. Network Security, 1997(11): 19–20, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897873308.

Anonymous:1997:SIV

[882] Anonymous. SGI IRIX vulnerabilities. Network Security, 1997 (12):2, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900787.

Anonymous:1997:VYN

[883] Anonymous. View your neighbours’ data using Wave. Network Security, 1997(12):2–3, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897900799.

Anonymous:1997:VGW

[884] Anonymous. Vulnerabilities in GlimpseHTTP and WebGlimpse. Network Security, 1997(12):3, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897900805.

Anonymous:1997:AEA

[885] Anonymous. Attacker executes arbitrary commands on WWW server. Network Security, 1997(12):3–4, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897900817.

Szweda:1997:FMC

[886] Roy Szweda. Firewall market continues impressive growth. Network Security, 1997(12):4, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900829.

Szweda:1997:BDP

[887] Roy Szweda. Billion dollar programme to fight cyber threats. Network Security, 1997(12):4–5, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897900830.

Anonymous:1997:DHE

[888] Anonymous. Diffie–Hellman encryption freely available. Network Security, 1997(12):5, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900842.

Anonymous:1997:NHC

[889] Anonymous. Net hijacker caught by FBI. Network Security, 1997 (12):5, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900854.

Anonymous:1997:VPS

[890] Anonymous. VPN provides secure remote access. Network Security, 1997 (12):5, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900866.

Anonymous:1997:SOB

[891] Anonymous. Secure online banking solution offered. Network Security, 1997(12):6, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900878.

Anonymous:1997:PRG

[892] Anonymous. Puerto Rican Government chooses TVD. Network Security, 1997(12):6, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589790088X.

Anonymous:1997:NMS

[893] Anonymous. Network monitoring with SessionWall-3. Network Security, 1997 (12):6, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900908.

Szweda:1997:VSB

[894] Roy Szweda. VistaNet security briefing issued. Network Security, 1997 (12):6, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900891.

Anonymous:1997:RNA

[895] Anonymous. Remote network access via local phone call. Network Security, 1997(12):7, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589790091X.

Anonymous:1997:ASG

[896] Anonymous. ActiveX scanned at the gateway. Network Security, 1997 (12):7, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897900921.

Anonymous:1997:ESW

[897] Anonymous. Enterprise security for Web-based applications. Network Security, 1997(12):7–8, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897900933.

Gengler:1997:EBF

[898] Barbara Gengler. EDI-based fraud prevention for cellular phones. Network Security, 1997(12):8, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485897885479.

Gengler:1997:AUU

[899] Barbara Gengler. Axent upgrades Unix Privilege Manager. Network Security, 1997(12):8–9, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897885480.

Gengler:1997:BRE

[900] Barbara Gengler. Bovine RC5 Effort. Network Security, 1997(12): 9–10, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897885492.

Wehde:1997:WAS

[901] Ed Wehde. War against spam. Network Security, 1997(12):10, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485897885509.

Cohen:1997:MNSl

[902] Fred Cohen. Managing network security — Part 14: 50 ways to defeat your intrusion detection system. Network Security, 1997(12): 11–14, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897885510.

Hopwood:1997:CBJ

[903] David Hopwood. A comparison between Java and ActiveX security. Network Security, 1997(12): 15–20, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897885522.

Anonymous:1998:EB

[904] Anonymous. Editorial board. Network Security, 1998(1):1, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898901519.

Anonymous:1998:SDS

[905] Anonymous. The ‘smurf’ denial-of-service attack. Network Security, 1998(1):2, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901520.

Anonymous:1998:YTH

[906] Anonymous. Yahoo! threatened by hackers. Network Security, 1998 (1):2–3, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901532.

Anonymous:1998:FBP

[907] Anonymous. The ftp bounce problem. Network Security, 1998 (1):3, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901544.

Anonymous:1998:SBR

[908] Anonymous. statd buffer over-run vulnerability. Network Security, 1998 (1):3–4, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901556.

Anonymous:1998:NEA

[909] Anonymous. No export approvals required for key recovery toolkit. Network Security, 1998(1):4, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898901568.

Anonymous:1998:APS

[910] Anonymous. Action plan for safe Internet use. Network Security, 1998(1):4, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589890157X.

Anonymous:1998:SEF

[911] Anonymous. Safety of encrypted files questioned. Network Security, 1998(1):4–5, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901581.

Anonymous:1998:FOA

[912] Anonymous. Firewall offers advanced filtering capabilities. Network Security, 1998(1):5, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901593.

Anonymous:1998:SSR

[913] Anonymous. Speedy, secure remote access offering. Network Security, 1998(1):5–6, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589890160X.

Anonymous:1998:SOP

[914] Anonymous. Suite offers protection against viruses, hostile applets and hackers. Network Security, 1998(1):6, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901611.

Anonymous:1998:NGV

[915] Anonymous. New Gauntlet has VPN capabilities. Network Security, 1998(1):6, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901623.

Anonymous:1998:CCF

[916] Anonymous. Credit card fraud on Internet explodes. Network Security, 1998(1):6–7, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898901635.

Madsen:1998:CPU

[917] Wayne Madsen. Cryptography protected under US constitution. Network Security, 1998(1):7–8, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485898901647.

Anonymous:1998:FII

[918] Anonymous. The future of Internet/ intranet security. Network Security, 1998(1):8–9, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898901659.

Schultz:1998:STP

[919] E. Eugene Schultz. Securing thirdparty connections. Network Security, 1998(1):10–13, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800875923.

Hancock:1998:AID

[920] Bill Hancock. Automated intrusion detection systems and network security. Network Security, 1998 (1):14–15, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800875935.

Cohen:1998:MNSa

[921] Fred Cohen. Managing network security — Part 15: Y2K alternative solutions. Network Security, 1998 (1):16–18, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800875947.

Penman:1998:TC

[922] Ian Penman. The Tasini case. Network Security, 1998(1):19, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898901660.

Anonymous:1998:Ea

[923] Anonymous. Events. Network Security, 1998(1):20, January 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901672.

Anonymous:1998:VCD

[924] Anonymous. Vulnerabilities in common desktop environment. Network Security, 1998(2):2, February 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898900794.

Anonymous:1998:VSA

[925] Anonymous. Vulnerability in ssh-agent. Network Security, 1998(2):2, February 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898900800.

Anonymous:1998:MVU

[926] Anonymous. Macro viruses updated. Network Security, 1998 (2):2–3, February 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900812.

Anonymous:1998:SCT

[927] Anonymous. SET compliance tested. Network Security, 1998(2):3, February 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898900824.

Anonymous:1998:CMG

[928] Anonymous. Companies merge to give UK network security. Network Security, 1998(2):3, February 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900848.

Anonymous:1998:KRH

[929] Anonymous. Key recovery heavily criticized by expert. Network Security, 1998(2):3, February 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900836.

Anonymous:1998:CNJ

[930] Anonymous. C2Net and nCipher join forces for security’s sake. Network Security, 1998(2):4, February 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348589890085X.

Anonymous:1998:CASa

[931] Anonymous. Cryptographic accelerators support key management. Network Security, 1998(2):4, February 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898900861.

Anonymous:1998:KUM

[932] Anonymous. Kill unwanted e-mail messages. Network Security, 1998 (2):4, February 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900873.

Anonymous:1998:PAA

[933] Anonymous. Preemptive action against network attacks. Network Security, 1998(2):4–5, February 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900885.

Gengler:1998:NDT

[934] Barbara Gengler. Netect’s detection tool. Network Security, 1998 (2):5, February 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897831315.

Wehde:1998:RC

[935] Ed Wehde. RSA challenge. Network Security, 1998(2):5–6, February 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485897831327.

Wehde:1998:HCS

[936] Ed Wehde. HP, Cisco secure. Network Security, 1998(2):6, February 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485897831339.

Cohen:1998:MNSb

[937] Fred Cohen. Managing network security — the management of fear. Network Security, 1998(2): 7–9, February 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897831340.

Wright:1998:LPK

[938] Marie A. Wright. A look at public key certificates. Network Security, 1998 (2):10–13, February 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897831352.

Wood:1998:ECI

[939] Charles Cresson Wood. Essential controls for Internet electronic commerce. Network Security, 1998(2): 13–18, February 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485897831364.

Miller:1998:SSW

[940] Nigel Miller and Claire Frater. Shetland squabbles or worldwide wrangles? Network Security, 1998(2): 18–19, February 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485897831376.

Anonymous:1998:Ec

[941] Anonymous. Events. Network Security, 1998(2):20, February 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900897.

Anonymous:1998:MWL

[942] Anonymous. Microsoft Windows long file names exploited. Network Security, 1998(3):2, March 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901131.

Anonymous:1998:IAT

[943] Anonymous. IBM AIX telnet denial-of-service. Network Security, 1998(3):2, March 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901143.

Anonymous:1998:VPV

[944] Anonymous. volrmmount program vulnerability. Network Security, 1998(3):3, March 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901155.

Anonymous:1998:CMS

[945] Anonymous. Confidentiality of e-mail sites in UK jeopardized. Network Security, 1998(3):3, March 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901167.

Anonymous:1998:UEC

[946] Anonymous. US exports constraints removed for TIS. Network Security, 1998(3):3–4, March 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901179.

Anonymous:1998:EWB

[947] Anonymous. Encryption for Web-based banking applications. Network Security, 1998(3):4, March 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901180.

Anonymous:1998:DPD

[948] Anonymous. Data protection and disaster recovery for enterprise computing. Network Security, 1998(3): 4, March 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485898901209.

Anonymous:1998:MWC

[949] Anonymous. E-mail and Web content security solution. Network Security, 1998(3):4, March 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901192.

Gengler:1998:MF

[950] Barbara Gengler. Microsoft fix. Network Security, 1998(3):4–5, March 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898901210.

Wehde:1998:SBB

[951] Ed Wehde. Security breakthrough from Bell Labs. Network Security, 1998(3):5–6, March 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901222.

Wehde:1998:LMF

[952] Ed Wehde. Lotus moves to fix security flaws. Network Security, 1998(3):6–7, March 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901234.

Madsen:1998:SDM

[953] Wayne Madsen. State Department misses boat on telecom snooping. Network Security, 1998(3):7, March 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898901246.

Madsen:1998:USC

[954] Wayne Madsen. Uncle Sam’s crypto road show. Network Security, 1998 (3):8–13, March 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800875959.

Cohen:1998:MNSc

[955] Fred Cohen. Managing network security — red teaming. Network Security, 1998(3):13–15, March 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800875960.

Schultz:1998:PWN

[956] E. Eugene Schultz. Planning for Windows NT 5.0 security. Network Security, 1998(3):15–20, March 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485800875972.

Anonymous:1998:SRC

[957] Anonymous. Solaris rpc.cmsd vulnerability. Network Security, 1998(4):2, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898900915.

Anonymous:1998:TAC

[958] Anonymous. The truth about cookies. Network Security, 1998(4):2, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485898900903.

Anonymous:1998:ACD

[959] Anonymous. Attackers cause denialof-service. Network Security, 1998 (4):2–3, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898900927.

Anonymous:1998:SYD

[960] Anonymous. Shore up your defences, warns Microsoft. Network Security, 1998(4):3, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900939.

Anonymous:1998:SFV

[961] Anonymous. Sun finds vacation vulnerability. Network Security, 1998(4):3, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900940.

Anonymous:1998:FMV

[962] Anonymous. FreeBSD mmap vulnerability. Network Security, 1998(4): 3, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485898900952.

Anonymous:1998:VFD

[963] Anonymous. Vulnerability in FreeBSD derived stacks. Network Security, 1998(4):3–4, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900964.

Anonymous:1998:BRA

[964] Anonymous. Buffer over-runs allow arbitrary commands. Network Security, 1998(4):4, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900976.

Anonymous:1998:ZAA

[965] Anonymous. Zergo announces acquisition of PKI company. Network Security, 1998(4):4, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900988.

Anonymous:1998:IMF

[966] Anonymous. IT managers fail to receive support for security policies. Network Security, 1998(4):5, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348589890099X.

Anonymous:1998:CDR

[967] Anonymous. The cryptographic debate rages on. Network Security, 1998(4):5, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901003.

Anonymous:1998:ABF

[968] Anonymous. Alliance brings full strength encryption to Europe. Network Security, 1998(4):5, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898901015.

Anonymous:1998:DRA

[969] Anonymous. Data recovery and antivirus companies unite. Network Security, 1998(4):5–6, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901027.

Anonymous:1998:FCA

[970] Anonymous. Fred Cohen & Associates introduces new NetScan service. Network Security, 1998(4):6, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898901039.

Anonymous:1998:FSS

[971] Anonymous. File security system for business networks. Network Security, 1998(4):6, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901040.

Anonymous:1998:CSI

[972] Anonymous. E-commerce security increased by encryption suite. Network Security, 1998(4):6–7, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898901052.

Anonymous:1998:SIIa

[973] Anonymous. Solving the insecurities of Internet commerce. Network Security, 1998(4):7, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901064.

Anonymous:1998:ADB

[974] Anonymous. Automatic detection for breaches. Network Security, 1998(4):7, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898901076.

Anonymous:1998:HOM

[975] Anonymous. Hackers own methods turned against them. Network Security, 1998(4):7, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898901088.

Anonymous:1998:VSC

[976] Anonymous. Virus security combined with network security. Network Security, 1998(4):8, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348589890109X.

Anonymous:1998:TCS

[977] Anonymous. Threat to e-commerce still significant. Network Security, 1998(4):8–9, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898901106.

Gengler:1998:SS

[978] Barbara Gengler. Shredder software. Network Security, 1998(4):9, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485898901118.

Madsen:1998:EFB

[979] Wayne Madsen. European firestorm over Brussel’s report. Network Security, 1998(4):10–11, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348589890112X.

Cohen:1998:MNSd

[980] Fred Cohen. Managing network security — the unpredictability of defence. Network Security, 1998 (4):12–14, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800875984.

Osen:1998:OPI

[981] Janet Osen. Online privacy issues: The other Tim McVeigh. Network Security, 1998(4):15–18, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800875996.

Miller:1998:EN

[982] Nigel Miller. Employees on the net. Network Security, 1998(4):18–20, April 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580087600X.

Anonymous:1998:MVB

[983] Anonymous. Multiple vulnerabilities in BIND. Network Security, 1998(5):2, May 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901684.

Anonymous:1998:OGR

[984] Anonymous. Open Group reveals vulnerabilities. Network Security, 1998(5):2–3, May 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901696.

Anonymous:1998:CPSa

[985] Anonymous. Cell phone system weakened to allow surveillance. Network Security, 1998(5):3–4, May 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898901702.

Anonymous:1998:CSCa

[986] Anonymous. Call for safe e-commerce exceeding predictions. Network Security, 1998(5):4, May 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901714.

Anonymous:1998:NCI

[987] Anonymous. Name change for Integralis. Network Security, 1998(5): 4, May 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485898901726.

Anonymous:1998:PSA

[988] Anonymous. Partnership to secure against cybercrime. Network Security, 1998(5):4, May 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901738.

Anonymous:1998:SAP

[989] Anonymous. Strong authentication protects Windows NT. Network Security, 1998(5):4–5, May 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589890174X.

Anonymous:1998:CSCb

[990] Anonymous. Cryptographic solution for e-commerce security. Network Security, 1998(5):5, May 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901751.

Gengler:1998:OBF

[991] Barbara Gengler. Outsiders break into Fortune 1000 companies. Network Security, 1998(5):5–6, May 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898901763.

Zizzo:1998:ARS

[992] Thomas Zizzo. Ascend router security concerns. Network Security, 1998(5):6, May 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901775.

Madsen:1998:KEF

[993] Wayne Madsen. Key escrow flawed. Network Security, 1998(5):6–7, May 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485898901787.

Madsen:1998:EDR

[994] Wayne Madsen. Encryption debate rages again. Network Security, 1998(5):8–9, May 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898801190.

Oosthuizen:1998:SIR

[995] Gilliaume Oosthuizen. Security issues related to e-commerce. Network Security, 1998(5):10–11, May 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485898801207.

Cohen:1998:MNSe

[996] Fred Cohen. Managing network security: Risk staging. Network Security, 1998(5):12–15, May 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898801219.

Forte:1998:JRS

[997] Dario Forte. Is Java really secure? Network Security, 1998 (5):15–19, May 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898801220.

Anonymous:1998:Ed

[998] Anonymous. Events. Network Security, 1998(5):20, May 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901799.

Anonymous:1998:SIV

[999] Anonymous. SGI, IRIX vulnerabilities. Network Security, 1998 (6):2–3, June 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900009.

Anonymous:1998:PS

[1000] Anonymous. 3Com passwords stolen.

Network Security, 1998(6):3, June 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898900010.

Anonymous:1998:VFA

[1001] Anonymous. Vulnerability found in Advanced File System Utility. Network Security, 1998(6):3, June 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898900022.

Anonymous:1998:MWS

[1002] Anonymous. More Web sites vulnerable to hackers. Network Security, 1998(6):3–4, June 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900034.

Anonymous:1998:SIIb

[1003] Anonymous. Smartcard integration to intranet security. Network Security, 1998(6):4, June 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898900046.

Anonymous:1998:DAS

[1004] Anonymous. Deploy applications securely over the Internet. Network Security, 1998(6):4–5, June 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898900058.

Anonymous:1998:DWI

[1005] Anonymous. Discovering weaknesses in IP networks. Network Security, 1998(6):5, June 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348589890006X.

Madsen:1998:FFE

[1006] Wayne Madsen. Family feud over encryption policy. Network Security, 1998(6):5–6, June 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898900071.

Gengler:1998:BES

[1007] Barbara Gengler. Blizzard Entertainment sued. Network Security, 1998(6):6–7, June 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900083.

Gengler:1998:NPF

[1008] Barbara Gengler. NetScreen plays in firewall arena. Network Security, 1998(6):7, June 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900095.

Hancock:1998:ISE

[1009] Bill Hancock. IPV6 security enhancements still not everything you need. Network Security, 1998 (6):8–10, June 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589880007X.

Lau:1998:MFSa

[1010] Oliver Lau. Meta-firewall: a sixth generation firewall — Part 1. Network Security, 1998(6):10–16, June 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898800081.

Cohen:1998:MNSf

[1011] Fred Cohen. Managing network security: Technical protection for the joint venture. Network Security, 1998(6):16–20, June 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898800093.

Anonymous:1998:MIV

[1012] Anonymous. More IRIX vulnerabilities. Network Security, 1998 (7):2, 3, July 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898900320.

Anonymous:1998:BON

[1013] Anonymous. Buffer overflow in NIS+. Network Security, 1998(7):3, July 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485898900332.

Anonymous:1998:PER

[1014] Anonymous. Parsing error reduces key length. Network Security, 1998(7):4, July 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898900344.

Anonymous:1998:FOL

[1015] Anonymous. Financial organizations leave Internet security lax. Network Security, 1998(7):4, July 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898900356.

Anonymous:1998:PMB

[1016] Anonymous. Partnership for migrating business securely on Internet. Network Security, 1998(7):4–5, July 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485898900368.

Anonymous:1998:CFP

[1017] Anonymous. Companies focus on PKI business. Network Security, 1998(7):5, July 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589890037X.

Anonymous:1998:ASW

[1018] Anonymous. Added security for Webbased Java applications. Network Security, 1998(7):5, July 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900393.

Anonymous:1998:SCSa

[1019] Anonymous. Shark’s cage security for NT. Network Security, 1998(7):5, July 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900381.

Anonymous:1998:CPP

[1020] Anonymous. Crypto pack protects electronic information. Network Security, 1998(7):6, July 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589890040X.

Madsen:1998:ELP

[1021] Wayne Madsen. Encryption legislation and policy. Network Security, 1998(7):6–7, July 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898800056.

Madsen:1998:NCK

[1022] Wayne Madsen. NAFTA has crypto key recovery agenda. Network Security, 1998(7):7–8, July 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900411.

Cohen:1998:MNSg

[1023] Fred Cohen. Managing network security: How does a typical IT audit work? Network Security, 1998(7):8–11, July 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898800068.

Rawlings:1998:ECIa

[1024] Jane Rawlings. Electronic commerce on the Internet — Part 1. Network Security, 1998(7):11–14, July 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348589880010X.

Lau:1998:MFSb

[1025] Oliver Lau. Meta-firewall: a sixth generation firewall — Part 2. Network Security, 1998(7):14–19, July 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898800111.

Anonymous:1998:Ee

[1026] Anonymous. Events. Network Security, 1998(7):20, July 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900423.

Anonymous:1998:HUF

[1027] Anonymous. HP-UX ftp problems. Network Security, 1998 (8):2, August 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900228.

Anonymous:1998:OVL

[1028] Anonymous. OpenVMS V7.1 LOGI-NOUT security vulnerability. Network Security, 1998(8):2, August 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348589890023X.

Anonymous:1998:CBT

[1029] Anonymous. Cracking of 56-bit DES not a threat. Network Security, 1998(8):2–3, August 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900241.

Anonymous:1998:SIS

[1030] Anonymous. Security integration to simplify user access. Network Security, 1998(8):3, August 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900253.

Anonymous:1998:CSP

[1031] Anonymous. Certified SET payment system integrated into IN-TERSHOP. Network Security, 1998 (8):3, August 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898900265.

Anonymous:1998:ESI

[1032] Anonymous. Enterprise security and intrusion detection come together. Network Security, 1998 (8):3–4, August 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898900277.

Anonymous:1998:PSF

[1033] Anonymous. Peapod and SeaChange form new company. Network Security, 1998(8):4, August 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898900289.

Anonymous:1998:WSB

[1034] Anonymous. Web security for browsers. Network Security, 1998 (8):4, August 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898900290.

Anonymous:1998:SPS

[1035] Anonymous. Smartcard provides strong private key functions. Network Security, 1998(8):5, August 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898900307.

Madsen:1998:CPH

[1036] Wayne Madsen. Crypto politics heating up. Network Security, 1998 (8):5–6, August 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898800706.

Wehde:1998:MBE

[1037] Ed Wehde. Moves to break encryption deadlock. Network Security, 1998(8):6–7, August 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898800718.

Gengler:1998:CFA

[1038] Barbara Gengler. CIA fears America is vulnerable. Network Security, 1998(8):7, August 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589880072X.

Rees:1998:CCA

[1039] Frank Rees. Computer crime and airforce information systems. Network Security, 1998(8):7–9, August 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898800731.

Cohen:1998:MNSh

[1040] Fred Cohen. Managing network security: The seedy side of security. Network Security, 1998 (8):9–13, August 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898800743.

Hancock:1998:UNH

[1041] Bill Hancock. Using network hacking techniques to disrupt military operational networks. Network Security, 1998(8):13–17, August 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898800755.

Rawlings:1998:ECIb

[1042] Jane Rawlings. Electronic commerce on the Internet — Part 2. Network Security, 1998(8): 17–19, August 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898800767.

Anonymous:1998:Ef

[1043] Anonymous. Events. Network Security, 1998(8):20, August 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898900319.

Anonymous:1998:SVM

[1044] Anonymous. Security vulnerabilities in Microsoft PPTP. Network Security, 1998(9):2, September 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901258.

Anonymous:1998:WEJ

[1045] Anonymous. ‘Window external’ jScript vulnerability in Internet Explorer 4. Network Security, 1998 (9):2–3, September 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589890126X.

Anonymous:1998:CIR

[1046] Anonymous. Cisco IOS remote router crash. Network Security, 1998 (9):3, September 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901271.

Anonymous:1998:SCSb

[1047] Anonymous. Support for child-safe Web sites. Network Security, 1998 (9):3, September 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901283.

Anonymous:1998:UAC

[1048] Anonymous. Unix access control software. Network Security, 1998 (9):3–4, September 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901295.

Anonymous:1998:CASb

[1049] Anonymous. Cryptographic accelerator for speedy digital signatures. Network Security, 1998 (9):4, September 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898901301.

Anonymous:1998:SCR

[1050] Anonymous. Secure card reader for Internet security, e-commerce. Network Security, 1998(9):4–5, September 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485898901313.

Gengler:1998:INC

[1051] Barbara Gengler. IBM’s new cryptosystem. Network Security, 1998 (9):5, September 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898800123.

Gengler:1998:ESR

[1052] Barbara Gengler. Encryption standard replaced. Network Security, 1998 (9):5–6, September 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898800135.

Cohen:1998:MNSi

[1053] Fred Cohen. Managing network security: What should I report to whom? Network Security, 1998 (9):6–12, September 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898800147.

Forte:1998:GSN

[1054] Dario Forte. Guaranteeing the safety of a network beyond the firewall. Network Security, 1998(9): 12–16, September 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898800159.

Schultz:1998:PMN

[1055] E. Eugene Schultz and Philip C. Cox.

Putting the Mime name vulnerability in outlook and messenger in perspective. Network Security, 1998(9): 16–19, September 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898800160.

Anonymous:1998:Eg

[1056] Anonymous. Events. Network Security, 1998(9):19–20, September 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898901325.

Anonymous:1998:VF

[1057] Anonymous. Vulnerabilities with FTP. Network Security, 1998 (10):2, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901805.

Anonymous:1998:UFR

[1058] Anonymous. US finally relaxes encryption policies. Network Security, 1998(10):2–3, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898901817.

Anonymous:1998:HSP

[1059] Anonymous. Hit squad put on trail of hackers. Network Security, 1998(10):3, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898901829.

Anonymous:1998:VEE

[1060] Anonymous. VPN encouraged in Europe. Network Security, 1998 (10):3, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898901830.

Anonymous:1998:CAP

[1061] Anonymous. Company acquisition to provide network protection. Network Security, 1998 (10):3–4, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898901842.

Anonymous:1998:COS

[1062] Anonymous. Companies offer secure file delivery. Network Security, 1998(10):4, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901854.

Anonymous:1998:DSR

[1063] Anonymous. Dell to sell Raptor firewalls. Network Security, 1998 (10):4, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901866.

Anonymous:1998:MMF

[1064] Anonymous. Maximize multiple firewalls. Network Security, 1998 (10):4–5, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901878.

Anonymous:1998:TMS

[1065] Anonymous. Threat management solution offered. Network Security, 1998(10):5, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589890188X.

Anonymous:1998:WRC

[1066] Anonymous. Weaknesses in RSA combated. Network Security, 1998 (10):5–6, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901891.

Anonymous:1998:NFN

[1067] Anonymous. Native firewall for NT.

Network Security, 1998(10):6, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898901908.

Anonymous:1998:JBM

[1068] Anonymous. Java-based management solution for global network security. Network Security, 1998 (10):6, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589890191X.

Anonymous:1998:ESC

[1069] Anonymous. Entegrity secures cash flows. Network Security, 1998 (10):6–7, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901921.

Gengler:1998:VE

[1070] Barbara Gengler. The Vulnerability Engine. Network Security, 1998 (10):7, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901933.

Gengler:1998:NIS

[1071] Barbara Gengler. New Internet security standard. Network Security, 1998(10):7–8, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901945.

Meears-White:1998:NI

[1072] Tristan Meears-White. NetRegulation.inconfusion@uk.

Network Security, 1998(10):8–9, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485800876011.

Cohen:1998:MNSj

[1073] Fred Cohen. Managing network security: Time-based security. Network Security, 1998(10):10–14, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800876023.

Wright:1998:ECC

[1074] Marie A. Wright. The elliptic curve cryptosystem: a synopsis. Network Security, 1998(10): 14–17, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800876035.

Hancock:1998:SIN

[1075] Bill Hancock. Skills inventory for network management. Network Security, 1998(10):17–19, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800876047.

Anonymous:1998:Eh

[1076] Anonymous. Events. Network Security, 1998(10):20, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901957.

Anonymous:1998:SLH

[1077] Anonymous. Simple lessons are the hardest to learn. Network Security, 1998(11):2, November 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900435.

Szweda:1998:CAF

[1078] Roy Szweda. China arrests first online bank robbers. Network Security, 1998(11):2, November 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900447.

Anonymous:1998:USP

[1079] Anonymous. Untrusted scripted paste in Internet Explorer 4.01. Network Security, 1998(11):2–3, November 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898900459.

Anonymous:1998:CIL

[1080] Anonymous. Cisco IOS login prompt vulnerability. Network Security, 1998 (11):3, November 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900460.

Anonymous:1998:BOVa

[1081] Anonymous. Buffer overflow vulnerability in mountd. Network Security, 1998(11):3, November 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900472.

Anonymous:1998:UGT

[1082] Anonymous. UK Government to toughen encryption regulations. Network Security, 1998(11):3–4, November 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898900484.

Anonymous:1998:QIS

[1083] Anonymous. Quantitative information security risk study launched. Network Security, 1998(11):4, November 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898900496.

Anonymous:1998:WSL

[1084] Anonymous. Web sites leave companies legally exposed. Network Security, 1998(11):4–5, November 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900502.

Anonymous:1998:CDS

[1085] Anonymous. Companies develop secure home banking solutions. Network Security, 1998(11):5, November 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898900514.

Anonymous:1998:DSI

[1086] Anonymous. Digital signatures incorporated into electronic forms. Network Security, 1998(11):5, November 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485898900526.

Anonymous:1998:PED

[1087] Anonymous. Protection for electronic documents. Network Security, 1998 (11):5–6, November 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898900538.

Anonymous:1998:LEE

[1088] Anonymous. Link encryptor with electronically loadable algorithms. Network Security, 1998(11):6, November 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348589890054X.

Anonymous:1998:NSE

[1089] Anonymous. NT security enhanced. Network Security, 1998 (11):6–7, November 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898900551.

Szweda:1998:MMS

[1090] Roy Szweda. Micro-miniature silicon lock makes computers ‘hackerproof’. Network Security, 1998 (11):7–8, November 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898900563.

Cohen:1998:MNSk

[1091] Fred Cohen. Managing network security: The real Y2K issue. Network Security, 1998(11):8–11, November 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800876059.

Hancock:1998:INS

[1092] Bill Hancock. Improving network security through port switching. Network Security, 1998(11): 12–13, November 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800876060.

Osen:1998:TSJ

[1093] Janet Osen. The thorny side of jurisdiction and the Internet. Network Security, 1998(11):13–17, November 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800876072.

Osen:1998:SBC

[1094] Janet Osen. Singing the blues: The case of the ‘little’ Blue Note. Network Security, 1998(11):18–19, November 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898900575.

Anonymous:1998:Ei

[1095] Anonymous.  Events.  Network

Security, 1998(11):19–20, November 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898900587.

Szweda:1998:USF

[1096] Roy Szweda. University switches to financial security system. Network Security, 1998(12):2, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898901337.

Anonymous:1998:BOVb

[1097] Anonymous. Buffer overflow vulnerabilities in rdist. Network Security, 1998(12):2, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901349.

Anonymous:1998:RGB

[1098] Anonymous. Reactor glitches blamed on software. Network Security, 1998 (12):2, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901350.

Anonymous:1998:MSS

[1099] Anonymous. E-mail security survey. Network Security, 1998(12):3, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485898901362.

Anonymous:1998:IHT

[1100] Anonymous. IRC hit by Trojan horse. Network Security, 1998 (12):3, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901374.

Anonymous:1998:IOP

[1101] Anonymous. Internet outages predicted at COMDEX. Network Security, 1998(12):3–4, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901386.

Anonymous:1998:CEC

[1102] Anonymous. Coming explosion in ecommerce. Network Security, 1998 (12):4, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901398.

Anonymous:1998:AAV

[1103] Anonymous. Alphabetic anti-virus guide. Network Security, 1998 (12):4, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901404.

Anonymous:1998:BIB

[1104] Anonymous. Brokat is Internet banking leader says report. Network Security, 1998(12):4, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485898901416.

Anonymous:1998:BPL

[1105] Anonymous. Baltimore partners logica for net security. Network Security, 1998(12):4–5, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898901428.

Anonymous:1998:TDP

[1106] Anonymous. Tritheim defender provides affordable security. Network Security, 1998(12):5, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348589890143X.

Anonymous:1998:VCP

[1107] Anonymous. VLSI chip provides secure e-commerce. Network Security, 1998(12):5, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898901441.

Anonymous:1998:CPSb

[1108] Anonymous. Commercial phone scanner. Network Security, 1998 (12):5, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485898901453.

Anonymous:1998:CHS

[1109] Anonymous. COMDEX highlights SecureZone. Network Security, 1998 (12):5–6, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901465.

Szweda:1998:CEM

[1110] Roy Szweda. Cyberflex enables mobile smart services via GSM. Network Security, 1998(12):6, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485898901477.

Gengler:1998:CHT

[1111] Barbara Gengler. Controversy hits Tristrata. Network Security, 1998 (12):7, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901489.

Wehde:1998:SCW

[1112] Ed Wehde. Security companies work together. Network Security, 1998 (12):7–8, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901490.

Hancock:1998:NSM

[1113] Bill Hancock. Network security:

Monitoring is the first step. Network Security, 1998(12):8–9, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800876084.

Cohen:1998:MNSl

[1114] Fred Cohen. Managing network security: Balancing risk. Network Security, 1998(12):10–14, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800876096.

Piper:1998:CSV

[1115] Fred Piper and Michael Walker. Cryptographic solutions for voice telephony and GSM. Network Security, 1998 (12):14–19, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800876102.

Anonymous:1998:Ej

[1116] Anonymous. Events. Network Security, 1998(12):20, December 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485898901507.

Anonymous:1999:EB

[1117] Anonymous. Editorial board. Network Security, 1999(1):OFC, January 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899901861.

Anonymous:1999:SIF

[1118] Anonymous. SGI IRIX fcagent daemon vulnerability. Network Security, 1999(1):2, January 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901873.

Anonymous:1999:CISa

[1119] Anonymous. Cisco IOS syslog denialof-service. Network Security, 1999 (1):2–3, January 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901885.

Anonymous:1999:HUVa

[1120] Anonymous. HP-UX vulnerabilities. Network Security, 1999 (1):3, January 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901897.

Anonymous:1999:SVS

[1121] Anonymous. Security vulnerability in sendmail. Network Security, 1999(1):3, January 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901903.

Anonymous:1999:SSD

[1122] Anonymous. Sun Solaris dtmail, passwd vulnerabilities. Network Security, 1999(1):3–4, January 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899901915.

Anonymous:1999:THV

[1123] Anonymous. Trojan horse version of TCP wrappers. Network Security, 1999(1):4, January 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901927.

Anonymous:1999:WNR

[1124] Anonymous. Windows NT remote explorer. Network Security, 1999 (1):4–5, January 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901939.

Anonymous:1999:EGC

[1125] Anonymous. Entrust gets contract to provide authenticity. Network Security, 1999(1):5, January 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901940.

Anonymous:1999:ITP

[1126] Anonymous. ICL/TimeStep partnership produces VPN. Network Security, 1999(1):5–6, January 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901952.

Anonymous:1999:INH

[1127] Anonymous. Integrated network and host-based intrusion detection solution. Network Security, 1999 (1):6, January 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901964.

Anonymous:1999:HAF

[1128] Anonymous. High availability firewall server solution. Network Security, 1999(1):6–7, January 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901976.

Anonymous:1999:IPSa

[1129] Anonymous. Intelligent probe simulates hacking and potential damage. Network Security, 1999(1):7, January 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899901988.

Gengler:1999:SGR

[1130] Barbara Gengler. Security guidelines released. Network Security, 1999(1):7–8, January 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589990199X.

Hancock:1999:ANR

[1131] Bill Hancock. Attacking network routers. Network Security, 1999 (1):8–9, January 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899800014.

Tantum:1999:LRIa

[1132] Mark Tantum and Touche Ross. Legal responsibility of Internet service providers: Part 1. Network Security, 1999(1):10–15, January 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899800026.

Cohen:1999:MNSa

[1133] Fred Cohen. Managing network security: Anatomy of a successful sophisticated attack. Network Security, 1999(1):16–19, January 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899800038.

Anonymous:1999:Ea

[1134] Anonymous. Events. Network Security, 1999(1):19–20, January 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899902003.

Anonymous:1999:WFM

[1135] Anonymous. W97M.Footprint macro virus discovered. Network Security, 1999(2):2, February 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899900387.

Anonymous:1999:FPP

[1136] Anonymous. FTP PASV ‘pizza thief’ exploit. Network Security, 1999 (2):2–3, February 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899900399.

Anonymous:1999:NVC

[1137] Anonymous. Network video conferencing needs security. Network Security, 1999(2):3, February 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900405.

Anonymous:1999:BFA

[1138] Anonymous. Budgets fail to address security needs. Network Security, 1999(2):3, February 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900417.

Anonymous:1999:SRA

[1139] Anonymous. Security research alliance to promote network security. Network Security, 1999(2): 3–4, February 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900429.

Anonymous:1999:PDP

[1140] Anonymous. Partnership to distribute PKI solutions. Network Security, 1999(2):4, February 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900430.

Anonymous:1999:TPS

[1141] Anonymous. Turnkey PI solution for VPN-1. Network Security, 1999 (2):4, February 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900442.

Anonymous:1999:TCF

[1142] Anonymous. Tool centralizes firewall configuration. Network Security, 1999(2):5, February 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899900454.

Anonymous:1999:FIR

[1143] Anonymous. Fraud on the Internet reduced. Network Security, 1999(2):5, February 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899900466.

Gengler:1999:NAG

[1144] Barbara Gengler. Network Associates gets contract. Network Security, 1999 (2):5–6, February 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899900478.

Wehde:1999:ASG

[1145] Ed Wehde. ATM security gets better. Network Security, 1999 (2):6, February 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348589990048X.

Hancock:1999:ARN

[1146] Bill Hancock. Assessing and reducing network risk. Network Security, 1999(2):7–8, February 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800239.

Hancock:1999:NSU

[1147] Bill Hancock. Network security: The unsolved mystery. Network Security, 1999(2):9–11, February 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800240.

Cohen:1999:MNSb

[1148] Fred Cohen. Managing network security: Returning fire. Network Security, 1999(2):11–15, February 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800800252.

Tantum:1999:LRIb

[1149] Mark Tantum and Touche Ross. Legal responsibility of Internet service providers: Part 2. Network Security, 1999(2):15–19, February 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800800264.

Anonymous:1999:Ec

[1150] Anonymous. Events. Network Security, 1999(2):19–20, February 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899900491.

Anonymous:1999:BOF

[1151] Anonymous. Buffer overflows in FTP servers. Network Security, 1999(3):2, March 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900119.

Anonymous:1999:SSV

[1152] Anonymous. Sun Solaris vulnerabilities. Network Security, 1999(3):2, March 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900120.

Anonymous:1999:MBV

[1153] Anonymous. Microsoft BackOffice vulnerability. Network Security, 1999(3):2–3, March 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900132.

Anonymous:1999:DLS

[1154] Anonymous. Debian Linux ‘Super’ package buffer overflow. Network Security, 1999(3):3, March 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900144.

Anonymous:1999:DUV

[1155] Anonymous. Digital Unix vulnerabilities. Network Security, 1999 (3):3–4, March 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900156.

Anonymous:1999:VSU

[1156] Anonymous. Virus sends users’ details to virus exchange site. Network Security, 1999(3):4, March 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900168.

Anonymous:1999:ECP

[1157] Anonymous. EU called to petition against unwanted e-mail. Network Security, 1999(3):4–5, March 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348589990017X.

Anonymous:1999:UGH

[1158] Anonymous. UK to get a head start in the digital marketplace. Network Security, 1999(3):5, March 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900181.

Anonymous:1999:PSM

[1159] Anonymous. PKI solution for the Microsoft commerce platform. Network Security, 1999(3):5–6, March 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899900193.

Anonymous:1999:SST

[1160] Anonymous. Security simplified through integration of adaptive security. Network Security, 1999(3): 6, March 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S135348589990020X.

Anonymous:1999:BPE

[1161] Anonymous. Baltimore to PKI enable Adobe Acrobat. Network Security, 1999(3):6, March 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899900211.

Anonymous:1999:PCE

[1162] Anonymous. Product checks and enforces security best practices. Network Security, 1999(3):7, March 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485899900223.

Gengler:1999:TPT

[1163] Barbara Gengler. Total privacy technology. Network Security, 1999(3): 7, March 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1353485899900235.

Gengler:1999:FMS

[1164] Barbara Gengler. Free e-mail services or disservices? Network Security, 1999(3):8, March 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899900247.

Gengler:1999:CSC

[1165] Barbara Gengler. E-commerce start-up has cash in its wallet. Network Security, 1999(3):8–9, March 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900259.

Madsen:1999:NCO

[1166] Wayne Madsen. NSA continues to oppose crypto export control relief. Network Security, 1999 (3):9–10, March 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900260.

Hancock:1999:NSM

[1167] Bill Hancock. Network security:

Monitoring is the first step. Network Security, 1999(3):10–11, March 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800800215.

Cohen:1999:MNSc

[1168] Fred Cohen. Managing network security: The millisecond fantasy. Network Security, 1999(3):12–19, March 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800800227.

Anonymous:1999:Ed

[1169] Anonymous. Events. Network Security, 1999(3):20, March 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900272.

Anonymous:1999:MSW

[1170] Anonymous. Melissa spreading widely in USA. Network Security, 1999(4):2, April 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899901794.

Hancock:1999:MAC

[1171] Bill Hancock. Melissa’s author caught. Network Security, 1999 (4):2–3, April 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899901800.

Anonymous:1999:HUVb

[1172] Anonymous. HP-UX vulnerabilities. Network Security, 1999 (4):3–4, April 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899901812.

Gengler:1999:ELM

[1173] Barbara Gengler. Encryption laws may slacken. Network Security, 1999(4):4, April 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899901824.

Gengler:1999:MQN

[1174] Barbara Gengler. Microsoft questions Novell’s NDS security. Network Security, 1999(4):4–5, April 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901836.

Madsen:1999:NRE

[1175] Wayne Madsen. NSA recommended for expanded Infosec and Infowar responsibilities. Network Security, 1999(4):5–6, April 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901848.

Cohen:1999:MNSd

[1176] Fred Cohen. Managing network security: Simulating network security. Network Security, 1999 (4):6–13, April 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800094.

Miller:1999:MAC

[1177] Nigel Miller. E-mail abuse and corporate policies. Network Security, 1999(4):13–17, April 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800100.

Schultz:1999:BRF

[1178] E. Eugene Schultz. Book review: Fighting computer crime: Author: Donn Parker, ISBN 0-471-16378-3, Publisher: John Wiley and Sons, 1998. Network Security, 1999 (4):17–19, April 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800800112.

Anonymous:1999:Ee

[1179] Anonymous. Events. Network Security, 1999(4):19–20, April 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348589990185X.

Anonymous:1999:BOS

[1180] Anonymous. Buffer overflow in SGI X server’s font path. Network Security, 1999(5):2, May 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899902799.

Anonymous:1999:VCTa

[1181] Anonymous. Vulnerabilities of Cisco 7xx TCP and HTTP. Network Security, 1999(5):2–3, May 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899902805.

Anonymous:1999:IFL

[1182] Anonymous. Implementation flaw in the Linux TCP/IP stack. Network Security, 1999(5):3–4, May 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485899902817.

Anonymous:1999:EPR

[1183] Anonymous. Entrust PKI receives validation to new global standard. Network Security, 1999(5):4, May 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899902829.

Anonymous:1999:PSI

[1184] Anonymous. PKI solution integrated with Web security software. Network Security, 1999(5):4–5, May 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899902830.

Anonymous:1999:UCB

[1185] Anonymous. UK companies benefit from VPNs. Network Security, 1999(5):5, May 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899902842.

Anonymous:1999:EPFa

[1186] Anonymous. Encryptors provide frame relay security. Network Security, 1999(5):5, May 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899902854.

Madsen:1999:BBG

[1187] Wayne Madsen. Big brother goes global. Network Security, 1999(5): 6, May 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485899902866.

Madsen:1999:MBF

[1188] Wayne Madsen. McCain’s bill further ‘Balkanizes’ US crypto export policy. Network Security, 1999(5):6–7, May 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899902878.

Cohen:1999:MNSe

[1189] Fred Cohen. Managing network security: Watching the world. Network Security, 1999(5):8–13, May 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485800800197.

Miller:1999:YMS

[1190] Nigel Miller. Y2K mediation success stories. Network Security, 1999(5):13, May 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348589990288X.

Miller:1999:PCC

[1191] Nigel Miller. Pirates and casual copiers: how to protect the content of your Web site. Network Security, 1999(5):13–14, May 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899902891.

Miller:1999:MDD

[1192] Nigel Miller. $107 million damages for disseminating unlawful information on a Web site. Network Security, 1999(5):14–15, May 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899902908.

Miller:1999:UCB

[1193] Nigel Miller. UK e-commerce bill — summary. Network Security, 1999(5):15–16, May 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589990291X.

Miller:1999:LIL

[1194] Nigel Miller. Liability for infringements on linked sites? Network Security, 1999(5):16–17, May 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899902921.

Miller:1999:TDP

[1195] Nigel Miller. The telecommunications (data protection and privacy) (direct marketing) regulations 1998. Network Security, 1999(5):17, May 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899902933.

Hancock:1999:ECI

[1196] Bill Hancock. Export of cryptographic information from the USA: A brief look at the problems. Network Security, 1999(5):17–19, May 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800800203.

Anonymous:1999:Ef

[1197] Anonymous. Events. Network Security, 1999(5):19–20, May 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899902945.

Anonymous:1999:HUS

[1198] Anonymous. HP-UX sendmail — denial of service failures. Network Security, 1999(6):2, June 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899900508.

Anonymous:1999:CISb

[1199] Anonymous. Cisco IOS software vulnerability. Network Security, 1999(6):2–3, June 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348589990051X.

Anonymous:1999:TIE

[1200] Anonymous. Telecommunications infrastructure to expand in Latin America. Network Security, 1999(6):3, June 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485899900521.

Anonymous:1999:PDS

[1201] Anonymous. Partnership delivers server-based anti-virus solution. Network Security, 1999 (6):3–4, June 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900533.

Anonymous:1999:IIV

[1202] Anonymous. IPSec/IKE VPN and X.509 CA device secures networks. Network Security, 1999(6):4, June 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899900545.

Anonymous:1999:CSS

[1203] Anonymous. Compact solution to simplified firewall deployment. Network Security, 1999(6):4–5, June 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899900557.

Anonymous:1999:LCV

[1204] Anonymous. Low-cost VPN solution for the remote office. Network Security, 1999(6):5, June 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900569.

Gengler:1999:NCW

[1205] Barbara Gengler. Networking companies won’t make kit Y2K compliant. Network Security, 1999 (6):5–6, June 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900570.

Gengler:1999:NCG

[1206] Barbara Gengler. Now cryptography gets the ‘open source’ treatment. Network Security, 1999(6):6, June 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485899900582.

Gengler:1999:CPO

[1207] Barbara Gengler. Check point outlines VPN strategy. Network Security, 1999(6):6–7, June 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899900594.

Gengler:1999:FAS

[1208] Barbara Gengler. Federal agency sites still lack privacy notices. Network Security, 1999(6):7, June 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899900600.

Cohen:1999:MNSf

[1209] Fred Cohen. Managing network security: The limits of awareness. Network Security, 1999 (6):8–10, June 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580080015X.

Forte:1999:FAE

[1210] Dario Forte. The future of the Advanced Encryption Standard. Network Security, 1999(6):10–13, June 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800800161.

Arden:1999:YWY

[1211] Michelle J. Arden. “Are you who you say you are?” the financial market is finding out with PKI. Network Security, 1999 (6):13–16, June 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800173.

David:1999:VAPa

[1212] Jon David. Vulnerabilities assessment — Part 1. Vulnerability basics. Network Security, 1999 (6):16–18, June 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800185.

Anonymous:1999:Eg

[1213] Anonymous. Events. Network Security, 1999(6):19–20, June 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900612.

Anonymous:1999:VCTb

[1214] Anonymous. Vulnerability in Compaq’s Tru64/DIGITAL UNIX software. Network Security, 1999(7):2, July 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899900004.

Anonymous:1999:RSA

[1215] Anonymous. rpc.statd and automountd

vulnerabilities allow intruder access. Network Security, 1999(7):2, July 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899900016.

Anonymous:1999:TSI

[1216] Anonymous. TROJ XPLOREZIP — stripping it from your system. Network Security, 1999(7):2–3, July 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899900028.

Anonymous:1999:EUD

[1217] Anonymous. EU-US data privacy dispute to continue. Network Security, 1999(7):3–4, July 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589990003X.

Anonymous:1999:OWD

[1218] Anonymous. Online Web domain registration protects against cybersquatting. Network Security, 1999(7):4, July 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900041.

Anonymous:1999:CTS

[1219] Anonymous. Companies team up to secure e-commerce in Hong Kong. Network Security, 1999(7):4, July 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899900053.

Anonymous:1999:EPFb

[1220] Anonymous. Entrust provide free PKI and multi-platform toolkits for download. Network Security, 1999(7):4–5, July 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900065.

Gengler:1999:IFC

[1221] Barbara Gengler. Internet Fraud Council. Network Security, 1999(7):5, July 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899900077.

Gengler:1999:T

[1222] Barbara Gengler. Twinkle. Network Security, 1999(7):5–6, July 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899900089.

Wehde:1999:MSE

[1223] Ed Wehde. Military strength encryption for consumers. Network Security, 1999(7):6–7, July 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900090.

Cohen:1999:MNSg

[1224] Fred Cohen. Managing network security: Attack and defence strategies. Network Security, 1999 (7):7–11, July 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800124.

Webb:1999:ANPa

[1225] Alison Webb. Auditing NT — Part 1. Network Security, 1999 (7):11–15, July 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800136.

Forte:1999:PSA

[1226] Dario Forte. Postfix — a secure alternative to Sendmail: A new frontier in the realm of free mailers. Network Security, 1999 (7):16–18, July 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800148.

Anonymous:1999:Eh

[1227] Anonymous. Events. Network Security, 1999(7):19–20, July 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900107.

Anonymous:1999:NES

[1228] Anonymous. Netscape enterprise server fails to properly process some URL’s. Network Security, 1999(8):2, August 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900284.

Anonymous:1999:MHR

[1229] Anonymous. Malformed .HTR request vulnerability. Network Security, 1999(8):2, August 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900296.

Anonymous:1999:DSV

[1230] Anonymous. Denial-of-service vulnerabilities in Windows NT 4.0. Network Security, 1999(8):2–3, August 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899900302.

Anonymous:1999:FPW

[1231] Anonymous. Finland paves the way for a digital society. Network Security, 1999(8):3, August 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900314.

Anonymous:1999:GMS

[1232] Anonymous. Global market for security products to grow to $8 billion in 2003. Network Security, 1999(8):3–4, August 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899900326.

Anonymous:1999:CDC

[1233] Anonymous. Companies develop Chinese language CA. Network Security, 1999(8):4, August 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899900338.

Anonymous:1999:CEV

[1234] Anonymous. Cylink expands its VPN offerings with ATM encryptor. Network Security, 1999(8):4–5, August 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348589990034X.

Anonymous:1999:EDB

[1235] Anonymous. Encrypt data at 6.7 billion bits per second. Network Security, 1999(8):5, August 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899900351.

Anonymous:1999:SSF

[1236] Anonymous. Secure solution for frame relay networks. Network Security, 1999(8):5, August 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899900363.

Wehde:1999:TVC

[1237] Ed Wehde. Third voice causes problems. Network Security, 1999 (8):5–6, August 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800800070.

Wehde:1999:SRM

[1238] Ed Wehde. Sniffing out rogue modems. Network Security, 1999 (8):6–7, August 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800082.

Wehde:1999:CLJ

[1239] Ed Wehde. Crime lab just for the Internet. Network Security, 1999(8):7, August 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800331.

Cohen:1999:MNSh

[1240] Fred Cohen. Managing network security: What’s happening out there. Network Security, 1999 (8):8–11, August 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800343.

David:1999:VAPb

[1241] Jon David. Vulnerabilities assessment — Part 2. Getting in through the ‘Net’. Network Security, 1999 (8):11–14, August 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800355.

Webb:1999:ANPb

[1242] Alison Webb. Auditing NT — Part 2. Network Security, 1999 (8):14–18, August 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800367.

Madsen:1999:BOC

[1243] Wayne Madsen. Barr offers Congressional oversight amendment on ECHELON. Network Security, 1999 (8):18–19, August 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800800379.

Anonymous:1999:Ei

[1244] Anonymous. Events. Network Security, 1999(8):20, August 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899900375.

Anonymous:1999:MWE

[1245] Anonymous. Microsoft Web engine vulnerability. Network Security, 1999 (9):2, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899902490.

Anonymous:1999:WNT

[1246] Anonymous. Windows NT Terminal Servers DOS vulnerability. Network Security, 1999(9):2, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485899902507.

Anonymous:1999:IAV

[1247] Anonymous. IBM AIX vulnerability in ptrace() system call. Network Security, 1999(9):2, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899902519.

Anonymous:1999:FSC

[1248] Anonymous. Free smart card software for digital certificates. Network Security, 1999(9):2–3, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899902520.

Anonymous:1999:WDE

[1249] Anonymous. Wireless data encryption for handhelds. Network Security, 1999(9):3, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899902532.

Anonymous:1999:IPSb

[1250] Anonymous. Internet postage system launched in the United States. Network Security, 1999(9):3, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899902544.

Anonymous:1999:CAC

[1251] Anonymous. Clinton administration calls for network monitoring. Network Security, 1999(9): 3, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899902556.

Anonymous:1999:SMB

[1252] Anonymous. Six million Brazilians file taxes over the Net. Network Security, 1999(9):4, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899902568.

Anonymous:1999:AEH

[1253] Anonymous. ADSL enables hacker access. Network Security, 1999 (9):4, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589990257X.

Anonymous:1999:FYI

[1254] Anonymous. Face your image on a bar code. Network Security, 1999 (9):4–5, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899902581.

Anonymous:1999:PYN

[1255] Anonymous. Protect your networks with data-security chips. Network Security, 1999(9):5, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899902593.

Gengler:1999:ARI

[1256] Barbara Gengler. Alarm raised over intrusion detection network. Network Security, 1999(9):5–6, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348589990260X.

Gengler:1999:HPS

[1257] Barbara Gengler. Hewlett Packard security strategy. Network Security, 1999(9):6–7, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899902611.

Gengler:1999:CSS

[1258] Barbara Gengler. E-commerce security study. Network Security, 1999 (9):7, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899902623.

Cohen:1999:MNSi

[1259] Fred Cohen. Managing network security: In your face information warfare. Network Security, 1999(9): 8–10, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800276.

Madsen:1999:SBA

[1260] Wayne Madsen. SAFE Bill “ambushed”. Network Security, 1999(9): 10–11, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800288.

Madsen:1999:JDP

[1261] Wayne Madsen. Justice Department, PCCIP ex-staffers offer critical infrastructure assurance proposals. Network Security, 1999(9):

11–12, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580080029X.

Madsen:1999:SBC

[1262] Wayne Madsen. SAFE Bill clears Telecommunications Subcommittee with amendments. Network Security, 1999(9):13–14, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800800306.

Wright:1999:OP

[1263] Marie A. Wright. An overview of PKI. Network Security, 1999(9): 14–17, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800318.

Heath:1999:WSE

[1264] Guy Heath. WIPO serves eviction notice on cyber-squatters. Network Security, 1999(9):18–19, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580080032X.

Anonymous:1999:Ej

[1265] Anonymous. Events. Network Security, 1999(9):19–20, September 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899902635.

Anonymous:1999:ACS

[1266] Anonymous. ActiveX controls not ‘safe for scripting’. Network Security, 1999(10):2, October 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901708.

Anonymous:1999:NVG

[1267] Anonymous. Netscape vulnerability gives attackers full control. Network Security, 1999(10):2, October 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899901691.

Anonymous:1999:UGR

[1268] Anonymous. Users gain root privilege with wu-ftpd. Network Security, 1999(10):2–3, October 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589990171X.

Anonymous:1999:ULE

[1269] Anonymous. US lifts export curbs on encryption. Network Security, 1999(10):3, October 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901721.

Anonymous:1999:JGS

[1270] Anonymous. Japanese group simplifies hacker detection. Network Security, 1999(10):3, October 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901733.

Anonymous:1999:CCL

[1271] Anonymous. Continuous compliance with legislation for UK companies. Network Security, 1999 (10):3–4, October 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901745.

Anonymous:1999:IPB

[1272] Anonymous. IBM’s PC 300PL to become standard technology. Network Security, 1999(10):4, October 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800882690.

Anonymous:1999:ISS

[1273] Anonymous. IP service switch prevents fragmentation. Network Security, 1999(10):4, October 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800882707.

Gengler:1999:UAT

[1274] Barbara Gengler. US Army thinks different. Network Security, 1999 (10):5, October 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800882719.

Wehde:1999:VWT

[1275] Ed Wehde. Vendors work towards LAN security. Network Security, 1999 (10):5–6, October 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800882720.

Gengler:1999:PSC

[1276] Barbara Gengler. Politicians speak out on cyberterrorism. Network Security, 1999(10):6, October 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800882732.

Cohen:1999:MNSj

[1277] Fred Cohen. Managing network security: Security education in the information age. Network Security, 1999 (10):7–10, October 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800882744.

Forte:1999:FST

[1278] Dario Forte. Free security tools: Pros and cons: a review of free security tools. Network Security, 1999 (10):10–13, October 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901757.

Hancock:1999:ANE

[1279] Bill Hancock. Auditing the network environment at a technical level: Why’s, how’s and aha!’s. Network Security, 1999(10):13–17, October 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899901769.

Hancock:1999:HCN

[1280] Bill Hancock. Health care and network security: Protecting patient privacy. Network Security, 1999 (10):17–19, October 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901770.

Anonymous:1999:Ek

[1281] Anonymous. Events. Network Security, 1999(10):19–20, October 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899901782.

Anonymous:1999:WNV

[1282] Anonymous. Windows NT 4.0 vulnerability. Network Security, 1999 (11):2, November 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901253.

Anonymous:1999:MID

[1283] Anonymous. Microsoft IE5: ‘download behaviour’ vulnerability. Network Security, 1999(11):2, November 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485899901265.

Anonymous:1999:JDE

[1284] Anonymous. Jet database engine vulnerability. Network Security, 1999 (11):2–3, November 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901277.

Anonymous:1999:MPC

[1285] Anonymous. Microsoft and Protek collaboration meets NATO messaging standards. Network Security, 1999 (11):3, November 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901289.

Anonymous:1999:SGS

[1286] Anonymous. Secure e-government for Spain. Network Security, 1999 (11):3, November 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901290.

Anonymous:1999:TUP

[1287] Anonymous. Tripwire 2.2 for the Unix platform. Network Security, 1999 (11):3–4, November 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901307.

Anonymous:1999:FOS

[1288] Anonymous. Free open-source code to result in global private network. Network Security, 1999 (11):4, November 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901319.

Anonymous:1999:SSW

[1289] Anonymous. Secure solution for the wireless market. Network Security, 1999(11):4–5, November 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901320.

Gengler:1999:IS

[1290] Barbara Gengler. IPSec specification. Network Security, 1999(11): 5–6, November 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580080001X.

Gengler:1999:SU

[1291] Barbara Gengler. SNMP update. Network Security, 1999 (11):6, November 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800021.

Cohen:1999:MNSk

[1292] Fred Cohen. Managing network security: The limits of cryptography. Network Security, 1999(11): 7–11, November 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800033.

Wright:1999:EAE

[1293] Marie A. Wright. The evolution of the Advanced Encryption Standard. Network Security, 1999(11): 11–14, November 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800045.

David:1999:IR

[1294] Jon David. Incident response. Network Security, 1999(11):15–18, November 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800800057.

Madsen:1999:TC

[1295] Wayne Madsen. Trust in cyberspace. Network Security, 1999 (11):18–19, November 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800800069.

Anonymous:1999:El

[1296] Anonymous. Events. Network Security, 1999(11):20, November 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899901332.

Anonymous:1999:MII

[1297] Anonymous. Microsoft improves ISN generation. Network Security, 1999 (12):2, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899903239.

Anonymous:1999:MVM

[1298] Anonymous. Microsoft virtual machine vulnerability. Network Security, 1999(12):2, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899903240.

Anonymous:1999:CPS

[1299] Anonymous. Companies promote secure wireless e-commerce. Network Security, 1999(12):2, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485899903252.

Anonymous:1999:FSM

[1300] Anonymous. Fraud screening for emerchants. Network Security, 1999 (12):3, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485899903264.

Anonymous:1999:SPSa

[1301] Anonymous. Survey places security as top barrier to e-commerce. Network Security, 1999(12):3, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485899903288.

Anonymous:1999:CPP

[1302] Anonymous. Companies provide PKI SSO integration for secure business. Network Security, 1999 (12):3, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899903276.

Anonymous:1999:ANZ

[1303] Anonymous. Australia and New Zealand join global security standard. Network Security, 1999(12): 3–4, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348589990329X.

Anonymous:1999:DOO

[1304] Anonymous. Distributor offers one stop shop for smart cards. Network Security, 1999(12):4, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800872414.

Anonymous:1999:FCP

[1305] Anonymous. Flexible copyright protection for Web content. Network Security, 1999(12):4, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/1353485899903306.

Anonymous:1999:ARN

[1306] Anonymous. Automatic repair for network security holes. Network Security, 1999(12):4, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800872426.

Anonymous:1999:SPSb

[1307] Anonymous. Secure payment solution for e-commerce. Network Security, 1999(12):4–5, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800872438.

Gengler:1999:FPP

[1308] Barbara Gengler. The future of the palm print. Network Security, 1999 (12):5, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580087244X.

Gengler:1999:EEL

[1309] Barbara Gengler. Encryption export laws. Network Security, 1999 (12):5–6, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800872451.

Wehde:1999:DHP

[1310] Ed Wehde. DVD hack program on Web. Network Security, 1999 (12):6–7, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800872463.

Cohen:1999:MNSl

[1311] Fred Cohen. Managing network security: Why it was done that way. Network Security, 1999(12): 7–9, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800872475.

Webb:1999:UAO

[1312] Alison Webb. User authentication: Options in Oracle. Network Security, 1999 (12):10–14, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800872487.

Miller:1999:DLD

[1313] Nigel Miller. Deep linking — deep trouble. Network Security, 1999(12): 15–16, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800872499.

Miller:1999:ILC

[1314] Nigel Miller. Internet lotteries considered — taking a spanner to the microchip. Network Security, 1999 (12):16–19, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800872505.

Anonymous:1999:Em

[1315] Anonymous. Events. Network Security, 1999(12):20, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485899903318.

Anonymous:2000:EB

[1316] Anonymous. Editorial board. Network Security, 2000(1):1, January 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800900230.

Anonymous:2000:PESa

[1317] Anonymous. PGP encryption software granted global export license. Network Security, 2000(1):1, January 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800900242.

Anonymous:2000:AET

[1318] Anonymous. Alliance enhances trust in PCs. Network Security, 2000 (1):2, January 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800900254.

Anonymous:2000:SMP

[1319] Anonymous. Secure mobile e-payment system for Europe. Network Security, 2000(1):2, January 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800900266.

Anonymous:2000:PSM

[1320] Anonymous. PKI solution for Microsoft smart card for Windows. Network Security, 2000 (1):2–3, January 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800900278.

Anonymous:2000:PAI

[1321] Anonymous. Protect against Internet intruders and hackers. Network Security, 2000(1):3, January 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580090028X.

Anonymous:2000:USV

[1322] Anonymous. Unix server vulnerability. Network Security, 2000 (1):3, January 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800900308.

Anonymous:2000:SSV

[1323] Anonymous. Solaris snoop vulnerability. Network Security, 2000 (1):3, January 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580090031X.

Anonymous:2000:NTI

[1324] Anonymous. New tool identifies the root cause of network vulnerability. Network Security, 2000 (1):3, January 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800900291.

Sterlicchi:2000:SCD

[1325] John Sterlicchi. Software companies disappointed by encryption draft. Network Security, 2000 (1):4, January 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800866477.

Gengler:2000:JBM

[1326] Barbara Gengler. JavaScript ban may happen. Network Security, 2000(1):4, January 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800866489.

Townsend:2000:CLL

[1327] Kevin Townsend. Carraig Ltd: Latches for Windows. Network Security, 2000 (1):5–7, January 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800866490.

Blatchford:2000:ISBa

[1328] Clive Blatchford. Information security, business and the Internet — Part 1. Network Security, 2000 (1):8–12, January 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800866507.

Humphreys:2000:SW

[1329] Ted Humphreys. Signing the eword. Network Security, 2000 (1):13, January 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800866519.

Barnett:2000:TCS

[1330] Steve Barnett. Top 10 challenges to securing a network. Network Security, 2000(1):14–16, January 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800866520.

Boni:2000:NMO

[1331] Bill Boni. New means — old crimes. Network Security, 2000 (1):17, January 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800866532.

Cohen:2000:DF

[1332] Fred Cohen. Digital forensics. Network Security, 2000(1):18–20, January 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800866544.

Anonymous:2000:ECa

[1333] Anonymous. Events calendar. Network Security, 2000(1):20, January 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800900321.

Anonymous:2000:TCD

[1334] Anonymous. TFN2K causes denial-of-service attack. Network Security, 2000 (2):1–2, February 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800020018.

Anonymous:2000:SSG

[1335] Anonymous. Security scanning guidelines for Singapore ISPs. Network Security, 2000(2):2, February 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580002002X.

Scarrott:2000:CDC

[1336] Paul Scarrott. Consortium to define common standards. Network Security, 2000(2):2, February 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800020031.

Anonymous:2000:NAP

[1337] Anonymous. Network Associates plans to divide. Network Security, 2000 (2):3, February 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800020043.

Anonymous:2000:RSW

[1338] Anonymous. Radicchio to secure wireless e-commerce. Network Security, 2000(2):3, February 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800020055.

Hough:2000:AHE

[1339] Jill Hough. Adapters help to enable LAN security. Network Security, 2000 (2):3, February 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800020079.

Anonymous:2000:HPS

[1340] Anonymous. HP provides security for Microsoft Web servers. Network Security, 2000(2):3, February 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800020067.

Wehde:2000:MBS

[1341] Ed Wehde. Microsoft to beef up SQL server security. Network Security, 2000 (2):4, February 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800876515.

Gengler:2000:UIR

[1342] Barbara Gengler. US ISP requires password changes. Network Security, 2000(2):4, February 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800876527.

Gengler:2000:CPC

[1343] Barbara Gengler. Check Point creates broadband Internet company. Network Security, 2000(2): 4–5, February 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800876539.

Wehde:2000:IME

[1344] Ed Wehde. IBM, Microsoft in encryption effort. Network Security, 2000 (2):5, February 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800876540.

Schultz:2000:IDR

[1345] E. Eugene Schultz. Intrusion detection revisited. Network Security, 2000 (2):6–9, February 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800876552.

Blatchford:2000:ISBb

[1346] Clive Blatchford. Information security, business and the Internet — Part 2. Network Security, 2000(2): 10–14, February 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800876564.

Boni:2000:RC

[1347] Bill Boni. The Russian connection. Network Security, 2000(2): 15–16, February 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800876576.

Anonymous:2000:CSS

[1348] Anonymous. E-commerce survey shows that security is the greatest concern. Network Security, 2000 (2):16, February 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800020158.

Cohen:2000:WM

[1349] Fred Cohen. Worker monitoring. Network Security, 2000(2):17– 19, February 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800876588.

Anonymous:2000:DSP

[1350] Anonymous. Denial-of-service with PMTU strategy. Network Security, 2000(3):1, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800030014.

Anonymous:2000:VMR

[1351] Anonymous. Vulnerability in Microsoft RTF files. Network Security, 2000(3):2, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800030026.

Anonymous:2000:SFR

[1352] Anonymous. Service failure resulting from attack. Network Security, 2000(3):2, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800030038.

Anonymous:2000:PA

[1353] Anonymous. Prevention of DoS attacks. Network Security, 2000 (3):2–3, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580003004X.

Anonymous:2000:EES

[1354] Anonymous. European encryption still safer than US versions. Network Security, 2000(3):3, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800030051.

Dierickx:2000:EGD

[1355] Jüergen Dierickx. European guidelines for digital signatures. Network Security, 2000(3):3–4, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485800030063.

Anonymous:2000:CPF

[1356] Anonymous. Companies provide firewall and VPN solution. Network Security, 2000(3):4, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800030075.

Anonymous:2000:NEF

[1357] Anonymous. National extranet further secured. Network Security, 2000 (3):4, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800030087.

Anonymous:2000:ALS

[1358] Anonymous. Aussies lack security for e-commerce. Network Security, 2000 (3):4–5, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800030099.

Anonymous:2000:FSM

[1359] Anonymous. Further security for mobile transactions. Network Security, 2000(3):5, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800030117.

Anonymous:2000:MCL

[1360] Anonymous. Mobile e-commerce links secured. Network Security, 2000 (3):5, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800030105.

Anonymous:2000:WCS

[1361] Anonymous. Wireless e-commerce security strengthened. Network Security, 2000(3):5–6, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800030129.

Anonymous:2000:VOS

[1362] Anonymous. VPN offers secure Internet connectivity. Network Security, 2000(3):6, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800030142.

Bunker:2000:SOS

[1363] Lucy Bunker. Symantec offers safer Internet. Network Security, 2000 (3):6, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800030154.

Crenshaw:2000:SPK

[1364] Scott Crenshaw. Speedy public key cryptography system. Network Security, 2000(3):6, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800030130.

Wehde:2000:ATT

[1365] Ed Wehde. AOL in trouble on two fronts. Network Security, 2000 (3):7, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800880228.

Sterlicchi:2000:BCA

[1366] John Sterlicchi. BugNet confirms Active Directory bug. Network Security, 2000(3):7–8, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580088023X.

Madsen:2000:RDU

[1367] Wayne Madsen. Revised draft US crypto export regulations leaked. Network Security, 2000(3):8, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800880241.

Townsend:2000:RML

[1368] Kevin Townsend. Reflex Magnetics Ltd: MailSafe. Network Security, 2000 (3):9–11, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800880253.

Forte:2000:ASP

[1369] Dario Forte. Auditing and security policy: The cornerstone of company information protection. Network Security, 2000(3):12–13, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800880265.

Hughes:2000:ECT

[1370] John Hughes. Enabling e-commerce through PKI. Network Security, 2000 (3):14–16, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800880277.

Cohen:2000:CDa

[1371] Fred Cohen. Collaborative defence. Network Security, 2000(3): 17–20, March 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800880289.

Anonymous:2000:SMW

[1372] Anonymous. ‘Screen-melting’ worm on the loose. Network Security, 2000(4):1, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800040010.

Anonymous:2000:THD

[1373] Anonymous. Teenage hacker denies Web attacks. Network Security, 2000(4):2, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800040022.

Anonymous:2000:CCS

[1374] Anonymous. Credit cards stolen through known weakness. Network Security, 2000(4):2, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485800040034.

Anonymous:2000:USR

[1375] Anonymous. Unsafe sites reported. Network Security, 2000 (4):2–3, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800040046.

Anonymous:2000:SAE

[1376] Anonymous. Sophos attempts to educate Internet users. Network Security, 2000(4):3, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800040058.

Anonymous:2000:EVA

[1377] Anonymous. Exploitable vulnerability in atsar. Network Security, 2000(4):3, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800040071.

Anonymous:2000:CAB

[1378] Anonymous. Clip Art buffer overflow vulnerability. Network Security, 2000(4):3, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580004006X.

Anonymous:2000:WLR

[1379] Anonymous. wmcdplay local root compromise. Network Security, 2000 (4):3–4, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800040083.

Anonymous:2000:MTF

[1380] Anonymous. Managing traffic flow to stop DOS attack. Network Security, 2000(4):4, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800040101.

Anonymous:2000:VF

[1381] Anonymous. Vulnerability in FireWall-1. Network Security, 2000(4):4, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485800040095.

Anonymous:2000:CRR

[1382] Anonymous. China relaxes rules on encryption products. Network Security, 2000(4):4–5, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800040113.

Anonymous:2000:CWSa

[1383] Anonymous. Chinese Web site hacked. Network Security, 2000 (4):5, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800040149.

Anonymous:2000:BBS

[1384] Anonymous. Baltimore buys stake in Japanese market. Network Security, 2000(4):5, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800040125.

Curran:2000:RID

[1385] John Curran. Rapid integration of digital certificates. Network Security, 2000(4):5, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800040137.

Anonymous:2000:CPSa

[1386] Anonymous. Companies provide secure online brokering. Network Security, 2000(4):5–6, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800040150.

Anonymous:2000:FVT

[1387] Anonymous. Firewall vendors target smaller companies. Network Security, 2000(4):6, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800040162.

Anonymous:2000:VSM

[1388] Anonymous. Virus scanner for email as it arrives. Network Security, 2000(4):6, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800886183.

Anonymous:2000:GSA

[1389] Anonymous. Gigabit security appliance launched. Network Security, 2000(4):6, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800886195.

Anonymous:2000:JBS

[1390] Anonymous. Java-based secure application environment created. Network Security, 2000(4):6, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485800886201.

Hickson:2000:GSS

[1391] Nigel Hickson. Government support for secure electronic commerce. Network Security, 2000(4): 7–10, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800886213.

Anonymous:2000:WWY

[1392] Anonymous. Where will your ecommerce disputes be resolved? Network Security, 2000(4):11–12, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485800886225.

Landrok:2000:CCV

[1393] Peter Landrok. Challenging the conventional view of PKI: Will it really work? Network Security, 2000 (4):12–15, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800886237.

Cohen:2000:CDb

[1394] Fred Cohen. Countering DCAs. Network Security, 2000(4):15–19, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485800886249.

Boni:2000:HCL

[1395] Bill Boni. Hackers, crackers, lawyers and other dangers. Network Security, 2000(4):19–20, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800886250.

Anonymous:2000:MCG

[1396] Anonymous. Malicious codes get more mobility. Network Security, 2000(5):1, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800050017.

Anonymous:2000:BCS

[1397] Anonymous. Bug could slow down servers. Network Security, 2000 (5):1–2, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800050029.

Anonymous:2000:TPW

[1398] Anonymous. Tighter permissions for Windows Registry key. Network Security, 2000(5):2, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800050030.

Anonymous:2000:CAR

[1399] Anonymous. Cyber attacks reported by 90% of survey respondents. Network Security, 2000 (5):2–3, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800050042.

McCathie:2000:ISF

[1400] Andy McCathie. Internet security fears prompt IT security measures. Network Security, 2000 (5):3, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800050054.

Anonymous:2000:IAW

[1401] Anonymous. Internet attacks will replace military conflicts. Network Security, 2000(5):3, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800050078.

Anonymous:2000:CLS

[1402] Anonymous. 2:20% of companies liable to suffer DoS attacks. Network Security, 2000(5):3, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800050066.

Anonymous:2000:GBE

[1403] Anonymous. Governments back down on encryption regulations. Network Security, 2000(5):3–4, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580005008X.

Smith:2000:CES

[1404] Tim Smith. Company expands to support PKI demands. Network Security, 2000(5):4, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800050108.

Anonymous:2000:PTD

[1405] Anonymous. Privacy threatened by digital signatures. Network Security, 2000(5):4, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800050091.

Zafar:2000:ACB

[1406] Naeem Zafar. Authentication company buys smart card firm. Network Security, 2000(5):4–5, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580005011X.

Decaux:2000:EOS

[1407] Christophe Decaux. European online stores secured. Network Security, 2000(5):5, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800050121.

Bridson:2000:CFB

[1408] David Bridson. Companies form ebusiness alliance. Network Security, 2000(5):5, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800050133.

Anonymous:2000:CPSb

[1409] Anonymous. E-commerce portals securely managed. Network Security, 2000(5):5–6, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800050145.

StPierre:2000:TFA

[1410] Jim St. Pierre. Two-factor authentication added to PKI solutions. Network Security, 2000 (5):6, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800050182.

Anonymous:2000:SES

[1411] Anonymous. Signing and encryption software system launched. Network Security, 2000(5):6, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800050170.

Anonymous:2000:EPS

[1412] Anonymous. Enhancement for PKI security. Network Security, 2000 (5):6, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800050169.

Anonymous:2000:VPS

[1413] Anonymous. Visitor’s PC scanned before trading. Network Security, 2000(5):6, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800050157.

Anonymous:2000:SEW

[1414] Anonymous. Security enhancements for wireless VPN connections. Network Security, 2000 (5):6–7, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800050194.

Ott:2000:C

[1415] Rachael Ott. Cybersquatting. Network Security, 2000(5):7, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800050212.

Anonymous:2000:PGO

[1416] Anonymous. Product guarantees origin of e-mail. Network Security, 2000(5):7, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800050200.

Helmich:2000:PKI

[1417] Paul Helmich. Public key infrastructures: a panacea solution? Network Security, 2000(5):8–11, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800050224.

Corell:2000:TRP

[1418] Simon Corell. Ten risks of PKI: In favour of smart card-based PKI. Network Security, 2000(5):12–14, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800050236.

Cohen:2000:EIA

[1419] Fred Cohen. Eliminating IP address forgery — five years old and going strong. Network Security, 2000 (5):15–18, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800050248.

Boni:2000:MTC

[1420] Bill Boni. The more things change, the more they stay the same!: Headline: “DTI Survey Finds Most UK Companies Experienced Security Breaches.” Headline: “FBI Study Finds U.S. Losses to Computer Crime Increased 100%.”. Network Security, 2000 (5):18–19, May 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580005025X.

Anonymous:2000:NVC

[1421] Anonymous. NewLove virus changes to avoid detection. Network Security, 2000(6):1, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800060013.

Anonymous:2000:VSU

[1422] Anonymous. Virus spreads unopened. Network Security, 2000 (6):1–2, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800060025.

Anonymous:2000:BAC

[1423] Anonymous. Bugs allow compromising of SSL sessions. Network Security, 2000(6):2, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800060037.

Anonymous:2000:SFH

[1424] Anonymous. Server forced to halt and then reload. Network Security, 2000(6):2–3, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800060049.

Anonymous:2000:EFP

[1425] Anonymous. Experts find potential for cookie stealing. Network Security, 2000(6):3, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800060050.

Anonymous:2000:IFR

[1426] Anonymous. IP fragment reassembly vulnerability. Network Security, 2000(6):3–4, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800060062.

Moore:2000:ZLP

[1427] Carol Moore. Zaxus launches PKI solution for e-commerce. Network Security, 2000(6):4, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800060086.

Parker:2000:CUS

[1428] Sara Parker. Companies unite to secure EBPP market. Network Security, 2000(6):4, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800060098.

Anonymous:2000:UIS

[1429] Anonymous. UK initiative to secure e-commerce. Network Security, 2000(6):4, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800060074.

Anonymous:2000:CLI

[1430] Anonymous. Convenience lost for increased security. Network Security, 2000(6):4–5, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800060104.

Smith:2000:ABS

[1431] Tim Smith. Authentication by biometric smart card. Network Security, 2000(6):5, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800060116.

Anonymous:2000:BSF

[1432] Anonymous. Backup to save files from virus. Network Security, 2000 (6):5, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580006013X.

Anonymous:2000:CVT

[1433] Anonymous. Creation of VPN tunnels simplified. Network Security, 2000(6):5, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800060128.

Bunker:2000:AVT

[1434] Lucy Bunker. Anti-virus technology for Yahoo! Network Security, 2000(6):5–6, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800060141.

Anonymous:2000:LB

[1435] Anonymous. Love bugged! Network Security, 2000(6):6, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800060153.

Spalding:2000:DWU

[1436] Michael Spalding. Deciding whether or not to use a third party certificate authority. Network Security, 2000(6):7–8, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800060165.

Tobias:2000:LBD

[1437] Henk Tobias. To be or not to be — legally binding digital certificates. Network Security, 2000 (6):9–11, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800060177.

Venter:2000:NSI

[1438] H. S. Venter and J. H. P. Eloff. Network security: Important issues. Network Security, 2000(6): 12–16, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800060189.

Cohen:2000:WCW

[1439] Fred Cohen. Why can’t we do DNS right? Network Security, 2000 (6):17–18, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800060190.

Boni:2000:ODD

[1440] Bill Boni. Ode to ‘digital defenders’. Network Security, 2000 (6):19, June 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800060207.

Anonymous:2000:FFC

[1441] Anonymous. Flaw in firewall could lead to DoS attack. Network Security, 2000(7):1, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580007001X.

Anonymous:2000:SBT

[1442] Anonymous. The Serbian Badman Trojan. Network Security, 2000 (7):1–2, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800070021.

Anonymous:2000:IDS

[1443] Anonymous. IPCS denial-of-service attack. Network Security, 2000 (7):2, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800070033.

Anonymous:2000:KRA

[1444] Anonymous. Kerberos at risk from DoS attacks. Network Security, 2000(7):3, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800070045.

Anonymous:2000:SCV

[1445] Anonymous. SSL certificate validation vulnerability. Network Security, 2000(7):3, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800070057.

Anonymous:2000:ACV

[1446] Anonymous. AIX cdmount vulnerability. Network Security, 2000 (7):3–4, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800070069.

Anonymous:2000:NVR

[1447] Anonymous. New viruses run without being opened. Network Security, 2000(7):4, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800070070.

Anonymous:2000:PVP

[1448] Anonymous. Powerful virus protection through partnership. Network Security, 2000(7):4, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800070082.

Anonymous:2000:BGA

[1449] Anonymous. BT gives anti-virus contract to NAI. Network Security, 2000(7):4, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800070094.

Anonymous:2000:APE

[1450] Anonymous. Alliance provides enhanced Internet security. Network Security, 2000(7):4–5, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800070100.

Anonymous:2000:ULC

[1451] Anonymous. UK launches code of best practice for e-business. Network Security, 2000(7):5, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800070112.

Anonymous:2000:PESb

[1452] Anonymous. Product enhances secure file transfers. Network Security, 2000(7):5, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800070124.

Anonymous:2000:ANT

[1453] Anonymous. Attempt number two to protect against ILOVEYOU. Network Security, 2000(7):5, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800070136.

Anonymous:2000:AVT

[1454] Anonymous. Anti-virus technology for the Palm OS platform. Network Security, 2000(7):6, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800070148.

Anonymous:2000:BUP

[1455] Anonymous. Businesses urged to protect performance. Network Security, 2000(7):6, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580007015X.

Anonymous:2000:ICF

[1456] Anonymous. Intelligent content filtering offering. Network Security, 2000(7):6, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800070161.

Webb:2000:AEM

[1457] Alison Webb. Auditing electronic mail. Network Security, 2000(7): 7–10, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800070173.

Wright:2000:VPN

[1458] Marie A. Wright. Virtual private network security. Network Security, 2000 (7):11–14, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800070185.

Cohen:2000:WDI

[1459] Fred Cohen. What does it do behind your back? Network Security, 2000(7):15–17, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800070197.

Boni:2000:WDW

[1460] Bill Boni. What to do when the worst happens! Network Security, 2000(7):18–19, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800070203.

Dwan:2000:TBS

[1461] Berni Dwan. Time-based security.

Network Security, 2000(7):19, July 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800070215.

Anonymous:2000:DUC

[1462] Anonymous. Don’t use the ‘Crayon of Doom’. Network Security, 2000 (8):1–2, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800080016.

Anonymous:2000:ASD

[1463] Anonymous. “Active Setup Download” vulnerability. Network Security, 2000(8):2, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800080028.

Anonymous:2000:LWA

[1464] Anonymous. LISTSERV Web archive remote overflow. Network Security, 2000(8):2, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580008003X.

Anonymous:2000:OUS

[1465] Anonymous. OpenSSH UseLogin security risk. Network Security, 2000 (8):2–3, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800080041.

Anonymous:2000:PFT

[1466] Anonymous. PIX firewall TCP reset vulnerability. Network Security, 2000(8):3, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800080053.

Anonymous:2000:PCF

[1467] Anonymous. Powergen customer finds debit card details. Network Security, 2000(8):3, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800080065.

Anonymous:2000:HVC

[1468] Anonymous. Hackers/viruses cost businesses $1.6 trillion. Network Security, 2000(8):3–4, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800080077.

Anonymous:2000:IPS

[1469] Anonymous. Internet poses serious threat to national security. Network Security, 2000(8):4, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800080089.

Anonymous:2000:IWH

[1470] Anonymous. Investigation into wireless hacks. Network Security, 2000 (8):4, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800080090.

Anonymous:2000:MCB

[1471] Anonymous. Mission-critical e-business security. Network Security, 2000 (8):4, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800080107.

Anonymous:2000:SWA

[1472] Anonymous. Secure wireless access to applications. Network Security, 2000(8):4, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800080119.

Anonymous:2000:MIU

[1473] Anonymous. Microsoft and ISS unite to secure ISA Server 2000. Network Security, 2000(8):5, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485800080120.

Anonymous:2000:NWS

[1474] Anonymous. New Web security solution. Network Security, 2000 (8):5, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800080132.

Anonymous:2000:TPE

[1475] Anonymous. Tool prevents employee computer misuse. Network Security, 2000(8):5, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800080144.

Anonymous:2000:FSL

[1476] Anonymous. Firewall suitable for large data centres. Network Security, 2000(8):5, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800080156.

Wheeler:2000:DSC

[1477] Graham Wheeler. Denial-of-service: courting disaster. Network Security, 2000(8):6, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800080168.

Madsen:2000:WUN

[1478] Wayne Madsen. Whitehouse unveils new cyber-crime, crypto export policies. Network Security, 2000 (8):7, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580008017X.

Cox:2000:EWI

[1479] Peter Cox. Exactly what it says on the Tin … says who? Network Security, 2000(8):8–10, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800080181.

Cohen:2000:UVB

[1480] Fred Cohen. Understanding viruses bio-logically. Network Security, 2000 (8):11–16, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800080193.

Boni:2000:UMG

[1481] Bill Boni. The US mob gets savvy to cyber crime. Network Security, 2000 (8):17–18, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580008020X.

Schultz:2000:BRI

[1482] E. Eugene Schultz. Book review:

Intrusion Detection: Rebecca Bace [Macmillan Technical Publishing, Indianapolis, IN, USA, ISBN 1-57870-185-6]. Network Security, 2000 (8):19, August 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800080211.

Anonymous:2000:BIW

[1483] Anonymous. 32-bit Internet worm ‘Sysid.exe’. Network Security, 2000 (9):1, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800090012.

Anonymous:2000:MPT

[1484] Anonymous. Mobile phone targeted by virus writers. Network Security, 2000 (9):1–2, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800090024.

Anonymous:2000:TGH

[1485] Anonymous. Trojan gives hackers control. Network Security, 2000(9): 2, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800090036.

Anonymous:2000:WSC

[1486] Anonymous. Weak security for ecommerce servers. Network Security, 2000(9):2–3, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485800090048.

Anonymous:2000:JAR

[1487] Anonymous. Java applets read protected resources. Network Security, 2000(9):3, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580009005X.

Anonymous:2000:IOM

[1488] Anonymous. IE opens Microsoft Access and executes code. Network Security, 2000(9):3, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800090061.

Anonymous:2000:RLC

[1489] Anonymous. Remote and local compromise allowed. Network Security, 2000 (9):3–4, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800090073.

Anonymous:2000:LCV

[1490] Anonymous. Linux capability vulnerability. Network Security, 2000 (9):4, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800090085.

Anonymous:2000:PAV

[1491] Anonymous. PGP ADKs vulnerability. Network Security, 2000(9): 4, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800090097.

Anonymous:2000:WWV

[1492] Anonymous. WorldView Wnn vulnerability. Network Security, 2000 (9):4, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800090103.

Anonymous:2000:FWG

[1493] Anonymous. FBI wants global Internet security organization. Network Security, 2000(9):4–5, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485800090115.

Anonymous:2000:BNS

[1494] Anonymous. Banking on network security basics. Network Security, 2000 (9):5, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800090127.

Anonymous:2000:OSH

[1495] Anonymous. Online shopping hindered by security fears. Network Security, 2000(9):5, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800090139.

Anonymous:2000:USI

[1496] Anonymous. Ultimo secures insurance group. Network Security, 2000 (9):5, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800090140.

Anonymous:2000:BNW

[1497] Anonymous. Bolero.net and WISekey join forces. Network Security, 2000 (9):5, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800090152.

Anonymous:2000:VPM

[1498] Anonymous. Virus protection for mobile phones and handheld computers. Network Security, 2000(9): 5–6, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800090164.

Anonymous:2000:SMT

[1499] Anonymous. Secure e-mail technology unveiled. Network Security, 2000 (9):6, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800090176.

Anonymous:2000:PPV

[1500] Anonymous. Plug-and-play VPN security box. Network Security, 2000 (9):6, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800090188.

Anonymous:2000:AIH

[1501] Anonymous. AES IP hardware encryptor introduced. Network Security, 2000 (9):6–7, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580009019X.

Anonymous:2000:UAS

[1502] Anonymous. User authentication via smart card. Network Security, 2000 (9):7, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800090206.

Anonymous:2000:MSSa

[1503] Anonymous. Managed security service enables outsourcing of security. Network Security, 2000(9): 7, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800090218.

Anonymous:2000:ICO

[1504] Anonymous. iSolve clamp down on online credit card fraud. Network Security, 2000(9):7, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580009022X.

Gengler:2000:UPC

[1505] Barbara Gengler. US President Clinton signs Digital Signature Bill. Network Security, 2000(9):7–8, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800090231.

Donnelly:2000:UNL

[1506] Allan Donnelly. US national lab fights viruses. Network Security, 2000 (9):8, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800090243.

Kaliski:2000:CNP

[1507] Burt Kaliski. Considerations for new public-key algorithms. Network Security, 2000(9):9–10, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800090255.

Syed:2000:CLA

[1508] Furqan Syed. Children of DES: a look at the Advanced Encryption Standard. Network Security, 2000(9):11– 12, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800090267.

Wright:2000:IQC

[1509] Marie A. Wright. The impact of quantum computing on cryptography. Network Security, 2000(9):13– 15, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800090279.

Cohen:2000:C

[1510] Fred Cohen. Chipping. Network Security, 2000(9):16–17, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800090280.

Boni:2000:CPC

[1511] Bill Boni. Carnivores, predators and civil liberties. Network Security, 2000 (9):18–19, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800090292.

Anonymous:2000:FSF

[1512] Anonymous. Federal security found lacking. Network Security, 2000 (10):1, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800100017.

Anonymous:2000:IIR

[1513] Anonymous. Indian IT out of reach. Network Security, 2000(10): 2, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800100029.

Anonymous:2000:ILL

[1514] Anonymous. International legislation at last. Network Security, 2000 (10):2, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800100030.

Anonymous:2000:PSA

[1515] Anonymous. Privacy suit against About.com. Network Security, 2000 (10):2, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800100042.

Anonymous:2000:CWSb

[1516] Anonymous. Czech W2K.Stream virus discovered. Network Security, 2000 (10):2–3, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800100054.

Anonymous:2000:VIS

[1517] Anonymous. Virus inoculation by satellite. Network Security, 2000 (10):3, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800100066.

Anonymous:2000:HNI

[1518] Anonymous. Home networking on increase. Network Security, 2000 (10):3, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800100078.

Anonymous:2000:EPE

[1519] Anonymous. Equant provides extranet. Network Security, 2000(10): 3–4, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580010008X.

Anonymous:2000:MIE

[1520] Anonymous. Music industry enlists help of hackers. Network Security, 2000(10):4, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800100091.

Anonymous:2000:CP

[1521] Anonymous. Crime pays. Network Security, 2000(10):4, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800100108.

Anonymous:2000:CCF

[1522] Anonymous. 43% of credit card fraud not reported. Network Security, 2000 (10):4, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580010011X.

Anonymous:2000:AID

[1523] Anonymous. Advanced intrusion detection. Network Security, 2000 (10):4, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800100121.

Anonymous:2000:MSD

[1524] Anonymous. E-mail scanner detects porn. Network Security, 2000(10): 4–5, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800100133.

Anonymous:2000:VAS

[1525] Anonymous. Voice authentication smart card. Network Security, 2000 (10):5, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800100145.

Anonymous:2000:SCA

[1526] Anonymous. Security a competitive advantage? Network Security, 2000 (10):5–6, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800100157.

Madsen:2000:CVA

[1527] Wayne Madsen. Carnivore’s voracious appetite. Network Security, 2000 (10):6–7, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800100169.

Wheeler:2000:RB

[1528] Graham Wheeler. RIP Bill. Network Security, 2000(10):7–8, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800100170.

Madsen:2000:EMA

[1529] Wayne Madsen. Employer Monitoring Act proposed. Network Security, 2000(10):8, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800100182.

Bond:2000:BBB

[1530] Colin L. Bond. Big Brother has been here all along. Network Security, 2000 (10):9, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800100194.

McDermott:2000:BTO

[1531] Paul McDermott. Building trust into online business. Network Security, 2000 (10):10–12, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800100200.

Osen:2000:TSJ

[1532] Janet Osen. The thorny side of jurisdiction and the Internet. Network Security, 2000(10):13–16, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485800100212.

Cohen:2000:RR

[1533] Fred Cohen. Revisiting risk. Network Security, 2000(10):17–18, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485800100224.

Boni:2000:LD

[1534] Bill Boni. Lights in the darkness. Network Security, 2000(10): 19–20, October 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800100236.

Anonymous:2000:MHL

[1535] Anonymous. Microsoft hacked … twice. Network Security, 2000(11): 1–2, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850011.

Anonymous:2000:HBI

[1536] Anonymous. Hackers boycott industry olive-branch. Network Security, 2000(11):2, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850023.

Anonymous:2000:CKC

[1537] Anonymous. Curiosity killed the CueCat. Network Security, 2000 (11):2, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850035.

Anonymous:2000:AHU

[1538] Anonymous. Anti-hackers to be unleashed. Network Security, 2000 (11):2–3, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850047.

Anonymous:2000:CMI

[1539] Anonymous. China muzzles Internet. Network Security, 2000 (11):3, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850059.

Anonymous:2000:MTG

[1540] Anonymous. Mexicans have trouble getting online. Network Security, 2000(11):3, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850060.

Anonymous:2000:B

[1541] Anonymous. In brief. Network Security, 2000(11):3, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800850084.

Anonymous:2000:IWC

[1542] Anonymous. Ireland wins cyber champion award. Network Security, 2000 (11):3–4, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850072.

Anonymous:2000:CUH

[1543] Anonymous. Columbia University host crystal ball gazing. Network Security, 2000(11):4, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850096.

Anonymous:2000:BKF

[1544] Anonymous. Branding key to financial services. Network Security, 2000(11):4, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850102.

Anonymous:2000:CWA

[1545] Anonymous. CA warns about email interception. Network Security, 2000(11):4, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850114.

Anonymous:2000:CWSc

[1546] Anonymous. Consumers want security. Network Security, 2000 (11):4, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850126.

Anonymous:2000:FWB

[1547] Anonymous. Free Web-based security assessment. Network Security, 2000(11):4, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850138.

Anonymous:2000:WPI

[1548] Anonymous. Working patterns impact security. Network Security, 2000 (11):4–5, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580085014X.

Anonymous:2000:BPD

[1549] Anonymous. BIND 9 plugs DNS. Network Security, 2000(11):5, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800850151.

Anonymous:2000:XMM

[1550] Anonymous. Xinetica monitor modems. Network Security, 2000 (11):5, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850163.

Anonymous:2000:PPE

[1551] Anonymous. Plug-and-play eToken. Network Security, 2000(11):5, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485800850175.

Anonymous:2000:CCW

[1552] Anonymous. Cost-cutting Web server. Network Security, 2000 (11):5, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850187.

Madsen:2000:FRC

[1553] Wayne Madsen. FBI releases Carnivore documents. Network Security, 2000(11):6, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850199.

Donnelly:2000:CBF

[1554] Allan Donnelly. Consumers bombard fraud reporting centre. Network Security, 2000(11):6–7, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800850205.

Gengler:2000:CS

[1555] Barbara Gengler. Cisco’s SAFE. Network Security, 2000(11):7, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800850217.

David:2000:VF

[1556] Jon David. The vanishing firewall. Network Security, 2000(11): 8–11, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850230.

Roberts:2000:TWT

[1557] Graham Roberts. Tangled Web — tales of digital crime from the shadows of cyberspace: Richard Power, Que Corporation, 2000. Network Security, 2000(11):8, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850229.

McDermott:2000:PFL

[1558] Paul McDermott. Personal firewalls … one more step towards comprehensive security. Network Security, 2000 (11):11–14, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850242.

Zieger:2000:BEC

[1559] Melissa Zieger. Boost in electronic courier traffic — traditional UK couriers left stranded in petrol shortage: Internet proved to be best vehicle for secure and trackable document transfer. Network Security, 2000(11): 14–15, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850254.

Cohen:2000:WEK

[1560] Fred Cohen. Why everything keeps failing. Network Security, 2000(11): 16–18, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850266.

Boni:2000:MM

[1561] Bill Boni. Meteors and managers. Network Security, 2000(11): 18–19, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800850278.

Anonymous:2000:ECb

[1562] Anonymous. Events calendar. Network Security, 2000(11):20, November 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580085028X.

Anonymous:2000:HHI

[1563] Anonymous. Hybris hits international scene. Network Security, 2000 (12):1, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580012001X.

Anonymous:2000:PMC

[1564] Anonymous. PKI may carry viruses.

Network Security, 2000(12):2, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485800120021.

Anonymous:2000:IAO

[1565] Anonymous. IBM to adopt open source. Network Security, 2000(12): 2, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800120033.

Anonymous:2000:SLH

[1566] Anonymous. Standards lobby on hold. Network Security, 2000(12): 2, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800120045.

Anonymous:2000:NCS

[1567] Anonymous. Non-contact smart cards for M-commerce. Network Security, 2000(12):2–3, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485800120057.

Anonymous:2000:MSSb

[1568] Anonymous. M-spam, spam, spam. Network Security, 2000(12):3, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1353485800120069.

Anonymous:2000:GGT

[1569] Anonymous. Graffiti growing trend. Network Security, 2000(12):3, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1353485800120070.

Anonymous:2000:GLR

[1570] Anonymous. German land registry goes electronic. Network Security, 2000 (12):3, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800120082.

Anonymous:2000:CCC

[1571] Anonymous. Cyber crime can hit directors. Network Security, 2000 (12):4, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800120094.

Anonymous:2000:CYT

[1572] Anonymous. Can you trust your bank? Network Security, 2000 (12):4, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800120100.

Anonymous:2000:BSN

[1573] Anonymous. Biometric screensaver for notebook. Network Security, 2000 (12):4, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800120112.

Anonymous:2000:LSP

[1574] Anonymous. Layered system preferred. Network Security, 2000(12): 4, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800120124.

Madsen:2000:HCI

[1575] Wayne Madsen. Health care industry debate: Electronic versus digital signatures. Network Security, 2000 (12):5, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800120136.

Wales:2000:DEP

[1576] Elspeth Wales. DNSSEC emerges to protect Internet infrastructure. Network Security, 2000(12):5, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485800120148.

Madsen:2000:DPC

[1577] Wayne Madsen. Data protection commissioners explore new ground. Network Security, 2000(12): 6, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580012015X.

McDermott:2000:SIN

[1578] Paul McDermott. Security in IP networks. Network Security, 2000(12): 7–9, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800120161.

Pemble:2000:WYL

[1579] Matthew Pemble. Washing your laundry in public — an analysis of recent high-publicity security incidents. Network Security, 2000(12): 10–12, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485800120173.

Kakalik:2000:PSW

[1580] John S. Kakalik and Marie A. Wright. Privacy and security in wireless computing. Network Security, 2000(12):

12–15, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800120185.

Cohen:2000:MAY

[1581] Fred Cohen. The Millennium article — yet again!: The bots are coming!!! The bots are coming!!! Network Security, 2000(12): 15–18, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485800120197.

Boni:2000:TTH

[1582] Bill Boni. Tricks and treats from the hacker underground! Network Security, 2000(12):18–19, December 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485800120203.

Anonymous:2001:TMC

[1583] Anonymous. Three million credit cards compromised. Network Security, 2001(1):1, January 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801001015.

Anonymous:2001:SBM

[1584] Anonymous. Singapore backs emoney. Network Security, 2001 (1):2, January 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801001027.

Anonymous:2001:PSL

[1585] Anonymous. Port scanning legal in US. Network Security, 2001 (1):2, January 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801001039.

Anonymous:2001:CPL

[1586] Anonymous. Canadian privacy law worries US. Network Security, 2001 (1):2, January 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801001040.

Anonymous:2001:CLN

[1587] Anonymous. Cybercrime laws need overhaul. Network Security, 2001 (1):3, January 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801001052.

Anonymous:2001:Ba

[1588] Anonymous. In brief. Network Security, 2001(1):3, January 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801001076.

Anonymous:2001:NWA

[1589] Anonymous. NIPC worried about DoS. Network Security, 2001(1): 3–4, January 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801001064.

Anonymous:2001:ITP

[1590] Anonymous. Indian teens to police Net. Network Security, 2001 (1):4, January 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801001088.

Anonymous:2001:SS

[1591] Anonymous. Security on scale of 1 to 10. Network Security, 2001 (1):4, January 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580100109X.

Anonymous:2001:FIW

[1592] Anonymous. FTC investigates wireless privacy. Network Security, 2001 (1):4, January 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801001106.

Madsen:2001:CDR

[1593] Wayne Madsen. Carnivore documents reveal enhanced tapping abilities. Network Security, 2001(1): 5, January 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801001118.

Wales:2001:MTC

[1594] Elspeth Wales. McAfee targets corporates with its ASP security software. Network Security, 2001(1): 5–6, January 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580100112X.

Groves:2001:SAS

[1595] Jackie Groves. Security for application service providers. Network Security, 2001(1):6–9, January 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801001131.

Pemble:2001:ATC

[1596] Matthew Pemble. Always trust content from Microsoft Corporation? Network Security, 2001(1): 10–12, January 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801001143.

Brown:2001:VOP

[1597] Arlene Brown. VPNs: Only part of the remote access security solution. Network Security, 2001(1): 12–14, January 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801001155.

Cohen:2001:MHF

[1598] Fred Cohen. Marketing hyperbole at its finest. Network Security, 2001(1): 15–17, January 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801001167.

Boni:2001:YJP

[1599] Bill Boni. The year just past and the one ahead. Network Security, 2001 (1):18–19, January 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801001179.

Anonymous:2001:ECa

[1600] Anonymous. Events calendar. Network Security, 2001(1):20, January 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801001180.

Anonymous:2001:TSB

[1601] Anonymous. TV station blasts hacked smartcards. Network Security, 2001 (2):1, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580100201X.

Anonymous:2001:CM

[1602] Anonymous. Champagne is on MIS. Network Security, 2001(2): 2, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801002021.

Anonymous:2001:MBR

[1603] Anonymous. More businesses reporting piracy. Network Security, 2001 (2):2–3, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801002033.

Anonymous:2001:CPM

[1604] Anonymous. CPRM to protect movies and music. Network Security, 2001 (2):3, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801002045.

Anonymous:2001:Bb

[1605] Anonymous. In brief. Network Security, 2001(2):3, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801002070.

Anonymous:2001:PPN

[1606] Anonymous. PDA policy needed. Network Security, 2001(2):3, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801002057.

Anonymous:2001:MCS

[1607] Anonymous. Marconi cracks secure M-commerce. Network Security, 2001 (2):3–4, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801002069.

Anonymous:2001:BPS

[1608] Anonymous. Buchanan to provide security blanket. Network Security, 2001 (2):4, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801002082.

Anonymous:2001:ESM

[1609] Anonymous. Entercept secures $33 million. Network Security, 2001 (2):4, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801002094.

Anonymous:2001:UAB

[1610] Anonymous. Unisys acquires Balance. Network Security, 2001(2): 4, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801002100.

Anonymous:2001:SGE

[1611] Anonymous. SCM get EMV seal of approval. Network Security, 2001 (2):4, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801002112.

Donnelly:2001:HRR

[1612] Allan Donnelly. How to reduce risks with ActiveX. Network Security, 2001 (2):5, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801002124.

Gengler:2001:WMT

[1613] Barbara Gengler. Workplace monitoring tops privacy hit-list. Network Security, 2001(2):5–6, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801002136.

Madsen:2001:CSB

[1614] Wayne Madsen. Computer security board criticized. Network Security, 2001(2):6, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801002148.

Gengler:2001:HNS

[1615] Barbara Gengler. HP and Nokia secure wireless transactions. Network Security, 2001(2):7, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580100215X.

Anonymous:2001:AVV

[1616] Anonymous. Anti-virus vendors urge Valentine’s day caution. Network Security, 2001(2):7, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801002161.

Roberts:2001:BRI

[1617] Graham Roberts. Book review: Inside Internet Security — What Hackers Don’t Want You To Know: Jeff Crume, Pearson Education Limited, £29.95. Network Security, 2001 (2):8, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801002173.

Forte:2001:PPF

[1618] Dario Forte. Peer-to-peer file sharing is here to stay: Napster’s definitive shift to a pay service means more problems for system administrators. Meanwhile, other similar systems are gaining a foothold. Network Security, 2001 (2):9–11, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801002185.

Pemble:2001:WHI

[1619] Matthew Pemble. W(h)ither regulation? Network Security, 2001(2): 11–12, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801002197.

Hruska:2001:VPG

[1620] Jan Hruska. Is the virus problem getting worse? Network Security, 2001 (2):13–16, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801002203.

Cohen:2001:TYS

[1621] Fred Cohen. Testing your security defence by breaking in?: Maybe not. Network Security, 2001(2):16– 18, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801002215.

Boni:2001:BS

[1622] Bill Boni. Balancing the scales. Network Security, 2001(2):19–20, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801002227.

Anonymous:2001:ECb

[1623] Anonymous. Events calendar. Network Security, 2001(2):20, February 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801002239.

Anonymous:2001:USC

[1624] Anonymous. US says 2600 can’t link to DeCSS. Network Security, 2001 (3):1–2, March 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801003014.

Anonymous:2001:ULA

[1625] Anonymous. UK legislates against hackers as cyber-terrorists. Network Security, 2001(3):2, March 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801003026.

Anonymous:2001:SPA

[1626] Anonymous. Swiss police arrest suspect in WEF hack. Network Security, 2001(3):2–3, March 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801003038.

Anonymous:2001:MBE

[1627] Anonymous. McDonalds in BIND exploitation. Network Security, 2001 (3):3, March 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580100304X.

Anonymous:2001:Bc

[1628] Anonymous. In brief. Network Security, 2001(3):3, March 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801003063.

Anonymous:2001:BPC

[1629] Anonymous. BT profits from cybercrime. Network Security, 2001 (3):3–4, March 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801003051.

Anonymous:2001:SMS

[1630] Anonymous. Stock-market sensitive data not secured. Network Security, 2001(3):4, March 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801003075.

Anonymous:2001:ESF

[1631] Anonymous. EU says fraud has doubled. Network Security, 2001 (3):4, March 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801003087.

Anonymous:2001:UCN

[1632] Anonymous. Users click on Naked Wife. Network Security, 2001 (3):4, March 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801003099.

Donnelly:2001:DSH

[1633] Allan Donnelly. DNS software hole allows Web attacks. Network Security, 2001(3):5, March 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801003105.

Anonymous:2001:SLI

[1634] Anonymous. Security left out of investment decisions. Network Security, 2001(3):5–6, March 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801003117.

Gengler:2001:TCP

[1635] Barbara Gengler. Trusted computing platform alliance. Network Security, 2001(3):6, March 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801003129.

Silver:2001:AFA

[1636] Oscar Silver. AOL files another lawsuit to stop junk e-mail. Network Security, 2001(3):7, March 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801003130.

Anonymous:2001:POE

[1637] Anonymous. Post Office enabled PKI helps kids shop online. Network Security, 2001(3):7, March 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801003142.

Madsen:2001:TWN

[1638] Wayne Madsen. Troubled Waters at NSA. Network Security, 2001 (3):8–10, March 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801003154.

Pemble:2001:PBI

[1639] Matthew Pemble. Poking at the borders: The internationally-minded toad. Network Security, 2001(3): 10–11, March 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801003166.

Cohen:2001:CSI

[1640] Fred Cohen. Corporate security intelligence: An oxymoron? Network Security, 2001(3):12–17, March 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801003178.

Boni:2001:TCS

[1641] Bill Boni. The threat of cybersabotage: The new Internet economy has recently seen a spate of layoffs. Ebusiness had better watch out for their ex-technical employees — and guard against the potentially crippling effects of cyber-sabotage. Network Security, 2001(3):18–19, March 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580100318X.

Anonymous:2001:ECc

[1642] Anonymous. Events calendar. Network Security, 2001(3):20, March 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801003191.

Anonymous:2001:VDI

[1643] Anonymous. VeriSign duped into issuing fraudulent certificates. Network Security, 2001(4):1–2, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801004019.

Anonymous:2001:MIT

[1644] Anonymous. Massive identity theft by NY dish washer. Network Security, 2001(4):2, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801004020.

Anonymous:2001:PFE

[1645] Anonymous. Processes at fault in extortion scam. Network Security, 2001(4):2–3, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801004032.

Anonymous:2001:Bd

[1646] Anonymous. In brief. Network Security, 2001(4):3, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801004056.

Anonymous:2001:OSM

[1647] Anonymous. Open source movement turns tables on IP laws. Network Security, 2001(4):3–4, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801004044.

Anonymous:2001:DPE

[1648] Anonymous. Defamatory poster exposed. Network Security, 2001 (4):4, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801004068.

Anonymous:2001:GRI

[1649] Anonymous. Germany to require ISPs to monitor Web. Network Security, 2001(4):4, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580100407X.

Anonymous:2001:UTF

[1650] Anonymous. UK trials first mobile electronic signatures. Network Security, 2001(4):4, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801004081.

Anonymous:2001:WWS

[1651] Anonymous. Wireless WAN standard gets industry backing. Network Security, 2001(4):5, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801004093.

Anonymous:2001:LEP

[1652] Anonymous. Lion eats penguin. Network Security, 2001(4):5, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580100410X.

Anonymous:2001:VCH

[1653] Anonymous. Virus can hit Linux or Windows. Network Security, 2001(4):5, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801004111.

Silver:2001:WNV

[1654] Oscar Silver. Wireless networks vulnerable to attack. Network Security, 2001(4):6, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801004123.

Gengler:2001:NSS

[1655] Barbara Gengler. Network solutions selling database information. Network Security, 2001(4): 6–7, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801004135.

Madsen:2001:ECE

[1656] Wayne Madsen. EU and Council of Europe showdown looming on Internet surveillance. Network Security, 2001(4):7–8, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801004147.

Anonymous:2001:CBFa

[1657] Anonymous. Cisco e-business forum ‘scared to death’. Network Security, 2001(4):8–9, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801004159.

Kirkby:2001:CCW

[1658] Andrea Kirkby. Cryptography and e-commerce: a Wiley tech brief. Network Security, 2001 (4):9, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801004160.

Pemble:2001:CEW

[1659] Matthew Pemble. Confidentiality:

From encryption, to where? Network Security, 2001(4):10–11, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801004172.

Pohlmann:2001:SCA

[1660] Norbert Pohlmann. Smart cards: The authenticated solution for ebusiness user. Network Security, 2001 (4):12–15, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801004184.

Cohen:2001:PP

[1661] Fred Cohen. To prosecute or not to prosecute? Network Security, 2001 (4):15–18, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801004196.

Boni:2001:BB

[1662] Bill Boni. E-biz blitz. Network Security, 2001(4):18–19, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801004202.

Anonymous:2001:MKS

[1663] Anonymous. My kingdom for a sandbox. Network Security, 2001 (4):20, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801004214.

Anonymous:2001:ECd

[1664] Anonymous. Events calendar. Network Security, 2001(4):20, April 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801004226.

Anonymous:2001:UCF

[1665] Anonymous. US and China fight it out online. Network Security, 2001(5):1–2, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801005013.

Anonymous:2001:MPB

[1666] Anonymous. Microsoft plasters big hole in W2K. Network Security, 2001(5):2, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801005025.

Anonymous:2001:HGP

[1667] Anonymous. Hacking group plays Peekabooty with censorship. Network Security, 2001(5):2–3, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801005037.

Anonymous:2001:LVL

[1668] Anonymous. L0phtCrack v. 3.0 at large. Network Security, 2001 (5):3, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801005049.

Anonymous:2001:Be

[1669] Anonymous. In brief. Network Security, 2001(5):3, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801005074.

Anonymous:2001:VGA

[1670] Anonymous. Vigilantes to go after hackers. Network Security, 2001 (5):3, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801005050.

Anonymous:2001:MOW

[1671] Anonymous. Microsoft offers worm to best US customers. Network Security, 2001(5):3–4, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801005062.

Anonymous:2001:FSS

[1672] Anonymous. FBI sting sets dangerous precedent. Network Security, 2001(5):4, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801005086.

Anonymous:2001:MRO

[1673] Anonymous. Medical records open book in UK. Network Security, 2001(5):4, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801005098.

Gengler:2001:GE

[1674] Barbara Gengler. UN global e security.

Network Security, 2001(5):5, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801005104.

Anonymous:2001:DLA

[1675] Anonymous. DeCSS linking appeal update. Network Security, 2001(5):5, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801005116.

Madsen:2001:CCS

[1676] Wayne Madsen. Cybercrime convention steams ahead. Network Security, 2001(5):6, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801005128.

Pemble:2001:IBT

[1677] Matthew Pemble. The importance of being timely. Network Security, 2001(5):7–8, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580100513X.

Barber:2001:MXC

[1678] Richard Barber. Managing Xcommerce: The importance of a security-based architecture when preparing for e-commerce. Network Security, 2001(5):9–12, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801005141.

Franklin:2001:SOS

[1679] Iain Franklin. Securing the operating system: Does lesser-known mean less at risk? Is the OS an example of security though obscurity? Network Security, 2001(5):12–13, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801005153.

Anscombe:2001:OSE

[1680] Tony Anscombe. Online secure enterprises and the Linux effect. Network Security, 2001(5):14, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801005165.

Cohen:2001:NCG

[1681] Fred Cohen. The new cyber gang — a real threat profile. Network Security, 2001(5):15–17, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801005177.

Boni:2001:BBS

[1682] Bill Boni. Building bridges, standing guard. Network Security, 2001 (5):18–19, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801005189.

Anonymous:2001:SSSa

[1683] Anonymous. SOAPBOX: Spider squishing. Network Security, 2001 (5):20, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801005190.

Anonymous:2001:E

[1684] Anonymous. Events. Network Security, 2001(5):20, May 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801005207.

Anonymous:2001:ADM

[1685] Anonymous. Attrition discontinues mirroring service. Network Security, 2001(6):1, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801006018.

Anonymous:2001:NUU

[1686] Anonymous. NIPC is understaffed and under-performing. Network Security, 2001(6):2, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580100602X.

Anonymous:2001:MMA

[1687] Anonymous. Microsoft make antihacker film. Network Security, 2001 (6):2, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801006031.

Anonymous:2001:PKN

[1688] Anonymous. Protecting kids on the Net is political magnet. Network Security, 2001(6):2–3, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801006043.

Anonymous:2001:GPR

[1689] Anonymous. GlobalSign to provide root signatures online. Network Security, 2001(6):3, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801006055.

Anonymous:2001:Bf

[1690] Anonymous. In brief. Network Security, 2001(6):3, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801006080.

Anonymous:2001:FIP

[1691] Anonymous. Free international privacy guide out on Internet. Network Security, 2001(6):3, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801006067.

Anonymous:2001:ACC

[1692] Anonymous. Amazon cleared of customer privacy violations. Network Security, 2001(6):3, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801006079.

Anonymous:2001:WLC

[1693] Anonymous. Worm launched to crash Echelon. Network Security, 2001(6):4, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801006092.

Anonymous:2001:SSSb

[1694] Anonymous. Security spending set to soar. Network Security, 2001 (6):4, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801006110.

Anonymous:2001:FSC

[1695] Anonymous. 70% fail to secure confidential data. Network Security, 2001(6):4, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801006122.

Anonymous:2001:DIF

[1696] Anonymous. DERA issues free virus protection. Network Security, 2001(6):4, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801006109.

Gengler:2001:USI

[1697] Barbara Gengler. US set to improve Federal Websites, establish CIO. Network Security, 2001 (6):5, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801006134.

Sterlicchi:2001:CSS

[1698] John Sterlicchi. CERT to sell security alerts. Network Security, 2001 (6):5–6, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801006146.

Donnelly:2001:FMP

[1699] Allan Donnelly. Flawed mobile protocol delays release of standard. Network Security, 2001 (6):6, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801006158.

Anonymous:2001:VDT

[1700] Anonymous. Vulnerability database tops 1000 flaws. Network Security, 2001(6):6, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580100616X.

Anonymous:2001:LC

[1701] Anonymous. The lowdown on CISSP. Network Security, 2001 (6):7, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801006171.

Pemble:2001:TIS

[1702] Matthew Pemble. In theory, it shouldn’t be difficult in practice. Network Security, 2001(6):7–8, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801006183.

Forte:2001:VTH

[1703] Dario Forte. Vampire techniques and how to escape them. Network Security, 2001(6):9–11, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801006195.

Graham:2001:FCB

[1704] Rob Graham. The fifth column:

The biggest security threats to financial institutions may come from within — whether employees cooperate or not. Network Security, 2001 (6):12–15, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801006201.

Cohen:2001:WR

[1705] Fred Cohen. The wireless revolution. Network Security, 2001(6): 16–17, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801006213.

Boni:2001:CCC

[1706] Bill Boni. Corporations caught in the crossfire? Network Security, 2001 (6):18–19, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801006225.

Anonymous:2001:ECe

[1707] Anonymous. Events calender. Network Security, 2001(6):20, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801006249.

Anonymous:2001:PPV

[1708] Anonymous. A partly political virus problem. Network Security, 2001 (6):20, June 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801006237.

Anonymous:2001:SOD

[1709] Anonymous. SANS.org defaced, down for three days. Network Security, 2001(7):1, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801007012.

Anonymous:2001:ATS

[1710] Anonymous. Attrition team shares mirror’s insights. Network Security, 2001(7):2, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801007024.

Anonymous:2001:WS

[1711] Anonymous. White hat sent down.

Network Security, 2001(7):2–3, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801007036.

Anonymous:2001:NFV

[1712] Anonymous. New free vulnerability list launched. Network Security, 2001(7):3, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801007048.

Anonymous:2001:Bg

[1713] Anonymous. In brief. Network Security, 2001(7):3, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801007085.

Anonymous:2001:PPS

[1714] Anonymous. Privacy private say US firms. Network Security, 2001 (7):3, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580100705X.

Anonymous:2001:MLS

[1715] Anonymous. Mac losing squeakyclean image. Network Security, 2001(7):3, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801007061.

Anonymous:2001:EAU

[1716] Anonymous. EC asks US to keep its spying ECHR compliant. Network Security, 2001(7): 3–4, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801007073.

Anonymous:2001:ESS

[1717] Anonymous. EU set to secure Internet. Network Security, 2001 (7):4, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801007097.

Anonymous:2001:IFP

[1718] Anonymous. Infosec feeling the pinch. Network Security, 2001 (7):4, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801007103.

Anonymous:2001:EMB

[1719] Anonymous. Encryption market bolstered by hackers. Network Security, 2001(7):4, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801007115.

Gengler:2001:CVT

[1720] Barbara Gengler. CERT victim of three day denial-of-service attack. Network Security, 2001 (7):5, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801007127.

Wales:2001:TPM

[1721] Elspeth Wales. Tap-proof mobile phone is launched. Network Security, 2001(7):5–6, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801007139.

Gengler:2001:MRL

[1722] Barbara Gengler. Micron–Rambus lawsuit delayed. Network Security, 2001(7):6, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801007140.

Pemble:2001:SPA

[1723] Matthew Pemble. A sceptical pigeon amongst the crypto cats: Report on the Edinburgh Financial Cryptographic Engineering Conference, 2001. Network Security, 2001 (7):7–9, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801007152.

Barber:2001:SEP

[1724] Richard Barber. Social engineering: a people problem? Network Security, 2001(7):9–11, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801007164.

Webb:2001:WAWa

[1725] Alison Webb. Why audit a Web server?: Auditing Web servers: IIS 4.0 — Part 1. Network Security, 2001 (7):11–14, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801007176.

Madsen:2001:PON

[1726] Wayne Madsen. Pentagon orders NSA to monitor US citizens: DoD panel recommends Echelon be turned on US citizens, increased domestic role for NSA. Network Security, 2001 (7):14–16, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801007188.

Cohen:2001:MP

[1727] Fred Cohen. A matter of power. Network Security, 2001(7):16–18, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580100719X.

Boni:2001:PA

[1728] Bill Boni. The price of admission. Network Security, 2001(7): 18–19, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801007206.

Anonymous:2001:DBS

[1729] Anonymous. Don’t buy security products. Network Security, 2001 (7):20, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801007218.

Anonymous:2001:ECf

[1730] Anonymous. Events calendar. Network Security, 2001(7):20, July 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580100722X.

Anonymous:2001:BAC

[1731] Anonymous. Browser attacks “can wipe out almost anything”. Network Security, 2001(8):1, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801008017.

Anonymous:2001:EER

[1732] Anonymous. Encryption expert released on bail. Network Security, 2001(8):2, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801008029.

Anonymous:2001:CRH

[1733] Anonymous. Code Red — hype or horror? Network Security, 2001 (8):2–3, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801008030.

Anonymous:2001:SSM

[1734] Anonymous. SirCam stampedes mail servers. Network Security, 2001 (8):3, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801008042.

Anonymous:2001:Bh

[1735] Anonymous. In brief. Network Security, 2001(8):3, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801008066.

Anonymous:2001:PNL

[1736] Anonymous. PDF no longer quite so Peachy. Network Security, 2001 (8):3–4, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801008054.

Anonymous:2001:WRM

[1737] Anonymous. WEP: ready in 15 minutes. Network Security, 2001 (8):4, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801008078.

Anonymous:2001:AUC

[1738] Anonymous. AI used to catch pirates on Net. Network Security, 2001 (8):4, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580100808X.

Anonymous:2001:PCS

[1739] Anonymous. Port cloaking saves zombies. Network Security, 2001 (8):4, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801008091.

Gengler:2001:PVS

[1740] Barbara Gengler. PoizonB0x vandalizes security sites. Network Security, 2001(8):5, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801008108.

Sterlicchi:2001:CEN

[1741] John Sterlicchi. California energy network under attack. Network Security, 2001(8):5–6, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580100811X.

Johns:2001:WHH

[1742] Gerald Johns. Watching hackers in the honeynet. Network Security, 2001 (8):6, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801008121.

Pemble:2001:YAP

[1743] Matthew Pemble. Yet another paradigm!: Trust: a web, a hierarchy, or something else? Network Security, 2001(8):7–9, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801008133.

Forte:2001:WFW

[1744] Dario Forte. Web filtering: Where, how and why: Control of Internet use: some considerations about the implications of this type of control in the light of the Italian experience. Network Security, 2001(8):9–10, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801008145.

Schwartau:2001:NSA

[1745] Winn Schwartau. Network security: It’s about time: An offer for a metric. Network Security, 2001 (8):11–13, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801008157.

Madsen:2001:FCS

[1746] Wayne Madsen. FBI at centre stage of Code Red. Network Security, 2001 (8):14–15, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801008169.

Boni:2001:FCW

[1747] Bill Boni. Fish, CHIPS and worms. Network Security, 2001(8): 15–17, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801008170.

Cohen:2001:BC

[1748] Fred Cohen. Bootable CDs. Network Security, 2001(8):17–19, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801008182.

Anonymous:2001:TTP

[1749] Anonymous. It’s time to take porn seriously! Network Security, 2001 (8):20, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801008194.

Anonymous:2001:ECg

[1750] Anonymous. Events calendar. Network Security, 2001(8):20, August 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801008200.

Anonymous:2001:WDT

[1751] Anonymous. War declared on terrorism. Network Security, 2001(9): 1, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801009011.

Anonymous:2001:HFN

[1752] Anonymous. Hole found in NAI’s Gauntlet. Network Security, 2001 (9):2, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801009023.

Anonymous:2001:HCR

[1753] Anonymous. Hackers carry out revenge attacks on Islamic sites. Network Security, 2001(9):2, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801009035.

Anonymous:2001:MGE

[1754] Anonymous. Mafiaboy gets eight months. Network Security, 2001(9): 2–3, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801009047.

Anonymous:2001:LTW

[1755] Anonymous. Linux Trojan in the wild. Network Security, 2001(9): 3, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801009059.

Anonymous:2001:Bi

[1756] Anonymous. In brief. Network Security, 2001(9):3, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801009072.

Anonymous:2001:CBFb

[1757] Anonymous. Code Blue fights back against Red. Network Security, 2001 (9):3–4, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801009060.

Anonymous:2001:AKA

[1758] Anonymous. Anna Kournikova author stands trial. Network Security, 2001 (9):4, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801009084.

Anonymous:2001:NVH

[1759] Anonymous. Nimda virus hits. Network Security, 2001(9):4, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801009096.

Anonymous:2001:ESB

[1760] Anonymous. Email survives, buildings crash. Network Security, 2001 (9):4, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801009102.

Anonymous:2001:MRC

[1761] Anonymous. Military remote controls foil highjackers. Network Security, 2001 (9):4, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801009114.

Anonymous:2001:LUD

[1762] Anonymous. Legal update: DMCA happenings. Network Security, 2001 (9):5–6, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801009126.

Kirkby:2001:BRI

[1763] Andrea Kirkby. Book review: Internet Trust And Security: James Essinger, Addison Wesley, £31.99. Network Security, 2001(9):6, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801009138.

Farrell:2001:XPW

[1764] Stephen Farrell and Michael Zolotarev. XML and PKI — What’s the story? Network Security, 2001(9): 7–10, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580100914X.

Webb:2001:WAWb

[1765] Alison Webb. Why audit a Web server?: Auditing Web servers: IIS 4.0

— Part 2. Network Security, 2001(9): 11–14, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801009151.

Pemble:2001:CHC

[1766] Matthew Pemble. Crying ‘havoc’, crying ‘wolf’ or just howling at the Moon? Network Security, 2001(9): 14–16, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801009163.

Cohen:2001:BSB

[1767] Fred Cohen. The best security book ever written. Network Security, 2001 (9):17–18, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801009175.

Boni:2001:CGC

[1768] Bill Boni. Creating a global consensus against cybercrime. Network Security, 2001(9):18–19, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801009187.

Anonymous:2001:WIB

[1769] Anonymous. When ignorance is not bliss! Network Security, 2001(9): 20, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801009199.

Anonymous:2001:ECh

[1770] Anonymous. Events calendar. Network Security, 2001(9):20, September 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801009205.

Anonymous:2001:GS

[1771] Anonymous. Going for a song? Network Security, 2001(10):1, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801010017.

Anonymous:2001:NBC

[1772] Anonymous. Napster back in court. Network Security, 2001(10):

2, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801010029.

Anonymous:2001:CST

[1773] Anonymous. Cyber-security time and people poor, Science Committee told. Network Security, 2001 (10):2, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801010030.

Anonymous:2001:BAI

[1774] Anonymous. Bush abandons infrastructure order. Network Security, 2001 (10):2–3, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801010042.

Anonymous:2001:MRB

[1775] Anonymous. Microsoft to rate bug severity. Network Security, 2001 (10):3, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801010054.

Anonymous:2001:Bj

[1776] Anonymous. In brief. Network Security, 2001(10):3, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801010078.

Anonymous:2001:NRS

[1777] Anonymous. NSA release Secure Linux. Network Security, 2001(10): 3–4, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801010066.

Anonymous:2001:SIS

[1778] Anonymous. Security incidents set to double. Network Security, 2001 (10):4, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580101008X.

Anonymous:2001:NFM

[1779] Anonymous. Nimda fix is malware. Network Security, 2001(10): 4, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801010091.

Anonymous:2001:FGS

[1780] Anonymous. Finance to get secure payments. Network Security, 2001 (10):4, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801010108.

Anonymous:2001:CCA

[1781] Anonymous. Censoring crypto not the answer says Schneier. Network Security, 2001(10):4, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580101011X.

Anonymous:2001:NWS

[1782] Anonymous. Netherlands to watch strong crypto. Network Security, 2001 (10):4, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801010121.

Madsen:2001:BCO

[1783] Wayne Madsen. Bush creates Office of Homeland Security. Network Security, 2001(10):5–6, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801010133.

Gengler:2001:PRS

[1784] Barbara Gengler. PDAs at risk, says report. Network Security, 2001 (10):6, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801010145.

Palmer:2001:CCR

[1785] Chloe Palmer. Cracker court to rule on computer misuse. Network Security, 2001(10):6, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801010157.

Pemble:2001:LLW

[1786] Matthew Pemble. Licensed to … well, to what? And, by whom? Network Security, 2001(10):

7–9, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801010169.

Gray:2001:WWC

[1787] J. J. Gray. What we can learn about malicious mobile code. Network Security, 2001(10):9–10, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801010170.

Wright:2001:AES

[1788] Marie A. Wright. The Advanced Encryption Standard. Network Security, 2001(10):11–13, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801010182.

David:2001:IOI

[1789] Jon David. The ins and outs of intrusion detection. Network Security, 2001 (10):13–15, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801010194.

Cohen:2001:DWI

[1790] Fred Cohen. The DMCA: Why it is bad for InfoSec professionals. Network Security, 2001(10):16– 18, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801010200.

Boni:2001:TFW

[1791] Bill Boni. The towers fall and war begins. Network Security, 2001(10): 18–19, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801010212.

Anonymous:2001:ARL

[1792] Anonymous. Abstracts of recent literature. Network Security, 2001(10): 20, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801010224.

Anonymous:2001:ECi

[1793] Anonymous. Events calendar. Network Security, 2001(10):20, October 31, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801010236.

Anonymous:2001:DLR

[1794] Anonymous. DeCSS legal, rules Californian court. Network Security, 2001 (11):1–2, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801011011.

Anonymous:2001:SNS

[1795] Anonymous. Security news site defaced twice. Network Security, 2001 (11):2, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801011023.

Anonymous:2001:DBG

[1796] Anonymous. Dutch bank goes mobile. Network Security, 2001(11): 2, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801011035.

Anonymous:2001:ETM

[1797] Anonymous. Email trust is made cheaper. Network Security, 2001(11): 2–3, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801011047.

Anonymous:2001:IAF

[1798] Anonymous. Internet attracts fraud compliants. Network Security, 2001 (11):3, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801011059.

Anonymous:2001:Bk

[1799] Anonymous. In brief. Network Security, 2001(11):3, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801011072.

Anonymous:2001:VBA

[1800] Anonymous. Visa buys in AI tech to fight fraud. Network Security, 2001 (11):3–4, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801011060.

Anonymous:2001:NFS

[1801] Anonymous. Negligent firms support terrorism. Network Security, 2001 (11):4, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801011084.

Anonymous:2001:IIJ

[1802] Anonymous. IAAC and ICAF join up thinking. Network Security, 2001 (11):4, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801011096.

Anonymous:2001:WFL

[1803] Anonymous. What have firms learnt post 9/11? Network Security, 2001 (11):4, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801011102.

Gengler:2001:XNT

[1804] Barbara Gengler. XP a national threat? Network Security, 2001 (11):5, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801011114.

Gengler:2001:PPS

[1805] Barbara Gengler. Princeton poll shows 56% favour crypto regulations. Network Security, 2001(11): 5–6, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801011126.

Gengler:2001:CGI

[1806] Barbara Gengler. CATT — government and industry team up in US. Network Security, 2001(11): 6, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801011138.

Szerszen:2001:WNN

[1807] Dennis Szerszen. Wireless networking: Nirvana or nightmare? Network Security, 2001(11):7, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580101114X.

Pemble:2001:MR

[1808] Matthew Pemble. A Moment’s reflection. Network Security, 2001(11): 8–10, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801011151.

Hunter:2001:BBP

[1809] Philip Hunter. Barclays Bank prosecutes rather than covering up. Network Security, 2001(11):10–11, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801011163.

Lang:2001:WWY

[1810] Katherine Lang. Who’s watching you? Network Security, 2001(11): 11–13, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801011175.

Forte:2001:CWG

[1811] Dario Forte. Can worms be good? Network Security, 2001(11): 14–15, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801011187.

White:2001:FPW

[1812] Ben White. Fighting the porn war: The rise of email pornography in the workplace. Network Security, 2001(11): 16–17, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801011199.

Cohen:2001:SWU

[1813] Fred Cohen. Should we use deception as an InfoSec defence? Network Security, 2001(11):18–19, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801011205.

Anonymous:2001:DWH

[1814] Anonymous. Do women hack?: Haxor chicks are go. Network Security, 2001 (11):20, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801011217.

Anonymous:2001:ECj

[1815] Anonymous. Events calendar. Network Security, 2001(11):20, November 30, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485801011229.

Anonymous:2001:CTS

[1816] Anonymous. Cybercrime treaty signed, Budapest. Network Security, 2001 (12):1, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801012016.

Anonymous:2001:MLF

[1817] Anonymous. Magic Lantern fries crypto keys. Network Security, 2001 (12):2, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801012028.

Anonymous:2001:CSC

[1818] Anonymous. China shuts down 18000 cyber-cafes. Network Security, 2001 (12):2, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580101203X.

Anonymous:2001:GBO

[1819] Anonymous. Germany backs open source. Network Security, 2001(12):

2, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801012041.

Anonymous:2001:PCD

[1820] Anonymous. playboy.com data probed by hackers. Network Security, 2001 (12):2–3, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801012053.

Anonymous:2001:SDA

[1821] Anonymous. DoS set to double again in 2002. Network Security, 2001 (12):3, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801012065.

Anonymous:2001:TSD

[1822] Anonymous. ‘Tis the season to defraud. Network Security, 2001(12): 3, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801012077.

Anonymous:2001:Bl

[1823] Anonymous. In brief. Network Security, 2001(12):3, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801012090.

Anonymous:2001:MDW

[1824] Anonymous. MS doesn’t want to ‘cry wolf’. Network Security, 2001 (12):3–4, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801012089.

Anonymous:2001:IIB

[1825] Anonymous. IBM invests in biometrics. Network Security, 2001 (12):4, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801012107.

Anonymous:2001:CTY

[1826] Anonymous. CCTV through your PC or PDA. Network Security, 2001 (12):4, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801012119.

Gengler:2001:UGB

[1827] Barbara Gengler. US Government to build own Net. Network Security, 2001 (12):5, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801012120.

Gengler:2001:NPP

[1828] Barbara Gengler. Network printers pose security risk, CERT. Network Security, 2001(12):5, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801012132.

Sterlicchi:2001:IED

[1829] John Sterlicchi. IBM establishes data privacy groups. Network Security, 2001 (12):6, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801012144.

Madsen:2001:PPN

[1830] Wayne Madsen. Pentagon pushes national identification card for US. Network Security, 2001(12):6, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801012156.

Duke:2001:HM

[1831] David Duke. Hack of the month. Network Security, 2001(12):7, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801012168.

Pemble:2001:NDN

[1832] Matthew Pemble. Nightmare on the Dot-NET, Part II (Bill’s back). Network Security, 2001(12):8–9, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580101217X.

Packer:2001:PNN

[1833] Ryon Packer. Protecting the network: NIDS: the logical first step in intrusion detection deployment. Network Security, 2001(12):

10–11, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801012181.

Forte:2001:FAP

[1834] Dario Forte. Fragmentation attacks: Protection tools and techniques: Called “true preliminaries to denial-of-service”, IpFrags are a tough nut to crack for some firewalls and intrusion detection systems. Network Security, 2001(12): 12–13, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801012193.

Jones:2001:WHS

[1835] Deri Jones. Web hosting — the security risks. Network Security, 2001(12): 14–15, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580101220X.

Palmer:2001:RYP

[1836] Chloë Palmer. Review of the year, Part I: 2001: a game of two halves. Network Security, 2001(12): 15–17, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485801012211.

Boni:2001:CTC

[1837] Bill Boni. Cyber-terrorists and counter spies. Network Security, 2001(12): 17–18, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801012223.

Cohen:2001:WDW

[1838] Fred Cohen. The world doesn’t want to be fixed. Network Security, 2001 (12):19, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485801012235.

Anonymous:2001:EH

[1839] Anonymous. ‘Ethical’ hacking. Network Security, 2001(12):20, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801012247.

Anonymous:2001:ECk

[1840] Anonymous. Events calendar. Network Security, 2001(12):20, December 1, 2001. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485801012259.

Anonymous:2002:LYU

[1841] Anonymous. Lock up your UPnP says Microsoft. Network Security, 2002(1):1, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802001010.

Anonymous:2002:SHC

[1842] Anonymous. Sklyarov home for Christmas. Network Security, 2002 (1):2, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802001022.

Anonymous:2002:MLS

[1843] Anonymous. Magic Lantern snooping update. Network Security, 2002 (1):2, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802001034.

Anonymous:2002:BFH

[1844] Anonymous. Bid farewell to HTML in email. Network Security, 2002 (1):2–3, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802001046.

Anonymous:2002:BBNa

[1845] Anonymous. Bunni bates news site. Network Security, 2002(1): 3, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802001058.

Anonymous:2002:USC

[1846] Anonymous. US “somewhat concerned” about Web security. Network Security, 2002(1):3, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580200106X.

Anonymous:2002:Ba

[1847] Anonymous. In brief. Network Security, 2002(1):3, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802001083.

Anonymous:2002:IGT

[1848] Anonymous. IDSs generate too many false alarms. Network Security, 2002 (1):3–4, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802001071.

Anonymous:2002:BBW

[1849] Anonymous. Biometrics boom in wake of terror. Network Security, 2002 (1):4, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802001095.

Anonymous:2002:IMW

[1850] Anonymous. IT managers work too hard say Intel. Network Security, 2002(1):4, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802001101.

Anonymous:2002:BFS

[1851] Anonymous. Baltimore founder sues old firm. Network Security, 2002 (1):4, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802001113.

Anonymous:2002:SSF

[1852] Anonymous. Supermarket has security free e-warehouse. Network Security, 2002(1):4, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802001125.

Gengler:2002:UVV

[1853] Barbara Gengler. Unix vendors vulnerable to malicious attacks. Network Security, 2002(1):5, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802001137.

Gengler:2002:UAC

[1854] Barbara Gengler. US Army chooses Smartcard technology. Network Security, 2002(1):5–6, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802001149.

Palmer:2002:TMQ

[1855] Chloë Palmer. Toshiba makes quantum crypto breakthrough. Network Security, 2002(1):6, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802001150.

Duke:2002:HM

[1856] David Duke. Hack of the month. Network Security, 2002(1):7, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802001162.

Palmer:2002:RYP

[1857] Chloë Palmer. Review of the year, Part II: 2001: a game of two halves. Network Security, 2002(1): 8–10, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802001174.

Pemble:2002:HSI

[1858] Matthew Pemble. Be happy? Share information?? Network Security, 2002 (1):10–11, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802001186.

Pollino:2002:HSO

[1859] David Pollino. How to secure an Office wireless network. Network Security, 2002(1):12–13, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485802001198.

Anonymous:2002:RCS

[1860] Anonymous. Ringing the changes in the security cupboard. Network Security, 2002(1):14–16, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485802001204.

Cohen:2002:EIW

[1861] Fred Cohen. The end of the Internet as we know it. Network Security, 2002 (1):16–18, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802001216.

Boni:2002:GR

[1862] Bill Boni. The Golden Rule? Network Security, 2002(1):18–19, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802001228.

Anonymous:2002:POC

[1863] Anonymous. 2002: a privacy odyssey:

Customer: Hello, I’d like to upgrade my life. Can I have some privacy please?”. Network Security, 2002 (1):20, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580200123X.

Anonymous:2002:ECa

[1864] Anonymous. Events calendar. Network Security, 2002(1):20, January 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802001241.

Anonymous:2002:HH

[1865] Anonymous. Hackers on holiday? Network Security, 2002(2): 1–2, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802002015.

Anonymous:2002:GAW

[1866] Anonymous. Gates admits: we need to improve our security. Network Security, 2002(2):2–3, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485802002027.

Anonymous:2002:GWR

[1867] Anonymous. Gigger wants to reformat C drives. Network Security, 2002 (2):3, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802002039.

Anonymous:2002:Bb

[1868] Anonymous. In brief. Network Security, 2002(2):3, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485802002052.

Anonymous:2002:LIR

[1869] Anonymous. Linux infected by Remote Shell Trojan b. Network Security, 2002 (2):3–4, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802002040.

Anonymous:2002:SHA

[1870] Anonymous. Shock horror — another virus. Network Security, 2002 (2):4, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802002064.

Anonymous:2002:IMV

[1871] Anonymous. Instant messaging vulnerability. Network Security, 2002 (2):4, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802002076.

Anonymous:2002:OIE

[1872] Anonymous. Online intruders enter through Solaris hole. Network Security, 2002(2):4, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802002088.

Anonymous:2002:CSS

[1873] Anonymous. Cross-site scripting closed e-commerce site. Network Security, 2002(2):4, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580200209X.

Anonymous:2002:UUD

[1874] Anonymous. Users unable to download security patches. Network Security, 2002(2):4, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802002106.

Sellors:2002:VH

[1875] Nick Sellors. Viral hoaxes. Network Security, 2002(2):5, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802002118.

Anonymous:2002:PLU

[1876] Anonymous. Passwords linked to user ‘obsessions’. Network Security, 2002 (2):6–7, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580200212X.

White:2002:DSB

[1877] Paul White. Data security: the backup backdoor. Network Security, 2002(2): 8–9, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802002131.

Anonymous:2002:SSG

[1878] Anonymous. Security spending to grow in 2002. Network Security, 2002 (2):9, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802002143.

Forte:2002:NWE

[1879] Dario Forte. New worm experiments — a real danger? Network Security, 2002 (2):10–11, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802002155.

Kinsella:2002:SBW

[1880] Richard Kinsella. Securing e-business in a wireless environment. Network Security, 2002(2):12–13, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802002167.

Anonymous:2002:UAA

[1881] Anonymous. Update on Amsterdam airport biometric project. Network Security, 2002(2):13, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802002179.

Cohen:2002:SSI

[1882] Fred Cohen. Strategic security intelligence — embedded security. Network Security, 2002(2):14–15, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802002180.

Anonymous:2002:ASV

[1883] Anonymous. Another SMS vulnerability discovered. Network Security, 2002 (2):15, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802002192.

Anonymous:2002:DTE

[1884] Anonymous. Does trust exist? Network Security, 2002(2):16, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802002209.

Anonymous:2002:ECb

[1885] Anonymous. Events calendar. Network Security, 2002(2):16, February 28, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485802002210.

Anonymous:2002:SSA

[1886] Anonymous. SNMP security alarms ringing– is your enterprise awake and ready? Network Security, 2002 (3):1–2, March 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580200301X.

Anonymous:2002:VPA

[1887] Anonymous. Virus poses as antivirus program! Network Security, 2002(3):2, March 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802003021.

Anonymous:2002:WLU

[1888] Anonymous. Wireless LANs unprotected in London. Network Security, 2002(3):2, March 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802003033.

Anonymous:2002:EHC

[1889] Anonymous. Ethical hackers — can we trust them? Network Security, 2002(3):3, March 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802003045.

Anonymous:2002:DFP

[1890] Anonymous. Don’t forget to protect your custom applications. Network Security, 2002(3):4–5, March 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802003057.

Anonymous:2002:IHB

[1891] Anonymous. The importance of hardware-based cryptography for added security. Network Security, 2002(3):5, March 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802003069.

Anonymous:2002:ATM

[1892] Anonymous. It’s about time: a metric for InfoSecurity 1. Network Security, 2002(3):6–9, March 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802003070.

Anonymous:2002:HTH

[1893] Anonymous. Hidden threats — how much goes on inside your LAN? Network Security, 2002(3): 9–11, March 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802003082.

Anonymous:2002:BBNb

[1894] Anonymous. Behaviour blockers — a new dimension in protection. Network Security, 2002(3): 11–13, March 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802003094.

Anonymous:2002:SAN

[1895] Anonymous. Semantic attacks — a new wave of cyber-terrorism. Network Security, 2002(3):13–15, March 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485802003100.

Anonymous:2002:CIT

[1896] Anonymous. The cost of internal threats. Network Security, 2002(3): 15–16, March 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802003112.

Anonymous:2002:HGA

[1897] Anonymous. ‘How to get around your ISP’. Network Security, 2002 (3):17–18, March 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802003124.

Anonymous:2002:CDH

[1898] Anonymous. A cool day in Hades? Network Security, 2002(3):18–19, March 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1353485802003136.

Anonymous:2002:ECc

[1899] Anonymous. Events calendar. Network Security, 2002(3):20, March 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802003148.

Anonymous:2002:OHH

[1900] Anonymous. Oracle holes hide in the shade of SNMP limelight. Network Security, 2002(3):20, March 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580200315X.

Anonymous:2002:MLR

[1901] Anonymous. Microsoft: the latest round of vulnerabilities. Network Security, 2002(4):1, April 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802004014.

Anonymous:2002:TCD

[1902] Anonymous. ‘Trustworthy computing’ — delivers a new vulnerability scanner to trust! Network Security, 2002(4):2, April 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802004026.

Anonymous:2002:RDN

[1903] Anonymous. Remote dial-up numbers free for all to see. Network Security, 2002(4):2, April 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802004038.

Anonymous:2002:NWS

[1904] Anonymous. New Web services specifications from Microsoft, IBM and VeriSign. Network Security, 2002 (4):2–3, April 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580200404X.

Anonymous:2002:SEU

[1905] Anonymous. Security embarrassment for UK business. Network Security, 2002(4):3, April 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802004051.

Anonymous:2002:NFI

[1906] Anonymous. New fingerprint identification technology for wireless devices. Network Security, 2002 (4):3, April 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802004063.

Anonymous:2002:HCV

[1907] Anonymous. HP CEO voicemail leaked to press. Network Security, 2002(4):3, April 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802004075.

Duke:2002:IE

[1908] David Duke. Image is everything! Network Security, 2002 (4):4–5, April 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802004087.

Gengler:2002:WPN

[1909] Barbara Gengler. Wireless PANs new standard. Network Security, 2002(4):5, April 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802004099.

Forte:2002:ORD

[1910] Dario Forte. Onion routers: a dangerous response to traffic analysis? Network Security, 2002 (4):6–7, April 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802004105.

Laurie:2002:MVA

[1911] Bruce Laurie. Managed Vulnerability Assessment (MVA) — improve security by understanding your own vulnerabilities! Network Security, 2002(4):8–9, April 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802004117.

Hunter:2002:CPV

[1912] Philip Hunter. Canal Plus versus NDS case. Network Security, 2002 (4):9–11, April 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802004129.

Davis:2002:B

[1913] Dai Davis. Bluetooth. Network Security, 2002(4):11–12, April 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485802004130.

Cohen:2002:MWN

[1914] Fred Cohen. Misimpressions we need to extinguish. Network Security, 2002 (4):12–17, April 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802004142.

Boni:2002:DSC

[1915] Bill Boni. The dark side of e-commerce — cracking the code or passing the Buck? Network Security, 2002 (4):18–19, April 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802004154.

Anonymous:2002:ECd

[1916] Anonymous. Events calendar. Network Security, 2002(4):20, April 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485802004166.

Anonymous:2002:UAO

[1917] Anonymous. US Army opts for vulnerability assessment software to defend its global network terrain. Network Security, 2002(4): 20, April 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802004178.

Anonymous:2002:TSB

[1918] Anonymous. Telecom security breaches. Network Security, 2002 (5):1–2, May 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802050018.

Anonymous:2002:KB

[1919] Anonymous. Klez is back. Network Security, 2002(5):2, May 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580205002X.

Anonymous:2002:UFH

[1920] Anonymous. US FAA — hackers demonstrate what could happen …. Network Security, 2002 (5):2–3, May 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802050031.

Anonymous:2002:URG

[1921] Anonymous. US retail giant closes wireless registers as data beams at hackers. Network Security, 2002 (5):3, May 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802050043.

Anonymous:2002:LTA

[1922] Anonymous. Linux is target for attacks in Germany. Network Security, 2002(5):3, May 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802050055.

Anonymous:2002:DSS

[1923] Anonymous. DHCP servers subject to remote takeover. Network Security, 2002(5):3, May 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802050067.

Anonymous:2002:NNI

[1924] Anonymous. New network IDS solutions from Cisco. Network Security, 2002(5):3, May 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802050079.

Duke:2002:WDB

[1925] David Duke. What is the difference between Denial-of-Service (DoS) and Distributed-Denial-of-Service (DDoS)? Network Security, 2002(5):4, May 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802050080.

Anonymous:2002:WSW

[1926] Anonymous. Wireless security — what is out there? Network Security, 2002(5):5–6, May 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802050092.

Jervis:2002:WCA

[1927] Julie Jervis. Worldwide cyberattacks. Network Security, 2002 (5):6, May 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802050109.

Wilson:2002:BHL

[1928] Piers Wilson. Biometrics: Here’s looking at you …. Network Security, 2002(5):7–9, May 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802050110.

Ionescu:2002:SDP

[1929] Ionut Ionescu. Secondary data — the poor relative of business continuity. Network Security, 2002 (5):9–11, May 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802050122.

Forte:2002:VMO

[1930] Dario Forte. Vulnerability management: One problem, several potential approaches. Network Security, 2002 (5):11–13, May 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802050134.

Webb:2002:SWPa

[1931] Alison Webb. Security and Windows 2000: Part 1. Network Security, 2002 (5):13–16, May 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802050146.

Cohen:2002:TC

[1932] Fred Cohen. Terrorism and cyberspace. Network Security, 2002 (5):17–19, May 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802050158.

Anonymous:2002:ECe

[1933] Anonymous. Events calendar. Network Security, 2002(5):20, May 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580205016X.

Anonymous:2002:IMS

[1934] Anonymous. Instant Messaging — serious and secure? Network Security, 2002(5):20, May 31, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802050171.

Anonymous:2002:LS

[1935] Anonymous. Linux — to be or not to be secure. Network Security, 2002(6):1–2, June 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802060014.

Anonymous:2002:PAS

[1936] Anonymous. Privacy activists scheme to dodge government monitoring. Network Security, 2002(6):2, June 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802060026.

Anonymous:2002:CSB

[1937] Anonymous. Cisco switches become even more intelligent. Network Security, 2002(6):2–3, June 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802060038.

Anonymous:2002:RPD

[1938] Anonymous. Rogue protocol detection now possible. Network Security, 2002(6):3, June 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580206004X.

Anonymous:2002:NVC

[1939] Anonymous. New virus can target more than one operating system. Network Security, 2002 (6):3, June 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802060051.

Anonymous:2002:MNV

[1940] Anonymous. More new virus tactics — JPEGS. Network Security, 2002(6):3, June 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802060063.

Duke:2002:HW

[1941] David Duke. Hybrid worms. Network Security, 2002(6):4, June 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485802060075.

Webb:2002:SWPb

[1942] Alison Webb. Security and Windows 2000: Part 2. Network Security, 2002(6):5–8, June 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802060087.

Watson:2002:HSY

[1943] David Watson. How to secure your wireless network. Network Security, 2002(6):8–11, June 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802060099.

Davies:2002:WNR

[1944] Kerry Davies. What next? — A response to security for business leaders. Network Security, 2002 (6):12–13, June 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802060105.

Hone:2002:WME

[1945] Karin Höne and J. H. P. Eloff. What makes an effective information security policy? Network Security, 2002 (6):14–16, June 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802060117.

Boni:2002:CFM

[1946] Bill Boni. Card fraud — more serious than given credit for. Network Security, 2002(6):16–17, June 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802060129.

Cohen:2002:AVR

[1947] Fred Cohen. Academia’s vital role in information protection. Network Security, 2002(6):17–20, June 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802060130.

Anonymous:2002:ECf

[1948] Anonymous. Events calendar. Network Security, 2002(6):20, June 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802060142.

Anonymous:2002:HDN

[1949] Anonymous. Hackers demonstrate a new show of strength. Network Security, 2002(7):1–2, July 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802070010.

Anonymous:2002:CCW

[1950] Anonymous. Changing the computer, as we know it. Network Security, 2002(7):2, July 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802070022.

Anonymous:2002:MSH

[1951] Anonymous. Major satellite hacked in China. Network Security, 2002 (7):2–3, July 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802070034.

Anonymous:2002:AFA

[1952] Anonymous. Apache flaws — another premature leak. Network Security, 2002(7):3, July 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802070046.

Anonymous:2002:SAT

[1953] Anonymous. Symantec acquire three more security companies. Network Security, 2002(7):3, July 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485802070058.

Duke:2002:PPS

[1954] David Duke. Peer-to-peer sharing. Network Security, 2002(7):4, July 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580207006X.

Midian:2002:GMI

[1955] Paul Midian. Getting the most out of Intrusion Detection Systems. Network Security, 2002 (7):5–7, July 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802070071.

Asenjo:2002:AES

[1956] Juan C. Asenjo. The Advanced Encryption Standard — implementation and transition to a new cryptographic benchmark. Network Security, 2002(7):7–9, July 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802070083.

Abimbola:2002:SHP

[1957] Abiola Abimbola, David Gresty, and Qi Shi. SubSeven’s honey pot program. Network Security, 2002(7): 10–14, July 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802070095.

Mitrakas:2002:CCI

[1958] Andreas Mitrakas. Citizen centric identity management: Chip tricks? Network Security, 2002 (7):15–16, July 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802070101.

Cohen:2002:OSM

[1959] Fred Cohen. Is open source more or less secure? Network Security, 2002(7):17–19, July 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802070113.

Anonymous:2002:ECN

[1960] Anonymous. Event: CSI NetSec, 2002, San Francisco. Network Security, 2002(7):19, July 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802070125.

Anonymous:2002:ECg

[1961] Anonymous. Event calendar. Network Security, 2002(7):20, July 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485802070137.

Anonymous:2002:UGR

[1962] Anonymous. US Government release security benchmarks. Network Security, 2002(7):20, July 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802070149.

Anonymous:2002:LWH

[1963] Anonymous. At last, a wireless honeypot. Network Security, 2002 (8):1–2, August 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802080017.

Anonymous:2002:SJD

[1964] Anonymous. Security job descriptions in disarray. Network Security, 2002(8):2, August 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802080029.

Anonymous:2002:HCP

[1965] Anonymous. Hollywood crack down on piracy. Network Security, 2002 (8):2–3, August 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802080030.

Anonymous:2002:UHS

[1966] Anonymous. US Homeland Security Bill. Network Security, 2002 (8):3, August 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802080042.

Anonymous:2002:HPR

[1967] Anonymous. Hewlett Packard — rage after Tru64 disclosure. Network Security, 2002(8):3, August 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802080054.

Anonymous:2002:Bc

[1968] Anonymous. In brief. Network Security, 2002(8):3, August 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802080066.

Anonymous:2002:ZDV

[1969] Anonymous. Zero day vulnerabilities. Network Security, 2002 (8):4, August 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802080078.

Anonymous:2002:PPT

[1970] Anonymous. Perspectives on penetration testing — what’s the deal with Web security? Network Security, 2002 (8):5–8, August 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580208008X.

Anonymous:2002:AV

[1971] Anonymous. Adventures in virusland. Network Security, 2002(8): 8–10, August 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802080091.

Anonymous:2002:AGF

[1972] Anonymous. Alexander Galitsky, Founder & Co-Chairman, Trust-Works. Network Security, 2002(8): 10–12, August 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802080108.

Anonymous:2002:MSN

[1973] Anonymous. Mobile security — new needs on new devices. Network Security, 2002(8):13–14, August 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580208011X.

Anonymous:2002:QCR

[1974] Anonymous. Quantum cryptography revisited. Network Security, 2002(8): 14–16, August 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802080121.

Anonymous:2002:AQV

[1975] Anonymous. All quiet on the virus front? Network Security, 2002(8): 16–18, August 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802080133.

Anonymous:2002:YB

[1976] Anonymous. You’re in a Bind! Network Security, 2002(8):18–19, August 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802080145.

Anonymous:2002:ECh

[1977] Anonymous. Events calendar. Network Security, 2002(8):20, August 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802080157.

Anonymous:2002:NTW

[1978] Anonymous. New trends but the world keeps spinning. Network Security, 2002(8):20, August 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802080169.

Anonymous:2002:SLA

[1979] Anonymous. Security landscape analysis. Network Security, 2002(9):1– 2, September 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802090013.

Anonymous:2002:NSD

[1980] Anonymous. NASA secret data hacked. Network Security, 2002(9): 2–3, September 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802090025.

Anonymous:2002:SVM

[1981] Anonymous. Security vendors move into the peer-2-peer space. Network Security, 2002(9):3, September 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485802090037.

Anonymous:2002:WBH

[1982] Anonymous. Who has been hit by the big bad DDoS? Network Security, 2002 (9):3, September 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802090049.

Anonymous:2002:EMI

[1983] Anonymous. EU to monitor Internet communications. Network Security, 2002(9):3, September 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802090050.

Anonymous:2002:SAS

[1984] Anonymous. Spyware, adware, systemware and cookies. Network Security, 2002(9):4–5, September 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802090062.

Anonymous:2002:PFQ

[1985] Anonymous. Promise from the future — quantum cryptography. Network Security, 2002(9):6, September 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802090074.

Anonymous:2002:ASS

[1986] Anonymous. Application security — a serious pitfall. Network Security, 2002 (9):7, September 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802090086.

Anonymous:2002:FID

[1987] Anonymous. Firewalls, intrusion detection systems and vulnerability assessment: a superior conjunction? Network Security, 2002(9): 8–11, September 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802090098.

Anonymous:2002:MSS

[1988] Anonymous. Managed security services — a buyer’s guide. Network Security, 2002(9):12–15, September 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485802090104.

Anonymous:2002:VVV

[1989] Anonymous. Virtual vigilantes vs digital desperados. Network Security, 2002 (9):15–16, September 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802090116.

Anonymous:2002:PD

[1990] Anonymous. Protection by deception. Network Security, 2002(9):17– 19, September 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802090128.

Anonymous:2002:ECi

[1991] Anonymous. Events calendar. Network Security, 2002(9):20, September 1, 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580209013X.

Anonymous:2002:HP

[1992] Anonymous. The hacker preacher. Network Security, 2002(10):1–2, October 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802100018.

Anonymous:2002:WNS

[1993] Anonymous. The Whitehouse national strategy to secure cyberspace. Network Security, 2002(10):2–3, October 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580210002X.

Anonymous:2002:NCT

[1994] Anonymous. New cybersecurity taskforce from EC. Network Security, 2002(10):3–4, October 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802100031.

Anonymous:2002:SLB

[1995] Anonymous. A slap for Linux, a bug for Windows. Network Security, 2002(10):4, October 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802100043.

Anonymous:2002:NOA

[1996] Anonymous. New organization to alleviate vulnerability slip-ups. Network Security, 2002(10):4, October 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802100055.

Anonymous:2002:ONV

[1997] Anonymous. Out with the old and in with the new vulnerabilities. Network Security, 2002 (10):4, October 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802100067.

Anonymous:2002:SGS

[1998] Anonymous. SUN guru shines on Linux security and Web services. Network Security, 2002 (10):5, October 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802100079.

Anonymous:2002:PW

[1999] Anonymous. The political worm. Network Security, 2002(10):5, October 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485802100080.

Anonymous:2002:CNS

[2000] Anonymous. Compsec — network security and disaster survival. Network Security, 2002 (10):6, October 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802100092.

Anonymous:2002:GNS

[2001] Anonymous. GPRS network security: Welcome to life on the frontier! Network Security, 2002 (10):7, October 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802100109.

Anonymous:2002:HCI

[2002] Anonymous. HTML code injection and cross-site scripting. Network Security, 2002(10):8–12, October 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802100110.

Anonymous:2002:FCN

[2003] Anonymous. The future of computer and network forensics. Network Security, 2002(10):13–15, October 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802100122.

Anonymous:2002:WRS

[2004] Anonymous. When to review security — timing is Everything … . Network Security, 2002(10): 15–17, October 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802100134.

Anonymous:2002:RYF

[2005] Anonymous. Reworking your firewalls. Network Security, 2002(10): 18–19, October 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802100146.

Anonymous:2002:HJS

[2006] Anonymous. Honeypots — not just sticking to research. Network Security, 2002(10):20, October 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802100158.

Anonymous:2002:ECj

[2007] Anonymous. Events calendar. Network Security, 2002(10):20, October 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580210016X.

Anonymous:2002:SID

[2008] Anonymous. Security industry — disorganized. Network Security, 2002 (11):1–2, November 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802110014.

Madsen:2002:UPA

[2009] Wayne Madsen. USA Patriot Act II — more snooping powers. Network Security, 2002(11):2, November 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485802110026.

Anonymous:2002:DSI

[2010] Anonymous. DDoS strikes — industry thinks twice. Network Security, 2002(11):3, November 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485802110038.

Anonymous:2002:RW

[2011] Anonymous. Replacement for WEP.

Network Security, 2002(11):3, November 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S135348580211004X.

Anonymous:2002:HN

[2012] Anonymous. Hacking for news. Network Security, 2002(11):3, November 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802110051.

Duke:2002:DT

[2013] David Duke. Downloader Trojans. Network Security, 2002(11): 4–5, November 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802110063.

Hunter:2002:VLS

[2014] Philip Hunter. VOIP the latest security concern: DoS attack the greatest threat. Network Security, 2002 (11):5–7, November 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802110075.

Forte:2002:LAP

[2015] Dario Forte. Logging architecture — problems and solutions. Network Security, 2002(11):7–10, November 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802110087.

Midian:2002:PPT

[2016] Paul Midian. Perspectives on penetration testing — black box vs. white box. Network Security, 2002 (11):10–12, November 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802110099.

Schultz:2002:DID

[2017] Eugene Schultz. Demystifying intrusion detection: Sorting through the confusion, hyperbole and misconceptions 1. Network Security, 2002 (11):12–17, November 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802110105.

Cohen:2002:BTS

[2018] Fred Cohen. Breaking in to test security? Network Security, 2002 (11):17–19, November 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802110117.

Anonymous:2002:GSP

[2019] Anonymous. A gentle security player? Network Security, 2002 (11):19–20, November 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802110129.

Anonymous:2002:ECk

[2020] Anonymous. Events calendar. Network Security, 2002(11):20, November 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802110130.

Hunter:2002:NSA

[2021] Philip Hunter. Network security architecture without perimeters. Network Security, 2002(12): 1–2, December 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802120010.

Madsen:2002:IAU

[2022] Wayne Madsen. Industry’s attitude to US cybersecurity plan. Network Security, 2002(12):2–3, December 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802120022.

Anonymous:2002:SAP

[2023] Anonymous. System administrators patch too late. Network Security, 2002(12):3, December 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802120034.

Anonymous:2002:IRV

[2024] Anonymous. ISS release vulnerability disclosure rules. Network Security, 2002(12):3, December 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802120046.

Anonymous:2002:IDA

[2025] Anonymous. .info domain attacked by DdoS. Network Security, 2002 (12):3, December 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802120058.

Anonymous:2002:Bd

[2026] Anonymous. In brief. Network Security, 2002(12):3, December 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580212006X.

Duke:2002:PPT

[2027] David Duke. The peer-to-peer threat. Network Security, 2002 (12):4, December 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802120071.

Regan:2002:DSN

[2028] Kevin Regan. Designing a secure network for e-business and beyond. Network Security, 2002(12): 5–8, December 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802120083.

Anonymous:2002:GSA

[2029] Anonymous. Guidelines for securing Apache Web servers. Network Security, 2002(12):8–14, December 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802120095.

Wilson:2002:ISS

[2030] Piers Wilson. IT security — state of the nation. Network Security, 2002 (12):15–17, December 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802120101.

Cohen:2002:BM

[2031] Fred Cohen. Back up a minute. Network Security, 2002(12):17–19, December 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485802120113.

Anonymous:2002:ECl

[2032] Anonymous. Events calendar. Network Security, 2002(12):20, December 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485802120125.

Anonymous:2002:SEG

[2033] Anonymous. Social engineering — gift of the gab. Network Security, 2002(12):20, December 2002. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485802120137.

Anonymous:2003:MSC

[2034] Anonymous. Microsoft security culture shock. Network Security, 2003 (1):1–3, January 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803001016.

Anonymous:2003:SBZ

[2035] Anonymous. Security benchmarks zap 91% of vulnerabilities. Network Security, 2003(1):3, January 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803001028.

Anonymous:2003:TWA

[2036] Anonymous. Top 10 Web application holes exposed. Network Security, 2003(1):3, January 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580300103X.

Anonymous:2003:UGS

[2037] Anonymous. US Government slashes security strategy. Network Security, 2003(1):3–4, January 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803001041.

Anonymous:2003:Ba

[2038] Anonymous. In brief. Network Security, 2003(1):4, January 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803001053.

Anonymous:2003:MOC

[2039] Anonymous. Microsoft opens code to governments. Network Security, 2003(1):4, January 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803001065.

Madsen:2003:NCC

[2040] Wayne Madsen. NIAC calls for changes to US cybersecurity plan. Network Security, 2003(1):4, January 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803001077.

Duke:2003:P

[2041] David Duke. Passwords. Network Security, 2003(1):5, January 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803001089.

Hunter:2003:LIU

[2042] Philip Hunter. Lack of integration undermines IT security. Network Security, 2003(1):5–7, January 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803001090.

Regan:2003:WLS

[2043] Kevin Regan. Wireless LAN security: Things you should know about WLAN security. Network Security, 2003(1):7–9, January 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803001107.

Lord:2003:TTW

[2044] Steve Lord. Trouble at the Telco: When GSM goes bad. Network Security, 2003(1):10–12, January 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485803001119.

King:2003:MTB

[2045] David L. King. Moving towards a business continuity culture. Network Security, 2003(1):12–17, January 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485803001120.

Cohen:2003:SPA

[2046] Fred Cohen. Security programmers — not all the same. Network Security, 2003(1):18–19, January 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803001132.

Anonymous:2003:ECa

[2047] Anonymous. Events calendar. Network Security, 2003(1):20, January 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485803001144.

Anonymous:2003:QG

[2048] Anonymous. 2002 — quick glance. Network Security, 2003(1):20, January 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803001156.

Anonymous:2003:SWL

[2049] Anonymous. Slammer Worm — lessons learned. Network Security, 2003(2):1, February 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803002010.

Everett:2003:VHF

[2050] Catherine Everett. Vulnerabilities — hitting faster and harder. Network Security, 2003(2):2, February 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803002022.

Anonymous:2003:EHS

[2051] Anonymous. European honeypot server — hacked sticky. Network Security, 2003(2):2, February 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803002034.

Anonymous:2003:CAP

[2052] Anonymous. CERT accused of profiting from research. Network Security, 2003(2):3, February 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803002046.

Anonymous:2003:HGC

[2053] Anonymous. Hacker group — cracked. Network Security, 2003 (2):3, February 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803002058.

Madsen:2003:NUI

[2054] Wayne Madsen. New US Internet monitoring centre. Network Security, 2003(2):3, February 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580300206X.

Anonymous:2003:Bb

[2055] Anonymous. In brief. Network Security, 2003(2):3, February 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803002071.

Duke:2003:FAV

[2056] David Duke. Fooling anti-virus systems. Network Security, 2003 (2):4, February 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803002083.

Gallo:2003:SST

[2057] Vince Gallo. Secret steganography techniques revealed. Network Security, 2003(2):4–8, February 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803002095.

Midian:2003:PPT

[2058] Paul Midian. Perspectives on penetration testing — finding the right supplier. Network Security, 2003 (2):9–11, February 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803002101.

Venter:2003:AVS

[2059] H. S. Venter and J. H. P. Eloff.

Assessment of vulnerability scanners. Network Security, 2003(2): 11–16, February 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803002113.

Forte:2003:SRN

[2060] Dario Forte. Slammer — the return of the network nightmare. Network Security, 2003(2):17–18, February 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803002125.

Cohen:2003:SYI

[2061] Fred Cohen. Switching your infrastructure. Network Security, 2003 (2):19–20, February 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803002137.

Anonymous:2003:ECb

[2062] Anonymous. Events calendar. Network Security, 2003(2):20, February 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485803002149.

Anonymous:2003:GBS

[2063] Anonymous. Getting back to the source. Network Security, 2003 (3):1–3, March 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803003015.

Anonymous:2003:Bc

[2064] Anonymous. In brief. Network Security, 2003(3):2–3, March 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803003027.

Anonymous:2003:SBF

[2065] Anonymous. Security breach forces closure of company. Network Security, 2003(3):3, March 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803003039.

Anonymous:2003:LOG

[2066] Anonymous. Large online gamer paid ransom to hacker extortionists. Network Security, 2003(3):3, March 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485803003040.

Anonymous:2003:INF

[2067] Anonymous. IDS — the new firewall? Network Security, 2003(3):3, March 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803003052.

Anonymous:2003:UPP

[2068] Anonymous. USB — plug and pray?

Network Security, 2003(3):4, March 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803003076.

Anonymous:2003:WLD

[2069] Anonymous. Wireless LANs — do they fundamentally change the conventional security model? Network Security, 2003(3):4–7, March 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803003088.

Anonymous:2003:IFP

[2070] Anonymous. Internet freedom pendulum turns again … and do people really care? Network Security, 2003(3):8–9, March 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580300309X.

Anonymous:2003:MSS

[2071] Anonymous. The MS-SQL Slammer Worm. Network Security, 2003 (3):10–14, March 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803003106.

Anonymous:2003:NAW

[2072] Anonymous. Network armies in a world without secrets. Network Security, 2003(3):14–15, March 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803003118.

Anonymous:2003:DID

[2073] Anonymous. Distributed intrusion detection systems (DIDS) can make security more adaptive. Network Security, 2003(3):16–18, March 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580300312X.

Anonymous:2003:ND

[2074] Anonymous. Novelty detection. Network Security, 2003(3):18–19, March 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485803003131.

Anonymous:2003:ECc

[2075] Anonymous. Events calendar. Network Security, 2003(3):20, March 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485803003143.

Anonymous:2003:UI

[2076] Anonymous. US insight. Network Security, 2003(3):20, March 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803003155.

Anonymous:2003:BAC

[2077] Anonymous. Bugtraq attacked for censorship and delay. Network Security, 2003(4):1–2, April 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580300401X.

Anonymous:2003:ESC

[2078] Anonymous. Enterprise security concerns in year ahead. Network Security, 2003(4):2, April 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803004021.

Anonymous:2003:HAW

[2079] Anonymous. Hacktivists attack websites in war protests. Network Security, 2003(4):2, April 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803004033.

Anonymous:2003:Bd

[2080] Anonymous. In brief. Network Security, 2003(4):2–3, April 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803004057.

Anonymous:2003:TVC

[2081] Anonymous. Top 10 vulnerabilites from CERT. Network Security, 2003(4):3, April 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803004045.

Anonymous:2003:YCS

[2082] Anonymous. Is your current security SECURE?: John Jessop, Cryptic Software. Network Security, 2003(4):3, April 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803004069.

Anonymous:2003:RAP

[2083] Anonymous. Rogue access points — threat to enterprise security: Bruce Potter. Network Security, 2003(4):4–5, April 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803004070.

Anonymous:2003:EAM

[2084] Anonymous. Early alerts — making sense of security information overload: Kevin Hawkins, Senior Principal Consultant, Symantec Corp. Network Security, 2003(4):5–7, April 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803004082.

Anonymous:2003:NSA

[2085] Anonymous. Network security in action — can a worm teach us anything?: Paul King, security consultant, Cisco Systems UK and Ireland. Network Security, 2003 (4):8–11, April 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803004094.

Anonymous:2003:PPT

[2086] Anonymous. Perspectives on penetration testing — everybody knows this, surely?: Paul midian, insight consulting. Network Security, 2003 (4):12–14, April 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803004100.

Hunter:2003:GC

[2087] Philip Hunter. Grid computing. Network Security, 2003(4):15–16, April 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803004112.

Anonymous:2003:CCC

[2088] Anonymous. Covert channels: Covering ‘malicious’ traffic: Dario Forte. Network Security, 2003 (4):16–18, April 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803004124.

Anonymous:2003:DSF

[2089] Anonymous. Documenting security:

Fred Cohen. Network Security, 2003 (4):18–19, April 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803004136.

Anonymous:2003:ECd

[2090] Anonymous. Events calendar. Network Security, 2003(4):20, April 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485803004148.

Anonymous:2003:MF

[2091] Anonymous. Mistakes in fixes. Network Security, 2003(4):20, April 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580300415X.

Anonymous:2003:API

[2092] Anonymous. Analyst predicts infosecurity market 2003–6. Network Security, 2003(5):1–2, May 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803005014.

Anonymous:2003:ISD

[2093] Anonymous. IDS switches from detection to prevention. Network Security, 2003(5):2, May 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803005038.

Anonymous:2003:Be

[2094] Anonymous. In brief. Network Security, 2003(5):2–3, May 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803005026.

Anonymous:2003:WSS

[2095] Anonymous. Windows Server 2003 secure? Network Security, 2003(5): 3, May 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485803005051.

Anonymous:2003:NGM

[2096] Anonymous. New group merge physical and cyber security. Network Security, 2003(5):3, May 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803005063.

Anonymous:2003:CVM

[2097] Anonymous. Corporates are vulnerable to minor flaws. Network Security, 2003(5):3–4, May 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580300504X.

Potter:2003:BSO

[2098] Bruce Potter. Bluetooth — security optional. Network Security, 2003(5):4–5, May 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803005075.

Regan:2003:SVD

[2099] Kevin Regan. Secure VPN design considerations. Network Security, 2003(5):5–10, May 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803005087.

Anonymous:2003:ILS

[2100] Anonymous. An inside look at Security Operation Centres: Dario Forte outlines the role of Security Operation Centres, the gatherers of intelligence, in the future of security. Network Security, 2003 (5):11–12, May 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803005099.

Hunter:2003:DDS

[2101] Philip Hunter. Distributed Denial of Service (DDOS) mitigation tools. Network Security, 2003 (5):12–14, May 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803005105.

Wilson:2003:WSS

[2102] Piers Wilson. Web services security. Network Security, 2003 (5):14–16, May 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803005117.

Eiram:2003:SSS

[2103] Carsten H. Eiram. Snort, Sendmail and Samba — recent risks. Network Security, 2003(5):16–17, May 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803005129.

Cohen:2003:OSR

[2104] Fred Cohen. Operations security for the rest of us. Network Security, 2003(5):17–19, May 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803005130.

Anonymous:2003:ECe

[2105] Anonymous. Events calendar. Network Security, 2003(5):20, May 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803005142.

Anonymous:2003:BFV

[2106] Anonymous. Blackhats find vulnerabilities first. Network Security, 2003(5):20, May 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803005154.

Anonymous:2003:DPM

[2107] Anonymous. DNS — prone to more attacks. Network Security, 2003(6):1–2, June 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803006019.

Anonymous:2003:CSP

[2108] Anonymous. Cisco self-protecting networks. Network Security, 2003(6): 2, June 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485803006032.

Anonymous:2003:Bf

[2109] Anonymous. In brief. Network Security, 2003(6):2–3, June 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803006020.

Anonymous:2003:HFA

[2110] Anonymous. Harmful ‘free’ applications cost corporates lots. Network Security, 2003(6):2–3, June 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485803006044.

Anonymous:2003:SPL

[2111] Anonymous. Security products lack integration. Network Security, 2003(6):3, June 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803006056.

Anonymous:2003:ULV

[2112] Anonymous. University launches virus writers course. Network Security, 2003(6):3, June 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803006068.

Potter:2003:WAO

[2113] Bruce Potter. Wireless authentication options for up and down the Stack. Network Security, 2003 (6):4–5, June 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580300607X.

Kristensen:2003:RIV

[2114] Thomas Kristensen. Rash of IE vulnerabilities. Network Security, 2003(6):6, June 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803006081.

Mockapetris:2003:DYD

[2115] Paul V. Mockapetris. Defending your DNS & DHCP — best practices. Network Security, 2003 (6):7–10, June 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803006093.

Morrison:2003:DS

[2116] Paul Morrison. Database security. Network Security, 2003 (6):11–12, June 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580300610X.

Ollmann:2003:BPM

[2117] Gunter Ollmann. Best practice in managing HTTP-based client sessions. Network Security, 2003 (6):13–17, June 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803006111.

Hunter:2003:DDP

[2118] Philip Hunter. Defence in depth — protecting the queen. Network Security, 2003(6):17–18, June 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803006123.

Cohen:2003:BCS

[2119] Fred Cohen. Background checks: Series introduction. Network Security, 2003(6):18–20, June 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803006135.

Anonymous:2003:ECf

[2120] Anonymous. Events calendar. Network Security, 2003(6):20, June 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485803006147.

Anonymous:2003:WCS

[2121] Anonymous. Web crawling spies hunt corporate pirates. Network Security, 2003(7):1–2, July 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803007013.

Anonymous:2003:HCM

[2122] Anonymous. Hackers control 3 million servers. Network Security, 2003(7):1–2, July 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803007025.

Anonymous:2003:Bg

[2123] Anonymous. In brief. Network Security, 2003(7):2–3, July 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803007037.

Anonymous:2003:SPH

[2124] Anonymous. Symantec push honeypot as add-on to IDS. Network Security, 2003(7):2–3, July 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803007049.

Anonymous:2003:MRP

[2125] Anonymous. Microsoft RMS to protect “digital rights”. Network Security, 2003(7):3, July 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803007050.

Anonymous:2003:VA

[2126] Anonymous. Vulnerability analysis. Network Security, 2003(7):4, July 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803007062.

Anonymous:2003:KYW

[2127] Anonymous. Know your wireless Gear.

Network Security, 2003(7):5–6, July 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803007074.

Anonymous:2003:DSP

[2128] Anonymous. Database security (Part II). Network Security, 2003(7):6– 8, July 2003. CODEN NTSCF5.

ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485803007086.

Anonymous:2003:PSP

[2129] Anonymous. The perils of security patch management. Network Security, 2003(7):9–12, July 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803007098.

Anonymous:2003:PDH

[2130] Anonymous. Part 1: Deploying honeypots: Project background and implications. Network Security, 2003 (7):13–14, July 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803007104.

Anonymous:2003:PIH

[2131] Anonymous. Part II: Honeypots in detail: the variations. Network Security, 2003(7):14–15, July 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803007116.

Anonymous:2003:VAT

[2132] Anonymous. Vulnerability assessment tools. Network Security, 2003 (7):15–17, July 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803007128.

Anonymous:2003:QAI

[2133] Anonymous. Questions to ask in information security. Network Security, 2003(7):18–19, July 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580300713X.

Anonymous:2003:ECg

[2134] Anonymous. Events calendar. Network Security, 2003(7):20, July 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803007141.

Anonymous:2003:WCP

[2135] Anonymous. 17 ways to crack a password. Network Security, 2003 (8):1–2, August 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803008018.

Anonymous:2003:FPW

[2136] Anonymous. First predicted worm arrives to exploit MS flaw. Network Security, 2003(8):1–3, August 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580300802X.

Anonymous:2003:Bh

[2137] Anonymous. In brief. Network Security, 2003(8):2–3, August 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803008031.

Anonymous:2003:NSE

[2138] Anonymous. New senior Editor appointed to network security. Network Security, 2003(8):3, August 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803008043.

Potter:2003:WDD

[2139] Bruce Potter. Wireless device discovery. Network Security, 2003 (8):4–5, August 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803008055.

Hunter:2003:SIO

[2140] Philip Hunter. Security issues with offshore outsourcing: Offshore coding booming, but is it safe? Answer is a qualified yes, but only if you do your homework. Network Security, 2003(8):5–6, August 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803008067.

Schultz:2003:SWV

[2141] E. Eugene Schultz. The Sobig Worm variants: Letter after letter from A–E. Network Security, 2003 (8):7–10, August 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803008079.

Chuvakin:2003:HHV

[2142] Anton Chuvakin. “Honeynets: High value security data”: Analysis of real attacks launched at a honeypot. Network Security, 2003 (8):11–15, August 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803008080.

Wilson:2003:RS

[2143] Piers Wilson. Rogue servers. Network Security, 2003(8):16–18, August 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803008092.

Kristenson:2003:BPR

[2144] Thomas Kristenson. A big picture on recent vulnerabilities: Active AX, Web server compromises and insecure documents. Network Security, 2003(8):19, August 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803008109.

Anonymous:2003:ECh

[2145] Anonymous. Events calendar. Network Security, 2003(8):20, August 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803008110.

Anonymous:2003:UCS

[2146] Anonymous. US calls for Spyware law. Network Security, 2003 (8):20, August 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803008122.

Anonymous:2003:MSP

[2147] Anonymous. Man steals passwords with keystroke logger. Network Security, 2003(8):20, August 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803008134.

Anonymous:2003:BVW

[2148] Anonymous. Blaster variant writers busted. Network Security, 2003 (9):1, 3, September 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803009024.

Anonymous:2003:SFSa

[2149] Anonymous. Sobig family set to get bigger and badder. Network Security, 2003(9):1–2, September 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803009012.

Anonymous:2003:Bi

[2150] Anonymous. In brief. Network Security, 2003(9):2–3, September 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485803009036.

Potter:2003:NGW

[2151] Bruce Potter. Next generation wireless security tools. Network Security, 2003(9):4–5, September 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803009048.

Anonymous:2003:BCJ

[2152] Anonymous. Black Hat Conference:

Not just hackers: Report from the Information Security Festival. Conference reporter. Network Security, 2003 (9):5–6, September 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580300905X.

Anonymous:2003:UCD

[2153] Anonymous. Unravelling crypto developments: Dr Nicko van Someren, founder and CTO of nCipher, sorts out fact from fiction when it comes to quantum encryption. Network Security, 2003(9):7–8, September 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803009061.

King:2003:TSW

[2154] Stuart King. Threats and solutions to Web services security. Network Security, 2003(9):8–11, September 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803009073.

Shaw:2003:SAR

[2155] Geoff Shaw. Spyware & adware:

the risks facing businesses. Network Security, 2003(9):12–14, September 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803009085.

Wales:2003:IST

[2156] Elspeth Wales. Industry sinks teeth into spam. Network Security, 2003 (9):15–17, September 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803009097.

David:2003:IR

[2157] Jon David. Incident response. Network Security, 2003(9):17–19, September 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485803009103.

Kristensen:2003:BPBa

[2158] Thomas Kristensen. The big picture on big flaws: RPC DCOM vulnerability — what went wrong? Network Security, 2003(9):19–20, September 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485803009115.

Anonymous:2003:ECi

[2159] Anonymous. Events calendar. Network Security, 2003(9):20, September 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485803009127.

Anonymous:2003:MBB

[2160] Anonymous. Microsoft bites back as blame deepens. Network Security, 2003 (10):1, 3, October 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580301002X.

Anonymous:2003:WIA

[2161] Anonymous. Why it all went wrong with Blaster. Network Security, 2003 (10):1–2, October 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803010018.

Anonymous:2003:VAF

[2162] Anonymous. Virus authors faster to the kill. Network Security, 2003 (10):2, October 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803010043.

Anonymous:2003:Bj

[2163] Anonymous. In brief. Network Security, 2003(10):2–3, October 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803010031.

Schultz:2003:MWG

[2164] E. Eugene Schultz. The MS-Blaster worm: going from bad to worse. Network Security, 2003 (10):4–8, October 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803010055.

Anonymous:2003:SWA

[2165] Anonymous. Summer of worms: August’s child is a virus writer. Network Security, 2003(10):8–10, October 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803010067.

Potter:2003:WSP

[2166] Bruce Potter. Wireless security policies. Network Security, 2003 (10):10–12, October 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803010079.

Stewart:2003:SSA

[2167] Joe Stewart. Spam & Sobig: arm in arm. Network Security, 2003 (10):12–16, October 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803010080.

Sullivan:2003:VAW

[2168] Paul Sullivan. Vulnerability analysis — what is important and what is not. Network Security, 2003 (10):17–19, October 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803010092.

Kristensen:2003:BPBb

[2169] Thomas Kristensen. The big picture on big holes: SSH and Sendmail challenge Microsoft’s top spot for flaws. Network Security, 2003 (10):19–20, October 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803010109.

Anonymous:2003:ECj

[2170] Anonymous. Events calendar. Network Security, 2003(10):20, October 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803010110.

Anonymous:2003:VTM

[2171] Anonymous. Virus tracking moves back to basics. Network Security, 2003 (11):1, 3, November 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803011024.

Madsen:2003:CST

[2172] Wayne Madsen. Computerized surveillance a top priority for Pentagon. Network Security, 2003(11): 1–2, November 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803011012.

Anonymous:2003:Bk

[2173] Anonymous. In brief. Network Security, 2003(11):2–3, November 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803011036.

Anonymous:2003:EDR

[2174] Anonymous. Expert debunks regulation in cybersecurity. Network Security, 2003(11):2–3, November 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803011048.

Potter:2003:WBL

[2175] Bruce Potter. Wireless-based location tracking. Network Security, 2003 (11):4–5, November 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580301105X.

Lobo:2003:SLM

[2176] Colin Lobo. Security log management. Network Security, 2003 (11):6–9, November 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803011061.

Ranum:2003:DS

[2177] Marcus J. Ranum. The death of spam? Network Security, 2003(11): 10–13, November 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803011073.

Anonymous:2003:PB

[2178] Anonymous. Placing blame. Network Security, 2003(11):14–15, November 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485803011085.

Kristenson:2003:BPBa

[2179] Thomas Kristenson. The big picture on big holes. Network Security, 2003 (11):15–17, November 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803011097.

Dwan:2003:BMP

[2180] Berni Dwan. Body mnemonics in PDA security. Network Security, 2003 (11):17, November 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803011103.

Ollmann:2003:IPS

[2181] Gunter Ollmann. Intrusion prevention systems (IPS) destined to replace legacy routers. Network Security, 2003 (11):18–19, November 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803011115.

Anonymous:2003:ECk

[2182] Anonymous. Events calendar. Network Security, 2003(11):20, November 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803011127.

Anonymous:2003:IMF

[2183] Anonymous. ISPs — more forceful in security. Network Security, 2003 (11):20, November 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803011139.

Anonymous:2003:PAW

[2184] Anonymous. Port 80 apps — wider attack avenues. Network Security, 2003(12):1, December 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803000023.

Everett:2003:FDB

[2185] Catherine Everett. Full disclosure — the biggest flaw? Network Security, 2003(12):1, 2, December 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803000011.

Anonymous:2003:Bl

[2186] Anonymous. In brief. Network Security, 2003(12):2–3, December 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803000035.

Anonymous:2003:SFSb

[2187] Anonymous. Sobig.F spams its way to top of 2003 chart. Network Security, 2003(12):3, December 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803000047.

Potter:2003:TWS

[2188] Bruce Potter. Trends in wireless security — the big picture. Network Security, 2003(12):4–5, December 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803000059.

Forte:2003:PDE

[2189] Dario Forte. Principles of digital evidence collection. Network Security, 2003(12):6–7, December 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803000060.

Starnes:2003:TD

[2190] Richard Starnes. The Trojan defence. Network Security, 2003 (12):8, December 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803000072.

Schultz:2003:MHS

[2191] E. Eugene Schultz. The myth of Homeland security. Network Security, 2003(12):9, December 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803000084.

David:2003:UVB

[2192] Jon David. Unpatched vulnerabilities — the big issues. Network Security, 2003(12):10–14, December 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485803000096.

Dwan:2003:LTP

[2193] Berni Dwan. Laptop thieves: Price on their head$. Network Security, 2003(12):15, December 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485803000102.

Conner:2003:WS

[2194] Bill Conner. Worldwide security. Network Security, 2003(12): 16, December 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803000114.

Wilson:2003:TVB

[2195] Piers Wilson. “Top-down” versus “bottom-up” — different approaches to security. Network Security, 2003 (12):17–19, December 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803000126.

Kristenson:2003:BPBb

[2196] Thomas Kristenson. The big picture on Blg holes. Network Security, 2003 (12):19–20, December 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485803000138.

Anonymous:2003:ECl

[2197] Anonymous. Events calendar. Network Security, 2003(12):20, December 2003. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580300014X.

Everett:2004:SEE

[2198] Catherine Everett. Social engineering emails get more devious. Network Security, 2004(1):1, January 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000157.

Madsen:2004:PDS

[2199] Wayne Madsen. Passenger data sharing continues to be a problem between US & EU. Network Security, 2004(1):2, January 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000170.

Anonymous:2004:Ba

[2200] Anonymous. In brief. Network Security, 2004(1):2–3, January 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804000169.

Anonymous:2004:STD

[2201] Anonymous. Sober tops December virus chart. Network Security, 2004(1):3, January 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000182.

Anonymous:2004:TLA

[2202] Anonymous. TruSecure launches application-based vulnerability assessment. Network Security, 2004 (1):3, January 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000194.

Anonymous:2004:CPG

[2203] Anonymous. Check Point goes inside to stop threat. Network Security, 2004(1):3, January 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804000200.

Potter:2004:WBM

[2204] Bruce Potter. Warchalking and Bluejacking: Myth or reality. Network Security, 2004(1):4–5, January 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485804000212.

Schultz:2004:WSP

[2205] E. Eugene Schultz. Windows 2000 security: a postmortem analysis. Network Security, 2004(1): 6–9, January 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804000224.

Ford:2004:GGW

[2206] Richard Ford and Helayne Ray. Googling for gold: Web crawlers, hacking and defense explained. Network Security, 2004(1):10–13, January 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485804000236.

Levenhagen:2004:TCV

[2207] Roger Levenhagen. Trends, codes and virus attacks — 2003 year in review. Network Security, 2004 (1):13–15, January 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000248.

Anonymous:2004:C

[2208] Anonymous. It’s a crime. Network Security, 2004(1):15–17, January 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580400025X.

hunter:2004:IVS

[2209] Philip hunter. IP V6: Security issues. Network Security, 2004(1): 17–19, January 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000261.

Kristenson:2004:BPBa

[2210] Thomas Kristenson. The big picture on big holes. Network Security, 2004 (1):19–20, January 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000273.

Anonymous:2004:ECa

[2211] Anonymous. Events calendar. Network Security, 2004(1):20, January 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804000285.

Anonymous:2004:AVA

[2212] Anonymous. ASN.1 vulnerability — all we can do is patch & wait. Network Security, 2004 (2):1, February 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000297.

Anonymous:2004:MSCa

[2213] Anonymous. Microsoft source code exploited. Network Security, 2004 (2):1–2, February 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000303.

McKenna:2004:ICS

[2214] Brian McKenna. IBM and Cisco step up security convergence. Network Security, 2004(2):2–3, February 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804000315.

Anonymous:2004:Bb

[2215] Anonymous. In brief. Network Security, 2004(2):2–3, February 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804000327.

Potter:2004:SMD

[2216] Bruce Potter. Securing the mobile device. Network Security, 2004 (2):4–5, February 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000339.

Miller:2004:MWY

[2217] Andrew Miller. Is malware wrecking your computer? Network Security, 2004(2):5–6, February 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000340.

Furnell:2004:HUH

[2218] Steven Furnell and Stamatis Bolakis.

Helping us to help ourselves: Assessing administrators’ use of security analysis tools. Network Security, 2004 (2):7–12, February 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000352.

Anonymous:2004:MDY

[2219] Anonymous. MyDoom: Do you “get it” yet? Network Security, 2004 (2):13–15, February 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000364.

Kristensen:2004:BPBa

[2220] Thomas Kristensen. The big picture on big holes. Network Security, 2004(2):15, February 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000376.

Dwan:2004:RRa

[2221] Berni Dwan. Research review. Network Security, 2004(2):16–17, February 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485804000388.

Hunter:2004:CVP

[2222] Philip Hunter. Combating video piracy. Network Security, 2004(2): 18–19, February 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580400039X.

Anonymous:2004:ATW

[2223] Anonymous. Not another teenager — where are the grown-up virus authors? Network Security, 2004 (2):20, February 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804000406.

Anonymous:2004:ECb

[2224] Anonymous. Events calendar. Network Security, 2004(2):20, February 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485804000418.

Anonymous:2004:NSN

[2225] Anonymous. Network security needs numbers say CSOs. Network Security, 2004(3):1, March 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580400042X.

Madsen:2004:HLC

[2226] Wayne Madsen. Highest level clearances planned for all US IT security pros. Network Security, 2004 (3):1–2, March 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000431.

Anonymous:2004:Bc

[2227] Anonymous. In brief. Network Security, 2004(3):2–3, March 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000443.

Anonymous:2004:CBB

[2228] Anonymous. Confidentiality breaches:

The biggest cost. Network Security, 2004(3):2–3, March 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000455.

Anonymous:2004:CFS

[2229] Anonymous. Cisco fattens security product line. Network Security, 2004(3):3, March 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000467.

Anonymous:2004:EFE

[2230] Anonymous. EU forks out EUR 50 million to clean Internet. Network Security, 2004(3):3, March 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000479.

Potter:2004:BV

[2231] Bruce Potter. Bluetooth “vulnerabilities”. Network Security, 2004 (3):4–5, March 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804000480.

Anonymous:2004:FEH

[2232] Anonymous. February: Eventful horizon. Network Security, 2004 (3):5–7, March 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804000492.

Ollmann:2004:SAT

[2233] Günter Ollmann. Securing against the ‘threat’ of instant. Network Security, 2004(3):8–11, March 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804000509.

Hunter:2004:CCI

[2234] Philip Hunter. Centrally controlled intrusions — the future threat. Network Security, 2004 (3):12–13, March 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804000510.

Forte:2004:MCT

[2235] Dario Forte. MyDoom & Co. is there really a problem? Network Security, 2004(3):14, March 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000522.

Hilley:2004:FCL

[2236] Sarah Hilley. The final countdown:

3, 2, 1 … zero. Network Security, 2004(3):15–16, March 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000534.

Dwan:2004:RRb

[2237] Berni Dwan. Research review. Network Security, 2004(3):17–18, March 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804000546.

Kristenson:2004:BPBb

[2238] Thomas Kristenson. The big picture on big holes. Network Security, 2004(3):19, March 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000558.

Hichert:2004:OFS

[2239] Jan Hichert. Opinion: The financial sector’s move to open source security products. Network Security, 2004(3):19–20, March 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580400056X.

Anonymous:2004:ECc

[2240] Anonymous. Events calendar. Network Security, 2004(3):20, March 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485804000571.

Madsen:2004:FFD

[2241] Wayne Madsen. Former FBI Director says encryption fuels terrorists. Network Security, 2004 (4):1, 3, April 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804000595.

Anonymous:2004:MSF

[2242] Anonymous. Microsoft squeezes 20 flaws into 4 patches. Network Security, 2004(4):1–2, April 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804000583.

Anonymous:2004:NCF

[2243] Anonymous. NetScreen combines firewall and intrusion protection. Network Security, 2004(4):2, April 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485804000613.

Anonymous:2004:Bd

[2244] Anonymous. In brief. Network Security, 2004(4):2–3, April 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804000601.

Anonymous:2004:UUH

[2245] Anonymous. US universities hit in planned attack. Network Security, 2004(4):3, April 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000625.

Potter:2004:WID

[2246] Bruce Potter. Wireless intrusion detection. Network Security, 2004(4):4–5, April 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000637.

Heasman:2004:MNP

[2247] John Heasman. Migrating to the .NET platform: an introduction. Network Security, 2004 (4):6–7, April 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000649.

Edelson:2004:SNA

[2248] Eve Edelson. Security in network attached storage (NAS) for workgroups. Network Security, 2004 (4):8–12, April 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000650.

Claassen:2004:SH

[2249] Pieter Claassen. The state of the hack. Network Security, 2004 (4):12–13, April 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804000662.

Forte:2004:ITS

[2250] Dario Forte. The importance of text searches in digital forensics. Network Security, 2004(4):13–15, April 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485804000674.

Morgan:2004:NSC

[2251] David Morgan. Network security and custom Web applications. Network Security, 2004(4):15–17, April 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485804000686.

Wood:2004:RIS

[2252] Peter Wood. Real IT security on a limited budget? Network Security, 2004(4):18, April 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804000698.

McKenna:2004:IRA

[2253] Brian McKenna. Interview — remote access at the BBC. Network Security, 2004(4):19–20, April 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804000704.

Anonymous:2004:ECd

[2254] Anonymous. Events calendar. Network Security, 2004(4):20, April 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804000716.

Anonymous:2004:MBH

[2255] Anonymous. Microsoft bounty hunt leads to German youths. Network Security, 2004(5):1, May 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580400073X.

Anonymous:2004:OBS

[2256] Anonymous. Online betting site beats blackmail DDoS. Network Security, 2004(5):1, 3, May 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000728.

Madsen:2004:HCS

[2257] Wayne Madsen. House considers security and privacy proposals. Network Security, 2004(5):2, May 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000753.

Anonymous:2004:Be

[2258] Anonymous. In brief. Network Security, 2004(5):2–3, May 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000741.

Everett:2004:ESB

[2259] Cath Everett. Eurovision secures borders. Network Security, 2004(5):3, May 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000765.

Potter:2004:GS

[2260] Bruce Potter. GSM security. Network Security, 2004(5):4–5, May 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804000777.

Wilson:2004:TCF

[2261] Piers Wilson. Technical challenges faced when companies merge …. Network Security, 2004(5):5–7, May 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804000789.

Morgan:2004:SHA

[2262] David Morgan. Securing high availability Web applications. Network Security, 2004(5):8–11, May 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000790.

Dwan:2004:OSV

[2263] Berni Dwan. Open source vs closed. Network Security, 2004 (5):11–13, May 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000807.

Furnell:2004:GCP

[2264] Steven M. Furnell. Getting caught in the phishing net. Network Security, 2004(5):14–18, May 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000819.

Balle:2004:BPB

[2265] Jakob Balle. The big picture on big holes. Network Security, 2004 (5):18–20, May 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000820.

Anonymous:2004:ECe

[2266] Anonymous. Events calendar. Network Security, 2004(5):20, May 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804000832.

Anonymous:2004:HEN

[2267] Anonymous. HP exploits new bugs to fix its systems. Network Security, 2004(6):1, June 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000856.

Anonymous:2004:UGA

[2268] Anonymous. US Government agencies struggle at patch management. Network Security, 2004 (6):1–2, June 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000844.

Anonymous:2004:Bf

[2269] Anonymous. In brief. Network Security, 2004(6):2–3, June 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000868.

McKenna:2004:BFA

[2270] Brian McKenna. Bug-fixed applications still insecure. Network Security, 2004(6):2–3, June 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580400087X.

McKenna:2004:CTM

[2271] Brian McKenna. Cisco and Trend Micro tighten collaboration around Sasser-like attacks. Network Security, 2004(6):3, June 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000881.

Potter:2004:FWS

[2272] Bruce Potter. Fixing wireless security. Network Security, 2004 (6):4–5, June 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000893.

Schultz:2004:RWS

[2273] E. Eugene Schultz. RPC in Windows systems: what you don’t know could hurt you. Network Security, 2004(6):5–8, June 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580400090X.

Paget:2004:HCA

[2274] Chris Paget. Hack chaining: Attacks that become possible when a server is compromised. Network Security, 2004(6):8–11, June 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000911.

Furnell:2004:WVR

[2275] Steven Furnell. When vulnerability reports can work against us. Network Security, 2004(6):11–15, June 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804000923.

Hunter:2004:ISN

[2276] Philip Hunter. Integrated security and network management remain elusive. Network Security, 2004(6):15–16, June 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000935.

Gordon:2004:SCA

[2277] Sarah Gordon. A short course in antivirus software testing: seven simple rules for evaluating tests. Network Security, 2004(6):17–18, June 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485804000947.

Balle:2004:BAF

[2278] Jakob Balle. Bruised Apple to fight back? Network Security, 2004(6):19–20, June 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804000959.

Anonymous:2004:ECf

[2279] Anonymous. Events calendar. Network Security, 2004(6):20, June 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485804000960.

McKenna:2004:UMC

[2280] Brian McKenna. UK MPs call for Computer Misuse Act upgrade: CMA in dire need of reform, says report. Network Security, 2004 (7):1, 3, July 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804000984.

Everett:2004:TTB

[2281] Cath Everett. Testing times for beleagured Internet Explorer. Network Security, 2004(7):1–2, July 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485804000972.

Anonymous:2004:Bg

[2282] Anonymous. In brief. Network Security, 2004(7):2–3, July 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804000996.

Anonymous:2004:MUU

[2283] Anonymous. Microsoft unleash 7 updates. Network Security, 2004(7):3, July 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580400100X.

Potter:2004:SGT

[2284] Bruce Potter. 802.16 security: getting there? Network Security, 2004(7):4–5, July 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001011.

Dunn:2004:AUR

[2285] Kevin Dunn. Automatic update risks: can patching let a hacker in? Network Security, 2004(7):5–8, July 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804001023.

Miller:2004:PSC

[2286] Andrew Miller. PDA security concerns. Network Security, 2004 (7):8–10, July 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001035.

Gorling:2004:PWY

[2287] Stefan Gorling. Parasites: what you can’t see can hurt you. Network Security, 2004(7):11–13, July 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804001047.

Dwan:2004:MPV

[2288] Berni Dwan. The mobile phone virus. Network Security, 2004 (7):14–15, July 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001059.

Papadaki:2004:IIW

[2289] Maria Papadaki and Steven Furnell. IDS or IPS: what is best? Network Security, 2004(7):15–19, July 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804001060.

Kristensen:2004:BPBb

[2290] Thomas Kristensen. The big picture on big holes. Network Security, 2004(7):19–20, July 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001072.

Anonymous:2004:ECg

[2291] Anonymous. Events calendar. Network Security, 2004(7):20, July 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804001084.

Anonymous:2004:SBS

[2292] Anonymous. Stealth bomber screen to block wireless hackers: BAE makes WLANs disappear. Network Security, 2004(8):1–2, August 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001096.

Anonymous:2004:PEV

[2293] Anonymous. Pfizer to exterminate Viagra spammers. Network Security, 2004(8):1–3, August 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001102.

OHalloran:2004:FSW

[2294] Joe O’Halloran. Fashion sharpens wireless risk. Network Security, 2004(8):2, August 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001126.

Anonymous:2004:Bh

[2295] Anonymous. In brief. Network Security, 2004(8):2–3, August 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804001114.

Anonymous:2004:MGG

[2296] Anonymous. MyDoom gags Google.

Network Security, 2004(8):3, August 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804001138.

Potter:2004:WUB

[2297] Bruce Potter. Wireless updates from Black Hat and DefCon. Network Security, 2004(8):4–5, August 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580400114X.

James:2004:MTS

[2298] Gareth James. Malicious threats to Smartphones. Network Security, 2004(8):5–7, August 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804001151.

Hunter:2004:LSS

[2299] Philip Hunter. Linux security: separating myth from reality. Network Security, 2004(8):8–9, August 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485804001163.

Chuvakin:2004:LID

[2300] Anton Chuvakin. Linux intrusion discovery: when security fails. Network Security, 2004(8):10–12, August 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485804001175.

Hilley:2004:TCP

[2301] Sarah Hilley. Trusted computing — path to security or road to servitude? Network Security, 2004 (8):12–15, August 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001187.

Lobo:2004:WMD

[2302] Colin Lobo. Where’s my data gone!? Network Security, 2004 (8):15–17, August 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001199.

Anonymous:2004:GGW

[2303] Anonymous. Google’s Gmail — will ads, privacy concerns or security fears spoil the free lunch? Network Security, 2004(8):18–19, August 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001205.

Sandblad:2004:BPB

[2304] Andreas Sandblad. The big picture on big holes. Network Security, 2004 (8):19–20, August 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001217.

Anonymous:2004:ECh

[2305] Anonymous. Events calendar. Network Security, 2004(8):20, August 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804001229.

Anonymous:2004:UGP

[2306] Anonymous. UK Government to push greater public awareness of Internet trust. Network Security, 2004 (9):1, 3, September 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001242.

Anonymous:2004:RPD

[2307] Anonymous. Research pinpoints 10 deadly sins. Network Security, 2004 (9):1–2, September 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001230.

Anonymous:2004:GPA

[2308] Anonymous. German police arrest eBay hi-jacker. Network Security, 2004(9):2, September 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001266.

Anonymous:2004:SAC

[2309] Anonymous. Sasser author charged. Network Security, 2004(9):2, September 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485804001278.

Anonymous:2004:Bi

[2310] Anonymous. In brief. Network Security, 2004(9):2–3, September 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804001254.

Anonymous:2004:VM

[2311] Anonymous. VPN for the masses. Network Security, 2004(9):3, September 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580400128X.

Potter:2004:SE

[2312] Bruce Potter. Security to the edge. Network Security, 2004(9): 4–5, September 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001291.

Wilson:2004:TSY

[2313] Piers Wilson. Tips to stop your users from being phished. Network Security, 2004(9):5–9, September 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001308.

Kemp:2004:GM

[2314] Michael Kemp. Ghosts in the machine. Network Security, 2004(9): 9–12, September 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580400131X.

David:2004:NTB

[2315] Jon David. New threats bring new treatments. Network Security, 2004 (9):12–15, September 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001321.

Dwan:2004:HAS

[2316] Berni Dwan. Honeynets aim to sting blackhats. Network Security, 2004 (9):15–18, September 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804001333.

Eiram:2004:BPB

[2317] Carsten Eiram. The big picture on big holes. Network Security, 2004 (9):18–20, September 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804001345.

Anonymous:2004:ECi

[2318] Anonymous. Events calendar. Network Security, 2004(9):20, September 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485804001357.

Anonymous:2004:TFL

[2319] Anonymous. Top 20 fear list unleashed. Network Security, 2004 (10):1–2, October 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804001369.

Anonymous:2004:IWB

[2320] Anonymous. Infosec wages buoyant. Network Security, 2004(10):2, October 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485804001370.

McKenna:2004:AIV

[2321] Brian McKenna. Accelerated IP v6 shift caught at Check Point. Network Security, 2004(10):3, October 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804001382.

McKenna:2004:CTO

[2322] Brian McKenna. Combination tools to oust point products. Network Security, 2004(10):3, October 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001394.

Anonymous:2004:Bj

[2323] Anonymous. In brief. Network Security, 2004(10):3, October 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804001400.

Potter:2004:SNS

[2324] Bruce Potter. Software & network security. Network Security, 2004 (10):4–5, October 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001412.

Anonymous:2004:UNS

[2325] Anonymous. US NSF to spend $30m to defeat malware. Network Security, 2004(10):5–6, October 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001424.

Schultz:2004:USD

[2326] Gene Schultz. Using ssh: Do security risks outweigh the benefits? Network Security, 2004 (10):7–10, October 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804001436.

Furnell:2004:MCA

[2327] Steven Furnell and Jeremy Ward. Malware comes of age: The arrival of the true computer parasite. Network Security, 2004(10): 11–15, October 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804001448.

Kemp:2004:BIR

[2328] Michael Kemp. Basic incident response for system administrators. Network Security, 2004(10): 16–18, October 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580400145X.

Kristensen:2004:LP

[2329] Thomas Kristensen. Lethal pictures. Network Security, 2004(10): 19–20, October 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804001461.

McKenna:2004:LAV

[2330] Brian McKenna. Lumeta adds visuals to network discovery tool. Network Security, 2004(10):20, October 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804001473.

Anonymous:2004:ECj

[2331] Anonymous. Events calendar. Network Security, 2004(10):20, October 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804001485.

Anonymous:2004:SFL

[2332] Anonymous. Security firms leap into Cahoot debate. Network Security, 2004 (11):1–2, November 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001497.

McKenna:2004:AED

[2333] Brian McKenna. American Express deploys appliance for secure file transfer. Network Security, 2004 (11):2, November 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001503.

Anonymous:2004:Bk

[2334] Anonymous. In brief. Network Security, 2004(11):3, November 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804001527.

Anonymous:2004:ISL

[2335] Anonymous. IBM’s support for the Liberty Alliance brings standard convergence for federated identity a step closer. Network Security, 2004(11): 3, 20, November 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001515.

Potter:2004:VIA

[2336] Bruce Potter. The value of Internet alerts. Network Security, 2004 (11):4, November 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001539.

Young:2004:HRV

[2337] Dale Young. Human resources have a vital role to play within employee identity and access management. Network Security, 2004 (11):5–7, November 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001540.

Morgan:2004:DIH

[2338] David Morgan. Does IT hang together? Network Security, 2004 (11):8–10, November 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001552.

Vidalis:2004:ACT

[2339] Stilianos Vidalis, Andy Jones, and Andrew Blyth. Assessing cyberthreats in the information environment. Network Security, 2004(11): 10–16, November 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001564.

Kristensen:2004:MLW

[2340] Thomas Kristensen. Microsoft leaves Win2000, XPSP1 users in lurch. Network Security, 2004(11): 16–17, November 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001576.

Shechtman:2004:DC

[2341] Katrin Shechtman. Divide and conquer. Network Security, 2004(11): 18–19, November 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001588.

Anonymous:2004:USA

[2342] Anonymous. USB sticks to be audited — no escape. Network Security, 2004(11):20, November 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580400159X.

Anonymous:2004:GRD

[2343] Anonymous. Government regulations drive job demand. Network Security, 2004(11):20, November 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804001606.

Anonymous:2004:MSCb

[2344] Anonymous. MS source code fence busted. Network Security, 2004 (11):20, November 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804001618.

Anonymous:2004:ECk

[2345] Anonymous. Events calendar. Network Security, 2004(11):20, November 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580400162X.

Anonymous:2004:VTS

[2346] Anonymous. Viruses in 2004 — technology stifles tenacious variants. Network Security, 2004(12): 1–2, December 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804001631.

Anonymous:2004:SAF

[2347] Anonymous. Skewed audit files scupper IDS. Network Security, 2004 (12):3, December 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485804001643.

Anonymous:2004:CSF

[2348] Anonymous. Company spotlights five flaws that will be worms. Network Security, 2004(12):3, December 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804001655.

Anonymous:2004:Bl

[2349] Anonymous. In brief. Network Security, 2004(12):3–4, December 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804001667.

Anonymous:2004:YR

[2350] Anonymous. Year in review. Network Security, 2004(12):4, December 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804001679.

Potter:2004:BBW

[2351] Bruce Potter. Big Boss is watching you — but should he? Network Security, 2004(12):5–6, December 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001680.

Fumy:2004:ISS

[2352] Walter Fumy. IT security standardisation. Network Security, 2004 (12):6–11, December 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001692.

Morgan:2004:DE

[2353] David Morgan. Deploying extranets? Network Security, 2004 (12):12–14, December 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001709.

Hunter:2004:TSP

[2354] Philip Hunter. Testing security products. Network Security, 2004(12): 15–16, December 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001710.

Knight:2004:MWD

[2355] William Knight. Mobile working drives switch to federated access rights. Network Security, 2004(12): 16–19, December 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001722.

Kristensen:2004:WSV

[2356] Thomas Kristensen. Winamp still vulnerable. Network Security, 2004 (12):19–20, December 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485804001734.

McKenna:2004:NYM

[2357] Brian McKenna. New York mayor, Giuliani, says “be prepared”. Network Security, 2004(12):20, December 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804001746.

Anonymous:2004:ECl

[2358] Anonymous. Events calendar. Network Security, 2004(12):20, December 2004. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485804001758.

Dudley-Gough:2005:WDC

[2359] Nova Dudley-Gough. Wave of deceit crashes on the charitable. Network Security, 2005(1):1, January 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805001765.

McKenna:2005:RSO

[2360] Brian McKenna. Reuters steps up outsourced monitoring. Network Security, 2005(1):2, January 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805001777.

McKenna:2005:SBV

[2361] Brian McKenna. Symantec buys Veritas to broaden enterprise appeal. Network Security, 2005 (1):2–3, January 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805001789.

Anonymous:2005:Ba

[2362] Anonymous. In brief. Network Security, 2005(1):3, January 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805001790.

Hunter:2005:TSG

[2363] Philip Hunter. Tsunami spares global IT but shakes up disaster recovery plans. Network Security, 2005 (1):4–5, January 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805001807.

Schultz:2005:IES

[2364] E. Eugene Schultz. Internet Explorer security: is there any hope? Network Security, 2005(1):6–10, January 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485805001819.

Ionescu:2005:ESD

[2365] Ionut Ionescu. Embedded security and the drivers behind it. Network Security, 2005(1):10–13, January 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805001820.

Morgan:2005:LLN

[2366] David Morgan. Low-level network assessment: firewalls. Network Security, 2005(1):14–16, January 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805001832.

Potter:2005:BBW

[2367] Bruce Potter. Is big boss watching you? Network Security, 2005 (1):17–18, January 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805001844.

Kristensen:2005:MHT

[2368] Thomas Kristensen. More holes than a phishing net. Network Security, 2005 (1):18–19, January 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805001856.

Dwan:2005:PS

[2369] Berni Dwan. Pervasive spyware. Network Security, 2005(1):19, January 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805001868.

McKenna:2005:BBS

[2370] Brian McKenna. Bullish British SMEs poised to splash on IT security. Network Security, 2005 (1):20, January 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580500187X.

Dudley-Gough:2005:JBF

[2371] Nova Dudley-Gough. Jail for bug finding researcher? Network Security, 2005(1):20, January 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805001881.

Anonymous:2005:ECa

[2372] Anonymous. Events calendar. Network Security, 2005(1):20, January 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805001893.

McKenna:2005:STP

[2373] Brian McKenna. Symantec’s Thompson pronounces old style IT security dead. Network Security, 2005 (2):1, 3, February 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805001947.

Anonymous:2005:MPF

[2374] Anonymous. Microsoft pushes further into anti-virus and anti-spyware territory. Network Security, 2005 (2):1–2, February 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580500190X.

Mathieson:2005:PDG

[2375] SA Mathieson. Profession drive gathers momentum. Network Security, 2005(2):2, February 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805001911.

Hilley:2005:CWS

[2376] Sarah Hilley. Cisco widens security span in network. Network Security, 2005(2):2–3, February 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805001923.

McKenna:2005:GAS

[2377] Brian McKenna. Gates announces security invasion. Network Security, 2005(2):3, February 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805001935.

Anonymous:2005:Bb

[2378] Anonymous. In brief. Network Security, 2005(2):3, February 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805001959.

Edelson:2005:VIS

[2379] Eve Edelson. Voice over IP: security pitfalls. Network Security, 2005 (2):4–7, February 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805001960.

Stanton:2005:GAS

[2380] Ray Stanton. Getting ahead in security. Network Security, 2005 (2):8–10, February 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805001972.

Kemp:2005:WIB

[2381] Mike Kemp. When ignorance is not bliss. Network Security, 2005 (2):10–12, February 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805001984.

Knight:2005:WCH

[2382] William Knight. Worm cure is a hard act. Network Security, 2005 (2):12–14, February 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805001996.

Potter:2005:BAS

[2383] Bruce Potter. Bluetooth attacks start to bite. Network Security, 2005 (2):14–15, February 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580500200X.

Hunter:2005:MTV

[2384] Philip Hunter. Microsoft tackles viruses and spyware at last, but will it be trusted on security? Network Security, 2005(2): 16–17, February 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805002011.

Pollard:2005:TFC

[2385] Craig Pollard. Telecom fraud: the cost of doing nothing just went up. Network Security, 2005(2): 17–19, February 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805002023.

Kristensen:2005:IPA

[2386] Thomas Kristensen. IE patched after ‘extremely critical’ warning. Network Security, 2005(2):19–20, February 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805002035.

Anonymous:2005:ECb

[2387] Anonymous. Events calendar. Network Security, 2005(2):20, February 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805002047.

Everett:2005:CEP

[2388] Cath Everett. CA exposure provokes disclosure debate. Network Security, 2005(3):1–3, March 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805002059.

Anonymous:2005:UEV

[2389] Anonymous. US election-voting systems blamed for Bush win. Network Security, 2005(3):3, March 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805002060.

Anonymous:2005:ILS

[2390] Anonymous. Ivy League schools bar wanna-B students. Network Security, 2005(3):3, March 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805002072.

Anonymous:2005:Bc

[2391] Anonymous. In brief. Network Security, 2005(3):3–4, March 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805002084.

McKenna:2005:CSJ

[2392] Brian McKenna. Close shave for Japanese bank. Network Security, 2005(3):4, March 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805002096.

Potter:2005:DUU

[2393] Bruce Potter. Dangerous URLs: Unicode & IDN. Network Security, 2005(3):5–6, March 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805002102.

Kemp:2005:BSC

[2394] Mike Kemp. Building security credibility. Network Security, 2005 (3):6–8, March 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805002114.

Furnell:2005:AOW

[2395] Steven Furnell. Authenticating ourselves: will we ever escape the password? Network Security, 2005 (3):8–13, March 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805002126.

Heywood:2005:KYD

[2396] Mark Heywood. Keeping your data available in 10 steps. Network Security, 2005(3):14–16, March 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805002138.

Gosling:2005:TSE

[2397] Paul Gosling. Trojans & spyware: an electronic Achilles. Network Security, 2005(3):17–18, March 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580500214X.

Mathieson:2005:MPT

[2398] SA Mathieson. Mobile phone tracking threatens privacy. Network Security, 2005(3):19–20, March 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805002151.

Anonymous:2005:ECc

[2399] Anonymous. Events calendar. Network Security, 2005(3):20, March 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805002163.

Anonymous:2005:TDD

[2400] Anonymous. Tips to defeat DDoS.

Network Security, 2005(4):1–2, April 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805702178.

McKenna:2005:QTC

[2401] Brian McKenna. Qualys ticks compliance box. Network Security, 2005(4):2, April 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580570218X.

McKenna:2005:RHW

[2402] Brian McKenna. Russian hackers are world class. Network Security, 2005(4):3, April 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702191.

Anonymous:2005:Bd

[2403] Anonymous. In brief. Network Security, 2005(4):3, April 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702208.

Stanton:2005:ISP

[2404] Ray Stanton. Inside out security: de-perimeterisation. Network Security, 2005(4):4–6, April 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580570221X.

Stewart:2005:CAN

[2405] Andrew Stewart. A contemporary approach to network vulnerability assessment. Network Security, 2005(4):7–10, April 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702221.

Hilley:2005:CRM

[2406] Sarah Hilley. Crypto race for mathematical infinity. Network Security, 2005(4):10–11, April 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702233.

Kemp:2005:BES

[2407] Mike Kemp. Biometrics: the eye of the storm. Network Security, 2005(4):11–13, April 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702245.

Hunter:2005:PSL

[2408] Philip Hunter. Proactive security latest: vendors wire the cage but has the budgie flown …. Network Security, 2005(4):14–15, April 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702257.

Dean:2005:MAS

[2409] Roger Dean. Management aspects of secure messaging between organizations. Network Security, 2005 (4):16–17, April 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702269.

Potter:2005:RMU

[2410] Bruce Potter. RFID: misunderstood or untrustworthy? Network Security, 2005(4):17–18, April 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702270.

Grimes:2005:NSM

[2411] Galen A. Grimes. Network security managers’ preferences for the Snort IDS and GUI add-ons. Network Security, 2005(4):19–20, April 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485805702282.

Anonymous:2005:VFY

[2412] Anonymous. Vulnerabilities so far this year — A-V, Oracle, media players and Microsoft. Network Security, 2005(5):1–2, May 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702294.

McKenna:2005:CUI

[2413] Brian McKenna. Cisco unveils integrated security appliance family. Network Security, 2005(5):2–3, May 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485805702300.

Anonymous:2005:CPU

[2414] Anonymous. Check Point upgrades security platform. Network Security, 2005(5):3, May 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702312.

Anonymous:2005:Be

[2415] Anonymous. In brief. Network Security, 2005(5):3, May 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702324.

Licari:2005:BPI

[2416] Joe Licari. Best practices for instant messaging in business. Network Security, 2005(5):4–7, May 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702336.

Ernest-Jones:2005:UNS

[2417] Terry Ernest-Jones. US National Science Foundation: $19m security research programme. Network Security, 2005(5):8–9, May 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702348.

Beal:2005:ISP

[2418] Barry Beal. IT security: the product vendor landscape. Network Security, 2005(5):9–10, May 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580570235X.

Marteau:2005:MXS

[2419] Laurent Marteau. Mac OS X & security — an overview. Network Security, 2005(5):11–13, May 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702361.

Potter:2005:ISN

[2420] Bruce Potter. IT security needs hierarchy. Network Security, 2005 (5):14–15, May 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702373.

Kemp:2005:WBT

[2421] Mike Kemp. For whom the bells toll: effective IDS deployment strategies. Network Security, 2005(5):16–18, May 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702385.

Sidaway:2005:RRB

[2422] Garry Sidaway. The rise and rise of bot networks. Network Security, 2005(5):19–20, May 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702397.

Anonymous:2005:ECd

[2423] Anonymous. Events calendar. Network Security, 2005(5):20, May 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805702403.

Mathieson:2005:UCR

[2424] SA Mathieson. UK crypto regulation option dies. Network Security, 2005(6):2, June 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702415.

Anonymous:2005:ISP

[2425] Anonymous. IT security professionals devil-may-care about USB sticks. Network Security, 2005(6):2, June 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805702427.

McKenna:2005:DDI

[2426] Brian McKenna. Dorothy Denning on infosec and physical security. Network Security, 2005(6):3, June 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805702439.

Anonymous:2005:Bf

[2427] Anonymous. In brief. Network Security, 2005(6):3, June 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702440.

Simmonds:2005:UFB

[2428] Paul Simmonds. Users fight back by breaking the boundaries. Network Security, 2005(6):4–6, June 2005.

CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805702452.

Rowan:2005:SWS

[2429] Lannon Rowan. Security in a Web services world. Network Security, 2005(6):7–10, June 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702464.

Kemp:2005:BIG

[2430] Mike Kemp. Barbarians inside the gates: addressing internal security threats. Network Security, 2005 (6):11–13, June 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702476.

Potter:2005:WVA

[2431] Bruce Potter. Wireless vulnerability assessment. Network Security, 2005(6):14–15, June 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702488.

Kay:2005:LVV

[2432] James Kay. Low volume viruses: new tools for criminals. Network Security, 2005(6):16–18, June 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580570249X.

Lambert:2005:BS

[2433] Martin Lambert. Blackberry security. Network Security, 2005 (6):18–20, June 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702506.

Anonymous:2005:ECe

[2434] Anonymous. Events calendar. Network Security, 2005(6):20, June 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805702518.

Anonymous:2005:USP

[2435] Anonymous. US state police leaders confront CardSystems. Network Security, 2005(7):1, July 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580570252X.

McKenna:2005:CTA

[2436] Brian McKenna. Cyber-terrorist assault will happen, says Conservative MP. Network Security, 2005(7):2, July 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702531.

McKenna:2005:CCD

[2437] Brian McKenna. Credit card details in the clear and up for sale in India. Network Security, 2005(7):2, July 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702543.

McKenna:2005:BSS

[2438] Brian McKenna. Bank starts to sell PKI. Network Security, 2005(7):3, July 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702555.

Anonymous:2005:Bg

[2439] Anonymous. In brief. Network Security, 2005(7):3, July 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702567.

Potter:2005:QCS

[2440] Bruce Potter. Quantum crypto: Star trek or real science? Network Security, 2005(7):4–5, July 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702579.

Furnell:2005:ITE

[2441] Steven Furnell. Internet threats to end-users: Hunting easy prey. Network Security, 2005(7):5–9, July 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485805702580.

Stuttard:2005:SO

[2442] Dafydd Stuttard. Security & obscurity. Network Security, 2005 (7):10–12, July 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702592.

Rowe:2005:VET

[2443] Paul Rowe. VOIP — extra threats in the converged environment. Network Security, 2005(7):12–16, July 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805702609.

OConor:2005:ISO

[2444] Mark O’Conor. The implications of Sarbanes–Oxley for non-US IT departments. Network Security, 2005 (7):17–20, July 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702610.

Anonymous:2005:ECf

[2445] Anonymous. Events calendar. Network Security, 2005(7):20, July 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805702622.

McKenna:2005:NSA

[2446] Brian McKenna. New Symantec to avoid “over integration”. Network Security, 2005(8):2, August 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805702634.

McKenna:2005:ECB

[2447] Brian McKenna. Enterasys CTO blasts Cisco and ISS over hole controversy. Network Security, 2005 (8):2, August 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702646.

Anonymous:2005:BWB

[2448] Anonymous. Bidding war for bug hunter rewards. Network Security, 2005(8):3, August 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702658.

Anonymous:2005:Bh

[2449] Anonymous. In brief. Network Security, 2005(8):3–4, August 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580570266X.

Everett:2005:GN

[2450] Cath Everett. Godfathers of the net. Network Security, 2005(8): 4–20, August 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702671.

Ernest-Jones:2005:BCS

[2451] Terry Ernest-Jones. Business continuity strategy — the life line. Network Security, 2005(8):5–9, August 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805702683.

Bertrand:2005:BCM

[2452] Christophe Bertrand. Business continuity and mission critical applications. Network Security, 2005 (8):9–11, August 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702695.

McKenna:2005:DDL

[2453] Brian McKenna. Dig deep and look sideways. Network Security, 2005 (8):12–13, August 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702701.

Kemp:2005:BTS

[2454] Mike Kemp and Michael Kemp. Beyond trust: security policies and defence-in-depth. Network Security, 2005(8):14–16, August 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702713.

Potter:2005:SNC

[2455] Bruce Potter. Securing the network:

changing priorities. Network Security, 2005(8):16–17, August 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702725.

Bradbury:2005:BBF

[2456] Danny Bradbury. Bits and bridles for free speech. Network Security, 2005(8):17–19, August 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702737.

Anonymous:2005:ECg

[2457] Anonymous. Events calendar. Network Security, 2005(8):20, August 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805702749.

Anonymous:2005:FAY

[2458] Anonymous. FBI arrests young Turk and Moroccan for Zotob. Network Security, 2005(9):1–2, September 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805702750.

OHalloran:2005:ZV

[2459] Joe O’Halloran. Zotob and VoIP. Network Security, 2005(9):2–20, September 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485805702762.

Anonymous:2005:URI

[2460] Anonymous. US researchers to invent new Internet. Network Security, 2005(9):3, September 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702774.

Anonymous:2005:Bi

[2461] Anonymous. In brief. Network Security, 2005(9):3, September 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805702786.

Forte:2005:LME

[2462] Dario Forte. Log management for effective incident response. Network Security, 2005(9):4–7, September 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805702798.

Potter:2005:PD

[2463] Bruce Potter. Are passwords dead? Network Security, 2005 (9):7–8, September 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702804.

Morgan:2005:FPF

[2464] David Morgan. Fool-proofing firewalls. Network Security, 2005(9): 9–11, September 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702816.

Wood:2005:IIM

[2465] Peter Wood. Implementing identity management security — an ethical hacker’s view. Network Security, 2005 (9):12–15, September 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702828.

Sullivan:2005:CFI

[2466] Roger K. Sullivan. The case for federated identity. Network Security, 2005 (9):15–19, September 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580570283X.

Anonymous:2005:ECh

[2467] Anonymous. Events calendar. Network Security, 2005(9):20, September 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485805702841.

Mathieson:2005:GPH

[2468] SA Mathieson. Gone phishing in Halifax. Network Security, 2005 (10):2, October 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702853.

Anonymous:2005:MOP

[2469] Anonymous. Microsoft’s October patch-fest causes pain. Network Security, 2005(10):2–20, October 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485805702865.

Anonymous:2005:Bj

[2470] Anonymous. In brief. Network Security, 2005(10):3, October 2005.

CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805702877.

Pemble:2005:ETB

[2471] Matthew Pemble. Evolutionary trends in bank customer-targeted malware. Network Security, 2005 (10):4–7, October 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702889.

Dunn:2005:DYD

[2472] Kev Dunn. Dig yourself out of the data crater — database security isn’t new, so why can’t we get it right? Network Security, 2005 (10):7–10, October 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702890.

Potter:2005:EZD

[2473] Bruce Potter. The end of zero days? Network Security, 2005 (10):10–11, October 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702907.

Hallam-Baker:2005:PSN

[2474] Phillip Hallam-Baker. Prevention strategies for the next wave of cyber crime. Network Security, 2005 (10):12–15, October 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702919.

Whitworth:2005:OSB

[2475] Martin Whitworth. Outsourced security — the benefits and risks. Network Security, 2005(10):16–19, October 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485805702920.

Anonymous:2005:QWC

[2476] Anonymous. QualysGuard to work with Cisco NAC. Network Security, 2005(10):20, October 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702932.

Anonymous:2005:LTH

[2477] Anonymous. Lloyds TSB hook up 30,000 customers to anti-phishing token. Network Security, 2005 (10):20, October 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805702944.

Anonymous:2005:ECi

[2478] Anonymous. Events calendar. Network Security, 2005(10):20, October 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485805702956.

Anonymous:2005:MIS

[2479] Anonymous. Microsoft investigates spammers with trick computer. Network Security, 2005(11):

1–2, November 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702968.

McKenna:2005:ELE

[2480] Brian McKenna. Espionage-linked exploits grow in threat potential. Network Security, 2005(11):2, November 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580570297X.

Mathieson:2005:AIS

[2481] SA Mathieson. DoS attack not illegal, says London judge. Network Security, 2005(11):2–3, November 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805702981.

Anonymous:2005:Bk

[2482] Anonymous. In brief. Network Security, 2005(11):3, 20, November 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805702993.

Crosbie:2005:BES

[2483] Mark Crosbie. Biometrics for enterprise security. Network Security, 2005 (11):4–8, November 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805703007.

Davis:2005:RSI

[2484] Adrian Davis. Return on security investment — proving it’s worth it. Network Security, 2005(11): 8–10, November 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805703019.

Morgan:2005:ESP

[2485] David Morgan. The evolution of security purchasing. Network Security, 2005 (11):11–12, November 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805703020.

Keanini:2005:PTI

[2486] Tim Keanini. Protecting TCP/ IP. Network Security, 2005(11): 13–16, November 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485805703032.

Viveros:2005:CMT

[2487] Sal Viveros. Changing malware threats — AV vendor’s view. Network Security, 2005(11):16–18, November 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485805703044.

Potter:2005:LSV

[2488] Bruce Potter. Layer 2 security: in vogue. Network Security, 2005(11): 18–20, November 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805703056.

Anonymous:2005:ECj

[2489] Anonymous. Events calendar. Network Security, 2005(11):20, November 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805703068.

Anonymous:2005:MSP

[2490] Anonymous. Microsoft start playing hardball with anti-virus. Network Security, 2005(12):1–2, December 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580570307X.

McKenna:2005:ISV

[2491] Brian McKenna. Internal system vulnerabilities poorly patched. Network Security, 2005(12):2, December 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805703081.

Anonymous:2005:SWR

[2492] Anonymous. Sober worm returns as largest outbreak of 2005. Network Security, 2005(12):2, December 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805703093.

Anonymous:2005:Bl

[2493] Anonymous. In brief. Network Security, 2005(12):3, December 2005.

CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580570310X.

Sunner:2005:ESB

[2494] Mark Sunner. Email security best practice. Network Security, 2005 (12):4–7, December 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805703111.

Forte:2005:SMT

[2495] Dario Forte. Spyware: more than a costly annoyance. Network Security, 2005(12):8–10, December 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805703123.

Starkins:2005:OSN

[2496] Graham Starkins. Outside the safety net — communicating beyond traditional network boundaries. Network Security, 2005(12): 10–12, December 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805703135.

Newson:2005:NTV

[2497] Alan Newson. Network threats and vulnerability scanners. Network Security, 2005(12):13–15, December 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805703147.

Voice:2005:OAM

[2498] Chris Voice. Online authentication: matching security levels to the risk. Network Security, 2005(12): 15–18, December 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805703159.

Potter:2005:SSI

[2499] Bruce Potter. Safe and secure infrastructure upgrades. Network Security, 2005(12):19–20, December 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805703160.

Anonymous:2005:SPC

[2500] Anonymous. Spam and phish are coming to town. Network Security, 2005(12):20, December 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485805703172.

Anonymous:2005:ECk

[2501] Anonymous. Events calendar. Network Security, 2005(12):20, December 2005. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485805703184.

Mathieson:2006:AUF

[2502] SA Mathieson. Apps under fire. Network Security, 2006(1):1–2, January 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806703191.

Anonymous:2006:Ba

[2503] Anonymous. In brief. Network Security, 2006(1):3, January 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806703208.

Furnell:2006:RPS

[2504] Steven Furnell and Leith Zekri. Replacing passwords: in search of the secret remedy. Network Security, 2006(1):4–8, January 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580670321X.

Rowan:2006:SCN

[2505] Lannon Rowan. Securing converged networks — pitfalls and procedures. Network Security, 2006 (1):8–10, January 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806703221.

Philpott:2006:ITD

[2506] Andrew Philpott. Identity theft — dodging the own-goals. Network Security, 2006(1):11–13, January 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485806703233.

Macleod:2006:PTO

[2507] Calum Macleod. The password that’s overstayed its welcome. Network Security, 2006(1):13–15, January 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806703245.

Potter:2006:OSF

[2508] Bruce Potter. Open source firewall alternatives. Network Security, 2006 (1):16–17, January 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703257.

Heasman:2006:RT

[2509] John Heasman. Rootkit threats. Network Security, 2006(1):18–19, January 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806703269.

OHalloran:2006:SAS

[2510] Joe O’Halloran. Skype — an attack surface here to stay. Network Security, 2006(1):20, January 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703270.

Anonymous:2006:ECa

[2511] Anonymous. Events calendar. Network Security, 2006(1):20, January 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806703282.

Anonymous:2006:SSN

[2512] Anonymous. Spyware spikes in northern Europe. Network Security, 2006 (2):1–2, February 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703294.

Mathieson:2006:FTO

[2513] SA Mathieson. Fine-tuning for Olympic IT security. Network Security, 2006(2):2, February 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703300.

McKenna:2006:GMC

[2514] Brian McKenna. Gates and Mc-Neally commit to sharing on Valentine’s Day. Network Security, 2006 (2):2, 20, February 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703312.

Anonymous:2006:Bb

[2515] Anonymous. In brief. Network Security, 2006(2):3, February 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806703324.

Potter:2006:NTA

[2516] Bruce Potter. New threat of Apple Mac OS X. Network Security, 2006(2):4–5, February 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703336.

Wood:2006:HTF

[2517] Peter Wood. The hacker’s top five routes into the network (and how to block them). Network Security, 2006(2):5–9, February 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703348.

Morgan:2006:RMG

[2518] David Morgan. Risk management — getting properly valued. Network Security, 2006(2):10–11, February 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580670335X.

Stevens:2006:UOS

[2519] Mark Stevens. UTM: one-stop protection. Network Security, 2006(2): 12–14, February 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703361.

Johnson:2006:SAS

[2520] Everett C. Johnson. Security awareness: switch to a better programme. Network Security, 2006 (2):15–18, February 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703373.

Cassidy:2006:NPS

[2521] Darren Cassidy. Network printing security — getting to grips with the multifunction device. Network Security, 2006 (2):19–20, February 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806703385.

Anonymous:2006:ECb

[2522] Anonymous. Events calendar. Network Security, 2006(2):20, February 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485806703397.

Mathieson:2006:JLM

[2523] SA Mathieson. John Lewis monitors moves on CCTV. Network Security, 2006(3):1–2, March 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806703403.

Anonymous:2006:ICB

[2524] Anonymous. ISS’s CTO on 2006 — botnet armies and security services online. Network Security, 2006 (3):2, 20, March 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806703415.

Anonymous:2006:Bc

[2525] Anonymous. In brief. Network Security, 2006(3):3, March 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806703427.

Furnell:2006:UPW

[2526] Steven Furnell and Bogdan Ghita. Usability pitfalls in wireless LAN security. Network Security, 2006 (3):4–8, March 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703439.

Morgan:2006:WIA

[2527] David Morgan. Web injection attacks. Network Security, 2006 (3):8–10, March 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703440.

Waite:2006:SOB

[2528] Steve Waite. Securing online business with SSL. Network Security, 2006(3):10–12, March 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703452.

McGurk:2006:SDS

[2529] Ian McGurk. Stealing data on small devices — tools to wipe out the iPod predator. Network Security, 2006 (3):13–16, March 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703464.

Frost:2006:VTG

[2530] Nick Frost. VoIP threats — getting louder. Network Security, 2006 (3):16–18, March 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703476.

Potter:2006:BSM

[2531] Bruce Potter. Bluetooth security moves. Network Security, 2006 (3):19–20, March 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806703488.

Anonymous:2006:ECc

[2532] Anonymous. Events calendar. Network Security, 2006(3):20, March 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580670349X.

Anonymous:2006:UBC

[2533] Anonymous. UK business careless with online data. Network Security, 2006(4):1–2, April 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806703506.

Hilley:2006:HSO

[2534] Sarah Hilley. HSBC to secure online business customers. Network Security, 2006(4):2, April 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806703518.

Anonymous:2006:Bd

[2535] Anonymous. In brief. Network Security, 2006(4):3, April 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580670352X.

Morgan:2006:WAS

[2536] David Morgan. Web application security — SQL injection attacks. Network Security, 2006 (4):4–5, April 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703531.

Khokhar:2006:SCB

[2537] Rafael Khokhar. Smartphones — a call for better safety on the move. Network Security, 2006 (4):6–7, April 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703543.

Rana:2006:SVOa

[2538] Omer Rana and Jeremy Hilton. Securing the virtual organization — Part 1: Requirements from grid computing. Network Security, 2006 (4):7–10, April 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703555.

Whitworth:2006:VCB

[2539] Martin Whitworth. VoIP — a call for better protection. Network Security, 2006(4):11–12, April 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703567.

delaCuadra:2006:ASB

[2540] Fernando de la Cuadra. The art of security budgeting — a new structure. Network Security, 2006 (4):12–15, April 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703579.

Potter:2006:UEH

[2541] Bruce Potter. User education — how valid is it? Network Security, 2006(4):15–16, April 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703580.

Forte:2006:KSI

[2542] Dario Forte. Keeping up to standard — incident management in ISO 17799. Network Security, 2006 (4):16–17, April 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703592.

Henry:2006:TFA

[2543] Paul A. Henry. Two-factor authentication — a look behind the headlines. Network Security, 2006 (4):18–19, April 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703609.

Anonymous:2006:FCI

[2544] Anonymous. Fix for critical IE flaw. Network Security, 2006(4):20, April 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806703622.

Anonymous:2006:SBT

[2545] Anonymous. Small botnets are threat.

Network Security, 2006(4):20, April 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806703634.

Anonymous:2006:ECd

[2546] Anonymous. Events calendar. Network Security, 2006(4):20, April 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806703646.

Hilley:2006:SSD

[2547] Sarah Hilley. Secret Service dismantles web forums. Network Security, 2006(4):20, April 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703610.

Anonymous:2006:NMS

[2548] Anonymous. New McAfee study reveals risk of using search engines. Network Security, 2006(5):1–2, May 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806703798.

Anonymous:2006:HE

[2549] Anonymous. Hacker to be extradited.

Network Security, 2006(5):2, May 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806703804.

Anonymous:2006:MWS

[2550] Anonymous. Mac and Windows systems at risk. Network Security, 2006(5):2, May 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703816.

Anonymous:2006:MMC

[2551] Anonymous. Money makes e-crime go round. Network Security, 2006 (5):2, 20, May 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703828.

Anonymous:2006:Be

[2552] Anonymous. In brief. Network Security, 2006(5):3, May 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580670383X.

Bergman:2006:SBN

[2553] Alisa Bergman and Agnieszka Verlet. Security breaches: to notify or not to notify — that is the question. Network Security, 2006(5):4–6, May 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703841.

Rana:2006:SVOb

[2554] Omer Rana and Jeremy Hilton. Securing the virtual organization, Part 2 — grid computing in action. Network Security, 2006 (5):6–10, May 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703853.

Spencer:2006:VAD

[2555] Dave Spencer. Vulnerability analysis — discovery to disclosure. Network Security, 2006(5):11–12, May 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806703865.

Surgenor:2006:SRS

[2556] Valerie Surgenor. Setting the records straight … and staying on the right side of the law. Network Security, 2006(5):13–15, May 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703877.

Walker:2006:BCO

[2557] Alan Walker. Business continuity and outsourcing — moves to take out the risk. Network Security, 2006(5):15–17, May 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703889.

Potter:2006:TTA

[2558] Bruce Potter. The times they are a’changing: trends that impact security strategy. Network Security, 2006(5):18–19, May 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806703890.

Murton:2006:CGH

[2559] Daniel Murton. Crypto goes to the heart. Network Security, 2006(5): 19, May 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1353485806703907.

Anonymous:2006:ECe

[2560] Anonymous. Events calendar. Network Security, 2006(5):20, May 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485806703919.

Anonymous:2006:SQE

[2561] Anonymous. Success for quantum encryption? Network Security, 2006(5):20, May 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806703920.

Hilley:2006:FYC

[2562] Sarah Hilley. Five years for Californian botmaster. Network Security, 2006(6):1–2, June 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806703932.

McKenna:2006:RWS

[2563] Brian McKenna. Real world security model for infosec, says Microsoft security guru. Network Security, 2006(6):2, June 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703944.

Anonymous:2006:TUP

[2564] Anonymous. Trojan uses pornography trap to steal bank details. Network Security, 2006(6):2, 20, June 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806703956.

Anonymous:2006:Bf

[2565] Anonymous. In brief. Network Security, 2006(6):3, June 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806703968.

Emm:2006:FTH

[2566] David Emm. Focus on Trojans — holding data to ransom. Network Security, 2006(6):4–7, June 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580670397X.

Forte:2006:EPS

[2567] Dario Forte. Endpoint and perimeter security: a new symbiosis. Network Security, 2006(6):7–8, June 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806703981.

Ernest-Jones:2006:PSP

[2568] Terry Ernest-Jones. Pinning down a security policy for mobile data.

Network Security, 2006(6):8–12, June 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485806703993.

Crosbie:2006:TCC

[2569] Mark Crosbie. Trusted computing — closing that lingering doubt. Network Security, 2006(6):13–15, June 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485806704007.

Morgan:2006:MSW

[2570] David Morgan. Maintaining state in Web applications. Network Security, 2006(6):16–18, June 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806704019.

McKewan:2006:BZG

[2571] Andy McKewan. Botnets — zombies get smarter. Network Security, 2006(6):18–20, June 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806704020.

Anonymous:2006:ECf

[2572] Anonymous. Events calendar. Network Security, 2006(6):20, June 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485806704032.

Anonymous:2006:ABS

[2573] Anonymous. Is Apple the best security of all? Network Security, 2006(7):1–2, July 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704044.

Anonymous:2006:GWE

[2574] Anonymous. Germany wins the European Cup. Network Security, 2006(7):2, July 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704056.

Anonymous:2006:TUW

[2575] Anonymous. Theft of 26.5m US war veterans’ details. Network Security, 2006(7):2, 20, July 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704068.

Anonymous:2006:Bg

[2576] Anonymous. In brief. Network Security, 2006(7):3, July 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580670407X.

OSullivan:2006:IMV

[2577] Shirley O’Sullivan. Instant Messaging vs. instant compromise. Network Security, 2006(7):4–6, July 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704081.

Alcorn:2006:CSS

[2578] Wade Alcorn. Cross-site scripting viruses and worms — a new attack vector. Network Security, 2006(7):7–8, July 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806704093.

Watson:2006:SPS

[2579] Ian Watson. Securing portable storage devices. Network Security, 2006(7):8–11, July 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580670410X.

Small:2006:USR

[2580] Mike Small. Unify and simplify: re-thinking identity management. Network Security, 2006 (7):11–14, July 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806704111.

Gordon-Smith:2006:VMP

[2581] Matt Gordon-Smith and Stephen Wing. VoIP and multimedia protocols in wireless and firewalled environments. Network Security, 2006 (7):14–17, July 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806704123.

Bech:2006:SLD

[2582] Soeren Bech. Shedding light on dark traffic attacks. Network Security, 2006(7):17–19, July 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704135.

Anonymous:2006:DSW

[2583] Anonymous. Data security worries still top the list. Network Security, 2006(7):20, July 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704147.

Anonymous:2006:JHD

[2584] Anonymous. Just 12% of handheld devices encrypted. Network Security, 2006(7):20, July 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704159.

Anonymous:2006:ECg

[2585] Anonymous. Events calendar. Network Security, 2006(7):20, July 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806704160.

Anonymous:2006:Bh

[2586] Anonymous. In brief. Network Security, 2006(8):2, August 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704287.

Lamb:2006:SFW

[2587] Stephen Lamb. Security features in Windows Vista and IE7 — Microsoft’s view. Network Security, 2006(8):3–7, August 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704299.

Broomfield:2006:NAD

[2588] Mike Broomfield. NTFS Alternate Data Streams: focused hacking. Network Security, 2006(8):7–9, August 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806704305.

Furnell:2006:SMD

[2589] Steven Furnell. Securing mobile devices: technology and attitude. Network Security, 2006 (8):9–13, August 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704317.

Emm:2006:PUH

[2590] David Emm. Phishing update, and how to avoid getting hooked. Network Security, 2006(8):13–15, August 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806704329.

Mueller:2006:WHB

[2591] Lance Mueller. Webjacking, and how to boot it out. Network Security, 2006(8):15–18, August 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704330.

Marsanu:2006:VCM

[2592] Michael Marsanu. VoIP: carrying the most sensitive data? Network Security, 2006(8):19–20, August 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704342.

Anonymous:2006:ECh

[2593] Anonymous. Events calendar. Network Security, 2006(8):20, August 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806704354.

Anonymous:2006:UTB

[2594] Anonymous. Upturn in Trojans and bots. Network Security, 2006 (9):1–2, September 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704172.

Anonymous:2006:NMV

[2595] Anonymous. No more the virus. Network Security, 2006(9):2, September 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806704184.

Anonymous:2006:Bi

[2596] Anonymous. In brief. Network Security, 2006(9):3, September 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806704196.

Powell:2006:TWW

[2597] Graeme Powell. ‘Tmobile’ — wising up to wireless exposure. Network Security, 2006(9):4–6, September 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704202.

Ernest-Jones:2006:TAS

[2598] Terry Ernest-Jones. Threat analysis — stretching the limits. Network Security, 2006(9):6–8, September 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704214.

Byrne:2006:AFD

[2599] Paul Byrne. Application firewalls in a defence-in-depth design. Network Security, 2006(9): 9–11, September 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704226.

Titmus:2006:SIT

[2600] Peter Titmus. Securing IP telephony systems — best practices. Network Security, 2006(9):11–13, September 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806704238.

Solterbeck:2006:PDR

[2601] Andy Solterbeck. Protecting data at rest and in motion. Network Security, 2006(9):14–17, September 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580670424X.

Braganza:2006:CSS

[2602] Richard Braganza. Cross-site scripting — an alternative view. Network Security, 2006(9):17–20, September 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806704251.

Anonymous:2006:IBI

[2603] Anonymous. IBM brings ISS into the family. Network Security, 2006 (9):20, September 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704263.

Anonymous:2006:ECi

[2604] Anonymous. Events calendar. Network Security, 2006(9):20, September 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806704275.

Anonymous:2006:MFI

[2605] Anonymous. Microsoft to the fore at ISSE 2006. Network Security, 2006 (10):1–2, October 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704366.

Anonymous:2006:SVB

[2606] Anonymous. Security vendors barred from Microsoft’s Vista. Network Security, 2006(10):2, October 2006.

CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806704378.

Anonymous:2006:Bj

[2607] Anonymous. In brief. Network Security, 2006(10):3, October 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580670438X.

Forte:2006:SIR

[2608] Dario Forte. Sharpening incident response. Network Security, 2006 (10):4–5, October 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704391.

Marsanu:2006:SSP

[2609] Michael Marsanu. Is size still a problem for wireless LANs? Network Security, 2006(10):6–8, October 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704408.

Surgenor:2006:PDS

[2610] Valerie Surgenor. The price of data sharing. Network Security, 2006 (10):8–10, October 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580670441X.

Boyd:2006:RBK

[2611] Chris Boyd. Rogue browsers — keeping Browsezilla & Co at bay. Network Security, 2006(10):11–12, October 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806704421.

Williams:2006:CMW

[2612] Paul Williams. Cappuccino, muffin, WiFi — but what about the security? Network Security, 2006 (10):13–17, October 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704433.

Murfitt:2006:CTI

[2613] Martin Murfitt. The case for a tiered internal network. Network Security, 2006(10):17–20, October 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806704445.

Anonymous:2006:ECj

[2614] Anonymous. Events calendar. Network Security, 2006(10):20, October 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806704457.

Anonymous:2006:BBC

[2615] Anonymous. BT buys counterpane to boost IT security services. Network Security, 2006(11): 1–2, November 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704469.

McKenna:2006:SIP

[2616] Brian McKenna. Senior infosec pros shift gears from technology to management. Network Security, 2006 (11):2, November 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806704470.

Anonymous:2006:SNS

[2617] Anonymous. ‘Social networking’ study shows cybercrime risk. Network Security, 2006(11):2, November 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485806704482.

Anonymous:2006:Bk

[2618] Anonymous. In brief. Network Security, 2006(11):3, November 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485806704494.

Emm:2006:MMN

[2619] David Emm. Mobile malware — new avenues. Network Security, 2006 (11):4–6, November 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806704500.

Furnell:2006:SHW

[2620] Steven Furnell and Network Research Group. Securing the home worker. Network Security, 2006(11): 6–12, November 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704512.

Wood:2006:SUW

[2621] Peter Wood. In search of the uncrackable Windows password. Network Security, 2006(11):12–13, November 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806704524.

Peacey:2006:TES

[2622] Angus Peacey. Teleworkers — extending security beyond the office. Network Security, 2006(11): 14–16, November 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704536.

Potter:2006:CFI

[2623] Bruce Potter. The changing face of IT security. Network Security, 2006 (11):16–17, November 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704548.

Dunn:2006:PLP

[2624] Kev Dunn. Poor lockdown practices pave the way to the domain controller. Network Security, 2006 (11):18–20, November 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580670455X.

Anonymous:2006:ECk

[2625] Anonymous. Events calendar. Network Security, 2006(11):20, November 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485806704561.

Anonymous:2006:IMR

[2626] Anonymous. 81% of IT managers report greynet incidents. Network Security, 2006(12):2, December 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485806704573.

Anonymous:2006:BBS

[2627] Anonymous. Botnets behind spam surge. Network Security, 2006(12): 2, 20, December 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806704585.

Anonymous:2006:Bl

[2628] Anonymous. In brief. Network Security, 2006(12):3, December 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485806704597.

Dean:2006:IMB

[2629] Roger Dean. Identity management — back to the user. Network Security, 2006(12):4–7, December 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806704603.

Neuenschwander:2006:IMM

[2630] Mike Neuenschwander. Identity management market shifts — who’s out there? Network Security, 2006 (12):7–10, December 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704615.

Perry:2006:NFI

[2631] Simon Perry. Network forensics and the inside job. Network Security, 2006 (12):11–13, December 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485806704627.

Potter:2006:CKM

[2632] Bruce Potter. Cryptographic key management for the masses. Network Security, 2006(12):13–14, December 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806704639.

Best:2006:EBT

[2633] Tim Best. Eying up biometrics — time to go multimodal? Network Security, 2006(12):15–17, December 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485806704640.

Matthews:2006:SCA

[2634] Adam Matthews. Side-channel attacks on smartcards. Network Security, 2006 (12):18–20, December 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485806704652.

Anonymous:2006:ECl

[2635] Anonymous. Events calendar. Network Security, 2006(12):20, December 2006. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485806704664.

Anonymous:2007:UGD

[2636] Anonymous. UK Government drops single-database ID registry. Network Security, 2007(1):1–2, January 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485807700016.

Anonymous:2007:PIU

[2637] Anonymous. Phishing increases as users get wise. Network Security, 2007(1):2, January 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700028.

Anonymous:2007:Ba

[2638] Anonymous. In brief. Network Security, 2007(1):3, January 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580770003X.

Watson:2007:HTC

[2639] David Watson. Honeynets: a tool for counterintelligence in online security. Network Security, 2007 (1):4–8, January 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700041.

Heron:2007:WBH

[2640] Simon Heron. Working the botnet:

how dynamic DNS is revitalising the zombie army. Network Security, 2007 (1):9–11, January 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700053.

Kinge:2007:FSU

[2641] Steve Kinge. Freedom of speech: Using speech biometrics for user verification. Network Security, 2007 (1):12–14, January 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700065.

Forte:2007:SSI

[2642] Dario Forte. Security standardization in incident management: the ITIL approach. Network Security, 2007 (1):14–16, January 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700077.

Potter:2007:RVB

[2643] Bruce Potter. Running Vista — big OS, big change. Network Security, 2007(1):17–18, January 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700089.

Casey:2007:BSI

[2644] Donal Casey. Building a secure instant messaging environment. Network Security, 2007(1): 18–20, January 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700090.

Anonymous:2007:ECa

[2645] Anonymous. Events calendar. Network Security, 2007(1):20, January 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485807700107.

Anonymous:2007:STU

[2646] Anonymous. Security through uncertainty. Network Security, 2007 (2):1, February 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700119.

Anonymous:2007:PTC

[2647] Anonymous. Pen test certifications prepare to compete. Network Security, 2007(2):1–2, February 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700120.

Anonymous:2007:RNS

[2648] Anonymous. Root name servers come under attack. Network Security, 2007(2):2, February 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700132.

Anonymous:2007:PWT

[2649] Anonymous. Peacomm worm takes net by surprise. Network Security, 2007(2):2, February 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700144.

Anonymous:2007:Bb

[2650] Anonymous. In brief. Network Security, 2007(2):3, February 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485807700156.

Harrison:2007:STU

[2651] Keith Harrison, Bill Munro, and Tim Spiller. Security through uncertainty. Network Security, 2007 (2):4–7, February 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700168.

Fafinski:2007:SRP

[2652] Stefan Fafinski. The security ramifications of the Police and Justice Act 2006. Network Security, 2007 (2):8–11, February 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580770017X.

Potter:2007:WDG

[2653] Bruce Potter. The widening disclosure gap. Network Security, 2007 (2):11–12, February 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700181.

Davis:2007:AAA

[2654] Adrian Davis. Authentication across the airwaves. Network Security, 2007 (2):13–19, February 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700193.

Forte:2007:NED

[2655] Dario Forte. The negative effects of ediscovery rules. Network Security, 2007 (2):19–20, February 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580770020X.

Anonymous:2007:ECb

[2656] Anonymous. Events calendar. Network Security, 2007(2):20, February 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485807700211.

Anonymous:2007:VSV

[2657] Anonymous. Vista security verdicts roll in. Network Security, 2007(3):1–2, March 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700223.

Anonymous:2007:EPE

[2658] Anonymous. Expert pushes envelope with passport RFID crack. Network Security, 2007(3):2, 20, March 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485807700235.

Anonymous:2007:Bc

[2659] Anonymous. In brief. Network Security, 2007(3):3, March 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700247.

Ritchie:2007:SRA

[2660] Paul Ritchie. The security risks of AJAX/web 2.0 applications. Network Security, 2007(3):4–8, March 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485807700259.

Westcott:2007:MRS

[2661] Ross Westcott. Maximizing the ROI of a security audit. Network Security, 2007(3):8–11, March 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700260.

deVries:2007:STS

[2662] Stephen de Vries. Software testing for security. Network Security, 2007(3):11–15, March 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700272.

Storey:2007:NDS

[2663] Dominic Storey. Network discovery and its security applications. Network Security, 2007 (3):15–18, March 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700284.

Potter:2007:OCN

[2664] Bruce Potter. Our changing network borders. Network Security, 2007 (3):18–20, March 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700296.

Anonymous:2007:ECc

[2665] Anonymous. Events calendar. Network Security, 2007(3):20, March 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485807700302.

Anonymous:2007:STC

[2666] Anonymous. Symantec tracks commercialisation of malware. Network Security, 2007(4):1–2, April 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700399.

Anonymous:2007:ACP

[2667] Anonymous. Animated cursor points to trouble for Microsoft. Network Security, 2007(4):2, April 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700405.

Anonymous:2007:UPP

[2668] Anonymous. UK plc pinpoints IM, VoIP as security risks. Network Security, 2007(4):2, April 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700417.

Anonymous:2007:Bd

[2669] Anonymous. In brief. Network Security, 2007(4):3, April 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700429.

Rowan:2007:AFF

[2670] Tom Rowan. Application firewalls:

filling the void. Network Security, 2007(4):4–7, April 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700430.

Owen:2007:FOR

[2671] Michael Owen. Fun with onion routing. Network Security, 2007 (4):8–12, April 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700442.

Heron:2007:BCC

[2672] Simon Heron. Botnet command and control techniques. Network Security, 2007(4):13–16, April 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700454.

Anonymous:2007:DBA

[2673] Anonymous. Data breach activity is getting worse. Network Security, 2007(4):16, April 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700466.

delaCuadra:2007:GM

[2674] Fernando de la Cuadra. The genealogy of malware. Network Security, 2007(4):17–20, April 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700478.

Anonymous:2007:ECd

[2675] Anonymous. Events calendar. Network Security, 2007(4):20, April 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580770048X.

Anonymous:2007:UIV

[2676] Anonymous. UK Internet voting comes under attack. Network Security, 2007 (5):1–2, 20, May 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700314.

Anonymous:2007:Be

[2677] Anonymous. In brief. Network Security, 2007(5):3, May 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700326.

Anderson:2007:CHA

[2678] Kent Anderson. Convergence: a holistic approach to risk management. Network Security, 2007(5):4–7, May 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485807700338.

Gleave:2007:MLI

[2679] Stephen Gleave. The mechanics of lawful interception. Network Security, 2007(5):8–11, May 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580770034X.

Potter:2007:BGS

[2680] Bruce Potter and Shmoo Group. Being a good security citizen. Network Security, 2007(5):11–12, May 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485807700351.

Everett:2007:TFO

[2681] Bernard Everett. Tapping into fibre optic cables. Network Security, 2007(5):13–16, May 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700363.

Ollmann:2007:WSC

[2682] Gunter Ollmann. Writing secure code. Network Security, 2007 (5):16–20, May 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700375.

Anonymous:2007:ECe

[2683] Anonymous. Events calendar. Network Security, 2007(5):20, May 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485807700387.

Anonymous:2007:FSL

[2684] Anonymous. FBI still lax on security. Network Security, 2007 (6):1–2, June 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700491.

Anonymous:2007:BBB

[2685] Anonymous. Better Business Bureau phishing attacks continue. Network Security, 2007(6):2, June 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700508.

Anonymous:2007:Bf

[2686] Anonymous. In brief. Network Security, 2007(6):3, June 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580770051X.

Heron:2007:RRK

[2687] Simon Heron. The rise and rise of the keyloggers. Network Security, 2007(6):4–6, June 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700521.

Munro:2007:WW

[2688] Ken Munro. Wireless worms. Network Security, 2007(6):6–8, June 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485807700533.

Owen:2007:NBC

[2689] Michael Owen and Colin Dixon.

A new baseline for cardholder security. Network Security, 2007 (6):8–12, June 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700545.

Dantu:2007:SMN

[2690] Ram Dantu, Herman Oosterwijk, Prakash Kolan, and Husain Husna. Securing medical networks. Network Security, 2007(6):13–16, June 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485807700557.

LeVine:2007:PSS

[2691] Richard LeVine. Putting the spirit into security. Network Security, 2007(6):17–20, June 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700569.

Anonymous:2007:SSI

[2692] Anonymous. Seattle spammer indicted. Network Security, 2007(6):20, June 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485807700570.

Anonymous:2007:ECf

[2693] Anonymous. Events calendar. Network Security, 2007(6):20, June 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485807700582.

Anonymous:2007:YHS

[2694] Anonymous. YouTube hit by security concerns. Network Security, 2007(7):1–2, July 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700594.

Anonymous:2007:MPN

[2695] Anonymous. MPack packs a nasty punch. Network Security, 2007(7): 2, July 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1353485807700600.

Anonymous:2007:Bg

[2696] Anonymous. In brief. Network Security, 2007(7):3, July 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700612.

Clark:2007:DMT

[2697] Andrew J. Clark. Document metadata, tracking and tracing. Network Security, 2007(7):4–7, July 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700624.

Potter:2007:AWB

[2698] Bruce Potter. Alternative web browsers: friend or foe? Network Security, 2007(7):7–8, July 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700636.

Badra:2007:AWC

[2699] Mohamad Badra. Alternative wireless client authentication and key distribution. Network Security, 2007 (7):9–13, July 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700648.

Mattsson:2007:DD

[2700] Ulf Mattsson. Defending the database. Network Security, 2007 (7):14–17, July 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580770065X.

deBorde:2007:STF

[2701] Duncan de Borde. Selecting a twofactor authentication system. Network Security, 2007(7):17–20, July 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485807700661.

Anonymous:2007:ECg

[2702] Anonymous. Events calendar. Network Security, 2007(7):20, July 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485807700673.

Anonymous:2007:WLV

[2703] Anonymous. WabiSabiLabi launches vulnerability market. Network Security, 2007(8):1–2, August 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485807700685.

Anonymous:2007:SFS

[2704] Anonymous. Spam formats shift again. Network Security, 2007 (8):2, August 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700697.

Anonymous:2007:Bh

[2705] Anonymous. In brief. Network Security, 2007(8):3, August 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700703.

Frith:2007:SAO

[2706] David Frith. Steganography approaches, options, and implications. Network Security, 2007 (8):4–7, August 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700715.

Coursen:2007:FMM

[2707] Shane Coursen. The future of mobile malware. Network Security, 2007 (8):7–11, August 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700727.

Williams:2007:EBR

[2708] Paul Williams. Executive and board roles in information security. Network Security, 2007(8): 11–14, August 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700739.

Hwang:2007:HDM

[2709] Tim Hwang. Herdict: a distributed model for threats online. Network Security, 2007(8): 15–18, August 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700740.

Potter:2007:MSR

[2710] Bruce Potter. Mobile security risks:

ever evolving. Network Security, 2007 (8):19–20, August 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700752.

Anonymous:2007:ECh

[2711] Anonymous. Events calendar. Network Security, 2007(8):20, August 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485807700764.

Anonymous:2007:LWB

[2712] Anonymous. Lords want banks, tech firms held responsible. Network Security, 2007(9):1–2, September 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485807700776.

Anonymous:2007:ABM

[2713] Anonymous. Attachment-based malware takes back seat to spam. Network Security, 2007(9):2, 20, September 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485807700788.

Anonymous:2007:Bi

[2714] Anonymous. In brief. Network Security, 2007(9):3, September 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580770079X.

Laurie:2007:PAA

[2715] Adam Laurie. Practical attacks against RFID. Network Security, 2007 (9):4–7, September 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700806.

Armstrong:2007:NIC

[2716] Neil Armstrong. Network infiltration with client-side attacks. Network Security, 2007(9):8–10, September 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485807700818.

Rowan:2007:IPS

[2717] Tom Rowan. Intrusion prevention systems: superior security. Network Security, 2007(9):11–15, September 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580770082X.

DeGuerre:2007:MVR

[2718] Jamie De Guerre. The mechanics of Vipul’s Razor technology. Network Security, 2007(9):15–17, September 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485807700831.

Potter:2007:SA

[2719] Bruce Potter. Security automation. Network Security, 2007(9): 18–19, September 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700843.

Anonymous:2007:MCS

[2720] Anonymous. monster.com suffers web attack. Network Security, 2007 (9):20, September 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700855.

Anonymous:2007:NSN

[2721] Anonymous. Network Security’s new look. Network Security, 2007 (9):20, September 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700867.

Anonymous:2007:ECi

[2722] Anonymous. Events calendar. Network Security, 2007(9):20, September 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485807700879.

Anonymous:2007:UDD

[2723] Anonymous. US DoD, DHS criticised for security flaws. Network Security, 2007(10):1–2, October 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700880.

Anonymous:2007:VPS

[2724] Anonymous. VMware patches security flaws. Network Security, 2007 (10):2, October 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700892.

Anonymous:2007:E

[2725] Anonymous. Erratum. Network Security, 2007(10):2, October 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485807700909.

Anonymous:2007:Bj

[2726] Anonymous. In brief. Network Security, 2007(10):3, October 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485807700910.

Skapinetz:2007:VBT

[2727] Kevin Skapinetz. Virtualisation as a blackhat tool. Network Security, 2007(10):4–7, October 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700922.

Clement:2007:SIG

[2728] Miles Clement. Securing the iPod generation. Network Security, 2007 (10):7–9, October 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700934.

Watson:2007:WAA

[2729] David Watson. Web application attacks. Network Security, 2007 (10):10–14, October 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700946.

Midian:2007:NRD

[2730] Paul Midian. Networks at risk: detection, defence and dealing with it. Network Security, 2007(10): 14–18, October 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700958.

Potter:2007:CWW

[2731] Bruce Potter. Converging wired and wireless authentication. Network Security, 2007(10):18–20, October 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580770096X.

Anonymous:2007:ECj

[2732] Anonymous. Events calendar. Network Security, 2007(10):20, October 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485807700971.

Anonymous:2007:ILS

[2733] Anonymous. IBM launches security offensive. Network Security, 2007 (11):1–2, November 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700357.

Anonymous:2007:HPS

[2734] Anonymous. Humans + porn = solved Captcha. Network Security, 2007(11):2, November 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700369.

Anonymous:2007:Bk

[2735] Anonymous. In brief. Network Security, 2007(11):3, November 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808700370.

Heron:2007:GCO

[2736] Simon Heron. Gang culture in the online world. Network Security, 2007 (11):4–7, November 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700382.

Watson:2007:EWA

[2737] David Watson. The evolution of web application attacks. Network Security, 2007(11):7–12, November 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808700394.

Moyle:2007:BTS

[2738] Steve Moyle. The blackhat’s toolbox:

SQL injections. Network Security, 2007 (11):12–14, November 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700400.

Frith:2007:VSH

[2739] David Frith. E-voting security: hope or hype? Network Security, 2007 (11):14–16, November 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700412.

Clement:2007:IES

[2740] Miles Clement. Issues in endpoint security. Network Security, 2007(11): 17–18, November 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700424.

Potter:2007:CP

[2741] Bruce Potter. From concept to production. Network Security, 2007(11): 19–20, November 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700436.

Anonymous:2007:ECk

[2742] Anonymous. Events calendar. Network Security, 2007(11):20, November 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485808700448.

Anonymous:2007:DBC

[2743] Anonymous. Data breach casts pall on UK ID card scheme. Network Security, 2007(12):1–2, December 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807700983.

Anonymous:2007:SPR

[2744] Anonymous. 2008 security predictions roll in. Network Security, 2007 (12):2, 20, December 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807700995.

Anonymous:2007:Bl

[2745] Anonymous. In brief. Network Security, 2007(12):3, December 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485807701009.

Sunner:2007:RTT

[2746] Mark Sunner. The rise of targeted Trojans. Network Security, 2007 (12):4–7, December 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807701010.

Newman:2007:HIS

[2747] Kenneth Newman. Home invasion: securing home access to business networks. Network Security, 2007(12): 8–10, December 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485807701022.

Abimbola:2007:ISI

[2748] Abiola Abimbola. Information security incident response. Network Security, 2007(12):10–13, December 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485807701034.

Rowan:2007:VTI

[2749] Tom Rowan. VPN technology: IPSEC vs SSL. Network Security, 2007(12): 13–17, December 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807701046.

Potter:2007:NFA

[2750] Bruce Potter. Network flow analysis. Network Security, 2007(12): 18–19, December 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807701058.

Anonymous:2007:PCS

[2751] Anonymous. Password crackers see bigger picture. Network Security, 2007 (12):20, December 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580770106X.

Anonymous:2007:NSG

[2752] Anonymous. Network Security goes online. Network Security, 2007 (12):20, December 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485807701071.

Anonymous:2007:ECl

[2753] Anonymous. Events calendar. Network Security, 2007(12):20, December 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485807701083.

Anonymous:2008:MDL

[2754] Anonymous. More data losses in UK government. Network Security, 2008(1):1, January 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700011.

Anonymous:2008:BTC

[2755] Anonymous. Banking Trojan causes account havoc. Network Security, 2008(1):2, January 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700023.

Anonymous:2008:MTI

[2756] Anonymous. Microsoft tests inky password solution. Network Security, 2008(1):2, January 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700035.

Anonymous:2008:Ba

[2757] Anonymous. In brief. Network Security, 2008(1):3, January 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700047.

Munro:2008:SCS

[2758] Ken Munro. SCADA — a critical situation. Network Security, 2008 (1):4–6, January 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700059.

Potter:2008:CJC

[2759] Bruce Potter. Controlling JavaScript with Caja. Network Security, 2008 (1):7–8, January 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700060.

Ullrich:2008:DWS

[2760] Johannes B. Ullrich and Jason Lam. Defacing websites via SQL injection. Network Security, 2008(1): 9–10, January 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700072.

Dunne:2008:ITB

[2761] Nollaig Dunne. ID theft for beginners. Network Security, 2008 (1):10–13, January 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700084.

Small:2008:ME

[2762] Mike Small. Managing entitlements. Network Security, 2008(1): 13–17, January 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700096.

Bridges:2008:CFM

[2763] Lloyd Bridges. The changing face of malware. Network Security, 2008 (1):17–20, January 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700102.

Anonymous:2008:ECa

[2764] Anonymous. Events calendar. Network Security, 2008(1):20, January 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808700114.

Anonymous:2008:MMW

[2765] Anonymous. Most malicious web sites are hacked. Network Security, 2008(2):1–2, February 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700126.

Anonymous:2008:BWA

[2766] Anonymous. Business worried about data loss — but probably not enough. Network Security, 2008 (2):2, February 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700138.

Anonymous:2008:OIG

[2767] Anonymous. Ontario issues guidance on RFID use. Network Security, 2008 (2):2, 20, February 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580870014X.

Anonymous:2008:Bb

[2768] Anonymous. In brief. Network Security, 2008(2):3, February 2008.

CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808700151.

Casey:2008:TLF

[2769] Donal Casey. Turning log files into a security asset. Network Security, 2008(2):4–7, February 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700163.

Ross:2008:EIS

[2770] Steven J. Ross. Enforcing information security: architecture and responsibilities. Network Security, 2008 (2):7–10, February 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700175.

LoIacono:2008:WSL

[2771] Luigi Lo Iacono and Jun Wang. Web service layer security (WSLS). Network Security, 2008(2):10–13, February 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808700187.

Forte:2008:IAS

[2772] Dario Forte. An integrated approach to security incident management. Network Security, 2008(2): 14–16, February 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700199.

Small:2008:EMP

[2773] Mike Small. Entitlement management — a practical approach. Network Security, 2008(2):16–19, February 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808700205.

Anonymous:2008:STS

[2774] Anonymous. Spammers turn to search engines. Network Security, 2008(2):20, February 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700217.

Anonymous:2008:ECb

[2775] Anonymous. Events calendar. Network Security, 2008(2):20, February 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808700229.

Anonymous:2008:SPM

[2776] Anonymous. Symantec prepares for managed security services. Network Security, 2008(3):1–2, March 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808700230.

Anonymous:2008:MDB

[2777] Anonymous. Microsoft downplays BitLocker claim. Network Security, 2008(3):2, March 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700242.

Anonymous:2008:GUD

[2778] Anonymous. Google unveils driveby download details. Network Security, 2008(3):2, March 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700254.

Anonymous:2008:PEC

[2779] Anonymous. Ponemon evaluates cost of UK breaches. Network Security, 2008(3):2, March 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700266.

Anonymous:2008:Bc

[2780] Anonymous. In brief. Network Security, 2008(3):3, March 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700278.

Dye:2008:WAS

[2781] Kerry Dye. Website abuse for search engine optimisation. Network Security, 2008(3):4–6, March 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580870028X.

Corrons:2008:EMR

[2782] Luis Corrons and Dominic Hoskins. Exploring MBR rootkits. Network Security, 2008(3):7–9, March 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700291.

Potter:2008:SRE

[2783] Bruce Potter. Is security really everyone’s responsibility? Network Security, 2008(3):9–10, March 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700308.

Fyffe:2008:AIT

[2784] George Fyffe. Addressing the insider threat. Network Security, 2008 (3):11–14, March 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580870031X.

Heron:2008:PMR

[2785] Simon Heron. Parasitic malware: The resurgence of an old threat. Network Security, 2008 (3):15–18, March 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700321.

Forte:2008:SST

[2786] Dario Forte. Selling security to top management. Network Security, 2008 (3):18–20, March 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700333.

Anonymous:2008:ECc

[2787] Anonymous. Events calendar. Network Security, 2008(3):20, March 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485808700345.

Anonymous:2008:SPE

[2788] Anonymous. SEO poisoning epidemic explodes. Network Security, 2008(4):1, April 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580870045X.

Anonymous:2008:ELP

[2789] Anonymous. EU launches PrimeLife project. Network Security, 2008(4): 2, April 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1353485808700461.

Anonymous:2008:ATE

[2790] Anonymous. Attackers target epilepsy site. Network Security, 2008(4):2, April 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485808700473.

Anonymous:2008:ATC

[2791] Anonymous. Adware tops charts in Q1. Network Security, 2008(4):2, April 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485808700485.

Anonymous:2008:Bd

[2792] Anonymous. In brief. Network Security, 2008(4):3, April 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700497.

Miller:2008:VWR

[2793] Charlie Miller. Virtual worlds, real exploits. Network Security, 2008(4):4–6, April 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700503.

Smith:2008:CFI

[2794] Don Smith. The challenge of federated identity management. Network Security, 2008(4):7–9, April 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700515.

Forte:2008:PAA

[2795] Dario V. Forte. The PTK: An alternative advanced interface for Sleuth Kit. Network Security, 2008 (4):10–13, April 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700527.

Howard:2008:MWA

[2796] Fraser Howard. Modern web attacks. Network Security, 2008 (4):13–15, April 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700539.

Fyffe:2008:MMC

[2797] George Fyffe. Managing multinational compliance efforts while addressing corporate security needs. Network Security, 2008(4):16–18, April 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808700540.

Potter:2008:SMA

[2798] Bruce Potter. Scalable malware analysis. Network Security, 2008 (4):18–20, April 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700552.

Anonymous:2008:ECd

[2799] Anonymous. Events calendar. Network Security, 2008(4):20, April 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808700564.

Anonymous:2008:SAS

[2800] Anonymous. SQL attackers strike again. Network Security, 2008(5): 1, May 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485808700576.

Anonymous:2008:RCB

[2801] Anonymous. Researchers crack bot net secrets. Network Security, 2008(5):2, May 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700588.

Anonymous:2008:BMW

[2802] Anonymous. BERR: More work to do on security. Network Security, 2008(5):2, May 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580870059X.

Anonymous:2008:Be

[2803] Anonymous. In brief. Network Security, 2008(5):3, May 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700606.

Sunner:2008:DWD

[2804] Mark Sunner. Developing world, developing problems. Network Security, 2008(5):4–7, May 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700618.

Colley:2008:MBC

[2805] John Colley. Managing both careers and risks. Network Security, 2008(5):7–9, May 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580870062X.

Roemer:2008:TET

[2806] Kurt Roemer. Treating employees as a threat. Network Security, 2008(5):9–11, May 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700631.

Arce:2008:VMCa

[2807] Iván Arce. Vulnerability management at the crossroads. Network Security, 2008(5):11–13, May 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700643.

Urzay:2008:CIA

[2808] Iñaki Urzay. Collective intelligence approaches to malware recognition. Network Security, 2008 (5):14–16, May 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700655.

Forte:2008:SSR

[2809] Dario V. Forte. Security for safety in railways. Network Security, 2008 (5):17–19, May 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700667.

Anonymous:2008:SSU

[2810] Anonymous. Symantec: Spammers using AdWords. Network Security, 2008(5):20, May 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700679.

Anonymous:2008:MSI

[2811] Anonymous. Mozilla serves up infected files. Network Security, 2008(5):20, May 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700680.

Anonymous:2008:ECe

[2812] Anonymous. Events calendar. Network Security, 2008(5):20, May 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808700692.

Anonymous:2008:OSS

[2813] Anonymous. Open source software getting better. Network Security, 2008(6):1–2, June 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700709.

Anonymous:2008:UPB

[2814] Anonymous. UK plans Big Brother database. Network Security, 2008(6): 2, June 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485808700710.

Anonymous:2008:RDC

[2815] Anonymous. Researcher demonstrates Cisco rootkit. Network Security, 2008(6):2, June 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700722.

Anonymous:2008:Bf

[2816] Anonymous. In brief. Network Security, 2008(6):3, June 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700734.

Munro:2008:BB

[2817] Ken Munro. Breaking into Bluetooth. Network Security, 2008 (6):4–6, June 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700746.

Frankland:2008:ISM

[2818] Jane Frankland. IT security metrics: implementation and standards compliance. Network Security, 2008(6):6–9, June 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700758.

Arce:2008:VMCb

[2819] Ivan Arce. Vulnerability management at the crossroads, part 2. Network Security, 2008(6):9–12, June 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580870076X.

Forte:2008:VDV

[2820] Dario V. Forte. Volatile data vs. data at rest: the requirements of digital forensics. Network Security, 2008(6):13–15, June 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700771.

Hardjono:2008:SEA

[2821] Thomas Hardjono. Strengthening enterprise applications using trusted platform modules. Network Security, 2008(6):15–18, June 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700783.

Potter:2008:PGI

[2822] Bruce Potter. Proxies: a good idea whose time has passed. Network Security, 2008(6):19–20, June 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808700795.

Anonymous:2008:ECf

[2823] Anonymous. Events calendar. Network Security, 2008(6):20, June 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808700801.

Anonymous:2008:KAD

[2824] Anonymous. Kaspersky asks for decryption help. Network Security, 2008(7):1–2, July 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700813.

Anonymous:2008:VEB

[2825] Anonymous. Verizon: External breaches more common. Network Security, 2008(7):2, July 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700825.

Anonymous:2008:RES

[2826] Anonymous. Researchers encode secret messages in VoIP calls. Network Security, 2008(7):2, July 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700837.

Anonymous:2008:Bg

[2827] Anonymous. In brief. Network Security, 2008(7):3, July 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700849.

DeMarines:2008:OHD

[2828] Victor DeMarines. Obfuscation — how to do it and how to crack it. Network Security, 2008(7):4–7, July 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808700850.

Nazario:2008:DAE

[2829] Jose Nazario. DDoS attack evolution. Network Security, 2008 (7):7–10, July 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700862.

Rowan:2008:SVC

[2830] Tom Rowan. The secure virtual computer on your keychain. Network Security, 2008(7):11–14, July 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808700874.

Ray:2008:SS

[2831] Ed Ray. It’s the software, stupid.

Network Security, 2008(7):14–17, July 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808700886.

Shaul:2008:IDS

[2832] Josh Shaul. Implementing database security: using attack analysis to improve your defences. Network Security, 2008(7):18–20, July 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700898.

Anonymous:2008:ECg

[2833] Anonymous. Events calendar. Network Security, 2008(7):20, July 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808700904.

Anonymous:2008:DFR

[2834] Anonymous. DNS flaw rocks web. Network Security, 2008 (8):1–2, August 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700916.

Anonymous:2008:CWH

[2835] Anonymous. City worker holds San Francisco to ransom. Network Security, 2008(8):2, August 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700928.

Anonymous:2008:OCN

[2836] Anonymous. Online crime networks mimic mafia. Network Security, 2008(8):2, August 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580870093X.

Anonymous:2008:Bh

[2837] Anonymous. In brief. Network Security, 2008(8):3, August 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700941.

Takanen:2008:FM

[2838] Ari Takanen. Fuzzing for the masses. Network Security, 2008 (8):4–6, August 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700953.

Lovet:2008:CSF

[2839] Guillaume Lovet. Cybercrime is in a state of flux. Network Security, 2008(8):6–8, August 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808700965.

Naumann:2008:PFE

[2840] Ingo Naumann and Giles Hogben. Privacy features of European eID card specifications. Network Security, 2008 (8):9–13, August 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700977.

Kim:2008:SSS

[2841] Gene Kim. Seven steps to a secure virtual environment. Network Security, 2008(8):14–18, August 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700989.

Forte:2008:SMI

[2842] Dario Forte. 2008: Six months of information security. Network Security, 2008(8):18–20, August 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808700990.

Anonymous:2008:ECh

[2843] Anonymous. Events calendar. Network Security, 2008(8):20, August 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808701004.

Anonymous:2008:HOP

[2844] Anonymous. Home Office: Pass go, do not collect £200. Network Security, 2008(9):1–2, September 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808701016.

Anonymous:2008:BHH

[2845] Anonymous. Black hats hit Red Hat. Network Security, 2008(9): 2, September 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808701028.

Anonymous:2008:GFG

[2846] Anonymous. Google firms up Gmail security. Network Security, 2008 (9):2, September 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580870103X.

Anonymous:2008:Bi

[2847] Anonymous. In brief. Network Security, 2008(9):3, September 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808701041.

Kaminsky:2008:CK

[2848] Dan Kaminsky. Catching up with Kaminsky. Network Security, 2008 (9):4–7, September 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808701053.

Mansfield-Devine:2008:XIT

[2849] Steve Mansfield-Devine. OS X — is it time to start worrying? Network Security, 2008(9):

7–9, September 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808701065.

Shaikh:2008:CIDa

[2850] Siraj A. Shaikh, Howard Chivers, Philip Nobles, John A. Clark, and Hao Chen. Characterising intrusion detection sensors. Network Security, 2008 (9):10–12, September 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808701077.

Potter:2008:DP

[2851] Bruce Potter. Document protection. Network Security, 2008(9): 13–14, September 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808701089.

May:2008:AUE

[2852] Clifford May. Approaches to user education. Network Security, 2008 (9):15–17, September 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808701090.

Forte:2008:DSS

[2853] Dario Forte. Do storage and security go arm in arm? Network Security, 2008 (9):18–19, September 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808701107.

Anonymous:2008:CRT

[2854] Anonymous. Customer records turn up on auctioned machine. Network Security, 2008(9):20, September 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485808701119.

Anonymous:2008:WSF

[2855] Anonymous. Web security flaws up. Network Security, 2008(9): 20, September 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808701120.

Anonymous:2008:ECi

[2856] Anonymous. Events calendar. Network Security, 2008(9):20, September 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485808701132.

Anonymous:2008:MSS

[2857] Anonymous. McAfee snaps up Secure Computing. Network Security, 2008(10):1–2, October 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808701144.

Anonymous:2008:SEU

[2858] Anonymous. Security experts: US trailing on cybersecurity. Network Security, 2008(10):2, October 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485808701156.

Anonymous:2008:Bj

[2859] Anonymous. In brief. Network Security, 2008(10):3, October 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808701168.

Hartley:2008:DWS

[2860] Dave Hartley. Defending Windows servers. Network Security, 2008 (10):4–8, October 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580870117X.

Shaikh:2008:CIDb

[2861] Siraj A. Shaikh, Howard Chivers, Philip Nobles, John A. Clark, and Hao Chen. Characterising intrusion detection sensors, part 2. Network Security, 2008(10):8–11, October 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808701181.

Sood:2008:HKB

[2862] Aditya K. Sood. Hiding a knife behind a smile — OBS hacking threats. Network Security, 2008 (10):12–16, October 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808701193.

Forte:2008:OBK

[2863] Dario Forte. An overview of the best known virtual solutions. Network Security, 2008(10):16–18, October 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580870120X.

Potter:2008:HBI

[2864] Bruce Potter. How bad is it? Network Security, 2008(10):18–20, October 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485808701211.

Anonymous:2008:ECj

[2865] Anonymous. Events calendar. Network Security, 2008(10):20, October 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485808701223.

Anonymous:2008:MIE

[2866] Anonymous. Microsoft issues emergency patch for RPC flaw. Network Security, 2008(11):1–2, November 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485808701235.

Anonymous:2008:CPM

[2867] Anonymous. CSI promotes more sophisticated security measures. Network Security, 2008(11):2, November 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485808701247.

Anonymous:2008:ALS

[2868] Anonymous. Alcatel Lucent ships laptop protection system. Network Security, 2008(11):2, November 2008.

CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808701259.

Anonymous:2008:Bk

[2869] Anonymous. In brief. Network Security, 2008(11):3, November 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808701260.

Mansfield-Devine:2008:ASN

[2870] Steve Mansfield-Devine. Anti-social networking: exploiting the trusting environment of Web 2.0. Network Security, 2008(11):4–7, November 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808701272.

Gebhardt:2008:PHB

[2871] Carlo Gebhardt, Chris I. Dalton, and Richard Brown. Preventing hypervisorbased rootkits with trusted execution technology. Network Security, 2008 (11):7–12, November 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808701284.

Shaikh:2008:NR

[2872] Siraj A. Shaikh, Howard Chivers, Philip Nobles, John A. Clark, and Hao Chen. Network reconnaissance. Network Security, 2008(11): 12–16, November 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485808701296.

Forte:2008:TAI

[2873] Dario Forte. Technological alternatives in incident response. Network Security, 2008(11):16–18, November 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485808701302.

Potter:2008:ES

[2874] Bruce Potter. An everyman SIM. Network Security, 2008(11):18–20, November 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1353485808701314.

Anonymous:2008:ECk

[2875] Anonymous. Events calendar. Network Security, 2008(11):20, November 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485808701326.

Anonymous:2008:MAO

[2876] Anonymous. Microsoft axes OneCare. Network Security, 2008(12):1–2, December 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1353485808701338.

Anonymous:2008:MM

[2877] Anonymous. Marshal and 8e6 merge. Network Security, 2008 (12):2, December 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580870134X.

Anonymous:2008:NMS

[2878] Anonymous. NetWitness makes software tool free. Network Security, 2008(12):2, December 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808701351.

Anonymous:2008:BMR

[2879] Anonymous. BNP members revealed by list leak. Network Security, 2008 (12):2, December 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808701363.

Anonymous:2008:Bl

[2880] Anonymous. In brief. Network Security, 2008(12):3, December 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808701375.

Munro:2008:DE

[2881] Ken Munro. Desktop encryption. Network Security, 2008(12): 4–6, December 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808701387.

Libeau:2008:ASE

[2882] Fabian Libeau. Automating security events management. Network Security, 2008(12):6–9, December 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808701399.

Mansfield-Devine:2008:DC

[2883] Steve Mansfield-Devine. Danger in the clouds. Network Security, 2008 (12):9–11, December 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808701405.

Shaikh:2008:FPR

[2884] Siraj A. Shaikh, Howard Chivers, Philip Nobles, John A. Clark, and Hao Chen. False positive response. Network Security, 2008(12): 11–15, December 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808701417.

Danahy:2008:PSG

[2885] Jack Danahy. The ‘phasing-in’ of security governance in the SDLC. Network Security, 2008(12):15–17, December 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808701429.

Forte:2008:DFS

[2886] Dario Forte. Dealing with forensic software vulnerabilities: is anti-forensics a real danger? Network Security, 2008 (12):18–20, December 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485808701430.

Anonymous:2008:ECl

[2887] Anonymous. Events calendar. Network Security, 2008(12):20, December 2008. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485808701442.

Anonymous:2009:DWT

[2888] Anonymous. Downadup worm takes net by storm. Network Security, 2009(1):1, January 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700017.

Anonymous:2009:NBP

[2889] Anonymous. New botnets on the prowl. Network Security, 2009 (1):2, January 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700029.

Anonymous:2009:PCA

[2890] Anonymous. Proof of concept attack further discredits MD5. Network Security, 2009(1):2, January 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700030.

Anonymous:2009:Ba

[2891] Anonymous. In brief. Network Security, 2009(1):3, January 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700042.

Erasmus:2009:AMA

[2892] Jacques Erasmus. Anatomy of a malware attack. Network Security, 2009(1):4–7, January 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700054.

Schipka:2009:DD

[2893] Maksym Schipka. Dollars for downloading. Network Security, 2009 (1):7–11, January 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700066.

Heron:2009:TSD

[2894] Simon Heron. Technologies for spam detection. Network Security, 2009 (1):11–15, January 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700078.

Potter:2009:MST

[2895] Bruce Potter. Microsoft SDL Threat Modelling Tool. Network Security, 2009(1):15–18, January 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580970008X.

Forte:2009:WGG

[2896] Dario Forte. So are we going to govern this security business? Network Security, 2009 (1):18–20, January 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700091.

Anonymous:2009:ECa

[2897] Anonymous. Events calendar. Network Security, 2009(1):20, January 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485809700108.

Anonymous:2009:COT

[2898] Anonymous. Cybersecurity: Obama takes care of business. Network Security, 2009(2):1–2, February 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580970011X.

Anonymous:2009:KCU

[2899] Anonymous. Kyrgyzstan comes under cyberattack. Network Security, 2009(2):2, February 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700121.

Anonymous:2009:CBR

[2900] Anonymous. Cost of breaches rises. Network Security, 2009 (2):2, February 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700133.

Anonymous:2009:Bb

[2901] Anonymous. In brief. Network Security, 2009(2):3, February 2009.

CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485809700145.

Rowan:2009:PPN

[2902] Tom Rowan. Password protection: the next generation. Network Security, 2009(2):4–7, February 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700157.

Broom:2009:PIE

[2903] Alastair Broom. De-perimeterisation: inevitable and essential. Network Security, 2009(2):7–10, February 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485809700169.

Biro:2009:WW

[2904] Tamas Biro. A window to the world? Network Security, 2009 (2):11–13, February 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700170.

Sood:2009:VPW

[2905] Aditya K. Sood. From vulnerability to patch: the window of exposure. Network Security, 2009(2): 14–16, February 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700182.

Potter:2009:TTY

[2906] Bruce Potter. Three tips for your network. Network Security, 2009 (2):16–18, February 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700194.

Forte:2009:DM

[2907] Dario Forte. The death of MD5. Network Security, 2009(2):18–20, February 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485809700200.

Anonymous:2009:ECb

[2908] Anonymous. Events calendar. Network Security, 2009(2):20, February 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485809700212.

Anonymous:2009:BBI

[2909] Anonymous. BBC botnet investigation causes controversy. Network Security, 2009(3):1–2, March 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700224.

Anonymous:2009:APO

[2910] Anonymous. Anti-phishing organisation launches common reporting standard. Network Security, 2009 (3):2, 20, March 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700236.

Anonymous:2009:Bc

[2911] Anonymous. In brief. Network Security, 2009(3):3, March 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700248.

Mansfield-Devine:2009:GH

[2912] Steve Mansfield-Devine. Google hacking 101. Network Security, 2009 (3):4–6, March 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580970025X.

Kamluk:2009:HLY

[2913] Vitaly Kamluk. Here’s looking at you Kido. Network Security, 2009 (3):6–8, March 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700261.

Eisen:2009:SPK

[2914] Ori Eisen. In-session phishing and knowing your enemy. Network Security, 2009(3):8–11, March 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700273.

Sood:2009:YSP

[2915] Aditya K. Sood. Is your system pwned? Network Security, 2009 (3):11–13, March 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700285.

Southern:2009:CRM

[2916] Steve Southern. Creating risk management strategies for IT security. Network Security, 2009 (3):13–16, March 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700297.

Forte:2009:SAM

[2917] Dario Forte. Security audits in mixed environments. Network Security, 2009 (3):17–19, March 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700303.

Anonymous:2009:MBS

[2918] Anonymous. McAfee: Beware the spam of March. Network Security, 2009(3):20, March 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700315.

Anonymous:2009:MDN

[2919] Anonymous. ‘Monroe Doctrine’ needed for cyberspace. Network Security, 2009(3):20, March 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700327.

Anonymous:2009:ECc

[2920] Anonymous. Events calendar. Network Security, 2009(3):20, March 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485809700339.

Anonymous:2009:CEO

[2921] Anonymous. Cyber-espionage operation discovered. Network Security, 2009(4):1–2, April 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700340.

Anonymous:2009:RCG

[2922] Anonymous. Rogueware continues to grow. Network Security, 2009 (4):2, 20, April 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700352.

Anonymous:2009:Bd

[2923] Anonymous. In brief. Network Security, 2009(4):3, April 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700364.

Jenik:2009:CEM

[2924] Aviram Jenik. Cyberwar in Estonia and the Middle East. Network Security, 2009(4):4–6, April 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700376.

Mansfield-Devine:2009:SWF

[2925] Steve Mansfield-Devine. Simple website footprinting. Network Security, 2009(4):7–9, April 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700388.

Dwyer:2009:CCA

[2926] Dennis Dwyer. Chinese cyber-attack tools continue to evolve. Network Security, 2009(4):9–11, April 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580970039X.

Heron:2009:GP

[2927] Simon Heron. A good policy. Network Security, 2009(4):12–14, April 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485809700406.

Mattsson:2009:RSV

[2928] Ulf Mattsson. Real security for virtual machines. Network Security, 2009(4):15–17, April 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700418.

Forte:2009:APA

[2929] Dario Forte. Anatomy of a phishing attack: a high-level overview. Network Security, 2009(4):17–19, April 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580970042X.

Anonymous:2009:RBL

[2930] Anonymous. Researcher blows lid off retail networks. Network Security, 2009(4):20, April 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700431.

Anonymous:2009:UEG

[2931] Anonymous. US electrical grid compromised. Network Security, 2009(4):20, April 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700443.

Anonymous:2009:ECd

[2932] Anonymous. Events calendar. Network Security, 2009(4):20, April 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485809700455.

Anonymous:2009:SGL

[2933] Anonymous. Security groups launch anti-malware effort. Network Security, 2009(5):1–2, May 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700467.

Anonymous:2009:GGA

[2934] Anonymous. Gumblar Google attack takes web by storm. Network Security, 2009(5):2, May 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700479.

Anonymous:2009:CFB

[2935] Anonymous. Conficker fails, but bots increase. Network Security, 2009(5):2, May 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700480.

Anonymous:2009:CSC

[2936] Anonymous. Cyber-swine cash in on real-world flu. Network Security, 2009(5):2, May 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700492.

Anonymous:2009:Be

[2937] Anonymous. In brief. Network Security, 2009(5):3, May 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700509.

Jenik:2009:FTM

[2938] Aviram Jenik. Fuzzing tools: making sense out of nonsense. Network Security, 2009(5):4–6, May 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700510.

Rodriguez:2009:CCT

[2939] Bruno Rodriguez. The cyber-crime threat to online transactions. Network Security, 2009(5):7–8, May 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485809700522.

Khan:2009:RTC

[2940] Shafiullah Khan and Kok-Keong Loo. Real-time cross-layer design for a largescale flood detection and attack traceback mechanism in IEEE 802.11 wireless mesh networks. Network Security, 2009(5):9–16, May 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700534.

Walters:2009:JD

[2941] Richard Walters. Joining the dots. Network Security, 2009 (5):16–19, May 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700546.

Forte:2009:PD

[2942] Dario Forte. Phishing in depth. Network Security, 2009(5):19–20, May 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485809700558.

Anonymous:2009:ECe

[2943] Anonymous. Events calendar. Network Security, 2009(5):20, May 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348580970056X.

Anonymous:2009:CPT

[2944] Anonymous. Chinese protection tool pegged as badware. Network Security, 2009(6):1–2, June 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700571.

Anonymous:2009:FUH

[2945] Anonymous. Feds unveil Hathaway review. Network Security, 2009(6):2, June 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700583.

Anonymous:2009:RPS

[2946] Anonymous. Researcher promises smart grid botnet. Network Security, 2009(6):2, 20, June 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700595.

Anonymous:2009:Bf

[2947] Anonymous. In brief. Network Security, 2009(6):3, June 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700601.

Heron:2009:OPB

[2948] Simon Heron. Online privacy and browser security. Network Security, 2009(6):4–7, June 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700613.

Mansfield-Devine:2009:MT

[2949] Steve Mansfield-Devine. A matter of trust. Network Security, 2009(6):7–9, June 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700625.

Gold:2009:SM

[2950] Steve Gold. Not-so-smart meters? Network Security, 2009 (6):9–11, June 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700637.

Shaikh:2009:TSI

[2951] Siraj A. Shaikh, Howard Chivers, Philip Nobles, John A. Clark, and Hao Chen. Towards scalable intrusion detection. Network Security, 2009(6):12–16, June 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700649.

Rexworthy:2009:IDS

[2952] Ben Rexworthy. Intrusion detections systems — an outmoded network protection model. Network Security, 2009(6):17–19, June 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700650.

Anonymous:2009:BVP

[2953] Anonymous. Browser vendors in patch blitz. Network Security, 2009(6):20, June 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700662.

Anonymous:2009:RBB

[2954] Anonymous. Researchers build browser-based darknet. Network Security, 2009(6):20, June 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700674.

Anonymous:2009:ECf

[2955] Anonymous. Events calendar. Network Security, 2009(6):20, June 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485809700686.

Anonymous:2009:RDE

[2956] Anonymous. In a recession, disgruntled employees could pose a security threat from within. Network Security, 2009(7):1, July 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700807.

Anonymous:2009:FPN

[2957] Anonymous. Finger pointed at North Korea after denial of service attacks on the South and the US. Network Security, 2009(7):1–2, July 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700819.

Anonymous:2009:NCV

[2958] Anonymous. NHS computer viruses impact on patient healthcare. Network Security, 2009(7):2, July 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485809700820.

Anonymous:2009:BDP

[2959] Anonymous. Bonn discovers partial solution to Conficker infections. Network Security, 2009(7):2, July 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485809700832.

Anonymous:2009:Bg

[2960] Anonymous. In brief. Network Security, 2009(7):3, July 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700844.

Mansfield-Devine:2009:PW

[2961] Steve Mansfield-Devine. The promise of whitelisting. Network Security, 2009(7):4–6, July 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700856.

Forte:2009:YCV

[2962] Dario Forte. Are you court validated? Network Security, 2009(7):6–8, July 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700868.

Caviglione:2009:UER

[2963] Luca Caviglione. Understanding and exploiting the reverse patterns of peer-to-peer file sharing applications. Network Security, 2009 (7):8–12, July 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580970087X.

Goucher:2009:EW

[2964] Wendy Goucher. The enemy within. Network Security, 2009 (7):12–14, July 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700881.

Potter:2009:RL

[2965] Bruce Potter. A review of L0phtCrack 6. Network Security, 2009(7):14– 17, July 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485809700893.

Forte:2009:ILF

[2966] Dario Forte. The importance of log files in security incident prevention. Network Security, 2009 (7):18–20, July 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580970090X.

Anonymous:2009:ECg

[2967] Anonymous. Events calendar. Network Security, 2009(7):20, July 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485809700911.

Anonymous:2009:CAI

[2968] Anonymous. Cyber attacks increasingly target wealthy smartphone owners, says Cisco. Network Security, 2009(8):1, August 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700698.

Anonymous:2009:FPV

[2969] Anonymous. Flaws and patching in Las Vegas. Network Security, 2009(8):1–2, August 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700704.

Anonymous:2009:RNM

[2970] Anonymous. Rogueware now a major menace. Network Security, 2009(8):2, August 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700716.

Anonymous:2009:Bh

[2971] Anonymous. In brief. Network Security, 2009(8):3, August 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700728.

Forte:2009:RIS

[2972] Dario V. Forte. The role of the Information Security Manager in cuttingedge companies. Network Security, 2009(8):4–5, August 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348580970073X.

Mortleman:2009:BA

[2973] Jim Mortleman. Behavioural analysis.

Network Security, 2009(8):5–7, August 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485809700741.

Barry:2009:SPS

[2974] Daniel Joseph Barry. Special problems of securing 10Gbps networks. Network Security, 2009 (8):8–11, August 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700753.

Frankland:2009:NDI

[2975] Jane Frankland. Numeric data integrity: piercing the corporate veil. Network Security, 2009(8): 11–14, August 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700765.

Sloan:2009:SVW

[2976] Kevin Sloan. Security in a virtualised world. Network Security, 2009 (8):15–18, August 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700777.

Gold:2009:SCS

[2977] Steve Gold. The SCADA challenge: securing critical infrastructure. Network Security, 2009(8): 18–20, August 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700789.

Anonymous:2009:ECh

[2978] Anonymous. Events calendar. Network Security, 2009(8):20, August 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485809700790.

Anonymous:2009:FFB

[2979] Anonymous. Failed firm banned from selling customers’ personal data. Network Security, 2009 (9):1, September 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700923.

Anonymous:2009:MWI

[2980] Anonymous. Massive website infection avoids detection by search engine security scanners. Network Security, 2009(9):1, September 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700935.

Anonymous:2009:ITM

[2981] Anonymous. ID theft malware on the increase. Network Security, 2009 (9):1–2, September 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700947.

Anonymous:2009:ULD

[2982] Anonymous. US lab develops federated model for defence against cyber attack. Network Security, 2009 (9):2, September 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700959.

Anonymous:2009:Bi

[2983] Anonymous. In brief. Network Security, 2009(9):3, September 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485809700960.

Abraham:2009:WC

[2984] Dave Abraham. Why 2FA in the cloud? Network Security, 2009 (9):4–5, September 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700972.

Garlick:2009:HBO

[2985] Nick Garlick and Nebulas Solutions Group. The hidden benefits of optimising your firewall. Network Security, 2009(9):6–9, September 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809700984.

Everett:2009:MAD

[2986] Cath Everett. Moving across to the dark side. Network Security, 2009 (9):10–12, September 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809700996.

Sridevi:2009:NAE

[2987] R. Sridevi, A. Damodaram, and S. V. L. Narasimham. A novel architecture for enhanced security through virtualisation and spoofing. Network Security, 2009(9): 12–16, September 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348580970100X.

Forte:2009:CVB

[2988] Dario Forte. Compliance vs business security. Network Security, 2009 (9):16–18, September 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809701011.

Goucher:2009:CSS

[2989] Wendy Goucher. Carpe sus: security in the time of swine flu. Network Security, 2009(9):18–20, September 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485809701023.

Anonymous:2009:ECi

[2990] Anonymous. Events calendar. Network Security, 2009(9):20, September 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485809701035.

Anonymous:2009:WRL

[2991] Anonymous. Wireless researchers learn to see through walls. Network Security, 2009(10):1–2, October 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809701047.

Anonymous:2009:DNF

[2992] Anonymous. DHS needs to fix web site security. Network Security, 2009(10):2, October 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809701059.

Anonymous:2009:SBM

[2993] Anonymous. Spam-borne malware is on the rise. Network Security, 2009 (10):2, 19, October 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809701060.

Anonymous:2009:Bj

[2994] Anonymous. In brief. Network Security, 2009(10):3, October 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485809701072.

Lowe:2009:SPB

[2995] Nick Lowe.  Shields up!  Pro-

tecting browsers, endpoints and enterprises against web-based attacks. Network Security, 2009 (10):4–7, October 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809701084.

Walsh:2009:BFC

[2996] Patrick J. Walsh. The brightening future of cloud security. Network Security, 2009(10):7–10, October 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485809701096.

Storey:2009:SPC

[2997] Dominic Storey. Securing process control networks. Network Security, 2009 (10):10–13, October 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809701102.

Agrawal:2009:MVO

[2998] A. Agrawal and R. A. Khan. Measuring the vulnerability of an objectoriented design. Network Security, 2009 (10):13–17, October 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809701114.

Forte:2009:SVR

[2999] Dario Forte. SSL VPN and return on investment: a possible combination. Network Security, 2009 (10):17–19, October 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809701126.

Anonymous:2009:AMS

[3000] Anonymous. Advertising malware scammers get smarter. Network Security, 2009(10):19–20, October 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485809701138.

Anonymous:2009:ELD

[3001] Anonymous. Europe leaves door open for three-strikes policy. Network Security, 2009(10):20, October 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348580970114X.

Anonymous:2009:ECj

[3002] Anonymous. Events calendar. Network Security, 2009(10):20, October 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485809701151.

Anonymous:2009:ADF

[3003] Anonymous. Adobe design flaw alleged. Network Security, 2009 (11):1–2, November 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809701163.

Anonymous:2009:IEF

[3004] Anonymous. IPS equipment fails to make grade. Network Security, 2009 (11):2, November 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809701175.

Anonymous:2009:WAV

[3005] Anonymous. Web applications vulnerabilities severe, says Cenzic. Network Security, 2009(11): 2, 20, November 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809701187.

Anonymous:2009:Bk

[3006] Anonymous. In brief. Network Security, 2009(11):3, November 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485809701199.

Iyer:2009:NAS

[3007] Sudha Iyer. New approaches to securing the database. Network Security, 2009(11):4–8, November 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809701205.

Jones:2009:RMT

[3008] Andy Jones. Recycling more than your IT equipment. Network Security, 2009 (11):8–9, November 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485809701217.

Xirasagar:2009:SIN

[3009] Sindhu Xirasagar and Masoud Mojtahed. Securing IP networks, Part I. Network Security, 2009(11): 10–14, November 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809701229.

Storey:2009:CFH

[3010] Dominic Storey. Catching flies with honey tokens. Network Security, 2009 (11):15–18, November 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809701230.

Forte:2009:IRI

[3011] Dario Forte. Is it really important to have a structured security budget? Network Security, 2009(11): 18–20, November 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485809701242.

Anonymous:2009:ECk

[3012] Anonymous. Events calendar. Network Security, 2009(11):20, November 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485809701254.

Anonymous:2009:WCT

[3013] Anonymous. WPA cracking tool launched. Network Security, 2009 (12):1–2, December 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700015.

Anonymous:2009:GLD

[3014] Anonymous. Google launches DNS server. Network Security, 2009 (12):2, December 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700027.

Anonymous:2009:MRD

[3015] Anonymous. Malware rebounds as data loss cause. Network Security, 2009 (12):2, 20, December 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700039.

Anonymous:2009:Bl

[3016] Anonymous. In brief. Network Security, 2009(12):3, December 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485810700040.

Lin:2009:AMD

[3017] Phillip Lin. Anatomy of the Mega-D takedown. Network Security, 2009 (12):4–7, December 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700052.

Heron:2009:AES

[3018] Simon Heron. Advanced Encryption Standard (AES). Network Security, 2009(12):8–12, December 2009.

CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810700064.

Yang:2009:CGA

[3019] Seung S. Yang and Hongsik Choi. A complement to the GridOne authentication method. Network Security, 2009 (12):12–18, December 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700076.

Forte:2009:ADP

[3020] Dario Forte. Application delivery: pros and cons both virtual and real. Network Security, 2009(12): 18–20, December 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700088.

Anonymous:2009:ECl

[3021] Anonymous. Events calendar. Network Security, 2009(12):20, December 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581070009X.

Anonymous:2010:AHS

[3022] Anonymous. Aurora hack spawns widespread concern. Network Security, 2010(1):1–2, January 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700106.

Anonymous:2010:GCS

[3023] Anonymous. Google challenged on search privacy. Network Security, 2010 (1):2, 20, January 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700118.

Anonymous:2010:Ba

[3024] Anonymous. In brief. Network Security, 2010(1):3, January 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581070012X.

Fergusonnam:2010:BF

[3025] Rik Fergusonnam. Back to the future. Network Security, 2010 (1):4–7, January 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700131.

Furnell:2010:MSA

[3026] Steven Furnell. Mac security: An Apple that can’t be bitten? Network Security, 2010(1):7–11, January 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485810700143.

Mansfield-Devine:2010:PS

[3027] Steve Mansfield-Devine. The perils of sharing. Network Security, 2010 (1):11–13, January 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700155.

Xirasagar:2010:SIN

[3028] Sindhu Xirasagar and Masoud Mojtahed. Securing IP networks. Network Security, 2010(1):13–17, January 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810700167.

Forte:2010:RIR

[3029] Dario V. Forte. The responsibilities of an incident responder. Network Security, 2010(1): 18–19, January 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700179.

Anonymous:2010:OCR

[3030] Anonymous. Online criminals replacing brawn with brain. Network Security, 2010(1):20, January 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700180.

Anonymous:2010:ECa

[3031] Anonymous. Events calendar. Network Security, 2010(1):20, January 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810700192.

Anonymous:2010:NIE

[3032] Anonymous. New Internet Explorer bug emerges. Network Security, 2010 (2):1–2, February 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700209.

Anonymous:2010:SCB

[3033] Anonymous. SpyEye continues battle of the botnets. Network Security, 2010(2):2, February 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700210.

Anonymous:2010:Bb

[3034] Anonymous. In brief. Network Security, 2010(2):3, February 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810700222.

Brandt:2010:WAA

[3035] Andrew Brandt and Zachary Wolff. When admins attack: 30 hours in the life of a Gumblar victim. Network Security, 2010(2): 4–8, February 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700234.

Rowan:2010:NWS

[3036] Tom Rowan. Negotiating WiFi security. Network Security, 2010 (2):8–12, February 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700246.

Goode:2010:MMS

[3037] Alan Goode. Managing mobile security: How are we doing? Network Security, 2010(2):12–15, February 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810700258.

McPherson:2010:CGC

[3038] Danny McPherson. Cybercrime — a game of cat and mouse in 2009. Network Security, 2010(2): 15–18, February 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581070026X.

Forte:2010:PIH

[3039] Dario Forte. Preventing and investigating hacking by auditing web applications. Network Security, 2010 (2):18–20, February 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700271.

Anonymous:2010:ECb

[3040] Anonymous. Events calendar. Network Security, 2010(2):20, February 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810700283.

Anonymous:2010:BBS

[3041] Anonymous. Botherders battle to stay online. Network Security, 2010 (3):1–2, March 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700295.

Anonymous:2010:RIC

[3042] Anonymous. Researcher issues code for yet another IE flaw. Network Security, 2010(3):2, March 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700301.

Anonymous:2010:SUS

[3043] Anonymous. SSD used to speed up password cracking. Network Security, 2010(3):2, March 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700313.

Anonymous:2010:Bc

[3044] Anonymous. In brief. Network Security, 2010(3):3, March 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700325.

Mansfield-Devine:2010:DCT

[3045] Steve Mansfield-Devine. Divide and conquer: the threats posed by hybrid apps and HTML 5. Network Security, 2010(3):4–6, March 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700337.

Wicherski:2010:PLI

[3046] Georg Wicherski. Placing a lowinteraction honeypot in-the-wild: A review of mwcollectd. Network Security, 2010(3):7–8, March 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700349.

Janssen:2010:VS

[3047] Robert Janssen. VDI and security. Network Security, 2010 (3):8–11, March 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700350.

Corrons:2010:RRN

[3048] Luis Corrons. The rise and rise of NDR. Network Security, 2010 (3):12–16, March 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700362.

Chandra:2010:CCO

[3049] S. Chandra and R. A. Khan. Confidentiality checking an object-oriented class hierarchy. Network Security, 2010(3):16–20, March 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700374.

Anonymous:2010:ECc

[3050] Anonymous. Events calendar. Network Security, 2010(3):20, March 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810700386.

Anonymous:2010:PARa

[3051] Anonymous. PDF attacks on the rise, says Symantec. Network Security, 2010(4):1–2, April 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700398.

Anonymous:2010:EMS

[3052] Anonymous. Extortion malware scares file sharers with fake copyright messages. Network Security, 2010(4):2, April 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700404.

Anonymous:2010:PARb

[3053] Anonymous. PAC attack redirects browsers to malicious sites using proxy hack. Network Security, 2010 (4):2, 20, April 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700416.

Anonymous:2010:Bd

[3054] Anonymous. In brief. Network Security, 2010(4):3, April 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700428.

Portnoy:2010:PWA

[3055] Aaron Portnoy. Pwn2Own wrap up and analysis. Network Security, 2010(4):4–5, April 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581070043X.

Wood:2010:BWS

[3056] Paul Wood. Bot wars: the spammers strike back. Network Security, 2010(4):5–8, April 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700441.

Cluley:2010:SMT

[3057] Graham Cluley. Sizing up the malware threat — key malware trends for 2010. Network Security, 2010 (4):8–10, April 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700453.

Eisen:2010:CFM

[3058] Ori Eisen. Catching the fraudulent Man-in-the-Middle and Manin-the-Browser. Network Security, 2010(4):11–12, April 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700465.

Carnie:2010:PSM

[3059] James Carnie. Preserving security in a multi-tenanted hosting environment. Network Security, 2010 (4):12–15, April 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700477.

Vlissidis:2010:TCS

[3060] Paul Vlissidis and Matthew Hickey. Thin clients: slim security? Network Security, 2010(4):16–19, April 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485810700489.

Anonymous:2010:NSF

[3061] Anonymous. Network security fixes WordPress configuration hole. Network Security, 2010(4):20, April 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485810700490.

Anonymous:2010:ECd

[3062] Anonymous. Events calendar. Network Security, 2010(4):20, April 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485810700507.

Anonymous:2010:FUA

[3063] Anonymous. Facebook under attack on all fronts. Network Security, 2010(5):1–2, May 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700519.

Anonymous:2010:GAS

[3064] Anonymous. Google ‘accidentally’ snooped on wifi data. Network Security, 2010(5):2, May 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700520.

Anonymous:2010:Be

[3065] Anonymous. In brief. Network Security, 2010(5):3, May 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700532.

Mansfield-Devine:2010:BB

[3066] Steve Mansfield-Devine. Battle of the botnets. Network Security, 2010(5):4–6, May 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700544.

Steiner:2010:WFN

[3067] Paul Steiner. Why FTP is no longer up to the job. Network Security, 2010(5):7–9, May 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700556.

Heron:2010:DSM

[3068] Simon Heron. Denial of service: motivations and trends. Network Security, 2010(5):10–12, May 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700568.

Davis:2010:MTP

[3069] Adrian Davis. Managing third parties — an information security perspective. Network Security, 2010 (5):13–15, May 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581070057X.

Bradbury:2010:SCC

[3070] Danny Bradbury. Shadows in the cloud: Chinese involvement in advanced persistent threats. Network Security, 2010(5):16–19, May 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810700581.

Anonymous:2010:BUI

[3071] Anonymous. Browsers uniquely identify users. Network Security, 2010(5):19, May 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700593.

Anonymous:2010:IFC

[3072] Anonymous. India fears cyberspying by China. Network Security, 2010(5):19–20, May 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581070060X.

Anonymous:2010:TCD

[3073] Anonymous. Twitter commands DIY botnets. Network Security, 2010(5):20, May 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700611.

Anonymous:2010:VW

[3074] Anonymous. Virtualisation worries.

Network Security, 2010(5):20, May 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810700623.

Anonymous:2010:ECe

[3075] Anonymous. Events calendar. Network Security, 2010(5):20, May 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810700635.

Anonymous:2010:RIS

[3076] Anonymous. Reduce IT security costs without increasing risk. Network Security, 2010(6):1–2, June 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700763.

Anonymous:2010:NVH

[3077] Anonymous. New ventures help developers in fight against security flaws. Network Security, 2010(6):2, June 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810700775.

Anonymous:2010:WFD

[3078] Anonymous. Windows flaw disclosure causes fierce debate. Network Security, 2010(6):2, June 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700787.

Anonymous:2010:Bf

[3079] Anonymous. In brief. Network Security, 2010(6):3, June 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700799.

Karrenberg:2010:DSG

[3080] Daniel Karrenberg. DNSSEC: Securing the global infrastructure of the Internet. Network Security, 2010(6):4–6, June 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700805.

Aljawarneh:2010:SWS

[3081] Shadi Aljawarneh, Shadi Masadeh, and Faisal Alkhateeb. A secure wifi system for wireless networks: an experimental evaluation. Network Security, 2010(6):6–12, June 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700817.

Bradbury:2010:HSW

[3082] Danny Bradbury. A hole in the security wall: ATM hacking. Network Security, 2010(6):12–15, June 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700829.

Hamelin:2010:PFM

[3083] Michael Hamelin. Preventing firewall meltdowns. Network Security, 2010(6):15–16, June 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700830.

Martin:2010:PCC

[3084] Luther Martin. Protecting credit card information: encryption vs tokenisation. Network Security, 2010 (6):17–19, June 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700842.

Anonymous:2010:ALP

[3085] Anonymous. AT&T leaks private data of Apple customers — twice. Network Security, 2010(6):20, June 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810700854.

Anonymous:2010:FSR

[3086] Anonymous. Facebook suffers from rash of clickjacking. Network Security, 2010(6):20, June 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700866.

Anonymous:2010:ECf

[3087] Anonymous. Events calendar. Network Security, 2010(6):20, June 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810700878.

Anonymous:2010:SV

[3088] Anonymous. Smartphone vulnerabilities. Network Security, 2010(7):2, July 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581070088X.

Anonymous:2010:MIS

[3089] Anonymous. Malware-infested spam on the rise. Network Security, 2010(7):2, July 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700891.

Anonymous:2010:Bg

[3090] Anonymous. In brief. Network Security, 2010(7):3, July 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700908.

Joffee:2010:CGE

[3091] Rodney Joffee. Cybercrime: the global epidemic at your network door. Network Security, 2010(7):4–7, July 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581070091X.

Bradbury:2010:HME

[3092] Danny Bradbury. Hands-on with Metasploit Express. Network Security, 2010(7):7–11, July 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700921.

Schiavo:2010:CSE

[3093] Jay Schiavo. Code signing for enduser peace of mind. Network Security, 2010(7):11–13, July 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700933.

Eisen:2010:OSN

[3094] Ori Eisen. Online security — a new strategic approach. Network Security, 2010(7):14–15, July 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700945.

Ward:2010:CSM

[3095] Des Ward. The cost of saving money: no longer the company reputation. Network Security, 2010 (7):15–19, July 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700957.

Anonymous:2010:DND

[3096] Anonymous. Disclosure and nondisclosure. Network Security, 2010 (7):19–20, July 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810700969.

Anonymous:2010:SFS

[3097] Anonymous. Software flaws set to double. Network Security, 2010(7):20, July 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810700970.

Anonymous:2010:TTT

[3098] Anonymous. Twitter told to tighten security. Network Security, 2010(7): 20, July 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1353485810700982.

Anonymous:2010:ECg

[3099] Anonymous. Events calendar. Network Security, 2010(7):20, July 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485810700994.

Anonymous:2010:IBM

[3100] Anonymous. Intel buys McAfee. Network Security, 2010(8):1–2, August 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485810701008.

Anonymous:2010:BDH

[3101] Anonymous. Black Hat and DefCon highlights. Network Security, 2010 (8):2, 20, August 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581070101X.

Anonymous:2010:Bh

[3102] Anonymous. In brief. Network Security, 2010(8):3, August 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701021.

Gold:2010:CP

[3103] Steve Gold. Cracking passwords. Network Security, 2010(8):4–7, August 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810701033.

Storey:2010:TCN

[3104] Dominic Storey. Ten consequences of network blindness. Network Security, 2010(8):7–9, August 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701045.

Kim:2010:SGF

[3105] Seong-Hwan Kim, Sundar Vedantham, and Pravin Pathak. SMB gateway firewall implementation using a network processor. Network Security, 2010(8):10–15, August 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701057.

Smith:2010:CDS

[3106] Graham Smith. Countering datacentre security pressures. Network Security, 2010(8):15–17, August 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701069.

Midgley:2010:SEE

[3107] Stephen Midgley. The state of encryption in Europe: some cultural comparisons. Network Security, 2010 (8):18–19, August 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810701070.

Anonymous:2010:DMC

[3108] Anonymous. Did malware cause Spanair crash? Network Security, 2010(8):20, August 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810701082.

Anonymous:2010:PZ

[3109] Anonymous. Phishing down, Zeus up. Network Security, 2010(8):20, August 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485810701094.

Anonymous:2010:ECh

[3110] Anonymous. Events calendar. Network Security, 2010(8):20, August 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485810701100.

Anonymous:2010:SMW

[3111] Anonymous. Stuxnet may be the work of state-backed hackers. Network Security, 2010(9): 1–2, September 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701112.

Anonymous:2010:WLB

[3112] Anonymous. Wins and losses in the botnet war. Network Security, 2010 (9):2, 19, September 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701124.

Anonymous:2010:Bi

[3113] Anonymous. In brief. Network Security, 2010(9):3, September 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810701136.

Bransfield-Garth:2010:MPC

[3114] Simon Bransfield-Garth. Mobile phone calls as a business risk. Network Security, 2010(9):4–11, September 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810701148.

Taylor:2010:FIP

[3115] Mark Taylor, John Haggerty, David Gresty, and Paul Fergus. Forensic investigation of peer-to-peer networks. Network Security, 2010(9): 12–15, September 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581070115X.

Gold:2010:PA

[3116] Steve Gold. Password alternatives. Network Security, 2010(9): 16–19, September 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701161.

Anonymous:2010:CAF

[3117] Anonymous. Corporate attacks focus on web. Network Security, 2010 (9):19–20, September 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701173.

Anonymous:2010:SPM

[3118] Anonymous. Smartphones present major threat. Network Security, 2010 (9):20, September 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701185.

Anonymous:2010:HPA

[3119] Anonymous. Hacking popular among students. Network Security, 2010 (9):20, September 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701197.

Anonymous:2010:ECi

[3120] Anonymous. Events calendar. Network Security, 2010(9):20, September 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810701203.

Anonymous:2010:SRI

[3121] Anonymous. Stuxnet: rumours increase, infections spread. Network Security, 2010(10):1–2, October 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810701215.

Anonymous:2010:TR

[3122] Anonymous. Threat round-up. Network Security, 2010(10):2, October 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810701227.

Anonymous:2010:Bj

[3123] Anonymous. In brief. Network Security, 2010(10):3, October 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810701239.

Papadaki:2010:VMA

[3124] Maria Papadaki and Steven Furnell.

Vulnerability management: an attitude of mind? Network Security, 2010(10):4–8, October 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701240.

Perche:2010:NLA

[3125] Patrice Perche. Network latency: avoid paying a tax on time. Network Security, 2010(10):9–12, October 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701252.

Zaitsev:2010:SKP

[3126] Oleg Zaitsev. Skeleton keys: the purpose and applications of keyloggers. Network Security, 2010(10): 12–17, October 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810701264.

Knight:2010:IDF

[3127] Eric Knight. Investigating digital fingerprints: advanced log analysis. Network Security, 2010(10): 17–20, October 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810701276.

Anonymous:2010:ECj

[3128] Anonymous. Events calendar. Network Security, 2010(10):20, October 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485810701288.

Anonymous:2010:BBT

[3129] Anonymous. Bredolab botnet taken down. Network Security, 2010(11): 1–2, November 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581070129X.

Anonymous:2010:ZMF

[3130] Anonymous. Zeus makes final appearance? Network Security, 2010 (11):2, November 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701306.

Anonymous:2010:AVM

[3131] Anonymous. Anti-virus for Macs. Network Security, 2010(11):2, November 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810701318.

Anonymous:2010:Bk

[3132] Anonymous. In brief. Network Security, 2010(11):3, November 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581070132X.

Bradbury:2010:AUH

[3133] Danny Bradbury. Avoiding URL hell. Network Security, 2010(11): 4–6, November 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701331.

Walters:2010:MPU

[3134] Richard Walters. Managing privileged user activity in the datacentre. Network Security, 2010(11): 6–10, November 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701343.

Gold:2010:SET

[3135] Steve Gold. Social engineering today: psychology, strategies and tricks. Network Security, 2010(11):

11–14, November 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810701355.

Boroditsky:2010:CCW

[3136] Marc Boroditsky. Click, click, who’s there? Network Security, 2010(11): 14–16, November 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810701367.

Warmack:2010:CTT

[3137] Rob Warmack. Compliance today — and tomorrow. Network Security, 2010 (11):16–18, November 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485810701379.

Cavalancia:2010:PDL

[3138] Nick Cavalancia. Preventing data loss by securing USB ports. Network Security, 2010(11):18–20, November 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485810701380.

Anonymous:2010:ECk

[3139] Anonymous. Events calendar. Network Security, 2010(11):20, November 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485810701392.

Anonymous:2010:PNY

[3140] Anonymous. Predictions for the New Year. Network Security, 2010 (12):1–2, December 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701409.

Anonymous:2010:MUM

[3141] Anonymous. More users means security challenges. Network Security, 2010(12):2, December 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701410.

Anonymous:2010:HPS

[3142] Anonymous. Higher priority for security. Network Security, 2010(12): 2, 20, December 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701422.

Anonymous:2010:Bl

[3143] Anonymous. In brief. Network Security, 2010(12):3, December 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810701434.

Dinoor:2010:PIM

[3144] Shlomi Dinoor. Privileged identity management: securing the enterprise. Network Security, 2010 (12):4–6, December 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701446.

Ridley:2010:OSF

[3145] Philip Ridley. Outsmarting the smartphone fraudsters. Network Security, 2010(12):7–9, December 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701458.

Gold:2010:PCA

[3146] Steve Gold. Protecting the cloud: attack vectors and other exploits. Network Security, 2010(12):10–12, December 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S135348581070146X.

Furnell:2010:UVC

[3147] Steven Furnell. Usability versus complexity — striking the balance in enduser security. Network Security, 2010 (12):13–17, December 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701471.

Ranum:2010:WMT

[3148] Marcus Ranum. The well-meaning threat. Network Security, 2010(12): 17–19, December 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485810701483.

Anonymous:2010:UPP

[3149] Anonymous. US public-private partnership for cyber-security. Network Security, 2010(12):20, December 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810701495.

Anonymous:2010:ECl

[3150] Anonymous. Events calendar. Network Security, 2010(12):20, December 2010. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485810701501.

Anonymous:2011:SPF

[3151] Anonymous. Security professionals focus on yesterday’s skills. Network Security, 2011(1):1–2, January 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700010.

Anonymous:2011:WSC

[3152] Anonymous. Wifi security cracked.

Network Security, 2011(1):2, 20, January 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485811700022.

Anonymous:2011:Ba

[3153] Anonymous. In brief. Network Security, 2011(1):3, January 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700034.

Mansfield-Devine:2011:AST

[3154] Steve Mansfield-Devine. Anonymous: serious threat or mere annoyance? Network Security, 2011 (1):4–10, January 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700046.

Amoroso:2011:CAA

[3155] Edward G. Amoroso. Cyber attacks: awareness. Network Security, 2011 (1):10–16, January 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700058.

Gold:2011:AET

[3156] Steve Gold. Advanced evasion techniques. Network Security, 2011 (1):16–19, January 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581170006X.

Anonymous:2011:PT

[3157] Anonymous. Passwords not up to the task. Network Security, 2011 (1):20, January 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700071.

Anonymous:2011:DLR

[3158] Anonymous. Data leaks result from malicious intent. Network Security, 2011(1):20, January 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700083.

Anonymous:2011:ECa

[3159] Anonymous. Events calendar. Network Security, 2011(1):20, January 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700095.

Anonymous:2011:STC

[3160] Anonymous. Sinister trends in cyberthreats. Network Security, 2011 (2):2, February 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700101.

Anonymous:2011:EGR

[3161] Anonymous. European guidelines for resilience and cyber-security. Network Security, 2011(2):2, 19–20, February 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700113.

Anonymous:2011:Bb

[3162] Anonymous. In brief. Network Security, 2011(2):3, February 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700125.

Durbin:2011:ISB

[3163] Steve Durbin. Information security without boundaries. Network Security, 2011(2):4–8, February 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700137.

Bradbury:2011:HWE

[3164] Danny Bradbury. Hacking wifi the easy way. Network Security, 2011 (2):9–12, February 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700149.

Gold:2011:FF

[3165] Steve Gold. The future of the firewall. Network Security, 2011(2): 13–15, February 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700150.

Ting:2011:TTA

[3166] David Ting. Thinking thin: addressing the challenges of client computing. Network Security, 2011(2): 16–17, February 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700162.

Hamelin:2011:PFA

[3167] Michael Hamelin. Preparing for a firewall audit. Network Security, 2011 (2):18–19, February 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700174.

Anonymous:2011:ZSH

[3168] Anonymous. Zeus and SpyEye hybrid emerges. Network Security, 2011(2):20, February 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700186.

Anonymous:2011:WMC

[3169] Anonymous. Waledac makes a comeback. Network Security, 2011 (2):20, February 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700198.

Anonymous:2011:ECb

[3170] Anonymous. Events calendar. Network Security, 2011(2):20, February 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700204.

Anonymous:2011:RZO

[3171] Anonymous. The resurgence of Zeus and other banking Trojans. Network Security, 2011(3):2, March 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700216.

Anonymous:2011:NHO

[3172] Anonymous. Nearly half of organisations are not protected. Network Security, 2011(3):2, 20, March 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700228.

Anonymous:2011:Bc

[3173] Anonymous. In brief. Network Security, 2011(3):3, March 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581170023X.

Taylor:2011:FIC

[3174] Mark Taylor, John Haggerty, David Gresty, and David Lamb. Forensic investigation of cloud computing systems. Network Security, 2011 (3):4–10, March 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700241.

Sanderson:2011:SDP

[3175] Russell Sanderson. A secure data protection strategy. Network Security, 2011(3):10–12, March 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700253.

Aljawarneh:2011:WES

[3176] Shadi Aljawarneh. A web engineering security methodology for elearning systems. Network Security, 2011(3):12–15, March 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700265.

Gold:2011:IMS

[3177] Steve Gold. IPv6 migration and security. Network Security, 2011 (3):15–18, March 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700277.

Lesnykh:2011:DLP

[3178] Alexei Lesnykh. Data loss prevention: a matter of discipline. Network Security, 2011(3):18–19, March 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700289.

Anonymous:2011:XTN

[3179] Anonymous. OS X Trojan now in beta. Network Security, 2011 (3):20, March 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700290.

Anonymous:2011:ECc

[3180] Anonymous. Events calendar. Network Security, 2011(3):20, March 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700307.

Anonymous:2011:RHL

[3181] Anonymous. RSA hack leaves status of SecurID uncertain. Network Security, 2011(4):1–2, April 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700319.

Anonymous:2011:FAS

[3182] Anonymous. Few aware of smartphone vulnerability. Network Security, 2011(4):2, 20, April 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700320.

Anonymous:2011:Bd

[3183] Anonymous. In brief. Network Security, 2011(4):3, April 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700332.

Anonymous:2011:BRSa

[3184] Anonymous. Book review: Security for Microsoft Windows Systems Administrators, by Derrick Rountree. Published by Syngress (ISBN 978-1-59749-594-3, e-ISBN 978-1-59749-595-0). Network Security, 2011(4):4, April 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700344.

Anonymous:2011:MWA

[3185] Anonymous. Microsoft Windows 7 administrator’s reference. Network Security, 2011(4):4, April 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700356.

Anonymous:2011:NBa

[3186] Anonymous. News in brief. Network Security, 2011(4):4, April 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700368.

Caldwell:2011:SS

[3187] Tracey Caldwell. Smart security.

Network Security, 2011(4):5–9, April 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581170037X.

Kemshall:2011:WMT

[3188] Andy Kemshall. Why mobile twofactor authentication makes sense. Network Security, 2011(4):9–12, April 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700381.

Gold:2011:CG

[3189] Steve Gold. Cracking GSM. Network Security, 2011(4):12–15, April 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700393.

Blandford:2011:ISC

[3190] Richard Blandford. Information security in the cloud. Network Security, 2011(4):15–17, April 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581170040X.

Jenkins:2011:LLS

[3191] Steve Jenkins. Learning to love SIEM. Network Security, 2011 (4):18–19, April 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700411.

Anonymous:2011:USU

[3192] Anonymous. Utilities still under threat. Network Security, 2011(4):20, April 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700423.

Anonymous:2011:Ca

[3193] Anonymous. Calendar. Network Security, 2011(4):20, April 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700435.

Anonymous:2011:UGT

[3194] Anonymous. US Government takes over Coreflood botnet. Network Security, 2011(5):1–2, May 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700447.

Anonymous:2011:NUF

[3195] Anonymous. NSS uncovers firewall shortcomings. Network Security, 2011(5):2, 19, May 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700459.

Anonymous:2011:Be

[3196] Anonymous. In brief. Network Security, 2011(5):3, May 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700460.

Anonymous:2011:BRCa

[3197] Anonymous. Book review: Computer Networks: a systems approach, by Larry Peterson and Bruce Davie. Fifth Edition. Morgan Kaufman. ISBN 978-0-12-385059-1. Network Security, 2011(5):4, May 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700472.

Anonymous:2011:NBb

[3198] Anonymous. News in brief. Network Security, 2011(5):4, May 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700484.

Bradbury:2011:RAC

[3199] Danny Bradbury. Routing around censorship. Network Security, 2011(5): 5–8, May 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485811700496.

Sood:2011:SBD

[3200] Aditya K. Sood and Richard J. Enbody. Spying on the browser: dissecting the design of malicious extensions. Network Security, 2011 (5):8–12, May 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700502.

Gold:2011:TBa

[3201] Steve Gold. Taking down botnets. Network Security, 2011 (5):13–15, May 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700514.

Jammalamadaka:2011:DSM

[3202] Ravi Chandra Jammalamadaka, Sharad Mehrotra, Nalini Venkatasubramanian, and Kent Seamons. DataVault: secure mobile access and data sharing. Network Security, 2011(5):16–19, May 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700526.

Anonymous:2011:EQS

[3203] Anonymous. EU questions stability of the Internet. Network Security, 2011(5):19–20, May 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700538.

Anonymous:2011:HMP

[3204] Anonymous. Hackers may be ‘pimping’ data. Network Security, 2011(5):20, May 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581170054X.

Anonymous:2011:ECd

[3205] Anonymous. Events calendar. Network Security, 2011(5):20, May 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700551.

Anonymous:2011:RRS

[3206] Anonymous. RSA replacing SecurID tokens. Network Security, 2011(6):1–2, June 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700563.

Anonymous:2011:GSP

[3207] Anonymous. Google spear-phishing targets governments and military. Network Security, 2011(6):2, 20, June 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700575.

Anonymous:2011:Bf

[3208] Anonymous. In brief. Network Security, 2011(6):3, June 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700587.

Anonymous:2011:BRI

[3209] Anonymous. Book review: IPv6 for Enterprise Networks, by Shannon McFarland, Muninder Sambi, Nikhil Sharma, and Sanjay Hooda. Cisco Press. ISBN 978-1-58714-227-

7. Network Security, 2011(6):4, June 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700599.

Anonymous:2011:BRPa

[3210] Anonymous. Book review: PKI Uncovered, Andre Karamanian, Srinivas Tenneti, François Dessart. Cisco Press.

ISBN 978-1-58705-916-2. Network Security, 2011(6):4, June 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700605.

Durbin:2011:TCT

[3211] Steve Durbin. Tackling converged threats: building a security-positive environment. Network Security, 2011(6):5–8, June 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700617.

Caldwell:2011:WWC

[3212] Tracey Caldwell. When worlds collide: the security of converged networks. Network Security, 2011 (6):8–12, June 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700629.

Hudson:2011:WMH

[3213] Jeff Hudson. Weaponised malware: how criminals use digital certificates to cripple your organisation. Network Security, 2011 (6):12–14, June 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700630.

Gold:2011:RP

[3214] Steve Gold. The rebirth of phreaking. Network Security, 2011(6):15–17, June 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700642.

Hocking:2011:TCS

[3215] Marc Hocking. Thin client security in the cloud. Network Security, 2011(6):17–19, June 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700654.

Anonymous:2011:MMA

[3216] Anonymous. More malware for Android. Network Security, 2011(6): 20, June 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485811700666.

Anonymous:2011:ECe

[3217] Anonymous. Events calendar. Network Security, 2011(6):20, June 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700678.

Anonymous:2011:HCH

[3218] Anonymous. High costs but higher profits with targeted attacks. Network Security, 2011(7):1–2, July 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581170068X.

Anonymous:2011:FTF

[3219] Anonymous. Firms turn down free data breach audits. Network Security, 2011(7):2, July 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700691.

Anonymous:2011:Bg

[3220] Anonymous. In brief. Network Security, 2011(7):3, July 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700708.

Anonymous:2011:BRCb

[3221] Anonymous. Book review: Cisco Firewalls, by Alexandre Moraes. Cisco Press. ISBN 978-1-58714-109-6. Network Security, 2011(7):4, July 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348581170071X.

Anonymous:2011:BRCc

[3222] Anonymous. Book review: Cisco Firewalls, by Alexandre Moraes. Cisco Press. ISBN 978-1-58714-109-6. Network Security, 2011(7):4, July 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485811700721.

Gold:2011:TBb

[3223] Steve Gold. Terrorism and Bluetooth. Network Security, 2011(7):5–7, July 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485811700733.

Gragido:2011:BZA

[3224] Will Gragido. Beyond zero: analysing threat trends. Network Security, 2011(7):7–9, July 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700745.

Caldwell:2011:EHP

[3225] Tracey Caldwell. Ethical hackers:

putting on the white hat. Network Security, 2011(7):10–13, July 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700757.

Sreenivas:2011:DKB

[3226] R. Sreeram Sreenivas and R. Anitha.

Detecting keyloggers based on traffic analysis with periodic behaviour. Network Security, 2011(7):14–19, July 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700769.

Anonymous:2011:WLW

[3227] Anonymous. Wins and losses in the war on botnets. Network Security, 2011(7):20, July 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700770.

Anonymous:2011:AQF

[3228] Anonymous. Al-Qaeda forum knocked offline. Network Security, 2011(7): 20, July 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485811700782.

Anonymous:2011:ECf

[3229] Anonymous. Events calendar. Network Security, 2011(7):20, July 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485811700794.

Anonymous:2011:BDH

[3230] Anonymous. Black Hat and DefCon highlights. Network Security, 2011 (8):2, 19–20, August 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700800.

Anonymous:2011:Bh

[3231] Anonymous. In brief. Network Security, 2011(8):3, August 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700812.

Anonymous:2011:BRPb

[3232] Anonymous. Book review: Practical Packet Analysis: using Wireshark to solve real-world network problems, by Chris Sanders. Second Edition. No Starch Press. ISBN 978-1-59327-266-1. Network Security, 2011(8):4, August 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700824.

Anonymous:2011:Cb

[3233] Anonymous. Correction. Network Security, 2011(8):4, August 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700836.

Mansfield-Devine:2011:HAD

[3234] Steve Mansfield-Devine. Hacktivism:

assessing the damage. Network Security, 2011(8):5–13, August 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811700848.

Turiel:2011:INT

[3235] Avi Turiel. IPv6: new technology, new threats. Network Security, 2011 (8):13–15, August 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581170085X.

Tankard:2011:APT

[3236] Colin Tankard. Advanced persistent threats and how to monitor and deter them. Network Security, 2011 (8):16–19, August 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700861.

Anonymous:2011:MUF

[3237] Anonymous. Mobile users face mounting threats. Network Security, 2011(8):20, August 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700873.

Anonymous:2011:WAA

[3238] Anonymous. Web apps attacked every two minutes. Network Security, 2011(8):20, August 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700885.

Anonymous:2011:ECg

[3239] Anonymous. Events calendar. Network Security, 2011(8):20, August 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485811700897.

Anonymous:2011:AHD

[3240] Anonymous. APT hype is a distraction from real security needs, says Gartner. Network Security, 2011 (9):1–2, September 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700903.

Anonymous:2011:AMA

[3241] Anonymous. Android is most-attacked mobile platform. Network Security, 2011(9):2, 20, September 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485811700915.

Anonymous:2011:Bi

[3242] Anonymous. In brief. Network Security, 2011(9):3, September 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485811700927.

Anonymous:2011:BRM

[3243] Anonymous. Book review: Metasploit: The Penetration Tester’s Guide, by David Kennedy, Jim O’Gorman, Devon Kearns, and Mati Aharoni. No Starch Press. ISBN 978-1-59327-288-3. Network Security, 2011 (9):4, September 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700939.

Gold:2011:CCN

[3244] Steve Gold. Cracking cellular networks via femtocells. Network Security, 2011(9):5–8, September 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700940.

Everett:2011:PNT

[3245] Cath Everett. Printers: the neglected threat. Network Security, 2011 (9):8–11, September 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700952.

Jammalamadaka:2011:PPD

[3246] Ravi Chandra Jammalamadaka, Sharad Mehrotra, and Nalini Venkatasubramanian. Protecting personal data from untrusted web-based data services. Network Security, 2011(9): 11–16, September 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811700964.

deJoode:2011:ECS

[3247] Alex de Joode. Effective corporate security and cybercrime. Network Security, 2011(9):16–18, September 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485811700976.

Creasey:2011:PD

[3248] Graeme Creasey. Protecting the datacentre. Network Security, 2011(9): 18–19, September 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811700988.

Anonymous:2011:ECh

[3249] Anonymous. Events calendar. Network Security, 2011(9):20, September 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348581170099X.

Anonymous:2011:WSU

[3250] Anonymous. Web security under threat. Network Security, 2011(10): 1–2, 20, October 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485811701003.

Anonymous:2011:Bj

[3251] Anonymous. In brief. Network Security, 2011(10):3, October 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485811701015.

Anonymous:2011:BRBa

[3252] Anonymous. Book review: The Book of Ruby, Huw Collingbourne. No Starch Press. ISBN 978-1-59327-294-4. Network Security, 2011 (10):4, October 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811701027.

Anonymous:2011:BRSb

[3253] Anonymous. Book review: Security Risk Management, Evan Wheeler. Syngress. ISBN 978-1-59749-615-5. Network Security, 2011(10):4, October 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811701039.

Gold:2011:AI

[3254] Steve Gold. Android insecurity. Network Security, 2011(10):5–7, October 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811701040.

Sood:2011:FFD

[3255] Aditya K. Sood and Richard J. Enbody. Frametrapping the framebusting defence. Network Security, 2011 (10):8–12, October 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811701052.

Harris:2011:DNS

[3256] James Harris. Defending the network several times over. Network Security, 2011(10):12–14, October 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811701064.

Nanda:2011:MDS

[3257] Rohan Nanda and P. Venkata Krishna. Mitigating denial of service attacks in hierarchical wireless sensor networks. Network Security, 2011 (10):14–18, October 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811701076.

Morrell:2011:CCN

[3258] Richard Morrell and Akash Chandrashekar. Cloud computing: new challenges and opportunities. Network Security, 2011(10):18–19, October 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811701088.

Anonymous:2011:SNW

[3259] Anonymous. Social networking in the workplace. Network Security, 2011(10):20, October 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581170109X.

Anonymous:2011:LLA

[3260] Anonymous. Lurid launches attack on Russia. Network Security, 2011 (10):20, October 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811701106.

Anonymous:2011:ECi

[3261] Anonymous. Events calendar. Network Security, 2011(10):20, October 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811701118.

Anonymous:2011:GWC

[3262] Anonymous. Governments warn of cyber-security. Network Security, 2011 (11):1–2, November 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581170112X.

Anonymous:2011:CES

[3263] Anonymous. Certificate ecosystem suffers further blows. Network Security, 2011(11):2, November 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811701131.

Anonymous:2011:NAT

[3264] Anonymous. Nitro attack targets chemical firms. Network Security, 2011(11):2, November 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811701143.

Anonymous:2011:Bk

[3265] Anonymous. In brief. Network Security, 2011(11):3, November 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811701155.

Anonymous:2011:BRCd

[3266] Anonymous. Book review: Comp-TIA Security + Deluxe Study Guide, by Emmett Dulaney. Second edition. Sybex ISBN 978-1-118-01474-5. Network Security, 2011(11):4, November 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811701167.

Andrus:2011:BSB

[3267] Frank Andrus. Beyond scan and block: an adaptive approach to network access control. Network Security, 2011 (11):5–9, November 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811701179.

Thomson:2011:APU

[3268] Gordon Thomson. APTs: a poorly understood challenge. Network Security, 2011(11):9–11, November 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811701180.

Gheri:2011:BAD

[3269] Klaus Gheri. The benefits of application detection. Network Security, 2011 (11):12–14, November 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811701192.

Gold:2011:CWN

[3270] Steve Gold. Cracking wireless networks. Network Security, 2011(11): 14–18, November 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811701209.

Facey:2011:WCS

[3271] Stuart Facey. Who’s in control: a six-step strategy for secure IT. Network Security, 2011(11): 18–20, November 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811701210.

Anonymous:2011:ECj

[3272] Anonymous. Events calendar. Network Security, 2011(11):20, November 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811701222.

Anonymous:2011:DUD

[3273] Anonymous. DDoS used to disrupt elections. Network Security, 2011 (12):1–2, December 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811701234.

Anonymous:2011:SMF

[3274] Anonymous. Security message fails to get through. Network Security, 2011 (12):2, 20, December 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811701246.

Anonymous:2011:Bl

[3275] Anonymous. In brief. Network Security, 2011(12):3, December 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811701258.

Anonymous:2011:BRPc

[3276] Anonymous. Book review: Penetration Tester’s Open Source Toolkit, by Jeremy Faircloth. Third edition. Syngress. ISBN 978-1-59749-627-8. Network Security, 2011 (12):4, December 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581170126X.

Anonymous:2011:BRBb

[3277] Anonymous. Book review: The Basics of Hacking and Penetration Testing, by Patrick Engebretson. Syngress. ISBN 978-1-59749-655-1. Network Security, 2011(12):4, December 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811701271.

Mansfield-Devine:2011:DTM

[3278] Steve Mansfield-Devine. DDoS: threats and mitigation. Network Security, 2011 (12):5–12, December 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811701283.

DHoinne:2011:CWS

[3279] Jérémy D’Hoinne. Could ‘wait and see’ be the best IPv6 strategy? Network Security, 2011(12): 12–14, December 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811701295.

Gold:2011:UHP

[3280] Steve Gold. Understanding the hacker psyche. Network Security, 2011(12): 15–17, December 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811701301.

Malecki:2011:DBL

[3281] Florian Malecki. A deeper, broader look at network security. Network Security, 2011(12):18–19, December 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485811701313.

Anonymous:2011:BBC

[3282] Anonymous. Big boost in cybersecurity spending. Network Security, 2011(12):20, December 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485811701325.

Anonymous:2011:ECk

[3283] Anonymous. Events calendar. Network Security, 2011(12):20, December 2011. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485811701337.

Anonymous:2012:HAS

[3284] Anonymous. Hackers attack security organisations. Network Security, 2012(1):1–2, January 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700016.

Anonymous:2012:HWM

[3285] Anonymous. Hackers warring in Middle East. Network Security, 2012(1):2, January 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700028.

Anonymous:2012:Ba

[3286] Anonymous. In brief. Network Security, 2012(1):3, January 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581270003X.

Anonymous:2012:BRBa

[3287] Anonymous. Book review: A Bug Hunter’s Diary, by Tobias Klein. No Starch Press. ISBN 978-1-59327-385-9. Network Security, 2012 (1):4, January 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700041.

Harley:2012:ATT

[3288] David Harley. AMTSO: the test of time? Network Security, 2012 (1):5–10, January 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700053.

Woodhead:2012:MBT

[3289] Simon Woodhead. Monitoring bad traffic with darknets. Network Security, 2012(1):10–14, January 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700065.

Ryan:2012:REB

[3290] Jaime Ryan. Rethinking the ESB:

building a secure bus with an SOA gateway. Network Security, 2012 (1):14–17, January 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700077.

Masadeh:2012:SMB

[3291] Shadi R. Masadeh, Nedal Turab, and Farhan Obisat. A secure model for building e-learning systems. Network Security, 2012(1): 17–20, January 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700089.

Anonymous:2012:ECa

[3292] Anonymous. Events calendar. Network Security, 2012(1):20, January 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485812700090.

Anonymous:2012:HHS

[3293] Anonymous. Hacktivists hit out at Symantec, police and anti-piracy groups. Network Security, 2012 (2):1–2, February 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700107.

Anonymous:2012:Bb

[3294] Anonymous. In brief. Network Security, 2012(2):3, February 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485812700119.

Anonymous:2012:BRW

[3295] Anonymous. Book review: The Web Application Hacker’s Handbook, 2nd Edition. Dafydd Stuttard and Marcus Pinto. Wiley. ISBN 978-1-118-02647-2. Network Security, 2012 (2):4, February 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700120.

Thomson:2012:BEC

[3296] Gordon Thomson. BYOD: enabling the chaos. Network Security, 2012 (2):5–8, February 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700132.

Wray:2012:SCN

[3297] Melvyn Wray. From server consolidation to network consolidation. Network Security, 2012(2):

8–11, February 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700144.

Caviglione:2012:EIS

[3298] Luca Caviglione and Alessio Merlo.

The energy impact of security mechanisms in modern mobile devices. Network Security, 2012(2):11–14, February 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485812700156.

Schneider:2012:SNS

[3299] David Schneider. The state of network security. Network Security, 2012 (2):14–20, February 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700168.

Anonymous:2012:TTS

[3300] Anonymous. Tough times for SSL. Network Security, 2012 (2):20, February 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581270017X.

Anonymous:2012:ECb

[3301] Anonymous. Events calendar. Network Security, 2012(2):20, February 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485812700181.

Anonymous:2012:AAL

[3302] Anonymous. Anonymous — arrests, leaks and infections. Network Security, 2012(3):2, March 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700399.

Anonymous:2012:Bc

[3303] Anonymous. In brief. Network Security, 2012(3):3, March 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700405.

Anonymous:2012:BRT

[3304] Anonymous. Book review: Thor’s Microsoft Security Bible, by Timothy ‘Thor’ Mullen. Syngress. ISBN 978-1-59749-572-1. Network Security, 2012(3):4, March 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700417.

Anonymous:2012:BRL

[3305] Anonymous. Book review: Low Tech Hacking: Street Smarts for Security Professionals, by Jack Wiles, Terry Gudaitis, Jennifer Jabbusch, Russ Rogers, and Sean Lowther. Syngress. ISBN 978-1-59749-665-0. Network Security, 2012(3):4, March 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700429.

Denman:2012:WML

[3306] Simon Denman. Why multi-layered security is still the best defence. Network Security, 2012(3):5–7, March 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485812700430.

Brewer:2012:PCC

[3307] Ross Brewer. Protecting critical control systems. Network Security, 2012(3):7–10, March 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700442.

Rouse:2012:MDM

[3308] Jason Rouse. Mobile devices — the most hostile environment for security? Network Security, 2012 (3):11–13, March 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700454.

Gold:2012:SV

[3309] Steve Gold. Securing VoIP. Network Security, 2012(3):14–17, March 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485812700466.

Dowling:2012:ERW

[3310] Mike Dowling. Enabling remote working: protecting the network. Network Security, 2012(3):18–20, March 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485812700478.

Anonymous:2012:ECc

[3311] Anonymous. Events calendar. Network Security, 2012(3):20, March 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348581270048X.

Anonymous:2012:MGF

[3312] Anonymous. Mac gets first serious malware. Network Security, 2012(4):1–2, April 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700193.

Anonymous:2012:CGT

[3313] Anonymous. Cybercrime a growing threat to financial services. Network Security, 2012(4):2, April 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348581270020X.

Anonymous:2012:Bd

[3314] Anonymous. In brief. Network Security, 2012(4):3, April 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700211.

Anonymous:2012:TW

[3315] Anonymous. The tangled Web. Network Security, 2012(4):4, April 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485812700223.

Anonymous:2012:CW

[3316] Anonymous. Cyber warfare. Network Security, 2012(4):4, April 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485812700235.

deCrespigny:2012:BCR

[3317] Michael de Crespigny. Building cyberresilience to tackle threats. Network Security, 2012(4):5–8, April 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485812700247.

Tankard:2012:TMP

[3318] Colin Tankard. Taking the management pain out of Active Directory. Network Security, 2012 (4):8–11, April 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700259.

Robertson:2012:SVL

[3319] Brian Robertson. Security: virtualisation’s last frontier. Network Security, 2012(4):12–15, April 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700260.

Everett:2012:EC

[3320] Bernard Everett. The encryption conundrum. Network Security, 2012 (4):15–18, April 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700272.

Serrato:2012:IAN

[3321] Christy Serrato. Identity assurance and network security. Network Security, 2012(4):19–20, April 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700284.

Anonymous:2012:ECd

[3322] Anonymous. Events calendar. Network Security, 2012(4):20, April 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485812700296.

Anonymous:2012:TIM

[3323] Anonymous. Trustworthy Internet Movement aims to solve SSL issues. Network Security, 2012(5):1–2, May 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485812700302.

Anonymous:2012:APM

[3324] Anonymous. Android’s popularity makes it a prime target. Network Security, 2012(5):2, May 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700314.

Anonymous:2012:Be

[3325] Anonymous. In brief. Network Security, 2012(5):3, May 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700326.

Anonymous:2012:BRBb

[3326] Anonymous. Book review: The Basics of Digital Forensics, by John Sammons. Syngress. ISBN 978-1-59749-661-2. Network Security, 2012(5):4, May 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700338.

Anonymous:2012:BRBc

[3327] Anonymous. Book review: The Basics of Digital Forensics, by John Sammons. Syngress. ISBN 978-1-59749-661-2. Network Security, 2012(5):4, May 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581270034X.

Mason:2012:CCS

[3328] Andrew Mason. Caught in the crosssite scripting fire. Network Security, 2012(5):5–9, May 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700351.

Gold:2012:WCT

[3329] Steve Gold. Wireless cracking: there’s an app for that. Network Security, 2012(5):10–14, May 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700363.

Hamamreh:2012:RPA

[3330] Rushdi Hamamreh. Routing path authentication in link-state routing protocols. Network Security, 2012 (5):14–20, May 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700375.

Anonymous:2012:ECe

[3331] Anonymous. Events calendar. Network Security, 2012(5):20, May 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485812700387.

Anonymous:2012:FCH

[3332] Anonymous. Flame creates heated debate. Network Security, 2012 (6):1–2, June 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700491.

Anonymous:2012:Bf

[3333] Anonymous. In brief. Network Security, 2012(6):3, June 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700508.

SM-D:2012:BRW

[3334] SM-D. Book review: Windows Forensic Analysis Toolkit, by Harlan Carvey. Third edition. Syngress. ISBN 978-1-59749-727-5. Price: $69.95, 272 pgs, paperback. Network Security, 2012(6):4, June 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581270051X.

Jones:2012:MMW

[3335] Greg Jones. Mobile menace: why SDR poses such a threat. Network Security, 2012(6):5–7, June 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700521.

Kenyon:2012:WAC

[3336] Paul Kenyon. What Australia can teach the world about least privilege. Network Security, 2012 (6):7–10, June 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700533.

Bager:2012:RAD

[3337] Kurt Bager. Remote access: don’t be a victim. Network Security, 2012(6):11–14, June 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700545.

Caldwell:2012:LV

[3338] Tracey Caldwell. Locking down the VPN. Network Security, 2012 (6):14–18, June 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700557.

Eisen:2012:CFM

[3339] Ori Eisen. Catching the fraudulent Man-in-the-Middle. Network Security, 2012(6):18–20, June 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700569.

Anonymous:2012:ECf

[3340] Anonymous. Events calendar. Network Security, 2012(6):20, June 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485812700570.

Anonymous:2012:AMG

[3341] Anonymous. Android malware growth and possible botnet. Network Security, 2012(7):1–2, July 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700582.

Anonymous:2012:BIC

[3342] Anonymous. Big increase in critical infrastructure attacks. Network Security, 2012(7):2, July 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700594.

Anonymous:2012:Bg

[3343] Anonymous. In brief. Network Security, 2012(7):3, July 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700600.

Anonymous:2012:BRPa

[3344] Anonymous. Book review: Privacy and Big Data: the players, regulators and stakeholders, by Terence Craig and Mary E. Ludloff. O’Reilly. ISBN 978-1-4493-0500-0. Network Security, 2012(7):4, July 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700612.

Anonymous:2012:BRPb

[3345] Anonymous. Book review: Privacy and Big Data: the players, regulators and stakeholders, by Terence Craig, and Mary E. Ludloff. O’Reilly. ISBN 978-1-4493-0500-0. Network Security, 2012(7):4, July 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700624.

Tankard:2012:BDS

[3346] Colin Tankard. Big data security.

Network Security, 2012(7):5–8, July 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485812700636.

Potts:2012:SIS

[3347] Mike Potts. The state of information security. Network Security, 2012(7):9–11, July 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700648.

Mansfield-Devine:2012:EWD

[3348] Steve Mansfield-Devine. Estonia: what doesn’t kill you makes you stronger. Network Security, 2012 (7):12–20, July 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581270065X.

Anonymous:2012:ECg

[3349] Anonymous. Events calendar. Network Security, 2012(7):20, July 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485812700661.

Anonymous:2012:BH

[3350] Anonymous. Black Hat highlights. Network Security, 2012 (8):1–2, August 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700673.

Anonymous:2012:Bi

[3351] Anonymous. In brief. Network Security, 2012(8):3, August 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700685.

SM-D:2012:BRPa

[3352] SM-D. Book review: Planning for IPv6, by Silvia Hagen. O’Reilly. ISBN 978-1-4493-0538-3 (e-book), 978-1-4493-0539-0 (print). Network Security, 2012(8):4, August 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700697.

SM-D:2012:BRPb

[3353] SM-D. Book review: Planning for IPv6, by Silvia Hagen. O’Reilly. ISBN 978-1-4493-0538-3 (e-book), 978-1-4493-0539-0 (print). Network Security, 2012(8):4, August 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700703.

Conrad:2012:SHI

[3354] James Conrad. Seeking help: the important role of ethical hackers. Network Security, 2012(8):5–8, August 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485812700715.

Caldwell:2012:PDW

[3355] Tracey Caldwell. The perimeter is dead — what next for the appliance? Network Security, 2012 (8):8–12, August 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700727.

Bradbury:2012:FBS

[3356] Danny Bradbury. Fighting botnets with sinkholes. Network Security, 2012 (8):12–15, August 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700739.

Gold:2012:CJ

[3357] Steve Gold. Cellular jamming. Network Security, 2012(8):15–18, August 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485812700740.

Malecki:2012:SWD

[3358] Florian Malecki. Simple ways to dodge the DDoS bullet. Network Security, 2012(8):18–20, August 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700752.

Anonymous:2012:ECh

[3359] Anonymous. Events calendar. Network Security, 2012(8):20, August 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485812700764.

Anonymous:2012:AHF

[3360] Anonymous. Anonymous hacks FBI laptop. Or maybe not. Network Security, 2012(9):1–2, September 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485812700776.

Anonymous:2012:MEO

[3361] Anonymous. Middle East oil firms hit by massive attacks. Network Security, 2012(9):2, 19, September 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485812700788.

Anonymous:2012:Bj

[3362] Anonymous. In brief. Network Security, 2012(9):3, September 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581270079X.

SM-D:2012:BRS

[3363] SM-D. Book review: Securing Cloud Services: a pragmatic, approach to security architecture in, the cloud, by Lee Newcombe. Published by IT Governance Publishing. ISBN 978-1-84928-398-4. Network Security, 2012 (9):4, September 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700806.

Mansfield-Devine:2012:PAJ

[3364] Steve Mansfield-Devine. Paranoid Android: just how insecure is the most popular mobile platform? Network Security, 2012(9): 5–10, September 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700818.

Tankard:2012:PMS

[3365] Colin Tankard. The promise of managed security services. Network Security, 2012(9):10–15, September 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581270082X.

Caldwell:2012:SD

[3366] Tracey Caldwell. Seek and destroy. Network Security, 2012(9): 15–19, September 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700831.

Anonymous:2012:AMJ

[3367] Anonymous. Another major Java panic. Network Security, 2012(9): 19–20, September 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700843.

Anonymous:2012:NRI

[3368] Anonymous. New research institute to study threats. Network Security, 2012(9):20, September 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700855.

Anonymous:2012:ECi

[3369] Anonymous. Events calendar. Network Security, 2012(9):20, September 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485812700867.

Anonymous:2012:ACE

[3370] Anonymous. Avoid Chinese equipment firms, says Congress. Network Security, 2012(10):1–2, October 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812700879.

Anonymous:2012:MEI

[3371] Anonymous. Mobile exploits increase. Network Security, 2012 (10):2, October 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700880.

Anonymous:2012:Bk

[3372] Anonymous. In brief. Network Security, 2012(10):3, October 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485812700892.

Anonymous:2012:BRO

[3373] Anonymous. Book review: Once More Unto the Breach, by Andrea C Simmons. IT Governance Publishing. ISBN 978-1-84928-390-8. Network Security, 2012(10):4, October 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700909.

Anonymous:2012:BRP

[3374] Anonymous. Book review: PCI Compliance, Branden R. Williams and Anton Chuvakin. Third edition. Syngress. ISBN 978-1-59749-948-4. Network Security, 2012 (10):4, October 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700910.

Mansfield-Devine:2012:AAA

[3375] Steve Mansfield-Devine. Android architecture: attacking the weak points. Network Security, 2012 (10):5–12, October 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700922.

Jordon:2012:CDD

[3376] Michael Jordon. Cleaning up dirty disks in the cloud. Network Security, 2012(10):12–15, October 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485812700934.

Bradbury:2012:SFF

[3377] Danny Bradbury. Spreading fear on Facebook. Network Security, 2012 (10):15–17, October 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700946.

Bella:2012:ISA

[3378] Madeleine A. Bihina Bella, J. H. P. Eloff, and Martin S. Olivier. Improving system availability with nearmiss analysis. Network Security, 2012 (10):18–20, October 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700958.

Anonymous:2012:ECj

[3379] Anonymous. Events calendar. Network Security, 2012(10):20, October 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581270096X.

Anonymous:2012:GHH

[3380] Anonymous. Georgia hacks hacker, allegedly. Network Security, 2012 (11):1–2, November 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700971.

Anonymous:2012:LAP

[3381] Anonymous. Latest Android problems and fixes. Network Security, 2012(11):2, November 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812700983.

Anonymous:2012:Bl

[3382] Anonymous. In brief. Network Security, 2012(11):3, November 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485812700995.

Anonymous:2012:BRR

[3383] Anonymous. Book review: Regular Expressions Cookbook, Jan Goyvaerts and Steven Levithan, Second edition, O’Reilly, ISBN 978-1-4493-1943-4. Network Security, 2012 (11):4, November 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812701009.

Anonymous:2012:BRH

[3384] Anonymous. Book review: Hacking Web Apps, Mike Shema. Syngress. ISBN 978-1-59749-951-4. Network Security, 2012(11):4, November 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485812701010.

Tankard:2012:CIS

[3385] Colin Tankard. Cultural issues in security and privacy. Network Security, 2012(11):5–8, November 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812701022.

Paessler:2012:MPC

[3386] Dirk Paessler. Monitoring private clouds. Network Security, 2012(11): 8–12, November 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812701034.

Mansfield-Devine:2012:AMM

[3387] Steve Mansfield-Devine. Android malware and mitigations. Network Security, 2012(11):12–20, November 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485812701046.

Anonymous:2012:ECk

[3388] Anonymous. Events calendar. Network Security, 2012(11):20, November 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485812701058.

Anonymous:2012:LB

[3389] Anonymous. A look back at 2012. Network Security, 2012(12):

1–2, December 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581270106X.

Anonymous:2012:HGA

[3390] Anonymous. How good is antivirus? Network Security, 2012 (12):2, December 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812701071.

Anonymous:2012:Bm

[3391] Anonymous. In brief. Network Security, 2012(12):3, December 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485812701083.

SM-D:2012:BRPc

[3392] SM-D. Book review: Practical Malware Analysis, by Michael Sikorski and Andrew Honig. No Starch Press. ISBN 978-1-59327-290-6. Network Security, 2012(12):4, December 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485812701095.

SM-D:2012:BRE

[3393] SM-D. Book review: Encyclopedia of Electronic Components, Volume 1, Charles Platt. Make/O’Reilly, ISBN 978-1-4493-3389-8. Network Security, 2012(12):4, December 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812701101.

Morrow:2012:BSC

[3394] Bill Morrow. BYOD security challenges: control and protect your most sensitive data. Network Security, 2012 (12):5–8, December 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812701113.

Sood:2012:AGP

[3395] Aditya K. Sood, Peter Greko, and Richard J. Enbody. Abusing Glype proxies: attacks, exploits and defences. Network Security, 2012(12): 8–15, December 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812701125.

Gold:2012:VJH

[3396] Steve Gold. Virtual jihad: how real is the threat? Network Security, 2012 (12):15–18, December 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812701137.

Malecki:2012:NGF

[3397] Florian Malecki. Next-generation firewalls: security with performance. Network Security, 2012(12): 19–20, December 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485812701149.

Anonymous:2012:ECl

[3398] Anonymous. Events calendar. Network Security, 2012(12):20, December 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485812701150.

Anonymous:2013:CER

[3399] Anonymous. CA error results in certificate blunder. Network Security, 2013(1):1–2, January 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700126.

Anonymous:2013:UBA

[3400] Anonymous. US banks attacked — but by whom? Network Security, 2013(1):2, January 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700138.

Anonymous:2013:Ba

[3401] Anonymous. In brief. Network Security, 2013(1):3, January 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581370014X.

SM-D:2013:BRC

[3402] SM-D. Book review: CUDA Programming, Shane Cook. Morgan Kaufmann. ISBN 978-0-12-415933-4. Network Security, 2013(1):4, January 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813700151.

SM-D:2013:BRT

[3403] SM-D. Book review: Think Like a Programmer, V Anton Spraul. No Starch Press. ISBN 978-1-59327-424-5. Network Security, 2013 (1):4, January 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700163.

Bradbury:2013:VSS

[3404] Danny Bradbury. Verifying software security — is it possible? Network Security, 2013(1):5–7, January 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485813700175.

Kurpjuhn:2013:ERU

[3405] Thorsten Kurpjuhn. The evolving role of the UTM appliance. Network Security, 2013(1):8–11, January 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485813700187.

Hart:2013:WTA

[3406] Jason Hart. Why the traditional approach to information security is no longer working. Network Security, 2013 (1):12–14, January 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700199.

Bhat:2013:UMS

[3407] Wasim Ahmad Bhat and S. M. K. Quadri. Understanding and mitigating security issues in Sun NFS. Network Security, 2013(1):15–18, January 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813700205.

Walker-Brown:2013:MVM

[3408] Andrew Walker-Brown. Managing VPNs in the mobile worker’s world. Network Security, 2013(1): 18–20, January 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700217.

Anonymous:2013:ECa

[3409] Anonymous. Events calendar. Network Security, 2013(1):20, January 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813700229.

Anonymous:2013:BBH

[3410] Anonymous. Bit9 bitten by hackers. Network Security, 2013(2): 1–2, February 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700230.

Anonymous:2013:MSC

[3411] Anonymous. Microsoft and Symantec cripple botnet. Network Security, 2013(2):2, February 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700242.

Anonymous:2013:Bb

[3412] Anonymous. In brief. Network Security, 2013(2):3, February 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485813700254.

Anonymous:2013:BRL

[3413] Anonymous. Book review: Logging and Log Management, Anton Chuvakin, Kevin Schmidt and Christopher Phillips. Syngress. ISBN 978-1-59749-635-3. Network Security, 2013(2):4, February 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700266.

Anonymous:2013:BRC

[3414] Anonymous. Book review: CISSP Study Guide, Eric Conrad, Seth Misenar and Joshua Feldman. Syngress. ISBN 978-1-59749-961-3. Network Security, 2013(2):4, February 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485813700278.

Jones:2013:PC

[3415] Greg Jones. Penetrating the cloud. Network Security, 2013(2):5–7, February 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S135348581370028X.

Hirst:2013:SYT

[3416] Mark Hirst. Securing your telemetry: making DCIM unsniffable. Network Security, 2013(2):8–9, February 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813700291.

Meulensteen:2013:SGM

[3417] Marc Meulensteen. A smart grid must also be streetwise. Network Security, 2013(2):10–12, February 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813700308.

Goh:2013:MST

[3418] Jane Goh. Moving security testing into the developer’s domain. Network Security, 2013(2):13–15, February 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581370031X.

Gold:2013:ECS

[3419] Steve Gold. Electronic countersurveillance strategies. Network Security, 2013(2):15–18, February 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813700321.

Smith:2013:LCD

[3420] Don Smith. Life’s certainties: death, taxes and APTs. Network Security, 2013(2):19–20, February 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813700333.

Anonymous:2013:ECb

[3421] Anonymous. Events calendar. Network Security, 2013(2):20, February 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485813700345.

Anonymous:2013:FAM

[3422] Anonymous. Facebook, Apple and Microsoft hit by Java exploit at watering hole. Network Security, 2013(3):1–2, March 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700357.

Anonymous:2013:URP

[3423] Anonymous. US report points finger at China’s military cyberwar unit. Network Security, 2013(3):2, March 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485813700369.

Anonymous:2013:Bc

[3424] Anonymous. In brief. Network Security, 2013(3):3, March 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700370.

Anonymous:2013:BRW

[3425] Anonymous. Book review: Windows PowerShell 3.0 Step by Step, by Ed Wilson. Microsoft Press. ISBN 978-0-7356-6339-8 (print), 978-0-7356-6337-4 (ebook). Network Security, 2013(3):4, March 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700382.

Pitt:2013:TCR

[3426] Dan Pitt. Trust in the cloud: the role of SDN. Network Security, 2013(3):5–6, March 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700394.

Nachreiner:2013:BSA

[3427] Corey Nachreiner. Beat security auditors at their own game. Network Security, 2013(3):7–11, March 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813700400.

Winter:2013:SVH

[3428] Robert Winter. SSD vs HDD — data recovery and destruction. Network Security, 2013(3):12–14, March 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813700412.

Bassill:2013:HAS

[3429] Peter Bassill. The holistic approach to security. Network Security, 2013 (3):14–17, March 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700424.

Fehr:2013:FNB

[3430] Shannon Fehr. Flexible networks for better security. Network Security, 2013(3):17–20, March 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700436.

Anonymous:2013:ECc

[3431] Anonymous. Events calendar. Network Security, 2013(3):20, March 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485813700448.

Anonymous:2013:SDA

[3432] Anonymous. Spamhaus DDoS attack fails to take down Internet. Network Security, 2013(4):1–2, April 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348581370045X.

Anonymous:2013:UBCa

[3433] Anonymous. US bans Chinese technology. Network Security, 2013(4):2, April 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485813700461.

Anonymous:2013:Bd

[3434] Anonymous. In brief. Network Security, 2013(4):3, April 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700473.

Anonymous:2013:BRI

[3435] Anonymous. Book review: Instant Penetration Testing: Setting Up a Test Lab How-to, by Vyacheslav Fadyushin. Packt Publishing. ISBN 1-84969-412-5. Network Security, 2013(4):4, April 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700485.

Walters:2013:BIS

[3436] Richard Walters. Bringing IT out of the shadows. Network Security, 2013(4):5–11, April 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700497.

Tokuyoshi:2013:SIB

[3437] Brian Tokuyoshi. The security implications of BYOD. Network Security, 2013(4):12–13, April 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700503.

Caldwell:2013:INP

[3438] Tracey Caldwell. Identity — the new perimeter. Network Security, 2013(4):14–18, April 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700515.

Newbold:2013:SMH

[3439] Gary Newbold. Secure mobility in healthcare networks for optimal patient care. Network Security, 2013 (4):18–20, April 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700527.

Anonymous:2013:ECd

[3440] Anonymous. Events calendar. Network Security, 2013(4):20, April 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485813700539.

Anonymous:2013:UBCb

[3441] Anonymous. US blames China for hacking, again. Network Security, 2013(5):1–2, May 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700540.

Anonymous:2013:UTS

[3442] Anonymous. US takes down Syrian domains. Network Security, 2013(5):2, May 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700552.

Anonymous:2013:Be

[3443] Anonymous. In brief. Network Security, 2013(5):3, May 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700564.

Anonymous:2013:BRT

[3444] Anonymous. Book review: Tallinn Manual on the International Law Applicable to Cyber Warfare, edited by Michael Schmitt. Cambridge University Press. ISBN 978-1-107-61377-5 (paperback), 978-1-107-02443-4 (hardback). Network Security, 2013(5):4, May 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700576.

McGregory:2013:PND

[3445] Steve McGregory. Preparing for the next DDoS attack. Network Security, 2013(5):5–6, May 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700588.

Caldwell:2013:SDL

[3446] Tracey Caldwell. Security at the data level. Network Security, 2013 (5):6–12, May 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581370059X.

Grafton:2013:AFP

[3447] Jane Grafton. Avoiding the five pitfalls of privileged accounts. Network Security, 2013(5):12–14, May 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813700606.

Gold:2013:BEM

[3448] Steve Gold. Black Hat Europe:

mobile attack vectors in the spotlight. Network Security, 2013 (5):14–18, May 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700618.

Hirst:2013:RSP

[3449] Mark Hirst. Rack security and protection. Network Security, 2013 (5):18–20, May 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581370062X.

Anonymous:2013:ECe

[3450] Anonymous. Events calendar. Network Security, 2013(5):20, May 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485813700631.

Anonymous:2013:PSF

[3451] Anonymous. PandaLabs says the fight against cybercrime is being won. Network Security, 2013 (6):1–2, June 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700643.

Anonymous:2013:MSB

[3452] Anonymous. McAfee says it’s business as usual. Network Security, 2013(6):2, June 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700655.

Anonymous:2013:Bf

[3453] Anonymous. In brief. Network Security, 2013(6):3, June 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700667.

Anonymous:2013:BRV

[3454] Anonymous. Book review: Violent Python, by T. J. O’Connor. Syngress. ISBN 978-1-59749-957-6. Network Security, 2013(6):4, June 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700679.

Barker:2013:SII

[3455] Keith Barker. The security implications of IPv6. Network Security, 2013(6):5–9, June 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700680.

Gold:2013:GLI

[3456] Steve Gold. Getting lost on the Internet: the problem with anonymity. Network Security, 2013 (6):10–13, June 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700692.

Brooks:2013:CEI

[3457] Tim Brooks. Classic enterprise IT:

the castle approach. Network Security, 2013(6):14–16, June 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700709.

Mansfield-Devine:2013:QCT

[3458] Steve Mansfield-Devine. Q&A: Colin Tankard — raising security awareness. Network Security, 2013 (6):16–19, June 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700710.

Anonymous:2013:UWI

[3459] Anonymous. US warns of increased cyber-attacks by Iran. Network Security, 2013(6):19–20, June 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485813700722.

Anonymous:2013:JTD

[3460] Anonymous. Journalists threatened with data protection laws after using exposed information. Network Security, 2013(6):20, June 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700734.

Anonymous:2013:NUA

[3461] Anonymous. NATO under attack. Network Security, 2013(6):20, June 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485813700746.

Anonymous:2013:ECf

[3462] Anonymous. Events calendar. Network Security, 2013(6):20, June 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813700758.

Anonymous:2013:WSW

[3463] Anonymous. Web is still the weakest point. Network Security, 2013(7):1–2, July 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581370076X.

Anonymous:2013:CFF

[3464] Anonymous. Crypto flaw found in Android. Network Security, 2013(7):2, July 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700771.

Anonymous:2013:Bg

[3465] Anonymous. In brief. Network Security, 2013(7):3, July 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700783.

Anonymous:2013:BRH

[3466] Anonymous. Book review: Hacking Web Apps, by Mike Shema. Syngress. ISBN 978-1-59749-951-4. Network Security, 2013(7):4, July 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700795.

Ring:2013:MSI

[3467] Tim Ring. IT’s megatrends: the security impact. Network Security, 2013(7):5–8, July 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700801.

Bradbury:2013:OD

[3468] Danny Bradbury. Offensive defence. Network Security, 2013 (7):9–12, July 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700813.

Scully:2013:ULK

[3469] Patrick Scully. Under lock and key: protecting the network from attack. Network Security, 2013 (7):12–15, July 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700825.

Begg:2013:SSA

[3470] Norman Begg. Securely sharing access to sensitive data. Network Security, 2013(7):15–18, July 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813700837.

Kerpan:2013:BCB

[3471] Patrick Kerpan and Sam Mitchell. Bringing control to the business application layer. Network Security, 2013(7):18–20, July 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700849.

Anonymous:2013:ECg

[3472] Anonymous. Events calendar. Network Security, 2013(7):20, July 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813700850.

Anonymous:2013:TAP

[3473] Anonymous. TOR attacked — possibly by the NSA. Network Security, 2013(8):1–2, August 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700862.

Anonymous:2013:MIM

[3474] Anonymous. Mobile insecurities multiply. Network Security, 2013 (8):2, 20, August 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700874.

Anonymous:2013:Bh

[3475] Anonymous. In brief. Network Security, 2013(8):3, August 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700886.

Anonymous:2013:BRBa

[3476] Anonymous. Book review: The Basics of Web Hacking, by Josh Pauli. Syngress. ISBN 978-0-12-416600-4. Network Security, 2013 (8):4, August 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700898.

Anonymous:2013:BRBb

[3477] Anonymous. Book review: The Basics of Web Hacking, by Josh Pauli. Syngress. ISBN 978-0-12-416600-4. Network Security, 2013 (8):4, August 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700904.

Thielens:2013:WAC

[3478] John Thielens. Why APIs are central to a BYOD security strategy. Network Security, 2013(8):5–6, August 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813700916.

Gaffney:2013:FFW

[3479] Tom Gaffney. Following in the footsteps of Windows: how Android malware development is looking very familiar. Network Security, 2013 (8):7–10, August 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700928.

Mansfield-Devine:2013:IME

[3480] Steve Mansfield-Devine. Interview: Mick Ebsworth — a matter of trust. Network Security, 2013 (8):11–13, August 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581370093X.

Gold:2013:SCT

[3481] Steve Gold. Subverting cellular technology: evolution, not revolution. Network Security, 2013 (8):14–19, August 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700941.

Anonymous:2013:FMI

[3482] Anonymous. Firms making it easy for attackers, says KMPG. Network Security, 2013(8):20, August 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813700953.

Anonymous:2013:ECh

[3483] Anonymous. Events calendar. Network Security, 2013(8):20, August 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813700965.

Anonymous:2013:NCI

[3484] Anonymous. NSA has cracked Internet encryption protocols. Network Security, 2013(9):1–2, September 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813700977.

Anonymous:2013:Bi

[3485] Anonymous. In brief. Network Security, 2013(9):3, September 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485813700989.

Anonymous:2013:IOH

[3486] Anonymous. Instant OSSEC hostbased intrusion detection. Network Security, 2013(9):4, September 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485813700990.

Anonymous:2013:SSD

[3487] Anonymous. Simple steps to data encryption. Network Security, 2013 (9):4, September 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813701004.

Arean:2013:DRC

[3488] Oscar Arean. Disaster recovery in the cloud. Network Security, 2013 (9):5–7, September 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813701016.

Kenyon:2013:RAW

[3489] Tony Kenyon. The role of ADCs within secure application delivery. Network Security, 2013(9):8–12, September 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485813701028.

Dunford:2013:MFT

[3490] Dan Dunford. Managed file transfer: the next stage for data in motion? Network Security, 2013(9): 12–15, September 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581370103X.

Hibbert:2013:CTC

[3491] Richard Hibbert. Calling time on compliance spreadsheet overload. Network Security, 2013(9): 15–17, September 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813701041.

Mansfield-Devine:2013:IJC

[3492] Steve Mansfield-Devine. Interview: Jon Callas, Silent Circle. Network Security, 2013(9): 17–20, September 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813701053.

Anonymous:2013:ECi

[3493] Anonymous. Events calendar. Network Security, 2013(9):20, September 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813701065.

Anonymous:2013:SRO

[3494] Anonymous. Silk Road online drug market taken down. Network Security, 2013(10):1–2, October 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813701077.

Anonymous:2013:MMT

[3495] Anonymous. Mobile malware tops one million, but Google says problem exaggerated. Network Security, 2013(10):2, October 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813701089.

Anonymous:2013:Bj

[3496] Anonymous. In brief. Network Security, 2013(10):3, October 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485813701090.

Anonymous:2013:ICW

[3497] Anonymous. Introduction to cyberwarfare. Network Security, 2013 (10):4, October 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813701107.

Anonymous:2013:ACS

[3498] Anonymous. Applied cyber-security and the Smart Grid. Network Security, 2013(10):4, October 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485813701119.

Small:2013:DBI

[3499] Mike Small. From data breach to information stewardship. Network Security, 2013(10):5–8, October 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813701120.

Erdheim:2013:DMN

[3500] Sam Erdheim. Deployment and management with next-generation firewalls. Network Security, 2013 (10):8–12, October 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813701132.

Chappell:2013:PMI

[3501] Brian Chappell. Privilege management — the industry’s best kept secret. Network Security, 2013(10): 12–14, October 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813701144.

Baker:2013:IBT

[3502] Scott Baker. IP-based technology — the future of the security industry. Network Security, 2013 (10):14–17, October 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813701156.

Mansfield-Devine:2013:ISH

[3503] Steve Mansfield-Devine. Interview: Seth Hallem, Mobile Helix — the right way to mobile security. Network Security, 2013(10): 18–20, October 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813701168.

Anonymous:2013:ECj

[3504] Anonymous. Events calendar. Network Security, 2013(10):20, October 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581370117X.

Anonymous:2013:WMF

[3505] Anonymous. Women may be the future of infosecurity. Network Security, 2013(11):1–2, November 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813701181.

Anonymous:2013:SOG

[3506] Anonymous. Security officers get greater voice in boardrooms. Network Security, 2013(11):2, November 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813701223.

Anonymous:2013:Bk

[3507] Anonymous. In brief. Network Security, 2013(11):3, November 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813701235.

Anonymous:2013:ICN

[3508] Anonymous. Introduction to computer and network security. Network Security, 2013(11):4, November 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813701247.

Anonymous:2013:MPT

[3509] Anonymous. Metasploit penetration testing cookbook [Book Review: Mobile Security: How to Secure, Privatize, and Recover Your Devices, by Abhinav Singh. Packt Publishing. ISBN 978-1-84951-742-3]. Network Security, 2013(11):4, November 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813701259.

Macrae:2013:ITR

[3510] Alistair Macrae. Identifying threats in real time. Network Security, 2013 (11):5–8, November 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813701193.

Shaw:2013:DE

[3511] John Shaw. Dealing with encryption. Network Security, 2013(11): 8–11, November 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581370120X.

Allen:2013:IDP

[3512] Phil Allen. The importance of data protection inside your enterprise. Network Security, 2013(11): 12–14, November 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813701211.

Hilbert:2013:LC

[3513] EJ Hilbert. Living with cybercrime. Network Security, 2013(11): 15–17, November 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813701260.

Mansfield-Devine:2013:CTD

[3514] Steve Mansfield-Devine. Colin Tankard, digital pathways: confusion in the cloud. Network Security, 2013 (11):17–20, November 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813701272.

Anonymous:2013:ECk

[3515] Anonymous. Events calendar. Network Security, 2013(11):20, November 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813701284.

Anonymous:2013:FFC

[3516] Anonymous. Firms failing to cope with targeted attacks. Network Security, 2013(12):1–2, December 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813701296.

Anonymous:2013:CRR

[3517] Anonymous. CryptoLocker runs rampant, but drops ransom price. Network Security, 2013(12):2, December 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813701302.

Anonymous:2013:Bl

[3518] Anonymous. In brief. Network Security, 2013(12):3, December 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813701314.

Anonymous:2013:BRM

[3519] Anonymous. Book review: Malware Forensics Field Guide for Windows Systems, James Aquilina, Syngress, ISBN 978-1-59749-472-4. Network Security, 2013(12):4, December 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813701326.

Anonymous:2013:IIS

[3520] Anonymous. Introduction to information security. Network Security, 2013(12):4, December 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813701338.

Johnson:2013:BIS

[3521] Steve Johnson. Bringing IT out of the shadows. Network Security, 2013 (12):5–6, December 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581370134X.

Hipgrave:2013:SFI

[3522] Shaun Hipgrave. Smarter fraud investigations with big data analytics. Network Security, 2013(12): 7–9, December 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813701351.

Meulensteen:2013:CIG

[3523] Marc Meulensteen. Critical infrastructure gets streetwise. Network Security, 2013(12):10–11, December 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813701363.

Mansfield-Devine:2013:JLI

[3524] Steve Mansfield-Devine. John Lyons, ICSPA: Resetting the clock on international co-operation. Network Security, 2013(12):12–15, December 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813701375.

Gold:2013:UDF

[3525] Steve Gold. Understanding the digital fingerprint. Network Security, 2013 (12):15–18, December 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813701387.

Austwick:2013:UOA

[3526] Tim Austwick. Using Oracle Apex securely. Network Security, 2013(12): 19–20, December 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485813701399.

Anonymous:2013:ECl

[3527] Anonymous. Events calendar. Network Security, 2013(12):20, December 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485813701405.

Anonymous:2014:MNR

[3528] Anonymous. More NSA revelations:

backdoors, snooping tools and worldwide reactions. Network Security, 2014 (1):1–2, 19–20, January 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700017.

Anonymous:2014:Ba

[3529] Anonymous. In brief. Network Security, 2014(1):3, January 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700029.

Anonymous:2014:CS

[3530] Anonymous. Cyber-security. Network Security, 2014(1):4, January 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700030.

Anonymous:2014:IIC

[3531] Anonymous. Investigating Internet crimes. Network Security, 2014 (1):4, January 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485814700042.

Watts:2014:PYI

[3532] Steve Watts. Protecting your identity when working remotely. Network Security, 2014(1):5–7, January 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485814700054.

Shteiman:2014:WCP

[3533] Barry Shteiman. Why CMS platforms are breeding security vulnerabilities. Network Security, 2014 (1):7–9, January 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485814700066.

Patterson:2014:INT

[3534] Mike Patterson. An index for network threat detection. Network Security, 2014(1):9–11, January 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485814700078.

Mansfield-Devine:2014:ICN

[3535] Steve Mansfield-Devine. Interview: Corey Nachreiner, WatchGuard — security visibility. Network Security, 2014 (1):11–15, January 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581470008X.

Moulds:2014:GDP

[3536] Richard Moulds. The global data protection conundrum. Network Security, 2014(1):16–17, January 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700091.

Constantine:2014:BDI

[3537] Conrad Constantine. Big data: an information security context. Network Security, 2014(1):18–19, January 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700108.

Anonymous:2014:YAS

[3538] Anonymous. Yahoo ads spread malware. Network Security, 2014 (1):20, January 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581470011X.

Anonymous:2014:CSL

[3539] Anonymous. CryptoLocker success leads to more malware. Network Security, 2014(1):20, January 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700121.

Anonymous:2014:ECa

[3540] Anonymous. Events calendar. Network Security, 2014(1):20, January 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700133.

Anonymous:2014:ERE

[3541] Anonymous. Encryption on the rise, but not enough. Network Security, 2014(2):1–2, February 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700145.

Anonymous:2014:DAR

[3542] Anonymous. DDoS attacks on the rise — by criminals and spies. Network Security, 2014(2):2, February 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700157.

Anonymous:2014:Bb

[3543] Anonymous. In brief. Network Security, 2014(2):3, February 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700169.

Anonymous:2014:MSH

[3544] Anonymous. Mobile security: How to secure, privatize and recover your devices. Network Security, 2014 (2):4, February 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700170.

Anonymous:2014:CIR

[3545] Anonymous. Computer incident response and forensics team management. Network Security, 2014 (2):4, February 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700182.

Kedgley:2014:FIM

[3546] Mark Kedgley. File integrity monitoring in the modern threat landscape. Network Security, 2014 (2):5–8, February 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700194.

Heim:2014:QCD

[3547] Peter Heim. The quest for clarity on data protection and security. Network Security, 2014(2): 8–10, February 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700200.

Allen:2014:MDS

[3548] Stephen Allen. Medical device software under the microscope. Network Security, 2014(2):11–12, February 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700212.

Zhang:2014:VCS

[3549] Hongwen Zhang. A vision for cloud security. Network Security, 2014 (2):12–15, February 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700224.

Gold:2014:SNP

[3550] Steve Gold. In search of a new perimeter. Network Security, 2014 (2):15–20, February 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700236.

Anonymous:2014:ECb

[3551] Anonymous. Events calendar. Network Security, 2014(2):20, February 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700248.

Anonymous:2014:TIM

[3552] Anonymous. Threats increase on mobile platforms — especially Android — as popularity grows. Network Security, 2014(3):1–2, March 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581470025X.

Anonymous:2014:NAW

[3553] Anonymous. Nearly all web and mobile applications have flaws. Network Security, 2014(3):2, March 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700261.

Anonymous:2014:Bc

[3554] Anonymous. In brief. Network Security, 2014(3):3, March 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700273.

Anonymous:2014:BRPa

[3555] Anonymous. Book review: Professional Penetration Testing, Second Edition. Thomas Wilhelm. Syngress. ISBN 978-1-59749-993-4. Network Security, 2014(3):4, March 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700285.

Caldwell:2014:CDF

[3556] Tracey Caldwell. Call the digital fire brigade. Network Security, 2014(3):5–8, March 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700297.

Barnes:2014:UDP

[3557] Pat Barnes. Using DNS to protect networks from threats within. Network Security, 2014(3):9–11, March 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700303.

vanLeeuwen:2014:BYO

[3558] Daniëlle van Leeuwen. Bring your own software. Network Security, 2014 (3):12–13, March 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485814700315.

Bradbury:2014:CWM

[3559] Danny Bradbury. Can we make email secure? Network Security, 2014(3):13–16, March 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485814700327.

Malecki:2014:CNB

[3560] Florian Malecki. The cost of networkbased attacks. Network Security, 2014(3):17–18, March 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485814700339.

Keightley:2014:LXD

[3561] Mike Keightley. The looming XP disaster in industrial environments. Network Security, 2014 (3):18–20, March 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485814700340.

Anonymous:2014:ECc

[3562] Anonymous. Events calendar. Network Security, 2014(3):20, March 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485814700352.

Anonymous:2014:HFL

[3563] Anonymous. ‘Heartbleed’ flaw leaves millions of websites, email servers and other services vulnerable to attack. Network Security, 2014(4):1–2, April 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700364.

Anonymous:2014:WXN

[3564] Anonymous. Windows XP now more vulnerable than ever as Microsoft support ends. Network Security, 2014(4):2, April 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700376.

Anonymous:2014:Bd

[3565] Anonymous. In brief. Network Security, 2014(4):3, April 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700388.

Anonymous:2014:CC

[3566] Anonymous. Cybersecurity and cyberwar. Network Security, 2014(4):4, April 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581470039X.

Brewer:2014:APT

[3567] Ross Brewer. Advanced persistent threats: minimising the damage. Network Security, 2014(4):5–9, April 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700406.

Sverdlove:2014:JVL

[3568] Harry Sverdlove. The Java vulnerability landscape. Network Security, 2014(4):9–14, April 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485814700418.

Bradbury:2014:UDW

[3569] Danny Bradbury. Unveiling the dark web. Network Security, 2014 (4):14–17, April 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581470042X.

Bird:2014:SLS

[3570] Keith Bird. Sandboxing: a line in the sand against malware. Network Security, 2014(4):18–20, April 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485814700431.

Anonymous:2014:ECd

[3571] Anonymous. Events calendar. Network Security, 2014(4):20, April 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485814700443.

Anonymous:2014:HBL

[3572] Anonymous. Heartbleed bug leads to forking and funding. Network Security, 2014(5):1–2, May 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485814700455.

Anonymous:2014:HPB

[3573] Anonymous. High-profile breaches failing to make financial firms tighten their security. Network Security, 2014(5):2, May 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700467.

Anonymous:2014:Be

[3574] Anonymous. In brief. Network Security, 2014(5):3, May 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700479.

Anonymous:2014:DFP

[3575] Anonymous. Digital forensics processing and procedures. Network Security, 2014(5):4, May 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700480.

Maisey:2014:MAL

[3576] Martin Maisey. Moving to analysisled cyber-security. Network Security, 2014(5):5–12, May 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700492.

Southam:2014:DWI

[3577] Mark Southam. DNSSEC: What it is and why it matters. Network Security, 2014(5):12–15, May 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700509.

Inns:2014:EAS

[3578] Jon Inns. The evolution and application of SIEM systems. Network Security, 2014(5):16–17, May 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700510.

Blizzard:2014:DFH

[3579] Sonia Blizzard. Don’t feed the hackers: how your attitude to data security could affect business. Network Security, 2014(5):18–20, May 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700522.

Anonymous:2014:ECe

[3580] Anonymous. Events calendar. Network Security, 2014(5):20, May 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700534.

Anonymous:2014:TPC

[3581] Anonymous. TrueCrypt project cancelled by its developers. Network Security, 2014(6):1–2, June 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700546.

Anonymous:2014:CTC

[3582] Anonymous. Cyber tensions with China heat up after US indicts members of the PLA. Network Security, 2014(6):2, June 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700558.

Anonymous:2014:Bf

[3583] Anonymous. In brief. Network Security, 2014(6):3, June 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581470056X.

Anonymous:2014:TCA

[3584] Anonymous. Targeted cyber-attacks.

Network Security, 2014(6):4, June 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700571.

Radford:2014:CSP

[3585] CJ Radford. Challenges and solutions protecting data within Amazon Web services. Network Security, 2014(6):5–8, June 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700583.

Bradbury:2014:TDB

[3586] Danny Bradbury. Testing the defences of bulletproof hosting companies. Network Security, 2014 (6):8–12, June 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700595.

Gold:2014:CAD

[3587] Steve Gold. Challenges ahead on the digital forensics and audit trails. Network Security, 2014 (6):12–17, June 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485814700601.

Marrison:2014:DAV

[3588] Chris Marrison. DNS as an attack vector — and how businesses can keep it secure. Network Security, 2014(6):17–20, June 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485814700613.

Anonymous:2014:GWB

[3589] Anonymous. Gartner warns of big data security problems. Network Security, 2014(6):20, June 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485814700625.

Anonymous:2014:ECf

[3590] Anonymous. Events calendar. Network Security, 2014(6):20, June 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485814700637.

Anonymous:2014:WEF

[3591] Anonymous. Western energy firms come under concerted and successful cyber-attack. Network Security, 2014(7):1–2, July 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700649.

Anonymous:2014:OFF

[3592] Anonymous. One in five firms hit by APTs. Network Security, 2014 (7):2, 20, July 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700650.

Anonymous:2014:Bg

[3593] Anonymous. In brief. Network Security, 2014(7):3, July 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700662.

Anonymous:2014:PT

[3594] Anonymous. Penetration testing.

Network Security, 2014(7):4, July 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700674.

Bisiaux:2014:DTM

[3595] Jean-Yves Bisiaux. DNS threats and mitigation strategies. Network Security, 2014(7):5–9, July 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700686.

Storey:2014:TNS

[3596] Allen Storey. There’s nothing ‘smart’ about insecure connected devices. Network Security, 2014 (7):9–12, July 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700698.

Bedwell:2014:FNA

[3597] Patrick Bedwell. Finding a new approach to SIEM to suit the SME environment. Network Security, 2014 (7):12–16, July 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700704.

Mansfield-Devine:2014:BS

[3598] Steve Mansfield-Devine. Building in security. Network Security, 2014 (7):16–19, July 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700716.

Anonymous:2014:CET

[3599] Anonymous. Cybercrime and espionage are a threat to Internet itself. Network Security, 2014(7):20, July 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700728.

Anonymous:2014:ECg

[3600] Anonymous. Events calendar. Network Security, 2014(7):20, July 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581470073X.

Anonymous:2014:RGA

[3601] Anonymous. Russian gang amasses biggest-ever haul of stolen credentials, claims security firm. Network Security, 2014(8):1–2, August 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700741.

Anonymous:2014:LSI

[3602] Anonymous. Lack of security in Internet of Things devices. Network Security, 2014(8):2, August 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700753.

Anonymous:2014:Bh

[3603] Anonymous. In brief. Network Security, 2014(8):3, August 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700765.

Anonymous:2014:KLA

[3604] Anonymous. Kali Linux — assuring security by penetration testing. Network Security, 2014(8):4, August 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700777.

Bridge:2014:EPP

[3605] Phil Bridge. EU puts pressure on businesses to erase data. Network Security, 2014(8):5–8, August 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700789.

Tang:2014:GPT

[3606] Andrew Tang. A guide to penetration testing. Network Security, 2014(8):8–11, August 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700790.

Lowe:2014:DAC

[3607] Mark Lowe. Defending against cyber-criminals targeting business websites. Network Security, 2014 (8):11–13, August 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700807.

Mansfield-Devine:2014:CC

[3608] Steve Mansfield-Devine. Not coping with change. Network Security, 2014(8):14–17, August 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700819.

Watts:2014:ICB

[3609] Steve Watts. Intelligent combination — the benefits of tokenless two-factor authentication. Network Security, 2014 (8):17–20, August 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700820.

Anonymous:2014:ECh

[3610] Anonymous. Events calendar. Network Security, 2014(8):20, August 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700832.

Anonymous:2014:NAN

[3611] Anonymous. NATO adopts new policy calling for collective defence against cyber-attacks. Network Security, 2014 (9):1–2, September 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700844.

Anonymous:2014:ODW

[3612] Anonymous. ‘One-day wonder’ websites used to launch attacks. Network Security, 2014(9):2, September 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700856.

Anonymous:2014:Bi

[3613] Anonymous. In brief. Network Security, 2014(9):3, September 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700868.

Anonymous:2014:BRPb

[3614] Anonymous. Book review: Python Forensics, Chet Hosmer, Syngress. ISBN 978-0-12-418676-7 (print), 978-0-12-418683-5 (e-book). Network Security, 2014(9):4, September 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348581470087X.

Anonymous:2014:BIS

[3615] Anonymous. The basics of information security. Network Security, 2014(9):4, September 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485814700881.

Saunders:2014:PAE

[3616] Simon Saunders. Protecting against espionage. Network Security, 2014 (9):5–7, September 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485814700893.

Beckett:2014:BPP

[3617] Phil Beckett. BYOD — popular and problematic. Network Security, 2014 (9):7–9, September 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581470090X.

Lockner:2014:DYK

[3618] Julie Lockner. Do you know where your sensitive data is kept? Network Security, 2014(9):10–12, September 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485814700911.

Mansfield-Devine:2014:HIS

[3619] Steve Mansfield-Devine. Hacking on an industrial scale. Network Security, 2014 (9):12–16, September 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700923.

Gonda:2014:UTS

[3620] Oded Gonda. Understanding the threat to SCADA networks. Network Security, 2014(9):17–18, September 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700935.

Andresen:2014:OSF

[3621] Lasse Andresen. Open sourcing the future of IAM. Network Security, 2014 (9):18–20, September 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814700947.

Anonymous:2014:ECi

[3622] Anonymous. Events calendar. Network Security, 2014(9):20, September 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700959.

Anonymous:2014:SBS

[3623] Anonymous. Shellshock bug has security experts in a panic and hackers searching for exploits. Network Security, 2014(10):1–2, October 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814700960.

Anonymous:2014:SCM

[3624] Anonymous. Spam creates major disruption. Network Security, 2014(10):2, October 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485814700972.

Anonymous:2014:Bj

[3625] Anonymous. In brief. Network Security, 2014(10):3, October 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485814700984.

Anonymous:2014:PNS

[3626] Anonymous. The practice of network security monitoring. Network Security, 2014(10):4, October 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485814700996.

Brazil:2014:SMM

[3627] Jody Brazil. Security metrics to manage change. Network Security, 2014 (10):5–7, October 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581470100X.

Boldyreva:2014:MEW

[3628] Alexandra Boldyreva and Paul Grubbs. Making encryption work in the cloud. Network Security, 2014 (10):8–10, October 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814701011.

Bradbury:2014:APG

[3629] Danny Bradbury. Anonymity and privacy: a guide for the perplexed. Network Security, 2014 (10):10–14, October 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814701023.

Martini:2014:SAW

[3630] Paul Martini. A secure approach to wearable technology. Network Security, 2014(10):15–17, October 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814701035.

Mansfield-Devine:2014:MSD

[3631] Steve Mansfield-Devine. Masking sensitive data. Network Security, 2014 (10):17–20, October 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814701047.

Anonymous:2014:ECj

[3632] Anonymous. Events calendar. Network Security, 2014(10):20, October 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814701059.

Anonymous:2014:ITB

[3633] Anonymous. Insider threat becoming more significant, says Kaspersky. Network Security, 2014(11): 1–2, November 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814701060.

Anonymous:2014:FFI

[3634] Anonymous. Four-fold increase in DDoS attacks. Network Security, 2014(11):2, November 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814701072.

Anonymous:2014:Bk

[3635] Anonymous. In brief. Network Security, 2014(11):3, November 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814701084.

Anonymous:2014:WFA

[3636] Anonymous. Windows forensic analysis toolkit. Network Security, 2014 (11):4, November 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814701096.

Anonymous:2014:SEP

[3637] Anonymous. Social engineering penetration testing. Network Security, 2014(11):4, November 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814701102.

Munro:2014:ASA

[3638] Ken Munro. Android scraping: accessing personal data on mobile devices. Network Security, 2014(11): 5–9, November 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814701114.

Taylor:2014:FIS

[3639] Mark Taylor, John Haggerty, David Gresty, Peter Almond, and Tom Berry. Forensic investigation of social networking applications. Network Security, 2014(11):9–16, November 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814701126.

Mansfield-Devine:2014:MSA

[3640] Steve Mansfield-Devine. Mobile security: it’s all about behaviour. Network Security, 2014(11):16–20, November 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814701138.

Anonymous:2014:RGL

[3641] Anonymous. Russian Government likely behind APTs, says Fire-Eye. Network Security, 2014(11): 20, November 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581470114X.

Anonymous:2014:ECk

[3642] Anonymous. Events calendar. Network Security, 2014(11):20, November 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485814701151.

Anonymous:2014:SSM

[3643] Anonymous. Sony suffers major data breach as attackers leak files and destroy hard drives. Network Security, 2014(12):1–2, December 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485814701163.

Anonymous:2014:RSS

[3644] Anonymous. Is Regin the son of Stuxnet? Network Security, 2014 (12):2, December 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485814701175.

Anonymous:2014:Bl

[3645] Anonymous. In brief. Network Security, 2014(12):3, December 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485814701187.

Anonymous:2014:HPT

[3646] Anonymous. Hacking and penetration testing with low power devices. Network Security, 2014 (12):4, December 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485814701199.

Tipping:2014:RTV

[3647] David Tipping. The rising threats from Voice over IP. Network Security, 2014 (12):5–6, December 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814701205.

Kalra:2014:TAE

[3648] Gursev Singh Kalra. Threat analysis of an enterprise messaging system. Network Security, 2014(12): 7–13, December 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814701217.

Jones:2014:TDD

[3649] Robert Jones and Tony Dearsley.

Time for a data detox. Network Security, 2014(12):13–15, December 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814701229.

Goldberg:2014:TUT

[3650] Joe Goldberg. Tackling unknown threats. Network Security, 2014(12): 16–17, December 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814701230.

Lane:2014:CDS

[3651] Ashley Lane. Cross domain solutions — and why they matter. Network Security, 2014(12): 18–20, December 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485814701242.

Anonymous:2014:ECl

[3652] Anonymous. Events calendar. Network Security, 2014(12):20, December 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485814701254.

Anonymous:2015:DRW

[3653] Anonymous. Doubts remain over whether North Korea was responsible for massive hack of Sony Pictures. Network Security, 2015 (1):1–2, January 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815700012.

Anonymous:2015:TUA

[3654] Anonymous. Tor under attack, again. Network Security, 2015 (1):2, January 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815700024.

Anonymous:2015:SG

[3655] Anonymous. Steve Gold — 1956– 2015. Network Security, 2015 (1):3, January 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815700036.

Anonymous:2015:Ba

[3656] Anonymous. In brief. Network Security, 2015(1):3, January 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815700048.

Anonymous:2015:BRP

[3657] Anonymous. Book review: Penetration Testing with the Bash Shell, by Keith Makan. Packt Publishing. ISBN 978-1-84969-510-7. Network Security, 2015(1):4, January 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581570005X.

Anonymous:2015:BRL

[3658] Anonymous. Book review: Learning PHP, MySQL, JavaScript, CSS & HTML5, by Robin Nixon. Third Edition. O’Reilly. ISBN 978-1-4919-4946-7. Network Security, 2015 (1):4, January 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815700061.

Densham:2015:TCS

[3659] Ben Densham. Three cyber-security strategies to mitigate the impact of a data breach. Network Security, 2015(1):5–8, January 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815700073.

Patel:2015:NPC

[3660] Ashish Patel. Network performance without compromising security. Network Security, 2015 (1):9–12, January 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485815700085.

Lafuente:2015:BDS

[3661] Guillermo Lafuente. The big data security challenge. Network Security, 2015(1):12–14, January 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485815700097.

Mansfield-Devine:2015:MIC

[3662] Steve Mansfield-Devine. Managing identity for a competitive edge. Network Security, 2015(1): 14–18, January 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485815700103.

Zhang:2015:BYO

[3663] Hongwen Zhang. Bring your own encryption: balancing security with practicality. Network Security, 2015 (1):18–20, January 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485815700115.

Anonymous:2015:ECa

[3664] Anonymous. Events calendar. Network Security, 2015(1):20, January 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815700127.

Anonymous:2015:EOS

[3665] Anonymous. European organisations are still not ready for forthcoming data protection rules. Network Security, 2015(2):1–2, February 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300015.

Anonymous:2015:LHA

[3666] Anonymous. Linux hit by another long-term flaw. Network Security, 2015(2):2, February 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300027.

Anonymous:2015:Bb

[3667] Anonymous. In brief. Network Security, 2015(2):3, February 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300039.

Anonymous:2015:BRK

[3668] Anonymous. Book reviewKali Linux Network Scanning Cookbook, by Justin Hutchens. Packt Publishing. ISBN 978-1-78398-214-1. Network Security, 2015(2):4, February 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300040.

Mansfield-Devine:2015:PD

[3669] Steve Mansfield-Devine. The privacy dilemma. Network Security, 2015 (2):5–10, February 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485815300052.

Nielsen:2015:ICK

[3670] Poul Nielsen. The importance of context in keeping end users secure. Network Security, 2015(2): 10–13, February 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485815300064.

Green:2015:SAC

[3671] John Green. Staying ahead of cyberattacks. Network Security, 2015 (2):13–16, February 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485815300076.

Millard:2015:CMT

[3672] Gavin Millard. Continuous monitoring for transient devices. Network Security, 2015(2):16–18, February 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485815300088.

Beckett:2015:IAS

[3673] Phil Beckett. An intelligent approach to security. Network Security, 2015 (2):18–20, February 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581530009X.

Anonymous:2015:ECb

[3674] Anonymous. Events calendar. Network Security, 2015(2):20, February 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300106.

Anonymous:2015:MSL

[3675] Anonymous. More Snowden leaks reveal hacking by NSA and GCHQ against communications firm. Network Security, 2015(3):1–2, March 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300118.

Anonymous:2015:NTH

[3676] Anonymous. NSA also targets hard drives. Network Security, 2015(3):2, March 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581530012X.

Anonymous:2015:Bc

[3677] Anonymous. In brief. Network Security, 2015(3):3, March 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300131.

Anonymous:2015:BRI

[3678] Anonymous. Book review: Industrial Network Security, by Eric Knapp and Joel Langill. Syngress. ISBN 978-0-12-420114-9. Network Security, 2015(3):4, March 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300143.

Anonymous:2015:BRGa

[3679] Anonymous. Book review: Google Earth Forensics, by Michael Harrington and Michael Cross. Syngress. ISBN 978-0-12-800216-2. Network Security, 2015(3):4, March 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300155.

Tankard:2015:HSY

[3680] Colin Tankard. How secure is your building? Network Security, 2015(3):5–8, March 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300167.

Cates:2015:ESI

[3681] Sol Cates. The evolution of security intelligence. Network Security, 2015(3):8–10, March 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300179.

Everett:2015:SDN

[3682] Cath Everett. Should the dark net be taken out? Network Security, 2015(3):10–13, March 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300180.

Buesnel:2015:TSN

[3683] Guy Buesnel. Threats to satellite navigation systems. Network Security, 2015(3):14–18, March 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300192.

Newe:2015:DIT

[3684] Gary Newe. Delivering the Internet of Things. Network Security, 2015(3):18–20, March 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300209.

Anonymous:2015:ECc

[3685] Anonymous. Events calendar. Network Security, 2015(3):20, March 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300210.

Anonymous:2015:MVO

[3686] Anonymous. Mobile vulnerabilities are opening the door to cyber-attacks on corporates. Network Security, 2015(4):1–2, April 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300222.

Anonymous:2015:BAA

[3687] Anonymous. British Airways among latest breaches. Network Security, 2015(4):2, 20, April 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300234.

Anonymous:2015:Bd

[3688] Anonymous. In brief. Network Security, 2015(4):3, April 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300246.

Anonymous:2015:BRBa

[3689] Anonymous. Book review: Black Hat Python, Justin Seitz. No Starch Press. ISBN 978-1-59327-590-7. Network Security, 2015(4):4, April 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300258.

Clay:2015:MTR

[3690] Peter Clay. A modern threat response framework. Network Security, 2015(4):5–10, April 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581530026X.

Kedgley:2015:IYC

[3691] Mark Kedgley. If you can’t stop the breach, at least spot the breach. Network Security, 2015 (4):11–12, April 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300271.

Auty:2015:AAP

[3692] Mike Auty. Anatomy of an advanced persistent threat. Network Security, 2015(4):13–16, April 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300283.

Amigorena:2015:RKU

[3693] François Amigorena. Relevance is the key to users’ security understanding. Network Security, 2015 (4):17–18, April 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300295.

Beckett:2015:GYB

[3694] Phil Beckett. Getting your back-up data back up to date. Network Security, 2015(4):18–20, April 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300301.

Anonymous:2015:ECd

[3695] Anonymous. Events calendar. Network Security, 2015(4):20, April 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300313.

Anonymous:2015:MMG

[3696] Anonymous. Mobile malware goes straight for the money, says Kaspersky. Network Security, 2015(5):1–2, May 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300325.

Anonymous:2015:DAG

[3697] Anonymous. DDoS attacks grow. Network Security, 2015(5):2, 20, May 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300337.

Anonymous:2015:Be

[3698] Anonymous. In brief. Network Security, 2015(5):3, May 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300349.

Anonymous:2015:BRBb

[3699] Anonymous. Book review: Building an Information Security Awareness Program, Bill Gardner and Valerie Thomas. Syngress. ISBN 978-0-12-419967-5. Network Security, 2015(5):4, May 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300350.

Anonymous:2015:BRGb

[3700] Anonymous. Book review: The GNU Make Book, John Graham-Cumming. No Starch Press. ISBN 978-1-59327-649-2. Network Security, 2015(5):4, May 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300362.

Brewer:2015:CTR

[3701] Ross Brewer. Cyber threats: reducing the time to detection and response. Network Security, 2015(5):5–8, May 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300374.

Tankard:2015:DCF

[3702] Colin Tankard. Data classification — the foundation of information security. Network Security, 2015(5):8–11, May 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300386.

Caldwell:2015:HGH

[3703] Tracey Caldwell. Hacktivism goes hardcore. Network Security, 2015 (5):12–17, May 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300398.

German:2015:DUV

[3704] Paul German. The dangers of unsecured voice. Network Security, 2015(5):18–19, May 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300404.

Anonymous:2015:IMR

[3705] Anonymous. IoT multiplies risk of attack. Network Security, 2015(5): 20, May 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1353485815300416.

Anonymous:2015:ECe

[3706] Anonymous. Events calendar. Network Security, 2015(5):20, May 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485815300428.

Anonymous:2015:DLS

[3707] Anonymous. Darknet less secure than many think, say researchers. Network Security, 2015 (6):1–2, June 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581530043X.

Anonymous:2015:LLL

[3708] Anonymous. Logjam is latest long-term flaw. Network Security, 2015(6):2, June 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485815300441.

Anonymous:2015:Bf

[3709] Anonymous. In brief. Network Security, 2015(6):3, June 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485815300453.

Anonymous:2015:BRAa

[3710] Anonymous. Book review: Android Security Internals, Nikolay Elenkov. No Starch Press. ISBN 978-1-59327-581-5. Network Security, 2015(6):4, June 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300465.

Ring:2015:KTT

[3711] Tim Ring. Keeping tabs on tracking technology. Network Security, 2015(6):5–8, June 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300477.

Hooson:2015:SYD

[3712] Stuart Hooson. Smarten your data security before new EU legislation or risk corporate loss. Network Security, 2015(6):8–10, June 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300489.

Stange:2015:DMA

[3713] Szilard Stange. Detecting malware across operating systems. Network Security, 2015(6):11–14, June 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300490.

Hald:2015:RRA

[3714] David Hald and Alex Udakis. Rethinking remote authentication: time to kiss tokens goodbye? Network Security, 2015(6):15–17, June 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485815300507.

Huard:2015:DQP

[3715] Boris Huard. The data quality paradox. Network Security, 2015 (6):18–20, June 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485815300519.

Anonymous:2015:ECf

[3716] Anonymous. Events calendar. Network Security, 2015(6):20, June 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485815300520.

Anonymous:2015:SSF

[3717] Anonymous. Surveillance software firm Hacking Team is hacked. Network Security, 2015(7):1–2, July 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485815300532.

Anonymous:2015:ESS

[3718] Anonymous. Europol sets sights on IS. Network Security, 2015(7):2, July 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485815300544.

Anonymous:2015:Bg

[3719] Anonymous. In brief. Network Security, 2015(7):3, July 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300556.

Anonymous:2015:BRM

[3720] Anonymous. Book review: The Mobile Application Hacker’s Handbook, by Dominic Chell, Tyrone Erasmus, Shaun Colley, and Ollie Whitehouse Wiley. ISBN 978-1-118-95850-6. Network Security, 2015(7):4, July 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300568.

Kingswood:2015:CCW

[3721] Matt Kingswood. Climate change will require more agile business continuity planning. Network Security, 2015(7):5–10, July 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581530057X.

Ring:2015:CSF

[3722] Tim Ring. Cloud security fears:

fact or FUD? Network Security, 2015(7):10–14, July 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300581.

McKenna:2015:HFA

[3723] Matthew McKenna. How federal agencies can secure their cloud migrations. Network Security, 2015 (7):14–16, July 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300593.

Larson:2015:ITW

[3724] Dave Larson and Stephen Gates. Is it true that what you can’t see can’t hurt you? Network Security, 2015(7):17–18, July 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581530060X.

Watts:2015:NDP

[3725] Steve Watts. NFC and 2FA: the death of the password? Network Security, 2015(7):19–20, July 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485815300611.

Anonymous:2015:ECg

[3726] Anonymous. Events calendar. Network Security, 2015(7):20, July 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485815300623.

Anonymous:2015:ASV

[3727] Anonymous. Android stagefright vulnerability threatens all devices — and fixing it isn’t that easy. Network Security, 2015(8):1–2, August 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485815300635.

Anonymous:2015:UPS

[3728] Anonymous. UK public sector weak on security. Network Security, 2015(8):2, August 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300647.

Anonymous:2015:Bh

[3729] Anonymous. In brief. Network Security, 2015(8):3, August 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300659.

SM-D:2015:BRH

[3730] SM-D. Book review: Hacking Web Intelligence, [by Sudhanshu Chauhan and Nutan Kumar Panda. Syngress. ISBN 978-0-12-801867-5]. Network Security, 2015(8):4, August 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300660.

SM-D:2015:BRB

[3731] SM-D. Book review: The Book of GNS3 [by Jason Neumann. No Starch Press. ISBN 978-1-59327-554-9]. Network Security, 2015 (8):4, August 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300672.

Smyth:2015:CSF

[3732] Vincent Smyth. Cyber-security fortresses built on quicksand. Network Security, 2015(8):5–8, August 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300684.

Simmonds:2015:DII

[3733] Paul Simmonds. The digital identity issue. Network Security, 2015 (8):8–13, August 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485815300696.

Richter:2015:MYD

[3734] Chris Richter. Managing your data risk: back to basics. Network Security, 2015(8):13–15, August 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485815300702.

Knights:2015:MID

[3735] Ricky Knights and Emma Morris. Move to intelligence-driven security. Network Security, 2015 (8):15–18, August 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485815300714.

Anstee:2015:PTT

[3736] Darren Anstee. Preparing for tomorrow’s threat landscape. Network Security, 2015(8):18–20, August 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485815300726.

Anonymous:2015:ECh

[3737] Anonymous. Events calendar. Network Security, 2015(8):20, August 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300738.

Anonymous:2015:DAR

[3738] Anonymous. DDoS attacks on the rise and being used to mask other, more significant breaches. Network Security, 2015(9):1–2, September 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581530074X.

Anonymous:2015:MHD

[3739] Anonymous. Malvertising hits dating websites. Network Security, 2015 (9):2, September 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300751.

Anonymous:2015:Bi

[3740] Anonymous. In brief. Network Security, 2015(9):3, September 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300763.

Anonymous:2015:BRSa

[3741] Anonymous. Book review: Securing SQL Server, by Denny Cherry. Syngress. ISBN 978-0-12-801275-8. Network Security, 2015(9):4, September 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300775.

Anonymous:2015:BRDa

[3742] Anonymous. Book review: Digital Identity Management, Maryline Laurent and Samia Bouzefrane. ISTE Press/Elsevier. ISBN 978-1-78548-004-1. Network Security, 2015 (9):4, September 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300787.

Bocek:2015:HEP

[3743] Kevin Bocek. Is HTTPS enough to protect governments? Network Security, 2015(9):5–8, September 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300799.

Mansfield-Devine:2015:AMA

[3744] Steve Mansfield-Devine. The Ashley Madison affair. Network Security, 2015 (9):8–16, September 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300805.

Kirk:2015:CFI

[3745] Richard Kirk. Cars of the future: the Internet of Things in the automotive industry. Network Security, 2015 (9):16–18, September 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300817.

Rozanski:2015:UCD

[3746] Shahaf Rozanski. Using cloud data to accelerate forensic investigations. Network Security, 2015(9): 19–20, September 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300829.

Anonymous:2015:ECi

[3747] Anonymous. Events calendar. Network Security, 2015(9):20, September 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300830.

Anonymous:2015:HOS

[3748] Anonymous. Healthcare organisations struggle to maintain security. Network Security, 2015 (10):1–2, October 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300842.

Anonymous:2015:CC

[3749] Anonymous. The cost of cybercrime. Network Security, 2015 (10):2, October 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300854.

Anonymous:2015:Bj

[3750] Anonymous. In brief. Network Security, 2015(10):3, October 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300866.

Anonymous:2015:BRSb

[3751] Anonymous. Book review: Success Strategies from Women in STEM, Edited by Peggy A Pritchard and Christine Grant. Second edition. Academic Press. ISBN 978-0-12-397181-4. Network Security, 2015(10):4, October 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300878.

Anonymous:2015:BRDb

[3752] Anonymous. Book review: Doing Math With Python, Amit Saha. No Starch Press. ISBN 978-1-59327-640-9. Network Security, 2015 (10):4, October 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581530088X.

Hof:2015:AND

[3753] Jan Hof. Addressing new demands on network security. Network Security, 2015(10):5–7, October 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300891.

Marrison:2015:UTD

[3754] Chris Marrison. Understanding the threats to DNS and how to secure it. Network Security, 2015 (10):8–10, October 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300908.

Harrison:2015:RCS

[3755] Reuven Harrison. Reducing complexity in securing heterogeneous networks. Network Security, 2015(10):

11–13, October 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581530091X.

Mansfield-Devine:2015:GED

[3756] Steve Mansfield-Devine. The growth and evolution of DDoS. Network Security, 2015(10):13–20, October 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300921.

Anonymous:2015:ECj

[3757] Anonymous. Events calendar. Network Security, 2015(10):20, October 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300933.

Anonymous:2015:UGB

[3758] Anonymous. UK Government battles tech firms over encryption. Network Security, 2015(11):1–2, November 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300945.

Anonymous:2015:RDN

[3759] Anonymous. Ransomware defeated but new forms emerge. Network Security, 2015(11):2, November 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300957.

Anonymous:2015:Bk

[3760] Anonymous. In brief. Network Security, 2015(11):3, November 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300969.

Anonymous:2015:BRH

[3761] Anonymous. Book review: How Software Works, by V Anton Spraul. No Starch Press. ISBN 978-1-59327-666-9. Network Security, 2015 (11):4, November 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300970.

Mansfield-Devine:2015:WAT

[3762] Steve Mansfield-Devine. When advertising turns nasty. Network Security, 2015(11):5–8, November 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815300982.

German:2015:CSC

[3763] Paul German. Counting the security cost of cheap calls. Network Security, 2015(11):9–11, November 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815300994.

Ring:2015:CCN

[3764] Tim Ring. Connected cars — the next target for hackers. Network Security, 2015(11):11–16, November 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815301008.

Beckett:2015:BRU

[3765] Phil Beckett. The business risks of using smartphones. Network Security, 2015(11):16–17, November 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581530101X.

Maule-ffinch:2015:KTI

[3766] Bradley Maule-ffinch. Key trends in information security. Network Security, 2015(11):18–20, November 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815301021.

Anonymous:2015:ECk

[3767] Anonymous. Events calendar. Network Security, 2015(11):20, November 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815301033.

Anonymous:2015:FUC

[3768] Anonymous. A fifth of UK consumers are victims of theft via data breaches. Network Security, 2015 (12):1–2, December 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815301045.

Anonymous:2015:FFW

[3769] Anonymous. Four out of five web apps fail key tests. Network Security, 2015 (12):2, 20, December 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815301057.

Anonymous:2015:Bl

[3770] Anonymous. In brief. Network Security, 2015(12):3, December 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485815301069.

Anonymous:2015:BRC

[3771] Anonymous. Book review: Cocoa Programming for OS X, by Aaron Hillegass, Adam Preble and Nate Chandler. Fifth edition. Big Nerd Ranch. ISBN 978-0-13-407695-9. Network Security, 2015(12):4, December 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815301070.

Anonymous:2015:BRAb

[3772] Anonymous. Book review: Android Programming, by Bill Phillips, Chris Stewart, Brian Hardy & Kristin Marsicano. Second edition, Big Nerd Ranch. ISBN 978-0-13-417145-6. Network Security, 2015 (12):4, December 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815301082.

Kirk:2015:TSN

[3773] Richard Kirk. Threat sharing — a neighbourhood watch for security practitioners. Network Security, 2015 (12):5–7, December 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815301094.

Caldwell:2015:TAD

[3774] Tracey Caldwell. Taking agile development beyond software — what are the security risks? Network Security, 2015 (12):8–11, December 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815301100.

Adato:2015:NOP

[3775] Leon Adato. Network outages — pick your battles. Network Security, 2015 (12):12–13, December 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815301112.

Mansfield-Devine:2015:UR

[3776] Steve Mansfield-Devine. Under the radar. Network Security, 2015(12): 14–18, December 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815301124.

Watts:2015:HGA

[3777] Steve Watts. The holy grail of authentication. Network Security, 2015 (12):18–19, December 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485815301136.

Anonymous:2015:WRD

[3778] Anonymous. Wide range of devices vulnerable to hacking. Network Security, 2015(12):20, December 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485815301148.

Anonymous:2015:ECl

[3779] Anonymous. Events calendar. Network Security, 2015(12):20, December 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348581530115X.

Anonymous:2016:JFC

[3780] Anonymous. Juniper firewalls contain deliberate weakness that provides back door. Network Security, 2016(1):1–2, January 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300010.

Anonymous:2016:MBE

[3781] Anonymous. More battles over encryption & surveillance. Network Security, 2016(1):2, January 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300022.

Anonymous:2016:Ba

[3782] Anonymous. In brief. Network Security, 2016(1):3, January 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300034.

Anonymous:2016:BRH

[3783] Anonymous. Book review: How to Attack and Defend Your Website, by Henry Dalziel. Syngress. ISBN 978-0-12-802754-7. Network Security, 2016(1):4, January 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300046.

Anonymous:2016:BRS

[3784] Anonymous. Book review: Securing Social Media in the Enterprise, by Henry Dalziel. Syngress. ISBN 978-0-12-804180-2. Network Security, 2016(1):4, January 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300058.

Shaulov:2016:BMS

[3785] Michael Shaulov. Bridging mobile security gaps. Network Security, 2016(1):5–8, January 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581630006X.

Fimin:2016:BBA

[3786] Michael Fimin. Breaking bad:

avoiding the 10 worst IT admin habits. Network Security, 2016 (1):8–11, January 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300071.

Touchette:2016:EM

[3787] Fred Touchette. The evolution of malware. Network Security, 2016 (1):11–14, January 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300083.

Wilkinson:2016:CDI

[3788] William Wilkinson. Cyber discovery in investigations. Network Security, 2016(1):15–17, January 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300095.

King:2016:KCS

[3789] Jason King and Dan Evans. Key criteria for selecting a secure cloud wireless net work. Network Security, 2016 (1):17–20, January 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300101.

Anonymous:2016:ECa

[3790] Anonymous. Events calendar. Network Security, 2016(1):20, January 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485816300113.

Anonymous:2016:IIP

[3791] Anonymous. Impending Investigatory Powers Bill comes under fire. Network Security, 2016(2): 1–2, February 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300125.

Anonymous:2016:FWP

[3792] Anonymous. Firms willing to pay ransoms. Network Security, 2016 (2):2, 20, February 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300137.

Anonymous:2016:Bb

[3793] Anonymous. In brief. Network Security, 2016(2):3, February 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485816300149.

Anonymous:2016:BRI

[3794] Anonymous. Book review: iOS Application Security, by David Thiel. No Starch Press. ISBN 978-1-59327-601-0. Network Security, 2016 (2):4, February 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300150.

Muscat:2016:WVI

[3795] Ian Muscat. Web vulnerabilities: identifying patterns and remedies. Network Security, 2016(2): 5–10, February 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300162.

Everett:2016:PFD

[3796] Cath Everett. Are passwords finally dying? Network Security, 2016 (2):10–14, February 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300174.

Mansfield-Devine:2016:SGB

[3797] Steve Mansfield-Devine. Security guarantees: building credibility for security vendors. Network Security, 2016 (2):14–18, February 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300186.

Ganesan:2016:SSP

[3798] Rajesh Ganesan. Stepping up security with password management control. Network Security, 2016(2): 18–19, February 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300198.

Anonymous:2016:ESF

[3799] Anonymous. Energy sector firms face physically damaging attacks. Network Security, 2016(2):20, February 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485816300204.

Anonymous:2016:ECb

[3800] Anonymous. Events calendar. Network Security, 2016(2):20, February 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485816300216.

Anonymous:2016:REA

[3801] Anonymous. Ransomware expands, attacks hospitals and local authorities, and moves to new platforms. Network Security, 2016 (3):1–2, March 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300228.

Anonymous:2016:Bc

[3802] Anonymous. In brief. Network Security, 2016(3):3, March 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581630023X.

SM-D:2016:BRA

[3803] SM-D. Book review: Automating Open Source Intelligence [edited by Robert Layton and Paul Watters, Syngress. ISBN 978-0-12-802916-9]. Network Security, 2016(3):4, March 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300241.

SM-D:2016:BRI

[3804] SM-D. Book review: Infosec Management Fundamentals, Henry Dalziel [Syngress. ISBN 978-0-12-804172-7]. Network Security, 2016(3):4, March 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485816300253.

Larson:2016:DDS

[3805] Dave Larson. Distributed denial of service attacks — holding back the flood. Network Security, 2016 (3):5–7, March 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300265.

Gross:2016:DDB

[3806] Garrett Gross. Detecting and destroying botnets. Network Security, 2016(3):7–10, March 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300277.

Haughey:2016:ANF

[3807] Hamish Haughey, Gregory Epiphaniou, and Haider M. Al-Khateeb. Anonymity networks and the fragile cyber ecosystem. Network Security, 2016(3):10–18, March 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300289.

German:2016:NMN

[3808] Paul German. A new month, a new data breach. Network Security, 2016(3):18–20, March 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300290.

Anonymous:2016:ECc

[3809] Anonymous. Events calendar. Network Security, 2016(3):20, March 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485816300307.

Anonymous:2016:HBM

[3810] Anonymous. Hospitals become major target for ransomware. Network Security, 2016(4):1–2, April 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300319.

Anonymous:2016:CAF

[3811] Anonymous. Car alarm at the FBI.

Network Security, 2016(4):2, April 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485816300320.

Anonymous:2016:Bd

[3812] Anonymous. In brief. Network Security, 2016(4):3, April 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300332.

Anonymous:2016:BRCa

[3813] Anonymous. Book review: The Car Hacker’s Handbook, by Craig Smith. No Starch Press. ISBN 978-1-59327-703-1. Network Security, 2016(4):4, April 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300344.

Anonymous:2016:BRN

[3814] Anonymous. Book review: Next Generation Red Teaming, by Henry Dalziel. Syngress. ISBN 978-0-12-804171-0. Network Security, 2016(4):4, April 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485816300356.

Furnell:2016:VMP

[3815] Steve Furnell. Vulnerability management: not a patch on where we should be? Network Security, 2016(4):5–9, April 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300368.

Harwood:2016:LPG

[3816] Will Harwood. Locking up passwords — for good. Network Security, 2016(4):10–13, April 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581630037X.

Mansfield-Devine:2016:YLY

[3817] Steve Mansfield-Devine. Your life in your hands: the security issues with healthcare apps. Network Security, 2016(4):14–18, April 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300381.

Rozanski:2016:PTC

[3818] Shahaf Rozanski. Peering through the cloud. Network Security, 2016 (4):19–20, April 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300393.

Anonymous:2016:ECd

[3819] Anonymous. Events calendar. Network Security, 2016(4):20, April 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581630040X.

Anonymous:2016:UGS

[3820] Anonymous. UK Government says two-thirds of firms under attack. Network Security, 2016(5):1–2, May 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485816300411.

Anonymous:2016:OTB

[3821] Anonymous. Outside threats are the biggest concern. Network Security, 2016(5):2, May 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300423.

Anonymous:2016:Be

[3822] Anonymous. In brief. Network Security, 2016(5):3, May 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300435.

Anonymous:2016:BRA

[3823] Anonymous. Book review: Automated Security Analysis of Android and iOS Applications with Mobile Security Framework, by Henry Dalziel and Ajin Abraham. Syngress. ISBN 978-0-12-805105-4. Network Security, 2016(5):4, May 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300447.

Anonymous:2016:BRCb

[3824] Anonymous. Book review: CISSP Study Guide, by Eric Conrad, Seth Misenar and Joshua Feldman. Third edition. Syngress. ISBN 978-0-12-802437-

9. Network Security, 2016(5):4, May 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485816300459.

Grimm:2016:PCU

[3825] John Grimm. PKI: crumbling under the pressure. Network Security, 2016(5):5–7, May 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300460.

Taylor:2016:AMS

[3826] Mark Taylor, John Haggerty, David Gresty, Chris Wren, and Tom Berry. Avoiding the misuse of social media by employees. Network Security, 2016(5):8–11, May 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300472.

Pittenger:2016:KYO

[3827] Mike Pittenger. Know your open source code. Network Security, 2016 (5):11–15, May 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300484.

Mansfield-Devine:2016:CSO

[3828] Steve Mansfield-Devine. Creating security operations centres that work. Network Security, 2016 (5):15–18, May 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300496.

Liu:2016:ABN

[3829] Cricket Liu. Actively boosting network security with passive DNS. Network Security, 2016(5):18–20, May 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485816300502.

Anonymous:2016:ECe

[3830] Anonymous. Events calendar. Network Security, 2016(5):20, May 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485816300514.

Anonymous:2016:MUC

[3831] Anonymous. Millions of user credentials for popular sites sold on dark markets. Network Security, 2016(6):1–2, June 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300526.

Anonymous:2016:Bf

[3832] Anonymous. In brief. Network Security, 2016(6):3, June 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300538.

Anonymous:2016:BRMa

[3833] Anonymous. Book review: Mobile Data Loss: Threats and Countermeasures, by Michael Raggo. Syngress. ISBN 978-0-12-802864-3. Network Security, 2016(6):4, June 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581630054X.

Anonymous:2016:BRT

[3834] Anonymous. Book review: Traffic Anomaly Detection, by Antonio Cuadra-Sánchez and Javier Aracil. Iste Press. ISBN 978-1-78548-012-6. Network Security, 2016(6):4, June 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300551.

Tankard:2016:WGM

[3835] Colin Tankard. What the GDPR means for businesses. Network Security, 2016(6):5–8, June 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300563.

Hall:2016:WPK

[3836] Mark Hall. Why people are key to cyber-security. Network Security, 2016(6):9–10, June 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300575.

Mansfield-Devine:2016:BP

[3837] Steve Mansfield-Devine. The battle for privacy. Network Security, 2016(6):11–15, June 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300587.

Dixit:2016:HFB

[3838] Sameer Dixit. Holding the fort:

a business case for testing security. Network Security, 2016 (6):16–18, June 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300599.

German:2016:SSF

[3839] Paul German. The SIP security fallacy. Network Security, 2016 (6):18–20, June 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300605.

Anonymous:2016:ECf

[3840] Anonymous. Events calendar. Network Security, 2016(6):20, June 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485816300617.

Anonymous:2016:ALB

[3841] Anonymous. Authorities losing the battle against cybercrime, says UK National Crime Agency. Network Security, 2016(7):1–2, July 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300629.

Anonymous:2016:PBT

[3842] Anonymous. Police breached thousands of times. Network Security, 2016(7):2, July 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300630.

Anonymous:2016:Bg

[3843] Anonymous. In brief. Network Security, 2016(7):3, July 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300642.

Anonymous:2016:BRMb

[3844] Anonymous. Book review: Malware Diffusion Models for Modern Complex Networks, by Vasileios Karyotis and M. H. R. Khouzani. Morgan Kaufman. ISBN 978-0-12-802714-1. Network Security, 2016(7):4, July 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300654.

Anonymous:2016:BRE

[3845] Anonymous. Book review: Essential Skills for Hackers, by Kevin Cardwell, Henry Dalziel. Syngress. ISBN 978-0-12-804755-2. Network Security, 2016(7):4, July 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300666.

Murphy:2016:NPH

[3846] Malcolm Murphy. No place to hide as DNS comes under attack. Network Security, 2016(7):5–7, July 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485816300678.

Hack:2016:IAB

[3847] Michael Hack. The implications of Apple’s battle with the FBI. Network Security, 2016(7):8–10, July 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581630068X.

Bush:2016:HDB

[3848] Don Bush. How data breaches lead to fraud. Network Security, 2016(7):11–13, July 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300691.

Mansfield-Devine:2016:SSM

[3849] Steve Mansfield-Devine. Securing small and medium-size businesses. Network Security, 2016 (7):14–20, July 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300708.

Anonymous:2016:E

[3850] Anonymous. Events. Network Security, 2016(7):20, July 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581630071X.

Anonymous:2016:RMG

[3851] Anonymous. Ransomware menace grows as new threats emerge. Network Security, 2016(8):1–2, August 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485816300721.

Anonymous:2016:Bh

[3852] Anonymous. In brief. Network Security, 2016(8):3, August 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300733.

Anonymous:2016:BRBa

[3853] Anonymous. Book review: Bitcoin and Cryptocurrency Technologies, Arvind Narayanan, Joseph Bonneau, Edward Felten, Andrew Miller and Steven Goldfeder. Network Security, 2016(8):4, August 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300745.

Eddolls:2016:MCP

[3854] Matt Eddolls. Making cybercrime prevention the highest priority. Network Security, 2016 (8):5–8, August 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300757.

German:2016:FFY

[3855] Paul German. Face the facts — your organisation will be breached. Network Security, 2016(8):9–10, August 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485816300769.

Rafferty:2016:DSG

[3856] Ben Rafferty. Dangerous skills gap leaves organisations vulnerable. Network Security, 2016(8): 11–13, August 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300770.

Everett:2016:SES

[3857] Cath Everett. Should encryption software be banned? Network Security, 2016(8):14–17, August 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300782.

Hughes:2016:SRN

[3858] Duncan Hughes. Silent risk: new incarnations of longstanding threats. Network Security, 2016(8):17–20, August 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485816300794.

Anonymous:2016:ECg

[3859] Anonymous. Events calendar. Network Security, 2016(8):20, August 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485816300800.

Anonymous:2016:UUG

[3860] Anonymous. US and UK government agencies criticised over security. Network Security, 2016(9): 1–2, September 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300812.

Anonymous:2016:TTF

[3861] Anonymous. Two-thirds of firms may break data laws. Network Security, 2016(9):2, September 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300824.

Anonymous:2016:Bi

[3862] Anonymous. In brief. Network Security, 2016(9):3, September 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485816300836.

Anonymous:2016:BRBb

[3863] Anonymous. Book review: The Book of R, by Tilman Davies. No Starch Press. ISBN 978-1-59327-651-5. Network Security, 2016 (9):4, September 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300848.

Anonymous:2016:BRCc

[3864] Anonymous. Book review: Cyber Guerilla, by Jelle van Haaster, Rickey Gevers and Martijn Sprengers. Syngress. ISBN 978-0-12-805197-9. Network Security, 2016(9):4, September 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581630085X.

Brewer:2016:RAD

[3865] Ross Brewer. Ransomware attacks: detection, prevention and cure. Network Security, 2016(9): 5–9, September 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300861.

Sollars:2016:RBS

[3866] Marc Sollars. Risk-based security: staff can play the defining role in securing assets. Network Security, 2016 (9):9–12, September 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300873.

Crawley:2016:HH

[3867] Adrian Crawley. Hiring hackers. Network Security, 2016(9): 13–15, September 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300885.

Burdette:2016:TA

[3868] Phil Burdette. Timeline of an attack. Network Security, 2016(9): 16–17, September 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300897.

Zhang:2016:HDS

[3869] Hongwen Zhang. How to disinfect and secure the Internet of Things. Network Security, 2016(9): 18–20, September 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300903.

Anonymous:2016:ECh

[3870] Anonymous. Events calendar. Network Security, 2016(9):20, September 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485816300915.

Anonymous:2016:UOA

[3871] Anonymous. US officially accuses Russia of DNC hack while election systems come under attack. Network Security, 2016(10):1–2, October 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300927.

Anonymous:2016:Bj

[3872] Anonymous. In brief. Network Security, 2016(10):3, October 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485816300939.

Anonymous:2016:BRDa

[3873] Anonymous. Book review: Data Breach Preparation and Response, by Kevvie Fowler. Syngress. ISBN 978-0-12-803451-4. Network Security, 2016(10):4, October 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300940.

Buratowski:2016:DSB

[3874] Michael Buratowski. The DNC server breach: who did it and what does it mean? Network Security, 2016 (10):5–7, October 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816300952.

Mansfield-Devine:2016:RTB

[3875] Steve Mansfield-Devine. Ransomware:

taking businesses hostage. Network Security, 2016(10):8–17, October 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485816300964.

Anonymous:2016:RTR

[3876] Anonymous. Ransomware: threat and response. Network Security, 2016 (10):17–19, October 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300976.

Tankard:2016:SBN

[3877] Colin Tankard. Smart buildings need joined-up security. Network Security, 2016(10):20, October 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816300988.

Anonymous:2016:ECi

[3878] Anonymous. Events calendar. Network Security, 2016(10):20, October 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348581630099X.

Anonymous:2016:UGL

[3879] Anonymous. UK Government launches new £1.9bn cyber-security strategy. Network Security, 2016(11): 1–2, November 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816301003.

Anonymous:2016:Bk

[3880] Anonymous. In brief. Network Security, 2016(11):3, November 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485816301015.

Anonymous:2016:BRDb

[3881] Anonymous. Book review: DNS Security, by Allan Liska and Geoffrey Stowe. Syngress. ISBN 978-0-12-803306-7. Network Security, 2016 (11):4, November 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816301027.

Fenton:2016:REC

[3882] Mike Fenton. Restoring executive confidence: Red Team operations. Network Security, 2016 (11):5–7, November 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816301039.

Mansfield-Devine:2016:DGM

[3883] Steve Mansfield-Devine. DDoS goes mainstream: how headlinegrabbing attacks could make this threat an organisation’s biggest nightmare. Network Security, 2016(11): 7–13, November 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816301040.

Snyder:2016:PHT

[3884] Patrick Snyder and Alien Vault. Playing hackers at their own game. Network Security, 2016(11):14–16, November 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485816301052.

Higgins:2016:AAL

[3885] David Higgins. Anatomy of an attack: ‘lights out’ in Ukraine. Network Security, 2016(11):17–19, November 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485816301064.

Anonymous:2016:ECj

[3886] Anonymous. Events calendar. Network Security, 2016(11):20, November 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485816301076.

Renaud:2016:NUI

[3887] Karen Renaud. No, users are not irrational. Network Security, 2016 (11):20, November 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816301088.

Anonymous:2016:MIT

[3888] Anonymous. Major ISPs targeted in Internet of Things botnet attacks. Network Security, 2016(12): 1–2, December 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581630109X.

Anonymous:2016:RCM

[3889] Anonymous. Ransomware claims more victims. Network Security, 2016 (12):2, December 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816301106.

Anonymous:2016:Bl

[3890] Anonymous. In brief. Network Security, 2016(12):3, December 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485816301118.

Anonymous:2016:BRU

[3891] Anonymous. Book review: Understanding ECMAScript 6, by Nicholas C Zakas. No Starch Press. ISBN 978-1-59327-757-4. Network Security, 2016 (12):4, December 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581630112X.

Anonymous:2016:BRW

[3892] Anonymous. Book review: Wicked Cool Shell Scripts, by Dave Taylor and Brandon Perry. Second edition. No Starch Press. ISBN 978-1-59327-602-7. Network Security, 2016 (12):4, December 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816301131.

Pittenger:2016:ASC

[3893] Mike Pittenger. Addressing the security challenges of using containers. Network Security, 2016(12): 5–8, December 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485816301143.

Dixit:2016:OVR

[3894] Sameer Dixit. Opportunity vs risk with the Internet of Things. Network Security, 2016(12):8–10, December 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485816301155.

Mansfield-Devine:2016:DCK

[3895] Steve Mansfield-Devine. Data classification: keeping track of your most precious asset. Network Security, 2016 (12):10–15, December 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816301167.

Wagner:2016:HTA

[3896] Mathias Wagner. The hard truth about hardware in cyber-security: it’s more important. Network Security, 2016 (12):16–19, December 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485816301179.

Tankard:2016:DW

[3897] Colin Tankard. The danger within. Network Security, 2016(12):20, December 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1353485816301180.

Anonymous:2016:ECk

[3898] Anonymous. Events calendar. Network Security, 2016(12):20, December 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485816301192.

Anonymous:2017:RBMa

[3899] Anonymous. Ransomware becomes most popular form of attack as payouts approach $1bn a year. Network Security, 2017(1):1–2, January 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300016.

Anonymous:2017:Ba

[3900] Anonymous. In brief. Network Security, 2017(1):3, January 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300028.

Anonymous:2017:BRUa

[3901] Anonymous. Book review: Understanding Social Engineering-Based Scams, edited by Markus Jakobsson. Springer. ISBN 978-1-4939-6455-0. Network Security, 2017 (1):4, January 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581730003X.

Graves:2017:DFM

[3902] Jamie Graves. Data flow management: why and how. Network Security, 2017(1):5–6, January 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300041.

Mansfield-Devine:2017:OSS

[3903] Steve Mansfield-Devine. Open source software: determining the real risk posed by vulnerabilities. Network Security, 2017(1):7–12, January 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485817300053.

Wood:2017:HMS

[3904] Michael Wood. How to make SD-WAN secure. Network Security, 2017 (1):12–14, January 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300065.

German:2017:YSB

[3905] Paul German. Is your session border controller providing a false sense of security? Network Security, 2017 (1):14–16, January 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300077.

Welch:2017:EWS

[3906] Bill Welch. Exploiting the weaknesses of SS7. Network Security, 2017 (1):17–19, January 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300089.

Renaud:2017:WCP

[3907] Karen Renaud. Webcam covering phenomenon. Network Security, 2017(1):20, January 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300090.

Anonymous:2017:ECa

[3908] Anonymous. Events calendar. Network Security, 2017(1):20, January 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300107.

Anonymous:2017:RBMb

[3909] Anonymous. Ransomware becomes the most prevalent form of malware and hits an ever-wider range of victims. Network Security, 2017 (2):1–2, February 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300119.

Anonymous:2017:Bb

[3910] Anonymous. In brief. Network Security, 2017(2):3, February 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300120.

Anonymous:2017:BRUb

[3911] Anonymous. Book review: Understanding the Digital World, by Brian Kernighan. Princeton University Press. ISBN 978-0-691-17654-3. Network Security, 2017(2):4, February 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300132.

Lilly:2017:ICH

[3912] Andy Lilly. IMSI catchers: hacking mobile communications. Network Security, 2017(2):5–7, February 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485817300144.

Bedell-Pearce:2017:WBD

[3913] Jack Bedell-Pearce. When big data and Brexit collide. Network Security, 2017(2):8–9, February 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300156.

Graves:2017:WIP

[3914] Jamie Graves. What is intellectual property and how do you protect it? Network Security, 2017 (2):9–11, February 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300168.

Schenk:2017:RMF

[3915] Gert-Jan Schenk. Retailers must focus on the growing cyberthreat landscape. Network Security, 2017(2): 12–13, February 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581730017X.

Mansfield-Devine:2017:PDS

[3916] Steve Mansfield-Devine. A process of defence — securing industrial control systems. Network Security, 2017 (2):14–19, February 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300181.

Erridge:2017:ORM

[3917] Tim Erridge. Optimising risk management. Network Security, 2017 (2):20, February 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300193.

Anonymous:2017:ECb

[3918] Anonymous. Events calendar. Network Security, 2017(2):20, February 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581730020X.

Anonymous:2017:FIB

[3919] Anonymous. Financial institutions become more confident about cyber-security but weaknesses remain. Network Security, 2017 (3):1–2, March 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300211.

Anonymous:2017:ZDL

[3920] Anonymous. Zero-days last for years.

Network Security, 2017(3):2, March 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300223.

Anonymous:2017:Bc

[3921] Anonymous. In brief. Network Security, 2017(3):3, March 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300235.

Anonymous:2017:BRM

[3922] Anonymous. Book review: The Mathematics of Secrets, by Joshua Holden. Princeton University Press. ISBN 978-0-691-14175-6. Network Security, 2017(3):4, March 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300247.

Tankard:2017:ECB

[3923] Colin Tankard. Encryption as the cornerstone of big data security. Network Security, 2017 (3):5–7, March 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300259.

Henderson:2017:NIN

[3924] Lewis Henderson. National infrastructure — the next step for seasoned hackers. Network Security, 2017(3):8–10, March 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300260.

Smyth:2017:SVM

[3925] Vincent Smyth. Software vulnerability management: how intelligence helps reduce the risk. Network Security, 2017(3):10–12, March 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300272.

Hawkins:2017:WCV

[3926] Nick Hawkins. Why communication is vital during a cyberattack. Network Security, 2017 (3):12–14, March 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300284.

Ring:2017:RCS

[3927] Tim Ring. The Russians are coming! Are security firms over-hyping the hacker threat? Network Security, 2017(3):15–18, March 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300296.

Green:2017:RG

[3928] Andy Green. Ransomware and the GDPR. Network Security, 2017 (3):18–19, March 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300302.

Tankard:2017:FEC

[3929] Colin Tankard. The firewall: The eSignature comes of age. Network Security, 2017(3):20, March 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300314.

Anonymous:2017:ECc

[3930] Anonymous. Events calendar. Network Security, 2017(3):20, March 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485817300326.

Anonymous:2017:RCN

[3931] Anonymous. The Russian connection: new attacks and arrests. Network Security, 2017(4):1–2, April 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300338.

Anonymous:2017:Bd

[3932] Anonymous. In brief. Network Security, 2017(4):3, April 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581730034X.

SM-D:2017:BRH

[3933] SM-D. Book review: Hiding Behind the Keyboard, by Brett Shavers and John Bair, Syngress. ISBN 978-0-12-803340-1. Network Security, 2017(4):4, April 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300351.

Adato:2017:MAE

[3934] Leon Adato. Monitoring and automation: it’s easier than you think. Network Security, 2017 (4):5–7, April 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300363.

Mansfield-Devine:2017:FAC

[3935] Steve Mansfield-Devine. Fileless attacks: compromising targets without malware. Network Security, 2017(4):7–11, April 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300375.

Zeichick:2017:EIO

[3936] Alan Zeichick. Enabling innovation by opening up the network. Network Security, 2017 (4):12–14, April 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300387.

Talwar:2017:AIN

[3937] Rohit Talwar and April Koury. Artificial intelligence — the next frontier in IT security? Network Security, 2017(4):14–17, April 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300399.

Hickman:2017:TI

[3938] Mark Hickman. The threat from inside. Network Security, 2017 (4):18–19, April 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300405.

Renaud:2017:FDC

[3939] Karen Renaud. Facts do not change minds. Network Security, 2017(4):20, April 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300417.

Anonymous:2017:ECd

[3940] Anonymous. Events calendar. Network Security, 2017(4):20, April 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485817300429.

Anonymous:2017:SES

[3941] Anonymous. Social engineering scams ensnare Google, Facebook and their users. Network Security, 2017(5):1–2, May 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300430.

Anonymous:2017:FSM

[3942] Anonymous. Firms struggle to manage networks. Network Security, 2017(5):2, May 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300442.

Anonymous:2017:Be

[3943] Anonymous. In brief. Network Security, 2017(5):3, May 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300454.

SM-D:2017:BRP

[3944] SM-D. Book review: Practical Packet Analysis, Chris Sanders. Network Security, 2017(5):4, May 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300466.

SM-D:2017:BRW

[3945] SM-D. Book review: We Know All About You, Rhodri Jeffreys-Jones. Network Security, 2017(5):4, May 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300478.

Newman:2017:SPG

[3946] Sean Newman. Service providers:

the gatekeepers of Internet security. Network Security, 2017(5):5–7, May 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581730048X.

Gajek:2017:MMD

[3947] Jacob Gajek. Macro malware:

dissecting a malicious Word document. Network Security, 2017 (5):8–13, May 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300491.

Mansfield-Devine:2017:THA

[3948] Steve Mansfield-Devine. Threat hunting: assuming the worst to strengthen resilience. Network Security, 2017 (5):13–17, May 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300508.

Farral:2017:API

[3949] Travis Farral. The attribution problem with information security attacks. Network Security, 2017 (5):17–19, May 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581730051X.

Erridge:2017:GS

[3950] Tim Erridge. Gaining strength. Network Security, 2017(5):20, May 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485817300521.

Anonymous:2017:ECe

[3951] Anonymous. Events calendar. Network Security, 2017(5):20, May 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485817300533.

Anonymous:2017:NLS

[3952] Anonymous. NSA leak shows Russian attack on US electoral system. Network Security, 2017 (6):1–2, June 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300545.

Anonymous:2017:AI

[3953] Anonymous. Attacks on industry.

Network Security, 2017(6):2, June 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300557.

Anonymous:2017:Bf

[3954] Anonymous. In brief. Network Security, 2017(6):3, June 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300569.

Anonymous:2017:BRPa

[3955] Anonymous. Book review: Practical Forensic Imaging, by Bruce Nikkel. No Starch Press. ISBN 978-1-59327-793-2. Network Security, 2017(6):4, June 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300570.

Anonymous:2017:BRPb

[3956] Anonymous. Book review: The Plot to Hack America, by Malcolm Nance. Skyhorse Publishing. ISBN 978-1-5107-2332-0. Network Security, 2017(6):4, June 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300582.

Shugrue:2017:FAT

[3957] Daniel Shugrue. Fighting application threats with cloud-based WAFs. Network Security, 2017 (6):5–8, June 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300594.

Zerlang:2017:GMC

[3958] Jesper Zerlang. GDPR: a milestone in convergence for cyber-security and compliance. Network Security, 2017(6):8–11, June 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300600.

Botelho:2017:HAD

[3959] Jay Botelho. How automating data collection can improve cybersecurity. Network Security, 2017 (6):11–13, June 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300612.

Mansfield-Devine:2017:LRK

[3960] Steve Mansfield-Devine. Leaks and ransoms — the key threats to healthcare organisations. Network Security, 2017(6):14–19, June 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300624.

Tankard:2017:SE

[3961] Colin Tankard. Securing emails. Network Security, 2017(6):20, June 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300636.

Anonymous:2017:ECf

[3962] Anonymous. Events calendar. Network Security, 2017(6):20, June 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300648.

Anonymous:2017:NAO

[3963] Anonymous. National Accounting Office demands action. Network Security, 2017(7):1–2, July 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581730065X.

Anonymous:2017:PA

[3964] Anonymous. Parliament attacked.

Network Security, 2017(7):2, July 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300661.

Anonymous:2017:Bg

[3965] Anonymous. In brief. Network Security, 2017(7):3, July 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300673.

Anonymous:2017:BRG

[3966] Anonymous. Book review: Gray Hat C#, by Brandon Perry. No Starch Press. ISBN 978-1-59327-759-

8. Network Security, 2017(7):4, July 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300685.

Parkinson:2017:UAC

[3967] Simon Parkinson. Use of access control to minimise ransomware impact. Network Security, 2017(7):5–8, July 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300697.

Wood:2017:TRS

[3968] Michael Wood. Top requirements on the SD-WAN security checklist. Network Security, 2017(7):9–11, July 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300703.

Taylor:2017:ACM

[3969] Mark Taylor, Denis Reilly, and Brett Lempereur. An access control management protocol for Internet of Things devices. Network Security, 2017(7):11–17, July 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300715.

Murphy:2017:ITT

[3970] Malcolm Murphy. The Internet of Things and the threat it poses to DNS. Network Security, 2017 (7):17–19, July 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300727.

Renaud:2017:IMY

[3971] Karen Renaud. It makes you Wanna Cry. Network Security, 2017(7):20, July 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300739.

Anonymous:2017:ECg

[3972] Anonymous. Events calendar. Network Security, 2017(7):20, July 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300740.

Anonymous:2017:TID

[3973] Anonymous. Two illicit dark web marketplaces taken down. Network Security, 2017(8):1–2, August 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300752.

Anonymous:2017:IBH

[3974] Anonymous. Italian bank hit with major breach. Network Security, 2017(8):2, August 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300764.

Anonymous:2017:Bh

[3975] Anonymous. In brief. Network Security, 2017(8):3, August 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300776.

Anonymous:2017:BRT

[3976] Anonymous. Book review: Threat Forecasting, by John Pirc, David DeSanto, Iain Davison, and Will Gragido. Syngress. ISBN 978-0-12-800006-9. Network Security, 2017 (8):4, August 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300788.

Anonymous:2017:BRI

[3977] Anonymous. Book review: Intercept: The Secret History of Computers and Spies [by Gordon Corera, Weidenfeld and Nicolson, London, UK 2015, ISBN 1-78022-784-1]. Network Security, 2017(8):4, August 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581730079X.

Luszcz:2017:HMD

[3978] Jeff Luszcz. How maverick developers can create risk in the software and IoT supply chain. Network Security, 2017(8):5–7, August 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300806.

Haggerty:2017:HDT

[3979] Eileen Haggerty. Healthcare and digital transformation. Network Security, 2017(8):7–11, August 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300818.

Major:2017:WTI

[3980] Tom Major. Weaponising threat intelligence data. Network Security, 2017(8):11–13, August 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581730082X.

German:2017:TBD

[3981] Paul German. Time to bury dedicated hardware-based security solutions. Network Security, 2017 (8):13–15, August 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300831.

Mansfield-Devine:2017:ADP

[3982] Steve Mansfield-Devine. Adapting to the disappearing perimeter. Network Security, 2017(8): 15–19, August 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300843.

Erridge:2017:GWH

[3983] Tim Erridge. Good will hunting.

Network Security, 2017(8):20, August 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300855.

Anonymous:2017:ECh

[3984] Anonymous. Events calendar. Network Security, 2017(8):20, August 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300867.

Anonymous:2017:SBH

[3985] Anonymous. State-backed hackers target US and European energy sectors. Network Security, 2017 (9):1–2, September 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817301071.

Anonymous:2017:RIA

[3986] Anonymous. Ransomware and IoT among leading threats. Network Security, 2017(9):2, September 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817301083.

Anonymous:2017:Bi

[3987] Anonymous. In brief. Network Security, 2017(9):3, September 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817301095.

Anonymous:2017:BRD

[3988] Anonymous. Book review: Deception in the Digital Age, by Cameron Malin, Terry Gudaitis, Thomas Holt, and Max Kilger. Academic Press. ISBN 978-0-12-411630-6. Network Security, 2017(9):4, September 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817301101.

Farral:2017:NSA

[3989] Travis Farral. Nation-state attacks:

practical defences against advanced adversaries. Network Security, 2017 (9):5–7, September 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817301113.

Condra:2017:ANS

[3990] Jon Condra. Assessing nation state threats. Network Security, 2017(9): 8–10, September 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817301125.

Vincent:2017:SSH

[3991] Adam Vincent. State-sponsored hackers: the new normal for business. Network Security, 2017(9): 10–12, September 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817301137.

Robinson:2017:DDG

[3992] Nick Robinson and Keith Martin.

Distributed denial of government: the Estonian Data Embassy Initiative. Network Security, 2017(9): 13–16, September 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817301149.

Beckett:2017:DIN

[3993] Phil Beckett. Data and IP are the new nuclear: facing up to state-sponsored threats. Network Security, 2017(9): 17–19, September 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817301150.

Tankard:2017:WA

[3994] Colin Tankard. Who are the attackers? Network Security, 2017 (9):20, September 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817301162.

Anonymous:2017:ECi

[3995] Anonymous. Events calendar. Network Security, 2017(9):20, September 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485817301174.

Anonymous:2017:NRM

[3996] Anonymous. NCSC reports more than 1,000 cyber-attacks in first year. Network Security, 2017 (10):1–2, October 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300983.

Anonymous:2017:Bj

[3997] Anonymous. In brief. Network Security, 2017(10):3, October 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485817300995.

Anonymous:2017:BRCa

[3998] Anonymous. Book review: Cybercrime and Business, by Sanford Moskowitz. Butterworth-Heinemann. ISBN 978-0-12-800353-4. Network Security, 2017(10):4, October 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817301009.

Lundbohm:2017:UNS

[3999] Eric Lundbohm. Understanding nation-state attacks. Network Security, 2017(10):5–8, October 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817301010.

Bambenek:2017:NSA

[4000] John Bambenek. Nation-state attacks:

the new normal. Network Security, 2017(10):8–10, October 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817301022.

Laing:2017:CGW

[4001] Brian Laing. Cyber global warming:

six steps towards meltdown. Network Security, 2017(10):11–13, October 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817301034.

Mansfield-Devine:2017:WIT

[4002] Steve Mansfield-Devine. Weaponising the Internet of Things. Network Security, 2017(10):13–19, October 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485817301046.

Renaud:2017:PPL

[4003] Karen Renaud and Gareth Renaud. To phish, or not to phish …. Network Security, 2017 (10):20, October 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817301058.

Anonymous:2017:ECj

[4004] Anonymous. Events calendar. Network Security, 2017(10):20, October 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348581730106X.

Anonymous:2017:BRR

[4005] Anonymous. Bad Rabbit ransomware attacks Russia and Ukraine. Network Security, 2017(11):1–2, November 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485817300879.

Anonymous:2017:WFA

[4006] Anonymous. Wifi flaw affects nearly all devices. Network Security, 2017 (11):2, November 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817300880.

Anonymous:2017:Bk

[4007] Anonymous. In brief. Network Security, 2017(11):3, November 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300892.

Anonymous:2017:BRCb

[4008] Anonymous. Book review: Computer and Information Security Handbook, Third edition, edited by John R Vacca. Morgan Kaufmann. ISBN 978-0-12-803843-7. Network Security, 2017(11):4, November 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300909.

Anonymous:2017:BRB

[4009] Anonymous. Book review: Big Data:

a very short introduction, by Dawn E. Holmes. Oxford University Press. ISBN 978-0-19-877957-5. Network Security, 2017(11):4, November 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300910.

Steffen:2017:SJB

[4010] Chris Steffen. Should jump box servers be consigned to history? Network Security, 2017(11):5–6, November 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300922.

Thompson:2017:PSH

[4011] Nick Thompson. Putting security at the heart of app development. Network Security, 2017 (11):7–8, November 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300934.

Bhardwaj:2017:SCC

[4012] Akashdeep Bhardwaj and Sam Goundar. Security challenges for cloud-based email infrastructure. Network Security, 2017(11):8–15, November 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300946.

Mansfield-Devine:2017:GCA

[4013] Steve Mansfield-Devine. Going critical: attacks against national infrastructure. Network Security, 2017 (11):16–19, November 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817300958.

Tankard:2017:BNK

[4014] Colin Tankard. BYOE: New kid on the block. Network Security, 2017 (11):20, November 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581730096X.

Anonymous:2017:ECk

[4015] Anonymous. Events calendar. Network Security, 2017(11):20, November 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817300971.

Anonymous:2017:NCT

[4016] Anonymous. NHS to create a £20m SOC and pen-testing operation. Network Security, 2017(12): 1–2, December 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817301186.

Anonymous:2017:UGA

[4017] Anonymous. UK Government announces interim strategy. Network Security, 2017(12):2, December 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817301198.

Anonymous:2017:B

[4018] Anonymous. In brief. Network Security, 2017(12):3, December 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485817301204.

Anonymous:2017:SWA

[4019] Anonymous. Safety of Web applications. Network Security, 2017 (12):4, December 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485817301216.

Furnell:2017:SEA

[4020] Steven Furnell and Ismini Vasileiou. Security education and awareness: just let them burn? Network Security, 2017(12):5–9, December 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817301228.

Lauria:2017:HFR

[4021] Filippo Lauria. How to footprint, report and remotely secure compromised IoT devices. Network Security, 2017 (12):10–16, December 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581730123X.

Mansfield-Devine:2017:RRH

[4022] Steve Mansfield-Devine. The right response: how organisations should react to security incidents. Network Security, 2017(12):16–19, December 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485817301241.

Erridge:2017:FT

[4023] Tim Erridge. A framework for threats. Network Security, 2017 (12):20, December 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485817301253.

Anonymous:2017:EC

[4024] Anonymous. Events calendar. Network Security, 2017(12):20, December 2017. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485817301265.

Anonymous:2018:NKB

[4025] Anonymous. North Korea blamed for WannaCry, PoS attacks and Bitcoin phishing. Network Security, 2018(1):1–2, January 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300011.

Anonymous:2018:Ba

[4026] Anonymous. In brief. Network Security, 2018(1):3, January 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300023.

Anonymous:2018:SITa

[4027] Anonymous. Securing the Internet of Things. Network Security, 2018(1):4, January 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300035.

Anonymous:2018:NR

[4028] Anonymous. Network routing. Network Security, 2018(1):4, January 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818300047.

Luszcz:2018:ASH

[4029] Jeff Luszcz. Apache Struts 2: how technical and development gaps caused the Equifax Breach. Network Security, 2018(1):5–8, January 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300059.

Boireau:2018:SBA

[4030] Olivier Boireau. Securing the blockchain against hackers. Network Security, 2018(1):8–11, January 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818300060.

Nicholson:2018:BBB

[4031] Dave Nicholson. Blurring the boundaries between networking and IT security. Network Security, 2018 (1):11–13, January 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300072.

Farha:2018:MRA

[4032] Fadi Farha and Hongsong Chen. Mitigating replay attacks with ZigBee solutions. Network Security, 2018 (1):13–19, January 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300084.

Tankard:2018:TPP

[4033] Colin Tankard. Tackling push payment scams. Network Security, 2018 (1):20, January 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300096.

Anonymous:2018:ECa

[4034] Anonymous. Events calendar. Network Security, 2018(1):20, January 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818300102.

Anonymous:2018:CCH

[4035] Anonymous. Crypto-currencies hit by hacking attacks, theft and fraud. Network Security, 2018(2):1–2, February 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818300114.

Anonymous:2018:Bb

[4036] Anonymous. In brief. Network Security, 2018(2):3, February 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818300126.

Anonymous:2018:STM

[4037] Anonymous. Seeking the truth from mobile evidence. Network Security, 2018(2):4, February 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300138.

Wickes:2018:COD

[4038] James Wickes. CCTV: an open door into enterprise and national infrastructure. Network Security, 2018 (2):5–8, February 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581830014X.

Al-Azzam:2018:SRC

[4039] Saad Al-Azzam, Ahmad Sharieh, Azzam Sleit, and Nedaa Al-Azzam. Securing robot communication using packet encryption distribution. Network Security, 2018(2):8–14, February 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818300151.

Mansfield-Devine:2018:OSI

[4040] Steve Mansfield-Devine. Open source and the Internet of Things. Network Security, 2018(2):14–19, February 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818300163.

Renaud:2018:GTC

[4041] Karen Renaud and Lynsay Shepherd. GDPR: its time has come. Network Security, 2018(2):20, February 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818300175.

Anonymous:2018:ECb

[4042] Anonymous. Events calendar. Network Security, 2018(2):20, February 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818300187.

Anonymous:2018:CBH

[4043] Anonymous. Cyber-breaches hit twice as hard in past year, says Cisco. Network Security, 2018 (3):1–2, March 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300199.

Anonymous:2018:UGL

[4044] Anonymous. UK Government launches IoT code. Network Security, 2018(3): 2, March 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485818300205.

Anonymous:2018:Bc

[4045] Anonymous. In brief. Network Security, 2018(3):3, March 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300217.

Anonymous:2018:SWA

[4046] Anonymous. Safety of Web applications. Network Security, 2018(3): 4, March 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1353485818300229.

Scott:2018:HZT

[4047] Barry Scott. How a zero trust approach can help to secure your AWS environment. Network Security, 2018(3):5–8, March 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300230.

Shipulin:2018:WNT

[4048] Kirill Shipulin. We need to talk about IDS signatures. Network Security, 2018(3):8–13, March 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818300242.

Ayrour:2018:MCA

[4049] Yassine Ayrour, Amine Raji, and Mahmoud Nassar. Modelling cyber-attacks: a survey study. Network Security, 2018(3):13–19, March 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818300254.

Tankard:2018:TFT

[4050] Colin Tankard. The threat of fileless trojans. Network Security, 2018(3):20, March 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818300266.

Anonymous:2018:ECc

[4051] Anonymous. Events calendar. Network Security, 2018(3):20, March 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818300278.

Anonymous:2018:UOS

[4052] Anonymous. UK organisations seriously unprepared for cyber-attacks. Network Security, 2018(4):1–2, April 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581830028X.

Anonymous:2018:Bd

[4053] Anonymous. In brief. Network Security, 2018(4):3, April 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300291.

Anonymous:2018:PC

[4054] Anonymous. Port cybersecurity. Network Security, 2018(4):4, April 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818300308.

Anonymous:2018:BBa

[4055] Anonymous. Bit by bit. Network Security, 2018(4):4, April 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581830031X.

Mackey:2018:BOS

[4056] Tim Mackey. Building open source security into agile application builds. Network Security, 2018(4):5–8, April 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818300321.

Bajpai:2018:AMI

[4057] Pranshu Bajpai, Aditya K. Sood, and Richard J. Enbody. The art of mapping IoT devices in networks. Network Security, 2018(4):8–15, April 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300333.

Bhardwaj:2018:RTS

[4058] Akashdeep Bhardwaj and Sam Goundar. Reducing the threat surface to minimise the impact of cyber-attacks. Network Security, 2018(4):15–19, April 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818300345.

Erridge:2018:TCR

[4059] Tim Erridge. True colours of red teaming. Network Security, 2018(4):20, April 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300357.

Anonymous:2018:ECd

[4060] Anonymous. Events calendar. Network Security, 2018(4):20, April 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818300369.

Anonymous:2018:OYA

[4061] Anonymous. One year after WannaCry: assessing the aftermath. Network Security, 2018(5):1–2, May 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818300370.

Anonymous:2018:EVS

[4062] Anonymous. Equifax vulnerability still a problem. Network Security, 2018(5):2, May 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300382.

Anonymous:2018:CSB

[4063] Anonymous. Cyber-security Breaches Survey: fraudulent emails present a major menace. Network Security, 2018(5):3, May 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300394.

Anonymous:2018:Be

[4064] Anonymous. In brief. Network Security, 2018(5):4, May 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300400.

Anonymous:2018:BRS

[4065] Anonymous. Book review: Simplifying JavaScript, by Joe Morgan. Published by Pragmatic Bookshelf. ISBN: 978-1-68050-288-6. Price: $39.95, 282 pgs, paperback. Network Security, 2018(5):5, May 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300412.

James:2018:MCS

[4066] Lee James. Making cyber-security a strategic business priority. Network Security, 2018(5):6–8, May 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818300424.

Aminzade:2018:CIA

[4067] Michael Aminzade. Confidentiality, integrity and availability — finding a balanced IT framework. Network Security, 2018(5):9–11, May 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818300436.

Sim:2018:DAM

[4068] Greg Sim. Defending against the malware flood. Network Security, 2018(5):12–13, May 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818300448.

Mansfield-Devine:2018:IWP

[4069] Steve Mansfield-Devine. The intelligent way to protect complex environments. Network Security, 2018 (5):13–17, May 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581830045X.

Wilkins:2018:AIH

[4070] Jonathan Wilkins. Is artificial intelligence a help or hindrance? Network Security, 2018(5):18–19, May 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818300461.

Renaud:2018:CSA

[4071] Karen Renaud. Cooking up security awareness & training. Network Security, 2018(5):20, May 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300473.

Anonymous:2018:ECe

[4072] Anonymous. Events calendar. Network Security, 2018(5):20, May 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818300485.

Anonymous:2018:SRC

[4073] Anonymous. Sharp rise in costs and damage from DNS-related attacks. Network Security, 2018 (6):1–2, June 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300497.

Anonymous:2018:HUA

[4074] Anonymous. Healthcare under attack. Network Security, 2018(6):2, June 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818300503.

Anonymous:2018:SOS

[4075] Anonymous. Synopsis: Open source security and risk analysis. Network Security, 2018(6):3, June 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300515.

Anonymous:2018:Bf

[4076] Anonymous. In brief. Network Security, 2018(6):4, June 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818300527.

Anonymous:2018:RMC

[4077] Anonymous. Research methods for cyber security. Network Security, 2018(6):5, June 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818300539.

Townsley:2018:BGC

[4078] Chris Townsley. Are businesses getting complacent when it comes to DDoS mitigation? Network Security, 2018(6):6–9, June 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818300540.

Simmonds:2018:ICD

[4079] Mike Simmonds. Instilling a culture of data security throughout the organisation. Network Security, 2018(6):9–12, June 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818300552.

Macy:2018:HBS

[4080] Jason Macy. How to build a secure API gateway. Network Security, 2018(6):12–14, June 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300564.

Khan:2018:NCC

[4081] Javid Khan. The need for continuous compliance. Network Security, 2018(6):14–15, June 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300576.

Mansfield-Devine:2018:FFH

[4082] Steve Mansfield-Devine. Friendly fire: how penetration testing can reduce your risk. Network Security, 2018(6):16–19, June 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300588.

Tankard:2018:SDD

[4083] Colin Tankard. Streamlining data discovery. Network Security, 2018(6): 20, June 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S135348581830059X.

Anonymous:2018:ECf

[4084] Anonymous. Events calendar. Network Security, 2018(6):20, June 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818300606.

Anonymous:2018:NCSa

[4085] Anonymous. New cyber-security guidelines for government departments.

Network Security, 2018(7):1–2, July 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818300618.

Anonymous:2018:HTS

[4086] Anonymous. HMRC takes down 20,000 sites. Network Security, 2018(7):2, 5, July 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581830062X.

Anonymous:2018:HNK

[4087] Anonymous. Harvey Nash & KPMG: CIO Survey 2018 & growing pains. Network Security, 2018(7):3, July 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818300631.

Anonymous:2018:Bg

[4088] Anonymous. In brief. Network Security, 2018(7):4, July 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818300643.

Anonymous:2018:BRM

[4089] Anonymous. Book review: Modern Vim. Network Security, 2018(7):5, July 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818300655.

Anonymous:2018:FCM

[4090] Anonymous. Firms collect more data than they can handle. Network Security, 2018(7):5, July 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300667.

Beecher:2018:EGN

[4091] Phil Beecher. Enterprise-grade networks: the answer to IoT security challenges. Network Security, 2018(7):6–9, July 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300679.

Dennis:2018:WPM

[4092] Colin Dennis. Why is patch management necessary? Network Security, 2018(7):9–13, July 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300680.

Chadd:2018:DAP

[4093] Anthony Chadd. DDoS attacks: past, present and future. Network Security, 2018(7):13–15, July 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300692.

Mansfield-Devine:2018:DFR

[4094] Steve Mansfield-Devine. DevOps: finding room for security. Network Security, 2018(7):15–20, July 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300709.

Anonymous:2018:ECg

[4095] Anonymous. Events calendar. Network Security, 2018(7):20, July 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818300710.

Anonymous:2018:RAB

[4096] Anonymous. Russian attackers breach US electricity network. Network Security, 2018(8):1–3, August 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818300722.

Anonymous:2018:Ta

[4097] Anonymous. Threatwatch. Network Security, 2018(8):3, August 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818300734.

Anonymous:2018:GRI

[4098] Anonymous. GDPR reduces insider threats. Network Security, 2018 (8):3, 19, August 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818300746.

Anonymous:2018:QIR

[4099] Anonymous. Quarterly Incident Response Threat Report, July 2018. Network Security, 2018(8):4, August 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818300758.

Anonymous:2018:Bh

[4100] Anonymous. In brief. Network Security, 2018(8):5, August 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581830076X.

Mason:2018:ORP

[4101] Mike Mason. Overcoming the risks of privileged user abuse in Salesforce. Network Security, 2018 (8):6–8, August 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300771.

Caldwell:2018:PIOa

[4102] Tracey Caldwell. Plugging IT/OT vulnerabilities — part 1. Network Security, 2018(8):9–14, August 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818300783.

Mansfield-Devine:2018:KWY

[4103] Steve Mansfield-Devine. Knowing what you have: the road to effective data discovery. Network Security, 2018 (8):15–19, August 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300795.

Anonymous:2018:VSS

[4104] Anonymous. Vulnerabilities stay the same. Network Security, 2018 (8):19, August 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818300801.

Fay:2018:BRC

[4105] Oliver Fay. BEC reality check. Network Security, 2018(8):20, August 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818300813.

Anonymous:2018:ECh

[4106] Anonymous. Events calendar. Network Security, 2018(8):20, August 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818300825.

Anonymous:2018:TFT

[4107] Anonymous. Tech firms tackle fake accounts and political hackers. Network Security, 2018(9): 1–3, September 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818300837.

Anonymous:2018:Tb

[4108] Anonymous. Threatwatch. Network Security, 2018(9):3, September 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818300849.

Anonymous:2018:BBb

[4109] Anonymous. BA breached. Network Security, 2018(9):3, September 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818300850.

Anonymous:2018:TMU

[4110] Anonymous. Trend Micro: Unseen threats, imminent losses: 2018 midyear security roundup. Network Security, 2018(9):4, September 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300862.

Anonymous:2018:Bi

[4111] Anonymous. In brief. Network Security, 2018(9):5, September 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818300874.

Macy:2018:ASW

[4112] Jason Macy. API security: whose job is it anyway? Network Security, 2018(9):6–9, September 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300886.

Caldwell:2018:PIOb

[4113] Tracey Caldwell. Plugging IT/OT vulnerabilities — part 2. Network Security, 2018(9):10–15, September 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818300898.

Keplinger:2018:QCB

[4114] Keegan Keplinger. Is quantum computing becoming relevant to cybersecurity? Network Security, 2018 (9):16–19, September 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818300904.

Renaud:2018:UON

[4115] Karen Renaud and Verena Zimmermann. A user by any other name. Network Security, 2018 (9):20, September 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818300916.

Anonymous:2018:ECi

[4116] Anonymous. Events calendar. Network Security, 2018(9):20, September 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818300928.

Anonymous:2018:CPS

[4117] Anonymous. China put spy chips in servers, claims Bloomberg. Network Security, 2018(10):1–3, October 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348581830093X.

Anonymous:2018:Tc

[4118] Anonymous. Threatwatch. Network Security, 2018(10):3, October 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818300941.

Anonymous:2018:FLG

[4119] Anonymous. Flaw leads to Google+ shutting down. Network Security, 2018(10):3, October 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300953.

Anonymous:2018:EIO

[4120] Anonymous. Europol: Internet Organised Crime Threat Assessment. Network Security, 2018 (10):4, October 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300965.

Anonymous:2018:Bj

[4121] Anonymous. In brief. Network Security, 2018(10):5, October 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818300977.

Sens:2018:RFN

[4122] Ronald Sens. Be ready to fight new 5G vulnerabilities. Network Security, 2018(10):6–7, October 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300989.

Hogewoning:2018:IRS

[4123] Marco Hogewoning. IoT and regulation — striking the right balance. Network Security, 2018(10): 8–10, October 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818300990.

Sollars:2018:LMW

[4124] Marc Sollars. Love and marriage: why security and SD-WAN need to go together. Network Security, 2018 (10):10–12, October 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818301004.

McKenna:2018:MCR

[4125] Benedict McKenna. Measuring cyberrisk. Network Security, 2018(10): 12–14, October 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818301016.

Mansfield-Devine:2018:HDA

[4126] Steve Mansfield-Devine. Hacking democracy: abusing the Internet for political gain. Network Security, 2018 (10):15–19, October 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818301028.

Tankard:2018:HTA

[4127] Colin Tankard. How trustworthy is AI? Network Security, 2018 (10):20, October 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581830103X.

Anonymous:2018:ECj

[4128] Anonymous. Events calendar. Network Security, 2018(10):20, October 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818301041.

Anonymous:2018:CSH

[4129] Anonymous. China steps up hacking as US issues indictments. Network Security, 2018(11):1–2, November 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818301053.

Anonymous:2018:TBR

[4130] Anonymous. Triton blamed on Russia. Network Security, 2018 (11):3, November 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818301065.

Anonymous:2018:Td

[4131] Anonymous. Threatwatch. Network Security, 2018(11):3, November 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818301077.

Anonymous:2018:NCSb

[4132] Anonymous. National Cyber Security Centre: Annual Review 2018. Network Security, 2018 (11):4, November 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818301089.

Anonymous:2018:Bk

[4133] Anonymous. In brief. Network Security, 2018(11):5, November 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485818301090.

Bedell-Pearce:2018:SDT

[4134] Jack Bedell-Pearce. Safe digital transformation for SMEs. Network Security, 2018(11):6–7, November 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818301107.

Roberts:2018:LLD

[4135] Steve Roberts. Learning lessons from data breaches. Network Security, 2018 (11):8–11, November 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818301119.

Claxson:2018:SVE

[4136] Nick Claxson. Securing VoIP: encrypting today’s digital telephony systems. Network Security, 2018(11): 11–13, November 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818301120.

Timms:2018:BCD

[4137] Paul Timms. Business continuity and disaster recovery — advice for best practice. Network Security, 2018 (11):13–14, November 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818301132.

Mansfield-Devine:2018:NSA

[4138] Steve Mansfield-Devine. Nation-state attacks: the start of a new Cold War? Network Security, 2018(11): 15–19, November 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818301144.

Green:2018:PTC

[4139] Ranulf Green. Pen-testing the cloud. Network Security, 2018 (11):20, November 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818301156.

Anonymous:2018:ECk

[4140] Anonymous. Events calendar. Network Security, 2018(11):20, November 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818301168.

Anonymous:2018:SITb

[4141] Anonymous. Securing the Internet of Things — part one. Network Security, 2018(12):1, December 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581830117X.

Anonymous:2018:FFD

[4142] Anonymous. Firms failing to disclose IoT vulnerabilities. Network Security, 2018(12):1–2, December 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818301181.

Anonymous:2018:UBR

[4143] Anonymous. Ukraine blames Russia for cyber attacks. Network Security, 2018(12):2–3, December 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818301193.

Anonymous:2018:Te

[4144] Anonymous. Threatwatch. Network Security, 2018(12):3, December 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348581830120X.

Anonymous:2018:ILG

[4145] Anonymous. IoT leads growth in endpoint security. Network Security, 2018(12):3, December 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818301211.

Anonymous:2018:NTI

[4146] Anonymous. Nokia Threat Intelligence Report — 2019. Network Security, 2018(12):4, December 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485818301223.

Anonymous:2018:Bl

[4147] Anonymous. In brief. Network Security, 2018(12):5, December 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818301235.

Baldin:2018:SWS

[4148] Andy Baldin. Securing the weak spots. Network Security, 2018(12): 6–8, December 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818301247.

Stone:2018:PCT

[4149] Tim Stone. Protecting connected transportation networks. Network Security, 2018(12):8–10, December 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818301259.

Raghu:2018:MDS

[4150] RV Raghu. The many dimensions of successful IoT deployment. Network Security, 2018(12): 10–15, December 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818301260.

Mansfield-Devine:2018:GGI

[4151] Steve Mansfield-Devine. Getting to grips with IoT. Network Security, 2018 (12):15–19, December 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818301272.

Tankard:2018:LAA

[4152] Colin Tankard. A layered approach to authentication. Network Security, 2018(12):20, December 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485818301284.

Anonymous:2018:ECl

[4153] Anonymous. Events calendar. Network Security, 2018(12):20, December 2018. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485818301296.

Anonymous:2019:SIT

[4154] Anonymous. Securing the Internet of Things — part two. Network Security, 2019(1):1, January 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300017.

Anonymous:2019:UCI

[4155] Anonymous. UK critical infrastructure is wide open to attack, warns parliamentary committee. Network Security, 2019(1):1–3, January 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300029.

Anonymous:2019:Ta

[4156] Anonymous. Threatwatch. Network Security, 2019(1):3, January 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300030.

Anonymous:2019:ARH

[4157] Anonymous. Avast report highlights IoT vulnerabilities. Network Security, 2019(1):3, January 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300042.

Anonymous:2019:NKD

[4158] Anonymous. North Korean defectors targeted. Network Security, 2019(1):3, January 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300054.

Anonymous:2019:ECE

[4159] Anonymous. ENISA: Cyber Europe 2018 — After Action Report. Network Security, 2019 (1):4, January 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300066.

Anonymous:2019:Ba

[4160] Anonymous. In brief. Network Security, 2019(1):5, January 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300078.

Perez:2019:PAS

[4161] Gabe Perez. A practical approach to securing IoT. Network Security, 2019(1):6–8, January 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581930008X.

Oliveri:2019:SUS

[4162] Andrea Oliveri and Filippo Lauria. Sagishi: an undercover software agent for infiltrating IoT botnets. Network Security, 2019(1): 9–14, January 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819300091.

Cox:2019:MRS

[4163] Gary Cox. Managing the risks of shadow IoT. Network Security, 2019 (1):14–17, January 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819300108.

Hodgson:2019:SSC

[4164] Roderick Hodgson. Solving the security challenges of IoT with public key cryptography. Network Security, 2019(1):17–19, January 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581930011X.

Goslin:2019:KYE

[4165] Ian Goslin. Know your enemy. Network Security, 2019(1):20, January 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300121.

Anonymous:2019:ECa

[4166] Anonymous. Events calendar. Network Security, 2019(1):20, January 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300133.

Anonymous:2019:UAH

[4167] Anonymous. US authorities hijack North Korea’s Joanap botnet. Network Security, 2019(2): 1–2, February 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300145.

Anonymous:2019:MDH

[4168] Anonymous. Massive data haul not all it seems. Network Security, 2019 (2):2–3, February 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300157.

Anonymous:2019:Tb

[4169] Anonymous. Threatwatch. Network Security, 2019(2):3, February 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300169.

Anonymous:2019:DTP

[4170] Anonymous. Digital transformation puts data at risk. Network Security, 2019(2):3, February 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300170.

Anonymous:2019:GSI

[4171] Anonymous. Gemalto: State of IoT security. Network Security, 2019(2):4, February 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300182.

Anonymous:2019:Bb

[4172] Anonymous. In brief. Network Security, 2019(2):5, February 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300194.

Rutherford:2019:ITS

[4173] Robert Rutherford. Internet of Things — striking the balance between competition and security. Network Security, 2019(2):6–8, February 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300200.

Keef:2019:UAD

[4174] Sean Keef. Using analytics-driven automation to target the biggest vulnerabilities. Network Security, 2019 (2):8–11, February 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300212.

Vincent:2019:DFP

[4175] Adam Vincent. Don’t feed the phish: how to avoid phishing attacks. Network Security, 2019(2): 11–14, February 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300224.

Bindley:2019:JDH

[4176] Phil Bindley. Joining the dots: how to approach compliance and data governance. Network Security, 2019 (2):14–16, February 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300236.

Yeoman:2019:HSB

[4177] Simon Yeoman. How secure are bare metal servers? Network Security, 2019 (2):16–17, February 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300248.

Newman:2019:URD

[4178] Sean Newman. Under the radar:

the danger of stealthy DDoS attacks. Network Security, 2019(2): 18–19, February 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581930025X.

Pogrebna:2019:MFA

[4179] Ganna Pogrebna, Karen Renaud, and Boris Taratine. The many faces of active cyber. Network Security, 2019(2):20, February 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819300261.

Anonymous:2019:ECb

[4180] Anonymous. Events calendar. Network Security, 2019(2):20, February 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485819300273.

Anonymous:2019:DBR

[4181] Anonymous. Data breaches reach record levels worldwide. Network Security, 2019(3):1–2, March 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819300285.

Anonymous:2019:FDB

[4182] Anonymous. Firms delay breach reports. Network Security, 2019 (3):2–3, March 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819300297.

Anonymous:2019:Tc

[4183] Anonymous. Threatwatch. Network Security, 2019(3):3, March 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485819300303.

Anonymous:2019:HBS

[4184] Anonymous. Huawei battles security concerns. Network Security, 2019(3):3, March 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300315.

Anonymous:2019:API

[4185] Anonymous. Accenture/Ponemon Institute: the cost of cybercrime. Network Security, 2019(3):4, March 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300327.

Anonymous:2019:Bc

[4186] Anonymous. In brief. Network Security, 2019(3):5, March 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300339.

Klein:2019:MSS

[4187] Dave Klein. Micro-segmentation:

securing complex cloud environments. Network Security, 2019 (3):6–10, March 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300340.

Canada:2019:RDW

[4188] Jesse Canada. Is reputational damage worse than a regulator’s fine? Network Security, 2019 (3):11–12, March 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819300352.

Schofield:2019:YWL

[4189] Greig Schofield. Has your wifi left you wide open to cybercrime? Network Security, 2019(3):13–14, March 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485819300364.

Haria:2019:GHS

[4190] Sam Haria. The growth of the hide and seek botnet. Network Security, 2019(3):14–17, March 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819300376.

Duncan:2019:HSY

[4191] Rory Duncan. How to secure your supply chain. Network Security, 2019 (3):18–19, March 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819300388.

Tankard:2019:IRF

[4192] Colin Tankard. IoT regulations: fact or fiction? Network Security, 2019(3):20, March 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581930039X.

Anonymous:2019:ECc

[4193] Anonymous. Events calendar. Network Security, 2019(3):20, March 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300406.

Anonymous:2019:FCB

[4194] Anonymous. Facebook confronted by a barrage of privacy issues. Network Security, 2019(4):1–3, April 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300418.

Anonymous:2019:Td

[4195] Anonymous. Threatwatch. Network Security, 2019(4):3, April 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581930042X.

Anonymous:2019:AIM

[4196] Anonymous. Asus issues malicious updates. Network Security, 2019(4):3, April 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300431.

Anonymous:2019:DCS

[4197] Anonymous. DCMS: Cyber Security Breaches Survey 2019. Network Security, 2019(4):4, April 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300443.

Anonymous:2019:Bd

[4198] Anonymous. In brief. Network Security, 2019(4):5, April 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300455.

Goldberg:2019:LDO

[4199] Daniel Goldberg. Living with decade-old vulnerabilities in datacentre software. Network Security, 2019(4):6–8, April 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300467.

Jewell:2019:ALA

[4200] Nick Jewell. The analytics lifecycle and the age of innovation. Network Security, 2019(4):8–11, April 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300479.

Giese:2019:EEI

[4201] Sascha Giese. The endpoint epidemic of IoT — just a bad dream? Network Security, 2019(4):11–12, April 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300480.

Herbert:2019:PSO

[4202] Sean Herbert. Preventing software from opening doors to the network. Network Security, 2019 (4):13–15, April 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300492.

Mathews:2019:WCM

[4203] Alex Mathews. What can machine learning do for information security? Network Security, 2019 (4):15–17, April 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300509.

Boughton:2019:PIC

[4204] Nick Boughton. Protecting infrastructure from cyber attack. Network Security, 2019(4):18–19, April 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300510.

Spence:2019:NFR

[4205] Dave Spence. The non-functional requirement fallacy. Network Security, 2019(4):20, April 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300522.

Anonymous:2019:ECd

[4206] Anonymous. Events calendar. Network Security, 2019(4):20, April 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300534.

Anonymous:2019:UGO

[4207] Anonymous. UK Government offers initiatives and assurances. Network Security, 2019(5):1–3, May 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485819300546.

Anonymous:2019:Te

[4208] Anonymous. Threatwatch. Network Security, 2019(5):3, May 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485819300558.

Anonymous:2019:PHV

[4209] Anonymous. Presidential hopefuls vulnerable. Network Security, 2019 (5):3, 19, May 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581930056X.

Anonymous:2019:HCR

[4210] Anonymous. Hiscox Cyber Readiness Report 2019. Network Security, 2019(5):4, May 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819300571.

Anonymous:2019:Be

[4211] Anonymous. In brief. Network Security, 2019(5):5, May 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819300583.

Yaacoubi:2019:REM

[4212] Omar Yaacoubi. The rise of encrypted malware. Network Security, 2019(5):6–9, May 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300595.

Gordon:2019:MT

[4213] Scott Gordon. A matter of trust.

Network Security, 2019(5):9–11, May 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300601.

Nguyen:2019:WAD

[4214] Van-Linh Nguyen, Po-Ching Lin, and Ren-Hung Hwang. Web attacks: defeating monetisation attempts. Network Security, 2019 (5):11–19, May 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300613.

Anonymous:2019:ROS

[4215] Anonymous. Risky open source.

Network Security, 2019(5):19, May 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300625.

Goslin:2019:CET

[4216] Ian Goslin. Cyber extortion is threatening Industry 4.0. Network Security, 2019(5):20, May 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300637.

Anonymous:2019:ECf

[4217] Anonymous. Events calendar. Network Security, 2019(5):20, May 2019.

CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485819300649.

Anonymous:2019:RBO

[4218] Anonymous. RDP bug in older Windows versions leaves millions vulnerable to ‘wormable’ attack. Network Security, 2019(6):1–2, June 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485819300650.

Anonymous:2019:UFM

[4219] Anonymous. UK fails to make progress. Network Security, 2019 (6):2–3, June 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819300662.

Anonymous:2019:Tf

[4220] Anonymous. Threatwatch. Network Security, 2019(6):3, June 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485819300674.

Anonymous:2019:ITG

[4221] Anonymous. Insider threat grows. Network Security, 2019(6):3, June 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485819300686.

Anonymous:2019:MSE

[4222] Anonymous. Mimecast: the State of Email Security Report 2019. Network Security, 2019(6):4, June 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300698.

Anonymous:2019:Bf

[4223] Anonymous. In brief. Network Security, 2019(6):5, June 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300704.

Team:2019:EEL

[4224] SophosLabs Research Team. Emotet exposed: looking inside highly destructive malware. Network Security, 2019(6):6–11, June 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300716.

Boughton:2019:PWC

[4225] Nick Boughton. Protecting the world of connected devices. Network Security, 2019(6):11–13, June 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300728.

Macy:2019:PVT

[4226] Jason Macy. Product vs toolkit: API and IAM security. Network Security, 2019(6):13–15, June 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581930073X.

Bhardwaj:2019:FET

[4227] Akashdeep Bhardwaj and Sam Goundar. A framework for effective threat hunting. Network Security, 2019 (6):15–19, June 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819300741.

Renaud:2019:EPM

[4228] Karen Renaud and Verena Zimmermann. Encouraging password manager use. Network Security, 2019(6):20, June 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819300753.

Anonymous:2019:ECg

[4229] Anonymous. Events calendar. Network Security, 2019(6):20, June 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485819300765.

Anonymous:2019:WBB

[4230] Anonymous. War breaks out between US and Iran in cyberspace. Network Security, 2019(7):1–3, July 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485819300777.

Anonymous:2019:Tg

[4231] Anonymous. Threatwatch. Network Security, 2019(7):3, July 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300789.

Anonymous:2019:CAM

[4232] Anonymous. China attacks major tech firms. Network Security, 2019(7):3, July 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300790.

Anonymous:2019:SMD

[4233] Anonymous. Sikich: 2019 Manufacturing and Distribution Report. Network Security, 2019(7):4, July 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300807.

Anonymous:2019:Bg

[4234] Anonymous. In brief. Network Security, 2019(7):5, July 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300819.

Roytman:2019:CPP

[4235] Michael Roytman and Jay Jacobs. The complexity of prioritising patching. Network Security, 2019(7):6–9, July 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300820.

Herbert:2019:WIS

[4236] Sean Herbert. Why IIoT should make businesses rethink security. Network Security, 2019(7):9–11, July 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300832.

Breitbarth:2019:IGO

[4237] Paul Breitbarth. The impact of GDPR one year on. Network Security, 2019(7):11–13, July 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300844.

Barker:2019:VHW

[4238] Peter Barker. Visual hacking — why it matters and how to prevent it. Network Security, 2019 (7):14–17, July 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300856.

Oetjen:2019:UAI

[4239] Jan Oetjen. Using artificial intelligence in the fight against spam. Network Security, 2019(7):17–19, July 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300868.

Tankard:2019:WYN

[4240] Colin Tankard. Why you need cyber essentials. Network Security, 2019(7):20, July 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581930087X.

Anonymous:2019:ECh

[4241] Anonymous. Events calendar. Network Security, 2019(7):20, July 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300881.

Anonymous:2019:MDB

[4242] Anonymous. More data breaches compromise millions of records. Network Security, 2019(8):1–3, August 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300893.

Anonymous:2019:Th

[4243] Anonymous. Threatwatch. Network Security, 2019(8):3, August 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581930090X.

Anonymous:2019:VP

[4244] Anonymous. Vulnerable PCs. Network Security, 2019(8):3a–3, August 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819300911.

Anonymous:2019:NAC

[4245] Anonymous. NCSC: Active Cyber Defence (ACD) — The Second Year. Network Security, 2019 (8):4, August 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300923.

Anonymous:2019:Bh

[4246] Anonymous. In brief. Network Security, 2019(8):5, August 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819300935.

Skelly:2019:WTW

[4247] Keenan Skelly. It’s a war out there: why malicious hackers are beating businesses to the punch — and what we can do about it. Network Security, 2019(8):6–8, August 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819300947.

Ghalaieny:2019:AIN

[4248] Dawood Ghalaieny. Achieving IoT nirvana by protecting cellular networks. Network Security, 2019 (8):8–12, August 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819300959.

OMalley:2019:SFA

[4249] Mike O’Malley. Securing the future of AgTech. Network Security, 2019 (8):12–14, August 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819300960.

Kurpjuhn:2019:DRA

[4250] Thorsten Kurpjuhn. Demystifying the role of AI for better network security. Network Security, 2019 (8):14–17, August 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300972.

Taylor:2019:DST

[4251] Adrian Taylor. Decrypting SSL traffic:

best practices for security, compliance and productivity. Network Security, 2019(8):17–19, August 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300984.

Spence:2019:RC

[4252] Dave Spence. Response in the cloud. Network Security, 2019 (8):20, August 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819300996.

Anonymous:2019:ECi

[4253] Anonymous. Events calendar. Network Security, 2019(8):20, August 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348581930100X.

Anonymous:2019:DRD

[4254] Anonymous. Dramatic rise in data breaches, with Yves Rocher, Monster and ECB among latest to be hit. Network Security, 2019(9): 1–3, September 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819301011.

Anonymous:2019:Ti

[4255] Anonymous. Threatwatch. Network Security, 2019(9):3, September 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485819301023.

Anonymous:2019:CGW

[4256] Anonymous. Cybercrime groups working together. Network Security, 2019 (9):3, September 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819301035.

Anonymous:2019:KSI

[4257] Anonymous. Kaspersky: the state of industrial cyber security. Network Security, 2019(9):4, September 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485819301047.

Anonymous:2019:Bi

[4258] Anonymous. In brief. Network Security, 2019(9):5, September 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485819301059.

Young:2019:ASO

[4259] Spencer Young. Automated systems only: why CISOs should switch off their dumb machines. Network Security, 2019(9):6–8, September 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485819301060.

Grimes:2019:MWH

[4260] Roger Grimes. The many ways to hack 2FA. Network Security, 2019 (9):8–13, September 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819301072.

Baldin:2019:BPF

[4261] Andy Baldin. Best practices for fighting the fileless threat. Network Security, 2019(9):13–15, September 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819301084.

Walker:2019:WDP

[4262] Paul Walker. Why do PAM projects fail? Network Security, 2019(9): 15–18, September 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819301096.

Bunting:2019:CSH

[4263] Charlene Bunting. Cloud security:

how to protect critical data and stay productive. Network Security, 2019 (9):18–19, September 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819301102.

Goodall:2019:IIS

[4264] Richard Goodall. The ideal industrial SOC. Network Security, 2019 (9):20, September 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819301114.

Anonymous:2019:ECj

[4265] Anonymous. Events calendar. Network Security, 2019(9):20, September 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485819301126.

Anonymous:2019:NWU

[4266] Anonymous. NCSC warns UK universities of cyberthreats. Network Security, 2019(10):1–2, October 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819301138.

Anonymous:2019:ULT

[4267] Anonymous. US looks to tighten aircraft security. Network Security, 2019(10):2, October 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348581930114X.

Anonymous:2019:FIM

[4268] Anonymous. FDA issues medical device warning. Network Security, 2019 (10):2–3, October 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819301151.

Anonymous:2019:Tj

[4269] Anonymous. Threatwatch. Network Security, 2019(10):3, October 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819301163.

Anonymous:2019:LCA

[4270] Anonymous. Loss of confidence among compliance professionals. Network Security, 2019(10):3, October 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819301175.

Anonymous:2019:BHa

[4271] Anonymous. Bitdefender: Hacked off! Network Security, 2019 (10):4, October 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819301187.

Anonymous:2019:Bj

[4272] Anonymous. In brief. Network Security, 2019(10):5, October 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819301199.

Heritage:2019:PIC

[4273] Ian Heritage. Protecting Industry 4.0:

challenges and solutions as IT, OT and IP converge. Network Security, 2019(10):6–9, October 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819301205.

Mansfield-Devine:2019:SOT

[4274] Steve Mansfield-Devine. The state of operational technology security. Network Security, 2019(10):9–13, October 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485819301217.

Bhardwaj:2019:CSA

[4275] Akashdeep Bhardwaj, Vinay Avasthi, and Sam Goundar. Cyber security attacks on robotic platforms. Network Security, 2019(10): 13–19, October 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819301229.

Tankard:2019:GSH

[4276] Colin Tankard. Goodbye SIEM, hello SOARX. Network Security, 2019 (10):20, October 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485819301230.

Anonymous:2019:ECk

[4277] Anonymous. Events calendar. Network Security, 2019(10):20, October 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485819301242.

Anonymous:2019:RPI

[4278] Anonymous. Russians pose as Iranians to attack Middle East targets. Network Security, 2019(11): 1–3, November 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819301254.

Anonymous:2019:CES

[4279] Anonymous. Child exploitation site is shut down. Network Security, 2019 (11):3, November 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819301266.

Anonymous:2019:Tk

[4280] Anonymous. Threatwatch. Network Security, 2019(11):3, November 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819301278.

Anonymous:2019:NAR

[4281] Anonymous. NCSC Annual Review 2019. Network Security, 2019 (11):4, November 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581930128X.

Anonymous:2019:Bk

[4282] Anonymous. In brief. Network Security, 2019(11):5, November 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819301291.

Knauer:2019:HCC

[4283] Chris Knauer. How contact centres can leave businesses exposed to cybercrime. Network Security, 2019 (11):6–9, November 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819301308.

Stitt:2019:OWR

[4284] Tom Stitt. Office walls and roadblocks: how workflows and terminology get in the way of visibility. Network Security, 2019(11): 9–13, November 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581930131X.

Homan:2019:IDS

[4285] Stu Homan. The importance of domains to security management. Network Security, 2019(11): 13–15, November 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819301321.

Aleem:2019:TWW

[4286] Azeem Aleem. Treading water: why organisations are making no progress on cyber security. Network Security, 2019 (11):15–18, November 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819301333.

Kunneke-Trenaman:2019:RFR

[4287] Nathalie Künneke-Trenaman. RPKI and the future of routing security. Network Security, 2019(11): 18–19, November 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819301345.

Renaud:2019:SWS

[4288] Karen Renaud and Marc Dupuis.

Should we scare people into behaving securely? Network Security, 2019 (11):20, November 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819301357.

Anonymous:2019:ECl

[4289] Anonymous. Events calendar. Network Security, 2019(11):20, November 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819301369.

Anonymous:2019:NWR

[4290] Anonymous. New wave of ransomware hits datacentre and healthcare targets. Network Security, 2019 (12):1–3, December 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819301370.

Anonymous:2019:Tl

[4291] Anonymous. Threatwatch. Network Security, 2019(12):3, December 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819301382.

Anonymous:2019:SP

[4292] Anonymous. Selfie phishing. Network Security, 2019(12):3, December 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819301394.

Anonymous:2019:BMY

[4293] Anonymous. Bitdefender: Mid-Year Threat Landscape Report 2019. Network Security, 2019(12):4, December 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819301400.

Anonymous:2019:Bl

[4294] Anonymous. In brief. Network Security, 2019(12):5, December 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819301412.

Callan:2019:CEV

[4295] Tim Callan and Chris Bailey. Changes to extended validation indicators put users at risk. Network Security, 2019 (12):6–8, December 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819301424.

Thompson:2019:AAC

[4296] Steve Thompson. Apprenticeships as the answer to closing the cyber skills gap. Network Security, 2019 (12):9–11, December 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819301436.

Al-kasassbeh:2019:WTD

[4297] Mouhammd Al-kasassbeh and Tariq Khairallah. Winning tactics with DNS tunnelling. Network Security, 2019 (12):12–19, December 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485819301448.

Spence:2019:RSC

[4298] Dave Spence. Resilience in the supply chain. Network Security, 2019 (12):20, December 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348581930145X.

Anonymous:2019:ECm

[4299] Anonymous. Events calendar. Network Security, 2019(12):20, December 2019. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485819301461.

Anonymous:2020:RUF

[4300] Anonymous. Ring under fire over weakness in video device security. Network Security, 2020(1): 1–2, January 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300015.

Anonymous:2020:CFT

[4301] Anonymous. Citrix flaw threatens large firms. Network Security, 2020 (1):2–3, January 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300027.

Anonymous:2020:Ta

[4302] Anonymous. Threatwatch. Network Security, 2020(1):3, January 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820300039.

Anonymous:2020:TD

[4303] Anonymous. TikTok dangers. Network Security, 2020(1):3, 19, January 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820300040.

Anonymous:2020:USG

[4304] Anonymous. Upstream security: 2020 Global Automotive Cyber security Report. Network Security, 2020(1):4, January 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300052.

Anonymous:2020:Ba

[4305] Anonymous. In brief. Network Security, 2020(1):5, January 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300064.

Reynolds:2020:TRD

[4306] Roy Reynolds. It’s time to rethink DDoS protection. Network Security, 2020(1):6–8, January 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300076.

Andreu:2020:OTS

[4307] Andres Andreu. Operational technology security — a data perspective. Network Security, 2020 (1):8–13, January 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300088.

Gordon:2020:SWB

[4308] Scott Gordon. Securing workers beyond the perimeter. Network Security, 2020(1):14–16, January 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348582030009X.

Bunker:2020:TCA

[4309] Guy Bunker. Targeted cyber attacks: how to mitigate the increasing risk. Network Security, 2020 (1):17–19, January 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300106.

Tankard:2020:PV

[4310] Colin Tankard. The power of voice. Network Security, 2020 (1):20, January 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300118.

Anonymous:2020:ECa

[4311] Anonymous. Events calendar. Network Security, 2020(1):20, January 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348582030012X.

Anonymous:2020:UNC

[4312] Anonymous. United Nations covered up hack of dozens of servers. Network Security, 2020(2):1–2, February 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820300131.

Anonymous:2020:CFR

[4313] Anonymous. Citrix flaw remains critical. Network Security, 2020 (2):2–3, February 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300143.

Anonymous:2020:Tb

[4314] Anonymous. Threatwatch. Network Security, 2020(2):3, February 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820300155.

Anonymous:2020:NFM

[4315] Anonymous. NSA finds major Windows bug. Network Security, 2020 (2):3, February 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300167.

Anonymous:2020:IOC

[4316] Anonymous. IBM/ObserveIT: Cost of insider threats 2020. Network Security, 2020(2):4, February 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300179.

Anonymous:2020:Bb

[4317] Anonymous. In brief. Network Security, 2020(2):5, February 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820300180.

Blake:2020:RRE

[4318] Cindy Blake. Reducing risk with end-to-end application security automation. Network Security, 2020 (2):6–8, February 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300192.

Bajpai:2020:DNR

[4319] Pranshu Bajpai and Richard Enbody. Dissecting .NET ransomware: key generation, encryption and operation. Network Security, 2020 (2):8–14, February 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300209.

Bhardwaj:2020:KSC

[4320] Akashdeep Bhardwaj and Sam Goundar.

Keyloggers: silent cyber security weapons. Network Security, 2020 (2):14–19, February 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300210.

MacMillan:2020:STC

[4321] Kate MacMillan. Signs of things to come? Network Security, 2020 (2):20, February 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300222.

Anonymous:2020:ECb

[4322] Anonymous. Events calendar. Network Security, 2020(2):20, February 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820300234.

Anonymous:2020:WVF

[4323] Anonymous. Worrying vulnerabilities found in Intel, AMD, Cypress and Broadcom chips. Network Security, 2020(3):1–3, March 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300246.

Anonymous:2020:Tc

[4324] Anonymous. Threatwatch. Network Security, 2020(3):3, March 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820300258.

Anonymous:2020:MTN

[4325] Anonymous. Microsoft takes down Necurs botnet. Network Security, 2020(3):3, March 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348582030026X.

Anonymous:2020:EVS

[4326] Anonymous. Edgescan: Vulnerability stats report. Network Security, 2020(3):4, March 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300271.

Anonymous:2020:Bc

[4327] Anonymous. In brief. Network Security, 2020(3):5, March 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300283.

Reynolds:2020:FBM

[4328] Roy Reynolds. The four biggest malware threats to UK businesses. Network Security, 2020(3):6–8, March 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820300295.

Borchani:2020:AMB

[4329] Yessine Borchani. Advanced malicious beaconing detection through AI. Network Security, 2020(3):8–14, March 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820300301.

Dyess:2020:MBB

[4330] Colby Dyess. Maintaining a balance between agility and security in the cloud. Network Security, 2020 (3):14–17, March 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300313.

Geenens:2020:HWS

[4331] Pascal Geenens. How worried should you be about nation-state attacks? Network Security, 2020 (3):17–19, March 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300325.

Renaud:2020:SPC

[4332] Karen Renaud. Strong passwords can’t keep children safe. Network Security, 2020(3):20, March 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300337.

Anonymous:2020:ECc

[4333] Anonymous. Events calendar. Network Security, 2020(3):20, March 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820300349.

Anonymous:2020:RPI

[4334] Anonymous. Ransomware payments increase as more attackers leak data. Network Security, 2020(4):1–3, April 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300350.

Anonymous:2020:Te

[4335] Anonymous. Threatwatch. Network Security, 2020(4):3, April 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820300362.

Anonymous:2020:CT

[4336] Anonymous. Cybercrime takedowns.

Network Security, 2020(4):3, April 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820300374.

Anonymous:2020:LSE

[4337] Anonymous. Linux servers exploited for a decade. Network Security, 2020(4):3, April 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300386.

Anonymous:2020:WIS

[4338] Anonymous. Watchguard: Internet Security Report Q4 2019. Network Security, 2020(4):4, April 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300398.

Anonymous:2020:Bd

[4339] Anonymous. In brief. Network Security, 2020(4):5, April 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300404.

Ray:2020:DDN

[4340] Terry Ray. DDoS defence: new tactics for a rising shadow industry. Network Security, 2020 (4):6–7, April 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300416.

Chapman:2020:YIS

[4341] Phil Chapman. Are your IT staff ready for the pandemic-driven insider threat? Network Security, 2020(4):8–11, April 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300428.

Wright:2020:ESN

[4342] Cary Wright. Essentials for selecting a network monitoring tool. Network Security, 2020(4):11–14, April 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348582030043X.

Kelf:2020:SRC

[4343] Simon Kelf. The security risks created by cloud migration and how to overcome them. Network Security, 2020(4):14–16, April 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300441.

Higgins:2020:IRW

[4344] David Higgins. Innovation and risk walk hand-in-hand with 5G and IoT. Network Security, 2020 (4):16–18, April 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300453.

Hockey:2020:UCS

[4345] Alyn Hockey. Uncovering the cyber security challenges in healthcare. Network Security, 2020 (4):18–19, April 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300465.

Fox:2020:LHT

[4346] David Fox. Losing the human touch.

Network Security, 2020(4):20, April 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820300477.

Anonymous:2020:ECd

[4347] Anonymous. Events calendar. Network Security, 2020(4):20, April 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820300489.

Anonymous:2020:CHA

[4348] Anonymous. Chinese hackers attacking Covid-19 researchers, US warns. Network Security, 2020(5):1–2, May 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820300490.

Anonymous:2020:IMT

[4349] Anonymous. InfinityBlack market taken down. Network Security, 2020(5):2–3, May 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300507.

Anonymous:2020:Tf

[4350] Anonymous. Threatwatch. Network Security, 2020(5):3, May 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820300519.

Anonymous:2020:NNK

[4351] Anonymous. New North Korean threats. Network Security, 2020(5):3, May 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300520.

Anonymous:2020:ACR

[4352] Anonymous. AT&T Cybersecurity: the relationship between security maturity and business enablement. Network Security, 2020(5):4, May 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300532.

Anonymous:2020:Be

[4353] Anonymous. In brief. Network Security, 2020(5):5, May 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300544.

Malecki:2020:OSP

[4354] Florian Malecki. Optimising storage processes to reduce the risk of ransomware. Network Security, 2020(5):6–8, May 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300556.

Meryem:2020:HID

[4355] Amar Meryem and Bouabid El Ouahidi. Hybrid intrusion detection system using machine learning. Network Security, 2020(5):8–19, May 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820300568.

Tankard:2020:PUN

[4356] Colin Tankard. Pandemic underpins need for SOAR. Network Security, 2020(5):20, May 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348582030057X.

Anonymous:2020:ECe

[4357] Anonymous. Events calendar. Network Security, 2020(5):20, May 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820300581.

Anonymous:2020:RNS

[4358] Anonymous. Russian nation-state attackers target Exim mail servers. Network Security, 2020(6):1–2, June 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820300593.

Anonymous:2020:NWA

[4359] Anonymous. Network and web app attacks increase. Network Security, 2020(6):2–3, June 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348582030060X.

Anonymous:2020:Tg

[4360] Anonymous. ThreatWatch. Network Security, 2020(6):3, June 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820300611.

Anonymous:2020:SEC

[4361] Anonymous. Scammers exploit Covid-19 measures. Network Security, 2020(6):3, June 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300623.

Anonymous:2020:TMS

[4362] Anonymous. Trend micro: Shifts in underground markets. Network Security, 2020(6):4, June 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300635.

Anonymous:2020:Bf

[4363] Anonymous. In brief. Network Security, 2020(6):5, June 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300647.

Powell:2020:HDC

[4364] Leila Powell. How data can be the lingua franca for security and IT. Network Security, 2020(6):6–7, June 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300659.

Hampton:2020:KSH

[4365] Paul Hampton. Keeping a secure hold on data through modern electronic content management. Network Security, 2020(6):8–11, June 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300660.

Evangelakos:2020:KCA

[4366] Gus Evangelakos. Keeping critical assets safe when teleworking is the new norm. Network Security, 2020(6):11–14, June 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300672.

Bulpett:2020:SAI

[4367] Ben Bulpett. Safeguarding against the insider threat. Network Security, 2020(6):14–17, June 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300684.

Deighton:2020:KST

[4368] Thomas Deighton and Michael Wakefield. Keep security top of mind when moving into the cloud. Network Security, 2020(6):17–19, June 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820300696.

MacMillan:2020:TC

[4369] Kate MacMillan. Trusting the CISO. Network Security, 2020(6):20, June 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820300702.

Anonymous:2020:ECf

[4370] Anonymous. Events calendar. Network Security, 2020(6):20, June 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820300714.

Anonymous:2020:AGCa

[4371] Anonymous. Australian Government claims ‘sophisticated’ attack by nation-state actors. Network Security, 2020(7):1–2, July 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300726.

Anonymous:2020:CRC

[4372] Anonymous. Call for revisions to Computer Misuse Act. Network Security, 2020(7):2–3, July 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300738.

Anonymous:2020:Th

[4373] Anonymous. Threatwatch. Network Security, 2020(7):3, July 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348582030074X.

Anonymous:2020:MDE

[4374] Anonymous. A million dodgy emails.

Network Security, 2020(7):3, July 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820300751.

Anonymous:2020:ICR

[4375] Anonymous. IBM: 2020 Cyber Resilient Organisation Report. Network Security, 2020(7):4, July 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820300763.

Anonymous:2020:Bg

[4376] Anonymous. In brief. Network Security, 2020(7):5, July 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300775.

Cope:2020:SSS

[4377] Rod Cope. Strong security starts with software development. Network Security, 2020(7):6–9, July 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300787.

Jartelius:2020:DBI

[4378] Martin Jartelius. The 2020 Data Breach Investigations Report — a CSO’s perspective. Network Security, 2020(7):9–12, July 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300799.

Lim:2020:AMC

[4379] Mieng Lim. Avoiding the most common vulnerability-management pitfalls. Network Security, 2020 (7):12–14, July 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300805.

Wood:2020:APD

[4380] Simon Wood. Adhering to privacy by design with identity-as-a-service. Network Security, 2020 (7):14–17, July 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300817.

Murray:2020:OSS

[4381] Dale Murray. Open source and security: why transparency now equals strength. Network Security, 2020 (7):17–19, July 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300829.

Renaud:2020:BRR

[4382] Karen Renaud, Zinaida Benenson, and Daniela Oliveira. Balancing resistance and resilience. Network Security, 2020(7):20, July 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300830.

Anonymous:2020:ECg

[4383] Anonymous. Events calendar. Network Security, 2020(7):20, July 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820300842.

Anonymous:2020:IOM

[4384] Anonymous. Intel and other major firms suffer source code leaks. Network Security, 2020(8):1–3, August 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820300854.

Anonymous:2020:Ti

[4385] Anonymous. Threatwatch. Network Security, 2020(8):3, August 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820300866.

Anonymous:2020:WIC

[4386] Anonymous. Warning issued over critical infrastructure. Network Security, 2020(8):3, August 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300878.

Anonymous:2020:NNO

[4387] Anonymous. Nozomi Networks: OT/ IoT security report. Network Security, 2020(8):4, August 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348582030088X.

Anonymous:2020:Bh

[4388] Anonymous. In brief. Network Security, 2020(8):5, August 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300891.

Anonymous:2020:HCCa

[4389] Anonymous. Hacking connected cars: Tactics, techniques and procedures. Network Security, 2020 (8):6, August 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300908.

Anonymous:2020:HCCb

[4390] Anonymous. Hunting cyber criminals: a hacker’s guide to online intelligence gathering tools and techniques. Network Security, 2020 (8):6, August 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348582030091X.

Rodbert:2020:WOR

[4391] Mark Rodbert. Why organisational readiness is vital in the fight against insider threats. Network Security, 2020(8):7–9, August 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300921.

Waterson:2020:MEW

[4392] Dave Waterson. Managing endpoints, the weakest link in the security chain. Network Security, 2020 (8):9–13, August 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300933.

Secara:2020:ZEE

[4393] Ion-Alexandru Secara. Zoombombing — the end-to-end fallacy. Network Security, 2020(8):13–17, August 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820300945.

Duncan:2020:WDS

[4394] Rory Duncan. What does ‘secure by design’ actually mean? Network Security, 2020(8):18–19, August 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820300957.

Fox:2020:DTD

[4395] David Fox. Dealing with technical debt. Network Security, 2020 (8):20, August 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820300969.

Anonymous:2020:ECh

[4396] Anonymous. Events calendar. Network Security, 2020(8):20, August 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820300970.

Anonymous:2020:UPE

[4397] Anonymous. US presidential election already under attack. Network Security, 2020(9):1–2, September 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820300982.

Anonymous:2020:UGF

[4398] Anonymous. UK Government funds better healthcare security. Network Security, 2020(9):2–3, September 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820300994.

Anonymous:2020:Tj

[4399] Anonymous. Threatwatch. Network Security, 2020(9):3, September 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820301008.

Anonymous:2020:AGCb

[4400] Anonymous. APT groups change tactics. Network Security, 2020 (9):3, September 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348582030101X.

Anonymous:2020:RUH

[4401] Anonymous. Rapid7: Under the hoodie. Network Security, 2020 (9):4, September 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301021.

Anonymous:2020:Bi

[4402] Anonymous. In brief. Network Security, 2020(9):5, September 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820301033.

Campfield:2020:PMN

[4403] Mike Campfield. The problem with (most) network detection and response. Network Security, 2020 (9):6–9, September 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301045.

Cavaliere:2020:SIQ

[4404] Fabio Cavaliere, John Mattsson, and Ben Smeets. The security implications of quantum cryptography and quantum computing. Network Security, 2020 (9):9–15, September 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301057.

Barak:2020:CIU

[4405] Israel Barak. Critical infrastructure under attack: lessons from a honeypot. Network Security, 2020(9): 16–17, September 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301069.

Eckel:2020:EIS

[4406] Michael Eckel and Tom Laffey. Ensuring the integrity and security of network equipment is critical in the fight against cyber attacks. Network Security, 2020(9): 18–19, September 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301070.

Cooke:2020:UYC

[4407] Andrew Cooke. A unique year for cyber security. Network Security, 2020 (9):20, September 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301082.

Anonymous:2020:ECi

[4408] Anonymous. Events calendar. Network Security, 2020(9):20, September 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820301094.

Anonymous:2020:ZFE

[4409] Anonymous. Zerologon flaw exploited in the wild. Network Security, 2020 (10):1–2, October 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301100.

Anonymous:2020:DAH

[4410] Anonymous. DDoS attacks hit hard and fast. Network Security, 2020 (10):2–3, October 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301112.

Anonymous:2020:Tk

[4411] Anonymous. Threatwatch. Network Security, 2020(10):3, October 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820301124.

Anonymous:2020:EIR

[4412] Anonymous. Enterprise IoT at risk. Network Security, 2020(10): 3b–3, October 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301136.

Anonymous:2020:MDD

[4413] Anonymous. Microsoft Digital Defense Report. Network Security, 2020(10):4, October 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301148.

Anonymous:2020:NB

[4414] Anonymous. News in brief. Network Security, 2020(10):5, October 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S135348582030115X.

Morag:2020:EHU

[4415] Assaf Morag. Exploitable hosts used in cloud native cyber attacks. Network Security, 2020 (10):6–10, October 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820301161.

Erez:2020:HTA

[4416] Nadav Erez. How threat actors abuse ICS-specific file types. Network Security, 2020(10):10–13, October 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820301173.

Hofmann:2020:HOC

[4417] Tom Hofmann. How organisations can ethically negotiate ransomware payments. Network Security, 2020 (10):13–17, October 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820301185.

Meyers:2020:DHD

[4418] Robert Meyers. Data highway and the digital transformation: arguments for secure, centralised log management. Network Security, 2020 (10):17–19, October 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301197.

Tankard:2020:RWR

[4419] Colin Tankard. Remote working reset now required. Network Security, 2020(10):20, October 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301203.

Anonymous:2020:ECj

[4420] Anonymous. Events calendar. Network Security, 2020(10):20, October 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820301215.

Anonymous:2020:GIS

[4421] Anonymous. Governments implement sanctions against nation-state attackers. Network Security, 2020 (11):1–3, November 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301227.

Anonymous:2020:Tl

[4422] Anonymous. Threatwatch. Network Security, 2020(11):3, November 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820301239.

Anonymous:2020:CSR

[4423] Anonymous. Control system report finds poor security. Network Security, 2020(11):3, November 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820301240.

Anonymous:2020:GRF

[4424] Anonymous. Guardicore: Rethink firewalls — security and agility for the modern enterprise. Network Security, 2020(11):4, November 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820301252.

Anonymous:2020:Bj

[4425] Anonymous. In brief. Network Security, 2020(11):5, November 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820301264.

Hodler:2020:CGM

[4426] Amy Hodler. Can graphs mitigate against coronavirus-related cybercrime? Network Security, 2020 (11):6–8, November 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820301276.

Still:2020:BMH

[4427] Andy Still. Bot mitigation — how gaps in understanding and ownership are exposing businesses to greater threats. Network Security, 2020 (11):8–11, November 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301288.

Wasicek:2020:FSH

[4428] Armin Wasicek. The future of 5G smart home network security is microsegmentation. Network Security, 2020 (11):11–13, November 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348582030129X.

Trzupek:2020:HDP

[4429] Brian Trzupek. How DIY PKI often negates the promise of public key encryption. Network Security, 2020 (11):14–17, November 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301306.

Seifert:2020:DIS

[4430] René Seifert. Digital identities — selfsovereignty and blockchain are the keys to success. Network Security, 2020 (11):17–19, November 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301318.

Renaud:2020:HNC

[4431] Karen Renaud and Verena Zimmermann. How to nudge in cyber security. Network Security, 2020 (11):20, November 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348582030132X.

Anonymous:2020:ECk

[4432] Anonymous. Events calendar. Network Security, 2020(11):20, November 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820301331.

Anonymous:2020:URE

[4433] Anonymous. UK reveals existence of its military cyber operations unit. Network Security, 2020 (12):1–2, December 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820301343.

Anonymous:2020:Tm

[4434] Anonymous. Threatwatch. Network Security, 2020(12):3, December 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820301355.

Anonymous:2020:SFB

[4435] Anonymous. Security firms breached. Network Security, 2020(12):3, December 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1353485820301367.

Anonymous:2020:SDP

[4436] Anonymous. Synopsys: DevSecOps practices and open-source management in 2020. Network Security, 2020(12):4, December 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301379.

Anonymous:2020:Bk

[4437] Anonymous. In brief. Network Security, 2020(12):5, December 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485820301380.

Wood:2020:HSD

[4438] Michael Wood. How SASE is defining the future of network security. Network Security, 2020(12): 6–8, December 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301392.

Sangster:2020:WIC

[4439] Mark Sangster. When it comes to cyber security, ignorance isn’t bliss — it’s negligence. Network Security, 2020 (12):8–12, December 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301409.

Mansfield-Devine:2020:NSA

[4440] Steve Mansfield-Devine. Nationstate attacks: the escalating menace. Network Security, 2020(12): 12–17, December 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485820301410.

Evangelakos:2020:WCS

[4441] Gus Evangelakos. Where conventional security control validation falls short when evaluating organisational threats. Network Security, 2020(12): 18–19, December 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485820301422.

Stone:2020:SWN

[4442] Paul Stone. Securing Windows networks against WSUS attacks. Network Security, 2020(12):20, December 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820301434.

Anonymous:2020:ECl

[4443] Anonymous. Events calendar. Network Security, 2020(12):20, December 2020. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485820301446.

Anonymous:2021:SSC

[4444] Anonymous. SolarWinds supply chain breach threatens government agencies and enterprises worldwide. Network Security, 2021(1):1–3, January 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485821000015.

Anonymous:2021:Ta

[4445] Anonymous. Threatwatch. Network Security, 2021(1):3, January 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485821000027.

Anonymous:2021:CEB

[4446] Anonymous. The cyber-elephant in the boardroom: Cyber-accountability with the five pillars of security framework. Network Security, 2021 (1):4, January 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485821000039.

Anonymous:2021:HMA

[4447] Anonymous. Hacking multifactor authentication. Network Security, 2021(1):4, January 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485821000040.

Anonymous:2021:Ba

[4448] Anonymous. In brief. Network Security, 2021(1):5, January 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485821000052.

Gent:2021:HCM

[4449] Andy Gent. How can mobile networks protect critical infrastructure? Network Security, 2021 (1):6–8, January 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485821000064.

Klein:2021:RFH

[4450] Dave Klein. Relying on firewalls? Here’s why you’ll be hacked. Network Security, 2021 (1):9–12, January 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485821000076.

Udipi:2021:EDM

[4451] Sudhir Udipi. The event data management problem: getting the most from network detection and response. Network Security, 2021 (1):12–14, January 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485821000088.

Morley:2021:SLU

[4452] Connor Morley. Shining a light on UEFI — the hidden memory space being exploited in attacks. Network Security, 2021(1): 14–17, January 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S135348582100009X.

Malecki:2021:NTM

[4453] Florian Malecki. Now is the time to move past traditional 3-2-1 backups. Network Security, 2021(1): 18–19, January 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485821000106.

Jones:2021:PSC

[4454] Phil Jones. Power and security in a common mission. Network Security, 2021(1):20, January 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485821000118.

Anonymous:2021:ECa

[4455] Anonymous. Events calendar. Network Security, 2021(1):20, January 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S135348582100012X.

Anonymous:2021:FFH

[4456] Anonymous. Florida facility hacked in attempt to poison water. Network Security, 2021(2):1–2, February 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485821000131.

Anonymous:2021:MFS

[4457] Anonymous. More fallout from SolarWinds hack. Network Security, 2021(2):2–3, February 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485821000143.

Anonymous:2021:Tb

[4458] Anonymous. Threatwatch. Network Security, 2021(2):3, February 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485821000155.

Anonymous:2021:VSS

[4459] Anonymous. Veracode: State of software security v11. Network Security, 2021(2):4, February 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485821000167.

Anonymous:2021:Bb

[4460] Anonymous. In brief. Network Security, 2021(2):5, February 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485821000179.

Mansfield-Devine:2021:WKD

[4461] Steve Mansfield-Devine. Who’s that knocking at the door? The problem of credential abuse. Network Security, 2021(2):6–15, February 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485821000180.

Sheridan:2021:SZT

[4462] Ollie Sheridan. The state of zero trust in the age of fluid working. Network Security, 2021(2): 15–17, February 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485821000192.

Walsh:2021:ACD

[4463] Brendan Walsh. Avoiding costly downtime — how MSPs can manage their networks. Network Security, 2021 (2):17–19, February 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485821000209.

Tankard:2021:CSN

[4464] Colin Tankard. Credential stuffing — the new hack. Network Security, 2021(2):20, February 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485821000210.

Anonymous:2021:DCP

[4465] Anonymous. Due to the Covid-19 pandemic, many conferences are being cancelled, postponed or converted into virtual events. The events listed here were still planned to proceed at the time of publication. Network Security, 2021(2):20, February 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485821000222.

Anonymous:2021:AVL

[4466] Anonymous. Accellion vulnerabilities lead to breaches of major organisations. Network Security, 2021 (3):1–2, March 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485821000234.

Anonymous:2021:NKT

[4467] Anonymous. North Korea targets defence firms and Pfizer. Network Security, 2021(3):2–3, March 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485821000246.

Anonymous:2021:Tc

[4468] Anonymous. Threatwatch. Network Security, 2021(3):3, March 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1353485821000258.

Anonymous:2021:IXF

[4469] Anonymous. IBM: 2021 X-Force threat intelligence index. Network Security, 2021(3):4, March 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S135348582100026X.

Anonymous:2021:Bc

[4470] Anonymous. In brief. Network Security, 2021(3):5, March 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485821000271.

Anderson:2021:ICA

[4471] Chad Anderson, John ‘Turbo’ Conwell, and Tarik Saleh. Investigating cyber attacks using domain and DNS data. Network Security, 2021(3):6–8, March 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1353485821000283.

Ali:2021:LFC

[4472] Rashid Ali. Looking to the future of the cyber security landscape. Network Security, 2021 (3):8–10, March 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485821000295.

Mansfield-Devine:2021:LDT

[4473] Steve Mansfield-Devine. Locking the door: tackling credential abuse. Network Security, 2021 (3):11–19, March 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485821000301.

Renaud:2021:LP

[4474] Karen Renaud. Learning from the past. Network Security, 2021 (3):20, March 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1353485821000313.

Anonymous:2021:ECb

[4475] Anonymous. Events calendar. Network Security, 2021(3):20, March 2021. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1353485821000325.